A Complete Bibliography of Publications in Applied Mathematics and Computation: 2020–2024

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

28 August 2024 Version 1.53

Title word cross-reference

[CHJW23]. (1, 2) [MW20]. (1, 6) [ZCW24b]. (2 + 1) [ÖY21, ZLA^{+}21]. (2, r) [HZ23b]. (3 + 1) [GLZB20]. (4, 6) [SZZ22]. (k + 1) [ZLL20]. (m, n) [ZZ23]. (n, k) [GLW22, Guo23]. (n, s) [CLL23]. (ω, Lδ) [ZZX20a]. (Q, S, R) − α [LR23]. (s, 0, 0, 0) [GGW21b]. (ξ1, ξ, ξ2) [HLSL22]. −1 [XLS21]. −2 [RS22b]. 0 < α < 1 [ZZ20a]. 1 [CCCdLMR23, DHZ23, GCGM21, LW20a, LLL23b, LZ20c, LZ22e, PKJB22, PW22a, SZZ24, TCZ^{+}20]. {1, 2, 3, 1^{m}} [WC22c]. 10 [PWL23]. 1 < β ≤ 2 [ZLZ24a]. 2 [ABPP23, AC20, ADS23, BS22a, BFS20, BFRS24, CG21, CHJW23, CW23, CCHZ23, CZ21b, FRW^{+}24, GPW24, HT23, IZ20, JLOR22, KWH^{+}23, LHET23, LL23b, LC23, Ma21a, MTRA22, MVCK24, MV22, PBM20, PW22a, RO21, jSZyW^{+}20, UGG^{+}20, WDG21, WZCF21, WGSS22, YLH22, YCW23, ZL22b, ZZLK23, ZLU21]. 2 − (v, k, 1) [TYL^{+}20]. 2 − (v, k, λ) [TCZ22]. 3 [AHM23, BRFLGH21, BT22, CG21, DS22a, DLM20, DK22c, Fre22, INT20, JLJ23, JLOR22, LQLF22, LR21, LLSY23, LTC^{+}23, LR20, LL20b, LZY^{+}23b, LHFH22, MZ22, MV22, RWGS22, SH21, SFC23, UGG^{+}20, WGSS22, WHJK23, WWY24, WWT24, YL20b, YLH22, ZHFL23a]. 312 [WLY20]. 4 [BHGPC21, LR22, LW23b, LHZ21d]. 4+ [YI22]. 4/2 [ZEA22]. 4 × 4 [HLZ22, LWL^{+}20]. 5 [SS24b, XJ20]. [1, 2] [HCSJ24]. [h, k] [YM22a]. ^{+} [SWCT21]. ^{1} [RKJ20]. _{18} [KHA22]. _{2} [AGC20, KHA22]. 2.5 [WWZ20b]. _{3} [AGC20]. _{4} [KHA22]. _{8} [KHA22]. _{σ} [LWNZ22]. A_{α} [CLZ23, HWLL21, HLX20]. α [BD20a, DM21, DW20, LYhH^{+}24, LZ20b, TCG20]. AX = C, XB = D [ZLLY22]. AXA ≤?A [WL21a]. AXB = C [TLDL21, TWWL24, LLFZ20]. AXB = D [ZLLY22]. B [NNKH21, Rom21, ZZL20b, GN20a]. Bl, s [WY22]. C [LM22, NNKH21]. C^{∗} [ZYW21b]. C^{0} [CWZZ22, DR22]. C^{1} [GKK^{+}24]. C^{2} [FKVZ21, Yun21]. C^{3} [LMS24]. C_{3} [WZ24]. C_{4} [WZ24]. d [KRE20, WS21, Zha24b]. D(1, 2, t) [HH23b]. D(1, t) [HH23b]. D_{α} [QLM21]. Δ [TZ21]. Δ_{ω} [ÖÇÖ23]. ℓ_{0} [WWZ^{+}23]. ℓ_{2} [WSG^{+}21, XWWC21]. ℓ_{∞} [XWWC21]. ℓ_{p} [WSG^{+}21]. ϵ [MZLS22]. F [LHW^{+}23, DW20, LYB^{+}21]. G [BIJ22, MW20, YC21, CS24c, ZLC^{+}24b]. G(ϵ) [BIJ22]. G^{1} [BCHRS23, MMM24]. G^{2} [BCHRS23, KPS22]. GCI [LL23c]. H [HLSL22, TCPF24, YWX20b, ZZS^{+}20, YZM24]. H^{1} [KPP22]. H^{∞} [LZZ^{+}23]. H_{+} [ZV20]. H_{2} [QLC^{+}24]. H_{2k} [YLY24]. H_{∞} [BRTR21, CZWS20, CWH24, CZK^{+}22, CZ24b, CWJF21, DLS^{+}23, DL20c, FRW^{+}24, FSS20, FLH^{+}21, GS21, GSRL20, HTF20, HCL^{+}23, HKCK24, HSL^{+}24, KJL20, LSPK20, LHZ^{+}23, LZZG23, LZS^{+}21, LCL22a, MZSW21, NK21a, NH22, QZH21, SZ24b, SC20b, SZP21, SZWL21, SLPH22, SDW^{+}22, TWL20, TW21b, WCZY20, WCM^{+}21, WZC^{+}22, WC22a, WCW22, WYZX22, WC23, WLHW23, WHCW20, WB23, XH22a, XZQX20, XZZC22, XWDK23, YGZ^{+}22, YLCZ20, YLK^{+}21, YCZ^{+}23, ZDGL21, ZZQ^{+}20, ZCW^{+}24a, ZLZ24b, ZM21c, ZXZ20, ZC21b, ZCZW23, ZLP21]. H∞ [DZZ^{+}20]. K [JWY20, SNMM21, AB20a, ALM20, CLL23, CZMX21, CCHZ23, CLB21, DG20b, HCJC21, JTW^{+}23, KRT21, LJAM24, LW21a, LHW^{+}23, LM21, MB24, QLZL20, SVW23, WLL^{+}22, WLZD24, XYZZ22, ZLL20, ZHFL23a]. K_{1,4} [WJY21]. K_{1,r} [CCHZ23]. K_{4}(7) [CFL^{+}20]. K_{4}^{−} [CY24]. K_{5,n+1}\e [HW20]. K_{5} − e [Xu24b]. K_{λ_{1}} ∪ K_{λ_{2}} [ZY24a]. K_{t,s} [WLL21]. k ∈ {1, 2, 3} [CLB21]. L [GI23, RKJ20, RVV24, LY22a]. L(p, q) [HJL^{+}23]. L1 [HHGZ23]. L^{1} [Kwo22b]. L^{α} [dSAOG21]. L^{p} [STA24]. L^{p}([0, 1], Ω) [LFLP21]. L_{1} [LSZ^{+}23, SZ24a, ZLJ^{+}21]. L_{2} [LZZ24a, MWS23, HLW^{+}20, MFS22, CZ20]. L_{∞} [HLW^{+}20, KKH23, MWS23, CZ20, MFS22]. l_{Q} [PX20]. M [DLQY20, DWZ20, HLX21, Zha23a, AA21b]. m − 3 [MZ23b]. m − 4 [MZ23b]. Z [SRZK22]. H∈ [OPL22]. L_{∞} [LXZ23]. Q [ZCT22]. Le′vy [KK23c]. W [Lóc22]. m ≠ 3 [ZZ23]. μ [ZLLX22]. N [BSCL20, GLLZ24, LWL^{+}20, ALM20, CZL20, DG20b, HPL21, LHW^{+}23, ZHFL23a]. n − 2 [WLZD21]. ω [LKSS21]. osp(2|2) [HCS20]. P [NGGD23, SK20, ÅCL24, Baz20, CT24, HBCA21, JZ22, LL22a, LGX22, RGAM20, WLW20, YLL22, ZYySX23]. p^{3} [ZDMW23]. P_{3} [LCL23]. P_{4} [JJC24]. P_{5} [Xu24b, SLL24]. P_{k} [CDH23, Kun24]. PGL_{2}(q) [TYL^{+}20]. π [CDD23]. ±1 [HT24]. ψ [BGRM21, URLC21]. Q [LLZ20c, WDZL24, WFKL24, YZPL24, ÅCL24, AM20a, LCL23, LLH21, RLX22, jSZyW^{+}20, WYCL24, ZZ23]. R [CCdlH21, ZIML20, CFL^{+}20, LLL23b, PP23, Rom21, SLL24, WWZ20a, WX20a, YZM24]. R^{2} [LQSZ21]. R^{n} [GC20]. r_{pot}(C_{n}, K_{t}^{−k}) [ZY23a]. r_{pot}(P_{n}, K_{t}^{−k}) [ZY23a]. ρ [CRDD21]. S [CZWS20, CZ24b, GZQ^{+}20, HM20a, HWC20, HLX21, HH23b, KJL20, LKSS21, LX24a, SWZS20, YW22a, YXZ^{+}24, HLF22]. S^{p} [HWC20]. S_{ℓ_{1}} ∪ S_{ℓ_{2}} [LLY20]. s ≥ 4 [HLF22]. σ_{t} [DS23]. spo(2|2) [HCS20]. ∑_{i=1}^{t} A_{i}X_{i}B_{i} = E [QXX21]. T [CZWS20, CZ24b, GZQ^{+}20, KJL20, LX24a, SWZS20, YXZ^{+}24, LW21a]. θ [LCL22c, QS20]. v < 100 [TCZ22]. C⃗_{3} [HLSL22]. W [ÅCL24, JMM20]. X^{∗}AX + B^{∗}X + D = 0 [YLZL21]. X^{∗}X = I_{p} [ZLLY22]. Xa = b [GC20]. Y [WYW^{+}20]. Z [LM22]. Z_{n} [SLJ23].

-algebras [ZYW21b]. -apex [CLB21]. -arboricity [CW23]. -ary [DG20b, ZHFL23a]. -asymptotic [HWC20, LKSS21]. -avoiding [WLY20]. -based [LSZ^{+}23]. -benzenoids [LYB^{+}21]. -beta [LJAM24]. -binomial [ÅCL24]. -bipartite [YM22a]. -body [BHGPC21, XJ20]. -bounded [MZLS22]. -bubble-sort [GLW22, Guo23]. -calculus [AM20a]. -Caputo [BGRM21]. -Catalan [ZZ23]. -choosability [HZ23b]. -circulant [PP23]. -colorable [PWL23, ZCW24b]. -colored [HLSL22]. -component [YZM24]. -conforming [TCPF24]. -connected [CCHZ23, DW20, LW21a]. -connectivity [LW23b, HM20a, Kun24]. -core [WLL^{+}22]. -cube [ZHFL23a]. -cubes [DG20b]. -cycles [LR22, WWY24]. -D [MTRA22, FRW^{+}24, LW20a, LL23b]. -deformation [LLH21, RLX22]. -degree [SZZ24]. -dim [LR21]. -dimensional [ALM20, CZL20, GLZB20, ÖY21, ZLA^{+}21, Zha24b]. -dissipative [LR23]. -Distance [YCW23]. -divergence [BD20a]. -DOF [WWT24]. -dominating [MW20]. -domination [CHJW23, HCSJ24]. -Drazin [CS24c]. -eccentricity [LLSY23, YL20b, YLH22]. -edge-coloring [SS24b]. -eigenpair [LM22]. -eigenpairs [LM22]. -eigenvalues [HLX21, HWLL21, Zha23a]. -Existential [TCG20]. -exponential [LZZ24a]. -extra [YZM24]. -factors [GPW24, CDH23]. -flag-transitive [HLF22]. -flow [LHZ21d]. -fold [LWL^{+}20]. -free [CCHZ23, CY24, SLL24, YLY24]. -fullerenes [SZZ22]. -gain [ZLJ^{+}21]. -Galerkin [KPP22]. -gon [HPL21]. -good-neighbor [LZ20c, ZLC^{+}24b]. -graph [LLZ20c]. -graphs [ZCT22]. -Hamming [WWZ20a]. -hop [JTW^{+}23]. -hued [CFL^{+}20, LLL23b, SLL24]. -hull [LCL23]. -independence [LW21a]. -induced [KKH23]. -inverses [WC22c]. -irregularity [DS23]. -join [MW20]. -junctions [CDD23]. -kernels [HLSL22]. -Kneser [LCL23]. -labelling [HJL^{+}23]. -Laguerre [ÖÇÖ23]. -Laplacian [Baz20, CT24, HBCA21, JZ22, LL22a, WLW20]. -learning [WYCL24, WDZL24, WFKL24, YZPL24]. -Lipschitz [ZZX20a]. -lucky [KRE20]. -matching [CLL23]. -matrices [ZZL20b, ZV20]. -matrix [HLX20]. -matrix-based [DWZ20]. -means [CZMX21, JWY20]. -method [LCL22c]. -metric [GN20a, KRT21, SVW23]. -minor [CFL^{+}20]. -Mittag-Leffler [LJAM24]. -moments [CRDD21]. -norm [ZYySX23, STA24]. -normal [LGX22, YLL22]. -optimal [KWH^{+}23]. -orthogonal [RGAM20]. -out-of- [LHW^{+}23]. -packing [HH23b, YW22a]. -partite [LY22a, XYZZ22, ZLL20]. -path-connectivity [ZHFL23a]. -path-quasi-transitive [DLM20]. -peaks [ZIML20]. -periodicity [LKSS21]. -planar [LLL23b]. -plane [DHZ23]. -player [GLLZ24]. -policy [BSCL20]. -power [CZ21b]. -property [LL23c]. -pseudo [HWC20]. -quasi-transitive [HCJC21]. -rainbow [DDS24]. -rational [ZZ23]. -regularization [PX20]. -representation [ZZS^{+}20]. -restricted [LM21]. -schemes [QS20]. -SDEs [YC21]. -smooth [GKK^{+}24]. -spectra [YWX20b]. -spectral [CLZ23, DW20, LZ20b, QLM21]. -spline [Rom21]. -splines [NGGD23]. -stable [RVV24, SK20]. -star [CLL23]. -state [ZLLX22]. -strategy [INT20]. -subconnected [QLZL20]. -subdivisions [HLSL22]. -subtree [LYhH^{+}24]. -symplectic [BIJ22]. -systems [GI23]. -tensors [DLQY20]. -Toeplitz [AB20a].

-token [ZZLK23]. -total-rainbow [MZ22]. -type [RKJ20, URLC21, HLX21]. -uniform [WX20a, WLZD24, ZLL20]. -vector [ALM20]. -wave [WS21]. -weak [SNMM21].

/fixed [GWZZ23]. /passive [WLHW23].

1 [Ano20a, Ano20b, Ano20c, Ano20d, Ano20f, Ano20e, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21g, Ano21f, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22g, Ano22f, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23g, Ano23f, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24g, Ano24f, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, GL23b, LWNZ22, SWCT21, YI22]. 12 [GNA22a]. 127775 [CF23b]. 127923 [ZBA^{+}23a]. 15 [Ano20l, Ano20m, Ano20n, Ano20o, Ano20q, Ano20p, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22s, Ano22r, Ano22u, Ano22t, Ano22v, Ano22w, Ano22x, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano24m, Ano24n, Ano24p, Ano24q, Ano24s, Ano24r, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x]. 19 [DKKJ21, KKS22b, MGFF21, NGC21, XYR^{+}23].

2- [XLL24]. 2/cancer [EA21]. 2020 [Ano20a, Ano20l, Ano20b, Ano20m, Ano20c, Ano20d, Ano20n, Ano20o, Ano20f, Ano20q, Ano20e, Ano20p, Ano20g, Ano20r, Ano20h, Ano20s, Ano20i, Ano20t, Ano20j, Ano20u, Ano20k, Ano20v]. 2021 [Ano21a, Ano21m, Ano21b, Ano21n, Ano21c, Ano21o, Ano21d, Ano21p, Ano21e, Ano21q, Ano21g, Ano21s, Ano21f, Ano21r, Ano21h, Ano21t, Ano21i, Ano21u, Ano21j, Ano21v, Ano21k, Ano21w, Ano21l, Ano21x]. 2022 [Ano22a, Ano22m, Ano22b, Ano22n, Ano22c, Ano22o, Ano22d, Ano22p, Ano22e, Ano22q, Ano22g, Ano22s, Ano22f, Ano22r, Ano22u, Ano22h, Ano22t, Ano22i, Ano22j, Ano22v, Ano22k, Ano22w, Ano22l, Ano22x]. 2023 [Ano23a, Ano23m, Ano23b, Ano23n, Ano23c, Ano23o, Ano23d, Ano23p, Ano23e, Ano23q, Ano23g, Ano23s, Ano23f, Ano23r, Ano23h, Ano23t, Ano23i, Ano23u, Ano23j, Ano23v, Ano23k, Ano23w, Ano23l, Ano23x]. 2024 [Ano24a, Ano24m, Ano24b, Ano24n, Ano24c, Ano24d, Ano24p, Ano24e, Ano24q, Ano24g, Ano24s, Ano24f, Ano24r, Ano24h, Ano24t, Ano24i, Ano24u, Ano24j, Ano24v, Ano24k, Ano24w, Ano24l, Ano24x]. 22 [LOVHC23]. 2D [SSF24]. 2D/3D [SSF24]. 2nd [BCS24]. 2nd-order [BCS24].

3-stars [XLL24]. 342 [RO21, WDG21]. 362 [PPS24, RRR20]. 367 [KWM20b]. 370 [Elg22]. 371 [WCY20a]. 394 [Oru22, SZ22]. 395 [AV21a]. 3D [SSF24].

400 [MV23, WHM^{+}21a]. 403 [KFM21b]. 404 [LPP21a]. 412 [GNA22a]. 417 [KA23]. 419 [CSG^{+}23]. 420 [NHK^{+}23]. 425 [MM24]. 427 [XH23]. 440 [AACGTB24]. 443 [CF23b]. 448 [ZBA^{+}23a]. 4th [HLX21]. 4th-order [HLX21].

800 [LOVHC23]. 800-22 [LOVHC23].

AA7072 [SIK22]. AA7072-methanol [SIK22]. AA7075 [SIK22]. AA7075-methanol [SIK22]. ABC [DD21c, DRS20a]. Abel [ACH^{+}22, MDSS20]. abelian [LWW23, WLJ22, WF23a]. abilities [LQS^{+}23]. Ability [LJN^{+}22, WHW^{+}21, ZRMD21]. Ability-based [LJN^{+}22]. ability-to-pay [ZRMD21]. Ablowitz [DYA21]. Abnormal [WGW^{+}21, WWW^{+}22, LLZ^{+}20a]. Abramov [AAL^{+}21]. abrupt [LDX20]. ABS [LY24]. abstract [YYF24]. Abundance [Sek23, Xia21]. accelerated [KKLJ20, LLF^{+}23, LW22g, LWZ22c, SSCW22]. accelerating [AEE21, SZCP21, YOF21]. Acceleration [THBPMC22, GZH22, PPS24, RRR19, RRR20, WLB22]. access [FLL^{+}22, HLKC22, YWWY23]. accessible [LL22b, LYL22]. accidental [RS24]. accounts [Chi22]. accumulation [ZZ24]. accuracy [ARÁSY21, DR22, GYC24, HL24b, MH22, MV21b, MV23, SW22, ZZL^{+}24]. accuracy-preserving [MV21b, MV23]. Accurate [GRMQ22, LLP23, ADB20, ADS21, CKL22, CTBB24, DYT21, EBJ20, HWCD22, KWZX23, KM20, KNEM23, KK24a, KY21, LQNA23, NS22, QPHC21, TBRB24, TO20, WKM24]. accurately [RS22a]. achieve [XZZW21]. acid [CS21b]. acid- [CS21b]. acknowledgments [SY21b]. Acoustic [ACC^{+}21, HR23, OKT24, WDYS21, YLC20b]. acoustic-structural [WDYS21]. acoustics [GLC23]. ACQUIRE [dSLV20]. acquired [BJ21]. action [AGGPO23, HR20b, RHN^{+}24]. activated [SWCT21]. activation [QNX^{+}23, SN21, WLHW23, YLCL23]. Active [CHK21, DW22, VBM20, XWWL23]. activity [HLD^{+}23, YR20b, YLH22]. activity-driven [HLD^{+}23]. actuated [LNZ^{+}22]. Actuator [HD20, DYX24, GSRL20, HLL22, HYT21, JZS^{+}24, KMLW22, KM22, LLZ20b, LGX22, LTL^{+}22, LWY^{+}24, MWZX21, SZSZ22, VSM22, WXP21, WYZX22, WX21a, XH22a, Yan24, ZZX^{+}22, ZWYL22]. actuators [LJW^{+}22]. acyclic [YXW23]. Adaptation [TZL24]. adapted [MM22b, BS23]. adapting [TGL^{+}24]. Adaptive [AHM23, AY22, CD20a, Cav22, CYHR22, CCZ24, GSWL23, GLP20, HRLH21, HLLS22, HLL22, HL23a, HJW^{+}23, IFHA21, JY21, JJSY22, KLH24, LLY^{+}23, LR24, LNZ^{+}22, LWH22, LSC^{+}23, LH24b, MWZX21, MWS23, MZFX20, MXY^{+}20, MN22, PX23, SX23a, SJ23a, SPZS20, SCHS22, WXLZ21, WXW^{+}21, WSS22, WA22, WPZY23, WM23, WSXL22, XD21, XYZY20, XZZW21, XLLH22, XWL23, YXYZ23, Yi23, YZDC22, YZY22b, ZSC^{+}21, ZLDY22, ZCZ22, ZD23, ZWW23, ZPDL20, ZZ20b, ZLP21, AS23a, Ahm21, AS21c, ADS23, BTL^{+}22, BRTR21, BMY24, CLCY23, CNZ^{+}21, CLWX23, CZGX22, CXY^{+}23, DDC22, FT24, FLLW23, GSS24, GSL^{+}22, GXZZ20, GL22, GWH24, GL24, GZLY24, HD23, HXZ24, HQK21, HZX22, HXF23, HMZZ20, IP23, JYMP^{+}20, JMC23, KM22, KMS^{+}23, KG22, KBL^{+}24, KY22b, KSW24, LLZ20b, LDX20, LWD21, LWZ21, LZ22c, LZCL23, LCZW23, LYX23, LZZ^{+}20, LZL^{+}21, LL23d, LTB23, LJWD21, LJW^{+}22, LZD23, LY22d, MPX^{+}20, MZLS22]. adaptive [OKN^{+}22, QZH21, QPHC21, RNVC21, RK22, SZSZ22, SYTC24, SY22b, TFL22, TLY24, TWB23, WLQX20, WZZX21, WYLQ24, WZA^{+}24, WM24a, WPX^{+}20, WZXH22, WSSX22, WWW24b, XLZ23, XHWM20, XWL^{+}22, XLZ^{+}24b, XRZH22, YLC20a, YCM^{+}22, YWC24, YWL20, Yoo21, YZ24, ZW22a, ZPY22, ZY23b, ZTW22, ZJLT23]. adaptive-event-triggered [SZSZ22]. adaptive-R [BTL^{+}22]. adaptivity [ZMS23a, ZXF20]. Adding [BBH22, GZCS23]. additive [DWZZ22, WCX^{+}22, XXL^{+}21]. ADER [GÖRT23, LLQG21, VMT24]. ADER-DG [GÖRT23]. adhesive [BM22]. ADI [QL21, JCSH24, LQNA23, QWD22, SM21b, WZ20b, ZZY24, ZC22]. adjacency [GLZA21, GZ20c, LW22a]. adjacency-distance [GZ20c]. adjacent [jSZyW^{+}20]. Adjoint [AT21b]. Adjustable [DXZ^{+}23, HLW^{+}21, HBWQ20, LBX^{+}23]. adjustment [ASV22, HZX21, LOS21, RSXH20, SO21b, SO21c, vdZGO21]. administration [MMG^{+}23]. Admissibility [JZP^{+}20, RFFZ21, SS20, FSLC22, TMW20]. admissibilization [LZZ^{+}23]. admitting [LYJM21]. ADMM [CZW24]. adoption [PPZW20, TTC^{+}23, YZT^{+}21]. ADP [CXGL22, ZSYX23]. ADT [HLW^{+}20, ZSM20, ZJLT23]. adult [ZYZ22b]. Advanced [DBVCP23, CCPT20]. advancement [YS23]. advancements [ZM21a]. Advances [DdAS22, HR20a]. advection [Alf24, DTL22, EFF23, FG24, KDB^{+}23, KK20, LZRF21, ZZX23]. advection-diffusion [KK20, LZRF21]. advection-diffusion-reaction [DTL22, KDB^{+}23]. advection-dispersion [ZZX23]. advertisement [DD21a]. advertising [CZS21]. aerial [XXJtC24, XWWL23]. aeroacoustics [LZLM20, LTL^{+}20, YGM^{+}22]. aerodynamic [IFHA21]. aerodynamics [LZLM20, LTL^{+}20]. affect [HSLS20, JLZ^{+}22]. affects [KLP22]. affine [BMDE21, FJV21, KJL20, LePP24, LDX20, LZHZ24, LLWW22, MFS22, PHL23, PKL23, RD24, XZZC22, YDXH22, ZPDL20, ZEA22]. affine-splitting [RD24]. after [JJL23, RS24]. Against [HD20, CLD^{+}24, DTWJ21, DS24a, DWS24, GJS^{+}23, GL24, JLL^{+}20, KKPS23, LLLS20, LLWW22, LJWD21, LJW^{+}22, LY22d, MFS22, NK22, SWS^{+}22b, SLPH22, WZXY20, XWZC21, YCZ^{+}23, ZW22a, ZY23b, ZGZ^{+}23, ZWZS22]. age [AALMLM22, GGW21a, JC22b, JKJC21, LFA22, NS20, RHNT21, Shi21]. age-dependent [GGW21a]. age-structured [AALMLM22, JC22b, JKJC21, Shi21]. Agent [GLP20, BRTR21, CZLH20, CDZW21, CMYF21, CZ22b, CM20b, CYG20, CYL^{+}22, CLC^{+}22, CTF24, DXS^{+}20, DWZZ22, FJLG22, GZCS23, GWZZ23, GS24, GHS^{+}21, GZLY24, HQS^{+}24, HLW^{+}20, JLP23, KKPS21, LJL21, LZWC22, LDXZ23, LDCZ24, LNC24, LSWS21, LXK23, LY20b, MTFD24, MP22, MPC^{+}24, MYY21, PHS^{+}20, QYLW24, SCLD21, SZG21, SGX21, UU21, WHH^{+}20, WS20, WLQX20, Wan22b, WDTF22, WLS22, WC23, WZW23, WJDL24, WYLQ24, WZXH22, WWW24b, XLZ23, XZ23b, XWL^{+}22, XLZ^{+}24b, YLC20a, YY20a, YFW21, YXMW22, YDXH22, YXW23, YMYR21, ZL20d, ZZCW21, ZDZS21, ZZLL21, ZCZ22, ZDC22, ZLZ^{+}21, ZPY22, ZZW^{+}21c, ZWY^{+}22, ZD24, ZPW22]. agents [BJ21, GLW24, KSZ23, LCS23, SX23a, Zha21d]. aggregate [RCCL21]. aggregation [CGSRG20, GPRV21]. aggressiveness [LCZ^{+}20]. agreement [WCY^{+}24]. AHP [BCCI20, FZG23]. AIDS [KEiS23]. AIDS/HIV [KEiS23]. air [HTZ^{+}23, QDL^{+}21]. aircraft [DYW^{+}20, HQK21, IFHA21]. Al [AGC20]. Albert [PPS20]. Algebra [ZYZ22c, CCC23, SM22]. Algebraic [CIHP^{+}21, GJWV24, LTWY24, WLB22, BPG24, KX23, THBPMC22, VBMB21, YXAY21, ZJX21, BCHRS23]. algebras [Bia21, FCeFZ21, ZYW21b]. Algorithm [AT21b, GLX22, KMS^{+}20, Agg22, ARISV20, AM20b, ASAK21, AHL21, BHK^{+}22, BS20, BD20a, BL22, Bia21, BZ21, BGRM21, BM21, DESZ23, CCG^{+}22, CB21, CWZ20, CQ24, CZZP24, CH20b, CSV20, DS21b, GZH22, GLWX24, HR20a, xHhPjW^{+}21, HMS20, HLLY20, IR22, JDLZ21, JWK22, LSC^{+}21, LePP24, LCD^{+}20, LJW^{+}21, LR20, LAH22, LHZ21c, LWL24, LS22a, LM20d, Lyu23, MGM21, Mar24, MGFF21, MQ22, OKN^{+}22, PKJB22, PZ20, RLK^{+}20, SL20a, SV23a, SRN^{+}23, SM22, SZCP21, SFL^{+}22, SL24, TZD22, TLPY20, TVY21, TKF23, WCG^{+}20, WLW20, WCP20, WHHD21, WJK20, WSZ20, Xia22, XBC22b, YM21, YLH20, YH24a, YL24b, ZIML20, ZDC22, ZTYZ23, ZS22c, ZS22b, ZZD22, ZZY24, ZWZ^{+}22, BS23]. algorithm-optimized [CZZP24]. Algorithmic [BBW21, Gas20]. Algorithms [GRMQ22, HS23a, LYhH^{+}24, BSCV24, Che21, CTV21, DTP21, Du24, ES23, FJJJ21, GAA^{+}24, GM21a, GJWV24, HLP24, HCY22, KRRS22, KSW24, LAS23, RST21, TLPT21, ZW21c, Zou23]. all-at-once [LW22g]. Allee [Gök22, LD20d, LW21b, PSE24]. Allen [AKMR21, LL22a, PW22a, TYCK23, UK23]. alliances [WBV24]. Allocating [ZLW^{+}21b]. allocation [BY23b, CZW24, HRZ^{+}23, JP20, JLL^{+}20, LLCS23, MMR23, SWX24a, ZYZ22a]. allow [QDL^{+}21]. Almost [CFWZ20, LJZL22, BMY24, GGY22a, LFLP21, Prz24b, RC22, RM23b, Zha23b]. along [GPRV21]. alternating [BD20a, CH24a, Che22, Dol21, DHZ^{+}20, HZT^{+}21, LYJM21, LZRG21, PMP21, TZLL24]. altruistic [JLWX21, NCY^{+}20]. Alzheimer [LZ22a, WHW^{+}21]. amensalism [LW21b]. American [HANP23, LYZ22b]. AMG [DVC23]. among [GAMP20, LC22b, NS20, NZLW22, SGL21]. amyloid [GPRV21]. Anacondensed [CDR22b]. analyses [HXL^{+}20]. Analysis [AKDR23, ADS21, AT21b, CLL22, GGW21a, GYC24, GRMQ22, HZJL24, Jan22, LRCTR22, LFLL21, LX23a, LZ24c, LD20d, LF24, LZ22f, MCSD21, MMS21, NWM23, PS21a, RCCL21, RC22, SIK22, SR23, WTCC20, XWM23, XHL20, ZL20c, AS20, AK21, AR22, Alf21, Alf24, AFGG20, ANS^{+}20, ASS^{+}23a, AVJ24, AIAFvH21, AHEGM23, Ask22, Azi21, BF22, BFQ20, BEe21, BM24, BJK^{+}22, BP23, BBM21, CFR22, CZ21a, CHY22a, CXYZ23, CLC24, CLZ24, CTBB24, DD24a, DKKJ21, DG21, DE21, DYA21, DS22b, DWZ20, DTL22, DMP21, ED20, FLZ23, FLWL21, FGM24, FdLPC^{+}22, FSLC22, FJ24, GZ20a, GNSZ22, GWSC23, Gas20, GTWL21, GOO20, GPRV21, GM21b, Has20, HX22, HWCD22, HR23, HD22a, HL24b, HQ22, HH22, HSY24, HVZZ23, JA23, JYMP^{+}20, JNY^{+}21, JW23, JLD^{+}24, JWK22, JM23a, JGXL23, JZP^{+}20, JQJY22, JJL23, KX23, KKR21, KET22, KHA22]. analysis [KIJK20, KG21, KMD21, KLNN23, KY21, LD21a, LSZ^{+}23, LMZ20, LSPK20, LKP22, LD20b, LHW21, LLS21a, LLY^{+}22b, LC22a, LW22d, LXL22, LW23a, LZ20a, LQ24, LLD21, LWL21, LCL22b, LWQW22, Liu22b, LCXJ22, LV20, LWJ21, LYN22, Lu22, LOVHC23, LLT^{+}20, MGS20, MGM21, MYC^{+}23, MF21, ME21a, MSDR21, MMUS22, MB24, MG21, Miz21, MJW21, NE21, NTLNX24, PKL23, PZS24, PW22a, QH21, QL20, QXSW22, RAB24, RMO22, RSVL21, RFFZ21, RCG21, SO21a, SG20a, SS24d, SNR^{+}22, SL20b, SW20, SGX21, SWW21a, SW22, SZ23, SSF24, SZ24a, SS22c, SK24b, SKL23, SS23b, SLHX21, SYX22, SFC23, TNTL21, TWR20, TZ23b, TM22b, TXZ24, VJA^{+}22, VBP22, VSM23, WHC20a, WZY20, WLW20, WGM^{+}21, WHW^{+}21, WWWH21, WHL21, WBV22, WJL22, WL23, WKM24, WOB21, WS21, WPJ22, XWLL20, XYH^{+}24, XAZ^{+}21, XLG^{+}21, XZQ^{+}22, XZ23c, XYR^{+}23, XYlW^{+}24, YWZT20, YZJ^{+}20, YWS24, YYF24]. analysis [YCWH23, YXX^{+}22, YDSW22, YD22b, YWJ23, Yu23, YLCL23, ZHL24, ZF24, ZLX20, ZZL^{+}20a, ZZCL20, ZMD^{+}21, ZWS^{+}22, ZSFS22, ZMJ^{+}23, ZLZ24a, ZDZ24, ZQ24, ZLJ20, ZLJ^{+}21, ZGG^{+}22, ZZS^{+}22, ZLZ^{+}23b, vdMvRS21]. analysis-based [XYH^{+}24]. Analytic [KRF20, wXhGZ20, HSM21, wX20b]. Analytical [GA20, HCX^{+}24, KWM20a, KWM20b, LNO23, QCARL20, RY24, WMQ21, XLZL23, ZyZW21, BPG24, CCH24, HAFM21, KSPC22, KY21, LQLF22, LR20, MV21a, MMMBRS20]. analyze [ACVHC22]. analyzing [ZMNR^{+}22]. ANCF [ZZL^{+}24]. and/or [SMW22]. angiogenesis [ALL23]. angiograms [CCG^{+}22]. angle [RCG21, SS22c, WWZ^{+}23]. angle-ply [SS22c]. angular [DVS21, GZRM21]. Anharmonic [NNB21]. Anisotropic [DH23, HSM21, VGR^{+}21, BHL23, CK21, CFPM23, DK22c, MKKG22, SSR23, WDC^{+}23, WC21, XCGN24, Yu23]. anisotropic-isotropic [XCGN24]. anisotropic-porous-piezoelectric [MKKG22]. anisotropy [AZD^{+}24, CS21b, Miz21, Qi22]. ANN [Hoj20]. annealing [WGJK21]. Annihilation [CLUR22, YZW20]. ANNs [CCdS^{+}23]. annular [KTEM22, YZC^{+}22]. annulus [XAZ^{+}21]. anomalous [AD21a, FTML22]. anonymous [YZY22a]. answers [CSCZ24]. antagonistic [MS24, ZL20d]. Antennae [KMS^{+}20]. anthracnose [FMHDB22]. Anti [HLLH23, LCLS20, LLZ22b, XLL21a, ZWWH21, ZW21a, BF20, CS24c, GZL^{+}23, HZW21, HRA^{+}20, JWMW23, LW22a, PLH22, RLY20, SZZ22, SS23b, WZLC21, XZQ^{+}22, YLH22, ZYSW22]. anti-adjacency [LW22a]. anti-bump [HZW21]. Anti-disturbance [LCLS20, ZWWH21, GZL^{+}23, XZQ^{+}22, ZYSW22]. Anti-Eulerian [ZW21a]. anti-forcing [SZZ22]. anti-HIV [YLH22]. Anti-Ramsey [HLLH23, LLZ22b, XLL21a, JWMW23, PLH22]. anti-surfactant [SS23b]. anti-swing [RLY20]. anti-synchronization [WZLC21]. anti-triangular [CS24c]. anti-tridiagonal [BF20]. anti-windup [HRA^{+}20]. antibiotic [AMA20]. anticipative [PKG^{+}20]. Antidirected [YLM24]. antigenicity [LXCZ21]. antimicrobial [IG21]. antimicrobial-resistant [IG21]. antiplane [CMMN20]. antiretroviral [VMCG20]. antisocial [QZZ^{+}23]. antistrong [YM23b]. antitriangular [Sin20a]. antiviral [NNKH21, YI22]. anxiety [ZTLW24]. any [LU24]. AO [HC21]. aperiodic [LWD21, LYXZ21, LYL22, YYWL23, YSL21]. Aperiodically [LK21b, CFWL24]. aperture [KL22]. apex [CLB21]. Apollonian [LX24b]. Appell [ÖÇÖ23]. Appl [AACGTB24, LPP21a, PPS24, WDG21]. applicability [ACGG20]. Application [BE22, CH20b, DDE21, DCR^{+}20, FQHS24, EVB^{+}21, HT23, KM20, KY22a, KAS21, LKP22, MYC^{+}23, TS23, WPE20, XAZ^{+}21, YXZ^{+}24, AFGG20, AS23b, AACGTB23, AACGTB24, ALM20, BHK^{+}22, BDKZ21, BL22, BSG23, CPS24, CSS^{+}22, CT24, CG21, CZS21, CLLA23, CCH24, DEHFF24, GHND22, GN20a, GWZ^{+}20, GLD21, GT22b, Har24, HQK21, HH22, Hua24, JYMP^{+}20, JSZ20, KMM^{+}24, KDB^{+}23, KKR21, KKS22b, KIB20, LaGBT22, LYN^{+}21, LZRF21, LZL^{+}21, LLX21, MZL24, Pap21, PSSU21, Pop20, PMG20, QXX21, RSA24, SKWW21, STH24, TLPT21, TM22a, Voe23, WLG^{+}21, WWQ22, WLZY21, XKW^{+}23, YSL24, YWS24, YTLV22, YWX20b, Zha21b, ZL22d, ZSJ23a, ZQ24, ZH22b, ZS22d, ZWYL22, ZZW^{+}21c, PGPCK21, VMZDV20]. Applications [GGBEZRI23, HD23, Ily23, IÖÜP23, JZZ^{+}22, PKL23, SG20a, ZYySX23, AB20a, ARÁSY21, BMDE21, BGB24, CLCY23, CP20, DP24, DK22a, FZB24, FZ20a, FRR^{+}22, GW22, GA22, GJWV24, HCS20, HWC20, HLX21, HW24, IR22, KRF20, KT20, LPS21, LLL21, LLP23, Liu22a, LY23b, MGS20, MJZ21, SYW21, SRN^{+}23, jSZyW^{+}20, ySyS22, TCA23, TITN20, WW22b, XLZL23, YGM^{+}22, ZY20, ZJWV24, ZSJ23b, ZEA22]. Applied [ABNS24b, AV21a, CSG^{+}23, CF23b, CGMFSHR21, Elg22, GNA22a, KWM20b, KA23, KFM21b, MV23, NHK^{+}23, Oru22, RO21, SZ22, WCY20a, WHM^{+}21a, XH23, ZBA^{+}23a, dSAOG21, ADB20, BKS21, CPMS22, CG20a, IS21, KLS20, Lee23, LD20c, MM24, PPS24, PAY22, RRR19, RRR20, SZWL21, dFdOFdSBC20]. appointed [DYX24]. appointed-time [DYX24]. Approach [HD20, WSXL22, ASR22, AFL24, AMM21, ACVHC22, AVJ24, AAL^{+}21, AO21a, AGCGSC21, BRTR21, BSST20, BJK^{+}22, BK21, CCC23, CS24b, CWH24, CJ20, CL20, CCH24, CCF24, CWJF21, CXGL22, DD24b, DFP^{+}24, DMSS20, DYA21, DD21b, DS24b, DZX23, DL20c, DS21c, DS21d, Ema21, FC21, FZ22, GE21, GZX20, GNSZ22, GDFNNR22, GXZ^{+}24, GBCP21, GZLY24, HTF20, HWC23, HYS24, HBCA21, HT23, HXL^{+}20, HSZ20, HJW^{+}23, HRA^{+}20, JLP23, JLD^{+}24, JWK22, JMC23, KAT22, Kop22, KY21, KLPL20, KJL20, LZH^{+}20, LW20a, LZSC22, LM24, LNC24, LMZ^{+}21, LD21b, LLCS23, LLW24, LLT^{+}20, ML22, Mar22, MDSS20, Mon20, MZYW23, MJW21, MJ23, NGGD23, NLP^{+}22, ON21, PLH^{+}24, QHC21, QLC^{+}24, RS23, RR07, SCD23, SCC24, SRK^{+}20, SM21c, Sun22, SL24, TYCK23, TMW20, TAU^{+}22, TXZ24, VSM23, WBY^{+}21, Wan21a, WHCW20, WHL22, XZZ22b, XYZY20, XLZ^{+}24b, YZC^{+}22]. approach [YXXZ23, YC21, YSW^{+}22, YL20c, Yu21, ZM21a, ZTLL24, ZWZ^{+}21, ZC23a, ZLW23, ZTLW24, ZM21c, ZL21c, ZeF21, ZEA22, dSLV20, vdZGO21]. approaches [DD20b, FFL23a, FZS21, GA20, LKK^{+}23, Mar24, SMM^{+}24]. appropriate [LLLC23]. Approximate [DHM23, JS20, MSN21, RT22, ADLU23, DMSS20, GBF21, HA20, Kum23, LJY22, MKP20, OiEF22, ZZL^{+}24]. approximates [GAR22]. Approximating [OK21, JS20, KKS22b]. Approximation [CKN21, DTP21, FKS23, GA22, KS21, LK21a, MBMA20, MPRS24, ARMM23, AM20a, AS21c, AA21b, AAL^{+}21, ADB20, Bon22, BE22, CN23, DMPP21, DN21, DMP21, EHVMM24, Har22, HKU21, Iza21, IS21, KNN20, KL21, KSR21, LPS21, LHZ^{+}21b, LY22b, LY23c, MY22, NWM23, NYHC24, OT20, PSSU21, PS21b, QY22, QCARL20, QQLZ22, RZZ21, RD22, RFFB23, VZ23, Wan21b, WSG^{+}21, WL24b, XGF20, YCEW21, ZXF20]. Approximations [LLT22, AM20b, BMY24, CEPZ22, LH22, LLP23, LLZ22c, PDK23, RM23a, RM23b, WXHZ23, YL22]. April [Ano20a, Ano20l, Ano21a, Ano21m, Ano22a, Ano22m, Ano23a, Ano23m, Ano24a, Ano24m, CSG^{+}23]. aquaculture [ZZCL20]. Arbitrarily [JL22, BBB24, BBK^{+}24, JY21, JDLZ21, YWJ20a]. Arbitrary [BDRM23, LNZ21, TM22a, CSQT21, ECS21, HSY24, LCD^{+}20, LMS24, MCD21, NZ20, PD23, TA21, WDC^{+}23, WZCF21, REC^{+}23]. Arbitrary-Lagrangian [BDRM23]. arboricity [CW23]. arc [CZ21b, FKVZ21, KP22, KPS22, SS24a]. arc-colourings [KP22]. arches [Bab22, DC21]. Archimedean [FC21]. arclength [Fre22]. arcs [GZCS23]. Area [xHhPjW^{+}21, GSL^{+}22, LHZ^{+}23, LLQW24]. argument [Bon22]. arguments [CGJ21, SGL21, Wu20a]. arid [XLL^{+}21b]. arising [ÅCL24, LQNA23, NA21, WJL22, YWZ23, ZV21]. Arithmetic [VPV^{+}21, ZSJ23b, LPL21a]. Arithmetic-geometric [ZSJ23b]. Arrangement [LM20a, LZX22a, ZW21b]. array [TJAJAG21]. arrays [AB20a, SRK^{+}20, TJIM20]. arrhythmia [BAL^{+}24]. arrival [QPHC21]. arrivals [ZRSK21]. ART [GZH22]. arterial [PPS24, RRR19, RRR20]. arteries [Kri22]. artery [NE21, PM20, RB21]. Article [CF23b]. articles [LKK^{+}23]. Artificial [LH22, LM23, AS21a, BAL^{+}24, LWZ22c]. ary [DG20b, ZHFL23a]. Asian [LSCS21]. aspect [RWGS22]. Aspects [GLX22, GLWX24]. Aspiration [DGL^{+}23, LZCL23, HZX21, WLZW21, WWZW23, YSW^{+}21]. aspiration-based [HZX21]. assemblage [Voe23]. Assessing [LKK^{+}23]. assessment [CCG^{+}22, HGKD22, KY21, LHW^{+}23, LZY^{+}23b, dDMMTH22]. asset [JP20, KNN20]. assets [DD21a]. assigned [YZDC22]. assignment [LaGBT22]. assimilation [DR22]. assisted [GS22]. associated [AJIR21, DFX24, FF21a, FX20, Gu20, HCS20, LWL^{+}20, ÖÇÖ23, UUS21, WC24, XZS20]. associative [WHW^{+}21, XLYA21]. assortative [LXF21]. assortment [JLWX21]. assumption [MMDD21]. assumptions [Ili22, NL20, ZFS22]. asteroid [FZG23]. Asymmetric [GLWW22, QQWJ22, YZY22a, Zha24a, Ask21, CHP23, GYC^{+}22, GDS24, JLP23, KTZ24, LMT24, LLX21, TG20, WLZW21, WGZZ22, WWT24, Wri24, ZXZ24]. asymmetric-exit [GYC^{+}22]. asymmetrical [LJN^{+}22]. asymptomatic [CRTVG^{+}23, CH22b, GM21b, HCY21, WMCZ21]. Asymptotic [CUA^{+}22, DF20, JS22, LC21, LLD21, Ma21a, RB21, SX24, VBP22, WBV22, XAZ^{+}21, ADS21, CKW22, GST22, HWC20, LKSS21, PR20, RSVL21, WSS22, WPX^{+}20, XD21, XZZW21, XYLF21, ZZS^{+}20, ZD23, ZLZ24a, ZZW21b]. Asymptotical [CLLA23, LHW21]. Asymptotically [FMP20, HW24, WNG20]. Asymptotics [KL21, XZ23a, XGF20, BPRR22]. Asynchronous [CL24b, LLZ20b, LTW22, LSM23, NK21b, NK21a, SM23, WZC^{+}22, WLL^{+}24, WCZ^{+}21, WCZ^{+}22b, YLK^{+}21, ZM21c, CMYF21, CWJF21, DLS^{+}23, FRW^{+}24, FFL23a, JWC20, LL22b, LYL22, RJHC24, TQZ^{+}24, WCZY20, XH22a, XCW^{+}22, XWZC21, YM23a, YM20, YD22b, ZZLL21, ZCZW23, ZWZS22]. asynchronously [QQY^{+}23]. atherosclerosis [HT23]. atmospheric [CKWY23]. atrial [UMT22]. attaching [LG24]. Attack [ZZX^{+}22, CC22, DXZ^{+}23, GZ20a, HQK21, KB21, LWM^{+}21, LLZB24, LSC^{+}23, SY21b, WYZX22, WM23, WS23b, WLL^{+}24, XRZH22]. attack-detection-based [LLZB24]. attack-resilient [HQK21]. Attacks [HD20, CSS^{+}22, CMYF21, CL24b, DZL^{+}22, DTWJ21, DS24a, FT24, FSS20, GY22a, GSL^{+}22, GHP24, GZWX23, GJS^{+}23, GL24, HXWC24, JLP23, KKPS23, LWD21, LZWC22, LHX^{+}22, LTL^{+}22, LL23b, LDCZ24, LLZB24, LGC^{+}24a, LLLS20, LYL22, LLWW22, LWY^{+}24, LY22d, MCD22, MZY^{+}22, MFS22, OPL22, PGYW23, QZH21, SAB23, SWS^{+}22b, SCHS22, WZXY20, WCM^{+}21, WLL^{+}24, WSDW24, WQT^{+}24, WCZ^{+}21, WYS23, XD23a, XZZ22b, XWZC21, XXJtC24, XZQ23, YCZ^{+}23, YXZ^{+}24, ZDGL21, ZQP^{+}22, ZW22a, ZCS^{+}22, ZCD23, ZLC24a, ZY23b, ZGZ^{+}23, ZWYL22, ZCZW23, ZWZS22, ZPW22]. attenuated [DVS21]. attenuation [AHS20, LX20a, MMK^{+}23]. Attitude [DYW^{+}20, CXY^{+}23, JZS^{+}24, XCLW21]. attitudes [HLD^{+}23, WPH20]. attraction [MFA21, TCZ^{+}20]. attractiveness [DDH22]. attractor [AGA20]. attractors [LX22, LWX24, YL22, YW21]. attribute [YCZ^{+}20, ZHLZ21]. attributed [DVC23]. attributes [dAGSFM22]. auction [MZW^{+}21]. augmentation [CGTU20]. Augmented [CG20b, KLPL20, LLZ22c, BS20, BM24, CH24b, CLB21, JL21, LY20b, MD24, WLD^{+}22, XYKL20, ZYC^{+}24, ZZW21b]. August [Ano20b, Ano20m, Ano21b, Ano21n, Ano22b, Ano22n, Ano23b, Ano23n, Ano24b, Ano24n]. auto [MSE24]. auto-convolution [MSE24]. autocorrelation [Boc22, WY24]. autoencoder [DWT^{+}21]. Automata [TZL24, AZD^{+}24, LCZ^{+}20, ZXC20a]. automatic [LSZ^{+}23]. automorphism [YF21b]. automorphisms [BDK23, ZZLK23]. autonomous [AMF20, FLZ23, FLC^{+}22, HT20, Kum23, LKSS21, LYX23, LCL22b, LWY^{+}24, QQ23, SCD23, WWW^{+}21]. autorun [KIJK20]. auxiliary [TYCK23, Wan23a, ZZLV23].

Average [HGD^{+}21, ARÁSY21, GNSZ22, HMZZ20, LLSY23, Xia21, YWWY23, YL20c, ZeF21]. Averaged [FRRS24]. aversion [HSLS20, HWZZ23, KLP22]. Avoid [LLWY22]. avoidance [WC22b, YP22, ZST22]. avoiding [WLY20]. aware [HDP^{+}21]. awareness [CGW^{+}20, DKK20, PLQ^{+}22, WMCZ21]. AWM [LCD^{+}20]. axial [ASS^{+}23a, ZZL^{+}24]. axial-thermal-electrical [ASS^{+}23a]. axially [CH24a]. axis [XJY24, ZZW^{+}21c]. axisymmetric [MMUS22]. Azzalini [HANP23].

b [DFP^{+}24, CJ21, EFJO20, KG22, LR20, RG22, WL24b, WYLZ24]. b-coloring [DFP^{+}24]. B-spline [KG22, LR20, RG22, WL24b]. background [TTCY23]. backlash [CZZP24]. backlash-like [CZZP24]. Bäcklund [ySyS22]. Backstepping [YW22b, LLY22a, WGZZ22, XYZY20, YCM^{+}22, YWJ20a]. backtrack [CL20]. backtracking [HR20a, LKX20]. Backward [AK21, LX22, RT20a, ZY20, BPR21, BK21, DD24a, HWWW23, Lee23, LZRF21, LLX21, MQ22, Ten22, WWY21, ZXZ24, ZRL^{+}24, ZJX22b]. bacterial [XZS20]. Bakhvalov [LZ22e, ZL20a, ZL21a]. Bakhvalov-type [LZ22e, ZL20a, ZL21a]. balance [GBCP21, JT24, LZ20a, RSZA23, ZZCL20]. balanced [BFLR23, CCCdLMR23, CG23, CZ22b, CKWY23, ECS21, GGL^{+}23, GdFPF24, GBCP21, DDCM23, KPG22, LLZ23, LLQG21, LC23, LCL22c, LZ22f, MZ23a, RM23a, RM23b, RYG21, TCGR24, TZ23b, WHC20b, YTC23b, ZLL20]. ball [OMS22]. balls [KH23]. BAM [AAJ23, ANS^{+}20, HZ23a, XWM23]. Bampton [GLKrH20]. banana [TCGTT21]. bandit [XZY^{+}23]. bands [VGdFCBA20]. Barabási [PPS20]. bargaining [KSZ23]. baroclinic [ZZY^{+}21]. barotropic [DGP^{+}23, ZZY^{+}21]. barotropic-baroclinic [ZZY^{+}21]. barrier [WJL^{+}24]. barriers [DG22b]. barycentric [BE20, LHZ21c]. BAS [KMS^{+}20]. Based [AHM23, BY23a, CNZ^{+}21, DJW20, DDGBGS^{+}24, GLP20, MCSD21, PGL^{+}23, TZL24, AF21, AS23a, Ahm21, ASUH21, AMM21, ARST21a, AFG21, BHK^{+}22, BS20, BDH^{+}23, BR20, BD20a, BCR22, BY23b, BCS24, BE22, BP23, Bri20, CDZW21, CGG23, CCC20a, CPMS22, Cav22, CDP23, CS21a, CL22, CMYF21, CZ22b, CJ22, CH24b, CSW20, CZL20, CZK^{+}22, CHH22, CWY^{+}24b, CTF24, CHL24, CHHC24, CSS^{+}24, CHW^{+}24, CFWL24, CZZP24, CH20b, CCC20b, CDD23, CFA^{+}23, CFPM23, CMR20, CWL^{+}21, CZGX22, CRTV22, DdlCS24, DD24b, DKKJ21, DXS^{+}20, DWT^{+}21, DZ21, DE21, DHM^{+}20, DL20a, DWZ20, DQC^{+}21, DL21b, DK22c, DXZ^{+}23, DBS^{+}23, DSZ^{+}23, DYX24, DYW^{+}20, DL21c, DH23, DH24b, FGMS21, FZG23, FLZZ24, FJFS21, FHFS24, FLH^{+}21, FZ22, FLHH23, GZX20, GGW21b, GWZ^{+}21, GPH21b, GYM21, GJ21, GGY22a, GNSZ22, GG24, GMH^{+}20, GLWL22, GW21a, GI23, GV23, GBCP21, GZ20b, GXZZ20, GLD21]. based [GX23, GWH24, GL24, GZLY24, HYÇK22, HC20, Han20, HLW^{+}21, HZX21, HWC23, HXZ24, HKL22a, HRLH21, xHhPjW^{+}21, HQK21, HP23, He23, HC24, HKL22b, HZX22, HM20b, HOM22, HDQP23, HXWC24, HSY24, HL23b, IÖÜP24, JA23, JC22b, JZS^{+}24, JGXL23, JTC^{+}24, JLOR22, KMM^{+}24, KKPS21, KGPS22, KET22, KWW22, Kun24, LHET23, LSZ^{+}23, LLZ^{+}20a, LPP21a, LPP21b, LKP22, LPK22, LZY^{+}23a, LZW^{+}20a, LDZ20, LW20a, LCZ^{+}20, LXLP20, LHZ^{+}21a, LLQG21, LJW^{+}21, LJN^{+}22, LZ22b, LLY^{+}22b, LZDX22, LHZ^{+}23, LWL^{+}23a, LL23a, LD23a, LYX23, LLZB24, LZZL21, LPL^{+}22, LLWY23, LZHZ24, LCY^{+}21, LC23, LKX20, LZZ^{+}20, LSWS21, LZLZ21, LHS21, LHZ21c, LY21b, LY21c, LZX^{+}22b, LGK^{+}22, LCL22b, LWZ22c, LSC^{+}23, LZZ^{+}23, LL23d, LF24, LLQW24, dSLP20, LZD23, LX24a, LWY^{+}24, MPX^{+}20, MZW^{+}21, MCD22, MSZ22, MZ24, MXF^{+}24, MKKG22, MWDD24, Mar22, MFK22, MP22, MR22, MJM20, MDT22, MC22, MT23, NHBJ20, NGC21, DDM24, NL20]. based [NK20, NK21a, NTLNX24, NA21, NSSA20, NZ20, ÖÇÖ23, PWH20, PGZ^{+}24, PW22a, QH21, QXG21, QZWY20, QSW23, QCCW23, RKJ20, RVV24, RF21, SR24, SJ23a, SHWW21, SGPZ22, SWW21b, SWSW21, SM22, SRK^{+}20, SPM22, SV23b, SKWW21, SL20d, SM21c, Sun22, SFC23, STH24, TWL20, TlS^{+}22, TY22, TQZ^{+}24, Ten22, TNTL21, TLPT21, TTC^{+}23, TEGP24, TZLL24, TVY21, TPTA24, TXZ24, UHZ^{+}20, VSM^{+}21, VSM22, VJKZ21, WS20, WCZY20, WCG^{+}20, WYW^{+}20, WZXY20, WGW^{+}21, WXLZ21, WCM^{+}21, WWW^{+}21, WXP21, WCW21, WWW^{+}22, WLB22, WGZZ22, WC22a, WYZX22, WMT22, WCY^{+}22, WQZD23, Wan23a, WY23, WYCL24, WWW24a, WFKL24, WC22b, WWT24, Wet21a, WX21b, WSG^{+}21, WCZ^{+}22b, WHL22, WSSX22, WYS23, XFW^{+}21, XBLW23, XYCG21, XCW^{+}22, XSP22, XYZY20, XWZC21, XYH^{+}24, XXW^{+}20, XDS^{+}20, XTNF22, XZ23c, XFY20, YSL24, YYG20, YCEW21, YWWY23, YDXH22, YL20a, YCZ^{+}23, YSW^{+}22, YYL^{+}23, YM20, YWXL21, YLH22, Yu23]. based [ZS21a, ZLF22, ZTLL24, ZYC^{+}24, Zha20, ZP20, ZLX20, ZZQ^{+}20, ZDZS21, ZZLL21, ZLS21, ZHLZ21, Zha21b, ZF21, ZLDY22, ZCZ22, ZH22a, ZQP^{+}22, ZL22c, ZZLV23, ZCD23, ZL23b, ZLW23, ZG24, ZLZ24b, ZXF20, ZLC20, ZY21b, ZJX22a, ZBFW22, ZJW^{+}22, ZCLZ22, ZQ24, ZV20, ZV21, ZGG^{+}22, ZKLL23, ZSZL20, ZTW22, ZWYL22, ZLC^{+}24b, ZLP21, ZZW^{+}21c, ZWZS22, ZD24, ZHZZ24, ZWZ^{+}22, ZLU21, ZXW20, ZLZ^{+}23b, dSLV20, LZX^{+}22b]. basic [BCS24, GMH^{+}20]. basin [MRR20]. Basis [TC20b, AHB20, BEIR23, CNYC23, DEHFF24, JCH^{+}20, Kar22, LD21a, Luo20, LMS24, MKK23, NK22, NA21, Oru21, Oru22, WL24a, XFW^{+}21, YLC20b, Zha22b, ZF23b, ZHZZ24]. batch [LLYQ20, ZRSK21]. batch-sizing [LLYQ20]. Bateman [Jor23, Jor24]. Bates [GM21a]. bathymetry [LC23]. Bayesian [AMBG^{+}20, BLNT22, JG21, LWL21, LF24, ON21, SCC24]. BC [YZM24]. BDF [YCWH23]. BDF-PPIFE [YCWH23]. be [DML21, GGBEZRI23, SM20, ZCW24b, WDYS21]. beam [AR21a, BSS21, FB20, HXF23, KG21, LC22a, RJM23, SHWW21, ZMD^{+}21, ZZL^{+}24, dFdOFdSBC20]. beam-like [SHWW21]. beams [SLHX21, WXHZ23]. bearings [ME21a]. Beddington [LW21b]. Beddington-DeAngelis [LW21b]. Beetle [KMS^{+}20]. behavior [ABD22, Alf23, AC20, AS21d, BLO22, BSCL20, CPS24, CT21, DK22b, GP23, HWC23, HH23a, HWZZ23, Kab24, KK22, LCZ^{+}20, LLD21, MZY^{+}22, MK21a, MJM20, NZLW22, PWH20, PPZW20, QDW22, SS20, SWFJ20, WJG^{+}20, WDZL24, WZA^{+}24, YZZ^{+}22, YYW^{+}23, ZTLW24, ZLN20, ZLZ^{+}23b]. behavior-based [LCZ^{+}20]. behavioral [TTC^{+}23]. Behaviors [MS24, DZL^{+}24, FW24, GW20b, JLWX21, KTEM22, LZY^{+}23a, LJZL22, NNB21, QZZ^{+}23, SZXL24, WPH20]. behaviour [BDPV22, CSZ24, LFA22, SS23d, Vyn23, WLB22]. behaviours [BBH22, NS20]. being [WLY20, WCXS22]. belief [GF23, YHSZ20]. Bell [FR23]. BEM [LQLF22]. Benard [SKL23, DMP21, KZS20, Ma21a, SNZC20, YW21]. benchmarked [SG24]. bending [Yu23]. Bene [CWZ22]. benefit [PZZF24, ZLZ^{+}23b]. benefit-cost [ZLZ^{+}23b]. benefits [WZY21, Wet23, ZZ24]. Benjamin [HA20, SL20c]. benzenoid [FRR^{+}22, MJZ21, RRS24, SD23]. benzenoids [LYB^{+}21]. Berge [WZ24, LHZ21d, WZ24]. Berge- [WZ24, WZ24]. Bergman [QQLZ22, YHD24]. Bergman-type [YHD24]. Bermudan [AO21a]. Bernoulli [BSS21, ZMD^{+}21, ZZL^{+}24]. Bernstein [BEIR23, HMEDVR24]. berry [FMHDB22]. Bertrand [Ask21, YY21c]. Bessel [BBM21, CP20, HH22, PKL23, TEGP24, WKM24, YI22]. Bessel-Legendre [CP20]. Bessel-quasilinearization [YI22]. Bessel-type [HH22]. Best [JTW^{+}23, MY22, QQLZ22]. Best-shot [JTW^{+}23]. beta [LJAM24, Zha24a]. beta-binomial [Zha24a]. betrayer [YTX^{+}24]. betrayer-prevailing [YTX^{+}24]. better [CNYC23]. betting [WTCC20]. between [AE21, BM22, BBB24, CH22a, CTWG24, CSV20, DE21, Dol21, Gök22, Guo22, HH20, KAT22, LCZ^{+}20, LZCL23, LLS^{+}22, LJH^{+}20, PLQ^{+}22, Rom21, ST24, SHT^{+}21, TWZ^{+}20, VSM23, WZW24, Wu20b, YZAM20]. betweenness [CRTV22]. Beverton [AGA20]. Beyond [Jor21b]. Bézier [HBWQ20, VZ23]. BFGS [NL20]. BGK [BBD^{+}22]. Bi [DESZ23, SW20, WS21, DM21, GLGLH24, HH20, LD21a, PY21, TNTL21, WF23a, XLYA21, SZ23]. bi-block [DM21]. bi-Cayley [WF23a]. bi-circulants [PY21]. bi-dimensional [GLGLH24]. bi-directional [LD21a, TNTL21, XLYA21]. Bi-flux [SZ23]. Bi-level [DESZ23]. bi-objective [HH20]. Bi-wave [SW20, WS21]. Bias [LWL^{+}23c, PGL^{+}23, LKK^{+}23, LSHH22]. biased [ALM22]. bibliometric [SO21a]. BiCGSafe [HS23b]. BiCGSTAB [CHZ24]. bichromatic [SS22a]. bicomplex [ADV23]. BiCRSTAB [CHZ24]. Bicyclic [LPL21b, Das24, LYhH^{+}24, TLCM24, YW20, YJL20]. Bidirectional [SWZZ23, SKK^{+}21, LDZ20]. BIE [BFS20]. bifunctions [FN22].

Bifurcation [AIAFvH21, RCG21, WW22a, ZWS^{+}22, ZT22b, Alf21, Alf24, AFGG20, HZ23a, JCCC20, LHW21, LCLH24, LM20c, MYW21, PGKM24, QLL24, SS24d, WLY21, Wen20, XLYA21, ZMNR^{+}22]. bifurcations [CWC20, KR22, KK23a, LL20b, PJMW22, ZH20a]. big [CLZ24, DD21c, JSCB20]. Bihari [WJC^{+}20]. Biharmonic [Kum24, FWBY23, WFHG20]. bijection [Kun24]. bijection-connected [Kun24]. Bilateral [YSYS24, BM21, JLOR22]. bilevel [BHL23, dDMMTH22]. Bilinear [GW21b, AM20b, RD22, ZyZW21]. bimatrix [IÖÜP23]. binarization [DH23, DH24b, Fen22, JC22a]. binary [BS22c, DS21c, FTML22, GJS^{+}23, KRT21, TLPY20, WWZ20a, Yu21]. binary-valued [GJS^{+}23]. binomial [ÅCL24, Xia22, Zha24a]. bio [CSG^{+}22, CSG^{+}23]. bio-convective [CSG^{+}22, CSG^{+}23]. biochemical [GP22]. biodiversity [HDQP23, ZYW21a]. Bioeconomics [BBAS21]. biofilm [BCLNV20, EE22]. biogeography [Cas23]. biological [SWW21a]. biomedical [GN21]. Biot [SC21]. bipanconnectivity [HZ20b, NX21]. Bipancyclicism [HZ20b]. bipancyclicity [NXL21, WHC20b]. Bipartite [GL22, LDXZ23, LG23, LH24a, WHH^{+}20, WLL21, CZLH20, DQC^{+}21, DWZZ22, GW20a, GWZZ23, GL24, HQS^{+}24, HZ20b, JY23, JMY22, LHX^{+}22, LW23b, LZ24b, MPC^{+}24, NX21, NXL21, PHS^{+}20, PK20, SMQ24, SWS^{+}22a, WM22, XW23, XLZ^{+}24a, YM22a, YLHZ22, ZWC23, ZHW22, ZPY22, ZHFL23b]. biphase [VBP22]. bipyramid [SSP24]. biquaternion [GJWV24]. birefringent [TTCY23, WMQ21, YTZ20]. Birkhoff [RPB23, XLD23]. Birkhoffian [KSW24, QQ23, WHWW21]. Birnbaum [NHBJ20]. Birnbaum-Saunders [NHBJ20]. birth [GN20b, GN22, SRZK22]. birth-death [GN20b, GN22]. Bisection [TPSP24]. bistable [WBV22]. Bit [ZH22b, TEGP24]. bit-level [TEGP24]. Bivariate [GI23, GTK^{+}22, KL21, KFÖ23, MPW23]. Black [SO21b, SO21c, SUN24]. BLF [WWT24]. BLF-based [WWT24]. blind [DHZ^{+}20]. Bloch [BZS21, YTLV22]. Block [ADF^{+}21, TYL^{+}20, BK21, CH20a, CLZY21, CS24c, DM21, DM24, DS21b, HK21, LGQ20, LMX20, LZSC22, MH20, Pap21, RS22a, TZLL24, XD23b, Zen21, ZJX21]. block-by-block [LZSC22]. block-Toeplitz [MH20]. Blocking [SC20c, WOB21]. blocks [BDH^{+}20]. blood [JSA21, Kri22, LGK^{+}22, PM20]. Blow [TZW22, Kan20, LLZ23, MSE24]. Blow-up [TZW22, Kan20, MSE24]. blow-ups [LLZ23]. Board [Ano20w, Ano20x, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40]. Board [Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24y, Ano24z, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano24-41, Ano24-42, Ano24-43, Ano24-44, Ano24-45, Ano24-46, Ano24-47, Ano24-48]. body [BHGPC21, PPS24, RRR19, RRR20, XJ20]. Bogoyavlenskii [Alj21]. Bohemian [CCGV^{+}22]. Boltzmann [BBD^{+}22, Ily23, JNY^{+}21, LZY20, LYN^{+}21, NKO^{+}23, SMW^{+}23, UZ20, WWQ22, YOF21, ZCS21]. Bona [HA20, SL20c]. bond [DS24c, HREB21, JSZ20]. bonding [PS21a]. bonds [LOS21]. books [LLW20]. boolean [XBC22a, DLL^{+}23, DZX23, FCeFZ21, GR21, JLD^{+}24, LLLL23, TKY23, Ued20, WYLL22, WLFX22, WZPL24, YLLL20, YSLK23, ZeFWW20, ZLLL21, ZeFWD23, ZeF21]. Boolean-type [FCeFZ21]. boosting [Ten22]. boosting-based [Ten22]. bootstrap [DVC23]. border [GS22]. border-collision [GS22]. Borel [DHR22]. Boris [SRN^{+}23]. borne [DCR^{+}20, EKKE20, RHNT21]. Bose [TCZ^{+}20, XHWM20, ZGHM24]. both [Fre22, HTF20, ZLZ20]. Bound [Xu21, BF22, CDLR24, CHY22b, CHK21, FX20, JOR^{+}21a, KRE20, KFR20, SRZK22, ZZL20b]. boundaries [KZS20, SNZC20]. Boundary [CMMN20, HXZ21, LJ23, XWM23, ZZL22, ACC^{+}21, AJIR21, AC20, ACGG20, Azi21, BH22, BBHM21, BM20, BPRR22, CM20a, CRT22, CM22, CGLR23, DTK24, DMSS20, EM21, EBJ20, Eva20, Far22, GP23, GC24, GK20, GY22b, HWD20, HCTS23, HZZ^{+}21, HBCA21, HD22b, IZ20, JX24, JCH^{+}20, Kar22, KM20, KK20, KR20, KSR21, LC21, LWL23b, MNSR22, MV21b, MV23, NE21, NWM23, PBKS23, QYLW24, RS22a, RNVC21, RPB23, RG22, SS24c, SY21a, SLHX21, SV23b, SHYS23, SB24, TC22, TZJ20, TZW22, Vyn23, XbYW20, XLZL23, YR20a, ZZX24, ZL20b, Zha21b, ZH22a, ZJX21, ZZW21b, ZJX22a, ZZZ^{+}23].

Bounded [AKDR23, KZ22, Çay20, KZS20, LK21a, LTW22, LQSZ21, LCS^{+}22b, LXK23, MZLS22, PLW^{+}20, QL20, RS22b, SNZC20, Tre23, WHF20, WWWK24, ZWXQ20, Zha24a]. Bounded-Degree [KZ22]. Boundedness [HX22, JSA21, Jor21b, MAAEF21]. bounding [CGZ^{+}22, Pop20, ZYC^{+}24, ZSM20]. Bounds [GGW21b, HPT22, Hu20, MMS23, SS21a, CCC20a, CX22, Cha23, DESM^{+}20, DFX24, FCX21, GA21, LWCD24, OPS20, PLH22, SRGS20, TT20, YJL20, ZRSK21, ZMS23b]. Boussinesq [DMP21, KM20, LQSZ21, LG22, ySyS22, Zha22b]. box [KP24, RY24, LKX20, YL24a]. box-counting [RY24]. BP [HD23]. bracketing [Kod21]. brain [AMM21, GMT20]. branch [ÅCL24, CRS21, LSC^{+}21]. Branch-and-Prune [LSC^{+}21]. branching [CdTPP22, Zha24b]. Bratu [TBRB24]. Bratu-type [TBRB24]. breakage [LZ20a]. Breakdown [Jan22]. breaking [Alj21, CTWG24, Tak21]. breakthrough [JRP22]. breakwater [MMK^{+}23]. breast [HGKD22]. Breather [ÖY21, LM20b, YTZ20]. Breather-type [ÖY21]. breathers [YTZ20]. Bregman [CDM21]. Bresse [AMF20]. bridge [DP21, MJW21, Rom21]. bridges [MGK23]. brief [WLWL22]. bright [TCZ^{+}20]. Brinkman [Pap21, PR20, ZF21]. Brownian [DdlCS24, DDLW21, FJV21, HANP23, LXK23, MYY21, MJW21, MJ23, NKA20]. BSDEs [HLM23]. BSLM [KK20]. bubble [GLW22, Guo23, WMT22]. bubble-sort [WMT22]. buckling [MCB^{+}21, QL20]. Budget [AA20, NTNT24]. Budget-constrained [AA20]. buffer [Chy22d, Chy22c]. build [ZL22d]. Building [TC20a, ZYY20]. bulk [JLFL20, dAGSFM22]. bump [HZW21].

Bumpless [LZ22b, LJZL22]. bundle [TLJO20]. Burger [KIB20]. Burgers [CU23, CZL20, PBM20, PKJB22, RVV24, SG21, ZSFS22, HA20, WP21b, ZF23a]. Burgers’-type [CU23]. burglary [CGMFSHR21]. Burnability [TT23]. burning [LQL21, TT20]. bursting [WGMJ23]. butterfly [LZC21]. BVPs [Cav22, MSB22, WLW20, XTNF22, YSHV^{+}23]. bystander [JGC^{+}21].

C [KHA22]. CABARET [SCMK23]. cacti [LH24a, SLH20]. cages [CDD23]. Cahn [AKMR21, ALL23, CWZZ22, CLZY21, CHY22a, DG21, DR22, JLFL20, Lee20, LYKK21, LL22a, LW22d, PW22a, TYCK23, UK23, ZXF20]. Calculating [LM22, ZYW21a, BS22b]. calculation [FB20, Hua24, LJAM24, LJN^{+}22, Xia22]. calculations [Zha23a]. calculus [AM20a, AKPK21, FF21a, Wu20b]. call [Wan24b]. cam [YHL^{+}20]. can [CWY^{+}24b, NCY^{+}20, RHR^{+}21, SP23, ZHWS24]. cancellation [JJL23]. cancer [BJ21, HGKD22, EA21]. cancer-immune [BJ21]. canine [CXLM21]. canonical [MF23]. cantilever [KET22]. capacity [CZW24]. capillary [Tre23]. capital [Xia22]. capsules [CdTPP22]. capturing [CU23, KRRS22, ZHQ^{+}23]. Caputo [BGRM21, FF21a, LWNZ22, TZ21, YWS24]. carcinoma [DTK24]. cardiac [BEe21, BAL^{+}24, CK20]. cardinal [Rom21]. careers [RYR^{+}23]. Carlo [NKA20, TS23, VS24]. Cartesian [CHK21, CSK23, TX21]. Cartoon [WZN22]. cascaded [LJ23]. cascading [WWZ^{+}23]. Case [FMHDB22, CWF^{+}21, FB20, GKK^{+}24, HLKC22, HCL^{+}23, HKCK24, HL23b, LHC^{+}22, LLY^{+}22b, LL22b, LX24a, LJHL21, TWB23, VBP22, WCZ^{+}22b, ZLZ24a]. cases [CRTVG^{+}23]. Casson [KNEM23, SN21]. catacondensed [BTP20]. Catalan [ZZ23]. catastrophes [Sch22, dOSR21]. categorical [YCZ^{+}20]. Catlin [LHZ21d]. Cattaneo [CSG^{+}23, CSG^{+}22, vdMvRS21]. Cattaneo-Christov [CSG^{+}23]. Cauchy [DHM^{+}20, LH22, WFHG20, XbYW20]. Caudrey [ZLA^{+}21]. Causality [LQLF22, GN21]. caused [Guo22, ZyZW21]. causing [EKKE20]. cavity [AAHC21, Pap21, RWGS22]. Cayley [HZF21, JT22, LWW23, LTWY24, LHL24, WF23a, XFK22, YWFL23, ZXC^{+}20b]. CCA [XYKL20]. CD [YI22]. CD4 [SWCT21]. CDNs [WHCW20]. Cell [ARÁSY21, CFR22, LGK^{+}22, LWJ21, MS22a, RMO22, SG20b, THW^{+}21, WSY21]. Cell-average [ARÁSY21]. cell-based [LGK^{+}22]. cell-centered [WSY21]. cell-to-cell [LWJ21]. cells [SWCT21, YI22]. Cellular [TZL24, AZD^{+}24, LCZ^{+}20]. censored [LWL22]. center [BZ23a, HQZC24, HZCY23, LM21, LMS23]. centered [WSY21]. centerline [ZZL^{+}24]. centers [dSAOG21]. central [DL20a, KRRS22, QZZ^{+}23, QQ23, XJ20, YTC23a, YTC23b]. Centrality [KKR21, SG20a, BS23, CS21a, CTV21, CRTV22, LRCTR22]. cerebral [KCT^{+}21]. cerebrospinal [SBT22]. certain [Baz20, CdFP24, CW23, FCX21, LSVA20, SMW22, YL24a]. certificateless [JSCB20]. Certificates [BSdCCGCV21]. CES [CH20b]. CEV [SMR22]. CF [LtTBPA20]. CFD [MCT^{+}22]. CG [LM20d]. chain [DD24b, DN21, FJJJ21, KNN20, MKZ22, SSP24, She20, VS24, YLX^{+}21, ZYLH21, Zha22c]. chained [ZZL20b]. chains [BHR20, LZ24a, LY23a, LLZ22c, MV21a, MY24, RST21, SD23, SY23]. Change [SCC24, AMM21, DCCM20, DN21, KR20, NS22, XAZ^{+}21, ZWNG21]. change-based [AMM21]. changes [RCG21]. changing [CCH24]. channel [BH22, CL24b, Has20, HCL^{+}23, KNEM23, KLNN23, LSL^{+}20, MUOF22, MMUS22, ME21b, MKZ22, NHK^{+}22, NHK^{+}23, SBB^{+}22, WHZ^{+}20, XLG^{+}21, ZCZW23]. Channels [ZCL^{+}23, CHW^{+}24, CdTPP22, DD21a, ECS21, GY22a, LL23b, NSSA20]. Chaos [GS22, TlS^{+}22, BRFLGH21, DP21, Jor24, KG22, LSCS21, MS21a, NS21, QLL24, ZHMZ21]. Chaotic [BY21, LM20c, YTZ20, Ahm21, BSM23, HKKL22, HB20, KTEM22, LW20b, LKX20, LLH21, RLX22, SKL23, SJTCdlFL21, XWMC20, YZAM20, YZ21, ZH22b]. Chaplygin [SSJ22]. characteristic [DWW24, LXL22, LY23b, MPW23, TNL23]. characteristics [AR23, DMNV20, Xia21]. Characterization [CdTPP22, QLZL20, UMT22, ZLJ^{+}21]. Characterizations [MGS20]. Characterizing [YWX20b]. charged [LW23a]. charged-particle [LW23a]. Charlier [MP20]. chattering [YSL21]. Chebyshev [AAS23, CdFP24, Lee20, OT20, OPS20, PP23, WZCF21]. check [RKRRL21]. Cheeger [FKK24]. chemical [BBH22, CSC20, CMR20, CGR21, DS24c, MD24, TY21, ZH20a, ZXF20, dAGSFM22]. chemistry [GHND22]. chemoimmunotherapy [RMCG20]. chemostat [ZY21a]. chemotaxis [KAS21]. children [RS24]. Chimera [HGJ21]. China [CWF^{+}21, WWZ20b, YWJ23]. Chinese [YXZ^{+}21]. Chlamydia [AMA20]. choice [CB21, LDZ20, LWZ22d, SR24, TO20, ZSZL20, dFdOFdSBC20]. cholera [LJH^{+}20]. Cholesky [CZMX21]. Choosability [CFR^{+}20, HZ20a, HZ23b, LW22b, YCW23]. choose [JP22]. chords [GW20a]. choroidal [CCG^{+}22]. Christov [CSG^{+}23, CSG^{+}22]. Chromatic [PK20, BGHX22, DYZC20, ZZB24]. chronic [AMA20]. chub [JC22b]. CI [XFK22]. CI-property [XFK22]. Ciarlet [WS21]. CIR [AMOS23]. circuit [EVB^{+}21]. circuits [GLD21, xHhPjW^{+}21, YZAM20, YXZ^{+}24, ZC23b]. Circulant [HV24, PNJ^{+}21, PP23, Son24]. circulants [PY21]. circular [GPHHC23, WYLW20, WPE20]. circulation [JSA21, TWZ^{+}20]. circumstances [YTX^{+}24]. circumventing [LL23d]. CIR’s [DDM24]. CIST [QH21]. City [CGMFSHR21]. civilian [Cro22]. CKOEs [KK22]. claim [ON21]. claims [CKW22, LBJ22]. Clairaut [OiEF22]. clamped [WSZ22]. Clark [JZ22]. Class [AHM23, GLX22, KSS23, AFGG20, ACGG20, BM20, BLS21, CLC24, CSQT21, DDGN23b, DDC22, GS21, GN20b, GLWX24, GT22b, HA20, HCX^{+}24, HM20b, HVZZ23, IR22, JDLZ21, JTC^{+}24, LLZ^{+}20a, LMX20, LLY^{+}21, LGX22, LLY22a, LG22, MXF^{+}24, MD22, MKP20, MZFX20, NT23, Ndo23, OiEF22, PX23, SCLD21, SX24, SM23, TLPT21, TZW22, WLW20, WHWW21, wXhGZ20, wX20b, YM21, YGZ^{+}22, YLC20a, YXMW22, YHD24, YF21a, YLL20, Yoo21, YZY22b, ZM20, ZCJL21, ZG24, ZH20b, ZJW^{+}22, ZZ20b, ZPLL20, ZC21c, ZZD22, ZWY^{+}22, ZDMW23, ZZZ^{+}23]. classes [AC22, HJY24, LLL^{+}23a, SPX20, TT24, TPSP24, XTNF22, YL24a]. classical [SV23a, ZLA^{+}21, Zha23b]. classical-quantum [SV23a]. classics [SM21a]. Classification [HSM21, RMO22, TCZ22, dAGSFM22, CZ21b, HGKD22, Ndo23, SH21]. classifiers [PSSU21]. claw [SPX24, WJY21, XT23]. claw-free [SPX24, WJY21, XT23]. clearance [Zha22c]. clearing [BSCL20]. Clifford [De 22, Gu20, MSDR21, RSVL21, ZQ24]. Clifford-valued [RSVL21, ZQ24]. Climate [DCCM20, ZWNG21]. close [ARÁSY21, AY22]. Closed [Fre22, Hon24, MKP20, SMR22, FKVZ21, EVB^{+}21, MRR20, YLM24]. Closed-form [Hon24, MKP20, SMR22]. closed-loop [EVB^{+}21]. closure [ISM^{+}23]. Cloth [CFA^{+}23]. clothoids [Fre22]. cloud [JSCB20, MMM24, YWWY23]. Cluster [WZLH22, PM21]. Clustering [GRMQ22, CZMX21, DVC23, YCZ^{+}20]. CN [QL21, ZZY24]. CNTRC [Bab22]. CNTs [EM20]. Co [BY23b, PWZ22, SWX24a, KP23, KG21, TGL^{+}24, YTX^{+}24, ZHWS24, ZWJ^{+}20, SL20d]. Co-design [PWZ22]. Co-evolution [BY23b, SWX24a, YTX^{+}24, ZHWS24, ZWJ^{+}20, SL20d]. co-evolutionary [TGL^{+}24]. co-kriging [KG21]. coalescence [KOP21]. coalition [YHSZ20]. coarsening [FGM24, YXAY21]. coating [BM22]. code [GDM^{+}20, IP23, LZLM20, LTL^{+}20, ZLU21]. codebooks [WNG20]. codes [BDMB20, GGW21b, GYM21, MGD21, NXG23, SNC22]. Coefficient [YCEW21, Cha22, EA24, Jor21b, JJSY22, LP22, RRS21, ST24, SD22a, TZ24, WW22a, ZT22a, ZJZD20, ZC22]. coefficients [ÅCL24, ASR22, BS22b, BD20b, CLWX23, CKV21, DS22a, FHM22, FRS22, GC24, LX22, LNC22, LCL22c, PSXX21, ST24, YLS23, ZGZ^{+}24, ZLW21a]. Coevolution [DZW21, CGW^{+}20]. coexistence [PJ21, PNOdV20, YSYS24]. coffee [FMHDB22]. cographs [BHP21]. Cohen [DWZ20, YWS^{+}20, YDSW22, ZYZ22c]. Coherent [ZZY^{+}21]. coin [KB21]. coinvariant [AA21a]. cold [WXWS20]. Cole [CZL20]. Collapse [CLC20]. Collatz [DH24a]. Collective [KXW24, BBEM24, CYW21, DS21c, JGC^{+}21, MRW21, RMO22, WDZL24]. collective-risk [JGC^{+}21]. collinear [WPJ22, XJ20]. collision [DZL^{+}24, GS22, WC22b]. collisional [PDK23]. Collocation [SXC21, ZHZZ24, AHB20, CD20a, CD20b, CDDP21, DK22b, HL24b, Lee20, LWL21, LYN22, Mar24, MSN21, MSB22, NE21, QXG21, SG21, SYX22, SFC23, TZJ20, WZCF21, Wan22a, XFW^{+}21, XTNF22, YWZT20, YWZ23, YY22, Zha22b, ZH20b, ZLJ20]. colonic [CFR22]. Color [WGM^{+}21, HD22a, JG21, WHM^{+}21a, WHM^{+}21b, WZN22]. colorable [BDK23, PWL23, ZCW24b]. colored [HLSL22, JWWL20, JMY22, LWX24]. Coloring [CFL^{+}20, WHKT23, Xu24b, BZ23b, CLCD23, DFP^{+}24, HLM^{+}21, LLL23b, MSS24, PK20, QZ21, SS24b, SLL24, WSC^{+}24]. colorings [BGHX22, GH20a, GZZ21, KT21b, MMOS24, YW22a]. colors [WLL21]. colouring [DDPS22]. colourings [HH23b, KP22]. column [WCZ22a]. columnwise [BT21]. combination [AMA20, GF23, PMG20, WH21, XFY20]. combinations [CET22]. Combinatorial [LY23b, LPL^{+}22, ZZ23]. combinatorics [ZGWZ22]. Combined [BGS22, CSG^{+}22, CSG^{+}23, SLK^{+}23, AM20b, AAHC21, LHH24, WPJ22]. Combining [YCZ^{+}20, LX23a, WWZW23]. combusting [SKL23]. combustion [SX23b]. coming [JOR^{+}21a]. Command [MPX^{+}20, HZX22, WXLZ21, WZXH22, XLZ23, XLLH22, YCM^{+}22]. Command-filter-based [MPX^{+}20, HZX22]. commensurate [ZLX20, ZLZ24a]. Comments [RO21, Woo20]. commitment [FH24]. commodity [LWZ22a]. Common [Ued20, CHY22b, HWWW23, OK21, Wet21a, ZQZ21]. Communication [HS23b, CZLH20, DZZ^{+}20, Fan21, FLH^{+}21, HTF20, KKPS21, MTFD24, NSSA20, QZH21, Wan22b, WPZY23, YWG24, YP22, ZZL^{+}20a]. Communication-hiding [HS23b]. communications [SCLD21, ZZCW21, ZZLL21]. communities [AGCGSC21, CDLR24, LJH^{+}20, NZ20]. Community [GRMQ22, LZZ21, LJW^{+}21, LLG^{+}24, WBY^{+}21, XKW^{+}23]. commutative [LNC22, ZJWV24]. commuting [WLJ22, YWX^{+}20a]. comonotonic [ARMM23]. Compact [DS22a, KFM21b, KFM21a, Alm23, CZL20, DTK24, FHM22, FZ23, HWCD22, HJY24, KM20, LePP24, LC22a, LZ24c, LTB23, PW22a, QWD22, SYX22, WW20, WGSS22, Zha21a, ZSFS22, ZZW21b, ZC22]. comparative [BM24, DJWX21]. Comparing [vdMvRS21]. Comparison [Kri22, CSV20, GGW21a, LZL20, SO21b, SO21c, SNS^{+}20, TM20, ZLDY22]. Comparisons [SLHX21, BKK24]. compartmental [ACBLC24, CS20]. compatibility [Azi21]. compatible [ABD23]. Compelling [BGHX22]. compendium [KP24]. Compensated [LHC21]. Compensation [HD20, WSXL22, GLKrH20, HRA^{+}20, LLZB24, WHW^{+}21, WCY^{+}24]. compensator [CHHC24, GZ20b]. compensator-based [CHHC24]. competence [WLS22]. Competing [LXP^{+}22, WBV24, GPRV21, MS21a, NZLW22, NLP^{+}22]. Competition [PLQ^{+}22, Alf24, BM22, DE21, HDQP23, KSZ23, MKZ22, PJ21, WBV22, WX22]. competition-diffusion [WBV22]. competitive [CL24a, HCM21, HPC24, LDXZ23, LL20b, MGM21, RSXH20, SGX21, WZLH22, XYLF21, YXW23, Zha21d, ZZ22a]. complementarity [AA24, GZLZ24, ZP20, Zha21b, ZZLV23, ZZVL23, ZZL20b, ZV20, ZV21, ZZS^{+}22]. complements [LLCL23, YML21]. Complete [CL24a, GST22, MD24, CLWZ21, CY21b, DZX23, JWMW23, JJC24, LW23b, LQ24, RD23, WM22, WL22, WB20, XYZZ22, XW23, ZLL20, ZHW22, ZHFL23b]. Completely [WS20]. completion [CYG20, PLH^{+}24, Zha21d]. Complex [CLZ24, HWC23, HHY20, LXCZ21, MYW21, QDL^{+}21, YXW23, YWJ23, AAJ23, ADV23, Ask21, BSRV23, BG20a, CQ24, CFWL24, CLH24b, CLLA22, DTWJ21, FW24, GGL^{+}23, GXZZ20, GX23, HZ23a, HHZZ20, HSZ20, HB20, HXWC24, JWC20, Kud20a, LXP^{+}22, LWG20, LK21b, LZH21, MMHX20, NPJ^{+}20, NLP^{+}22, PPZW20, RD24, RJHC24, SD22a, THW^{+}21, VHB23, WMWS21, WLJ^{+}21, WZY21, WCWZ23, WZLC21, WCZ^{+}22b, XYH^{+}24, YLHZ22, YYF24, YWX^{+}20a, YWJ20a, ZP21, ZYZ22a, ZZPF23, ZGZ^{+}23, ZWFW22, ZKLL23, ZXB20]. Complex-Valued [HHY20, AAJ23, GX23, HZ23a, HHZZ20, HSZ20, HB20, LZH21, RJHC24, WZLC21, XYH^{+}24, YYF24]. complexes [SG20a, ZLP^{+}22]. complexities [GV23, GV24]. complexity [AMM21, CGB20, HLP24, RSZA23, SVW23, ZZY20]. Component [LMS23, CLL24, DGP^{+}23, LMX22, LZ23, UZ20, Xu24a, XYlW^{+}24, YZJ^{+}20, YZM24]. component-wise [LZ23]. compose [dSAOG21]. Composite [KBL^{+}24, XCLW21, AS20, DD20a, GAA^{+}24, LAS23, MKKG22, MK21a, MSK21, SSC22, SMW22, TITN20]. composites [BSRV23]. compositions [BCS24, CET22]. Compound [STH24, Xia22, CX22, YCL24]. comprehensive [YY21c]. compressed [LPL^{+}22, MJ23]. compressible [BT22, CGLR23, HL23b, JM23b, KRRS22, LZLM20, LTL^{+}20, MCT^{+}22, TKF23, XLG^{+}21, YGM^{+}22, YS23]. compressing [CDBR^{+}22]. compression [ADS23]. comprising [SSJ22]. compromise [LXF21]. Compt [LPP21a]. Comput [AACGTB24, PPS24, WDG21]. Computation [ABNS24b, AV21a, CSG^{+}23, CF23b, Elg22, FWM^{+}22, GNA22a, KWM20b, KA23, KFM21b, MV23, NV21, NHK^{+}23, Oru22, RRR20, RO21, SSP24, SPX20, SPX24, SZ22, WCY20a, WZ20a, WHM^{+}21a, XH23, ZBA^{+}23a, BFGGM24, BKS21, CJ20, DTL22, FLR23, FdLPC^{+}22, Grz22, HZT^{+}21, KFM21a, LOG20, LL20a, MM24, Miy23, OG20, PW22b, SV21b, SKJM21, SĆL^{+}21, SY22b, THZ23, XL24a, ZMS23b]. Computational [CCC20a, CEPZ22, LKK^{+}23, MM22a, MM24, MMMBRS20, MMUS22, VS24, VSM23, ASR22, ABD23, ASAK21, BJK^{+}22, DBVCP23, GA20, GNZ20, GMN22, KZ22, LBX^{+}23, MPRS24, vdZGO21, Yua20]. Computationally [BKK24]. computations [AO21a, AO21b, BSCV24, Ma21b, MVO21]. compute [HMEDVR24, WWY22]. computer [GS22, FGMS21]. computer-assisted [GS22]. Computing [HHR21, JWMW23, SRGS20, TCG20, AAS23, BCVV22, CSQT21, DK22a, KKH23, LLWY23, SPM22, SMM^{+}24, Wri24, ZMS23b]. concentration [JLFL20, WWZ20b, YWT20]. concentration-dependent [JLFL20]. concentrations [GP22]. concept [AAHC21, TVY21]. concordance [FZ20a]. condensate [TCZ^{+}20, ZGHM24]. condensates [XHWM20]. condensation [KSPC22]. Condition [AKDR23, AR21a, BM20, BPRR22, CWZZ22, DM23, DLQY20, DML21, GW20a, Har22, KR20, LQLF22, LH20b, LL23d, MV23, SSA^{+}23, Wan21b, Wu20a, ZH22a, ZCW24b]. Conditional [dJCGVM24, Mef21, SCT^{+}20, ZAD21, DHS23, GLL^{+}22, JLZ^{+}22, SMR22]. conditioning [ACS24, BGS22]. Conditions [Zha23a, AzR22, ACC^{+}21, Azi21, BH22, CD23, CCH24, DMSS20, ESA20, GJ21, GGY22a, GGBEZRI23, HKL22b, HBWQ20, HD22b, INOP20, JCH^{+}20, KM20, KK20, KEiS23, KSR21, LPP21a, LPP21b, LPK22, LL21, LJY22, MV21b, SM20, Sin20b, TZJ20, TZW22, VBMB21, WWK24, XZeF20, XHL23, XLZL23, YM23b, ZM21b, ZZW21a, ZL20b, dOS20]. conducting [ABD22]. conduction [BSRV23, BD20b, Lee23, MM22b, XLZL23, vdMvRS21]. conductivity [KR20]. conductor [LSV21]. cone [MJK22, TZ22, ZZS^{+}22]. cones [GS22]. Configuration [XZ23b, BHGPC21, KL22]. configurations [XJ20]. configured [CSG^{+}22, CSG^{+}23]. confined [GK20]. conflict [DDH^{+}24, JL20, RSR21]. conflict-free [DDH^{+}24, JL20]. conflicts [SGJ^{+}21]. Conformable [YFW21]. conformal [NV21]. Conforming [WSZ22, TCPF24, YJ20]. conformity [QDW22, YLW20]. conformity-driven [YLW20]. congestion [LCD^{+}20, ZSZL20]. Congruence [Gas23]. conic [NGGD23, WHLL20]. conic-type [WHLL20]. conics [BMDE21]. conjecture [CWZ22, DYZC20, DD21c, DH24a, GWW23, LW23b, LHZ21d]. Conjugate [Hua24, KRS20, CET22, GLP22, NL20, SSI23, THW^{+}21, YM21, You20, ZC21a, ZZW20]. Conjugate-mixed [KRS20]. Connected [HK21, AHL21, CLL21, CCHZ23, DW20, HAK20, Kun24, LW21a, LAH22, LLZC21, LX24b, SM20, SLL24, SZWL21]. connectedness [SM21a]. Connecting [NPE22, AMZ21]. connection [DDH^{+}24, DDS24, FSS21, JL20, LW22e, MZ23b, MFK22, NMS23, SRKA23]. Connective [YL20b, YLL22]. connectivity [BZ23a, Che23, CHL24, DXS^{+}20, DG20b, GKT^{+}23, GMT20, GLW22, Guo23, HM20a, JGW21, Kun24, LSWL20, LZRG21, LW23b, LM21, LMS23, Xu24a, YZM24, Yi23, ZHFL23a, ZHFL23b, ZT24]. connectivity-preserving [Yi23]. conquer [CJ20]. consecutive [LHW^{+}23]. consecutive- [LHW^{+}23]. Consensus [LZX^{+}22b, WLQX20, Wan22b, YFW21, ZZCW21, ZYXL23, BHK^{+}22, BRTR21, CZLH20, CDZW21, CMYF21, CZ22b, CLC^{+}22, CTF24, DWZZ22, FJLG22, GZCS23, GWZZ23, GHS^{+}21, GH21, GZWX23, GW23b, HQS^{+}24, HLW^{+}20, HJW^{+}23, JLP23, LSWL20, LJL21, LHX^{+}22, LDCZ24, LNC24, LXK23, MPC^{+}24, PHS^{+}20, PGZ^{+}24, QYLW24, SCLD21, TLHH21, WS20, WWW^{+}21, WDTF22, WC23, WZW23, WYLQ24, WWW24b, XWL^{+}22, XFY20, YLC20a, YWWY23, YXW23, Yi23, YL24b, ZZLL21, ZCZ22, ZDC22, ZCD23, ZLZ^{+}21, ZPY22, ZLP21, ZD24, ZPW22, ZDD^{+}22]. Consensus-based [LZX^{+}22b, WWW^{+}21, ZLP21]. conservation [AB23, CS24a, FCH21, FZ23, GN21, HCTS23, HC21, KK22, MS22a, MCD21, YTC23a, YF21a, ZZL22]. conservatism [ZCPZ20]. Conservative [CG20a, DGP^{+}23, LPKC22, LXL22, LZ24c, MC20, RO21, Sun22, TYCK23, WDG21, WL23, YTC23b, YLG21, YCEW21, Zha21a]. conserved [QLG^{+}22]. conserving [AR24, KTEM22, YZLS20]. conserving/stable [AR24]. Considering [TZL24, Ach21, BYCX24, BJ21, DZL^{+}24, GLKrH20, GW20b, JM23a, LXH24, MGFF21, MMDD21, MKZ22, RCG21, SSW22, WPH20, YWX^{+}20a, YI22]. consistency [BCCI20, LPS23, LZL20, PGYW23, Zha22a]. consistent [BKU20, BY23a, WWY^{+}20]. constant [FKK24, HKU21, Lee20, NXG23, PM21, SO21a, ZF24]. constants [KL21]. Constrained [BCCI20, LLL21, AA20, BHK^{+}22, CWZ20, CG20b, FLZZ24, GLY24, GWH22, HL23a, HZX22, HYT21, JLOR22, LWZ20, LLZ21, LD23a, LWH22, LH24b, LWL24, MD22, NGGD23, SZCP21, TZJ20, Wan21b, WC22a, WNZ21, wXhGZ20, wX20b, ZMD^{+}21, ZSYX23]. constraining [SSW22]. constraint [CCE^{+}22, DMD21, HCBY21, LePP24, LAS23, LR24, LD21b, NTNT24, SX24, WW23, WX21b, WSSX22, XYY23, XJY24, YL24b, ZWW23, ZZL^{+}24, ZZX^{+}22]. constraint-coupled [LAS23]. Constraints [MXY^{+}20, AKY21, AFG22a, AFG22b, BBB24, CGTU20, CGZ^{+}22, CLSN24, DSZ^{+}23, FSLZ23, GSWL23, HRLH21, JPZ^{+}21, KLP22, LD20a, LLL21, LXWH22, LZZL21, LNZ^{+}22, LFM^{+}22, LLCS23, LYN22, LWY^{+}24, MXF^{+}24, MCB^{+}21, PX23, PGZ^{+}24, QXX21, SYL21, SY21b, SLPH22, WZZX21, WSS22, WGZZ22, WPZY23, WPX^{+}20, WWW24b, XWL^{+}22, YWG24, YMYR21, ZL20d, ZYPL20]. construct [YHL^{+}20]. constructed [RY24]. constructing [LKX20, PD23, YAX^{+}21]. Construction [BIJ22, Çay20, HBWQ20, KPS22, RAB24, ABD23, ED20, FKVZ22, LPL^{+}22, NXG23, QZZ^{+}23, SUN24, XBC22a, YLPC22, ZHMZ21]. Constructions [CDD23, NXG23]. constructive [GC20]. Consumers [YY20b]. consumption [HH20, JP20, LWZ22d]. contact [AR21a, CH22b, CMMN20, DKKJ21, DSY22, GWH22, HWC23, HHW23, KRRS22, LSV21, LXH24, LJHL21, TAU^{+}22, TCZ^{+}20]. contact-discontinuity [KRRS22]. Contagion [KK23b, YCL24, LFY^{+}24, TNT23]. contagions [LLG^{+}24]. containing [LZ24a, MY24, NNKH21, SO21a]. Containment [LSL^{+}21, MYY21, XBLW23, YDXH22, YMYR21, CDZW21, GL24, MCD22, MTFD24, XWWC21, YXMW22, ZPLL20, ZWY^{+}22]. Contaminant [AMBG^{+}20]. contaminated [DM23, ON21]. content [YCZ^{+}20]. contest [ZWXQ20]. context [DKKJ21, JM23a, MZW^{+}21, WCY^{+}24]. contingent [LBJ22]. continuation [CD22, LD20a]. Continuity [LWX24, HBWQ20, Mar22]. Continuity-sets [LWX24]. Continuous [DR22, GLX22, ZZ22b, CNZ^{+}21, CCF24, DO21, GLWX24, Guo22, HKKL22, JDZH22, KNN20, LZZ24a, LSM23, LZHZ24, LZW^{+}20b, LYXZ21, LNC22, LCS^{+}22b, LCL22c, LXZ23, LS22a, MZLS22, PPS20, RST21, SGL21, SKJM21, WS20, WCG^{+}20, WWY^{+}20, Wan22b, WYLQ24, WZZ24, XZH^{+}20, XYZY20, XTS^{+}22, YL20a, Zha22c, ZLW21a, ZL22e]. continuous-Time [CNZ^{+}21, MZLS22, Guo22, LZZ24a, LSM23, LZHZ24, LYXZ21, LXZ23, PPS20, SKJM21, WS20, Wan22b, WZZ24, XTS^{+}22, YL20a, Zha22c, ZL22e]. continuous/periodic [WCG^{+}20]. continuum [KOP21, RCG21]. Contour [SSR23]. contourlet [WWZ^{+}23]. contracted [FMT^{+}21]. contractive [JSA21]. contracts [BO21]. contrast [FHM22]. contributions [LOG20]. Control [DDGBGS^{+}24, GLX22, GLP20, MWS23, MXY^{+}20, WYLW20, WSXL22, XZ23b, XWDK23, YLX^{+}21, YXYZ23, ZCL^{+}23, AS23a, AMA20, AS21d, AFG22a, AFG22b, BZZ21, BLS21, CZWS20, CSS^{+}22, CL22, CZ22b, CJ22, CWZL23, CWH24, CNZ^{+}21, CZS21, CYHR22, CYL^{+}22, CZK^{+}22, CLWX23, CCL^{+}24, CHHC24, CSS^{+}24, CFWL24, CZ24b, CZZP24, CLH24b, CMMN20, CHST20, CZWX21, CZGX22, CXY^{+}23, DLL^{+}23, DCR^{+}20, DG23, DS24a, DWS24, DZZ^{+}21, DXZ^{+}23, DLS^{+}23, DSZ^{+}23, DYW^{+}20, FKS23, FJLG22, FT24, FSS20, FLC^{+}22, FFL23a, FZS21, FLH^{+}21, FLLW23, GZ20a, GYC^{+}22, GSWL23, GWZZ23, GZL^{+}23, GSS24, GLWL22, GC24, GS21, GS24, GBCP21, GLT^{+}21, GLTY22, GLWX24, GSRL20, GZ20b, GSL^{+}22, GHP24, GLD21, GW23b, GX23, GL24, GZLY24, HYÇK22, HTF20, HWD20, HZW21, HZL21, HKL22a, HL24a, HRLH21, HQK21, HLL^{+}21, HTZ^{+}23, HXZ21, HCBY21, HLLS22, HLL22, HL23a, HCL^{+}23, HCX^{+}24, HKCK24, HZX22, HB20, HYT21, HSL^{+}24, HB24, HMZZ20]. control [HLLY20, IFHA21, JZS^{+}24, JDLZ21, JDZH22, JCCC20, JPZ^{+}21, JMC23, JTC^{+}24, JJSY22, KMLW22, KGCC20, KTZ24, KKS^{+}22a, KZWS22, KAT22, KBL^{+}24, KIJK20, LaGBT22, LLZ22a, LLZ20b, LCLS20, LDX20, LGL20, LCD^{+}20, LLL21, LWZ21, LJL21, LCXM21, LJZL22, LTW22, LZ22b, LGX22, LZC22b, LZWC22, LLY^{+}22b, LZDX22, LLY22a, LXWH22, LTL^{+}22, LDXZ23, LHZ^{+}23, LWL^{+}23a, LZZG23, LWL23b, LYX23, LLLL23, LR24, LM24, LZZL21, LSM23, LZHZ24, LCY^{+}21, LM20c, LZW^{+}20b, LLLS20, LWG20, LK21b, LHFW21, LTL^{+}21, LYXZ21, LZLZ21, LD21b, LNZ^{+}22, LWH22, LZX^{+}22b, LCL22b, LFM^{+}22, LCL22a, LZZ^{+}23, LLWK23, LL23d, LXZ23, LXK23, LH24b, LZ24d, LLQW24, LLS21b, LLWW22, LJWD21, LJW^{+}22, LR23, LY20b, LLT^{+}20, LWY^{+}24, MPX^{+}20, MCD22, MZY^{+}22, MTFD24, MXF^{+}24, MTRA22, MMR23, MS20, MNSR22, MZFX20, MPC^{+}24, MC22, MYY21, MKK23, NK20, NK21b, NK21a, NK22, NH22, NSSA20, OPL22, OG20, PHL23, PHS^{+}20, PLRD21, PWZ22, PX23, QQWJ22]. control [QQY^{+}23, RHR^{+}21, RST21, RLY20, RFFZ21, RDL23, RMCG20, SCLD21, SKLS22, SX23a, SJ23a, SYL21, SMH^{+}20, SWS^{+}22a, SX24, SWX^{+}21, SWX22, SPZS20, SC20b, SZP21, SWSW21, SZZL22, SZSZ22, SV23b, SYTC24, SWZS20, SCG^{+}20, SZWL21, SCHS22, SDW^{+}22, SRZ22, SSW22, SHYS23, TlS^{+}22, TWR20, TZD22, TWB23, TPTA24, VSM^{+}21, VSM22, WLQX20, WCZY20, WXWS20, WZXY20, WCZ20, WGW^{+}21, WXLZ21, WXW^{+}21, WFYL21, WCM^{+}21, WLZF21, WZZX21, WWY21, WWW^{+}21, WXP21, WSS22, WWW^{+}22, WA22, WZC^{+}22, WGZZ22, WYLL22, WLD^{+}22, WCW22, WLPZ22, WXF22, WLS22, WLFX22, WM23, WZW23, WS23b, WLL^{+}24, WJDL24, WM24a, WZPL24, WQT^{+}24, WJL^{+}24, WFKL24, WC22b, WWT24, WPX^{+}20, WZXH22, WSSX22, WYS23, WWW24b, XD21, XH22a, XBLW23, XLZ23, XD23a, XWLL20, XCW^{+}22, XW22, XYZY20, XXW^{+}20, XZZW21, XCLW21, XLLH22, XWL^{+}22, XZZC22, XZQ^{+}22, XWL23, XWWL23, XZQ23, XYR^{+}23, XLZ^{+}24b, XRZH22, YYG20, YWY^{+}22, YLLL20, YY20a, YLK^{+}21, YCL^{+}21, YLL22]. control [YXMW22, YCM^{+}22, YLHZ22, YWWY23, YM23a, YW22b, YDXH22, YCZ^{+}23, Yoo21, YP21, YYWL23, YM20, YZDC22, YMYR21, YZY22b, YZ24, YSL21, Zen21, ZDGL21, ZWH23, ZY20, ZZQ^{+}20, ZYPL20, ZZL^{+}20a, ZJZD20, ZWWH21, ZZCW21, ZC21a, ZDZS21, ZZLL21, ZLLL21, ZLS21, ZCZ22, ZYSW22, ZQP^{+}22, ZXZC22, ZW22a, ZCS^{+}22, ZD23, ZWW23, ZYXL23, ZCD23, ZXZ24, ZWY^{+}24, ZLC24a, ZCW^{+}24a, ZDZ24, ZPDL20, ZM21c, ZLZ^{+}21, ZJW^{+}22, ZCLZ22, ZGLZ22, ZY23b, ZGZ^{+}23, ZZ20b, ZPLL20, ZWHY20, ZC21b, ZZX^{+}22, ZTW22, ZWZX21, ZeF21, ZZW^{+}21c, ZCJ22, ZWY^{+}22, ZD24, ZPW22, ZDD^{+}22, ZXW20]. control-based [GBCP21]. control-interval-dependent [LHFW21]. control-localized [LaGBT22]. Controllability [AS21d, KSS23, LSWS21, SWF20, JS20, KMD21, Kum23, LLL21, LFA22, VJA^{+}22, WLZD21, ZZX24]. controlled [GC24, VGR^{+}21, Zha23b]. controller [Ahm21, CZXyS24, DMD21, HDP^{+}21, KKPS21, LNS^{+}21, LZS^{+}21, LXW^{+}20, MMHD20, QLC^{+}24, SC20b, SM23, WXWS20, WSDW24, XZeF20, YWJ20a, ZY20, ZeFWW20, ZLZ24b, ZZS^{+}22, ZZZ^{+}23]. controllers [GZX20, GL22, TKY23]. controlling [CWL^{+}21, CLD^{+}24, LDZ21]. convection [AAHC21, BS22c, DS22a, DS22b, DL21b, DMP21, GBF21, KZS20, KSPC22, KRS20, KTEM22, LF21, LGC24b, LR20, LW22g, LZ22e, LtTBPA20, MJK22, Mef21, PGF20, Que22, SKS21, SNZC20, SKL23, TM22b, VN20, WW20, WZ20b, Yu21, ZL21a, ZL22b]. convection-diffusion [LW22g, LZ22e, LtTBPA20, PGF20, Que22, VN20, WZ20b, ZL21a]. convection-diffusion-reaction [DL21b, LF21]. convection-reaction-diffusion [WW20]. convective [CSG^{+}22, CSG^{+}23, MM22a, MM24, SN21]. Convergence [AYY24, AT21a, DTL22, GL23b, KK20, LLT22, LNC22, LW24, LYN22, LMÖ23, LZ22g, MG21, OH24, SMW^{+}23, SNMM21, SGX21, SS23c, WL23, YWZT20, ZHL24, ZL20a, Zha22a, ZJX21, AA21b, AC20, BM24, BPR21, BMY24, CRT22, DHZ^{+}20, EBJ20, GZH22, GZCS23, GLL23, GYC24, GGY23, IÖÜP23, LLS21a, LW22d, LCXJ22, MPM23, NL20, RM23b, RG20, SSI23, SRZK22, Shi21, Sin20b, SYX22, SPM22, THBPMC22, WLW20, XTNF22, YOF21, YCWH23, You20, ZL22b, ZYX21, ZV20, dSTV21]. convergent [CG20a, RC22, SR23, SUN24]. convertible [DDGBGS^{+}24]. Convex [GF23, HS23a, AHL21, Fen22, HPL21, HLP24, HMS20, JOR^{+}21a, LFLH23, LHZ^{+}21b, LAH22, LMZ^{+}21, Mit24, PHL23, TLJO20, TCA23, TWR20, VSM23, WHZ^{+}20, WHPQ23, XZY^{+}23, ZS21a, ZL22c, ZGCZ22]. convexification [WLPZ22]. convexity [BLOP23, LOS21, WYZ24]. conveying [MYC^{+}23]. convolution [BFS20, But21, MSE24]. convolutional [LFLL21, OKT24]. Conway [Xu21]. Cook [CWZZ22]. coolant [Hoj20]. Cooperation [GWZ^{+}21, KTZ24, YTX^{+}24, ZHWS24, BBEM24, BY23a, CYW21, CWY^{+}21, CZW23, CTWG24, CWY^{+}24b, CWY^{+}24a, CCC20b, DWL^{+}24, DZ21, DCCM20, DW22, FZG23, FFL^{+}23b, FLHH23, GPH21b, GPH21a, GL23a, GLL^{+}22, GDS24, HZX21, HXZ24, HH24, HJJL24, HRZ^{+}23, JPZH23, JLZ^{+}22, LCS23, LXF20, LJP21, LJN^{+}22, LCJ^{+}22, LWZW20, LGD^{+}21, LC22b, LZD23, LHH24, LW20c, LS22b, LLZ24, MQW23, MZ24, MRW21, MH24, PZZF24, QZWY20, QYL^{+}20, QDW22, QSW23, QCCW23, QZC^{+}24, SHT^{+}21, SCT^{+}20, SYZ^{+}23, SWX24a, TGL^{+}24, WYW^{+}20, WWY^{+}20, WP21a, WCW21, WLZW21, WDZG22, WDW^{+}22, WCY^{+}22, WWZW23, WCWZ23, WCY^{+}24, WDZL24, WZW^{+}21, XL24b, YLW20, YSYS24, YWC24, YZPL24, YY24, YWJ^{+}20b, YSW^{+}21, ZLF22, ZLDY22, ZB23, ZH23a, ZRZL24, ZWNG21, ZPZH23]. Cooperative [GW20b, LHQ^{+}23, WJDL24, YYW^{+}23, ZDZS21, CWY^{+}24a, DYX24, GL24, HZJL24, INT20, JWL^{+}24, LDXZ23, SGX21, WCWZ23, WYCL24, XLZ23, YXW23, YZZ^{+}22, ZHQ^{+}24, ZTLW24, ZPLL20]. cooperative-competitive [LDXZ23, SGX21, YXW23]. Cooperator [INT20, ZAD21]. cooperators [LCS24]. Coordinate [SH21, TZLL24]. coordinated [LZX^{+}22b]. coordinates [ACS24]. coordination [LCL22b, MP22, MKZ22, XWWC21]. Copies [LG23, WJY21]. copolymers [CLZY21]. copula [HYS24]. copulas [BKMO22, HMEDVR24]. core [MSMK21, MS22b, WLL^{+}22, WC22c]. core-EP [MSMK21]. Coriolis [TCGR24]. corner [JRP22]. corona [KT21b]. Correct [Sha21, CRTVG^{+}23]. Corrected [SS23c, NKO^{+}23, ZBA^{+}23a, ZBA^{+}23b]. Correction [PGL^{+}23, SWS20, Wan23a, XHL20, YWL20, ZMS23a, SRN^{+}23]. corrections [CGLR23, OLW23]. correlated [ARMM23, CJC20, KNN20]. correlation [WW22a, ZH23a]. correspondence [SH21]. corresponding [SGL21]. Corrigendum [AV21a, AACGTB24, AFG22a, CF23b, Elg22, GNA22a, KWM20b, KA23, KFM21b, MM24, MV23, Oru22, PPS24, RRR20, SO21b, SZ22, WCY20a, WDG21, WHM^{+}21a, XH23, ZBA^{+}23a]. corrupted [Yoo21]. corruption [HH23a, LC22b, SWZZ23, ZBL^{+}23]. cos [WY24, BO21, JP22]. cosine [CWZ20, YLCL23]. cospectral [AA21a]. Cost [MB24, CYG20, CXGL22, Est21, HSL^{+}24, KMS^{+}20, KMS^{+}23, KLH24, LZZL21, LJR20, LPL21b, LCWZ24, PK20, QYL^{+}20, RSR21, SZXL24, Xia22, ZLZ^{+}23b]. Costa [LJHL21]. costly [INT20, NCY^{+}20, WCY^{+}22]. costs [HH20, KMS^{+}20, LYZ22b, TZ23a, Wet23]. cotangent [DF20]. Couette [AS22, DG20a]. Coulomb [Woo20, IHF13]. Coulomb-like [Woo20, IHF13]. count [ZL21b]. counter [KB21, XDS^{+}20]. counter-attack [KB21]. counter-game [XDS^{+}20]. Counting [BBW21, LY23a, SD23, FR23, GARD24, RY24]. Couple [DM23, Has20, MS22c, PJMW22]. Coupled [HHY20, ZJG21, BS22a, BZZ21, BK21, CK20, CYHR22, CWB22, DG21, DTL22, ESA20, GH21, HZZ^{+}21, HXZ21, HLZ22, LZY20, LWNZ22, LAS23, LZYL21, LM20d, MNSR22, MKP20, MVO21, NE21, PKJB22, RJHC24, SWS^{+}22a, SZ24a, SWW22, SY21a, STA24, jSZyW^{+}20, TWB22, TZW22, TTCY23, WMQ21, WMCZ21, WSZ20, YM21, YLHZ22, Yan23, YFM^{+}24, YL23b, YL24b, ZZX24, ZCS21, ZHYD20, ZW22b, ZTW22, ZZD22, ZWJ^{+}24, ZB22]. Coupling [AT24, SWP^{+}22, CH24a, CLSN24, CLH24b, FZS21, JA23, LZLM20, LTL^{+}20, LYN^{+}21, LXH24, LCL22b, LLT^{+}20, NPJ^{+}20, VBP22, VMZDV20, WJG^{+}20, WP21b, WDYS21, WGMJ23, XPP^{+}20, ZWHY20]. couplings [HB24, JWC20]. coupon [JSZ20, LOS21]. Cournot [Ask21, Ask22, YY21c]. Cournot-Bertrand [Ask21]. course [LaGBT22]. CoV [EA21]. Covariance [MXY^{+}20, YL23a, Wan21b]. cover [CLL22, HQZC24, HZCY23, LPL21b, LCWZ24, NXL21, RZZ21, SYW21, WHC20b, WF23b]. coverage [ZHJ^{+}21]. Covering [BBW21, TJAJAG21, JY23, TJIM20, WCS22]. covers [CF23a]. COVID [DKKJ21, KKS22b, MGFF21, NGC21, XYR^{+}23]. COVID-19 [DKKJ21, KKS22b, MGFF21, NGC21, XYR^{+}23]. Cowell [CY21a]. Cox [TZ23a]. Coxeter [Gas20]. CPSs [WSDW24, ZLC24a]. crack [GP23, YSL24]. cracks [WPJ22]. Craig [GLKrH20]. crane [AC20, AHM23]. Crank [ZL22a]. Creating [Rom21]. credibility [ZHLZ21]. credit [JJL23, LOG20, MKZ22, vdZGO21]. Creep [XYCG21, WGLL20]. crime [CGMFSHR21, Mon20]. crimes [Cro22]. criminal [RYR^{+}23, SS23d]. Criminals [SS23d]. crimo [IMOR24]. crimo-taxis [IMOR24]. criteria [AKY21, CCPT20, CGJ21, CZXyS24, DCR^{+}20, DSS22, GLWL22, JD20, LPKC22, MM20b, PP21, QLKW20, SS20, SZCC23, SWX22, WWWZ21, WHPQ23, WCX^{+}22, YL20c, ZYC^{+}24, ZCPZ20, ZYZ22c]. criterion [DL21c, LPP21a, LPP21b, WC22a, ZWW21a, ZWW21b]. critical [CHJW23, HBCA21, HAK20, HK21, OH24, TS22, WJL^{+}24, YWT20]. cross [BDKZ21, BLO22, DWT^{+}21, JM23a, KMUY21, MM22a, MM24, Mar24, MK21b, SBB^{+}22, XZSS21, ZCS21]. cross-diffusion [BLO22, KMUY21, MM22a, MM24]. cross-diffusion-fluid [ZCS21]. cross-domain [DWT^{+}21]. cross-efficiency [BDKZ21]. cross-immunity [XZSS21]. cross-regional [JM23a]. cross-section [SBB^{+}22]. crossing [CDD23, HW20, SS22a]. crosstalk [QCW24]. crowd [DZL^{+}24]. crowed [dJCGVM24]. Crowley [Lu22]. CRRA [ZY24b]. crypt [CFR22]. Cryptanalysis [LKX20]. Cryptocurrency [SNR^{+}22]. cryptography [BMDE21, TITN20]. cryptosystems [HLZ21]. crystal [LZY20, ZSS24]. crystallization [LTC^{+}23]. CT [CPMS22, WWZ^{+}23]. CTL [NNKH21, SWCT21]. CTL-mediated [NNKH21]. CTLs [BKR22]. cubature [LY22d, SV21b]. cube [YXFL23, ZHFL23a]. cubes [CLL23, DG20b, QH21]. cubic [AKP23, CHG23, IZDR24, LR20, SPX24, WHY20, XLL24, XT23]. cubics [ZN20]. CUDA [AFG21]. CUDA-based [AFG21]. cumulant [DD21b]. CUPL [FJJJ21, ZJJB22]. Curl [WCY20a, WCY20b, Han20]. currency [ASV22]. current [SN21]. curvature [Fre22, HYÇK22, SSR23, ZMD^{+}21]. Curve [ZZZ24, HLZ21, WL24b, YHL^{+}20]. Curved [SDD22, CGLR23, KA22, KA23, RT22]. curves [BCHRS23, FKVZ21, LPS21, MKK23, SS23a, SS24a, THZ23, Tre23]. curvilinear [RFFB23]. cushion [SSW22]. cuspidal [QCARL20]. Customer [SZXL24]. cut [HK21, HLP24, MS22a, YLH20]. CVA [AARS23, AO21b]. CVaR [BLNT22]. CVNs [LL22b]. Cyber [HD20, LWM^{+}21, SWS^{+}22b, CSS^{+}22, CC22, DZL^{+}22, DXZ^{+}23, FT24, FSS20, GSL^{+}22, GHP24, HLLS22, LLZB24, LLLS20, LJWD21, OPL22, QZH21, SAB23, SY21b, SCHS22, WLL^{+}24, WSDW24, XRZH22, YWC24, YCZ^{+}23, YXZ^{+}24, ZCS^{+}22, ZGZ^{+}23, ZWZX21, ZWZS22]. Cyber-attacks [SWS^{+}22b, CSS^{+}22, FSS20, WLL^{+}24, WSDW24, YXZ^{+}24, ZCS^{+}22]. Cyber-Physical [HD20, DXZ^{+}23, FT24, GHP24, HLLS22, LWM^{+}21, LLLS20, LJWD21, OPL22, SY21b, SWS^{+}22b, SCHS22, WLL^{+}24, XRZH22, YWC24, YCZ^{+}23, ZGZ^{+}23, ZWZX21]. cyberattacks [HLLS22]. cycle [CLW21, CLWZ21, HQZC24, HLLH23, LWZ22d, NXL21, QMS22, WHC20b, XZZC22, ZLL20]. cycles [CFR^{+}20, Che22, GT22a, GW20a, HZ20a, HZ23b, HJL^{+}23, JY23, KKB23, LR22, LLZ22b, QMS22, SS21a, WWY24, XLL21a, YXFL23, YML21, ZQ23]. Cyclic [SBV23, DWL^{+}24, HDQP23, PJ21, SK24b, WLJ22, WBV24, ZH22b]. cyclomatic [SS21b, SS22b]. cylinder [DMNV20, RT22]. cylinders [SC21]. cylindrical [IMIH22, KTEM22].

D [CG21, LQLF22, MTRA22, RO21, WDG21, ABPP23, AHM23, AC20, ADS23, BS22a, BRFLGH21, BFS20, BT22, BFRS24, CCCdLMR23, CG21, DS22a, DK22c, FRW^{+}24, Fre22, GCGM21, HT23, IZ20, JLJ23, JLOR22, LHET23, LW20a, LTC^{+}23, LL23b, LR20, LC23, LL20b, LZ22e, LZY^{+}23b, LHFH22, Ma21a, MVCK24, MV22, PBM20, PKJB22, PW22a, RWGS22, RLK^{+}20, SH21, jSZyW^{+}20, SFC23, TCZ^{+}20, UGG^{+}20, WZCF21, WGSS22, WHJK23, ZL22b, ZLU21]. D/ [CG21]. Dai [RO21]. daily [AS21c, WWZ20b]. damage [SHWW21]. damped [KMD21, SM21b, TTL20]. damping [AS21d, JZ20, LLS21a, SSCW22, ZLU21]. Darboux [LM20b, LWL^{+}20, RGAM20, ZJG21]. Darcy [Azi21, HKL22b, LQSZ21, Pap21, PR20, RM20, SN21, SKL23, XHL20, YL23b, ZyZW21]. dark [WHZ^{+}20]. Data [GR21, HREB21, ML22, XWDK23, ZTLL24, ABPP23, AS21c, BZ23a, BTL^{+}22, BDPV22, CSS^{+}22, CQ24, Chy22c, DDNS20, DR22, DZZ^{+}20, DS21d, FT24, GGW21a, GY22a, GTWL21, GLWL22, HKKL22, HQZC24, HZCY23, HCBY21, HB20, ISM^{+}23, JSCB20, JDLZ21, JOR^{+}21a, JLOR22, JHSL21, KYU22, KKPS23, KNW21, KKH23, LSPK20, LLY^{+}22b, LXW^{+}20, LHFW21, LM21, LWZ22c, LMS23, LLT^{+}20, NHBJ20, NK22, ON21, PHL23, QDL^{+}21, SG20a, SWX^{+}21, SZCP21, TVY21, VSM22, WZXY20, WWZ20b, WGL^{+}21, WHCW20, WCX^{+}22, XWMC20, XbYW20, XSP22, XXW^{+}20, XYlW^{+}24, XRZH22, YCL^{+}21, YL23a, Yoo21, YCZ^{+}20, ZZH^{+}20, ZDGL21, ZrLkL^{+}23, ZLZ^{+}21, ZY23b, ZGZ^{+}23]. Data-driven [HREB21, ML22, ZTLL24, CQ24, HCBY21, ISM^{+}23, KYU22, WWZ20b]. databases [AR20]. datacenter [YLPC22]. dataset [LKK^{+}23]. Davison [BBH21]. Davison-Maki [BBH21]. DC [ZBA^{+}23a, MP22, ZBA^{+}23b]. DCA [TLPT21]. DCell [LM21, LMS23]. DDEs [FY23]. DEA [MDJW20]. DEA-R [MDJW20]. Dead [GLP20, HL23a, JZS^{+}24, LNZ^{+}22, PX23, WM24a, WPX^{+}20, WWW24b, ZPDL20]. Dead-Zone [GLP20, HL23a, LNZ^{+}22, PX23, WM24a, WPX^{+}20]. dead-zones [WWW24b]. DeAngelis [LW21b]. death [GN20b, GN22, KWM20a, KWM20b, SRZK22, WZY21]. deblurring [DHZ^{+}20, GZG^{+}21, LMZ^{+}21]. Decay [BY21, FLC^{+}22, Ma21a, Ma23, WP21b, XZeF20, YYF24]. decay-rate-dependent [XZeF20]. December [Ano20c, Ano21c, Ano21o, Ano22c, Ano22o, Ano23c, Ano23o, Ano24c, Ano24o]. Decentralized [LLG22, WCXS22, YLL22, YXYZ23, Yoo21, ZXZ24, FT24, ZW22a]. deception [DTWJ21, JLP23, LHX^{+}22, LL23b, LGC^{+}24a, LLWW22, MFS22, SCHS22, WCM^{+}21, WYZX22, WCZ^{+}21, XZZ22b, XWZC21, ZQP^{+}22, ZCD23]. decimal [WGM^{+}21]. decimal-order [WGM^{+}21]. decision [BBEM24, Bri20, JP23, WJG^{+}20, WLS22, ZHLZ21]. decision-making [WLS22]. decisions [ZLZ^{+}23b]. decline [WHW^{+}21]. decoding [ZKLL23]. Decomposability [Prz24a]. decomposing [GW23a, Prz24b]. Decomposition [CFL^{+}20, BPT23, EV21, FMN21, FX20, GB21, HV24, HHR21, HLLY20, JS22, JLOR22, LM24, LFLH23, PLH^{+}24, TGDPS22, VGR^{+}21, WF23b, WZN22, YC21, ZLL20, ZJWV24, ZM21c, BFGGM24]. Decompositions [TNL23, DFX24, ZXC^{+}20b]. decoupled [CLZY21, CHY22a, Yu21]. decoupling [ZeFWD23]. decreasing [JGC^{+}21]. Deep [AO21b, KMM^{+}24, XCGN24, AACJ^{+}22, AO21a, CCZ24, DWT^{+}21, GW22, HH24, JW23, LJW^{+}21, PSSU21, WGW^{+}21, WHJ^{+}24, YL23b, AM24]. defection [LCS24, SWZZ23, YSYS24]. Defective [BZ23b, LZ22d]. defector [INT20, SC20c]. defense [CSS^{+}22, SWS^{+}22b]. defensive [GPH21b, GPH21a]. Deferred [SRN^{+}23, DSZ^{+}23, XHL20, YWL20, ZMS23a]. deficient [SS21c]. Defined [PZZF24, FKVZ21, Gas23, MBMA20]. definite [SD22a]. deformation [CdTPP22, LLH21, RLX22]. deformed [Bia21]. deg [DS22c]. Degenerate [TTCY23, CGSRG20, EFF23, HLZ21, HVZZ23, LLZW22, RKJ20]. degradation [CSC20]. degraded [DH23, WZZ23].

Degree

[KZ22, WY22, AKY21, BEIR23, BCR22, CMR20, DRS20a, DL21a, DDH^{+}24, DQC^{+}21, DBS^{+}23, DS24c, GG24, GMH^{+}20, GTWL21, GKK^{+}24, HPT22, KFR20, LLSY23, LZ24b, LTWY24, SS22b, Sha22, SZZ24, SS24e, WDC^{+}23, WHF20, WW23, WWWK24, WZL22, YCW23, ZH23a, ZCM^{+}24]. degree-based [DQC^{+}21, DBS^{+}23]. degree-degree [ZH23a]. degree-related [Sha22]. degree-sum [DDH^{+}24]. degrees [DGO22, FZ22, JY23, ST22, YG24]. dehazing [WHZ^{+}20]. Delay [AKDR23, AKR21, BY21, DCR^{+}20, LCLH24, LFY^{+}24, QLZL22, TLZ20, XWM23, XZeF20, ZLZ20, Alf21, ANS^{+}20, AS21b, ASAK21, AC20, AB21a, AS21d, AB20b, BB20, CZ21a, CZ22a, CJCL21, CHG23, CGJ21, CP20, CZ24b, CD22, CZGX22, DJWX21, DZDL22, DE21, DRS20b, DDH22, DL20b, DL21c, EW21, FMP20, FZLF23, FZW23, FZLJ24, FZ20b, FSLC22, FZ22, GZQ^{+}20, GJ21, GGY22a, GYC^{+}22, GGY22b, GSWL23, GWSC23, GLL23, GS21, Gök22, GTSV23, GNA22a, GNA22b, GH20b, HZJ^{+}21, HKKL22, Har22, HX22, HXZ21, HHZZ20, HYT21, HSY24, IS21, JD20, JMM20, JCCC20, Jor21c, JL24b, Kan20, KK23a, KKPS21, KGPS22, KÖY24, KLPL20, LKP22, LPKC22, LH20b, LGX22, LZZ24a, LSM23, LZHZ20, LXW^{+}20, LLD21, LLWK23, LV20, LZZ24b, MYW21, MM20b, NP20, NTT21, NH22, PMTT24, QLL24, RM23a, RM23b, RBPN20, SL20a, SZCC23, SR23, SZZL22, SLPH22, SZXL24, TMW20]. delay [UHZ^{+}20, WHC20a, WZXY20, WNP20, WJ20, WZZX21, WWWZ21, WHL21, WLPZ22, WM23, WHPQ23, Wu20a, XD21, XBLW23, XCLW21, XYLF21, XLL^{+}21b, YLCZ20, YYF24, YD22b, ZPSP22, ZZH^{+}20, ZZW21a, ZWH23, ZYC^{+}24, ZF24, ZH20a, ZLX20, ZCPZ20, ZYZ22c, ZLC20, ZJX21, ZZ20b, ZC21b, ZL22e, ZWW21a, ZWW21b, dOS20].

Delay-dependent [DCR^{+}20, LCLH24, XZeF20, CZ24b, FZ20b, GWSC23, LPKC22, RBPN20, SLPH22]. delay-difference [LZZ24a]. Delay-driven [TLZ20]. Delay-probability-dependent [QLZL22]. Delayed [XWM23, Alf23, BKR22, CZWS20, CZ22a, CH24b, CWB22, CLC^{+}22, CL24a, CLC24, CSL24, CLLA22, DCR^{+}20, DHS23, DWZ20, DZZ^{+}21, GLY24, GXZZ20, GL22, GZLY24, HP23, HSZ20, HLW^{+}20, HB20, LZH^{+}20, LCZW23, LCY^{+}21, LK21b, LCS^{+}22b, LZZ^{+}23, LZH21, PCN^{+}24, PLRD21, PN24, QYG^{+}20, QWD22, SJ23a, SWX22, SKH20, SKSC21, TWL20, WCG^{+}20, WCM^{+}21, WWWH21, WZC^{+}22, Wet21b, WHCW20, XZY^{+}23, YDSW22, ZSC^{+}21, ZXZC22, ZC23a]. delays [Alf24, BPG24, BMS21, CL22, CTZ21, CLZ^{+}22, CGZ^{+}22, CLLA23, CXYZ23, CTF24, CZ24b, CWWZ24, CLH24b, DHS23, EW21, FLZ23, FJFS21, FFL23a, GLTY22, GWY20, GX23, GW24, EVB^{+}21, HL24a, HZ23a, HQ22, HWL24, HH22, JDLZ21, JDZH22, JPZ^{+}21, LPP21a, LPP21b, LX20a, LGL20, LZS^{+}21, MR22, MS21a, MZSW21, NTT21, PP21, PKL23, PGZ^{+}24, QLZL22, RSVL21, SO21a, SAB23, SPZS20, SXC21, SWW21b, SWZS20, SCG^{+}20, SZWL21, SDW^{+}22, TDZ^{+}24, TW21b, TXZ24, VJA^{+}22, WCGW20, WXW^{+}20, WGL^{+}21, WWW^{+}21, Wan22b, WDTF22, WZLC21, Wet23, WCX^{+}22, XXL^{+}21, Xy20, XLYA21, XLZ^{+}24a, YWS24, YMW24, YTZ22, ZZL^{+}20a, ZLK^{+}22, ZLD22, ZLC24a, ZSP24, ZG24, ZLZ24a, Zha23b, ZLZ^{+}23a, ZW22b, ZZS^{+}22, ZTW22, ZLLX22]. deleting [LLW20]. delivery [CDL^{+}20]. demand [BBAS21, GP22]. demand-price [BBAS21]. demographic [CRTVG^{+}23, CLC20]. demographics [YWX^{+}20a]. demography [PLQ^{+}22]. dendrimer [XY22]. dengue [GGW21a, XZSS21]. denial [GY22a, GJS^{+}23, LWD21, LZWC22, LJWD21, LWY^{+}24, XXJtC24, YCZ^{+}23]. denial-of-service [GY22a, GJS^{+}23, LWD21, LZWC22, LWY^{+}24, YCZ^{+}23]. denoiser [WHJ^{+}24]. denoisers [CPMS22]. denoising [LMX22, LMZ^{+}21, UUS21, WHJK23]. dense [BBK^{+}24, CRTV22]. densities [GK20, GN22, HP20, RYG21, ŚW21, XCZ22, dAGSFM22]. Density [CFR22, KHNT22, ABTR21, BTL^{+}22, GYC^{+}22, GDB21, HZJ^{+}21, JNY^{+}21, JH21, Liu22b, LWJ21, MN22, Nak23, SJZA22]. Density-pressure [CFR22]. depend [RT20b]. depended [XD21]. Dependence [JX24, ZL20b, DP21, Guo22, JS22, LL23d, MJM20, STH24, WLW20]. dependent [AzR22, AIAFvH21, BFLR23, Cav22, CZ24b, CKW22, DCR^{+}20, DRS20b, FN22, FZ20b, GGW21a, GWSC23, GW24, HKKL22, HH20, JLFL20, JQJY22, JSM20, JHSL21, KMM^{+}24, KZC23, KPP22, LPKC22, LTW22, LZ22b, LM23, LCLH24, LLWY23, LHFW21, MP23, NK20, OLW23, PWZ22, QLZL22, QZC^{+}24, RW21, RDL23, RBPN20, SUN24, SLPH22, TYCK23, TWR20, VSM22, WNP20, WL21b, WLL^{+}22, XZeF20, XW22, XZZC22, YL20c, YZ23, YL23b, ZL20b, ZCH20, ZJZD20, ZYXL23, ZC21c]. depletion [LSL^{+}20, SWW21a]. depolarising [RMO22]. depth [CS21a, ZZY^{+}21]. derangements [ZGWZ22]. Derivation [DD20b, GMT20, MCD21, PR20]. Derivative [AEL21, CFA^{+}23, YYG20, YFW21, BA20, DZZ23, ELOY22, KL24, LK21a, MC22, RS22a, SZ24b, WL21b, ZMS23a, ZJG21, ZJZD20, ZL22d]. Derivative-based [YYG20]. derivative-term [ZJZD20]. derivatives [AS21a, AO21b, ADB20, BGRM21, KK23c, LWNZ22, LLP23, TZ21, WLB22, XL24a]. derived [LXLP20, TAU^{+}22]. Deriving [LZL20, Wen20]. descent [HD23, JW23, LZL^{+}21, SH21, TZLL24, WJK20]. descent-based [TZLL24]. described [Tre23]. Describing [GPHHC23, MVCK24, NWM23, SS23d]. description [KPPK23, MUOF22, WJDL24, ZKLL23]. descriptor [CL22, CZ24b, DZZ23, FSLC22, GZX20, HLW^{+}21, LLY^{+}21, LR23, SZ24b, SM23, ZZ20a]. descriptors [HMR23]. Design [LDCZ24, TKY23, WXWS20, ZHMZ21, ZZZ^{+}23, CZXyS24, DZL^{+}22, DMD21, GZX20, GT22b, HTF20, HDP^{+}21, HKL22a, HL24a, LGX22, LCZW23, LZS^{+}21, MSZ22, MMHD20, MZYW23, NK20, NK21b, NK22, PWZ22, SAB23, SNRA20, SZ24b, SC20b, TWR20, VSM^{+}21, VSM22, WSDW24, WWW24a, XZeF20, XZZW21, YLCZ20, YSLK23, YA21a, YWJ20a, ZD23, ZWY^{+}24, ZXZ20, ZZS^{+}22, ZC21c, ZQC24]. designed [DKLV23]. Designing [TZLL24, WYLQ24, JLL^{+}20, ZeFWW20]. designs [HLF22, LPL^{+}22, TYL^{+}20, TCZ22, Zho20]. Destabilization [PSE24]. destabilizing [LJZL22]. detect [BAL^{+}24]. detectability [GNSZ22]. detecting [LLZ^{+}20a]. Detection [GRMQ22, AGCGSC21, CHW^{+}24, DZL^{+}22, HHW23, KM22, LLZ22a, LWM^{+}21, LJW^{+}21, LL22b, LLZB24, LCC^{+}22, LYL22, ML22, SZG21, WHLL20, WGW^{+}21, WBY^{+}21, WWW^{+}22, WYLL22, YL20a, ZIML20, ZHG22, ZGLZ22]. deteriorating [CYG20]. determinant [AS23b, Kop22, RR07, TG20, ZJCA20]. Determination [AR23, KMZ22, CHG23, LW22a, LLS21b, RD23, WHL21]. determine [ST24]. determined [YWX20b]. Determining [BDPV22, AKP23, LAH22]. deterministic [GWSC23, WGJK21, WCZ22a, YL22, ZY20, ZYZ22b]. DETMs [WC22b]. DETMs-based [WC22b]. Developing [ZW21c, LKK^{+}23]. Development [LS22a, MGK23, SWFJ20, YWJ23]. developments [BKS21]. deviation [Zha24b]. deviations [CSW20]. devices [FN22]. DG [GÖRT23]. diagnosability [LZ20c]. diagnosis [DWT^{+}21]. diagnostic [AHS20]. diagonalizable [TTL20]. diagonalization [LD20c]. diagonally [ZZL20b]. diagram [RLK^{+}20]. Diameter [ADKS24, FLL^{+}20, GGS21, HCJC21, JL21, JT22, KWH^{+}23, LWZ22b, LD23b, LCWZ24, TX21, WZ20a, WHKT23, WLZD21]. diameters [HPL21, WMF23]. diamond [DK22c, FN22]. diamond-shaped [FN22]. dichotomic [SS20]. dichotomy [CS21b]. dictionary [LQ24]. dicyclic [LTWY24, YWFL23]. difference [AH21, Alm23, AT21a, Bon22, BZS21, CCPT20, CZL20, CKV21, CJ21, DX24, DTK24, FMP20, FHM22, GK21, HWCD22, JZ20, JCSH24, KÖ21, KM20, KG22, LC22a, LWZ22a, LZ24c, LZZ24a, LLX21, LHFH22, NE21, Oru21, Oru22, PW22a, QPHC21, QQ23, RO21, SJP22, SKR22, Sug21, WZ20a, WW20, WJC^{+}20, WDG21, WGSS22, WL23, WZ20b, Wu20a, XYM22, Zha21a, ZSFS22, ZHQ^{+}23, ZXZ24, ZGZ^{+}24]. difference/collocation [NE21]. difference/finite [BZS21]. differences [AF21, AA21b, LZ21, LtTBPA20, UGG^{+}20]. differencing [XCSW21]. different [AS22, BCLNV20, FN22, GGL^{+}23, GK20, GW20b, HLD^{+}23, LZCL23, LY22b, LZ24e, QQWJ22, RYR^{+}23, SZ24b, SRKA23, ST22, Wet23, YM22b, ZYZ22b, ZWY^{+}24]. differentiability [Sin20b]. Differential [JPY23, KWHS23, LRS20, ASR22, AS21b, AKR21, ASAK21, AFA24, AB21a, Baz20, BR20, BBH21, BB20, BMS21, But21, CFF20, CZ21a, Cha22, CGJ21, CdR22a, CD22, DS21a, DEHFF24, DDLW21, DO21, DYT21, DG23, ER22, ELOY22, ESA20, EW21, Ema21, ED20, FF21a, FLWL21, FRS22, FZ20b, GJ21, GGY22a, GGY22b, GLL23, GB24, GH20b, Har22, HX22, HXZ21, HY24, HWWW23, HAFM21, HVZZ23, INOP20, IS21, JC22a, JD20, JMM20, JS20, JGXL23, JT20, JSZ20, Jor21b, KK23a, KG22, KGPS22, KÖY24, KKLJ20, KMD21, LRCTR22, LK21a, LH20b, LHC21, LLQG21, LZSC22, LLY^{+}23, LFLP21, Liu21, LHZ21c, LY21b, LY21c, LNC22, LY22c, LFY^{+}24, MYC^{+}23, MWDD24, MDSS20, MKP20, MZLS22, MSN21, MV21b, MV23, MM20b, MN22, PMTT24, QXG21, QXSW22, RW21, RBPN20, RRS21, RG20, SO21a, SL20a, SS24d, SV23a, SWF20, SV21a]. differential [SJ20, SGL21, SV23b, SJ23b, SL24, TLC22, TY22, Ten22, TLY24, TZ23b, UHZ^{+}20, VJA^{+}22, VMZDV20, WHC20a, WWZ20b, WNP20, WZCF21, WWY21, WZGL21, WY23, WL24a, Wen20, WHL22, WHY22, XZ23c, YZC^{+}22, YZ21, YY22, ZPSP22, ZZX20a, ZF24, ZCJL21, ZS22a, ZLW23, ZG24, ZJX21, ZYX21, ZJX22b, ZS22d, ZLZ20, ZWW21a, ZWW21b]. differentially [SKS21]. differentials [TCA23]. Differentiated [Jan22]. Differentiating [ŚW21]. differentiation [EGMY23, FGMS21, Luo20, ZY21b, ZJX22b]. diffusants [NWM23]. Diffuse [GE21, CZ22c, SMW^{+}23]. Diffusion [HHY20, LWZW20, WGLS20, AD21a, Alf23, Alf24, AH21, dSAOG21, BS22c, BLO22, BFLR23, CGSRG20, CZLW24, CRT22, CGB20, CYHR22, CSG^{+}22, CSG^{+}23, CSC20, DS22a, DL21b, DTL22, DK22c, DH23, DH24b, EHA20, EFF23, GC24, GN23, GYC24, GK21, HWD20, HLZ20, HSM21, HXZ21, HPC24, HB24, Ily23, IMOR24, JHO^{+}21, JCSH24, KMUY21, KZWS22, KDB^{+}23, KK20, KHNT22, KSR21, LLZ22a, LF21, LWL23b, LGC24b, LR20, LZRF21, LCH20, LZZ^{+}20, LK21b, LGS22, LW22g, LZ22e, LtTBPA20, LOS21, LLS21b, LYN22, Ma21a, MZ23a, MM22a, MM24, MK21b, Nak23, NP20, NKA20, PSXX21, PGF20, PS21b, QL21, QSL21, Que22, RC22, SZ23, SLS24, SMTT23, SWW21b, SM22, SLZ20c, SSW21, SFL^{+}22, SHYS23, SKH20, SKSC21, TZ24, TA21, TZJ20, TZW22, VMZDV20, VN20, WW20, WSY21, WL21b, WZY21, WBV22, WJL22, WXLY23, WZZ23, WZW24, WZ20b, XbYW20, XYM22]. diffusion [XCZ22, XZZ20, XLL^{+}21b, YMRJ22, YGZ^{+}22, YCEW21, YLS23, YH24a, YAX^{+}21, ZT22a, ZCS21, ZL20a, ZZY20, ZYY20, ZL21a, ZFS22, ZL22b, ZZ24, ZGZ^{+}24, ZZX20b, ZJC21, ZJX22a, ZLS23, ZWZ^{+}22, vdZGO21]. diffusion-driven [ZZY20]. diffusion-wave [TZ24, XbYW20]. diffusions [HCM21, KNN20]. diffusive [Alf21, LCML20, Mef21, PGKM24, TC22, TM22b, Yu21]. diffusivity [LLS21b, Ma21a, TA21, WC21]. digital [CSV20, YLX^{+}21, ZHMZ21, ZH22b]. digraph [CZ22b, DML21]. digraphs [BDMB20, DLM20, FSS21, HCJC21, HLSL22, KP22, LW22e, LZ24b, LZC21, NMS23, WS24, XFK22, YW20, YM23b, ZW21a, CF23a]. dihedral [HZF21, LTWY24, LHL24, SNC22, YWFL23, ZXC^{+}20b]. Dilemma [FFL^{+}23b, BY23a, CWY^{+}21, CTWG24, DZ21, Elg20, Elg22, FLHH23, GPH21b, GPH21a, HZX21, HXZ24, HWY^{+}22, HC24, HZJL24, JGC^{+}21, LCS23, LCS24, LDZ20, LJN^{+}22, LSHH22, LWZW20, LJR20, LL23e, LZD23, LHH24, MRW21, MT21, PWH20, QSW23, QCW24, SHT^{+}21, SCT^{+}20, Tak21, Ued23, UU21, WWY^{+}20, WLZW21, YZPL24, YSW^{+}21, YZZ^{+}22, ZHQ^{+}24, ZWNG21, ZPZH23, ZWJ^{+}20, GLWW22, SK24b, WDW^{+}22, YSYS24, MHF23, YYW^{+}23]. dilemmas [DGL^{+}23, HH24, LJP21, LCJ^{+}22, SWP^{+}22, WP21a, ZLF22]. dilute [ZWJ^{+}24]. dim [LR21]. Dimension [BBW21, AKP23, CPPL23, FLR23, HLW^{+}21, KT20, KRT21, KMT^{+}21, Lia21, MS22a, NX23, NXG23, PMAK22, RY24, SVW23, SS21b, SJTCdlFL21, Wu20b, WWY22]. Dimensional [BBW21, WGLS20, AAS23, AFGG20, AM20b, ASS^{+}23a, AO21a, ALM20, AM24, BCLNV20, BF22, BLO22, BZS21, CCdlH21, CZL20, CJ21, DdAS22, DMSS20, ECS21, GLGLH24, GBF21, GLZB20, GZG^{+}21, HZJ^{+}21, HL23b, JCSH24, KY22b, Kri22, LHZ21c, LV20, LQNA23, MUOF22, Nak23, NS22, ÖY21, PH20, PSXX21, QHC21, QWD22, QL21, SM21b, SUN24, SSCG20, SMTT23, Sun22, SFL^{+}22, TCGR24, TBRB24, Ten22, TMA^{+}22, VBM20, Vyn23, WFHG20, WLY21, WMS^{+}22, WSZ22, XYY23, XLG^{+}21, XYlW^{+}24, YLG21, YHW^{+}23, YFHC21, YXXZ23, YK22, ZLA^{+}21, ZJJ24, Zha24b, ZRL^{+}24, ZGHM24, ZZX20b, ZyZW21, ZZY24, ZWJ^{+}24, ZQC24, ZL21c, ZLU21, dlCJ20]. dimensionality [DS21d, LLWY23]. dimensionally [RM20]. dimensions [AF24, GGL^{+}23, JL22, SS21a, SS22b, YL24a]. diode [YXZ^{+}24]. Diophantine [LY22b]. dipole [TCZ^{+}20]. Dirac [GST22, Gu20, LWCD24, WL23, ZJJ24]. Direct [LZH^{+}20, LL23e, Ahm21, BS22a, BA20, CXYZ23, KP23, QL21, WSC^{+}24, YF21a, SZCP21]. direct-co-direct [KP23]. direct-space-time [YF21a]. Directed [GRMQ22, HZF21, CZLH20, LLZL22, LHL24, WZLH22, XD23a, XWWC21, YW20]. direction [Dol21, HZT^{+}21, HLL22, WJK20]. directional [CG21, LD21a, TNTL21, XLYA21]. directions [LH24b, MZFX20, YZ24, ZW22a, dSTV21]. Dirichlet [ACC^{+}21, DMSS20, KM20, KK20, Pap21]. DIRK [ELOY22]. disabilities [FQHS24]. Disbandment [Kur21]. discard [HCL^{+}23, LX24a]. discontinuities [ARÁSY21, AY22, CLX22, MV21b, MV23, ZHQ^{+}23]. discontinuity [CU23, KRRS22]. discontinuity-capturing [CU23]. Discontinuous [MM20a, ASR22, ABPP23, CG20a, CG21, CWWZ24, CKV21, DK22b, DL21b, EFF23, FHM22, FWBY23, FY23, GN23, GHS^{+}21, Iza21, JWC20, LLQG21, LMÖ23, PS21b, QLG^{+}22, SLZ20c, WC21, WHWW21, ZMS23a, ZC23a, ZZX21, ZZX23]. discounting [PZ20]. discovery [ACBLC24].

Discrete

[JT24, LFLP21, MSDR21, OLW23, SZZL22, THW^{+}21, WJC^{+}20, BKU20, BDPV22, CFP22, CWH24, CHZ^{+}20, CL24a, CSS^{+}24, CFWL24, CSL24, CC24, CPLH20, CWJF21, DG21, DRS20b, DHS23, DWZ20, FLZ23, FM23, FLC^{+}22, GWY20, GHP24, GW23b, HD22a, HKCK24, HH22, HSY24, JC22b, JKJC21, JZ22, LSV21, LCXM21, LL22b, LWX24, LM20c, LWG20, LZLZ21, LLWK23, LR23, MSZ22, MTRA22, Nak23, NWM23, OPL22, PRP24, PKL23, PW22b, PN24, SAB23, SZG21, SGX21, SSF24, SCG^{+}20, SRZ22, TLC22, TZ24, TM22a, TGDPS22, TXZ24, VHB23, WW22a, WXF22, WSDW24, WWW24a, WFKL24, XXL^{+}21, XZZ22b, XSP22, XZQX20, XZQ23, YMW24, YW22b, YL20c, YMYR21, ZTLL24, ZZY20, ZSM20, ZS21b, ZLK^{+}22, ZDC22, ZWW23, ZGLZ22, ZLZ^{+}23a, ZHJ^{+}21, ZDX21, ZPW22]. discrete-interval [TGDPS22]. Discrete-time [OLW23, SZZL22, GW23b, CWH24, CHZ^{+}20, CL24a, CSS^{+}24, CFWL24, CC24, CWJF21, DWZ20, FLZ23, GWY20, GHP24, HH22, HSY24, JKJC21, LCXM21, LL22b, LM20c, LWG20, LZLZ21, LLWK23, MSZ22, MTRA22, OPL22, PKL23, PN24, SAB23, SZG21, SCG^{+}20, SRZ22, TXZ24, WXF22, WSDW24, WWW24a, WFKL24, XXL^{+}21, XZZ22b, XSP22, XZQX20, XZQ23, YMW24, YW22b, YL20c, YMYR21, ZTLL24, ZSM20, ZS21b, ZLK^{+}22, ZDC22, ZWW23, ZGLZ22, ZLZ^{+}23a, ZDX21]. discretely [HL22]. discretizable [LSC^{+}21]. Discretization [VBMB21, AP20, BM24, CCdS^{+}23, DS22a, GMN22, HKL22b, IR22, LLLC23, MHAT21, NRP22, Shi21, Zha20, ZBA^{+}23a, ZBA^{+}23b]. discretizations [GDB21, ZHYD20]. discretized [ZJX22a]. discrimination [TNT23]. discs [RiTZ24]. Disease [Zha22c, FMHDB22, LP22, LZ22a, Liu22b, TAU^{+}22, WHW^{+}21, WMCZ21, WZW24, YFM^{+}24]. disease-awareness [WMCZ21]. diseases [EKKE20, HTZ^{+}23, HCY21, RS23]. Disjoint [JY23, Che22, HQZC24, HZCY23, JJC24, LG23, LYhH^{+}24, NXL21, QMS22, SS21a, WJY21, WHC20b, YXFL23, ZHW22]. disk [Ach21, DJWX21, RLK^{+}20]. disks [HPL21]. disordered [UH24]. Disorders [LWL^{+}23c, LHC^{+}22, LWLA21]. dispersability [SYL23]. dispersal [EE22, NPE22, SJZA22, ZZ22a]. dispersion [CLX22, EKKE20, LSZ^{+}23, MJK22, OKT24, RC21, ZZX23]. dispersionless [ZHYD20]. dispersive [Kud20b, LMZ20]. displacement [WPE20]. DisPredict3.0 [UH24]. Disproof [GWW23]. dissemination [WZW24, ZJ23]. dissipation [JLJ23, KA22, KA23, Ma21a, PR20]. Dissipative [NSSA20, TW21a, ZC23b, CL24b, GLWL22, JZ20, LTL^{+}21, LR23, LMÖ23, MSZ22, NK21b, PAY22, SAB23, SDD22, SSW21, WCZY20, WP21b, WSL^{+}21, ZWZS22]. Dissipative-based [NSSA20]. dissipativity [FSLC22, GLD21, LPK22, LLT^{+}20, MR22, XXW^{+}20]. dissipativity-based [GLD21, XXW^{+}20]. dissociation [DM24, TT24, TLD22]. dissolution [OBP21]. dissolved [GP22]. Distance [LLCL23, MS21c, RGS23, YCW23, Dol21, DGMP22, Gan21, GRD23, GYM21, GV23, GN20a, GZ20c, HH23b, Kwo22a, LSC^{+}21, LXLP20, LYhH^{+}24, RD23, SSP24, SPX20, SPX24, WWZ20a, WM22, WWY24, XD23b, XLS21, YWFL23, ZL23c]. distance-based [GV23, LXLP20]. distance-gap-preserving [WWZ20a]. Distance-unbalancedness [MS21c]. distances [CY21b, KH23]. distancing [HCY21, KLP22]. distemper [CXLM21]. distinct [PS21a, WMWS21, ZL23c]. Distinguishing [NCY^{+}20, DDPS22, KP22]. distorted [WSY21]. distortion [NT24]. Distributed [CZLH20, CZW24, FSLZ23, GZWX23, GW23b, HCBY21, KY22b, LWD21, LHC^{+}22, LWL^{+}23c, MTFD24, MWS23, SCLD21, WDTF22, WC22b, XLZ23, XD23a, YXZ^{+}24, YL24b, ZZLL21, ZCZW23, ZWZX21, ZWY^{+}22, ZDD^{+}22, AD21a, CTZ21, CYL^{+}22, CLSN24, CZWX21, CXGL22, CXY^{+}23, DL20c, FTML22, GY22a, HZJ^{+}21, HLKC22, HH22, JL24b, KK23a, LX20a, LZWC22, LAS23, LGC^{+}24a, MAAEF21, MZYW23, PZ20, SWW21b, SFL^{+}22, Ued20, WYS23, XZZ22b, XWWC21, XZY^{+}23, YLC20a, YWR^{+}24, YP22, YTLV22, ZC21a, ZDZS21, ZL22c, ZCD23, ZLC20, Zha23b, ZQC24, ZLP21, ZD24]. distributed-order [AD21a, FTML22, SFL^{+}22, YWR^{+}24, YTLV22]. Distribution [IMS23, ASUH21, COPW23, GLL23, HYS24, KSH22, LW20b, LJH^{+}20, LWJ21, DDM24, NPE22, NK20, NK21b, ON21, RP20, RW21, SK24a, SJZA22, TITN20, Wri24, ZLH21, ZHJ^{+}21]. Distributions [KZZ20, CX22, Har24, JS22, VMR23]. Disturbance [LCL22b, SSKS21, SKLS22, WGZZ22, YSW^{+}22, ZeFWD23, ZQC24, GZL^{+}23, HKL22a, HCX^{+}24, HSL^{+}24, JY21, LCLS20, LX20a, LJ23, LCL22a, MXF^{+}24, SWFJ20, SSW22, WLQX20, WWWH21, XCLW21, XZQ^{+}22, XWWL23, YYWL23, ZWWH21, ZYSW22, ZL22e, ZXW20]. disturbances [ASS^{+}23b, CHZ^{+}20, CHHC24, CLSN24, DWS24, HQS^{+}24, xHhPjW^{+}21, KKS^{+}22a, LCS^{+}22b, MPX^{+}20, MC22, TDZ^{+}24, WC22b, WYS23, YLLL20, Yi23, ZWWH21, ZYSW22, ZSP24, ZQ24]. disturbed [GHS^{+}21, KEiS23, LSWL20, XHL23]. div [Agg22, AR21b, BS22b, TCPF24]. divergence [BD20a]. Diverse [CTWG24, LQS^{+}23, LWZW20]. diversity [MBLL24, Zhu23]. divide [CJ20]. divide-and-conquer [CJ20]. divided [AA21b]. Dividend [DD24b, LCH20, XZ21, ZCJL21]. dividing [WLHW23]. division [DS22c, GC20, QDL^{+}21, YTZ20]. divisors [HLZ21]. DMP [MGS20]. dn [qLgZ21]. DNA [CDD23, ZWZ^{+}22]. document [DH23, Fen22, JC22a, WZZ23]. DoD [MS22a]. Dodd [ZLA^{+}21]. DOF [WWT24]. domain [CJ21, DWT^{+}21, HLLY20, Lee20, LZY^{+}23b, MVCK24, MBMA20, MFA21, NNN24, SDD22, WYLW20, WHLL20, WJ20, WWZ^{+}23, YZC^{+}22, ZL22a, ZH22a, ZrLkL^{+}23]. domains [ACS24, BSRV23, CGLR23, KTEM22, VBMB21, WZCF21]. dominance [SK24b]. dominant [ZZL20b]. Dominated [KT21b]. dominating [CLH24a, DHZ23, MW20, SSA^{+}23]. Domination [BHR20, BGH24, RiTZ24, ACS20, AÁC^{+}21, ACSVT22, BHP21, BNR20, CHJW23, GLW20, GH20a, HCSJ24, HAK20, HK21, HPT22, IZDR24, KF23, LZC21, Lyu23, SD22b, YZW20, YS20]. dominator [BGHX22, KT21b]. donation [SWZZ23]. donations [WCW21]. dormants [AYKS22]. dose [XLS21, dFdOFdSBC20]. doses [LDZT24]. Double [LY20b, RSA24, YH24a, ACS20, ACSVT22, BD20b, CH20a, CSG^{+}22, CSG^{+}23, GW24, HOM22, IZDR24, LX22, MJK22, Mef21, MT24, PGYW23, QZH21, QXG21, SSA^{+}23, SK24a, TT23, TM22b, WSL^{+}21, Yu21, ZYXL23, Zhu23]. double-diffusive [Yu21]. double-edged [Zhu23]. double-layer [WSL^{+}21]. double-Layered [PGYW23]. doubled [CLW21]. Doublet [EM20]. doubly [DS21b, LLZW22, TLJO20]. Douglas [BZ21]. Dove [YM22b]. down [DGV22]. drafting [GK20]. drag [BS22b]. Drazin [CS24c]. drift [FJV21, GN23, PS21b, SMR22, WYLW20]. drive [CFWZ20, YLLL20]. drive-response [CFWZ20, YLLL20]. driven [ATDJ22, CXLM21, CYW21, CL23, CQ24, DDLW21, DGL^{+}23, ER22, HREB21, HGD^{+}21, HCBY21, HLD^{+}23, ISM^{+}23, KYU22, Kum23, LY23c, LZ24d, MYW21, ML22, Pap21, RWGS22, SK24a, SZP21, TLZ20, WHC20a, WWZ20b, YLW20, YWC24, YDSW22, ZTLL24, ZZY20, ZLZ^{+}23b]. driver [WLZD24]. drives [LZCL23, MP22]. driving [JLL^{+}20]. droplet [LYKK21, LYN^{+}21]. dropout [CZK^{+}22]. dropouts [CZWS20, HDP^{+}21, YCL^{+}21]. dropping [CA20, FSLZ23]. drug [CTZ21, KKR21, Liu22b, YI22]. drug-resistant [CTZ21]. drug-sensitive [CTZ21]. drugs [WOB21]. drunk [JLL^{+}20]. DSC [LD23a]. Dual [CH24a, GLX22, MKZ22, CL24b, DHM23, GLWX24, JJL23, LAS23, LQS^{+}23, LGC^{+}24a, LCH20, QH21, Vis23, WWW24a, YL24b, vdMvRS21]. Dual-channel [MKZ22, CL24b]. dual-CIST [QH21]. Dual-mode [CH24a]. dual-noise [WWW24a]. dual-relationship [LQS^{+}23]. duality [ASS24, YY21c]. ductal [DTK24]. due [CYG20]. Duffing [KLS20, MYW21]. duopoly [Ask22, YY21c]. duplex [ZJ23]. during [LTC^{+}23]. dwell [CZK^{+}22, GLD21, HMZZ20, JSM20, LLWW22, MFS22, WLHW23, YCL^{+}21, YL20c, ZGLZ22]. dwell-time [CZK^{+}22, GLD21, LLWW22, MFS22, WLHW23, YCL^{+}21]. DY [ZZW20]. Dynamic [ANS^{+}20, CXGL22, FLLW23, FW24, GWH22, HD20, JYMP^{+}20, JP20, JM23a, KSS23, LD21a, LGX22, LHZ^{+}23, LM24, LZHZ24, LYL22, LWL^{+}23c, MK21a, WCZY20, WHW^{+}21, WFYL21, WSDW24, WCZ^{+}22b, WSXL22, XZ23b, YWWY23, ZWXQ20, ZCL^{+}23, ZLC24a, ZCJ22, ZWZS22, AR21a, ADF^{+}21, Alf23, BBHM21, CH24a, CZ22b, Che21, CZS21, CC22, CLZ24, CWL^{+}21, DSY22, DXZ^{+}23, DW22, DH24b, GS21, Gök22, GNA22a, GNA22b, GZ20b, HREB21, HRLH21, HBCA21, HL23a, HZX22, HRZ^{+}23, JJL23, JLOR22, KM22, KET22, KLH24, LLZ20b, LCLS20, LXWH22, LZS^{+}21, LCC^{+}22, LWG20, LWL21, LFM^{+}22, LXZ23, LY20b, LWY^{+}24, MSZ22, MZY^{+}22, MTFD24, MZLS22, MUMS22, NNB21, QYL^{+}20, RDL23, SC20a, SZCP21, Tak21, TS23, WCW22, WYZX22, XYZY20, XJY24, YLW20, YXZ^{+}21, YHL^{+}20, YY21c, ZZCL20, ZWWH21, ZHLZ21, ZCW^{+}24a, ZRZL24, ZLN20, ZCLZ22, ZWHY20, Zhu23, ZXW20]. Dynamic-memory [ZWZS22]. Dynamical [KTEM22, KK22, LSYX21, Lu22, NS21, ZMJ^{+}23, BK21, CLCY23, CLLA22, DYA21, DFG22, DSS22, FR24, GMN22, HC20, HLL^{+}21, HGZ24, HXWC24, LZC22b, LLY^{+}22b, LWL^{+}20, LWG20, LK21b, LYXZ21, Mon20, SPM22, WLG^{+}21, XYH^{+}24, YWX^{+}20a, ZMNR^{+}22, ZWFW22]. Dynamics [BKR22, KZWS22, LGS22, PLW^{+}20, SY21a, TNT23, XGRM20, AKY21, AMM21, AKH^{+}22, ANG^{+}24, Ask21, Ask22, BLO22, BE22, CFR22, CLL22, CXLM21, CD23, CGW^{+}20, CWF^{+}21, CTZ21, CYL^{+}22, CWY^{+}24b, DJWX21, DNR21, DW22, DS21c, EHA20, EA21, Fan21, FHFS24, FMHDB22, FQHS24, GWZ^{+}21, GG22, GOO20, GS24, GWZ^{+}20, GSW^{+}20, HCH24, HZXT24, IG21, INT20, JRP22, KT21a, KKS22b, LCZ^{+}20, LXCZ21, LXP^{+}22, LW23a, LCC^{+}24, LW20b, LD21b, LW22f, LZ24d, LZYL21, LW21b, MWZX21, MHAT21, MBLL24, NA21, PJ21, PAY22, QZ22, RS23, RLX22, RK22, SG24, TS23, VS24, VBM20, WXW^{+}21, WMCZ21, WLJ^{+}21, WLWL22, WW22a, Wan22b, WMS^{+}22, WJL22, WWZW23, Wet21b, Wet23, WOB21, WX22, WYLZ24, XLZ23, XWL23, XZSS21, YZY22a, Yan23, YCL24, YS23, YOF21, YP21, YFM^{+}24, YM22b, Zar21, ZYW21a, ZSJ23a, ZJJ24, ZYySX23, ZLH21, ZT22b, ZBL^{+}23, ZZY24, ZJ23]. dynamo [DJWX21, Miz21]. Dynkin [Bia21].

E6 [Bia21]. E7 [Bia21]. E8 [Bia21]. each [HZCY23, QMS22]. early [HL22, JNY^{+}21]. early-exercise [HL22]. earth [YXX^{+}22]. Ebola [KWM20b, KWM20a]. eccentric [GKT^{+}23]. eccentricity [CWZ21, LLSY23, SLH20, WL22, YL20b, YLH22]. ECG [BAL^{+}24]. Echo [LZLZ21]. eco [CXLM21, KLH24, LW22f]. eco-epidemiological [CXLM21]. eco-evolutionary [KLH24, LW22f]. Ecological [SBV23, AGCGSC21, KB21, ZZCL20, ZT22b]. Economic [KRNG24, LC22b]. economical [PKJB22]. ecosystem [SP23, WBV24, Zha23b]. Edge [CZ22b, LG23, TPSP24, AS20, CHL24, DDPS22, FMT^{+}21, GW20a, GZZ21, HLM^{+}21, JWWL20, JMY22, KT21b, KMT^{+}21, LYhH^{+}24, LM21, PY21, SS21a, SS24b, WCY20a, WCY20b, WL20, WP20, WHY20, Xu24a, YW22a, YXFL23, YZM24, YJL20, ZZLL21, ZWJ^{+}20, Zhu23, ZZB24]. Edge-based [CZ22b, ZZLL21]. edge-colored [JWWL20, JMY22]. edge-coloring [HLM^{+}21]. edge-colorings [GZZ21, YW22a]. edge-connectivity [LM21, Xu24a, YZM24]. edge-contracted [FMT^{+}21]. Edge-disjoint [LG23, LYhH^{+}24, YXFL23]. edge-event [ZZLL21]. edge-primitive [WP20]. edge-strategy [ZWJ^{+}20]. edge-transitive [PY21]. edge-Wiener [WL20]. edged [Zhu23]. edges [GWSC23, WM22]. edit [KH23]. Editorial [Ano20w, Ano20x, Ano20y, Ano20z, Ano20-27, Ano20-28, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano20-34, Ano20-35, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano20-42, Ano20-43, Ano20-44, Ano20-45, Ano20-46, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-39, Ano21-40, Ano21-41, Ano21-42, Ano21-43, Ano21-44, Ano21-45, Ano21-46, Ano21-47, Ano21-48, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22-40, Ano22-41, Ano22-42, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano23-32, Ano23-33, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano23-38, Ano23-39, Ano23-40]. Editorial [Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano24y, Ano24z, Ano24-27, Ano24-28, Ano24-29, Ano24-30, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ano24-35, Ano24-36, Ano24-37, Ano24-38, Ano24-39, Ano24-40, Ano24-41, Ano24-42, Ano24-43, Ano24-44, Ano24-45, Ano24-46, Ano24-47, Ano24-48]. education [NS20]. Effect [DMNV20, GYC^{+}22, MRW21, Miz21, QSW23, RS23, RHN^{+}24, RDKR22, WDZL24, WZA^{+}24, ZLN20, AGA20, BS22b, DWL^{+}24, DL20a, DBS^{+}23, Gök22, HZXT24, JGC^{+}21, KHA22, LD21a, LG24, LD20d, LZD23, LW21b, MJK22, PSE24, RHR^{+}21, SRKA23, SLK^{+}23, TTMU20, WCWZ23, WHM22, ZLF22]. Effective [Fen22, GMN22, JLL^{+}20, MCT^{+}22, SWZZ23, TLPY20, WBY^{+}21]. effectively [CWY^{+}24a]. effectiveness [DKKJ21, KIJK20]. Effects [Alf23, BBEM24, CGW^{+}20, CWY^{+}21, CWF^{+}21, GPH21a, GZCS23, LJR20, LC22b, PR20, RHNT21, SYZ^{+}23, WMCZ21, WM22, BM22, BS22c, BP23, CLLA22, DDLW21, GDFNNR22, GWH22, JNY^{+}21, KET22, KY21, LSL^{+}20, LXF20, NNB21, RSVL21, SDD22, YWJ^{+}20b, YSW^{+}21, ZST22]. efficacy [WZA^{+}24]. Efficiency [FJ24, BDKZ21, LHET23, VS24, VMT24]. Efficient [Alm23, CLZY21, FdLPC^{+}22, HZT^{+}21, JCSH24, KWZX23, MFK22, RNVC21, SG20b, SS21c, YL23b, ZSS24, ZZ24, BDH^{+}23, BKK24, BFS20, BZS21, CY21a, ER22, GDB21, GdFPF24, GM21a, Grz22, HWXC20, HOM22, HL22, IZ20, IS21, JLFL20, KKK22, Kar22, LMX20, LZ23, LAH22, LMZ^{+}21, LX23b, LWL24, LG22, LQNA23, MDSS20, MDJW20, PSXX21, PGPCK21, QDL^{+}21, SL20a, SW20, SG21, TYCK23, TBRB24, VPK20, Wan21b, WHJK23, WKM24, YTC23a, ZHQ^{+}23, ZXF20]. effort [JC22b]. EFGM [LLLC23]. eigen [GJWV24, WQZD23]. eigen-problems [GJWV24]. eigendata [SRK23]. eigenpair [LM22, XL24a]. eigenpairs [LM22]. eigenparameter [ZL20b]. eigenproblem [GGY23]. eigensolvers [HR23]. Eigenvalue [XH22b, XH23, AK21, BSCV24, CWY^{+}20, COPW23, FWBY23, GLZA21, Han20, JLD^{+}24, LW22a, LZ22d, MV21a, MM20a, SY22b, TLCM24, TTL20, WC24, XLS21, YHB^{+}20, YWX20b]. Eigenvalue-free [XH22b, XH23]. eigenvalues [BBT21, Cas23, CFLL21, FLS20, GRD23, HT24, HV24, HLX21, HWLL21, HCS23, HDZ22, JX24, MH20, MBMA20, PW22b, SY22a, SLJ23, Sta20, Wan22c, XY22, ZL20b, ZL23c, Zha23a]. eigenvector [LDZ21]. eighth [Alm23]. eighth-order [Alm23]. eikonal [Qi22]. Einstein [TCZ^{+}20, XHWM20, ZGHM24]. Einstein-condensate [ZGHM24]. elastic [Bab22, DSY22, LW22c, MK21a, SFC23, WPE20, WSZ22, YHB^{+}20]. Elasticity [DLQY20, AP22, PBKS23, PWWZ23, QHC21, TCPF24]. elastomeric [ME21a]. elastoplastic [CK21, LWQW22]. election [CM20b]. electric [LLQW24]. electrical [ASS^{+}23a, RHR^{+}21]. electrically [ABD22]. electricity [BO21, CJC20, HXL^{+}20]. Electro [KY24, LSL^{+}20, NHK^{+}22, NHK^{+}23, PPS24, RRR19, RRR20, VBM20]. electro-osmotic [LSL^{+}20]. electro-osmotically [NHK^{+}22, NHK^{+}23]. electrocardiograms [ZIML20]. electromagnetic [BL22, LSV21, LW20b, RDKR22]. electromechanical [MKKG22]. electromyography [XFW^{+}21]. electron [HLLW21]. electronic [GLD21, RCG21]. Electroosmotic [ATDJ22]. electrothermal [MS22c]. Element [WGLS20, AB23, AP22, AKH^{+}22, BZS21, BRMVCD21, BDRM23, CLL22, CHY22a, CZ22c, CCE^{+}22, DG21, DSY22, DL21b, Eft21, GP23, HHGZ23, Iza21, JGXL23, KMN22, KAS21, KPP22, Kum24, KK24a, LLF^{+}23, LF21, LW22c, LLLC23, LGC24b, LGK^{+}22, LZ22e, LHFH22, LZ22f, LZ22g, MZX22, ME21a, PBM20, PWWZ23, QZ22, SM21b, SL20c, SL20b, SW22, SZ24a, SLZ20a, SWX24b, TLC22, TCPF24, TM20, TZ23b, Voe23, WCY20a, WCY20b, WSZ22, XZZ22a, XTT22, XHWM20, YHW^{+}23, YJ20, YFHC21, YXXZ23, ZL22a, ZT22a, ZL20c, ZL20a, ZSL20, ZL21a, ZMD^{+}21, ZF21, ZL22b, ZF23a, ZXF20, ZS20a, ZS22c, ZZL^{+}24, ZWJ^{+}24]. element-finite [AB23]. element-free [DSY22, LLLC23, MZX22, SWX24b, ZL20c]. elements [DDGN23a, DDGN23b, GKK^{+}24, GLC23, LePP24, MV22, PT21, SY22b, VBM20, WDC^{+}23, WHY20]. ellipsoidal [AS21c, RWGS22]. ellipsoids [Dol21]. elliptic [AMPS20, ACS24, AP20, BFGGM24, FLZZ24, FHM22, HL24b, HLLY20, IP23, LLP23, RRS24, SJP22, WZCF21, XZZ22a, ZL20c]. elliptical [SBB^{+}22]. ellipticity [DLQY20, LL21, Zha23a]. Elo [WYW^{+}20]. elongated [TMA^{+}22]. emanating [BHGPC21]. embeddable [GPW24]. Embedding [QMS22, YHD24, Zhu24, WM22]. embeddings [ZDMW23]. Emden [Iza21, IS21, SS24c]. emerge [ZHWS24]. emerged [YTX^{+}24]. Emergence [MUMS22, WP21a, NCY^{+}20, PZZF24, WYW^{+}20, WCY^{+}24, ZRMD21]. emergency [YYW^{+}20]. Emotion [FLHH23, CWY^{+}21]. Emotion-based [FLHH23]. emotional [CWY^{+}24b, LL23e, YXZ^{+}21]. emotions [BYCX24]. empirical [HMEDVR24]. employing [RKRRL21]. Empty [JLZ^{+}22].

Enabling [PNOdV20]. ENATE

[LtTBPA20]. encoding [WJDL24, ZKLL23]. encoding-decoding-based [ZKLL23]. encroachment [HWZZ23]. encryption [JYMP^{+}20, LHET23, jSZyW^{+}20, TEGP24, WLZY21, ZQ24, ZWYL22, ZWZ^{+}22]. endpoints [OK21]. ends [BSS21]. energies [AHS20, MV22]. Energy [BSdCCGCV21, LW22c, LTW24, LX24a, Ma24, VGdFCBA20, YZAM20, Alm23, CZW24, CCE^{+}22, DG21, Gan21, GGS21, GLZA21, GA21, GWH24, GT22b, HH20, JJL23, JT24, KTEM22, LHYL24, LZX^{+}22b, LH24a, MCD21, SN21, SZ24a, SY21b, TZ23b, UKA23, WM22, XCSW21, XY22, YW20, YLG21, YWXL21, ZSS24]. energy-based [GWH24]. energy-conservative [YLG21]. energy-conserving [KTEM22]. energy-harvesting [LTW24]. Energy-preserving [LW22c, Alm23, XCSW21]. energy-stable [LHYL24, SZ24a]. Energy-to-peak [LX24a, GT22b]. enforced [KKPG20]. enforcement [BCCI20]. engineering [CWZ20]. engines [KM22]. Enhanced [LGD^{+}21, CWZ20, LTC^{+}23, NYHC24, PLH^{+}24, QMS22]. enhancement [AGC20, CPMS22, ME21b]. enhances [HXZ24, HDQP23, KLH24, MT21, ZAD21, ZLDY22]. Enhancing [CCF24, CFPM23, LZ24d, WYCL24]. enlargements [HP20]. ENO [MT23]. ENO-ET [MT23]. enriched [CLL22, DDGN23b, GLC23, KK24a]. Enrichment [DDGN23a]. ensemble [CQ24, YCZ^{+}20]. entanglement [FZ20a]. enterprise [RSXH20]. entities [LMT24]. entries [TJIM20]. entropic [ASUH21]. entropies [DESTP22, DQC^{+}21]. Entropy [AR24, SN21, BAL^{+}24, Bri20, CSW20, DBS^{+}23, FZG23, FWM^{+}22, GÖRT23, Has20, MS22c, PD23, SIK22, UMT22, VMR23, WZZ23, YS23]. Enumerating [TT24, LYhH^{+}24]. Enumeration [AA21a, LZ24a, LQW20, MY24, SLL^{+}22]. envelope [AC22, WC22a]. envelope-constrained [WC22a]. environment [AEE21, Alf21, ES23, GCLHT23, LCML20, Lu22, MQW23, PRP24, PNOdV20, Xia22, ZLDY22, Zhu23]. Environmental [GPH21b, LW22f, ZWNG21, HZXT24, IG21, JPZH23, LLZ24, SCD23, SK24b, WPH20]. Environmental-based [GPH21b]. environments [XLL^{+}21b, ZYZ22b, ZHWS24]. EP [MSMK21]. Epidemic [CS20, JL24b, SH22, WMS^{+}22, WQZD23, ASAZ23, AFG22a, AFG22b, CRTVG^{+}23, CGW^{+}20, CH22b, CWL^{+}21, CLD^{+}24, GCLHT23, GSW^{+}20, HSLS20, HZJ^{+}21, HWC23, HHW23, HLD^{+}23, HCM21, KAT22, KKS22b, LXH24, MCSD21, MK21b, PLQ^{+}22, RP20, RHNT21, SK24a, WJG^{+}20, WLG^{+}21, WW22a, XWLL20, XW22, YWX^{+}20a, ZYPL20, ZYZ22a, ZLD22, ZHJ^{+}21, ZLW^{+}21b, ZLZ^{+}23b]. epidemics [DS24b, KWM20a, KWM20b, KLP22, WCWZ23]. epidemiological [CXLM21, LDZT24, SS23d, VMZDV20, ZM21a]. epitaxial [HXF23]. Epstein [TS22, ZH20a]. equal [GA20, ZS20a]. equal-order [ZS20a]. equalities [LSPK20]. equality [CG20b, FSLZ23, KLPL20, LD20a, YL24b]. Equation [IMS23, WYD20, WGLS20, AD21a, AF21, AKMR21, AGGPO23, AzR22, AEL21, AH21, Alj21, Alm23, ARST21b, AJIR21, AFA24, AF24, ACS24, ADS21, ACH^{+}22, AB20b, BKU20, BL22, BM20, BFS20, BSS21, BD20b, BZS21, CPS24, CGSRG20, CS24a, CRT22, CWZZ22, CLX22, CM22, CHZ24, CSZ24, CZR21, DdAS22, DS22a, DX24, DMSS20, DYA21, DR22, DH24b, ED20, FKS23, FLR23, FZ20b, FCX21, GA20, GP23, GC20, GLZB20, HLLW21, Har22, HWCD22, HSM21, HXZ21, HZXZ20, HCY22, HWWW23, HAFM21, Ily23, JLFL20, JCSH24, JL24a, JSZ20, JT24, Jor21b, Jor21c, KMN22, KYU22, KM20, Kud20a, LGQ20, Lee20, LC21, LYKK21, LL22a, LM20b, LF21, LW22c, LW22d, LWCD24, LGC24b, LLZW22, LLWY23, qLgZ21, LLFZ20, LY21b, LY21c, LLZ22c, LZ22e, LY23c, LV20, LLS21b, LZYL21, LFY^{+}24, Ma23, MWDD24, MFK22, MZL24, MP23]. equation [MCD21, NP20, NI21, NA21, OPP22, Oru22, OKT24, ÖY21, PZS24, PGF20, PW22a, QNX^{+}23, QWD22, QXG21, QXSW22, QYD22, QXX21, RVV24, RS21, RO21, SM21b, SL20c, SL20b, SZ23, SJP22, SG21, SV21a, SV23b, SLZ20c, Sun22, SX23b, TLC22, TY21, TA21, TLDL21, TWWL24, TO20, UZ20, UUS21, UGG^{+}20, UK23, VPK20, VBMB21, WZY20, WDG21, WWY21, WC21, WXLY23, WH21, WP21b, WZ20b, XbYW20, XZ23a, XLJL22, XYM22, XZZ20, XZ23c, YSL24, YY21a, YZLS20, YLG21, YAX^{+}21, YTZ20, YLZL21, ZJG21, ZL20a, ZZY20, ZHYD20, ZSL20, ZC21a, ZLA^{+}21, Zha21a, Zha22b, ZSFS22, ZLLY22, ZF23a, ZSS24, ZJJ24, ZXF20, ZZX20b, ZLW21a, ZLJ20, ZZY24, ZXB20, ZLS23, ZC22]. equation-based [MFK22]. Equations [CUA^{+}22, KWHS23, ASR22, ABPP23, AAS23, AMPS20, AM20b, Alj21, AS21b, ARST21a, ASAK21, AHB20, AB21a, AP21, AHEGM23, BS22a, BDPV22, BDH^{+}23, Baz20, BR20, BBH21, BG20a, BS22c, BG20b, BB20, BM24, BRB24, BFRS24, BLS21, BCVV22, BDRM23, CZ21a, CG23, CG20a, CU23, CKN21, CN23, CCPT20, Cha22, CGJ21, CHY22a, CdR22a, CKV21, CKWY23, CD22, CPPL23, CCSPMV23, DS21a, DDLW21, DP24, DO21, DDE21, DDR21, DYT21, DG23, DDH22, DFG22, DNR21, DL21b, DS21b, ER22, ELOY22, ESA20, EHVMM24, FF21a, FMP20, FRRS24, FRS22, FG24, GJ21, GGY22a, GGY22b, GLL23, GJJC^{+}23, GN22, GTSV23, GB24, GNA22a, GNA22b, DDCM23, GWH24, HYÇK22, HA20, HLZ20, Har24, HWC20, HX22, He23, HVYMS21, HXZ21, HM20b, HJY24, HD22b, HVZZ23, Iza21, JC22a, JD20, JPY23]. equations [JMM20, JS20, JZ20, JLJ23, JHO^{+}21, JCH^{+}20, JT20, Jor23, Jor24, JZ22, Kan20, KÖ21, KK23a, KG22, KY22a, KGPS22, KÖY24, KKLJ20, Kod21, KPG22, KKS22b, KIB20, KCT^{+}21, Kud20b, KK24b, LLT22, LK21a, LaGBT22, LLZ22a, LH20b, LHC21, LLS21a, LLQG21, LWNZ22, LZSC22, LM23, LLY^{+}23, LZ23, LWX24, LHYL24, LZ20a, LC23, LFLP21, Liu21, LHZ21c, LLX21, LNC22, LY22c, LRS20, LYN22, LHFH22, LMÖ23, LM20d, Ma21a, MPM23, MMK^{+}23, MCT^{+}22, Mar22, MDSS20, MKP20, MSN21, MV21b, MV23, MM20b, MVO21, MN22, MSE24, NNN24, NSD^{+}23, OiEF22, Oru21, PH20, PSXX21, PBM20, PKJB22, PMTT24, Qi22, QL21, QSL21, Que22, RKJ20, RD24, RW21, RRS21, RG20, RSA24, RT22, SO21a, SL20a, SS24c, SS24d, SWS20, SW22, SSF24, SZ24a, SD22a, SLZ20a, SV23a, SWF20, SWW22, SJ20, SXC21, SGL21, SYX22, Sug21, SFL^{+}22]. equations [SL24, TZ24, TY22, TC22, Ten22, TZJ20, TZ23b, UHZ^{+}20, VJA^{+}22, VMZDV20, WHC20a, WJC^{+}20, WFHG20, WWZ20b, WNP20, WMQ21, WSY21, WZCF21, WZGL21, Wan22a, WGSS22, Wan23a, WL24a, WM24b, Wen20, WSZ20, Wu20a, WHY20, WHY22, Xy20, XTT22, XHL20, YGM^{+}22, YM21, YTC23b, YWZT20, YL22, YLS23, YJ20, YTZ22, YYF24, YLL20, YK22, YY22, ZHL24, ZL22a, ZT22a, ZZX20a, ZF24, ZM20, ZW21c, ZF21, ZS22a, ZFS22, ZLLY22, ZLW23, ZG24, ZRL^{+}24, ZGZ^{+}24, ZO21, ZH20b, ZJX21, ZZX21, ZYX21, ZJC21, ZJX22a, ZJX22b, ZZX23, ZZL^{+}24, ZLZ20, ZWW21a, ZWW21b, ZZZ^{+}23]. equidistributed [KSR21, MWDD24]. equidistribution [GK21]. equilateral [TPSP24]. equilibria [BG20b, HY24, PZ20]. equilibrium [CLSN24, GH21, JNY^{+}21, Ma23, PJMW22, TTMU20, Ued20, Ued23, UU21, ZSZL20]. equitable [AACI21]. equity [KLP22, WZY21]. equity-linked [WZY21]. equivalence [SGL21]. equivalences [AE21]. Equivalent [AKDR23, DHS23, RKRRL21, YXXZ23, YSW^{+}22]. equivalent/check [RKRRL21]. Erdös [MMMBRS20]. Ergodic [RP20, ZHJ^{+}21, Kut20]. Ergodicity [ZRSK21]. Eringen [LD21a]. Erlang [MB24]. Erlang- [MB24]. Erratum [LPP21a]. Error [CD20b, CHY22a, EBJ20, FGM24, HYÇK22, HLZ20, LD20b, NE21, SSF24, TZ23b, WDC^{+}23, ZMS23b, AD21a, AK21, BFQ20, BFLR23, CCC20a, CH20a, CZ22c, DD24a, GBH^{+}21, JW23, KX23, Kwo22a, LXL22, LWCD24, MF21, OPS20, SRGS20, SY22b, SM23, WGZZ22, WLD^{+}22, WKM24, WWT24, YL23a, YP22, ZSFS22, ZJJ24, ZZL20b]. error-feedback [SM23]. errors [Ily23]. essentially [TFL22]. estimable [QYLW24]. estimate [CH20a, Ma23, ZSFS22]. estimated [SSW22]. estimates [BB20, CZ22c, CKW22, EBJ20, GH20b, HYÇK22, HLZ20, KHNT22, SLCS20, SY22b, Wri24, ZJJ24]. Estimating [ES23, ZS22a, EA24]. Estimation [AT21b, FZLF23, MD23, WHM22, AD21a, ADS23, BTL^{+}22, BFLR23, CH24b, DDLW21, DZZ^{+}20, FZLJ24, FSLZ23, GGY22b, GY22a, GXZ^{+}24, GTK^{+}22, HLW^{+}21, HSZ20, JPZ^{+}21, KMM^{+}24, KDB^{+}23, LLZ22a, LWM^{+}21, LWD21, LLY^{+}21, LL22b, LL23b, Liu21, LY22c, LWZ22c, LCS^{+}22b, LWL24, LJWD21, MGFF21, MFA21, MZYW23, NGS22, PCN^{+}24, QLZL22, RW21, SCC24, SWS^{+}22b, SWZS20, SSW21, TWL20, TY22, TQZ^{+}24, TW21b, WDC^{+}23, WXW^{+}20, WFYL21, WLB22, WC22a, WB23, XD21, YL23a, ZR21, ZFJ^{+}20, ZrLkL^{+}23, ZWZX21, ZL22e, ZLLX22]. Estimator [GL24, HKL22a, JTC^{+}24, KKPS21, LZDX22, TWL20, XLZ^{+}24b]. Estimator-based [GL24, HKL22a, JTC^{+}24, KKPS21]. Estimators [CDP23]. Estrada [LPL21a, WX20a]. ethylene [AGC20]. Euler [BSS21, FJLG22, GW23b, JTC^{+}24, LSWL20, LH20b, LHC21, LHX^{+}22, LZ23, LMÖ23, PGZ^{+}24, SGL21, WJL^{+}24, ZMD^{+}21, ZJW^{+}22, ZZL^{+}24]. Eulerian [ABPP23, BDRM23, HM20a, MCD21, TM22a, ZW21a]. European [ARISV20, ASV22, LH20a, Wan24b]. evacuation [DZL^{+}24, FQHS24, GYC^{+}22, GW20b, RYG21, TS23, YYW^{+}20]. evaluated [AHS20]. Evaluating [SM22]. Evaluation [MZ24, MCT^{+}22, BDKZ21, CL20, GMH^{+}20, GN22, HR20a, HZJL24, Kun24, LZX22a, Lóc22, MV22, RKRRL21, WMT22, XSP22, Xu24a]. Evaporation [BS22c, LYKK21, SGI20]. even [BTP20, CLW21, CLWZ21]. even-cycle-free [CLW21, CLWZ21]. Event [CMYF21, CLWX23, CZWX21, DYX24, DWZZ22, GC24, GSL^{+}22, HZX22, HMZZ20, KMLW22, LDX20, LZC22b, LZW^{+}20b, LZ24d, MWS23, PLRD21, QQY^{+}23, SYL21, SYTC24, TWB22, TQZ^{+}24, WCG^{+}20, WCM^{+}21, WLZF21, WC22a, WPX^{+}20, WYS23, WWW24b, XXL^{+}21, XWZC21, XZY^{+}23, XZQX20, XWWL23, XRZH22, YXMW22, YM23a, YL20a, YCZ^{+}23, YZ21, YZ24, ZeFWW20, ZYSW22, ZCL^{+}23, ZCW^{+}24a, ZPLL20, ZB22, ZLLX22, AS23a, BZZ21, CZLH20, CDZW21, CZ22b, CNZ^{+}21, CC22, CTF24, CXGL22, DSZ^{+}23, FSS20, FLLW23, GZLY24, HRLH21, HLL22, HL23a, HJW^{+}23, HSL^{+}24, JY21, JLP23, JJSY22, KY22b, KJL20, LWD21, LZ22b, LGX22, LHC^{+}22, LZWC22, LLY22a, LHZ^{+}23, LWL^{+}23a, LCZW23, LR24, LM24, LGC^{+}24a, LSM23, LTL^{+}21, LYXZ21, LNZ^{+}22, LYL22, LSC^{+}23, LLQW24, LJW^{+}22, LR23, LX24a, LY20b, MSZ22, MZY^{+}22, MTFD24, QZH21, RJHC24, SKLS22, SWS^{+}22a, SWX22, SZSZ22, SLPH22, TLHH21, WCZY20, WFYL21, WWW^{+}21, WCW22, WYZX22, WPZY23, WYLQ24, WSDW24, WCZ^{+}22b, XBLW23, XCW^{+}22, XLZ^{+}24b, YYG20]. event [YWY^{+}22, YCM^{+}22, YWWY23, Yan24, Yoo21, YP22, YZY22b, ZZCW21, ZDZS21, ZZLL21, ZYXL23, ZLZ^{+}21, ZCLZ22, ZCZW23, ZCJ22, ZWY^{+}22, ZWZS22, ZXW20]. Event-based [CMYF21, DYX24, TQZ^{+}24, WCG^{+}20, WCM^{+}21, WC22a, WYS23, XWZC21, MSZ22, WCZY20, ZWZS22]. Event-driven [LZ24d]. Event-Triggered [CLWX23, MWS23, QQY^{+}23, ZCL^{+}23, ZLLX22, CZWX21, DWZZ22, GC24, GSL^{+}22, HZX22, HMZZ20, LDX20, LZC22b, LZW^{+}20b, PLRD21, SYL21, SYTC24, TWB22, WLZF21, WPX^{+}20, WWW24b, XXL^{+}21, XZY^{+}23, XZQX20, XWWL23, XRZH22, YXMW22, YM23a, YL20a, YZ21, YZ24, ZeFWW20, ZYSW22, ZCW^{+}24a, ZPLL20, ZB22, CNZ^{+}21, AS23a, BZZ21, CZLH20, CDZW21, CZ22b, CC22, CTF24, CXGL22, DSZ^{+}23, FSS20, FLLW23, GZLY24, HRLH21, HLL22, HL23a, HJW^{+}23, HSL^{+}24, JY21, JLP23, JJSY22, KY22b, KJL20, LWD21, LZ22b, LGX22, LHC^{+}22, LZWC22, LLY22a, LHZ^{+}23, LWL^{+}23a, LCZW23, LR24, LM24, LGC^{+}24a, LSM23, LTL^{+}21, LNZ^{+}22, LYL22, LLQW24, LJW^{+}22, LR23, LY20b, MZY^{+}22, RJHC24, SKLS22, SWS^{+}22a, SWX22, SLPH22, TLHH21, WFYL21, WWW^{+}21, WCW22, WYZX22, WPZY23, WYLQ24, WSDW24, WCZ^{+}22b, XBLW23, XCW^{+}22, XLZ^{+}24b, YYG20, YWY^{+}22, YCM^{+}22, YWWY23, Yan24, Yoo21, YP22, YZY22b, ZZCW21, ZDZS21, ZYXL23, ZLZ^{+}21, ZCLZ22, ZCZW23, ZCJ22, ZWY^{+}22]. Event-triggered-based [YCZ^{+}23, LX24a]. event-triggering [LSC^{+}23, MTFD24, QZH21, ZXW20]. events [LZZ21]. every [FMT^{+}21]. evidence [HGD^{+}21]. evil [XL24b, ZPZH23]. Evolution [CYW21, CLX22, FH24, GDS24, GLLZ24, HZX21, HRZ^{+}23, MQW23, Wan24a, WZW^{+}21, YLW20, BY23b, CRTVG^{+}23, CWY^{+}21, CZW23, CC24, CCSPMV23, Cro22, DCCM20, GPH21a, GLWW22, GL23a, GW20b, HH24, JPZH23, Kur21, LXF20, LZZ21, LHQ^{+}23, LC22b, LF24, LHH24, LW20c, LS22b, MZ24, MRW21, NKO^{+}23, QYL^{+}20, QCCW23, QZC^{+}24, SYZ^{+}23, SL20d, SWX24a, WZY20, WYCL24, WDZL24, YYF24, YTX^{+}24, ZHQ^{+}24, ZHWS24, ZZY24, ZWJ^{+}20]. Evolutionary [DJW20, DW22, GWZ^{+}20, INT20, JJL23, LXF21, MBLL24, Sch22, WLJ^{+}21, WWZW23, YXX^{+}22, AT24, BYCX24, FHFS24, GH21, HGD^{+}21, HJJL24, HZXT24, JTW^{+}23, KT21a, KAT22, KLH24, LJR20, LW22f, LHH24, LLZ24, MRW21, dDMMTH22, MCSD21, MJM20, PJ21, TGL^{+}24, Wet23, Xia21, YY24, YYL^{+}23, ZLF22, ZYW21a, ZLDY22, ZB23, ZZPF23, ZSJ23a, ZTLW24]. evolving [HXF23, YLW20]. Exact [Alj21, GB24, Jor21c, Kwo22a, LZ20a, SPM22, dlCJ20, BSS21, CHY22b, GLGLH24, KRE20, LLT22, LFA22, Ndo23, SS23b, SJ20, VSM23, Wen20, Zha20]. exactly [CCCdLMR23, GT22a, YWX20b]. examine [BJK^{+}22]. exceed [XLS21]. Exceptional [SBV23]. Exchange [AMOS23, AR21a, ME21b, WLS22, ZLH21, ZWZ^{+}22, ZBA^{+}23a, ZBA^{+}23b]. exchanger [Hoj20]. excitable [MUMS22]. excitation [RX20]. exclusion [LW20c, QYL^{+}20, ZRMD21]. exercise [HL22, LYZ22b]. exhibiting [YR20a]. Existence [BBHM21, CPS24, HY24, Kan20, LLZW22, MSE24, PS21b, SB24, YWT20, Ach21, DdAS22, EHVMM24, MW20, MT21, YSHV^{+}23, YK22]. Existential [TCG20]. existing [ZHJ^{+}21]. exit [DGL^{+}23, GYC^{+}22]. exit-option [DGL^{+}23]. exits [RYG21]. exogenous [ZYSW22]. exosystem [JWL^{+}24]. expanded [AP20]. expanding [GS22]. expansion [DP21, HC20, LWZ22a, DDM24, TM22a, WZY21, WWW24a, ZY21b]. expansions [GST22, Jor24, LSCS21, WJ20, XCZ22, ZGG^{+}22]. expansiveness [DSS22]. Expected [OG20, Bri20, TZ23a, ZYLH21]. experience [KXW24, LYX23, SZCP21]. experienced [YZZ^{+}22, YYW^{+}23]. experiment [MJ23]. experimental [MGK23]. expert [XFY20]. experts [ZHLZ21]. Explainable [LBJ22]. Explaining [LZZ21]. explains [GLL^{+}22]. explanation [LY23b]. Explicit [BMY24, DX24, GL23b, GH20b, HAFM21, RRS21, XCSW21, ABNS24a, ABNS24b, ABNS24c, CCCdLMR23, DDCM23, JH21, KKLJ20, MDT22, NI21, SK20, WHJK23, ZF24, ZYX21]. exploration [WYCL24, XLYA21]. exploratory [SJ23b]. Exploring [LLW24, SK24b, ZLZ^{+}23b, LCW^{+}23]. Exponential [AC20, CHG23, CKV21, CLH24b, FZ20b, FZS21, KNW21, LWL^{+}23a, LYXZ21, RBPN20, STA24, YWS^{+}20, ZH20b, ZPW22, BMY24, CFWZ20, CXYZ23, CLUR22, DTP21, DWZ20, GZH22, GG24, HWL24, HOM22, JY20, JHSL21, KZC23, KBL^{+}24, LX20a, LZC22b, LZZ24a, LWCD24, LCY^{+}21, LL20b, LXW^{+}20, LHFW21, LtTBPA20, LW24, MY22, MD24, NT24, ON21, QXG21, RM23b, SRK^{+}20, WW20, WHM^{+}21a, WHM^{+}21b, WHCW20, WX21a, XZeF20, XD23b, XCSW21, XZZC22, YGZ^{+}22, ZF24, ZS21b, ZYZ22c, ZSS24, ZLC20, Zha23b]. exponential-SAV [ZSS24]. Exponentially [ELOY22, BM22, MRVC20, SSC22, WXHZ23]. exponents [ES23, Kan20]. exposed [NWM23]. exposure [AO21a]. exposures [Grz22]. Express [DGO22]. expression [FCeFZ21]. Expressions [MS22b, Hon24]. expressivity [ZLS23]. expulsion [WP21a]. extend [PZS24]. Extended [GLD21, IÖÜP23, KLS20, LM20b, LTL^{+}21, NBTP23, PKL23, SAB23, WSL^{+}21, XXW^{+}20, BTL^{+}22, CWZL23, CL24b, DMP21, GLZA21, HCX^{+}24, LHZ21c, LG22, MT23, SC21, SM21c, WWW24a, wXhGZ20, wX20b, ZYCY22, ZSP24]. EXtender [TJAJAG21]. Extending [ACGG20, BS23, DVC23, GBF21]. Extension [EBGN23, GN20a, LJAM24, LW22b, MZS24, WLJ22, WWK24]. extensional [CFLL21]. Extensions [AACJ^{+}22, DH24a, LJAM24, LPL^{+}22, ZJCA20]. external [FLZ23, HQS^{+}24, RHR^{+}21, ZBL^{+}23, ZXW20]. Extinction [Liu22b, AB20b, HZJ^{+}21, KEiS23, RP20, ZHJ^{+}21]. extortion [SWX24a]. Extra [ZT24, YZM24]. extracorporeal [JSA21]. Extrapolation [PH20, BCS24, BSCV24, SSCW22]. extravasation [IMIH22]. Extremal [BNR20, CLWZ21, CZ24a, CMR20, DS22c, LLZ23, LLSY23, LY24, QZ21, SS21b, SY23, YW20, DL21a, FQ22, GG24, LWZ22b, LH24a, LPL21b, SHW21, WX20a]. Extremality [LYB^{+}21]. Extreme [BKMO22, WCP20, Yu23]. extremum [PGF20]. extremum-preserving [PGF20].

F16 [IFHA21]. facial [ZQ23]. facilitate [NCY^{+}20]. facilitates [LJP21, YSYS24]. facility [AA20, GLW24]. facing [MMS21]. factor [JHO^{+}21, QLKW20, TX21, ZSYX23]. factored [Qi22]. factorization [BD20a, DESZ23, CZMX21, Sin20a]. factorized [Du24]. factors [CDH23, GPW24, Guo22, LC22b, SSCW22, YWL20]. Fading [ZCL^{+}23, HCL^{+}23]. failed [LNC24]. failure [GSRL20, KMLW22]. failures [HLL22, KM22, LGX22, SZSZ22]. fairness [DZW21, HRZ^{+}23, QSW23]. fairness-based [QSW23]. fake [DMPP21]. fall [WDZG22]. falls [RS24]. false [FT24, KKPS23, WZXY20, XRZH22, ZY23b, ZGZ^{+}23]. Families [EFF23, BHGPC21, KZZ20, KFÖ23, LSVA20, TJIM20, WCS22]. family [CRT22, Cha22, GLKrH20, JY20, KKK22, RAB24, SK20, THBPMC22, YWY^{+}22, ZHL24]. fan [MJZ21]. fan-like [MJZ21]. fans [LLW20]. Fantope [LWL24]. Fantope-constrained [LWL24]. far [WPE20]. far-field [WPE20]. farm [CS20]. Fast [BSCV24, BZS21, CZMX21, FJJJ21, HMEDVR24, HLZ21, JHO^{+}21, LWNZ22, LCXJ22, Miy23, MV22, QL21, TO20, WXLY23, YHW^{+}23, YLS23, ZYY20, BL22, CS24a, DGP^{+}23, GdFPF24, LAH22, Mar24, Qi22, RKRRL21, SFL^{+}22, Voe23, WHL21, WHJK23, XYKL20, YH24a, ZFS22, ZGZ^{+}24, ZYySX23, ZZY24, Zhu24]. fast-slow [Zhu24]. fast-transient [DGP^{+}23]. fast-varying [WHL21]. Faster [THZ23]. fastest [YZ23]. fatality [RS24]. fatigue [YFM^{+}24]. Fault [DZL^{+}22, HZL21, KM22, LM20a, LCC^{+}22, NGS22, PCN^{+}24, WYLL22, Xu24a, ZHG22, ZLC^{+}24b, AKP23, CL22, CHW^{+}24, DWT^{+}21, GNSZ22, GXZ^{+}24, GHP24, HLW^{+}21, HJW^{+}23, KMLW22, LLZ22a, LLY^{+}21, LZDX22, LXWH22, LD23a, LZHZ24, LYL22, LZZ22, LLQW24, LJW^{+}22, MWZX21, ML22, MSZ22, MZFX20, MZYW23, PMAK22, SCLD21, SZG21, SWZS20, WXWS20, WHLL20, WXLZ21, WFYL21, WC22a, WMT22, WJDL24, XH22a, XD23a, XCLW21, YWY^{+}22, YLK^{+}21, YZM24, YL20a, ZW21b, ZGLZ22, ZTW22, ZWZX21, ZZW^{+}21c, ZWY^{+}22]. Fault-tolerance [LM20a]. Fault-tolerant [HZL21, NGS22, AKP23, CL22, GHP24, HJW^{+}23, LXWH22, LD23a, LLQW24, LJW^{+}22, MWZX21, PMAK22, SCLD21, WXWS20, WXLZ21, WFYL21, WJDL24, XH22a, XD23a, YWY^{+}22, YLK^{+}21, ZTW22, ZWZX21, ZWY^{+}22]. faults [GHP24, JZS^{+}24, KKS^{+}22a, KBL^{+}24, LLZ20b, LDX20, LNS^{+}21, MWZX21, NGS22, VSM22, WFYL21, WX21a, XH22a, XLZ23]. faulty [KKPS21, LJW^{+}22, ZZX^{+}22]. favor [SWFJ20]. favors [WCY^{+}22]. FBm [XWDK23]. FBSDEs [CGG23]. FCM [HGKD22]. FD [Oru21, Oru22]. FDI [DXZ^{+}23]. FE [Pop20, WDYS21]. fear [PGKM24, QLL24, TTMU20]. fear-induced [QLL24]. feasible [dDMMTH22, MBL20]. feature [BMMV23, CKSP20, YY21b]. features [AHS20, KA22, KA23]. February [Ano20d, Ano20n, Ano21d, Ano21p, Ano22d, Ano22p, Ano23d, Ano23p, Ano24d, Ano24p]. Feedback [WSXL22, CJ22, CLWX23, DSZ^{+}23, FLC^{+}22, GSS24, GS21, HZL21, HL23a, HKCK24, HZXT24, HMZZ20, JY21, JDZH22, JJSY22, KY22b, LLZ20b, LCLS20, LLZ21, LJ23, LZZG23, LZHZ24, LCY^{+}21, LZS^{+}21, LLLS20, LZZ^{+}23, LH24b, LLWW22, MQW23, MTRA22, MNSR22, MD22, NH22, OPL22, RFFZ21, RDL23, SK24b, SPZS20, Sug21, SYTC24, SM23, TLHH21, WYLW20, WXWS20, WZZX21, WQT^{+}24, WZXH22, XD23a, XZY^{+}23, XZZW21, XLZ^{+}24b, YWY^{+}22, YXMW22, Yi23, Yoo21, YP21, YP22, YZY22b, ZeFWW20, ZZL^{+}20a, ZZCL20, ZXC20a, ZCS^{+}22, ZLC24a, ZWNG21, ZCLZ22, ZZ20b, ZDD^{+}22]. feedback-based [DSZ^{+}23]. feedbacks [LLZ24]. Feller [DD21b]. FEM [CFP22, CU23, IP23, LW22d, LXL22, MV22]. FEMs [LWNZ22, PZS24]. fence [MT21]. fence-sitters [MT21]. Fermi [GST22, SS24c]. ferroelectrics [MVCK24]. ferrofluid [DS22b, LYN^{+}21, ME21b]. ferrofluids [NKO^{+}23]. ferrohydrodynamic [Yan23]. ferromagnetic [CH24a, SDD22]. fertility [WY24]. FFC [LSZ^{+}23]. FFT [EGMY23, LLF^{+}23, ZMNR^{+}22]. FFT-accelerated [LLF^{+}23]. FG [ASS^{+}23a, Bab22]. FGM [KKPG20]. fiber [TTCY23, WMQ21, YTZ20, ZJG21]. fibrillation [UMT22]. fictitious [ZH22a, ZDKC20]. fidelity [KG21, SMTT23]. Field [PGL^{+}23, ALL23, AM24, BP23, CH24a, CLZY21, CLLA23, DG21, GGW21a, HXF23, HWWW23, LZY20, LYN^{+}21, LTC^{+}23, LSL^{+}21, ME21b, NKO^{+}23, OBP21, PPS24, RRR19, RRR20, SRKA23, TDZ^{+}24, WHJK23, WPE20, WHM22, YWXL21, ZSC^{+}21, ZSS24, ZT22b]. fields [GYM21, LWW23, RKRRL21, SĆL^{+}21, WPE20, Yu21]. Filippov [FAEZ24, LHW21, VMCG20, ZT22b]. Filippov-type [LHW21]. filled [AAHC21, PWSE21]. filling [LPS21]. film [SS23b]. films [DGV22]. Filon [GW21a]. filter [dSAOG21, BTL^{+}22, CC22, CL24b, DZL^{+}22, GMH^{+}20, GT22b, HZX22, KJL20, LLZB24, LY22d, MPX^{+}20, MSZ22, NBTP23, SAB23, SR24, SZ24b, WXLZ21, WWW24a, YAX^{+}21, ZYCY22]. filter-based [SR24]. filtered [WZXH22, XLLH22, YCM^{+}22]. Filtering [LWL^{+}23c, CZ20, CWJF21, DL20c, FRW^{+}24, HLKC22, Ily23, KKPS23, LHC^{+}22, LWL22, LCC^{+}22, LWLA21, LX24a, MFS22, NSSA20, QZH21, SLPH22, TW21a, WPZY23, WZZ23, WCZ^{+}21, WSL^{+}21, XXL^{+}21, XWZC21, XZQX20, YGZ^{+}22, YXZ^{+}24, ZKLL23, ZCZW23, ZLP21, ZWZS22, CSCZ24]. filters [AMZ21, XLZ23]. final [GLT^{+}21]. finance [BHK^{+}22]. financial [AO21b, CSW20, Guo22, NHBJ20, PMG20, ZY24b]. Finding [HS23a, KKB23, MDJW20, AHL21, Dol21, FJJJ21, HLP24, KKK22, Kop22, RR07]. findings [BPG24]. Finite [ASKA20, BBW21, Bon22, CWB22, CJ21, CLLA22, DLS^{+}23, DL20b, DL20c, FRW^{+}24, GWZZ23, GS21, GB21, GHP24, GW24, HWD20, HZW21, HHZZ20, JSA21, KGCC20, LWZ21, LLZ21, LZHZ20, MMHD20, MFS22, PY21, PBM20, QNX^{+}23, REC^{+}23, RH20, RDL23, TDZ^{+}24, TZ23a, TPSP24, WGLS20, WZZX21, WGL^{+}21, WLD^{+}22, WP20, WB20, WZXH22, WHL22, XH22a, XZ21, XYM22, XWL^{+}22, XZQ23, YZC^{+}22, YWY^{+}22, YCM^{+}22, YYWL23, ZPSP22, ZLLL21, ZWHY20, ZJLT23, AB23, AF21, AP22, ABNS24a, ABNS24b, ABNS24c, AKH^{+}22, Alm23, AHL21, ASS^{+}23b, AT21a, BF22, BBD^{+}22, BMY24, BZS21, BRMVCD21, BDRM23, CCCdLMR23, CZLH20, CG23, CLL22, CZL20, CHY22a, CZ22c, CLLA23, CH24c, CL24b, CZXyS24, CKV21, CCE^{+}22, Chy22d, CZGX22, DG21, DDGN23a, DDGN23b, DX24, DTK24, DL21b, DDC22, DYW^{+}20, DL21c, Eft21, FJLG22, FHM22, FCeFZ21, GYM21, GP23, GSRL20, GLC23, HHGZ23, HWCD22, HLP24]. finite [HGD^{+}21, HLL22, Iza21, JZ20, JWC20, JCSH24, JPZ^{+}21, KMN22, KG22, KAS21, KPP22, Kum24, KK24a, LLF^{+}23, LD20b, LLY^{+}21, LHZ^{+}21a, LW22c, LJY22, LWZ22a, LZ24c, LGC24b, LC23, LSL^{+}21, LD21b, LGK^{+}22, LZ22e, LL23d, LTB23, dSLP20, LHFH22, LX24a, LZ22f, LZ22g, Ma21b, MH20, ME21a, MR22, MMDD21, MD22, MVO21, NE21, NZ20, Oru21, Oru22, PDK23, PGF20, PWWZ23, PT21, QPHC21, QLC^{+}24, Que22, Sek23, SZCC23, SGI20, SL20c, SL20b, SW22, SZ24a, SLZ20a, SCHS22, SLPH22, SY22b, SDW^{+}22, TLC22, TCPF24, TM20, TZ23b, TPTA24, TKF23, UGG^{+}20, Voe23, WHLL20, WSY21, WC21, WXF22, WL23, WTCC20, WSZ22, XLJL22, XTT22, XHL23, XYH^{+}24, XHWM20, XZQ^{+}22, XWL23, YLL22, YHW^{+}23, Yan23, YJ20, YXZ^{+}24, YZDC22, ZL22a, ZT22a, ZL20a, ZCH20, ZSL20, ZZQ^{+}20, ZXC20a, ZL21a, ZLS21, ZL22b, ZHQ^{+}23, ZLZ24b, ZGZ^{+}24, ZXF20, ZZW21b, ZYySX23, ZS20a]. finite [ZS22c, ZC21b, ZCZW23, ZWJ^{+}24, ZWZX21]. Finite- [GWZZ23]. Finite-/fixed-time [GWZZ23]. finite-buffer [Chy22d]. Finite-difference [Bon22, QPHC21]. finite-element [KMN22]. finite-field [CLLA23, LSL^{+}21]. Finite-frequency [RDL23, QLC^{+}24]. finite-horizon [WXF22]. Finite-region [FRW^{+}24]. Finite-time [ASKA20, CWB22, CLLA22, DLS^{+}23, DL20b, DL20c, GS21, GB21, GHP24, GW24, HWD20, HZW21, HHZZ20, JSA21, KGCC20, LWZ21, LLZ21, LZHZ20, MMHD20, MFS22, QNX^{+}23, RH20, TDZ^{+}24, TZ23a, WZZX21, WGL^{+}21, WLD^{+}22, WZXH22, WHL22, XH22a, XZ21, XWL^{+}22, XZQ23, YZC^{+}22, YWY^{+}22, YCM^{+}22, YYWL23, ZLLL21, ZWHY20, ZJLT23, JPZ^{+}21, ASS^{+}23b, CZLH20, CL24b, CZXyS24, CZGX22, DDC22, DYW^{+}20, DL21c, FJLG22, GSRL20, HLL22, JWC20, LLY^{+}21, LHZ^{+}21a, LJY22, LD21b, LL23d, LX24a, MR22, MD22, SCHS22, SLPH22, SDW^{+}22, TPTA24, XLJL22, XHL23, XZQ^{+}22, XWL23, YLL22, YXZ^{+}24, YZDC22, ZZQ^{+}20, ZLS21, ZC21b, ZCZW23, ZWZX21]. finite-volume [TKF23]. finite/fixed [XYH^{+}24]. finite/fixed-time [XYH^{+}24]. fins [KY21]. FIR [GJS^{+}23]. fire [LPS23]. firearm [Cro22]. firebringent [WMQ21]. fireworks [xHhPjW^{+}21]. firing [LSYX21]. firms [Ask22]. First [Kud20a, YWXL21, dJCGVM24, CZ22c, DD21b, GWZZ23, GLL23, GN22, JT20, Liu22a, LW21b, MC20, MRVC20, DDM24, QPHC21, QYD22, SGX21, Sug21, TCA23, TO20, Wu20a, XCZ22, YD22a, ZPSP22, ZWW21a, ZWW21b]. first-arrival [QPHC21]. first-order [GWZZ23, GLL23, SGX21, Sug21, ZWW21a, ZWW21b]. first-ordered [TCA23]. first-passage [dJCGVM24]. first-passage-time [DD21b, GN22]. fish [CPS24, CLZ24]. fish-small [CLZ24]. Fisher [AF24, DX24, PZS24, ZZY20]. fishery [BBAS21]. fishing [JC22b]. fission [YTZ20]. fission/fusion [YTZ20]. fistula [NS20]. fit [KNW21]. Fitness [HDQP23, LJN^{+}22]. Fitness-based [HDQP23]. fitted [AT21a, ELOY22, MRVC20]. fitting [KWW22, MMM24, ZH20b, ZZZ24]. fittings [WL24b]. FitzHugh [HGJ21, LX22, SY21a, ZS20b, AFGG20]. FitzHugh-Nagumo [ZS20b]. Fixation [LLZ24]. Fixed [AAJ23, BZZ21, CDZW21, CWWZ24, CXY^{+}23, GHS^{+}21, JWL^{+}24, LSWL20, LNC24, LZH21, WZW23, WM24a, XLZ^{+}24a, ZWH23, BSS21, DDE21, DZZ^{+}21, DS22c, DS24c, DDGBGS^{+}24, GP23, GZRM21, HQ22, KY22a, LX23a, LSL^{+}21, LNZ^{+}22, LLCL23, LWY^{+}24, NS22, OMS22, RDL23, RY24, SMH^{+}20, SLZ^{+}20b, SD22b, TlS^{+}22, THBPMC22, WA22, XLLH22, YY20a, ZC23a, ZJW^{+}22]. fixed-free [BSS21]. fixed-grid [NS22]. fixed-moments [LX23a]. fixed-order [RDL23]. fixed-point [THBPMC22]. fixed-supported [BSS21]. Fixed-time [BZZ21, CDZW21, CXY^{+}23, GHS^{+}21, JWL^{+}24, LSWL20, LNC24, LZH21, WZW23, WM24a, XLZ^{+}24a, ZWH23, DZZ^{+}21, GWZZ23, LNZ^{+}22, LWY^{+}24, SMH^{+}20, SLZ^{+}20b, TlS^{+}22, WA22, XYH^{+}24, XLLH22, YY20a, ZJW^{+}22]. fixed-wing [DDGBGS^{+}24]. Fixed/preassigned [CWWZ24]. Fixed/preassigned-time [CWWZ24]. Fixed/predefined [AAJ23, ZC23a]. Fixed/predefined-time [AAJ23, ZC23a]. flag [HLF22, TCZ22]. flag-transitive [TCZ22]. flat [KRS20]. flexibility [PV24]. Flexible [GDB21, LPL^{+}22, AKY21, CdR22a, HH20, XCLW21, ZLC20, ZTW22]. flexural [JA23]. flexural-torsional [JA23]. flipped [DZX23]. Flips [ZQ23]. floating [SSCG20]. flood [MGK23]. Flory [JLFL20]. Flow [FF21b, WLZ21, AS22, AOSO20, ATDJ22, Azi21, AAHC21, BH22, BPRR22, CCH24, CKWY23, CSG^{+}22, CSG^{+}23, Chy22d, FWM^{+}22, GAR22, GWH24, HH20, Has20, JRP22, JNY^{+}21, JT24, KRS20, KHA22, KA22, KA23, KTEM22, Kri22, KNEM23, LSL^{+}20, LZLM20, LTL^{+}20, LZY20, LHZ21d, LGK^{+}22, LCXJ22, MM22a, MM24, MMDD21, MJK22, MMUS22, ME21b, MVO21, MS22c, NHK^{+}22, NHK^{+}23, PPS24, Pap21, PR20, PM20, QLG^{+}22, RRR19, RRR20, RMSP22, RM20, SN21, SIK22, SKK^{+}21, SSR23, SWX24b, VBP22, WGW^{+}21, WWQ22, WHD^{+}23, XYCG21, XLG^{+}21, YZAM20, YFHC21, YXXZ23, Yu21, ZH22a, ZL23a, ZyZW21]. flow-guided [YZAM20]. flow-particle [ZH22a]. flows [AR21b, BT22, BRMVCD21, CGLR23, CF23b, CF23c, DG20a, EBGN23, GCGM21, HL23b, KRRS22, MCT^{+}22, MGK23, MUOF22, SBB^{+}22, SSJ22, TC20a, TKF23, YOF21, ZS20a, ZWJ^{+}24]. flowshop [AAA20]. fluctuation [AP21]. fluctuations [HR23, LNS^{+}21, TWZ^{+}20]. fluid [ABD23, Agg22, AR21b, AKH^{+}22, Azi21, BH22, BJK^{+}22, CJ21, DGP^{+}23, DM23, EFJO20, GAR22, Has20, HZZ^{+}21, HL23b, KA22, KA23, KNEM23, LSL^{+}20, LZLM20, LTL^{+}20, LQSZ21, LCXJ22, Ma21a, MYC^{+}23, MMDD21, MMUS22, MMS21, NHK^{+}22, NHK^{+}23, NA21, PPS24, PJMW22, PGR23, RRR19, RRR20, RC21, RM20, SBT22, SBB^{+}22, SKL23, TM22b, UZ20, VBP22, VGR^{+}21, ZCS21]. fluid-structure [DGP^{+}23, HZZ^{+}21]. fluids [ABD22, ATDJ22, Eva20, Fab20, GAR22, KSPC22, KPPK23, TYCK23]. Flux [CHK21, CSG^{+}22, CSG^{+}23, GB20, LNO23, SZ23]. fluxes [QLG^{+}22]. Fock [VPK20]. focusing [SC20c]. Fokker [AF24, LM23, UGG^{+}20]. fold [LWL^{+}20]. folded [CHL24, Xu24a, YZM24]. follower [CLC^{+}22, GZCS23, LNC24]. followers [WY23]. Following [WDTF22, BRTR21, GL22, LHX^{+}22, LNC24, SWS^{+}22a, TLHH21, WZW23, ZLZ^{+}21]. following/containment [CDZW21]. forbidden [GPHHC23, GPHC24, TNL23, ZQ23]. Force [GRMQ22, DZL^{+}24, FQHS24, WWT24, ZZL^{+}24]. Force-Directed [GRMQ22]. force/position [WWT24]. forced [JGXL23, KSPC22, KAT22, KWW22, ZLZ20]. Forchheimer [Pap21]. forcing [SZZ22, ZZ22b]. forecast [CQ24]. forecasting [AMBG^{+}20, CJC20, Ema21, STH24]. Foreign [AMOS23]. foreseeable [CWJF21]. forest [HGKD22]. forests [CLZ23, LY21a, TT20, TT23, WHF20, WWWK24]. form [ABD22, Fre22, GS22, Hon24, HM20b, JY21, LPP21a, LPP21b, LLW24, MKP20, NI21, SMR22, SP23, YLL22, YLC20b]. Formation [XZ23b, CG21, CWZL23, DXS^{+}20, DYX24, GS24, HCBY21, KY22b, LLT22, LZWC22, LTL^{+}22, LDXZ23, LGS22, LWX^{+}22, MTFD24, MVCK24, PGKM24, RSZA23, SMTT23, WHH^{+}20, WGZZ22, WJL^{+}24, WC22b, XXJtC24, XLZ^{+}24b, YY20a, YA21a, YP22]. formations [YMRJ22]. forming [LWZW20, Ued23]. forms [AFGG20, LD20c]. formula [KH23, LX20b, RVV24, SJ20, SMR22, WZ20b]. formulae [BT21, GH20b, KX23, OPS20, ZJX22b]. Formulas [TCA23, AOSO20, Cha22, DHS23, DF20, MH20]. formulation [CU23, CCE^{+}22, PAY22, QZ22, SMW^{+}23, SLHX21, YLH20]. formulations [ASS^{+}23a]. Forward [RT20a, BPR21, BK21, LLX21, MMS21, ZXZ24]. forward-backward [BK21]. forward-facing [MMS21]. foundation [Bab22, Yu23]. foundations [SSC22]. four [DWW24, GT22a, HJL^{+}23, WBV24, YSL24]. four-species [WBV24]. Fourier [De 22, JZ20, KWZX23, KX23, KGPS22, KS21, Lia21, SS23c, SV21a, WZY21, WL23, vdMvRS21]. Fourier-type [KX23]. fourth [Baz20, DYA21, KM20, LZ24c, LLZW22, MM20b, MSB22, NP20, PZS24, QXG21, RC22, RG20, SG21, WS21, YWZ23, Zha21a, ZSFS22]. fourth-Order [DYA21, Baz20, LZ24c, LLZW22, MM20b, MSB22, QXG21, RC22, YWZ23, Zha21a, ZSFS22]. Fractal [HVZZ23, AKP23, GTSV23, GB24, RY24, WLWL22, Wu20b, YL24a]. Fractional [DC21, KSS23, LLH21, MGJ21, TZ21, UMT22, WHM^{+}21a, WHM^{+}21b, Wan23b, XWM23, AD21a, AKMR21, ANS^{+}20, AH21, Alj21, AS21b, AARS23, ASAK21, ATDJ22, ADB20, AS21d, BL22, BLO22, BSS21, BGRM21, BZS21, BMJ20, BCVV22, CS24a, CZLW24, CRT22, CCdlH21, CYHR22, CLC^{+}22, CM22, CL24a, DEHFF24, DDLW21, DZZ23, DL20b, DL21c, ER22, ES24, EBJ20, ED20, FF21a, FZB24, FRS22, GC24, GH20a, GLT^{+}21, GLTY22, GWY20, GZG^{+}21, GL22, HD23, HHGZ23, HLZ20, HP23, HW24, HSM21, HD22a, HXZ21, HZ23a, HHZZ20, HSZ20, HLM23, HAFM21, JYMP^{+}20, JLJ23, JWC20, JHO^{+}21, JCH^{+}20, JL22, JCSH24, JSZ20, Jor23, Jor24, JZ22, KR22, KKLJ20, KIB20, KMD21, LSL^{+}20, LZH^{+}20, LKSS21, LHZ^{+}21a, LCXM21, LMX22, LZSC22, LC22a, LWL23b, LCLH24, Lia21, LM20c, LFLP21, LZL^{+}21, LCXJ22, LY23c, LXK23, LLS21b, ML20, MS20, MNZ20, MYY21, MMS21, MJW21, MJ23]. fractional [NP20, NH22, NBTP23, OPP22, PSXX21, PWZ22, QYLW24, QL21, QSL21, RLX22, RD24, RRS21, RG20, SM21b, SZCC23, SZ24a, SLS24, SJTCdlFL21, SPZS20, dOSR21, SKR22, SLTL20, SLZ20c, SFL^{+}22, SHYS23, SB24, TWB22, TZ24, TA21, THBPMC22, URLC21, UHZ^{+}20, VJA^{+}22, VPK20, VGdFCBA20, VHB23, WW20, WJC^{+}20, WLW20, WGLS20, WL21b, WXP21, WZGL21, WLB22, WLWL22, Wen20, WB20, Wu20b, WGLL20, WZ20b, XBLW23, XbYW20, XYCG21, XYM22, XZZ20, XLYA21, XYLF21, XYR^{+}23, YZLS20, YWZT20, YCEW21, YHW^{+}23, YLS23, YWS24, YH24a, YHD24, YW22b, YYF24, YLL20, YZ21, YTLV22, YMYR21, YW21, YI22, ZT22a, ZLX20, ZZ20a, ZSC^{+}21, ZCZ22, ZLK^{+}22, ZL22d, ZSFS22, ZFS22, ZF23a, ZrLkL^{+}23, ZJJ24, ZLZ24b, ZLZ24a, ZGZ^{+}24, ZZX20b, ZZX21, ZJX22b, ZZX23, ZLZ^{+}23a, ZLJ20, ZWFW22, ZW22b, ZQC24, ZL21c, ZLS23, ZHZZ24, ZC23b, ZLU21]. Fractional-order [WHM^{+}21a, WHM^{+}21b, ANS^{+}20, CYHR22, CLC^{+}22, CL24a, DZZ23, GWY20, GL22, HHZZ20, HSZ20, JYMP^{+}20, JWC20, KR22, LZH^{+}20, LKSS21, LHZ^{+}21a, LCXM21, MS20, NH22, QYLW24, SZCC23, SJTCdlFL21, SPZS20, SKR22, SLTL20, SHYS23, TWB22, URLC21, UHZ^{+}20, WB20, XLYA21, XYLF21, XYR^{+}23, YWS24, YZ21, YI22, ZLX20, ZLK^{+}22, ZrLkL^{+}23, ZLZ24a, ZLZ^{+}23a, ZWFW22, ZW22b, ZL21c]. fractional-orders [HD22a]. fracture [YXXZ23]. fractured [RM20]. fragile [CLC^{+}22, JWC20, KKPS21, LL22b, LR24, LLWK23, NSSA20, PN24, SSW21, VSM22, WLZF21, WM23, XZeF20, XXW^{+}20, ZZQ^{+}20, ZC21b]. fragmentation [PDK23]. frame [MCD21, WHJ^{+}24]. frames [DHM23, TZLL24]. Framework [AMOS23, ABD23, AMBG^{+}20, BSST20, HR20a, HOM22, KA22, KA23, KLS20, KKH23, LBX^{+}23, PKJB22, SC21, WLPZ22, YY20a, YL23a, YA21a, ZL22d, ZPW22]. Frank [DMRR22]. Fredholm [ARST21a, AHB20, BG20a, CKN21, FRRS24, KK24b, MWDD24, QYD22, SL24, Wan22a, XZ23a]. Free [Bab22, BSS21, CLW21, CLWZ21, CH24b, CFL^{+}20, CCHZ23, CY24, CSK23, DTK24, DSY22, DDH^{+}24, Eft21, GLT^{+}21, HLLW21, JRP22, JL20, JMC23, KZS20, LKP22, LOG20, LJL21, LLLC23, MZX22, NE21, PHS^{+}20, SPX24, SLL24, SLHX21, SWX24b, TLD22, WS20, WLQX20, WJY21, WXF22, WS23b, WWT24, XT23, XH22b, XH23, YFHC21, YXXZ23, YLY24, YTX^{+}24, ZL20c, Zha21b, ZPY22, ZCJ22]. free-stream [JRP22]. free-weighting-matrices [CH24b]. freedom [FZ22]. freedoms [AKY21]. frequencies [CSCZ24]. Frequency [LQ24, ZCH20, CCC23, LHZ^{+}23, LLQW24, Oru21, Oru22, QLC^{+}24, RDL23, WHLL20, WCM^{+}21, ZLZ24b, ZLC20]. Frequency-dependent [ZCH20]. Fresnel [LSVA20]. friction [CMMN20, DSY22]. Frobenius [FCX21]. fronts [AZD^{+}24]. frozen [Bon22, Mit24, SSCG20]. FTC [CXGL22]. Fulkerson [LHZ21d]. Full [MXY^{+}20, RD22, HRLH21, LWH22, LFM^{+}22, MXF^{+}24, PX23, Shi21, ZZX^{+}22]. Full-State [MXY^{+}20, PX23]. full-state-constrained [LWH22]. fullerene [DTDFT22]. Fullerenes [GARD24, SZZ22].

Fully [CYL^{+}22, LZWC22, YLC20a, ZJX22a, CGG23, CFP22, DG21, GYC24, MVO21, NWM23, SSF24, SC21, TLC22, TZ24, XYH^{+}24, ZCD23]. Function [HP23, ÅCL24, BA20, Boc22, BK21, CZ21a, Cha23, CHG23, CNYC23, CH20b, CA20, CWJF21, EB22, GSWL23, GW21a, HZJ^{+}21, JMM20, Kar22, KKPG20, KJL20, LJAM24, LPP21a, LPP21b, LT22, Liu22b, Lóc22, LWJ21, Ma24, MPW23, MB24, NT24, NA21, NYHC24, Oru21, Oru22, PWSE21, QNX^{+}23, SJZA22, WHL21, WLHW23, WY24, Xia21, ZY21a, Zha22b, ZLW23, ZF23b, ZZL^{+}24, ZLZ20]. function-based [GW21a]. Functional [CHZ^{+}20, LZW^{+}20a, AKR21, AIAFvH21, BM24, BMS21, CPS24, DDE21, DDR21, DFG22, ESA20, FLWL21, GH20b, HKKL22, Har24, JLP23, KY22a, LPKC22, LHFW21, LZ24e, WHCW20, WHY22, YL23a, YWG24, ZHL24]. functional-differential [BMS21]. functionally [DMNV20, JGXL23, KK24a, LD21a, LWQW22, MZX22, NTLNX24, PS21a, TNTL21]. functionals [LSPK20]. functions [ABTR21, AzR22, AHB20, ADLU23, BBM21, CLL22, DO21, FF21a, GAA^{+}24, GN20b, GKK^{+}24, HA20, JL22, JOR^{+}21a, KDB^{+}23, KFÖ23, LK21a, LQLF22, LW20a, Liu24, LSVA20, MMM24, MP23, Nak23, NK21b, NK22, OPS20, RRS21, RY24, Rom21, SO21a, SSA^{+}23, SKR22, TC20b, TO20, TS22, VGdFCBA20, WJL^{+}24, YLC20b, YLCL23, ZYZ22b, ZHZZ24]. functions-orientated [LQLF22]. Fundamental [ACS24, Rom21, SFC23, TTCY23]. fungal [EKKE20]. Funnel [YXYZ23, JTC^{+}24, LCD^{+}20]. Further [AC20, CLZ^{+}22, FSS20, LOVHC23, WZL22, XWMC20, XLYA21, dOS20]. fusion [GTWL21, QZWY20, YTZ20]. fuzzification [VJKZ21]. Fuzzy [FT24, GLP20, SZSZ22, WHLL20, WSXL22, XWDK23, ZY23b, ANS^{+}20, CZWS20, CSS^{+}22, CJ22, CFWL24, CZ24b, DH23, DL20c, GZQ^{+}20, GLWL22, GWZ^{+}20, HKL22a, HL23a, HCL^{+}23, HY24, HMZZ20, IKÖ24, KMS^{+}23, KKS^{+}22a, KJL20, LPK22, LR24, LM24, LR23, LX24a, LLT^{+}20, MCD22, MZSW21, NK22, NH22, NSSA20, PHL23, SAB23, SKLS22, SJ23a, SPZS20, SYTC24, SWZS20, SZWL21, SRZ22, VSM22, WZC^{+}22, XWMC20, XYZY20, XXW^{+}20, XZQ^{+}22, XWL23, Yan24, YL20a, YXZ^{+}24, Yi23, ZZL^{+}20a, ZJZD20, ZQP^{+}22, ZC23a, ZWY^{+}24, ZWZS22, ZJLT23]. fuzzy-based [DH23]. fuzzy-basis [NK22]. fuzzy-power [KMS^{+}23]. FV [LtTBPA20, QL21]. FXTZNN [MZL24].

G [GL23b]. GA [Zha23b]. GA-predation [Zha23b]. Gaddum [FKK24, HWLL21, HLX20]. gain [CSL24, JW21, LNS^{+}21, LXW^{+}20, TWB23, ZLJ^{+}21]. gains [YZDC22]. Galerkin [ML20, AD21a, AACJ^{+}22, BS22a, BBH21, CG20a, CG21, CWZZ22, DG21, DYT21, DSY22, DL21b, EFF23, FWBY23, FY23, HLZ20, KAS21, KPP22, Kum24, LLQG21, LWNZ22, LLLC23, LGC24b, LMÖ23, Luo20, LQNA23, MZX22, MHAT21, MM20a, PBM20, PWWZ23, QS20, QZ22, RD22, SW22, SLZ20c, SWX24b, TZ23b, WGLS20, WL24a, WHWW21, XTT22, XZZ20, ZMS23a, ZL20c, ZL22b, ZZX20b, ZZX21, ZLW21a, ZZX23]. gallery [AAL^{+}21]. Game [DJW20, JLL^{+}20, LLCS23, SXS22, Wet21a, Ask21, Ask22, BBEM24, BYCX24, BY23a, CT24, CWY^{+}21, CCC20b, CHP23, DZW21, DZ21, DS21c, FFL^{+}23b, FHFS24, FC21, FH24, GPH21b, GWZ^{+}20, GH21, GLLZ24, HZX21, HC24, HGD^{+}21, HQ22, HZJL24, HWWW23, INT20, JJL23, JTW^{+}23, KT21a, Kab24, KLH24, LXF20, LXF21, LJN^{+}22, LSHH22, LWZW20, LJR20, LL23e, LZD23, LW20c, LS22b, MQW23, MBLL24, MCSD21, MJM20, PGYW23, QZWY20, QDW22, QCCW23, SCT^{+}20, SLK^{+}23, Ued23, WPH20, WWY^{+}20, WJG^{+}20, WBY^{+}21, WCW21, WLZW21, WCXS22, WS22, WDW^{+}22, WY23, WYCL24, WFKL24, Wet21b, Xia21, XL24b, XDS^{+}20, YZY22a, YSYS24, YZPL24, YSW^{+}21, YZZ^{+}22, YYW^{+}23, YYL^{+}23, YHSZ20, YM22b, ZCH20, ZAD21, ZZPF23, ZSJ23a, ZHQ^{+}24, ZTLW24, ZRMD21, ZS22d, ZSZL20, ZWJ^{+}20, ZHZH24]. Game-based [Wet21a, WFKL24]. Game-theoretic [SXS22]. Game-theoretical [LLCS23]. games [AT24, AM24, CFF20, CWY^{+}24b, FLHH23, GLW24, HWY^{+}22, HY24, HJJL24, HRZ^{+}23, IÖÜP23, IÖÜP24, IKÖ24, KTZ24, KLP22, LLW24, LLZ24, MRW21, MZLS22, PZZF24, PZ20, QYL^{+}20, QSW23, QZC^{+}24, QCW24, Sek23, SWZZ23, SL20d, SJ23b, TG20, TGL^{+}24, TLY24, WLJ^{+}21, WCY^{+}22, WS23a, Wet23, WTCC20, XTS^{+}22, YY24, ZCJL21, ZLDY22, ZB23, ZH23a, ZRZL24]. gamma [DDM24]. gamma-based [DDM24]. gap [AS22, WWZ20a]. GARCH [Zha24a]. Gardner [ÖY21, WZY20]. gas [ABNS24a, ABNS24b, ABNS24c, GG22, THW^{+}21, YS23]. Gauss [CFA^{+}23, DS21b, KKS22b, MKK23, OPS20, WY24]. Gauss-cos [WY24]. Gaussian [AS23b, CLM20, ER22, GTK^{+}22, Ily23, LHET23, LZ22c, QYD22, RHN^{+}24, RJM23, WW22b, ZG24]. Gautschi [CM20a]. gBBMB [HA20]. Gegenbauer [HD22a, UHZ^{+}20, YSL24]. Gelfand [SX23b]. gene [KR22]. General [Che23, FFL23a, SO21a, XZZ20, ABD23, BIJ22, CKW22, CSQT21, DLL^{+}23, DRS20a, DDGN23b, DHS23, DW20, FLC^{+}22, GKK^{+}24, HZJ^{+}21, KK23a, KNN20, KR21, Kud20a, LPKC22, LHW21, LY24, LZ20b, LW24, PGF20, SV21b, TX21, TLD22, WCXS22, ZL20c, ZM20, ZYLH21, ZF23b]. Generalised [PSSU21, ÅCL24, MP20, SS24c]. generalization [BGHX22, GC20, GDM^{+}20, HJN^{+}24, LPP24, TCA23, UMT22, WC22c].

Generalizations [SM21a, SMW22]. Generalized [AM20a, Azi21, BFGGM24, BCS24, CSW20, CY21a, DVS21, Jor24, KT20, KKH23, Liu21, TA21, WNG20, WLL^{+}22, ZLA^{+}21, AA24, AYY24, BHL23, CLC24, CCE^{+}22, CJ21, CSQT21, DS22a, DE21, DKLV23, DFX24, EBJ20, FX20, FRR^{+}22, GLW20, GA20, GLZB20, HA20, HBWQ20, HWXC20, HM20b, HB24, Ili22, JMM20, KKK22, LKP22, LM20b, LLW20, LQL21, LWNZ22, LW23b, qLgZ21, LLZ22b, LSVA20, LLT^{+}20, LM20d, MSE24, NHBJ20, NK21a, NXL21, OK21, PMG20, PN24, RVV24, RBPN20, SNRA20, SWS21, SLZ20a, SĆL^{+}21, SMR22, UGG^{+}20, WHL21, WM24b, Wen20, WW22b, WX21b, XFK22, XKW^{+}23, XYR^{+}23, XYlW^{+}24, YM21, ZYC^{+}24, ZSL20, ZZS^{+}20, ZJWV24, ZJX22b, ZDX21, ZMS23b]. generally [QYG^{+}20, TW21a]. generated [HLM23]. Generating [BRFLGH21, BDPV22, ÖC24, Tre23]. Generation [TS22, FWM^{+}22, Has20, SN21, SIK22, SZWL21]. generative [ZC23b]. generator [Ahm21, ZH22b]. generators [BKMO22]. generic [HY24]. genetic [RST21, SXS22]. genetic-priced [SXS22]. Geodesic [MBK23, MBK22]. geodesic-transversal [MBK22]. Geodetic [BLOP23]. Geometric [DS21d, FJV21, HMR23, VSM23, XZ23b, CDBR^{+}22, Ema21, HBWQ20, LCLH24, LPL21a, ÖC24, VPV^{+}21, ZSJ23b]. geometric-arithmetic [LPL21a]. Geometrically [DKLV23, ABPP23, ZMD^{+}21]. geometries [THW^{+}21]. geometry [LSC^{+}21, LPP21a, LPP21b]. geometry-based [LPP21a, LPP21b]. Gerber [PRP24]. Germain [MP20]. get [dSTV21]. ghost [HL23b, THW^{+}21, UZ20]. ghost-cell [THW^{+}21]. Gibbon [ZLA^{+}21]. Ginzburg [Kud20a, RD24, SL20b, ZXB20]. Girsanov [Zha21c]. girth [GZZ21, ZCW24b]. give [XLL^{+}21b]. given [BNR20, COPW23, DRS20a, DM21, DM24, DL21a, DD20b, EW21, FLL^{+}22, GGS21, JL21, LYB^{+}21, LWZ22b, LLSY23, LY24, LD23b, LRS20, LCWZ24, SM24, YD22a, YDXH22]. glasses [AHS20]. GLce [SZCP21]. glioma [CS21b]. Global [CTZ21, EHA20, EA21, GM21b, HM20b, JDLZ21, JDZH22, KZC23, LZH^{+}20, LHW21, LKSS21, LLY^{+}22b, LCY^{+}21, LW21b, Ma23, SSI23, SLZ^{+}20b, UKA23, XYLF21, XZZC22, YGHW21, YK22, ZLK^{+}22, ZLC20, AGA20, CWZ20, CWF^{+}21, CXYZ23, DWZ20, ES23, FZG23, HWL24, LH20b, LHQ^{+}23, LZZ22, LHH24, NNN24, PWSE21, SZCP21, TCPF24, WLG^{+}21, WWWZ21, YXX^{+}22, ZYZ22c, ZLJ20, dSTV21]. globally [LNC22, LCL22c]. GLSPIA [ZZZ24]. glucose [AS21a]. glucose-insulin [AS21a]. glycol [AGC20]. glycolysis [LGS22]. GMRES [SSCW22, Zou23]. GNN [CLH24a]. go [CS21b, WHD^{+}23]. go-or-grow [CS21b]. goal [Wet21a]. Godunov [GG22, ZL23a]. Godunov-type [GG22, ZL23a]. Gompertz [EB22]. gon [HPL21]. good [KM20, LZ20c, XL24b, Zha22b, ZPZH23, ZLC^{+}24b]. goods [BBEM24, CWY^{+}24b, CCC20b, FH24, HWY^{+}22, HRZ^{+}23, JTW^{+}23, LW20c, LS22b, MQW23, PZZF24, QZWY20, QYL^{+}20, QDW22, QCCW23, QZC^{+}24, SLK^{+}23, WPH20, WCW21, WCXS22, WCY^{+}22, WS23a, WYCL24, Xia21, XL24b, YZY22a, ZAD21, ZB23, ZRZL24, ZRMD21, ZHZH24]. Gordon [JZ20, LZ24c, LWCD24, SW22, WL23, Alm23, CL23, NNB21, SM21b, VPK20, YL22]. governed [HLLY20, SWF20, SV23b]. GPU [AFG21, GM21a]. grad [Agg22, AR21b, BS22b]. grad-div [AR21b]. grade [LSL^{+}20, NHK^{+}22, NHK^{+}23]. graded [BM22, CFPM23, DMNV20, JGXL23, KK24a, LD21a, LWQW22, MZX22, NTLNX24, PS21a, SSC22, TNTL21, XYM22].

Gradient [BD20b, DP24, Ten22, ASS^{+}23a, BP23, CPMS22, CYZ23, GLP22, GWH24, HD23, Hua24, JA23, JW23, KET22, LAS23, LZL^{+}21, MCB^{+}21, NL20, NTLNX24, RHN^{+}24, SSI23, SPM22, TNTL21, You20, ZC21a, ZZW20, dSTV21]. gradient-based [CPMS22, JA23, NTLNX24]. gradient-type [Hua24]. gradients [UBSU21]. gradually [YWL20]. Graham [AHL21]. Gramian [QLC^{+}24]. Gramians [LJY22]. granular [LYhH^{+}24]. Graph [GMT20, GXZZ20, LLZ^{+}20a, LG23, TT20, Che21, Che23, Che22, COPW23, DESM^{+}20, DNR21, DBS^{+}23, FKK24, GH20a, HREB21, HPL21, HXL^{+}20, HCS23, JMY22, JWMW23, KFR20, LLL^{+}23a, LLZ20c, LXLP20, LZC22a, LR22, LW23b, LY23b, LLW24, PPS20, RSR21, SS24b, SLJ23, SM24, WLJ22, Wan22c, WMT22, XW23, XZ24, YZW20, ZTYZ23, ZHW22, ACBLC24]. Graph-based [LLZ^{+}20a]. Graph-Supported [ACBLC24]. Graph-theoretical [GMT20]. graphene [NTLNX24]. Graphical [HS23a, WF23b]. Graphs [DDS24, GV23, HM20a, KMT^{+}21, LM20a, MZ23b, TCG20, ZL23c, AA21a, AÁC^{+}21, ACSVT22, AYS22, AV21a, AV21b, AG20, BBB24, BBK^{+}24, BLOP23, BBT21, BDH^{+}20, BD21, BZ23b, BDK23, BY24, BTP20, BDTP23, BAH^{+}21, CLCD23, CLW21, CLWZ21, CGTU20, CDLR24, CS21a, CRDD21, CLL23, CFR^{+}20, CFL^{+}20, CLL21, CY21b, CDH23, CHJW23, CW23, CLZ23, CCHZ23, CZ24a, CY24, CFLL21, CDD23, CSK23, CGR21, DVC23, DSAY24, DRS20a, DM21, DM24, Das24, DESTP22, DYZC20, DFP^{+}24, Di 22, DHZ23, DDH^{+}24, DS20, DQC^{+}21, DTDFT22, DGMP22, DS22c, FMT^{+}21, FLL^{+}20, FLS20, GRD23, GW20a, GLW20, GM24, GKT^{+}23, GV24, GNZ20, GLMT21, GZ20c, GZZ21, HT24, HCSJ24, HZF21, HWLL21, HAK20, HK21, HPT22, HH23b, HZ20a, HZ23b, HJL^{+}23, HZ20b, HDZ22, IZDR24, JL20, JY23, JWWL20, JT22, JJC24, KF23, KCJ22, KKB23, KT20, KRT21]. graphs [KWH^{+}23, LW22a, LQW20, LLW20, LXLP20, LQL21, LW21a, LYJM21, LZRG21, LZX22a, LY22a, LW23b, LYhH^{+}24, LY24, LLL23b, LCL23, LNZ21, LPL21a, LLZ22b, LZZ22, LWW23, LD23b, LTWY24, LHL24, LPL21b, LCWZ24, MV21a, MJZ21, MMOS24, MGD21, MBK23, MW20, MS21c, NX23, PM21, PWL23, PLH22, PK20, Prz24a, Prz24b, QLZL20, QLL20, QLM21, QZ21, QLLY21, RD23, RGS23, RiTZ24, RS22b, SM20, SMQ24, SPX24, SS21a, SS22b, SY22a, SLL24, SW24, SHW21, SNC22, Son24, Spi22, Sta20, SZZ24, TLCM24, TNL23, TT24, TLD22, WZ20a, WHF20, WLZ21, WJY21, WY22, WM22, WHKT23, WF23a, WMF23, WWY24, WWWK24, WLZD21, WL22, WSC^{+}24, WP20, WZL22, WWY22, WYG23, XLL24, XT23, XD23b, XH22b, XH23, XYZZ22, Xu24b, XLS21, YD22a, YW22a, YLHZ22, YWFL23, YG24, YJL20, YF21b, YCW23, YWX20b, YML21, YLY24, YS20, ZW21b, ZCT22, ZZLK23, ZQ23, ZCM^{+}24, ZCW24b, ZSJ23b]. graphs [ZXC^{+}20b, ZDMW23, ZHFL23b, ZT24, ZZB24]. Grassmann [BFGGM24, LCL23]. gravitation [MM22a, MM24]. gravitation-induced [MM22a, MM24]. gravitational [GDFNNR22, GY22b]. gravity [AKH^{+}22, BS22c, CKWY23, TM22b]. Gray [SMTT23]. greedy [WCZ22a]. green [HWZZ23, KDB^{+}23, KK24b, KKPG20]. Gregory [ZBBK23]. grey [AR22]. Grid [AZD^{+}24, CHK21, Grz22, GK21, JLFL20, KKB23, LU24, MWDD24, NS22, PZS24, SL20c, SL20b, SW20, SZWL21, TLC22, TZ24, TT24, ZT22a]. grid-based [MWDD24]. grid-connected [SZWL21]. grids [DS22a, HLM^{+}21, SYL23]. Gronwall [YZC^{+}22]. Gross [CS24a]. Grossberg [DWZ20, YWS^{+}20, YDSW22, ZYZ22c]. Ground [HLLW21, XHWM20, YWT20].

Group [Azi21, BMDE21, LV20, MMS23, Ndo23, XFY20, ZJ23, BJK^{+}22, Che22, DHM^{+}20, GH21, JGC^{+}21, LYJM21, LZRG21, LDXZ23, LZ20a, Miz21, MS21d, MZS24, SWP^{+}22, WLJ22, YXW23, ZYPL20]. group-based [DHM^{+}20]. groups [AV21a, AV21b, GH21, HZF21, LL23c, LTWY24, LHL24, SS24e, TYL^{+}20, TJIM20, WDZG22, WF23a, XFK22, YWFL23, ZN20, ZXC^{+}20b]. grow [CS21b]. Growth [AP21, ZSZ20, ALL23, HWC23, HXF23, KH23, LZY20, LTC^{+}23, LZ20a, LL20b, MFA21, NE21, RP20]. GRP [HL23b]. GRP-based [HL23b]. guarantee [YP21]. Guaranteed [HSL^{+}24, Lóc22, CXGL22, LDX20, LZZL21, ZZ24]. Guaranteed- [Lóc22]. guarantees [DHZ^{+}20]. guesses [CB21]. guidance [GYC^{+}22]. Guide [YYW^{+}20]. guided [YZAM20]. guider [YZZ^{+}22]. guiders [YYW^{+}23]. guiding [RYG21, YXZ^{+}21]. Gummel [Que22]. gun [FR24]. gun-related [FR24]. guns [Mon20]. Guo [RO21]. gyrotactic [SKK^{+}21].

H [KHA22, Kop22, WWWZ21, SRK^{+}21]. H-naphtalenic [SRK^{+}21]. h-stability [WWWZ21]. habitat [FBT21]. Hadamard [CS24b]. Hager [SSI23]. hairtail [JKJC21]. Haldane [AKR21, ZY21a]. half [CGJ21, CZ21b, JD20, OPP22, PS22]. half-arc-transitive [CZ21b]. half-line [OPP22]. half-linear [CGJ21, JD20]. half-space [PS22]. Halin [WLZ21]. Hall [SN21]. Hamilton [HJY24, KLS20, LLWY23, ZLL20, ZXC^{+}20b]. hamiltonian [FMT^{+}21, DD24a, FMT^{+}21, KKB23, KFR20, LJY22, LW22e, LZZ22, LLZL22, MC20, PAY22, TZ21, XCSW21, YM21, YXFL23]. Hammerstein [MQ22]. Hamming [KWHS23, WWZ20a]. handed [ZZY20]. Hankel [AK21, BBM21, WC24]. haptotaxis [KAS21]. HARA [ZEA22].

Harary [FLL^{+}20]. harbour [MRR20]. hardening [CK21]. Harmonic [YH24b, GLC23, HLLW21, MYC^{+}23, SWS^{+}22a, SFC23, WGM^{+}21, XTT22]. harmonics [GAMP20]. Hartfiel [ZJCA20]. Hartwig [CIHP^{+}21]. harvest [JC22b]. harvesting [CLC20, JKJC21, LTW24, LB20, SY21b]. Hasegawa [KMN22, MN23]. hash [TJIM20, WCS22]. Hausdorff [YL24a, ZM21b]. having [HK21, Vyn23, YF21b]. Hawk [YM22b]. Hawk-Dove [YM22b]. HBV [MHAT21]. HDG [JPY23]. HDQ [MYC^{+}23]. health [NS20, ZM21a]. healthcare [PPS23]. heart [AGGPO23]. Heat [IMS23, ME21b, AOSO20, Ach21, AKH^{+}22, AR23, AGC20, ARST21b, AJIR21, Azi21, AAHC21, BSRV23, BD20b, CSG^{+}22, CSG^{+}23, GBF21, HR23, Hoj20, LC21, Lee23, LCXJ22, LQNA23, MM22b, ST24, SWW22, THW^{+}21, TMA^{+}22, Voe23, WWQ22, XLZL23, YWZ23, YWL20, ZZZ^{+}23, vdMvRS21, ZZX24]. heated [SKS21]. heating [KA22, KA23, LC21]. heatline [AAHC21]. heavy [OMS22]. Hebbian [AFL24]. Hebbian-like [AFL24]. hedge [CKL22]. hedging [DK22b, LBJ22, vdZGO21]. helices [DD20b]. helicopter [LCL22b, RLY20]. Helmholtz [ACS24, BL22, LGQ20, VGR^{+}21, WGSS22, ZC21a]. help [PGKM24]. hemivariational [MNZ20]. hepatitis [NNKH21, WYLZ24]. herd [BLO22]. hereditary [GPHHC23]. Hermite [AF21, ACH^{+}22, BCHRS23, CS24b, DDNS20, JY20, KPS22, LL23a, MZX21, MZX22]. Hermite-based [AF21]. Hermite-Hadamard [CS24b]. Hermitian [MSDR21, SWS21, TTL20, WX21b]. Hessian [THZ23]. Heston [AMOS23, LH20a, RSA24]. heteroclinic [WLY21]. Heterogeneity [SZXL24, HWC23, PPZW20, SLCS20, WS22, YZT^{+}21, YWJ^{+}20b]. Heterogeneous [CWY^{+}24b, GS24, Alf21, CZLH20, CJCL21, CGW^{+}20, CYHR22, CEPZ22, DG20a, DK22c, FBT21, GZWX23, HQS^{+}24, HJW^{+}23, LTL^{+}22, LDXZ23, LCML20, LW20c, MK21a, MMDD21, MCSD21, MPC^{+}24, NZLW22, QPHC21, RCG21, SH22, SWS^{+}22a, TTC^{+}23, XD23a, XWLL20, XWWC21, Yu23, ZDZS21, ZRMD21, ZLZ20, ZHZH24]. hexadecagonal [HLP24]. hexagonal [BHR20, CDR22b, LY23a, MY24, ZZ22b]. hexahedron [SSP24]. hexavalent [WP20]. Hidden [VSM22, LYL22, RY24, WZC^{+}22, XH22a, ZM21c]. hiding [HS23b]. Hierarchical [LTW24, ZZW21a, HJW^{+}23, KRNG24, KT20, KT21b, LNZ^{+}22, WCZ20, WF23b]. hierarchies [HCS20]. hierarchy [KL24, LWL^{+}20]. High [BT22, CET22, DYT21, Eva20, FZ23, GÖRT23, GBCP21, KWHS23, LHET23, MXF^{+}24, MD23, SW22, SMTT23, TCGR24, WGLS20, XYY23, YY20b, ZL21a, AH21, AMZ21, AO21a, AM24, CU23, CZL20, CLSN24, CGLR23, CTBB24, DEHFF24, ECS21, FHM22, FLR23, FY23, GJJC^{+}23, GYC24, HZJ^{+}21, HWCD22, HVYMS21, HGKD22, HL23b, JL22, KRRS22, LLZ21, LZSC22, LY22c, Lóc22, LTB23, LZ24e, MH22, Oru21, Oru22, PD23, PBKS23, QLG^{+}22, RG20, SGPZ22, SUN24, SKSC21, Ten22, TPTA24, TKF23, WWWZ21, WM24a, WZ20b, XCSW21, XYlW^{+}24, YY20a, YLL20, YHL^{+}20, ZWH23, ZM20, ZMD^{+}21, ZD23, ZF23a]. high-contrast [FHM22]. High-dimensional [XYY23, AO21a, AM24, HZJ^{+}21, SUN24, Ten22, XYlW^{+}24]. High-efficiency [LHET23]. High-fidelity [SMTT23]. High-Order [KWHS23, GBCP21, MXF^{+}24, TCGR24, ZL21a, AH21, CZL20, CLSN24, DEHFF24, GJJC^{+}23, HWCD22, HVYMS21, HGKD22, JL22, LLZ21, LY22c, QLG^{+}22, SKSC21, TPTA24, TKF23, WWWZ21, WM24a, XCSW21, YY20a, YLL20, YHL^{+}20, ZWH23, ZM20, ZD23]. high-precision [Lóc22, ZMD^{+}21]. high-speed [YHL^{+}20]. Higher [BEe21, DTK24, TZL24, ZO21, ZLU21, AF24, AGCGSC21, AS21d, BRTR21, DP23, DS22a, DWZ20, FGMS21, FG24, IP23, JX24, KEiS23, LLG^{+}24, Lia21, LG22, LFY^{+}24, MY22, NZLW22, PR20, PW22a, RDKR22, SR23, VN20, WMS^{+}22, WLL^{+}22, YR20a, ZJG21, ZC22]. higher-dimensional [WMS^{+}22]. Higher-Order [TZL24, BEe21, DTK24, AGCGSC21, BRTR21, DWZ20, IP23, KEiS23, LLG^{+}24, LFY^{+}24, MY22, NZLW22, PR20, PW22a, RDKR22, WLL^{+}22, ZJG21, ZC22]. higher-ordered [LG22]. higher-ordered/extended [LG22]. highest [MT23]. highest-order [MT23]. Highly [GM21a, HOM22, KM20, Kud20b, CFWL24, GW21a, Grz22, KX23, ML20, MMDD21, PM21, SZZL22, XZ23a, XGF20, YLPC22, ZL21b, ZYX21]. Hilbert [Gu20, HZXZ20, HLZ22, HGZ24, KKS22b, MGJ21, XGF20]. Hilliard [CWZZ22, CHY22a, DG21, ALL23, CLZY21, DR22, JLFL20, Lee20, LYKK21, LW22d, ZXF20]. hinders [CWY^{+}24a]. Hirota [HLZ22, TTCY23, WMQ21]. histogram [VJKZ21]. historical [JPZH23, LZY^{+}23a, QDL^{+}21, WCW21]. history [DZ21]. Hitting [YWFL23]. HIV [ADB20, CTZ21, KEiS23, LWJ21, SWCT21, VMCG20, YLH22, YI22]. HIV-1 [SWCT21, YI22]. HJB [CW22, FKS23]. HJB-POD [FKS23]. HMM [HQK21]. HMM-based [HQK21]. HNNs [SKSC21]. HOC [LLP23]. Hodgkin [BY21]. hodograph [FKVZ21, SS23a, BCHRS23]. Hohenberg [LHYL24, WYD20]. Hölder [AOSO20, MG21, Wu20b]. holes [WPE20]. Holling [AKR21, GP22]. hollow [DMNV20]. holonomic [HCBY21]. Holt [AGA20]. home [CGMFSHR21]. homogeneous [DG20a, HD22b, LT22, MH24, OMS22, PLQ^{+}22, PGR23, PS22, RDL23, SRZK22, ST22, ZLZ20]. homogenization [LLF^{+}23]. homologous [ZQ23]. homomorphism [WSC^{+}24]. homomorphisms [BBDVPV20]. Homophily [NZLW22]. Homotopy [CUA^{+}22, Yu23]. homotopy-based [Yu23]. Honey [YA21a]. hop [JTW^{+}23]. Hopf [Alf21, Alf24, AFGG20, AFGG21, CZL20, CWC20, HZ23a, JCCC20, LL20b, PGKM24, ZH20a]. Hopf-zero [AFGG20, AFGG21]. Hopfield [ACMT23, RM23a, RM23b]. horizon [HWW20, WXF22]. horizons [BMY24]. horizontal [JLJ23, KSPC22, ZZVL23, ZV20, ZV21]. Hosoya [AYKS22, GHND22]. hospital [RS24]. hospitalized [CRTVG^{+}23]. host [SWCT21, Zha22c]. hot [DWS24]. Hotelling [PLW^{+}20]. Hough [BSST20, RFFB23]. hp [FWBY23, CCC23]. hp-mixed [FWBY23]. HP-splines [CCC23]. HS [LWW23]. HS-splitting [LWW23]. HSDC [HZCY23]. HTLV [BKR22]. HTLV-I [BKR22]. Hua [MS21b]. Huber [CKSP20]. hued [CFL^{+}20, LLL23b, SLL24]. Huggins [JLFL20]. Hull [HS23a, AHL21, HLP24, LCL23]. hulls [LAH22, VSM23]. human [SWFJ20, WJG^{+}20]. humans [ID21]. hunting [ZCH20]. Huxley [BY21]. Hybrid [AKH^{+}22, BD20a, ISM^{+}23, PGL^{+}23, AB23, AP22, BRMVCD21, BDRM23, CFF20, CG20a, CYL^{+}22, CC22, CTF24, GWSC23, HLL^{+}21, HGKD22, IÖÜP24, LM20b, LZ22c, LX23a, LDXZ23, LD20d, LCL22a, QLZL22, RS22a, SO21a, SMH^{+}20, SJP22, SV23a, SPZS20, SZWL21, TW21a, THBPMC22, VN20, WHH^{+}20, WDTF22, WLL^{+}24, WSDW24, WWT24, WYS23, YLC20a, YTZ22, YXZ^{+}24, ZM20, ZCS^{+}22, ZHQ^{+}23, ZZW21b, ZHMZ21, ZLJ^{+}21, ZHJ^{+}21, ZZD22, ZL21c]. hybrid-order [CYL^{+}22].

Hybridizable [ZZX23]. hybridized [EFF23]. hydraulic [CM21a]. hydro [PKG^{+}20]. hydro-thermal [PKG^{+}20]. hydrodynamic [BS22a, PPS24, RRR19, RRR20, WM24b, ZV21]. hydrodynamics [LMZ20, LTL^{+}22, KPPK23]. hydromagnetic [Has20, KNEM23, Miz21]. hydrostatic [DL20a, EBGN23]. hydrotropism [LWX^{+}22]. Hyper [ZMJ^{+}23]. Hyper-SIR [ZMJ^{+}23]. Hyperbolic [DGV22, DBVCP23, MF23, AR24, DGP^{+}23, DMSS20, HCTS23, HC21, LL23a, MS22a, NT23, OK21, ÖC24, PGPCK21, STA24, SWSW21, TLC22, YTC23a, YF21a, ZZL22]. hyperbolical [ÖC24]. hyperchaos [YLX^{+}21]. hyperchaotic [JYMP^{+}20, LHET23, WA22]. hypercomplex [AHEGM23]. hypercomplex-variable [AHEGM23]. hypercube [CHL24, NX21, NXL21, QMS22]. hypercube-like [NX21]. hypercubes [LZ20c, WHC20b, YZM24]. hypercubic [TS22]. hypercycles [DWW24]. hyperelliptic [HLZ21]. hypergraph [GLPS22]. hypergraphs [BBDVPV20, DW20, HP20, LZ20b, ZLL20, ZHZH24]. hyperparameters [DESZ23]. hyperspectral [LHZ^{+}21b, WHHD21]. hypersurface [ZYCY22]. hypertournaments [YM22a]. hypertree [WLZD24]. hypertrees [WX20a]. hypothesis [Jor21b]. hysteresis [CZZP24, LWZ21]. HZ [MH22].

IBVP [CFR22, CFP22]. IC [PWL23].

IC-planar [PWL23]. ICU

[AFG22a, AFG22b]. idempotent [Çay20]. identical [DG22b, ZPLL20]. Identifiability [QXSW22]. Identification [BGRM21, GJS^{+}23, KSH22, KPP22, LZL^{+}21, MQ22, MC20, RNVC21, SHWW21, WQZD23, XbYW20, ZZW^{+}21c].

Identifying [BDMB20, CTV21, ZWFW22, GDM^{+}20, MGD21, SNC22]. identities [AS23b, CIHP^{+}21]. identity [WBY^{+}21]. idomatic [GH20a]. IFWENO [ZHQ^{+}23]. IGABEM [LZY^{+}23b]. II [WL21b, AKR21, BSM23, CF23a, GLWX24, LTL^{+}20, SNMM21, ZQZ21]. III [GP22, ZJG21]. ill [CPLH20]. ill-posed [CPLH20]. illiquid [ARISV20]. Image [LFLH23, PGL^{+}23, ALL23, dSAOG21, ADS23, BMMV23, CT24, CPMS22, CLCY23, CSV20, DEHFF24, DHZ^{+}20, Fen22, GZG^{+}21, HKU21, HD22a, Ily23, JC22a, JLOR22, LHET23, LMX22, LZ22c, LHZ^{+}21b, LMZ^{+}21, SMW22, jSZyW^{+}20, TEGP24, WHZ^{+}20, WGM^{+}21, WLZY21, WZN22, WSG^{+}21, WHJ^{+}24, XCGN24, ZQ24, ZWYL22, ZWZ^{+}22, dSLV20]. image-informed [ALL23]. images [DH23, UBSU21, UUS21, WHM^{+}21a, WHM^{+}21b, WZZ23]. imaginary [MP23, SW24]. imaginary-time [MP23]. imaging [dSAOG21, GN21, YTLV22]. IMEX [ADS21, BBD^{+}22]. imitation [FHFS24, HGD^{+}21, LXF20, NZ20, WWZW23]. imitation-mutation [FHFS24]. immanantal [WYG23]. immersed [GK20, GY22b, HZZ^{+}21, PBKS23]. immobile [YLL20]. immobilization [Vyn23]. immune [Alf23, BKR22, BJ21, CZZP24, EHA20, EA21, LXCZ21, NNKH21]. immunity [ID21, WLG^{+}21, XZSS21, ZHJ^{+}21]. immunization [LZW^{+}20a]. Impact [AAHC21, HWC23, HLD^{+}23, JLWX21, LDZT24, NS20, PPZW20, YFM^{+}24, CSG^{+}22, CSG^{+}23, Cro22, DKK20, GDS24, HH23a, JJL23, KMZ22, LCS24, LJHL21, LW20c, MJM20, SCD23, WYW^{+}20, WOB21, YI22, ZM21a, ZYZ22a, ZTLW24]. Impacts [HHW23, HCY21, CH22b, HCM21]. imperfect [EM20, GLGLH24, SY21b, WC23]. imperfections [MKKG22].

Implementation

[GRMQ22, LZLM20, LTL^{+}20, SKR22, AFG21, PGPCK21, THBPMC22, ZRMD21]. Implications [HWZZ23, HH24]. Implicit [CCCdLMR23, ZZX20b, ABNS24a, ABNS24b, ABNS24c, BT22, BRMVCD21, CZ22c, CCSPMV23, FZ23, FG24, GDFNNR22, GYC24, DDCM23, JHO^{+}21, LKK^{+}23, MCT^{+}22, MDT22, PBKS23, QS20, UK23, YZLS20, YWL20, Zha21a]. implicit-explicit [CCCdLMR23, DDCM23, MDT22]. Importance [GRMQ22, dRST23]. Important [Ma21b]. improve [PSSU21]. Improved [FFL23a, GTWL21, GLWL22, Kod21, KWHS23, LSPK20, LWCD24, NK22, PLH22, TJIM20, WSXL22, XLJL22, ZFJ^{+}20, ZCL^{+}23, ZJJ24, AFG21, CH20b, xHhPjW^{+}21, HGKD22, HCBY21, JMC23, KÖ21, KÖY24, LPP21a, LPP21b, LCC^{+}22, LWL21, LZX^{+}22b, NXG23, RX20, SWX24b, TWL20, WZXY20, WHPQ23, WDZL24, YL20c, ZL22a, Zha20, ZCPZ20]. Improvement [CCG^{+}22, ZZL^{+}24, KEiS23, ZJJB22]. Improvements [CIHP^{+}21, ADLU23]. improves [SX23a, UMT22]. Improving [ACS24, JOR^{+}21a, LSC^{+}21, LLZB24, MH20, VMT24]. impulse [CWB22, JSM20, Kum23, LZ22c, WCGW20, ZS21b, ZLC20]. impulses [CZ22a, FZW23, FZ20b, GWSC23, JQJY22, KZC23, KMD21, LCY^{+}21, WHH^{+}20]. Impulsive [KSS23, RSVL21, WWWH21, CXLM21, CSL24, CZXyS24, CLH24b, CD22, CLLA22, FZW23, FZLJ24, FZS21, FW24, GWZZ23, GLY24, GL22, GW24, Har22, HLL^{+}21, JS20, KZC23, KKPG20, LKSS21, LZC22b, LWL^{+}23a, LZZG23, LZZ^{+}23, PLRD21, RBPN20, SKH20, TWB22, WDTF22, WLZY21, XW22, XHL23, XYH^{+}24, XLZ^{+}24a, XLZ^{+}24b, YLLL20, YLHZ22, YSLK23, YWG24, YYWL23, YZ21, YDSW22, ZPSP22, ZZCL20, Zha22a, ZLC20, ZJC21, ZQ24, ZSZ20, ZL22e, ZD24, ZPW22]. in-host [SWCT21, Zha22c]. In-Parameter-Order [WCS22]. in-plane [AS20, KMZ22]. in-vehicle [ZSZL20]. incentives [FH24, HJJL24, WCXS22]. incidence [BZ23b, CTZ21, GM21b, LPL^{+}22, MHAT21, RP20, RHNT21, SK24a, XW22]. incident [DS24c]. incipient [GNSZ22]. inclined [Ach21, AKH^{+}22, DGV22, KRS20]. including [BA20, CRTVG^{+}23, WZXH22, XLZ23, YXXZ23]. inclusion [HLX21, ZS21a]. Inclusions [RT20a]. incoherent [TZLL24]. incommensurate [HAFM21]. incompatible [FAEZ24]. incomplete [CZMX21, FMN21, NK21a, QYG^{+}20, SRK23, XXW^{+}20, YL23a, ZM21c]. incompressible [AR21b, CK21, JLJ23, Ma23, MMDD21, SWS20, SSF24, SWW22, TCPF24, WWQ22, YK22, ZF21, ZRL^{+}24, ZS20a]. inconsistent [LWZ22d]. incorporating [AB20b, LD21a, LCML20, SWCT21]. increase [VJKZ21]. increased [ZB23]. increasing [LCS24]. Incremental [TZD22, HYT21]. indefinite [XTT22, ZL23b]. indeg [CLL21]. indemnity [CWL^{+}21]. Independence [GM24, ACVHC22, COPW23, DM21, LW21a, LOVHC23]. Independent [GH20a, ACS20, GW20a, MW20, Wet21a, XT23, YLPC22, Zha21c]. index [AD21b, BPG24, BNR20, BDH^{+}20, CH22a, CLL21, Che23, CHJW23, CZ24a, CLB21, CGR21, CRS21, DRS20a, DG22a, Das24, DYZC20, DL21a, DS22c, DS24c, FLL^{+}20, GKT^{+}23, GWW23, HCS23, JL21, LLZ23, LWZ22b, LY24, LZL20, LPL21a, LLZL22, MZ22, MMMBRS20, MD24, PM21, RRS24, SSP24, SPX20, SPX24, Sha22, Spi22, SD22b, SY23, VPV^{+}21, WL20, XZ24, YD22a, YWFL23, YJL20, YL20b, ZCM^{+}24, ZJX21]. index-1 [BPG24, ZJX21]. index-2 [ZJX21]. India [DKKJ21]. indicators [CD20b, RKJ20]. indices [BCR22, BDTP23, CM21a, CYG20, CMR20, GG24, LYB^{+}21, LD23b, RRS23, SLH20, SHW21, WZ20a, WX20a, YLH22, ZYLH21]. indirect [CS20, ES23, LXH24]. indistinguishability [ZeFWD23]. individual [BBEM24, BE22, GW20b, HSLS20, HH23a, LMT24, WCY^{+}22]. individual-based [BE22]. individuals [CJCL21, CWY^{+}24a, MGFF21]. induce [LLS^{+}22]. Induced [BY21, WW23, BKMO22, CWC20, CTWG24, Chy22c, DCCM20, EE22, FFL^{+}23b, GS22, KKH23, LXZ23, MM22a, MM24, MY22, QLL24, TM22a, WLY21, WLZW21, WWZW23, Xu24b, ZH20a, ZS20b, Zhu24]. induction [RHR^{+}21, RDKR22]. Industrial [Yua20, WGW^{+}21, WWW^{+}22]. industry [LLYQ20]. Inequalities [ES24, AOSO20, CS24b, DMRR22, HLX20, LPS21, MNZ20, PKL23]. Inequality [XWM23, AKPK21, CP20, HH22, HSY24, KLP22, LKP22, LD20a, PN24, TWR20, WJC^{+}20, WL21a, WHPQ23, WHL22, YZC^{+}22, YWS24, ZCPZ20, ZWZ^{+}21, ZJCA20, ZS22b, ZSZ20]. Inertia [WS23a, CdFP24, PBKS23]. inertial [DZZ^{+}21, FJFS21, FZS21, GX23, LGL20, LCY^{+}21, LZH21, MCD21, SSW21, WWWZ21, WZLC21, ZC23a]. inevitable [WCY^{+}22]. inexact [BPR21, Sin20b, dSLV20]. Infected [ASAZ23, WQZD23]. infection [AMA20, BBAS21, BKR22, CTZ21, CH22b, GSW^{+}20, HCY21, IG21, LWJ21, NNKH21, SWCT21, WMCZ21, XZS20, YI22, Zha22c]. infectious [GM21b, HCY21, RS23]. inference [NHBJ20]. Infinite [TCG20, AALMLM22, BMY24, CJ21, ESA20, FMP20, KZZ20, Liu24, MBMA20, SC21, XCZ22, YM22b, ZLC20, ZJC21]. infinitely [CL22]. Infinity [FGMS21]. inflation [HWW20]. Influence [CWL^{+}21, CLD^{+}24, KSZ23, LCW^{+}23, SLCS20, TTC^{+}23, BYCX24, CS21b, Fan21, HCH24, JPZH23, LHQ^{+}23, MRW21, SL20d, YZT^{+}21, YZZ^{+}22, YYW^{+}23]. influence-based [SL20d]. Influences [LW20b]. Information [ABTR21, DJW20, QZWY20, YWC24, AF24, Ask21, CL20, GZX20, GGBEZRI23, HWWW23, JPZH23, JTW^{+}23, LDZ20, LCW^{+}23, LXH24, LZZ^{+}20, LLX21, LYL22, QQWJ22, QDL^{+}21, SZXL24, Ued20, WWY21, WQT^{+}24, WZW24, XD21, XDS^{+}20, YZT^{+}21, YXZ^{+}21, YFM^{+}24, YCZ^{+}20, ZP21, ZXZ24, ZLP21, ZDD^{+}22, ZLZ^{+}23b]. information-based [GZX20, LDZ20]. information-disease [YFM^{+}24]. Information-driven [YWC24, ZLZ^{+}23b]. Informational [Est21]. informed [ALL23, CQ24, NYHC24]. inhibit [PZZF24]. inhomogeneity [DMNV20]. Inhomogeneous [Sha21, DN21, GN22, GN23, LJY22, STA24, YK22]. initial [BDPV22, CRT22, CB21, CWZZ22, CLX22, CWY^{+}24a, Far22, GTSV23, HWC23, LJY22, LL23d, NL20, SS24c, WMWS21, WWT24]. initialisation [JW23]. Initialization [JZZ^{+}22, CHH22]. initially [SSC22, Vyn23]. initiated [WMCZ21]. initiatives [CDL^{+}20]. injection [FT24, KKPS23, LY22d, MS22c, SCHS22, WZXY20, WQT^{+}24, XRZH22, Yoo21, ZY23b, ZGZ^{+}23]. Injective [ZZB24, CLCD23]. injunctive [AT24]. Inner [CCGV^{+}22, HH22]. innovation [YWJ23]. innovative [XJY24, ZHLZ21]. inpainting [CT24, LFLH23]. Input [ASS^{+}23b, DSS22, GLY24, GLP20, JW21, LPS23, YWG24, ZDX21, AS23a, ASKA20, CL22, CXY^{+}23, GSWL23, GSS24, GXZ^{+}24, GZWX23, HLW^{+}21, HL24a, HL23a, HRA^{+}20, JDLZ21, JQJY22, JPZ^{+}21, LSWL20, LZZL21, LZHZ20, MPX^{+}20, MTRA22, MMHD20, MB24, NTT21, SSKS21, SPZS20, SLPH22, WZXY20, WZZX21, WSS22, Wan22b, WS23b, WM24a, WZXH22, WSSX22, XLZ23, XCLW21, XLLH22, YSW^{+}22, YD22b, ZWH23, ZYSW22, ZXZC22, ZWY^{+}24, ZSP24, ZPDL20, ZY23b]. input-delay [LZHZ20]. Input-output [ASS^{+}23b, DSS22]. input-saturated [AS23a]. Input-to-state [GLY24, JW21, YWG24, ZDX21]. inputs [DCR^{+}20, FLZ23, HLL^{+}21, HLLY20, WL24a]. insect [FAEZ24]. inserting [NXG23]. insights [ZZH^{+}20]. inspired [CT21]. Instability [WLZY21, AS20, Bab22, BH22, Fab20, JNY^{+}21, KZS20, KNEM23, ZZY20, ZS20b, ZLU21]. instantaneous [FW24, GW24, KMD21, Kum23]. institutional [FH24, LC22b, ZAD21]. insulin [AS21a]. insurance [CWL^{+}21, Guo22, KMS^{+}20, KMS^{+}23, ON21, ZLW23]. insurances [CLD^{+}24]. Integer [CLL23, Gas23, KR21, KRT21]. integers [TITN20]. integrability [AFGG21, MP20]. Integrable [ZHYD20, HCS20, HLZ22, LWL^{+}20, LZYL21]. Integral [CUA^{+}22, CFA^{+}23, LZZL21, AAS23, AM20b, ARST21a, AJIR21, AHB20, ACH^{+}22, BPG24, BG20a, BM20, BM24, BRB24, CKN21, CN23, CM22, DMRR22, DDE21, DDR21, DVS21, DWS24, DXZ^{+}23, ESA20, ES24, EHVMM24, FJV21, FRRS24, GP23, GW21a, GST22, GN22, HZL21, HWC20, HVYMS21, IFHA21, JCH^{+}20, JSZ20, KY22a, KK24b, Kut20, LKP22, LHZ^{+}23, Liu22a, LSVA20, Mar22, MDSS20, MC22, MSE24, NNN24, NRP22, PH20, QYD22, RRS23, RS21, RSA24, SJ23a, SXC21, SYX22, Son24, TWL20, VSM^{+}21, Wan22a, WH21, WW22b, XZ23a, Xy20, YSL24, YHW^{+}23, YWS24, YTZ22, ZHL24, ZCPZ20, ZLW23, ZH20b, ZSYX23, ZSZ20, ZHZZ24]. integral-algebraic [BPG24]. integral-differential [ZLW23]. integral-sliding-mode [IFHA21]. integrals [CS24b, Hon24, KWZX23, KX23, Kud20a, LX23b, LPP24, ML20, MC20, WLB22]. integrate [LPS23]. integrated [CZW24, KET22]. Integrating [WBY^{+}21, CM20a]. integration [AALMLM22, BCP23, CZL20, EA24, GJJC^{+}23, JHO^{+}21, LX20b, MCT^{+}22, MDT22, RVV24, RX20, SS24c, TM20]. integrator [DdlCS24, DHR22, LWCD24, MMHD20, RS22a]. integrators [CET22, DD24a, HC20, KWW22, SG20b, WHWW21]. integro [AFA24, BR20, But21, DO21, DYT21, ESA20, JPY23, JS20, LHZ21c, MWDD24, MDSS20, MSN21, QXG21, SL24, TLC22, YY22]. integro-Differential [JPY23, AFA24, BR20, But21, DO21, DYT21, ESA20, JS20, LHZ21c, MWDD24, MDSS20, MSN21, QXG21, SL24, TLC22, YY22]. integrodifferential [DDH22]. intelligence [BAL^{+}24, LS22b, TGL^{+}24]. Intelligent [XZ23b, WYW^{+}20]. intensities [CTWG24, NK20, ZRSK21]. intensity [GN20b, HOM22, dFdOFdSBC20]. Interaction [SHT^{+}21, YZPL24, Agg22, CTWG24, CWY^{+}24b, DGP^{+}23, HZX21, HXZ24, HLW^{+}20, LL23e, LZD23, Oru21, Oru22, TGL^{+}24, WDYS21, ZZY^{+}21, ZH22a]. Interactions [XLL^{+}21b, AIAFvH21, AGCGSC21, BF22, HZZ^{+}21, IHF13, LMT24, LZ22a, LLS^{+}22, LFY^{+}24, MS24, NZLW22, RDKR22, Woo20, ZL20d]. interactive [MJM20, Zhu23]. intercity [YWX^{+}20a]. Interconnected [YXYZ23, CLWX23, CZWX21, CXGL22, DS24a, FT24, LPS23, MZYW23, SAB23, Yoo21, ZW22a]. interconnection [PMAK22, ZCW^{+}24a]. interconnections [LC22b]. interdependence [CYW21]. interdependent [BY23b, DZ21, GDS24, LGD^{+}21, MZ24, SGJ^{+}21, SYZ^{+}23, SL20d, WLJ^{+}21, YTGC20]. interest [LW24, Wan24b]. interface [CZ22c, CLL24, FHM22, GE21, JM23b, Kum24, LW20a, LGD^{+}21, PWWZ23, SMW^{+}23, SGI20, SLHX21, THW^{+}21, VBMB21, YCWH23]. interfaces [PS21a]. interfacial [MKKG22]. interfacing [LQSZ21]. interfering [TTMU20]. interior [DR22, EFF23, MWDD24, TlS^{+}22]. Intermediate [LZDX22]. Intermediate-estimator-based [LZDX22]. Intermittent [SZG21, CFWL24, GHP24, GXZZ20, EVB^{+}21, HB24, LWL23b, LK21b, LHFW21, LYXZ21, NGS22, QQWJ22, RJHC24, SCLD21, SWZS20, WZW23, YYWL23, ZWHY20, ZPW22]. internal [LZY^{+}23b, Pap21, SSW22, WC22b, ZZX24]. Internally [ZHW22]. interplay [Gök22, WPH20, WZW24]. Interplays [BBB24]. interpolant [LHZ21c]. interpolanting [ADLU23]. interpolants [BEIR23, FRRS24, KPS22]. interpolating [ARST21a, Vis23]. Interpolation [AKY21, AM20b, AS21c, AY22, ADS23, BCHRS23, BE20, BSCV24, CDLR24, CLL22, CWY^{+}20, CM22, CNYC23, CSV20, DDNS20, KG21, LS20, MZX22, MJ23, OT20, QY22, RY24, Rom21, SS23a, SLHX21, Sun22, WDC^{+}23, XLD23, Yun21]. interpolation-extrapolation [BSCV24]. interpolations [CLM20]. interpolators [DP21]. interpolatory [CG21]. interpretation [SLS24]. intersection [HPL21, WYW^{+}20]. Interval [DZDL22, LZZG23, MSS24, ZLS21, CHZ^{+}20, CLZ^{+}22, DZZ23, HHR21, HSZ20, HCL^{+}23, INOP20, LSC^{+}21, LePP24, LaGBT22, LWD21, LCZW23, LHFW21, LWL21, Liu24, LS22a, MMOS24, SKLS22, SRZ22, TGDPS22, WLPZ22, Wan24b, XZZ22b, XH22b, XH23, ZZS^{+}20, ZZ20a, ZSJ23a, ZL21c]. interval-valued [Wan24b]. intervention [FR24]. interventions [LX23a, SGJ^{+}21]. intraspecific [AIAFvH21, DWL^{+}24]. intrinsic [ABPP23, BFGGM24]. intrinsically [UH24]. Introduce [YY20b]. Intrusive [KYU22]. invariance [CM21a, DYA21]. Invariant [YY21b, AKDR23, AGGPO23, GS22, YTC23a]. invariant-region-preserving [YTC23a]. invariants [LXLP20, NV21]. invasion [CS21b, FBT21, SC20c]. invasive [MMR23].

Inverse [MMS23, VMZDV20, AK21, AACI21, AFGG21, AA20, AR23, BT21, Bon22, BF23, BMS21, But21, CLL21, DS23, Han20, HCTS23, Hua24, HGZ24, KKK22, KCT^{+}21, LH22, LLS21b, MGS20, MSMK21, MM20a, MT24, MZL24, Mit24, MS21d, MS22b, NTNT24, NH21, NYHC24, SRGS20, SMM^{+}24, WFHG20, WW22b, XFW^{+}21, XCGN24, Xu20, ZJWV24, ZMS23b]. Inverses [AB20a, CCGV^{+}22, CJ20, CS24c, CSQT21, DK22a, Ili22, MSK21, MZS24, SĆL^{+}21, SPM22, SMW22, WC22c, ZYW21b, ZMS23b]. inversion [AE21, CCC20a, GZRM21, LY20a, MPW23, TO20]. Inversions [MS21b]. inversive [BSG23]. Investigating [JNY^{+}21, KAT22, ZWJ^{+}20]. Investigation [BSM23, CUA^{+}22, HYS24, AD21a, ABD22, AGC20, CM21a, FTML22, GP23, LLLC23, SKS21]. Investigations [Ask22, ZLH21, LMZ20]. investment [CKW22, HWW20, LWZ22a, WPH20, ZHZH24]. investments [LW20c]. investor [LXF20, ZBFW22]. inviscid [Kri22]. Involution [WS22]. involving [AS21d, FQHS24, MNZ20, TX21, YFW21]. irredundance [KF23]. irregular [AYS20, AYS22, BS22a, BD21, BAH^{+}21, DSAY24, DS20, GW23a, KKS22b, Prz24a, Prz24b, RB21, WKM24]. irregularity [AG20, DS23]. Irreversibility [KHA22]. ISD [ZJLT23]. ISD-ADT [ZJLT23]. Islets [MD23]. Isochronous [LR21]. isogeny [THZ23]. Isogeometric [WDYS21, GKK^{+}24, NTLNX24]. isolated [SM24]. isolation [KM22]. isometries [ZYW21b]. isometry [FX20]. isothermal [AR21b, KZS20, KSPC22, LTC^{+}23, SNZC20]. isotropic [XCGN24]. issue [BKS21, DBVCP23, PHL23, Yua20]. issues [ASS24]. IT2 [KKS^{+}22a, SZWL21]. Itô [AAS23, RS21, SDW^{+}22, TZD22, WH21, YZC^{+}22]. Italian [BHP21]. item [CDL^{+}20]. iterated [Mit24, MHF23]. Iteration [AHM23, AA24, CDM21, GGY23, GZLZ24, Han20, KKK22, LGQ20, MG21, TLDL21, TWWL24, WX21b, Zen21, ZZLV23, ZZVL23, ZV20, ZV21, ZWW21a, AM24]. Iteration-Based [AHM23]. iterations [ZO21]. Iterative [CCE^{+}22, CSCZ24, ABNS24a, ABNS24b, ABNS24c, AM20b, CY21a, CHHC24, CSQT21, DK22a, Du24, EBJ20, Far22, HVYMS21, HCBY21, IÖÜP24, LH20a, LLFZ20, MGS20, OBP21, QQY^{+}23, SNMM21, SD22a, TZD22, WLQX20, WCP20, Wan21a, WHHD21, WLPZ22, WSZ20, WCZ22a, WL24b, XL24a, YM21, ZIML20, ZHL24, ZW21c, ZL23b, ZPY22, ZTW22, ZMS23b]. Iteratively [KKS22b, MG21, dSLV20]. Ito [HANP23]. IVPs [MRVC20, SK20].

Jacobi [AFGG21, DEHFF24, HJY24, LK21a, LLWY23, LYN22, ML20, RS21, WC24, XBC22b, YWZT20]. Jacobian [KFM21b, ES23, KFM21a]. janbu [ZSYX23]. January [Ano20o, Ano21e, Ano21q, Ano22e, Ano22q, Ano23e, Ano23q, Ano24e, Ano24q]. japonicus [JC22b]. Jarratt [ZO21]. Jarratt-like [ZO21]. Jaulent [CLX22]. Jeffrey [BJK^{+}22, MMUS22]. jellyfish [CT21]. jet [JRP22, KM22]. job [JP23]. jobs [DTP21]. Johnson [CLW21]. join [MW20]. joining [BSCV24]. joint [HDP^{+}21, LHQ^{+}23, WC22a, XJY24]. joint-space [XJY24]. jointly [LCJ^{+}22]. Jordan [GB21]. Josephson [WGMJ23]. Joule [KA23, KA22]. judges [YY24]. Julia [Wan23b]. July [Ano20e, Ano20p, Ano21f, Ano21r, Ano22f, Ano23f, Ano24r, Ano22r, Ano23r, Ano24f]. Jump [MWS23, XWDK23, ARMM23, ASKA20, CH24b, CZS21, CLZ^{+}22, CZ24b, CL24b, DLS^{+}23, FRW^{+}24, FLWL21, FLC^{+}22, FZLF23, FFL23a, GZQ^{+}20, GXZZ20, HKL22a, HL24a, HRLH21, HQK21, HOM22, JWK22, KGCC20, KKPS23, KHNT22, LLZ20b, LNS^{+}21, LM24, LSM23, LYL22, LZZ^{+}23, LOS21, NK20, NK21b, NK21a, PHL23, PS21b, QYG^{+}20, RFFZ21, SSKS21, SWX^{+}21, SSW21, SDW^{+}22, SRZ22, TQZ^{+}24, TMW20, TWR20, TW21a, VSM^{+}21, WCZY20, WXW^{+}20, WGL^{+}21, WZY21, WZC^{+}22, WPZY23, WM23, WC23, WLL^{+}24, WSDW24, WZZ24, WFKL24, WHCW20, XXL^{+}21, XH22a, XZeF20, XCZ22, XTS^{+}22, XWZC21, XXW^{+}20, XZQX20, XZQ^{+}22, YCZ^{+}23, YM20, ZJZD20, ZYSW22, ZCS^{+}22, ZZ24, ZM21c, ZC21b, ZeF21, ZWZS22, vdZGO21, SWX22]. jump-diffusion [LOS21, PS21b, vdZGO21]. jumping [RH20, SKSC21, WCZ^{+}21, WCX^{+}22, YLK^{+}21, ZWY^{+}24, ZLLX22]. Jumps [KOP21, KK23c, LPK22, LHC21, RM23a, XCZ22, YZC^{+}22, YM23a, ZLD22]. jumps-differential [YZC^{+}22]. junctions [CDD23, WGMJ23]. June [Ano20f, Ano20q, Ano21g, Ano21s, Ano22g, Ano22s, Ano23g, Ano23s, Ano24g, Ano24s]. juvenile [KB21, ZYZ22b].

K.L [PM20]. Kaczmarz [CH20a, GZH22, ZL21b]. Kadomtsev [GLZB20]. Kalman [BTL^{+}22, GMH^{+}20, LWL22, LLZB24, LY22d, NBTP23, WWW24a]. Kansa [Kar22]. Kansa-radial [Kar22]. Kantorovich [CCG^{+}22, CSV20, GA22]. Kantorovich-type [GA22]. Kaplan [SJTCdlFL21]. Karman [Kan20]. Kaup [DYA21]. KdV [BKU20, PW22b]. Kekulé [LY23a]. Keller [ALL23]. Kelly [WTCC20]. Kemeny [PM21]. Kepler [TO20]. Kernel [EA24, AD21a, AS21b, ADB20, Cav22, CZMX21, GW21a, HVYMS21, HSY24, KIB20, LW20a, OKN^{+}22, QQLZ22, SMW22, XZ23a, XTNF22, YTZ22]. kernel-based [XTNF22]. kernel-related [HSY24]. kernels [DYT21, HLSL22, HSY24, LQ24, LHS21, NPE22, QXG21, RSA24, ZZY24]. Kerr [ZXB20]. Khalatnikov [MVCK24].

Kind

[CUA^{+}22, AAS23, ARST21a, CKN21, CM22, EHVMM24, FRRS24, HWW20, PH20, QYD22, SXC21, SYX22, Wan22a, ZJX22b]. kinds [HWL24, WMF23, XLZL23]. kinematic [MC22]. kinematics [EFJO20]. Kinetic [DS21c, AR24, BDH^{+}23, GB20, GP22, LLT22, THW^{+}21, ZLH21]. kinetics [AS21a]. kink [MY24]. Kirchhoff [HCS23, PM21, SSP24, SPX20, SPX24, SY23, YWFL23]. kissing [GK20]. kite [BHGPC21]. kKr [LY22a]. Klein [JZ20, LZ24c, LWCD24, NNB21, SM21b, SW22, VPK20, WL23]. knapsack [HMS20, WJK20]. Kneser [BLOP23, BBDVPV20, JD20, JWWL20, LCL23]. Kneser-type [JD20]. knot [DKLV23]. knowledge [HCM21, HPC24, NS20, SM22, Ued20, WMWS21, ZJ23]. known [DLS^{+}23, QYLW24]. knuckleball [NS21]. Kobayashi [Hon24]. Kolmogorov [FLR23, PZS24]. Koopman [CCL^{+}24]. Korea [JC22b, JKJC21]. Korteweg [AF21, YZLS20]. Kotera [ZLA^{+}21]. kovian [SWX22]. Krasovskii [JLP23]. Kriging [CDP23, KG21]. Kronecker [Miy23]. Kronig [VGdFCBA20]. Kronig-Penney [VGdFCBA20]. Kundu [HZXZ20]. Küppers [KZS20]. Kuramoto [GBH^{+}21, LHH24]. Kutta [BMJ20, CCSPMV23, DS21a, RAB24, ZZX20b].

L1 [JCSH24]. L1-ADI [JCSH24]. L2 [AH21, LWNZ22]. L2-1 [LWNZ22]. label [YCZ^{+}20]. Labeling [Lyu23]. labelling [HJL^{+}23]. labor [HH23a]. laceability [SMQ24]. laden [ZWJ^{+}24]. lag [AAJ23, YWR^{+}24, vdMvRS21]. Lagrange [CCCdLMR23, CM22, DMSS20, FJLG22, DDCM23, GW23b, JTC^{+}24, LSWL20, LHX^{+}22, LHZ21c, PGZ^{+}24, SO21a, WJL^{+}24, YMW24, ZZX21, ZJW^{+}22]. Lagrange-hybrid [SO21a]. Lagrange-projection [CCCdLMR23, DDCM23]. Lagrangian [CG20b, ABPP23, BDRM23, CWZ22, CXYZ23, GLPS22, HP20, KWW22, MDSS20, PKJB22, PBKS23, TM22a, XYKL20, ZWJ^{+}24]. Lagrangian-Trust [CG20b]. Laguerre [CJ21, ÖÇÖ23]. Lakzian [GN20a]. Lambert [ÅCL24, JMM20, Lóc22]. Laminar [LZLM20]. laminated [AS20]. Landau [Kud20a, MVCK24, RD24, SL20b, ZXB20, ZLU21]. landslides [GdFPF24]. Landweber [MG21]. Lane [Iza21, IS21, SS24c]. lanes [ZL23a]. Langevin [GAA^{+}24]. language [UH24]. Laplace [BCVV22, CCC20a, DHR22, KGPS22, LPP24, MT24, UUS21, WFHG20, Wri24, ZrLkL^{+}23]. Laplacian [Baz20, BOT24, CT24, CCdlH21, COPW23, FKK24, Gan21, GRD23, HBCA21, HCS23, HDZ22, JZ22, LLZ23, LL22a, LLZ20c, LY23b, MNZ20, RD23, RGS23, SLJ23, WLW20, WYG23, YW20, YTGC20, YHW^{+}23, YH24a, YW21, ZCT22, ZHZZ24]. lapse [ZZ24]. Large [KFM21b, YXYZ23, Zha24b, AR22, AO21b, CSW20, CWY^{+}20, CY21a, CLLA23, DDS24, HCJC21, HS23b, KBL^{+}24, KFM21a, LLG22, LWL21, ML22, Miy23, PGPCK21, ST24, SNS^{+}20, SLCS20, XL24a, YLL22, YC21, Yu23, ZZVL23]. Large-Scale [YXYZ23, CWY^{+}20, CY21a, CLLA23, KBL^{+}24, LLG22, ML22, SNS^{+}20, YLL22, YC21]. larger [LTW22]. larger-scale [LTW22]. largest [XLS21]. laser [HLLW21, TM22a]. laser-induced [TM22a]. late [FR24, Zha21d]. latent [CCF24]. latest [MH24]. latex [RSZA23]. Lattice [Ily23, JNY^{+}21, JLWX21, LQW20, LZY20, LYN^{+}21, LX22, LWL^{+}20, NKO^{+}23, jSZyW^{+}20, WWQ22, WX22, YL22, YOF21, ZCS21]. lattice-Boltzmann [YOF21]. lattices [BFMGA20, Çay20, KR21, WYZ24, ZJC21]. law [BMDE21, CMMN20, CIHP^{+}21, KK22, PWZ22, ZC21c, ZXB20]. laws [AB23, FCH21, FZ23, GBCP21, HCTS23, HC21, MS22a, YTC23a, YF21a, ZZL22]. Lawson [LWCD24]. Lawson-type [LWCD24]. Lax [HCTS23, HLZ22, LL23a]. layer [DS22b, Eva20, DDCM23, LSL^{+}20, LXH24, MMK^{+}23, MWDD24, MK21a, PS22, RJHC24, SSCG20, Tak21, TM22b, WZW24, WZW^{+}21, WSL^{+}21, ZZY^{+}21, ZGZ^{+}23]. Layered [PGYW23, ZP21, JGXL23, PM20]. layers [BS22c, BD20b, KET22, YR20a]. LDPC [GYM21]. lead [CDL^{+}20]. lead-time [CDL^{+}20]. Leader [CLC^{+}22, LHX^{+}22, SWS^{+}22a, WDTF22, ZLZ^{+}21, BRTR21, CDZW21, GZCS23, GL22, LNC24, TLHH21, WZW23, WYLQ24, ZDD^{+}22]. Leader-follower [CLC^{+}22]. Leader-Following [WDTF22, LHX^{+}22, SWS^{+}22a, ZLZ^{+}21, BRTR21, CDZW21, GL22, LNC24, TLHH21, WZW23]. leader-following/containment [CDZW21]. leaderless [WWW24b]. leadership [KXW24]. leading [LDZ21]. leaf [LYhH^{+}24]. leaf-distance [LYhH^{+}24]. leakage [ANS^{+}20]. Leap [GKT^{+}23]. Learn [CLH24a]. Learning [AHM23, CDP23, CCL^{+}24, LHS21, WWY^{+}20, WWY22, ALM22, ACMT23, AO21a, AO21b, AGCGSC21, CCZ24, CTWG24, CLH24a, CHHC24, CZWX21, DWT^{+}21, DZ21, DWS24, GLL^{+}22, GS24, HCBY21, HXWC24, JLZ^{+}22, KMM^{+}24, LJL21, LJW^{+}21, LSHH22, LD23a, LYX23, LTW24, LZZL21, LF24, MPC^{+}24, PSSU21, PHS^{+}20, QQY^{+}23, QLKW20, QDW22, QZC^{+}24, SJ23b, TKY23, TLPT21, Ued23, UU21, WLQX20, WCP20, WGW^{+}21, WWW^{+}22, WYCL24, WDZL24, WFKL24, WHJ^{+}24, XTS^{+}22, XKW^{+}23, YLC20a, YZPL24, ZG24, ZPY22, ZSZL20, ZTW22, dAGSFM22]. learning-based [HXWC24, KMM^{+}24, LZZL21]. Least [REC^{+}23, RW21, BD20a, CHY22b, DK22c, HPT22, HM20b, MZX21, MSN21, SS22b, SS21c, TLCM24, WWY24, YCEW21, ZH22a, ZL23b, ZQZ21]. least-square [MZX21, ZH22a]. least-square/fictitious [ZH22a]. least-Squares [YCEW21]. leaves [CCHZ23, DS24c]. Lebesgue [Gu20, KL21]. Lecar [BSM23]. Leer [TC20a]. Leer-type [TC20a]. Leffler [BA20, KIB20, KFÖ23, LJAM24, LZH^{+}20, LWL23b, ZLK^{+}22, ZW22b, ZQC24]. Legendre [CP20, CJ21, Jor21b, MKK23, SRK^{+}20, ZY21b]. Legendre-truncated-exponential-based [SRK^{+}20]. Leibniz [ZL22d]. Leibniz-type [ZL22d]. lemma [CHG23, WHPQ23]. Length [LDZ20, DWW24, FKVZ21, HZ20a, HJL^{+}23, SS24a]. lengths [Che22, KPS22, WWY24]. Lengyel [ZH20a]. Lepturus [JKJC21]. Less [LPKC22, BSCV24, Mon20]. level [DESZ23, CLL24, CSC20, CFPM23, Fen22, FG24, SLCS20, TEGP24, WHJ^{+}24, ZS20a]. level-set [CSC20]. levels [RYR^{+}23]. Levenberg [AYY24, BGB24]. Levin [GW21a, ML20]. Lévy [AMOS23, CFWL24, HX22, WHC20a, XZ21, YDSW22, YD22b, ZCLW22, ZLD22]. Lexicographic [AS21a]. LFC [LSC^{+}23, WYZX22]. lid [Pap21, RWGS22]. lid-driven [Pap21, RWGS22]. Lidstone [MSB22]. Lidstone-collocation [MSB22]. lie [DYA21, AEL21, BJK^{+}22, HSM21, ZN20, ZL22d]. life [AALMLM22, LWZ22d]. life-cycle [LWZ22d]. lift [BS22b]. like [AFL24, Baz20, CZZP24, CW22, DESTP22, GHP24, GGY23, HKCK24, IHF13, LTB23, MJZ21, NX21, SHWW21, Sin20b, Woo20, YZW20, Zen21, ZLA^{+}21, ZO21]. likelihood [DDLW21, NHBJ20, Wri24]. Limit [BOT24, GT22a, GH20b, LZ24c, PPS20, XZZC22]. Limited [ASS24, GCLHT23, JM23a, KL22, LJP21, SL20d, WWZ^{+}23, YZT^{+}21]. limited-angle [WWZ^{+}23]. limited-aperture [KL22]. Limiting [CSZ24, CD23, CHK21]. Lindley [STH24]. line [BDMB20, BBK^{+}24, CDH23, Che23, CZR21, OPP22, SS20, YFHC21, YXXZ23, You20, ZP20, ZHW22, ZyZW21, ZHFL23b]. Linear [CHK21, DS22b, GZX20, VSM23, WWY21, WZ24, WPE20, YFW21, ZZ22a, AP22, AEL21, AA24, BRFLGH21, BDPV22, BOT24, BB20, BE20, BIJ22, BCVV22, CZLH20, CCC23, CFP22, CET22, CWZZ22, CKN21, CGJ21, CYG20, CWY^{+}20, CLZY21, CHY22a, CGZ^{+}22, CHH22, CW23, CLZ23, CL24b, CGLR23, CLC20, DdlCS24, DP24, DDR21, DDGN23a, DDGN23b, DZDL22, DHS23, DKLV23, DFG22, DYX24, DS21b, Du24, EW21, ED20, FMP20, FRS22, Fre22, FJJJ21, GMH^{+}20, GN21, GB21, GNA22a, GNA22b, GZ20b, GZLZ24, GH20b, HHGZ23, HSM21, HT20, HLP24, HL24b, HJW^{+}23, HWWW23, Hua24, HSL^{+}24, HGZ24, HS23b, IS21, JD20, JPY23, JT20, KGPS22, KKB23, KRT21, KLPL20, LZ24a, LaGBT22, LMX20, LH22, LJZL22, LZ22b, LY23a, LZZG23, LHYL24, LZW^{+}20b, LCL22a, LZ22e, LLWK23, LU24, LY20b, MWZX21, MS22a, MHAT21, MF23, NP20, NNN24]. linear [PKL23, PDK23, PWWZ23, PMG20, QZ22, QXSW22, QQ23, RBPN20, RRS21, RO21, SS24d, SG24, SZG21, SD22a, SK20, SWF20, SKK^{+}21, SZSZ22, SYX22, SKJM21, SM23, SJ23b, SL24, TLY24, UGG^{+}20, WX20a, WDG21, WLY21, WZGL21, WXF22, WWW24a, WLZD24, Wu20a, XD23a, XZeF20, XW22, XCGN24, XTS^{+}22, XZZ20, YZLS20, YAX^{+}21, YCWH23, Zen21, ZTLL24, ZY20, ZZS^{+}20, ZCPZ20, ZDZS21, ZL21b, ZLS21, ZS21b, ZWZ^{+}21, ZJJB22, ZZLV23, ZrLkL^{+}23, ZZVL23, ZDZ24, ZZL20b, ZV20, ZV21, ZS22d, ZWY^{+}22, ZLU21, ZZL22]. linear-deteriorating [CYG20]. linear-implicit [YZLS20]. linear-quadratic [HWWW23, SJ23b, ZY20, ZS22d]. Linearisation [GLX22, GLWX24]. linearity [Wu20b]. Linearization [GBH^{+}21, Ach21, MVO21, XLD23, ZZS^{+}22]. Linearized [HHGZ23, HZT^{+}21, LMZ20, QWD22, YLG21, YCWH23]. Linearly [SRK23, UK23, WSZ22, YLC20a, Zha21a]. linearly-implicit [Zha21a]. Link [LDZ21, WCWZ23, YZM24, DZW21, HWL24]. linked [LJH^{+}20, WLY20, WZY21]. Linking [CFA^{+}23]. links [CFWL24, LWG20, XYLF21]. Liouville [BBHM21, Bon22, HBCA21, MBMA20, SLTL20, ZL20b, ZJC21]. Lipschitz [ZZX20a, BRTR21, CPS24, CSL24, DL20b, GJ21, GGY22a, HRA^{+}20, LNC22, LCL22c, Mar22, RFFZ21, SNRA20, WJC^{+}20, YLCZ20]. Lipschitz-continuity [Mar22]. Lipschitzian [LLY22a]. liquid [CSG^{+}22, CSG^{+}23, LZY^{+}23b, PS22, SNZC20]. liquidations [ZCLW22]. List [CLCD23, LLL23b]. Listeria [BCLNV20]. little [FR24]. Liu [CL23, GLP22]. LKF [ZYC^{+}24]. LMI [CWH24, HSZ20, JMC23, ZL21c]. LMIs [HDP^{+}21]. load [AEE21, AS20, CJC20, DD20a, LHZ^{+}23, LCL22b, LLQW24, RLY20, WCM^{+}21, WPJ22]. loaded [ASS^{+}23a, WPE20]. loading [PV24]. Local [HLL^{+}21, KT20, LFA22, SLZ20c, Zha23b, ABD22, AA21b, ADS23, BS22a, BA20, BS23, CG21, CZR21, DRS20b, DL21b, EKKE20, ES23, FGM24, FCH21, HMR23, KK23c, KWW22, LF21, LHQ^{+}23, LMS24, MPM23, MTRA22, RKRRL21, SLZ20a, WZZ23, WX22, Yoo21, YLC20b, Zha22b, ZF23a, ZHQ^{+}24, ZHWS24, ZHMZ21]. localisation [NA21]. localization [Kwo22a, Kwo22b, MMDD21]. Localized [WFHG20, WZCF21, LaGBT22, SFC23, XFW^{+}21]. locally [BRTR21, EFF23, GW23a, LZC22a, Prz24a, Prz24b]. Locating [BBT21, SNC22]. Location [LSVA20, AA20, GLW24, MBL20, NTNT24]. locations [Wan21a]. logarithm [AB21b]. Logarithmic [AOSO20, JLFL20, KX23, Vyn23, WJK20]. Logarithmic-Sobolev [AOSO20]. logic [xHhPjW^{+}21]. Logical [GR21, CLLA23, LLL21]. Logistic [AB20b, LHET23, Alf21, PMG20, RP20]. loner [INT20]. Long [GTK^{+}22, LW23a, MH24, YYF24, AC20, BFRS24, CGMFSHR21, DSAY24, DR22, JDLZ21, qLgZ21, RO21, WDG21, ZJJ24]. long-time [AC20, BFRS24, DR22, ZJJ24]. long-wave [RO21, WDG21]. Longest [TPSP24]. longitudinal [KY21]. LOOCV [Cav22]. LOOCV-based [Cav22]. look [Chi22]. loop [BSCV24, EVB^{+}21, HY24]. looped [LSPK20, WHCW20]. looped-functional [WHCW20]. looped-functionals [LSPK20]. loops [FKVZ21, QCARL20]. Lortz [KZS20]. Löschian [KRNG24]. loss [RCCL21, WHW^{+}21, XSP22]. losses [Chy22d, Chy22c]. lossy [HKCK24]. Lotka [TC22, WBV22, XKW^{+}23]. Love [MKKG22]. Love-type [MKKG22]. Low [BEIR23, YZJ^{+}20, YY20b, ADS21, DHM^{+}20, FMN21, JDZH22, KTEM22, LFLH23, LHZ^{+}21b, Oru21, Oru22, PLH^{+}24]. Low-degree [BEIR23]. low-order [JDZH22, KTEM22]. Low-rank [YZJ^{+}20, DHM^{+}20, FMN21, LHZ^{+}21b, PLH^{+}24]. Lower [BY24, FCX21, KRE20, KFR20]. lowest [WCY20a, WCY20b]. loyalty [CCC20b, ZHQ^{+}24]. LPV [JHSL21, WXW^{+}20, XWZC21, XZQ23, ZWZ^{+}21]. LRBF [LLY^{+}23]. LS [SL24]. LS-SVM [SL24]. LSM [WPE20]. LSTM [CCF24]. LTI [AKDR23, ZSP24]. Lü [WA22]. lubrication [ZV21]. Lucas [YY22]. lucky [KRE20]. lump [GLZB20, ZLA^{+}21]. lumped [DWS24, WC22b]. lung [LLZ^{+}20a, LFLL21]. Lur’e [HB20]. LVI [KMS^{+}23]. LVI-PDNN [KMS^{+}23]. Lyapunov [Ahm21, CZ21a, CWJF21, ES23, HKKL22, JLP23, LT22, QNX^{+}23, XHL23, YWG24]. Lyapunov-based [Ahm21]. Lyapunov-Krasovskii [JLP23].

M [SCC24, GL23b, SCC24]. M/G/1 [GL23b]. M/M/1/m [SCC24]. M1 [EHA20]. mach [ADS21, BRMVCD21]. Machine [AGCGSC21, ZG24, DTP21, HH20, LLZ^{+}20a, MGM21, QLKW20, TLPT21, WCP20, dAGSFM22]. machine-dependent [HH20]. mackerel [JC22b]. macro [VS24, WHD^{+}23]. macroeconomic [GT22b]. macroscopic [CSZ24]. Mader [LLL^{+}23a]. Mader-perfect [LLL^{+}23a]. Magnetic [RHR^{+}21, UUS21, ABD22, BP23, CH24a, LYN^{+}21, Ma21a, ME21b, NKO^{+}23, PPS24, RRR19, RRR20, WHM22, WGMJ23, YTLV22]. magnetization [Yan23]. magnetized [FWM^{+}22]. magneto [BS22a, MK21a]. magneto-elastic [MK21a]. magneto-hydrodynamic [BS22a]. Magnetohydrodynamic [BH22, AKH^{+}22, CZ22c, CCSPMV23, LCXJ22, XLG^{+}21]. Magnetohydrodynamics [BDH^{+}23, SWS20]. Magnus [DdlCS24, HC20]. Magnus-based [DdlCS24]. Mahoney [SL20c]. Mahony [HA20]. main [FLS20, SY22a]. maintenance [HDQP23]. makespan [AAA20]. Maki [BBH21]. making [MMHX20, WLS22, ZHLZ21]. malaria [FM23, GOO20, ID21, MF21, WOB21]. malicious [HXWC24, ZLC24a]. Malmquist [SKR22]. Management [YY20b, AVJ24, AGCGSC21, LTW24, LZX^{+}22b, TCGTT21]. Managing [ASUH21]. Mandelbrot [WLWL22]. maneuvering [IFHA21]. manifold [AEL21]. manifolds [BFGGM24, CG23, ŚW21]. Manipulate [YHL^{+}20]. manipulation [RSR21]. manipulator [CZZP24, WWT24]. manufacturer [HWZZ23]. manufacturing [YLX^{+}21]. many [KX23, WLL21]. map [BE20, CLZ24, LHET23, LKX20, RLX22, jSZyW^{+}20, TEGP24]. mapping [HKU21, WZW24]. mappings [OK21, WWZ20a]. MAPS [GBF21, AGA20, GS22, LLH21, Ma24, SS24e, ZH22b]. Mar [SWX22]. marble [CSC20]. March [Ano20g, Ano20r, Ano21h, Ano21t, Ano22h, Ano22t, Ano23h, Ano23t, Ano24h, Ano24t]. marching [DK22b, Qi22, YGM^{+}22, YMRJ22]. margin [GLLZ24, QL20]. marginally [JSM20]. marine [DSZ^{+}23, HZL21]. market [Ask21, HXL^{+}20, SG24, Ued20, ZY24b]. markets [ARISV20]. Markov [CZ20, CL24b, DP23, DD24b, DN21, DLS^{+}23, FJFS21, FRW^{+}24, FLWL21, FLC^{+}22, FJJJ21, GZQ^{+}20, GXZZ20, HRLH21, HQK21, HKCK24, KMM^{+}24, KKS^{+}22a, KNN20, LPK22, LSM23, LZZ^{+}20, LHS21, LLZ22c, LYL22, LZ24d, LOS21, MWS23, PHL23, QYG^{+}20, RCCL21, SSKS21, SWX^{+}21, SSW21, SRZ22, TQZ^{+}24, TMW20, VS24, VSM^{+}21, VSM22, WCZY20, WXW^{+}20, WQZ^{+}20, WPZY23, WM23, WC23, WLL^{+}24, WSDW24, WFKL24, WHCW20, WCZ^{+}21, XH22a, XZeF20, XCW^{+}22, XTS^{+}22, XWZC21, XZQ23, YCZ^{+}23, YM20, Zha22c, ZCS^{+}22, ZM21c, ZHJ^{+}21, ZWZS22]. Markov-based [LZZ^{+}20]. Markov-modulated [LOS21]. Markovian [ASKA20, CH24b, CLZ^{+}22, CZ24b, FFL23a, GCLHT23, HWD20, HKL22a, HL24a, JWK22, KGCC20, KKPS23, LLZ20b, LNS^{+}21, LL22b, LM24, LLD21, LW22f, NK20, NK21b, NK21a, NLP^{+}22, RH20, RFFZ21, SDW^{+}22, SKSC21, TWR20, TW21a, WCZ20, WGL^{+}21, WZC^{+}22, WZZ24, WCX^{+}22, XXL^{+}21, Xia22, XXW^{+}20, XZQX20, XZQ^{+}22, XWDK23, YMW24, YD22b, ZRSK21, ZJZD20, ZYSW22, ZWY^{+}24, ZYX21, ZLJ^{+}21, ZC21b, ZeF21, ZLLX22]. Marquardt [AYY24, BGB24]. Martin [Lu22]. Maruyama [LH20b, LHC21, SGL21]. MAS [ZYXL23]. mask [GZLY24]. mass [dSAOG21, BJK^{+}22, CS24a, He23, HCH24, LXL22, MCD22, RB21, YTC23b, YXXZ23, XBLW23, XD23a, Yi23, ZCD23]. mass-based [He23]. Master [HB20]. Master-slave [HB20]. Match [MMHX20, CFF20, ZLC^{+}24b]. matched [KJL20, LJ23, SLHX21, ZWZX21]. matching [CJCL21, CLL23, JJC24, LZ24a, LH24a, SM24, WWK24, XBC22a, YD22a, ZLL20]. matchings [BGH24, HP20, JMY22, MY24, PLH22, SS22a, SD23]. material [BSRV23, CSG^{+}22, CSG^{+}23, DMNV20, KR20, ZYY20]. materials [DSY22, LWQW22, NS22, XAZ^{+}21]. Math [AACGTB24, LPP21a, PPS24, WDG21]. Mathematical [CS21b, CGMFSHR21, EKKE20, GOO20, KIJK20, LZ22a, LD20c, LGS22, MS22c, PPS24, RRR19, RRR20, SSCG20, VMCG20, WOB21, CM21a, DE21, DNR21, DBVCP23, Gök22, HT23, HCY21, HPC24, RMCG20, RT22, SBT22]. Mathematics [ABNS24b, AV21a, CSG^{+}23, CF23b, Elg22, GNA22a, KWM20b, KA23, KFM21b, MV23, NHK^{+}23, Oru22, RRR20, RO21, SZ22, WCY20a, WHM^{+}21a, XH23, ZBA^{+}23a, BKS21, MM24, Yua20]. Mathieu [MP23, PMP21]. MATLAB [MV22, SZCP21, IP23]. matrices [AACI21, AB20a, AC22, BF20, BGS22, BF23, CdFP24, CH24b, CS24c, CIHP^{+}21, DEHFF24, DZZ23, DFX24, EW21, FMN21, Gas23, HV24, HHR21, Kop22, KLNN23, LKP22, LZCL23, LPL^{+}22, LZL20, MH20, Miy23, ÖC24, PP23, RR07, SZ24b, SD22a, Sin20a, SS21c, TLPY20, UHZ^{+}20, Voe23, WC24, WC22c, WYG23, XH22b, XH23, XBC22b, ZWY^{+}24, ZZL20b, ZV20]. matrices-based [UHZ^{+}20]. Matrix [BFMGA20, CTF24, FCeFZ21, IÖÜP24, IKÖ24, Kut20, SS23c, ADF^{+}21, AA24, AS23b, BT21, BR20, BD20a, BSCV24, DESZ23, Cas23, Cha22, CPLH20, DP24, DFG22, DWZ20, DGMP22, ER22, FLZZ24, FZ20a, FB20, FJ24, Gan23, Gas23, GLZA21, GJWV24, HV24, HHR21, HP23, HLZ22, HLX20, IÖÜP23, KKK22, LK21a, LZ22d, LWL^{+}20, LLFZ20, LSWS21, Ma21b, MPM23, MDSS20, MS24, MZL24, QXX21, QLC^{+}24, RS21, SM22, TM20, TWR20, TLDL21, TWWL24, Wan21b, WL21a, XL24a, XD23b, wXhGZ20, wX20b, Xu20, YM21, YLZL21, ZN20, ZWZ^{+}21, ZLLY22, ZK22, ZJWV24, ZV20, ZSJ23b]. matrix-less [BSCV24]. matrix-weight-based [LSWS21]. matrix-weighted [MS24]. matter [TCZ^{+}20]. Max [LT22, GA22, NTNT24, YLH20]. max-cut [YLH20]. max-min [GA22]. Max-separable [LT22]. max-sum [NTNT24]. Maximal [ACSVT22, BDH^{+}20, DHZ23, JL21, TLD22, YML21, CLL21, DRS20a, LFLP21, RiTZ24, SD23, TT24, WZ20a, WHKT23, ZL21b]. Maximisers [GLPS22]. Maximization [DM24, LCW^{+}23, LYZ22b, PV24, wXhGZ20, wX20b, ZEA22]. Maximizing [XT23]. Maximum [DDLW21, DQC^{+}21, YD22a, AY22, BAH^{+}21, CYG20, DRS20a, DX24, HK21, LD23b, Mef21, Nak23, NH21, SM24, TLD22, VMR23, WSY21, Wri24, XYZZ22, YCW23, ZCM^{+}24, KZ22]. maximum-principle-satisfying [DX24]. Maximum-Weight [KZ22]. Maxwell [CSG^{+}23, CSG^{+}22, LCXJ22, MMS21, NI21, WHY20, XTT22, YJ20, ZL22a, ZLU21]. May [Ano20h, Ano20s, Ano21i, Ano21u, Ano22u, Ano22i, Ano23i, Ano23u, Ano24i, Ano24u]. McKean [CdR22a, HANP23, dRST23]. McKean-Vlasov [CdR22a]. mean [AM24, BCVV22, CZ22b, CDM21, GJ21, HWWW23, RAB24, SSR23, TDZ^{+}24, XZeF20, YMW24, ZY21a, REC^{+}23]. mean-field [HWWW23, TDZ^{+}24]. mean-reverting [ZY21a]. mean-square [RAB24]. Means [SS23c, CZMX21, CTV21, JWY20, MHF23, RSZA23]. measure [CTF24, CRTV22, DN21, Kum23, YL20b]. measure-based [CTF24]. measurement [LWL22, LPL^{+}22, LTL^{+}21, LWLA21, TLPY20, WSL^{+}21]. measurements [CSL24, FMN21, HLKC22, Kwo22b]. measures [CDBR^{+}22, DD24b, LRCTR22, SG20a]. measuring [WYLW20]. mechanical [SS22c, WZY20, WPJ22, KY24]. mechanics [ABD23, EM20]. mechanism [CDZW21, CC22, DW22, JZ20, LNS^{+}21, LTL^{+}21, LCL22a, MZY^{+}22, MTFD24, RSXH20, SWZZ23, WLZF21, WCW21, WZC^{+}22, WCW22, WF23b, WSL^{+}21, Xia21, XBLW23, XSP22, YY24, ZeFWW20, ZWY^{+}22]. mechanisms [LGC^{+}24a, LGS22, LY20b, RJHC24, SGX21, YXW23]. media [CKV21, CEPZ22, DKK20, HCH24, JT24, KRS20, LKK^{+}23, LWX24, MMDD21, OBP21, PJMW22, QPHC21, RM20, WYLW20, YFHC21, ZHJ^{+}21]. median [AA20, CM21b, NTNT24]. mediated [CS21b, JLWX21, NNKH21]. Medical [WHJ^{+}24, AHS20, JM23a, LHET23]. medium [Azi21, CK21, FTML22, GE21, GK20, HT23, HL23b, Ma23, MJK22, OPP22, RHN^{+}24, SN21, SGI20, Sha21, ZyZW21]. meet [CGTU20]. Meissner [YY21a]. melodies [CDBR^{+}22]. melts [LTC^{+}23]. membership [KJL20, LL23b]. Memory [CYZ23, DZ21, HXZ24, LLQW24, SWW21b, Ued23, WL21b, XCW^{+}22, AMF20, CK20, DZZ^{+}20, Gök22, GTK^{+}22, LSHH22, LHZ^{+}23, LWL23b, LSC^{+}23, LZZ^{+}23, LLT^{+}20, RCG21, WHW^{+}21, WGLL20, XLYA21, YWR^{+}24, ZZY24, ZWZS22]. Memory-based [DZ21, HXZ24, LLQW24, SWW21b, XCW^{+}22, LSC^{+}23]. Memory-dependent [WL21b]. Memory-two [Ued23]. memoryless [NL20]. Memristive [HHY20, SRKA23, BZZ21, CFWZ20, CXYZ23, DZZ^{+}21, FZS21, HWL24, LCXM21, LSYX21, TQZ^{+}24, WCG^{+}20, WCGW20, WHM22, WGMJ23, YWS^{+}20, ZLZ^{+}23a, ZWYL22, ZB22]. memristor [ANS^{+}20, DL21c, HP23, JYMP^{+}20, LCY^{+}21, TXZ24, ZQ24]. memristor-based [DL21c, HP23, LCY^{+}21, TXZ24, ZQ24]. Menger [Guo23]. mental [Chi22, ZM21a]. Menten [LNO23]. Mercenary [LCS22a]. merging [LYN^{+}21]. Merton [ARMM23]. mesh [BS20, GKK^{+}24, KZZ20, LZ22e, LTB23, LZ22g, MZ23a, WHJK23, ZL20a, ZL21a, ZL22b]. meshes [BBD^{+}22, BRMVCD21, BDRM23, CH24c, CGLR23, CFPM23, DK22c, GKK^{+}24, KSR21, MS22a, PGF20, Que22, SCMK23, TKF23, WSY21, XYM22, ZM20, ZZY24]. meshfree [CD20a, LWQW22, Mar24, WXLY23, YLC20b]. meshless [BS22a, BSRV23, GBF21, HL24b, LD20b, LZRF21, MZX21, SS22c, SV21b, Zha22b]. Mesoscale [LTC^{+}23]. message [LZW^{+}20a, YFM^{+}24]. metacirculants [CZ21b]. metadata [AR20]. metaheuristic [CT21]. metaheuristics [SJTCdlFL21]. metapopulation [SH22, ZLW^{+}21b]. Metastable [AKMR21, TWZ^{+}20]. methanizer [TWB23]. methanol [SIK22]. Method [CUA^{+}22, KWHS23, WGLS20, AD21a, ABD22, AB23, AP22, AS23a, AACJ^{+}22, AEL21, AMPS20, AA24, AM20b, AS21b, AHB20, AFG21, AACGTB23, AACGTB24, ACGG20, AHEGM23, AT21a, BS22a, BSRV23, BKK24, BPR21, BO21, BGB24, BD20b, BMY24, BZS21, BDRM23, CS24a, CM20a, CZ21a, CG21, CCdlH21, CLL22, CWZZ22, CN23, CH24b, CH20a, CZL20, CHY22a, CZ22c, CXYZ23, CYZ23, CLL24, CKWY23, CCE^{+}22, CJ21, CKL22, CEPZ22, CG20b, DLL^{+}23, DDR21, DK22a, DWT^{+}21, DK22b, DTK24, DR22, DSY22, DFG22, Dol21, DL21b, DHZ^{+}20, EM21, EGMY23, EA24, Far22, FWBY23, FZLJ24, FB20, FY23, FG24, FQHS24, GGY22b, GE20, GC20, GK20, GY22b, GW22, GBF21, Grz22, GWH24, GK21, HD23, HLZ20, HZZ^{+}21, Hon24, HGKD22, HL24b, HZT^{+}21, HWL24, HC21, HL22, HXF23, Hua24, HGZ24, HL23b, IÖÜP23, JPY23, JNY^{+}21]. method [JZ20, JLFL20, JCH^{+}20, JGXL23, JCSH24, JL24a, JT24, JM23b, JSM20, JP22, KSH22, Kar22, KG22, KGPS22, KAS21, KPP22, Kod21, KG21, LHET23, LP22, LSZ^{+}23, Lee20, LD20b, LZLM20, LTL^{+}20, LH20a, LZY20, LH20b, LHC21, LF21, LHZ^{+}21a, LJL21, LLQG21, LWZ22a, LYZ22a, LW22d, LLLC23, LL23b, LTW24, LWCD24, LGC24b, LT22, LLWY23, LZRF21, LWL21, LZL^{+}21, LZX^{+}22b, LM22, LNC22, LCL22b, LCXJ22, Liu22a, LZ22e, LCL22c, LPP24, LW24, LHFH22, Luo20, LZ22f, LZ22g, MGS20, MZX21, MZX22, MYC^{+}23, MZ23a, MPM23, MM22b, MH22, MMDD21, MGJ21, MMG^{+}23, MSN21, MG21, MSB22, MVO21, NS22, NE21, NL20, NKO^{+}23, Oru21, Oru22, PH20, PP21, PBM20, PBKS23, PWWZ23, PGPCK21, QS20, Qi22, QPHC21, QCARL20, QWD22, QXG21, QXX21, QSL21, Que22, REC^{+}23, RC22, RS21, RKRRL21, RX20, RMSP22, RPB23, RG20, SM21b, SSI23, SR23, SSCW22, SL20c]. method [SL20b, SW20, SHWW21, SW22, SWZZ23, SUN24, SJP22, SD22a, SLZ20a, SS22c, SWW22, SV21a, SJ20, SLHX21, SWX24b, TS23, TLHH21, TLC22, TLJO20, THW^{+}21, TZJ20, TWWL24, TM22a, TZ23b, UHZ^{+}20, VS24, WW20, WZXY20, Wan21b, WZCF21, WWWZ21, WHL21, WC21, WLD^{+}22, Wan22a, WM23, WWZ^{+}23, WLHW23, Wan23a, WKM24, WZZ24, WFKL24, WPE20, Wen20, WH21, WDYS21, WX21b, WGJK21, XZZ22a, XTT22, XL24a, XYKL20, XHWM20, XCSW21, XTNF22, XHL20, YWZT20, Yan23, YWZ23, YWL20, YTZ22, YXW23, YOF21, YCWH23, YLC20b, You20, YWXL21, Yu23, YY22, ZMS23a, Zen21, ZL22a, ZL20c, ZL20a, ZM20, Zha20, ZP20, ZSL20, ZSM20, ZC21a, ZL21b, ZL21a, ZHLZ21, Zha21b, ZF21, ZH22a, ZFS22, ZL22b, ZF23a, ZZVL23, ZRL^{+}24, ZZW21b, ZZX21, ZY21b, ZLW21a, ZeFWD23, ZSYX23, ZV20, ZLJ20, ZS20a, ZV21, ZH22b, ZGG^{+}22, ZF23b, ZWJ^{+}24, ACS24, CHK21]. method-based [LHZ^{+}21a]. method-How [JP22]. methodologies [CDL^{+}20]. methodology [FGMS21]. methods [ABNS24a, ABNS24b, ABNS24c, AZD^{+}24, ARST21a, AS21c, AYY24, BDKZ21, BBHM21, BBH21, BCS24, BCP23, BIJ22, BMJ20, CG20a, CD20a, Cav22, Çay20, CKN21, CWY^{+}20, CY21a, CCZ24, CHZ24, CW22, CDDP21, CSQT21, CSV20, CIHP^{+}21, DS21a, DO21, DDGN23b, EV21, ELOY22, EA24, EBJ20, EFF23, GP23, GW21a, GLKrH20, GBCP21, GLP22, GZLZ24, Han20, HHGZ23, HM20b, HS23b, IÖÜP24, IKÖ24, JHO^{+}21, KL22, KWZX23, KKS22b, Kum24, LGQ20, LW22c, LLP23, LW23a, LCLH24, LD20c, LLFZ20, LYN22, MM20a, MRVC20, MN23, NT23, NI21, NSD^{+}23, NKA20, PX20, PGPCK21, RAB24, RD24, SWS20, SK20, Sin20b, SXC21, SGL21, SYX22, SV23b, SLZ20c, TZ24, Ten22, TM20, TLDL21, THBPMC22, Tre23, TKF23, UZ20, UK23, VMT24, WXLY23, WL24a, WSZ22, WCZ22a, XLD23, XZZ20, YGM^{+}22, YMRJ22, YHB^{+}20, YLS23, YS23, YLL20, ZT22a]. methods [ZF24, Zha22a, ZJJB22, ZZLV23, ZL23b, ZGZ^{+}24, ZH20b, ZZX20b, ZJX21, ZJX22a, ZZX23, ZZW20, ZHZZ24, ZMS23b, ZZL22, dSTV21]. Metric [SS22b, AKP23, BPR21, GN20a, KP23, KT20, KRT21, KMT^{+}21, NX23, PMAK22, SVW23, SS21a, SS21b, WWY22]. Mexico [CGMFSHR21]. MFS [RNVC21, WFHG20, ZDKC20]. MHD [NHK^{+}23, BDH^{+}23, NHK^{+}22, SSF24, SWW22, Wan23a, YS23, YK22]. MIB [SLHX21]. Michaelis [LNO23]. micro [CdTPP22, JA23, LSL^{+}20, MMUS22, NHK^{+}22, NHK^{+}23, VS24, WHD^{+}23, YZY22a, KY24]. micro-capsules [CdTPP22]. micro-dynamics [YZY22a].

Micro-Electro-Mechanical [KY24]. micro-macro [VS24]. micro-rectangle [LSL^{+}20]. micro-rotors [JA23]. microbeam [KET22]. microbeams [AEE21]. microbial [CD23, XKW^{+}23]. microblog [YXZ^{+}21]. microchannel [MS22c]. microchannels [ATDJ22]. microgrid [LZX^{+}22b, LZ24d]. microgrids [SXS22, WYS23]. microorganisms [SKK^{+}21]. microplastic [RWGS22]. micropolar [AKH^{+}22, BPRR22]. microscale [IMIH22]. microscopic [TAU^{+}22, Yan23]. microscopy [MJW21]. microshells [ASS^{+}23a]. Microstate [WQZD23]. microstates [BAL^{+}24, FdLPC^{+}22]. migration [DCCM20, LJH^{+}20, RMO22]. mild [PM20, Sin20b]. mills [WXWS20]. Milne [KWHS23]. Milstein [GLL23, LNC22, TY22]. Mima [KMN22, MN23]. mimetic [AT21a]. MIMO [HMZZ20, MXF^{+}24, WXW^{+}21]. min [GA22]. Minimal [FLL^{+}20, LLW20, WZPL24, AG20, CLB21, DD21c, FKVZ22, GGS21, Kwo22a, LLLL23, LHL24, NTLNX24, SMM^{+}24, ZZ22a]. minimal-ABC [DD21c]. minimality [BBB24]. minimally [CCSPMV23]. minimax [LS22a]. minimization [AACGTB23, AACGTB24, DHZ^{+}20, Kwo22b]. minimize [AAA20, CYG20, Zha21d]. minimizing [CCE^{+}22, TLCM24]. Minimum [SYW21, DTP21, DDH^{+}24, GWW23, HPT22, KMS^{+}20, KMS^{+}23, KFR20, LL20a, SS22b, WLZD24, ZZ24, ZCM^{+}24]. minimum-cost [KMS^{+}20, KMS^{+}23]. mining [FZG23]. Minkowski [XBC22b]. minmax [NH21]. minor [CFL^{+}20]. minorities [LJP21]. minus [CHST20]. Miodek [CLX22]. Mirakjan [AM20a]. misalignment [JZS^{+}24]. mismatched [DMD21, NK22, SMH^{+}20]. mismatching [MPX^{+}20]. missing [BTL^{+}22, HLKC22]. mitigate [SGJ^{+}21]. mitigating [DKKJ21]. Mittag [BA20, KIB20, KFÖ23, LJAM24, LZH^{+}20, LWL23b, ZLK^{+}22, ZW22b, ZQC24]. Mittag-Leffler [BA20, KIB20, KFÖ23, LZH^{+}20, LWL23b, ZLK^{+}22, ZW22b, ZQC24]. Mixed [GLX22, MJK22, NX23, WLHW23, YHB^{+}20, AP20, BG20a, BBD^{+}22, CTZ21, CLZ^{+}22, CGZ^{+}22, CXYZ23, CCE^{+}22, CDDP21, FWBY23, GLW24, GWZ^{+}21, GLWX24, GKK^{+}24, GX23, JL24b, KRS20, KPP22, LGL20, LW22c, LXL22, LCH20, LWW23, Ma21a, MYW21, ML22, MZY^{+}22, PP21, QXX21, SS21b, SKS21, SL20c, SY22b, SHYS23, WCZY20, WWY^{+}20, WGL^{+}21, WZC^{+}22, XZQX20, YWS24, YMW24, YWX20b, ZL22a, ZY20, ZW22b, ZZS^{+}22, ZLLX22]. mixed-mode [MYW21]. mixing [LXF21, LL20b, ZB23]. mixture [AGC20, HYS24, JG21, NHBJ20]. mixtures [BS22c, BBD^{+}22, ZGHM24]. MKM [ASS^{+}23a]. MNNs [LCC^{+}22]. MnZnFe [KHA22]. mobile [CWZL23, CHHC24, GZ20a, HCBY21, MC22, XJY24, YLL20, YP21, ZDD^{+}22]. mobile/immobile [YLL20]. mobility [CTV21, HDQP23, JLFL20, Lee20, LJP21, LCJ^{+}22, SH22]. modal [HR23, LS22a, vdMvRS21]. mode [CH24a, CZ24b, CLL24, DMD21, DWS24, GZX20, GLKrH20, GHP24, GW23b, GX23, HZL21, HCX^{+}24, IFHA21, JQJY22, JLOR22, JSM20, KZC23, KKS^{+}22a, LLZ20b, LCLS20, LLY^{+}21, LTW22, LZWC22, LL22b, LR24, LSM23, LTL^{+}21, LNZ^{+}22, LXK23, LR23, MYW21, MS20, NK20, RMCG20, SJ23a, SWSW21, SCG^{+}20, SRZ22, TMW20, VSM22, WCG^{+}20, WFYL21, WCW22, WM23, WZZ24, XRZH22, YLK^{+}21, YCL^{+}21, YL20c, ZJG21, ZZQ^{+}20, ZJZD20, ZCZ22, ZQP^{+}22, ZYXL23, ZWY^{+}24, ZJW^{+}22, ZC21b]. mode-dependent [CZ24b, JQJY22, JSM20, KZC23, LTW22, VSM22, ZJZD20, ZYXL23]. mode-switching [LZWC22]. mode-transition-dependent [NK20]. Model [ACBLC24, AMOS23, CNZ^{+}21, DJW20, Jan22, LOG20, LJY22, TZL24, WXF22, WS23b, AALMLM22, ARMM23, ALL23, AMA20, AGA20, ACMT23, Alf21, Alf23, ASAZ23, AKR21, AR24, AIAFvH21, ADB20, Azi21, BSdCCGCV21, BBAS21, BCLNV20, BKR22, BJK^{+}22, BLO22, BBD^{+}22, BGRM21, BHL23, BS23, Bri20, BJ21, CRTVG^{+}23, CLZY21, CTZ21, CZS21, CZ22c, CL23, CCL^{+}24, CH20b, CKW22, CSZ24, CC24, CGMFSHR21, CZR21, DWL^{+}24, DP24, DKKJ21, DGP^{+}23, DG21, DE21, DG23, DZL^{+}24, EBGN23, EA21, EB22, FRW^{+}24, Fen22, FZ22, FW24, FR24, GLGLH24, GGW21a, GZ20a, Gök22, GCLHT23, GM21b, GSW^{+}20, GZG^{+}21, Guo22, HZJ^{+}21, HQK21, HT23, HCY21, HXF23, HPC24, HJN^{+}24, HSY24, ID21, Ily23, IS21, JC22b, JKJC21, JH21, JLL^{+}20, JSZ20, JMC23, JG21, Kab24, KWM20a, KWM20b, KMN22, KZWS22, KK23c, KEiS23, LH20a, LCZ^{+}20, LHW21, LYN^{+}21]. model [LJL21, LMX22, LLY^{+}22b, LZ22a, LZ22c, LX23a, LTC^{+}23, LZ20c, LHZ^{+}21b, LNO23, LD20d, LCH20, LB20, LJH^{+}20, LY21b, LY21c, LGK^{+}22, LGS22, Liu22b, LW24, LWJ21, Lu22, LQNA23, MKKG22, MVCK24, MP22, MCSD21, MZL24, MK21b, MUMS22, MMS21, MN23, MFA21, NHBJ20, NGC21, NT23, NS21, NH22, OBP21, PJMW22, PDK23, PR20, PLW^{+}20, PHS^{+}20, PM20, PSE24, QLL24, QLKW20, QLG^{+}22, RMO22, RP20, RCCL21, RHNT21, RST21, RF21, RSR21, RCG21, RMCG20, RSA24, RT22, RSZA23, RM20, SRK23, SN21, SO21b, SO21c, SBT22, SK24a, SJ23a, SMW^{+}23, SWW21a, SJZA22, SLS24, SS23b, dOSR21, SWCT21, TZ23a, TWZ^{+}20, TC20a, TLZ20, UH24, UUS21, UKA23, VGdFCBA20, VMZDV20, VMCG20, VSM22, WS20, WHZ^{+}20, WLQX20, WJ20, WBY^{+}21, WLZF21, WZY21, WZC^{+}22, WW22a, WWQ22, WJL22, WQZD23, WHJK23, WDZL24, WZW24, WWT24, WHD^{+}23, WGLL20, WX22]. model [WY24, WYLZ24, Xia21, Xia22, XWLL20, XZ21, XW22, XCZ22, XDS^{+}20, XZS20, XKW^{+}23, XYR^{+}23, YWZ23, YCL24, YL20a, YFHC21, YWX^{+}20a, YYL^{+}23, YY21c, YTLV22, ZCS21, ZWXQ20, ZYPL20, ZZY^{+}21, ZP21, ZY21a, ZWS^{+}22, Zha22c, ZYCY22, ZYZ22b, ZST22, ZLD22, ZMJ^{+}23, ZL23a, ZLN20, ZPY22, ZGCZ22, ZS20b, ZHMZ21, ZS22c, ZLH21, ZHJ^{+}21, ZWJ^{+}24, ZCJ22, ZJ23, dFdOFdSBC20]. Model-Based [CNZ^{+}21, DKKJ21, NGC21]. Model-free [LOG20, WXF22, JMC23, LJL21, PHS^{+}20, WS20, WWT24, ZPY22, ZCJ22]. modeled [WP21b, ZT24]. Modeling [CH22b, Cro22, DD20a, DDGBGS^{+}24, FBT21, GP22, HCM21, LYKK21, NHBJ20, RYR^{+}23, RLY20, SNS^{+}20, SKWW21, TNTL21, TTMU20, AGGPO23, ABNS24a, ABNS24b, ABNS24c, CCH24, CS21b, DNR21, EE22, GE21, Hoj20, ISM^{+}23, LFY^{+}24, MF21, MS22c, NE21, ON21, PPS24, RRR19, RRR20, Wet21a, WS21, Zar21, ZZCL20, ZLZ20]. Modelling [AVJ24, HTZ^{+}23, IG21, MPRS24, RC21, VHB23, XZ23b, YXZ^{+}21, CCF24, DC21, DS24b, EKKE20, FQHS24, KMUY21, KYU22, KK24a, LFA22, MGK23, XYCG21, ZCLW22]. Models [HS23a, ACBLC24, ADF^{+}21, AARS23, BE22, CJC20, DGV22, DKLV23, DBVCP23, FM23, FTML22, GG22, GDB21, GDFNNR22, HW24, HQ22, IMOR24, JL24b, KMM^{+}24, KAS21, KTEM22, KK20, Kri22, LRCTR22, LP22, LC22a, LOS21, MMMBRS20, MFK22, MQ22, MDJW20, NPE22, NNB21, OLW23, PGR23, Pop20, RDKR22, RD22, RS24, RT20b, Shi21, SS23d, SG20b, WLG^{+}21, WSZ22, ZLW23, Zha24a, ZZ24, ZZD22, ZEA22, vdMvRS21]. modes [AAL^{+}21, AF24, BFGGM24, LSYX21, SHWW21, YZ23, ZS21b, ZCZW23]. modification [BPT23, Liu22a]. Modified [AT21b, CRTVG^{+}23, WYD20, WMT22, AD21a, AGGPO23, AHL21, BKU20, CM20a, CZL20, CH20b, CZW24, CPLH20, DZL^{+}24, FB20, LYKK21, LM20d, OPS20, WM24b, XBC22a, YYL^{+}23, ZHYD20, ZWW23, ZZW20]. modifying [PNOdV20]. Modular [AR21b]. modulated [LOS21]. modulation [AKH^{+}22, ZZZ^{+}23]. modulo [TITN20]. Modulus [ZZLV23, ZP20, Zha21b, ZV20, ZV21]. Modulus-based [ZZLV23, ZP20, Zha21b, ZV20, ZV21]. Moisil [MP23]. Moisil-Teodorescu [MP23]. molecular [DS22c, HXF23, LSC^{+}21, VS24]. mollification [Lee23]. Moment [BE22, LY22c, SV21b, TY22, KPG22, Liu21, ZM21b, ZZS^{+}20]. Moments [PDK23, dJCGVM24, CRDD21, DVS21, GGY22b, HD22a, LX23a, SMR22, WHM^{+}21a, WHM^{+}21b]. momentum [HD23, YZLS20]. momentum-conserving [YZLS20]. monitored [HL22]. monitoring [YY24]. monoclinic [PS22]. monocytogenes [BCLNV20]. Monod [AKR21, ZY21a]. Monod-Haldane [ZY21a]. monodomain [CK20]. monodromy [HR20b]. Monostatic [KL22]. Monotone [GGY23, RT20a]. monotonic [BMMV23]. Monotonicity [NRP22, AOSO20, FZ20a, LH20b]. Monte [NKA20, TS23, VS24]. monthly [JC22b]. MooAFEM [IP23]. MOOD [LTB23, MDT22, TKF23]. MOOD-like [LTB23]. Moore [AACI21, BT21, BSST20, BF23, CJ20, Hua24, SRGS20, Xu20, ZYW21b, ZMS23b]. Moran [GWZ^{+}20, ZSJ23a]. morbidities [DCR^{+}20]. Moreau [FZB24]. morphing [AF24]. morphology [RHN^{+}24, RSZA23]. Morris [BSM23]. mortality [LWZ22d]. mortar [AP20]. Mosquito [FAEZ24, DCR^{+}20, LFA22]. mosquito-borne [DCR^{+}20]. most [CCHZ23, CRS21, FLL^{+}20, SC20c]. Mostar [AD21b, BDTP23, DL21a, LD23b]. motile [SKK^{+}21]. motion [BS22c, DDLW21, FJV21, KXW24, MYY21, MCD21, MJW21, PBKS23, WDZL24, YHL^{+}20, ZZW^{+}21c]. Motions [HANP23, LXK23]. movement [SWW21b]. movements [ANG^{+}24]. Moving [BS20, REC^{+}23, BDRM23, CH24a, GNSZ22, KR20, LSV21, MSN21, NS22, NWM23, TGDPS22, Vyn23]. moving-boundary [NWM23]. MP [MZS24]. MPD [MZS24]. MPRM [xHhPjW^{+}21]. MPS [ZDKC20]. Multi [GLP20, JCH^{+}20, KG21, LXH24, PMP21, Qi22, XZ23b, XWM23, YFW21, ZHLZ21, ZBFW22, AAS23, ASV22, BCLNV20, BSRV23, BRTR21, BZS21, CZLH20, CDZW21, CZ22b, CM20b, CWZ20, CYL^{+}22, CLC^{+}22, CTF24, CFWL24, CDM21, CDL^{+}20, DXS^{+}20, DFG22, DWZZ22, ER22, FJLG22, GZCS23, GWZZ23, GS24, GZ20b, GSL^{+}22, GHS^{+}21, GZLY24, HQS^{+}24, HP23, Hoj20, HWL24, HLW^{+}20, HL23b, JLP23, JWC20, JDLZ21, KKPS21, KNN20, LJL21, LZWC22, LDXZ23, LHZ^{+}23, LDCZ24, LNC24, LLYQ20, LSWS21, LZX^{+}22b, LZY^{+}23b, LXK23, LLW24, LLQW24, LY20b, MTFD24, MK21a, MPC^{+}24, MYY21, NS22, OKN^{+}22, ÖY21, PHS^{+}20, PGR23, QYLW24, QXG21, QSW23, QCW24, RYG21, RJHC24, SCLD21, SL20a, SZG21, SGX21, SMTT23, SL20d, SZXL24, SP23, TLJO20, UU21, UKA23, WHH^{+}20, WW20, WS20, WLQX20, WCGW20, WLPZ22, WDTF22, WLS22, WC23, WZW23, WY23, WJDL24, WYLQ24, WZXH22, WWW24b, WHJ^{+}24, Xia21, XLZ23, XWWC21, XLYA21]. multi [XWL^{+}22, XJY24, XLZ^{+}24b, XZSS21, YLC20a, YY20a, YXMW22, YDXH22, YXW23, YMYR21, YK22, ZL20d, ZZCW21, ZDZS21, ZZLL21, ZCZ22, Zha22b, ZDC22, ZG24, ZLZ^{+}21, ZPY22, ZZY24, ZQC24, ZL21c, ZZW^{+}21c, ZWY^{+}22, ZWZS22, ZD24, ZPW22, dlCJ20, XDS^{+}20]. Multi-Agent [GLP20, XZ23b, YFW21, BRTR21, CZLH20, CDZW21, CZ22b, CM20b, CYL^{+}22, CLC^{+}22, CTF24, DXS^{+}20, DWZZ22, FJLG22, GZCS23, GWZZ23, GS24, GHS^{+}21, GZLY24, HQS^{+}24, HLW^{+}20, JLP23, KKPS21, LJL21, LZWC22, LDXZ23, LDCZ24, LNC24, LSWS21, LXK23, LY20b, MTFD24, MPC^{+}24, MYY21, PHS^{+}20, QYLW24, SCLD21, SZG21, SGX21, UU21, WHH^{+}20, WS20, WLQX20, WDTF22, WLS22, WC23, WZW23, WJDL24, WYLQ24, WZXH22, WWW24b, XLZ23, XWL^{+}22, XLZ^{+}24b, YLC20a, YY20a, YXMW22, YDXH22, YXW23, YMYR21, ZL20d, ZZCW21, ZDZS21, ZZLL21, ZCZ22, ZDC22, ZLZ^{+}21, ZPY22, ZZW^{+}21c, ZWY^{+}22, ZD24, ZPW22]. multi-area [GSL^{+}22, LHZ^{+}23, LLQW24]. multi-asset [KNN20]. Multi-attribute [ZHLZ21]. multi-axis [XJY24, ZZW^{+}21c]. multi-currency [ASV22]. multi-cyber [ZWZS22]. Multi-Delay-Boundary [XWM23]. multi-delays [XLYA21, ZG24]. multi-dimensional [AAS23, NS22, SMTT23, YK22, ZQC24, ZL21c, dlCJ20]. multi-domain [LZY^{+}23b]. multi-exits [RYG21]. Multi-fidelity [KG21]. multi-fluid [PGR23]. multi-followers [WY23]. multi-fractional [ER22]. multi-games [SL20d]. multi-group [LDXZ23]. Multi-information [LXH24]. multi-interval [WLPZ22]. multi-item [CDL^{+}20]. multi-layer [MK21a, RJHC24]. multi-level [WHJ^{+}24]. multi-link [HWL24]. multi-links [CFWL24]. multi-material [BSRV23]. multi-medium [HL23b]. multi-memory [ZZY24]. multi-microgrid [LZX^{+}22b]. multi-objective [GZ20b, Hoj20]. Multi-parameter [PMP21]. Multi-period [ZBFW22, CDM21]. multi-player [QSW23, QCW24, Xia21]. multi-potential [LLW24]. multi-product [LLYQ20]. multi-proportional [WCGW20]. multi-rate [JDLZ21]. multi-resolution [OKN^{+}22]. Multi-scale [JCH^{+}20, SL20a]. multi-server [SZXL24]. multi-species [BCLNV20, SP23]. multi-stage [LLYQ20]. Multi-stencils [Qi22]. multi-step [TLJO20]. multi-strain [XZSS21]. multi-strategy [CWZ20]. multi-symplectic [UKA23, Zha22b]. multi-term [BZS21, DFG22, QXG21, WW20]. multi-time [HP23]. Multi-UAV [XDS^{+}20]. multi-unmanned [XWWC21]. multi-wave [ÖY21]. multi-weights [JWC20]. multiagent [DYX24, GZWX23, GL24, HJW^{+}23, JWL^{+}24, PGYW23, TLHH21, YP22, ZCLZ22]. multibody [AKY21]. Multicolor [ZWC23]. Multicolorful [JGW21]. multicomponent [TKF23]. multidelay [JHO^{+}21]. multidimensional [AR20, DS21d, Jor21c, Kut20, PT21]. multidirectional [LFLH23]. multifactor [LY21b]. multigame [BY23b, SWX24a, WZW^{+}21]. multigraph [LLZL22]. multigraphs [GW23a]. multigrid [BM21, CSC20, Han20, He23, IZ20, PSXX21, YXAY21, Zha21b]. multilayer [CYW21, GDB21, GL22, LSYX21, ME21a, SWP^{+}22, SRKA23, YWR^{+}24]. Multilevel [LS20]. multilinear [AOSO20, FKS23]. multiobjective [CYZ23]. multipartite [CY21b, RD23, WL22]. multiphase [ALL23]. multiphysics [HREB21]. multiplayer [XTS^{+}22]. Multiple [AT21b, HD20, SWZS20, SLPH22, TGDPS22, ZH20a, ZKLL23, AB20a, ASS^{+}23b, CJCL21, CXY^{+}23, DHS23, DS24a, EW21, Fan21, FZW23, FCH21, GY22a, GLTY22, GW23b, xHhPjW^{+}21, HWY^{+}22, HCBY21, HZ23a, HB24, JDZH22, JQJY22, LKP22, LSWL20, LYN^{+}21, LHX^{+}22, LTL^{+}22, LYhH^{+}24, LDZT24, LGC^{+}24a, LQ24, MMHD20, OG20, PGYW23, RB21, SN21, SSCW22, SLS24, SZWL21, SXS22, TDZ^{+}24, TTC^{+}23, VJA^{+}22, WCZ20, Wan21a, WGZZ22, WDZG22, WJDL24, WPE20, XYLF21, YLPC22, YSW^{+}21, YYW^{+}23, Yu21, ZDGL21, ZZL^{+}20a, ZYZ22c, ZDD^{+}22]. multiple-ferrofluid-droplet [LYN^{+}21]. multiple-integral [LKP22]. multiples [CHY22b, ZQZ21]. multiplex [CTV21, HPC24, JL24b, Tak21, TVY21, WMCZ21, WB20, WB23]. multiplexed [YTZ20]. multiplication [HLZ21]. Multiplicative [ÖC24, DWZZ22, WC23]. multiplicities [Wan22c]. multiplicity [LZ22d]. multiplier [AFGG21]. multipliers [Dol21, KS21]. multipoint [AKY21]. multipole [RKRRL21]. multiquadric [ADS23, Sun22]. multirate [LCZW23]. Multiscale [AP20, AMPS20, CCE^{+}22, CEPZ22, LHS21, MMDD21, RMSP22, YY21b, ZS20a]. multisplitting [GZLZ24, ZZLV23]. Multistability [YLCL23, FBT21, YWS^{+}20]. multistage [PMAK22, RT20b]. multistate [CL20]. Multistep [Wan22a, ZZL22]. Multivalue [CDDP21]. multivalued [AB21b, OK21]. Multivariable [NPJ^{+}20, DXZ^{+}23]. Multivariate [DMPP21, ARMM23, CKW22, HYS24, KS21, ZZD22]. multiwavelets [ACH^{+}22]. Müntz [DEHFF24]. mutation [FHFS24, Liu22b]. mutations [RS23]. muti [CMYF21]. muti-agent [CMYF21]. Mutiple [URLC21]. Mutual [Di 22, BY24, CSK23, MZ24, Ued23]. mutual-visibility [BY24]. mutualism [GLWW22, LB20, RK22]. mutually [TTMU20]. Mycielskian [HLLH23]. myocardial [SG20b]. myopic [LCS23].

N [Wet21b]. N-player [Wet21b]. nabla [WJC^{+}20]. Nagumo [AFGG20, HGJ21, LX22, SY21a, ZS20b]. Nakazawa [CLZY21]. naming [JG21]. nano [CSG^{+}22, CSG^{+}23]. nano-material [CSG^{+}22, CSG^{+}23]. nanobeam [LD21a]. nanobeams [BP23, TNTL21]. nanocontact [BM22]. nanofluid [Ach21, FWM^{+}22, IMIH22, KRS20, KHA22, MS22c, SN21, SKK^{+}21, TMA^{+}22, Yu21]. Nanofluids [Hoj20, AKH^{+}22, AAHC21]. nanoliquid [MM22a, MM24]. nanoparticles [AGC20]. nanoplates [NTLNX24]. nanosheet [SRK^{+}21]. nanostructures [SDD22]. nanotori [ADKS24]. nanotubes [WL20]. naphtalenic [SRK^{+}21]. narrow [AS22]. narrow-gap [AS22]. Nash [BG20b, CLSN24, GH21, HY24, KLP22]. Nash-type [BG20b]. natural [ABNS24a, ABNS24b, ABNS24c, DXS^{+}20, KTEM22]. Navier [CHY22a, DG21, Azi21, BFRS24, BDRM23, FKS23, HYÇK22, JLJ23, MCT^{+}22, MVO21, SLZ20a, YGM^{+}22, ZH22a, ZRL^{+}24, ZZX21, ZS22b]. navigability [Est21]. navigation [AR20]. NCSs [DS24a]. ND [CYG20]. ND-agent [CYG20]. Near [YG24, GP23, Ma23, SIK22]. nearest [HZT^{+}21]. nearly [TCPF24, TPSP24]. Necessary [GGBEZRI23, Wu20a]. negative [CHG23, WHL21, ZM21a]. negative-determination [CHG23, WHL21]. negativity [DX24]. negotiation [YSYS24]. Neighbor [DG20b, LZ20c, ZLC^{+}24b]. neighborhood [AS21c, Zho20]. neighbors [LZD23]. neighbour [DDPS22]. nerve [SBT22]. nested [LAS23, ZXZ24]. network [CC22, CHHC24, CZZP24, CZGX22, DP24, DGL^{+}23, DL20c, GGL^{+}23, GPRV21, GLW22, HWC23, HZCY23, HXL^{+}20, HKCK24, HPC24, HGJ21, KKR21, KRF20, LCD^{+}20, LJW^{+}21, LH22, LCJ^{+}22, LM23, LQS^{+}23, LCLH24, LW20b, LWG20, LSWS21, LZLZ21, LM21, LWZ22c, LMS23, LBJ22, ML22, MT21, NYHC24, PGZ^{+}24, PPS23, QY22, QSL21, SG20a, She20, SWP^{+}22, SWX^{+}21, SRKA23, SMM^{+}24, SYZ^{+}23, TLZ20, TPTA24, VJKZ21, WGW^{+}21, WLJ^{+}21, WZLH22, WDW^{+}22, WZW^{+}21, WGJK21, YLW20, YWJ^{+}20b, YL20b, YWJ23, ZCZ22, ZTLW24, ZJW^{+}22, ZGZ^{+}23, ZS20b, ZF23b, ZTW22, ZCZW23]. network-based [CZGX22, HWC23, LZLZ21, PGZ^{+}24, ZJW^{+}22]. network-organized [TLZ20]. networked [CSS^{+}22, CZK^{+}22, DZL^{+}22, DW22, EB22, FSS20, GH21, HTF20, HDP^{+}21, JTW^{+}23, KJL20, LLY22a, LTL^{+}21, MCD22, MZY^{+}22, MP22, NSSA20, NZ20, PGYW23, PGZ^{+}24, SC20b, SLPH22, WZXY20, WFYL21, WCXS22, XZQX20, YYG20, YWG24, YL20a, ZDGL21, ZZL^{+}20a, ZCW^{+}24a, ZCJ22]. Networks [HHY20, LWL^{+}23c, XWM23, AAJ23, AFL24, ABNS24a, ABNS24b, ABNS24c, ADF^{+}21, ANS^{+}20, AA20, ADV23, AKP23, BZ23a, BBH22, BZZ21, BY23b, BS23, CHG23, CH24b, CFWZ20, CL20, CYW21, CYHR22, CWB22, CHH22, CLLA23, CXYZ23, CL24a, CLC24, CFWL24, CLH24b, CWL^{+}21, CLLA22, CLD^{+}24, CTV21, CRTV22, DLL^{+}23, DZ21, DTWJ21, DS24b, DWZ20, DZZ^{+}20, DZZ^{+}21, DL21c, DZX23, Est21, Fan21, FJFS21, FSS20, FZS21, FW24, GZ20a, GWY20, GXZZ20, GL22, GX23, Guo23, GDS24, HKKL22, HD23, HQZC24, HLL^{+}21, HP23, HHW23, HLD^{+}23, HZ23a, HHZZ20, HSZ20, HCH24, HWL24, HB24, HXWC24, JW23, JLD^{+}24, JWC20, JL24b, KSH22, KSZ23, KR22, KK23b, Kun24, LFLL21, LPP21a, LPP21b, LPKC22, LZW^{+}20a, LZH^{+}20, LGL20, LLL21, LWD21, LXF21, LZZ21, LKSS21, LHZ^{+}21a, LCXM21, LDXZ23, LLLL23, LDZT24, LXH24, LLG^{+}24, LCC^{+}24, LCY^{+}21, LZZ^{+}20, LXW^{+}20, LK21b, LHFW21, LSL^{+}21, LSYX21, LGD^{+}21, LZH21, LJHL21, LX24b]. networks [MZ24, MAAEF21, MMHX20, MCSD21, MS24, NPJ^{+}20, NX21, NGS22, PP21, PPZW20, PLQ^{+}22, PMAK22, PN24, QNX^{+}23, QLZL22, QDL^{+}21, QMS22, RSVL21, RM23a, RM23b, RH20, SSKS21, SPX20, SS24d, SNR^{+}22, Sha22, SZCC23, SJ23a, SH22, SWP^{+}22, SWX22, SY21a, SGJ^{+}21, SM22, SL20d, SCG^{+}20, SSW21, SLL^{+}22, SHYS23, Tak21, TKY23, TWL20, TWB22, TQZ^{+}24, TAU^{+}22, TW21b, TTC^{+}23, TVY21, TXZ24, URLC21, WCG^{+}20, WCGW20, WXW^{+}21, WMWS21, WMCZ21, WGL^{+}21, WWWH21, WWWZ21, WHL21, WW22a, WMS^{+}22, WYLL22, WLL^{+}22, WMT22, WLFX22, WCWZ23, WLHW23, WZPL24, WZLC21, WLZD24, WB20, WX21a, WCZ^{+}22b, WCX^{+}22, WB23, XLJL22, XBC22a, XWLL20, XYH^{+}24, XLYA21, XYLF21, XLZ^{+}24a, Xu24a, YGZ^{+}22, YLLL20, YTGC20, YLPC22, YLHZ22, YSLK23, YZM24, YMW24, YWR^{+}24, YWS^{+}20, YXXZ23, YWX^{+}20a, YFM^{+}24, YXX^{+}22, YDSW22, YLCL23, YL23b, YTX^{+}24, ZYC^{+}24, ZeFWW20, ZSC^{+}21, ZLLL21, ZP21, ZYZ22c, ZLK^{+}22, ZYZ22a, ZCS^{+}22]. networks [ZB23, ZZPF23, ZC23a, ZH23a, ZeFWD23, ZLZ^{+}23a, ZQ24, ZWFW22, ZW22b, ZKLL23, ZPLL20, ZWHY20, ZWYL22, ZCZW23, ZLC^{+}24b, ZLZ20, ZLW^{+}21b, ZeF21, ZB22, ZJ23, ZT24, ZLLX22, ACBLC24]. networks-based [SJ23a]. Neumann [HD22b, KM20]. Neural [ACBLC24, CHHC24, CZGX22, HHY20, PGZ^{+}24, QSL21, SS24d, WSSX22, XWM23, ZJW^{+}22, ZTW22, AAJ23, AFL24, ANS^{+}20, ADV23, BZZ21, CHG23, CH24b, CFWZ20, CYHR22, CYL^{+}22, CWB22, CHH22, CXYZ23, CL24a, CLC24, CZZP24, DP24, DWZ20, DZZ^{+}20, DZZ^{+}21, DL21c, FJFS21, FSS20, FZS21, GWY20, GX23, HKKL22, HD23, HP23, HZ23a, HHZZ20, HSZ20, HZX22, HWL24, HB24, JW23, LFLL21, LPP21a, LPP21b, LPKC22, LZH^{+}20, LGL20, LWZ21, LKSS21, LHZ^{+}21a, LCXM21, LH22, LM23, LCLH24, LW20b, LCY^{+}21, LXW^{+}20, LHFW21, LSYX21, LWZ22c, LBJ22, LZH21, MAAEF21, PP21, PN24, QNX^{+}23, QLZL22, QY22, RSVL21, RM23a, RM23b, RH20, SSKS21, SZCC23, SJ23a, SWX^{+}21, SWX22, SY21a, SRKA23, SMM^{+}24, SCG^{+}20, SSW21, SHYS23, TWL20, TWB22, TQZ^{+}24, TW21b, TPTA24, TXZ24, URLC21, VJKZ21, WCG^{+}20, WCGW20, WXW^{+}21, WZZX21, WGL^{+}21, WWWH21, WWWZ21, WHL21, WLHW23, WM24a, WZLC21, WX21a, WGJK21]. neural [WZXH22, WCX^{+}22, XLJL22, XLYA21, XYLF21, YGZ^{+}22, YMW24, YWR^{+}24, YWS^{+}20, YDSW22, YLCL23, YL23b, ZYC^{+}24, ZSC^{+}21, ZYZ22c, ZCZ22, ZLK^{+}22, ZCS^{+}22, ZC23a, ZLZ^{+}23a, ZQ24, ZW22b, ZF23b, ZWYL22, ZB22, ZLLX22]. Neural-based [WSSX22]. neural-network [TPTA24]. Neuroadaptive [DS24a]. Neuron [BY21, MUMS22, RDKR22, WHM22, XGRM20, XLYA21]. neuronal [HGJ21]. neurons [BSM23, LPS23, LSYX21]. neutral [AFA24, CZ21a, CLZ^{+}22, CL24b, DL20b, FLWL21, FZ20b, GNA22a, GNA22b, JMM20, JS20, JPZ^{+}21, KGPS22, LLD21, LZZ24b, LZ24e, RM23a, SCT^{+}20, SZZL22, WHC20a, WGL^{+}21, WZGL21]. neutral-type [LLD21, WGL^{+}21]. neutron [XYM22]. neutrophils [XZS20]. Newborn [Kab24]. Newell [DYA21]. Newell-Segur [DYA21]. news [LKK^{+}23]. Newton [ACGG20, AHEGM23, CB21, CWY^{+}20, GGY23, KKS22b, MPM23, MVO21, MN23, Sin20b, TZD22, XYKL20, ZMS23a, dSTV21]. Newton-adaptivity [ZMS23a]. Newton-like [GGY23, Sin20b]. Newton-type [dSTV21]. Newtonian [PPS24, RRR20, Azi21, EFJO20, LGK^{+}22, MMS21, PM20, RRR19, SNZC20, ZM20]. Nicolson [ZL22a]. NIPG [MZ23a]. NIST [LOVHC23]. nitrogen [LNO23]. NiZnFe [KHA22]. NN [GZLY24]. NNs [ZCD23]. NNs-observer-based [ZCD23]. No [AAA20, Gan23, WWY24]. No-wait [AAA20]. Nodal [MV22, ZZX21]. Node [CDLR24, WLZD24, ZWJ^{+}20]. Node-bound [CDLR24]. nodes [DMPP21, DTWJ21, GGL^{+}23, JLZ^{+}22, OT20, TVY21]. Noether [SC20a, TZ21]. Noise [FFL^{+}23b, FZW23, JQJY22, MXY^{+}20, BMMV23, CFWL24, DRS20b, DHM^{+}20, ER22, HX22, HLM23, HWWW23, KT21a, LZ22c, LWX24, LY23c, RX20, SGPZ22, WC23, WWW24a, YZC^{+}22, YDSW22, YD22b, Zhu24]. Noise-induced [FFL^{+}23b, Zhu24]. Noise-input-to-state [JQJY22]. Noise-to-state [FZW23]. noises [DWZZ22, SHWW21, WHC20a, YM22b, ZL22c]. noisy [ES23, WLB22, ZIML20]. Non [CZR21, DL20c, FMT^{+}21, FC21, JWC20, LHZ^{+}21a, LL22b, LLWK23, MP20, NNN24, PWSE21, PN24, jSZyW^{+}20, SSW21, UGG^{+}20, WWWZ21, WSXL22, ZZQ^{+}20, ZC23a, ZC21b, AS20, ABNS24a, ABNS24b, ABNS24c, AR21b, AEL21, AMF20, ADB20, AG20, Azi21, BG20b, BS23, CL22, CLC^{+}22, CWY^{+}24a, CLC20, DHR22, DDR21, DX24, DS22b, DHS23, DKLV23, DL20a, EBGN23, EKKE20, EW21, ECS21, FLZ23, FLC^{+}22, FMN21, FMHDB22, FW24, GI23, GTK^{+}22, GNA22a, GNA22b, GW24, HSM21, HVYMS21, HCBY21, HD22b, IS21, JY20, KKPS21, KZWS22, KMD21, Kum23, LPK22, LDX20, LKSS21, LLY22a, LR24, LDCZ24, LFLH23, LHZ^{+}21b, LMZ^{+}21, LGK^{+}22, LNC22, LCL22c, LF24, LU24, MS22a, MHAT21, MM20a, MMS21, MCD21, NP20, NSSA20, PPS24, PJMW22, PNJ^{+}21, PD23, PDK23, PKG^{+}20, RRR19, RRR20, RC22, RBPN20, SN21, SCD23, SRZK22, SS22a, SG24, SWS21]. non [SYW21, SWW22, SKK^{+}21, SYX22, SYTC24, TCZ22, TFL22, TMA^{+}22, TITN20, TTMU20, TWB23, VBMB21, VSM22, WW20, WJC^{+}20, WLZF21, Wan21a, WM23, WX21b, WZXH22, WX22, XZeF20, XYM22, XW22, XTS^{+}22, XXW^{+}20, YLC20a, YLS23, YDXH22, YF21a, ZLX20, ZCJL21, ZDC22, ZPDL20, ZSYX23, ZPLL20, ZyZW21, ZZY24, ZWY^{+}22]. non- [DKLV23]. Non-adjacent [jSZyW^{+}20]. non-affine [LDX20, YDXH22, ZPDL20]. non-anticipative [PKG^{+}20]. Non-Archimedean [FC21]. non-autonomous [AMF20, FLZ23, FLC^{+}22, Kum23, LKSS21, SCD23]. non-Bayesian [LF24]. non-circulant [PNJ^{+}21]. non-commensurate [ZLX20]. non-compact [SYX22]. non-convex [LFLH23, LHZ^{+}21b, LMZ^{+}21]. non-cooperative [CWY^{+}24a]. non-crossing [SS22a]. non-Darcy [SN21, ZyZW21]. non-delayed [DHS23]. non-equilibrium [PJMW22, TTMU20]. Non-fragile [JWC20, LL22b, LLWK23, PN24, SSW21, ZZQ^{+}20, ZC21b, CLC^{+}22, KKPS21, LR24, NSSA20, VSM22, WLZF21, WM23, XZeF20, XXW^{+}20]. non-Gaussian [GTK^{+}22]. non-globally [LNC22, LCL22c]. Non-hamiltonian [FMT^{+}21]. non-Hermitian [SWS21, WX21b]. non-holonomic [HCBY21]. non-homogeneous [HD22b, SRZK22]. non-hydrostatic [EBGN23]. non-identical [ZPLL20]. non-inertial [MCD21]. non-infinitely [CL22]. non-instantaneous [FW24, GW24, KMD21, Kum23]. Non-integrability [MP20]. non-integral [ZSYX23]. non-isothermal [AR21b]. non-iterative [ABNS24a, ABNS24b, ABNS24c, Wan21a]. Non-linear [NNN24, UGG^{+}20, AEL21, CLC20, DDR21, DS22b, GNA22a, GNA22b, HSM21, IS21, LU24, MS22a, MHAT21, NP20, PDK23, RBPN20, SG24, SKK^{+}21, XW22, ZWY^{+}22]. non-linearly [YLC20a]. non-Lipschitz [WJC^{+}20]. non-Lipschitzian [LLY22a]. Non-local [CZR21, BS23, EKKE20, WX22]. non-negativity-preserving [DX24]. non-Newtonian [PPS24, RRR20, Azi21, LGK^{+}22, MMS21, RRR19]. non-orthogonal [FMN21]. non-oscillatory [PD23, TFL22, YF21a]. Non-PDC [DL20c, LPK22]. non-permutable [EW21]. non-potential [BG20b]. non-prismatic [ECS21]. Non-reduced [WWWZ21, ZC23a]. non-regular [AG20]. non-residues [TITN20]. non-selfadjoint [MM20a]. non-separable [HVYMS21, YLS23]. Non-separation [LHZ^{+}21a]. non-singular [ADB20]. non-smooth [KZWS22, LDCZ24, RC22, WW20, XYM22, ZDC22]. non-spatial [FMHDB22]. non-staggered [DL20a]. non-stationary [JY20, SWW22]. non-stiff [DHR22]. Non-Strict [WSXL22, SYTC24, WZXH22]. non-submodular [SYW21]. non-symmetric [TCZ22]. non-uniform [AS20, GI23, LU24, TMA^{+}22, ZZY24]. non-zero [XTS^{+}22]. non-zero-sum [ZCJL21]. nonaffine [WS23b]. Nonautonomous [TCZ^{+}20, WYD20, SS20, WYLQ24]. noncanonical [PMTT24]. noncausal [CSS^{+}24]. Nonclassical [SS23b]. Nonconforming [HKL22b, ZSL20, PZS24, SL20c, SZ24a, SY22b, XZZ22a]. nonconvex [BZ21, TLPT21, WNZ21, WSG^{+}21, YMYR21, ZH23b]. noncooperative [HY24]. nondegenerate [BBH22]. nondeterministic [LLWY22, ZXC20a]. nonexpansive [OK21]. Nonfragile [ZXZ20]. nonholonomic [YP21, ZDD^{+}22]. nonhomogeneous [LPK22, LZRF21, NK21a, WCZ^{+}21, ZC21a]. nonideal [ZCW^{+}24a]. nonidentical [LSYX21, ZYXL23]. Nonlinear [AHM23, ADLU23, CFP22, DSZ^{+}23, DH24b, EM20, FMHDB22, GLP20, JTC^{+}24, KSS23, LLS21b, MXY^{+}20, TM22b, WWW^{+}21, WSXL22, YXYZ23, ZL20d, ZCL^{+}23, ACC^{+}21, AM20b, AR23, ASS^{+}23a, AHB20, AP21, AHEGM23, AP20, ACH^{+}22, BSdCCGCV21, Bab22, BR20, BG20a, BM20, BM24, CDZW21, CM20a, CZ22a, CG20a, CJ22, CWY^{+}20, CNZ^{+}21, CYL^{+}22, CLC^{+}22, CZK^{+}22, CC22, CLWX23, CCL^{+}24, CHHC24, CFWL24, CSL24, CZ20, CZXyS24, CW22, CSC20, CZWX21, CXGL22, CZGX22, DX24, DYA21, DRS20b, DS24a, DMP21, ER22, FT24, FWM^{+}22, FLWL21, FZLJ24, FLR23, FZS21, FLLW23, GSWL23, GSS24, GTWL21, GS21, GS24, GM21b, GLTY22, GYC24, GMN22, GHS^{+}21, GZWX23, GGY23, GL24, HC20, HREB21, HLW^{+}21, HCTS23, HLL^{+}21, HVYMS21, HR23, HL24b, HLKC22, HL23a, HZX22, HYT21, HZXT24, HJY24, HMZZ20, IP23, IZ20, IR22, JC22a, JA23, JY21, JDLZ21]. nonlinear [JDZH22, JHO^{+}21, JCH^{+}20, JCSH24, JQJY22, JOR^{+}21a, JW21, JMC23, JWL^{+}24, JJSY22, KT21a, KMM^{+}24, KZC23, Kar22, KMUY21, KAS21, KBL^{+}24, KÖY24, Kod21, KKS22b, Kud20b, KL24, KK24b, LLZ22a, LPS21, LM20b, LDX20, LWZ21, LTW22, LGX22, LLG22, LWNZ22, LWL22, LWL^{+}23a, LD23a, LM24, LGC24b, LZZL21, LT22, LCY^{+}21, LTL^{+}21, LZLZ21, LWH22, LFM^{+}22, LZZ^{+}23, LH24b, LZYL21, LY22d, MPX^{+}20, MCD22, MQW23, Ma24, MXF^{+}24, MTRA22, ME21a, MS20, MMHD20, MKP20, MZFX20, MZLS22, MM20b, MV22, NE21, NL20, Ndo23, NTT21, PZS24, PCN^{+}24, PWZ22, PX23, PKG^{+}20, PMTT24, QCARL20, QWD22, RP20, RKJ20, RX20, RFFZ21, RY24, RG22, SCLD21, SM21b, SDD22, SJ23a, SMH^{+}20, SL20c, SW20, SZ23, SX24, SZ24a, SV23a, SWF20, SKL23, STA24, SPZS20, SLZ^{+}20b, SZZL22, SYX22, SLTL20, Sug21, SYTC24, SZWL21, SMR22, TLHH21, TlS^{+}22, TLC22, TZ24, TTCY23]. nonlinear [TO20, THBPMC22, TPTA24, VPK20, Voe23, WHH^{+}20, WLQX20, WHLL20, WXLZ21, WXW^{+}21, WZZX21, WSS22, WZLH22, WCW22, WLPZ22, WDTF22, WXLY23, WS23b, WYLQ24, WM24a, Wen20, WH21, WPX^{+}20, WHL22, WSSX22, WWW24b, XD21, XWMC20, XH22a, Xy20, XSP22, XHL23, XZZW21, XCSW21, XLLH22, XWL23, XZQ23, XLZ^{+}24a, XLZ^{+}24b, YSHV^{+}23, YLCZ20, YY20a, YLG21, YXMW22, YWL20, YJ20, YGHW21, YTZ22, Yi23, YTZ20, YWJ20a, Yoo21, YP21, YP22, YYWL23, YZDC22, YZY22b, YZ24, ZT22a, ZJG21, ZWH23, Zha20, ZP20, ZSL20, ZSM20, ZZCW21, ZZY^{+}21, ZMD^{+}21, Zha21b, Zha21a, ZHG22, ZZ22a, ZXZC22, ZW22a, ZD23, ZWW23, ZYXL23, ZJJ24, ZO21, ZPDL20, ZZX20b, ZZW21b, ZYX21, ZJW^{+}22, ZCLZ22, ZPY22, ZYySX23, ZY23b, ZXZ20, ZZ20b, ZWHY20, ZDX21, ZC21c, ZZD22, ZL22e, ZJLT23]. nonlinearities [CHW^{+}24, CZ20, DL20b, WJC^{+}20, WZXH22, ZZL^{+}20a]. nonlinearity [KY21, LZ24e, PBM20, WP21b, XBLW23, ZZS^{+}22, ZXB20]. nonlinearly [HLL22]. Nonlocal [IMS23, LLS^{+}22, NTLNX24, AzR22, BP23, GSW^{+}20, INOP20, JS20, LD21a, LZYL21, LYN22, LQNA23, ÖY21, TNTL21, TZJ20, TZW22, XLL^{+}21b, YWZ23, ZYY20, ZZ22a, ZZY24]. nonmonotone [ZP20]. nonnegative [BD20a, DESZ23, GH20b, LYZ22a, TNL23, XYY23]. nonoscillation [YY21a]. nonparametric [JG21]. nonpolitical [LKK^{+}23]. nonpunishing [NCY^{+}20]. Nonseparation [XYH^{+}24]. nonshifted [SC20a]. nonsingular [DWZ20, ZPSP22]. nonsmooth [AS21a, BK21, GAA^{+}24, TLJO20, ZH23b]. Nonstationary [CZ20, Mit24, EV21]. nonstrict [HL23a, LH24b, WZZX21, XZZW21, ZCLZ22]. nonstrict-feedback [HL23a, LH24b, WZZX21, XZZW21, ZCLZ22]. nonsubsampled [WWZ^{+}23]. nontoxic [ZST22]. nontoxic-phytoplankton [ZST22]. nontrivial [AACI21, Ma23]. nonuniform [DS22a, SS20]. nonuniqueness [Kwo22b]. Nonzero [BPRR22, MZLS22, TTCY23, TLY24]. Nonzero-sum [MZLS22, TLY24]. nonzero-zero [TTCY23]. Nordhaus [FKK24, HWLL21, HLX20]. Norin [CWZ22]. norm [BFLR23, DFX24, FCX21, IÖÜP23, IÖÜP24, IKÖ24, KKH23, MZ23a, SGPZ22, STA24, TZLL24, ZYySX23, dSLV20]. Normal [SS24b, XFK22, AS21a, ASUH21, AFGG20, GGY22a, GS22, LGX22, YLL22]. normal-Rayleigh [ASUH21]. normalization [MZSW21]. normalized [FKK24, GRD23, HDZ22]. normals [FKVZ22]. norms [AT24, LZ22f, TZ23b]. Note [AG20, BPT23, BS22b, GLMT21, Ili22, LZX22a, LW23b, wX20b, YS20, ADV23, BBDVPV20, BFRS24, CMT20, CWZ22, DYZC20, Far22, GPW24, Hu21, KY24, Kop22, LSCS21, NT24, Prz24b, SVW23, SM24, SW24, WLWL22, Wri24]. notice [ABNS24b, CSG^{+}23, NHK^{+}23]. notion [CS21a, MUOF22]. Novel [GHND22, He23, HD22a, PP21, TFL22, TLPT21, UHZ^{+}20, Wan23a, ZYC^{+}24, AMM21, BRTR21, CHL24, CZXyS24, CT21, ER22, FZB24, HRA^{+}20, LPP21a, LPP21b, LPKC22, LWZ22a, LZRF21, LHZ^{+}21b, LZL20, MZYW23, PHL23, PHS^{+}20, QNX^{+}23, RCG21, SMH^{+}20, SJP22, SL24, TPTA24, WGM^{+}21, WWQ22, WSZ20, WHL22, XFW^{+}21, XSP22, XHWM20, YZLS20, YLH20, YWX^{+}20a, ZIML20, ZS21a, ZYCY22, ZLZ^{+}21, ZGCZ22, ZWZ^{+}22]. November [Ano20i, Ano20t, Ano21j, Ano21v, Ano22v, Ano23j, Ano23v, Ano24v, Ano22j, Ano24j]. nozzle [GCGM21]. NSFD [RVV24]. nuclear [Ahm21, SGPZ22, Zar21]. Null [ZZX24]. nullnorms [Çay20]. Number [CF23b, Sug21, AR22, ADS21, BGHX22, BNR20, BDH^{+}20, BY24, BRMVCD21, CMT20, COPW23, DM21, DM24, DTP21, DDH^{+}24, DGO22, DS22c, DS24c, Eva20, GN21, GH20a, HK21, HPT22, HW20, JWMW23, KCJ22, KRE20, KR21, LLY20, LQL21, LW21a, LY21a, LR22, LZC21, LX24b, MW20, RT20b, SS21b, SS22b, SM24, SZ21, SZ22, SD22b, TX21, TPSP24, Wan21b, WLZ21, WS24, XT23, XW23, YD22a, YZW20, YS20, ZY24a, ZH22b, ZZB24]. numbers [CU23, DDS24, GLGLH24, GLW20, GM24, IKÖ24, KRNG24, KF23, LY22a, LCL23, LZ21, LLZ22b, MZ23b, PLH22, QLL20, QLLY21, TT20, WLL21, WZ24, ZF23a, ZWC23, ZY23a]. Numerical [AD21a, ABD22, AALMLM22, AAS23, AGC20, BSdCCGCV21, BPG24, BS22a, BG20a, CGSRG20, CFR22, CUA^{+}22, CCSPMV23, DG21, FLR23, FCH21, GCGM21, GJJC^{+}23, GLT^{+}21, GLWX24, GZRM21, GRMQ22, GR21, HC20, HWCD22, IMIH22, Kum24, LMZ20, LJAM24, LL22a, LZY20, LC22a, LLLC23, LHZ21c, LY23c, MT24, MSB22, NP20, PGR23, PW22a, QXG21, RST21, RVV24, RS21, SM21b, SKS21, SX23b, XZ23b, YL22, ZZX20a, ZRL^{+}24, ZBBK23, ARISV20, ASAK21, AFA24, BCLNV20, BR20, BBH21, BCP23, BMY24, CZL20, CKWY23, CKL22, CSC20, DdAS22, DC21, DDR21, EGMY23, EA24, EFJO20, FC21, FY23, FZ23, GdFPF24, GP23, GP22, Har24, Har22, HSM21, HT23, HL22, JNY^{+}21, JHO^{+}21, Kod21, KOP21, LPS21, LZSC22, LG22, Lu22, Luo20, MCT^{+}22, MVCK24, MPW23, MV21b, MV23, MRVC20, NI21, Oru21, Oru22, QWD22]. numerical [QXX21, RC22, RM23b, RNVC21, RG20, RJM23, SO21a, SS24c, SK20, SG20b, TCGR24, Ten22, TC20a, WLW20, Wan21b, WKM24, WH21, WZ20b, WHY22, Xy20, XGF20, YR20a, YSL24, YR20b, YJ20, YLL20, ZHYD20, Zha22a, ZG24, ZY21b, ZLU21]. numerics [DBVCP23]. nutrient [CD23, SCD23]. nutrient-limiting [CD23]. nutritional [ZZCL20]. Nyquist [ZLX20]. Nyquist-based [ZLX20]. Nyström [BMJ20, FRRS24].

O [AGC20, KHA22, Kop22]. object

[IP23, LTW24, MFK22]. object-oriented [MFK22]. objective [GZ20b, HH20, Hoj20]. objectives [AFG22a, AFG22b]. Oblique [MMS23, DHM23]. Obrechkoff [SK20]. observable [CL22, GYC^{+}22, GMH^{+}20, GTWL21, ZY24b]. observation [CFWL24, TZ23a, XZ21, XZZ22b]. observations [GJS^{+}23, LHS21, QQWJ22]. observed [ZS22d]. Observer [AS23a, CL22, CJ22, CZK^{+}22, GLP20, GX23, JZS^{+}24, LWY^{+}24, MCD22, MR22, MNSR22, MC22, NK20, SWSW21, WXP21, WYZX22, YM20, ZQP^{+}22, ZCLZ22, ZWYL22, ZD24, CWZL23, CHZ^{+}20, CSL24, CLH24b, DXZ^{+}23, FMHDB22, FLH^{+}21, GXZ^{+}24, HDP^{+}21, HLW^{+}21, HSL^{+}24, JY21, LPK22, LCZW23, LZHZ24, LWG20, LCL22b, LCL22a, MXF^{+}24, MP22, MZYW23, NK21a, SNRA20, SM21c, TPTA24, VSM^{+}21, WGZZ22, WCZ^{+}22b, YLCZ20, ZZQ^{+}20, ZLS21, ZHG22, ZCD23, ZLZ24b, ZXZ20, ZTW22, ZQC24]. Observer-Based [GLP20, AS23a, CL22, CJ22, CZK^{+}22, GX23, JZS^{+}24, LWY^{+}24, MCD22, MR22, MC22, NK20, SWSW21, WXP21, WYZX22, YM20, ZQP^{+}22, ZCLZ22, ZWYL22, ZD24, DXZ^{+}23, FLH^{+}21, LPK22, LCL22b, MXF^{+}24, NK21a, TPTA24, WGZZ22, WCZ^{+}22b, ZZQ^{+}20, ZLS21, ZLZ24b, ZTW22]. observers [HCX^{+}24, ZWZ^{+}21]. obstacle [BM21]. obstacles [LLWY22]. obstetric [NS20]. obtain [PZ20]. obtained [CET22]. occurrence [RMCG20]. occurring [CZ20, KMLW22, KGCC20, KKPS23, LY22d, WB23]. Ocean [Boc22, CT21]. OCT [CCG^{+}22]. Octagonal [HLP24, QHC21]. October [Ano20j, Ano20u, Ano21k, Ano21w, Ano22k, Ano22w, Ano23k, Ano23w, Ano24k, Ano24w]. octree [dSLP20]. odd [CDD23, JX24, PWL23, SYL23, YML21, ZD24]. odd- [PWL23]. odd-order [JX24]. ODE [BCP23, CCF24, LJ23, MNSR22, ZZX24]. ODE-heat [ZZX24]. ODE-LSTM [CCF24].

ODE-transport [MNSR22]. ODEs

[KFM21b, AT21b, GJJC^{+}23, KFM21a]. off [HH20, KAT22, TS22, WFKL24]. off-critical [TS22]. off-policy [WFKL24]. offence [RYR^{+}23]. Offline [TZL24]. offs [VSM23]. Ohta [CLZY21]. Oldroyd [CJ21, EFJO20]. Oldroyd-B [CJ21, EFJO20]. Olver [Alj21]. on-line [BBK^{+}24]. once [HCL^{+}23, LW22g, LX24a]. oncolytic [ALJ24, EHA20]. One [CSCZ24, HL23b, AKR21, BPT23, BF22, CCdlH21, CPPL23, DdAS22, DGP^{+}23, Elg20, Elg22, ECS21, HRA^{+}20, Kri22, LLZ^{+}20a, LW22a, LV20, MS22a, MUOF22, Nak23, PGKM24, RFFZ21, SH21, SNRA20, SGI20, TCGR24, TZ22, Vyn23, XLG^{+}21, YLCZ20, YFHC21, YWX20b, ZRSK21, Zha22a, ZyZW21, ZDKC20]. One-dimensional [HL23b, BF22, CCdlH21, DdAS22, ECS21, Kri22, MUOF22, Nak23, TCGR24, Vyn23, YFHC21, ZyZW21]. one-parameter [Elg20, Elg22]. one-predator [AKR21]. one-pressure [DGP^{+}23]. one-sided [HRA^{+}20, RFFZ21, SNRA20, YLCZ20]. one-step [Zha22a, ZDKC20]. one-to-one [SH21]. one-velocity [DGP^{+}23]. Online [AHM23, TZL24, XTS^{+}22, ZIML20, GMH^{+}20, LZLZ21, MGM21, XSP22, XZY^{+}23]. Online-Offline [TZL24]. only [FZ22, ZDD^{+}22]. onset [KNEM23]. onto [KSPC22]. Open [Das24, HY24, MUOF22, MD24]. open-loop [HY24]. operating [TZ23a]. Operational [FF21a, HR20a, BR20, DMSS20, ER22, LK21a, LNS^{+}21, RS21, UHZ^{+}20]. operations [DBS^{+}23, ZK22]. Operator [MF21, RT20a, Baz20, But21, CS24c, CCL^{+}24, Gu20, JS20, KKS22b, WLW20, YLG21]. operators [AM20a, ARST21a, BF22, CLUR22, CIHP^{+}21, GA22, GH20b, KS21, Kut20, LG22, MP23, NRP22, QY22, SYX22, ZBA^{+}23a, ZBA^{+}23b]. Opinion [HCH24, Fan21, LF24]. opponent [WZN22]. opposite [KB21]. Oppressed [SP23]. optic [SBT22]. optical [YTZ20]. Optimal [AFG22a, AFG22b, BLS21, CZS21, CUA^{+}22, DYA21, GLTY22, GR21, HLKC22, JC22b, JP23, JKJC21, LJL21, LB20, LYZ22b, MMR23, MPC^{+}24, PNJ^{+}21, PKG^{+}20, SR24, SS23a, TCGTT21, VZ23, WLS22, WP21b, YCWH23, ZYPL20, ZY24b, ZSYX23, ARMM23, AMA20, BHL23, CK20, CZ22c, CSS^{+}24, CZW24, CMMN20, DKK20, DZ21, FKS23, FX20, GS24, GLT^{+}21, HYÇK22, HTZ^{+}23, HWW20, HLLY20, KWH^{+}23, LLF^{+}23, LLY^{+}22b, LXL22, LZLZ21, Ma23, PHS^{+}20, RST21, RSZA23, SJ20, SV23b, TZD22, WNG20, WLD^{+}22, WLPZ22, WXF22, WS23b, WFKL24, XWL^{+}22, XTNF22, XYR^{+}23, YHL^{+}20, YZ23, Zen21, ZTLL24, ZY20, ZC21a, ZeF21]. Optimality [LL20a, LCML20]. optimisation [HR20a]. Optimistic [CSS^{+}24]. Optimization [LD20a, RF21, SJTCdlFL21, dFdOFdSBC20, BHK^{+}22, BZ21, BGRM21, BHL23, BSCL20, CWZ20, CYZ23, Chi22, CDM21, CFPM23, CG20b, FLZZ24, GLP22, GZ20b, xHhPjW^{+}21, Hoj20, LWZ20, LXF20, LAS23, MZW^{+}21, ML22, MM22a, MM24, dDMMTH22, MT24, MCB^{+}21, NGGD23, PWSE21, RNVC21, SH21, SG24, SZCP21, SXS22, SZXL24, TLJO20, Wan21a, WYCL24, WZZ24, XZY^{+}23, YYW^{+}20, YA21a, YWXL21, YL24b, ZL22c, ZZD22, ZZW20]. Optimized [KA22, Vis23, ZC21c, CZZP24, GZG^{+}21, KG22, MZY^{+}22, YHL^{+}20, KA23]. optimizer [CT21]. Optimizing [RSR21, SY21b, DESZ23]. Optimum [XPP^{+}20]. Option [DP21, ARISV20, CKL22, DGL^{+}23, GW22, JP22, KNN20, LSCS21]. optional [ZBL^{+}23, SK24b]. Options [AMOS23, HANP23, LCJ^{+}22, AO21a, ASV22, AT21a, GM21a, HOM22, HL22, LH20a, LWZ22a, Wan24b]. orb [KMZ22]. orb-web [KMZ22]. Orbital [AFGG20, qLgZ21]. Orbits [MD23, BHGPC21, DESM^{+}20, WLY21, Zhu24]. Order [DYA21, KWHS23, REC^{+}23, TZL24, WCS22, XWM23, AD21a, ANS^{+}20, AMPS20, AH21, Alm23, ARÁSY21, ASAK21, AFA24, AA21b, AMZ21, AY22, AGCGSC21, ADB20, AS21d, BRTR21, Baz20, BEe21, BCS24, BT22, BMJ20, BCVV22, BMS21, CM20a, CET22, CCPT20, CGJ21, CLCY23, CZL20, CY21a, CYHR22, CYL^{+}22, CLC^{+}22, CZ22c, CL24a, CLSN24, CGLR23, CHST20, CSQT21, CTBB24, CIHP^{+}21, DP23, DEHFF24, DS22a, DTK24, DZZ23, DG23, DMD21, DWZ20, DDC22, DL20b, DL21c, DS24c, ECS21, FGMS21, FHM22, FTML22, FY23, FG24, GÖRT23, GB20, GWZZ23, GLL23, GJJC^{+}23, GBCP21, GZ20b, GWY20, GL22, HCTS23, HLX21, HWCD22, HP23, HW24, HBCA21, HVYMS21, HT20, HGKD22, HHZZ20, HSZ20, HGZ24, HJY24, IP23, ISM^{+}23, JD20, JYMP^{+}20, JX24, JWC20, JDZH22, JL22, JT20, JSZ20, JWL^{+}24, JT24, KMUY21, KYU22, KR22, KM20]. order [KKPS23, KTEM22, KKLJ20, KEiS23, LZH^{+}20, LYB^{+}21, LLZ21, LKSS21, LHZ^{+}21a, LJL21, LCXM21, LZSC22, LJY22, LZ24c, LLG^{+}24, LHYL24, LLZW22, LZHZ24, LZL^{+}21, LY22c, LTB23, LQNA23, LFY^{+}24, MXF^{+}24, MAAEF21, MWDD24, MY22, MS20, MM20b, MSB22, MT23, MRVC20, NP20, NS22, NI21, NZLW22, NH22, OPP22, PD23, PR20, PZS24, PWZ22, PW22a, PMTT24, QS20, QLG^{+}22, QYLW24, QXG21, RDKR22, RS22a, RC22, RF21, RD22, RDL23, RG20, RG22, ST24, SZCC23, SR23, SGX21, SGPZ22, SK20, SLS24, SJTCdlFL21, SG21, SPZS20, SHW21, SKR22, SV23b, SLTL20, Sug21, SFL^{+}22, SHYS23, SKSC21, TCGR24, TKY23, TWB22, TZ24, THBPMC22, TPTA24, TKF23, URLC21, UHZ^{+}20, VJA^{+}22, VN20, WHH^{+}20, WCY20a, WCY20b, WHM^{+}21a, WGM^{+}21, WHM^{+}21b, WWWZ21, WXP21, WZGL21, WLL^{+}22, WGSS22, WM23, WM24a, WB20, Wu20b, WGLL20, WZ20b, Wu20a, WHY20, WS21]. order [XBLW23, XCSW21, XLYA21, XYLF21, XYR^{+}23, YR20a, YSHV^{+}23, YSL24, YY20a, YWZ23, YWS24, YWR^{+}24, YJ20, YCWH23, YLL20, YHL^{+}20, YZ21, YWXL21, YTLV22, YZ23, YI22, ZPSP22, Zar21, ZJG21, ZWH23, ZL20c, ZM20, ZLX20, ZL20d, ZZ20a, ZSC^{+}21, ZL21a, Zha21a, ZCZ22, ZLK^{+}22, ZL22d, ZSFS22, ZFS22, ZD23, ZC23a, ZrLkL^{+}23, ZLZ24b, ZLZ24a, ZO21, ZY21b, ZLZ^{+}23a, ZS20a, ZWFW22, ZW22b, ZZL^{+}24, ZL21c, ZDMW23, ZWW21a, ZWW21b, ZC22]. ordered [GN20a, TCA23]. ordered/extended [LG22]. ordering [Gan21, RGS23]. orderings [ES24, GPHHC23]. orders [CZ21b, HD22a, WL21a]. ordinal [SNR^{+}22]. ordinary [Cha22, FMP20, HXZ21, HVZZ23, QXSW22, SV23a, VMZDV20, WWZ20b]. Ore [DML21, YM23b]. organized [TLZ20]. orientably [ZDMW23]. orientably-regular [ZDMW23]. orientated [LQLF22]. orientation [GLMT21]. Orientations [GPHC24, LKK^{+}23, YD22a, ZQ23]. oriented [BCR22, IP23, MFK22, Son24, XYZZ22]. Ornstein [AB20b, WYLZ24, ZY21a, ZEA22, dlCJ20]. Orthogonal [DDM24, AHL21, ALM20, DK22b, FMN21, Hu20, JCH^{+}20, LAH22, RGAM20, WJ20, XBC22a]. orthogonalization [BKK24]. Orthonormal [Liu24]. Orthotope [WWW24a]. Orthotope-search-expansion-based [WWW24a]. orthotropic [KKPG20, PS21a, Sha21, Yu23]. Oscillation [CCPT20, YY21a, Baz20, CGJ21, JD20, KÖ21, KÖY24, KLS20, MM20b, PMTT24, WHD^{+}23, Wu20a, ZWW21a, ZWW21b]. oscillations [MRR20, RT20b]. oscillator [JYMP^{+}20, KWZX23, MYW21, WKM24]. oscillators [DdlCS24, GBH^{+}21, Ma24, PNJ^{+}21, SWS^{+}22a, SJTCdlFL21, Zha20]. oscillatory [DG20a, DS22b, ELOY22, GW21a, GNA22a, GNA22b, KX23, ML20, PD23, TFL22, WYLW20, XZ23a, XGF20, YF21a]. osmotic [LSL^{+}20]. osmotically [NHK^{+}22, NHK^{+}23]. Ostrowski [AKPK21]. out-of-range [NGGD23]. outbreak [DKKJ21]. outcome [CM20b, RSXH20]. outcomes [HZXT24]. outdegree [XYZZ22]. Outer [ACS20, KKK22, LLL23b, LK21b, MSK21, PLH22, SPM22, SMM^{+}24, ZMS23b]. outer- [LLL23b]. outer-planar [PLH22]. outerplanar [AV21a, AV21b, DYZC20, RiTZ24]. Outlier [FLH^{+}21, LCZW23]. Outlier-resistant [FLH^{+}21, LCZW23]. outliers [ON21, WCP20]. Output [GZL^{+}23, GZLY24, HKCK24, HB24, LGC^{+}24a, LLLS20, MD22, SM21c, TLHH21, ZXW20, ASS^{+}23b, CZLH20, CLWX23, DMD21, DSZ^{+}23, DSS22, GZX20, GSS24, GS21, HQS^{+}24, HZL21, HSL^{+}24, HMZZ20, JWL^{+}24, JJSY22, KY22b, LPS23, LLZ20b, LCLS20, LWZ21, LLZ21, LJZL22, LJ23, LZHZ24, LZS^{+}21, LD21b, LLWW22, MWZX21, MTRA22, MNSR22, NH22, PGZ^{+}24, RFFZ21, RDL23, SX24, SPZS20, SWSW21, SM23, TPTA24, WXWS20, WZZX21, WGZZ22, WSSX22, WWW24b, XD23a, XWL^{+}22, YXMW22, YZY22b, ZTLL24, ZZL^{+}20a, ZXC20a, ZDZS21, ZLLL21, ZCS^{+}22, ZLC24a, ZPDL20, ZZ20b, ZJLT23, ZDD^{+}22]. Output-based [ZXW20]. output-constrained [LLZ21]. Output-feedback [MD22, DSZ^{+}23, KY22b, RDL23, WXWS20, XD23a]. Output-mask-based [GZLY24]. outward [XAZ^{+}21]. Overall [JW23]. overdetermined [ZL21b]. overflowed [Chy22c]. Overhead [AHM23, AC20]. overlapping [EA24, GLC23, WBY^{+}21]. overlying [PS22]. overrelaxation [ZV21]. oversampling [CCE^{+}22]. overview [KRF20]. own [JPZH23]. ownership [Cro22]. oxygen [GP22, KCT^{+}21].

PACF [LBX^{+}23]. Packet [LWL^{+}23c, CZWS20, CZK^{+}22, FSLZ23, HDP^{+}21, KK24a, LHC^{+}22, LWLA21]. Packing [XLL24, HH23b, MBK23, RLK^{+}20, YW22a]. packings [ASS24, YG24]. Padé [DHR22]. Page [GNA22a]. PageRank [BS23, SSCW22, TVY21, ZTYZ23]. pair [HLZ22, LZYL21, SP23, WW22a]. pair-transition-coupled [LZYL21]. paired [HPT22]. pairs [HYS24, LYZ22a]. pairwise [BKK24, KT21a, LZL20]. Pancreatic [MD23, AS21a]. Pancyclicity [HCJC21, HQZC24]. pandemic [JM23a, MGFF21, MBLL24, XYR^{+}23]. pantograph [DYT21, GLTY22, IS21, WHY22]. pantograph-delays [GLTY22]. paper [DS21c, KT21a, Kop22, KLH24, RO21, SK24b]. parabolic [AzR22, BM20, CCdS^{+}23, CFP22, CEPZ22, DG23, HHGZ23, IR22, JPY23, KM20, KPP22, Kod21, KSR21, LLZW22, MR22, MNZ20, MV21b, MV23, NE21, QS20, RKJ20, SIK22, SR23, SZSZ22, Vyn23, YR20a, YCWH23, YY22, Zen21]. parabolic-logarithmic [Vyn23]. parabolic-type [YY22]. paradigms [SNS^{+}20]. paradise [CPS24]. Parallel [DTP21, RZZ21, YLPC22, AAS23, GM21a, LW22g, MGM21, SLZ20a, WFKL24, ZTYZ23, ZZVL23]. parallel-in-time [LW22g]. parallel-machine [MGM21]. parallelism [ZMS23a]. parallelization [SZCP21]. Parameter [AT21b, DE21, GGY22b, KSR21, MWS23, MC20, VSM23, WCS22, ZrLkL^{+}23, ASR22, ADS23, CCC23, CNYC23, Elg20, Elg22, HP23, HSZ20, HCX^{+}24, JT20, JHSL21, KMLW22, Kar22, LJAM24, LD20a, LWLA21, LLS21b, LZ22f, LZ22g, MT24, PP21, PWSE21, PMP21, PCN^{+}24, QNX^{+}23, RC22, RDL23, SR24, SJ20, TWR20, TZ23b, WWW24a, Wri24, XYCG21, YWS^{+}20, ZLS21, ZZD22, ZQC24]. parameter-dependent [RDL23, TWR20]. parameter-filled [PWSE21]. Parameter-uniform [KSR21]. parameter-uniformly [RC22]. Parameter-Varying [VSM23, PCN^{+}24, WWW24a, ZLS21]. parameterization [KNW21, ZTLL24]. parameterized [BZ21, CYL^{+}22, HLL22, LMX20, Pop20, TWWL24, YLC20a]. parameters [BHL23, BGH24, CCL^{+}24, CHHC24, CLL24, CKL22, DK22b, DHS23, KMM^{+}24, KPP22, LLY^{+}23, LD23a, LCLH24, LY24, LY22c, LLCL23, QYG^{+}20, TDZ^{+}24, XbYW20, YHL^{+}20, ZL21a, ZS22a, ZLZ^{+}23a, ZWFW22]. Parametric [AS20, GZ20b, BFGGM24, BCHRS23, DdlCS24, GW22, LSVA20, SV23b]. parametrisation [Fre22]. parametrization [SSR23]. parametrized [DG23]. parareal [XHL20]. parasite [CGB20]. Pareto [GR21, WY23]. Pareto-based [WY23]. Pareto-Optimal [GR21]. parity [CGTU20, KCJ22, ZZ23]. part [LZLM20, XbYW20, GLX22, GLWX24, LTL^{+}20]. Partial [KF23, ASR22, BR20, CHST20, DG23, FX20, GHND22, HT20, HWWW23, ID21, JC22a, JY23, KÖ21, KG22, LLY^{+}23, LWL^{+}23a, LFM^{+}22, LR23, Ma21a, MV21b, MV23, QXG21, RZZ21, RG20, SRK23, SS24d, SV21a, SV23b, WZCF21, WWY21, WL21a, WL24a, WZW24, Wen20, YY22, ZG24, ZYW21b, ZLP21]. Partially [ZZD22, DLS^{+}23, FLZ23, GYC^{+}22, HLX21, LL21, LTW22, LYL22, YTGC20, ZY24b, Zha23a, ZS22d]. Partially-coupled [ZZD22]. Particle [GN21, KPPK23, LS22b, TGL^{+}24, AD21a, BGRM21, CTBB24, DFG22, GY22b, LW23a, MJ23, PAY22, WYCL24, ZH22a, ZYCY22, ZWJ^{+}24, ZBA^{+}23a, ZBA^{+}23b]. particle-laden [ZWJ^{+}24].

Particle-number [GN21]. particles [GK20, KXW24, MJW21, RWGS22, RSZA23]. particular [GBF21]. partite [LY22a, SZ21, SZ22, XYZZ22, ZLL20]. Partition [CDP23, CD20b, CDLR24, CHL24, CM21b, FLL^{+}22, NA21, SNR^{+}22, WHF20, YXW23]. partitionability [LNZ21]. partitionable [BBB24, BBK^{+}24]. partitioned [BT21, SG20b]. Partitioning [WWWK24]. partitions [ASS24, AACI21, CF23a, WLY20]. partly [LL22b, ZWW21b]. partner [SC20c]. parts [EW21]. party [SWZZ23, SGJ^{+}21, YY24]. passage [dJCGVM24, DD21b, GN22, DDM24, XCZ22]. passification [WCGW20]. passing [LZW^{+}20a]. Passive [XZQ^{+}22, LR24, RFFZ21, WZC^{+}22, WLHW23, XZQX20]. Passivity [CZLW24, DDGBGS^{+}24, FJFS21, HHY20, WCGW20, TXZ24, YMW24]. Passivity-Based [DDGBGS^{+}24, FJFS21]. past [DM23]. patches [VZ23]. Path [ZHFL23b, CF23a, DP21, DLM20, HZCY23, KWW22, RX20, RW21, TT20, TT23, XZ24, Zha21c, ZHFL23a]. path-distribution [RW21]. path-independent [Zha21c]. path-index [XZ24]. pathogen [PPS23]. pathogens [EKKE20, NLP^{+}22]. Paths [YM22a, CMT20, DSAY24, QLLY21, Sch22, SS24a]. pathways [GPRV21]. pathwise [dlCJ20]. patient [RS24]. patrolling [CGMFSHR21]. Pattern [LWX^{+}22, BLO22, CG21, CWF^{+}21, LGS22, MGS20, MSMK21, SMTT23, WMT22, YMRJ22, ZST22, ZLC^{+}24b]. patterning [LLS^{+}22]. Patterns [GR21, ALM22, ACMT23, AE21, CWC20, CTV21, GPHC24, JG21, KZZ20, LY20a, LWZW20, LJHL21, MVCK24, MUMS22, PGKM24, TLZ20, TTMU20, XLL^{+}21b]. pay [ZRMD21]. payoff [HGD^{+}21, JTW^{+}23, LZCL23, MH24, QZWY20, YHSZ20]. payoff-driven [HGD^{+}21]. payoffs [GWZ^{+}20, IKÖ24, Wet21b, ZSJ23a]. PB [SMW^{+}23, SMW^{+}23]. PD [LLWK23]. PDC [DL20c, LPK22]. PDE [DBVCP23, BSdCCGCV21, FLZZ24, GW22, LWZ20, LJ23, MNSR22, SUN24, SWSW21, SZSZ22]. PDE-constrained [FLZZ24, LWZ20]. PDEs [CCdS^{+}23, CCZ24, CW22, HLLY20, IP23, IR22, LH22, LLP23, LW22g, UKA23, XCSW21, ZHZZ24]. PDFs [LZDX22]. PDNN [KMS^{+}23]. PDT [WLZF21, WCW22]. PDT-switched [WCW22]. peak [GT22b, LX24a]. peaks [LSZ^{+}23, ZIML20]. pedestrian [GW20b, LCZ^{+}20, RYG21, TS23, YYW^{+}20]. pedestrians [FQHS24]. peer [QZZ^{+}23]. Peirce [TCG20]. Pell [YY22]. penalized [DESZ23]. Penalty [CW22, Agg22, DR22, EFF23, HMS20, Mit24, SJ20, ZL20c]. penalty-like [CW22]. pencils [BMS21, Gan23]. pendant [DS22c, LG24]. pendent [DSAY24]. penetrable [Azi21]. Penney [VGdFCBA20]. Penrose [AACI21, BT21, BSST20, BF23, CJ20, Hua24, SRGS20, Xu20, ZYW21b, ZMS23b]. pension [PZZF24]. pentagonal [SY23]. pentagraphs [Xu23]. pepper [SGPZ22]. Per-flow [Chy22d]. per-spectrum [SW24]. perception [WZA^{+}24]. Percolation [ZLP^{+}22, WLL^{+}22]. Perfect [BHP21, WF23a, ZLL20, DTDFT22, LZ24a, LLL^{+}23a, MY24, SD23, TJIM20, WCS22, YS20]. perfectly [SS23b]. perforated [AEE21, DG22b]. Performance [DHR22, LD23a, LZY^{+}23b, RD24, BDKZ21, DZZ^{+}20, FZLJ24, GMH^{+}20, GL24, HTF20, HR20a, KLH24, LDX20, LX20a, LXWH22, LHS21, LFM^{+}22, LL23d, dDMMTH22, SYL21, SYTC24, TWL20, TW21b, VJKZ21, WZXH22, XZZW21, YP21, YHL^{+}20, YZ24, ZWW23, ZY24b]. Performance-constrained [LD23a]. perfusion [CCG^{+}22]. Period [MD23, CDM21, ZBFW22]. Periodic [MD23, WLY21, AS20, AGA20, AB21a, ASS^{+}23b, BKU20, BHGPC21, BE20, CD22, DH24a, HW24, ID21, Kab24, KY24, KR20, qLgZ21, LGS22, MV21a, NTLNX24, PPS24, RRR19, RRR20, RC21, Sug21, TZ23a, WCG^{+}20, XZ21, XBC22b, XLZL23, Yi23, Zen21, Zha23b]. Periodically [CJC20, CYL^{+}22, ZPW22]. periodicity [Bia21, HWC20, LKSS21]. permanental [LQW20]. permeable [Ach21, MMUS22]. permutable [EW21]. permutation [AGA20, CL20]. permutation-and-backtrack [CL20]. permutations [WWZ20a, WLY20, ZGWZ22]. Perron [FJ24, LYZ22a]. Persistence [AB21a, AB20b, HWY^{+}22, HC24, Liu22b, WMWS21, ZLDY22]. persistence-based [HC24]. persistent [CZK^{+}22, GLD21, LLWW22, MFS22, WLHW23, YCL^{+}21, ZGLZ22]. person [SJ23b]. Personalized [LZY^{+}23a]. perspective [HRZ^{+}23, YR20b, ZJ23]. perspectives [AD21b, BBAS21]. Perturbation [Gan23, GL23b, MMS23, TTL20, AMPS20, FX20, LP22, MGS20, MSMK21, RHNT21, RMSP22, SW20, Xu20]. perturbations [AAJ23, AACGTB23, AACGTB24, LXW^{+}20, LJW^{+}22, SGI20, TZ22, YWS^{+}20]. perturbed [ASR22, Ahm21, AP21, CSL24, DLS^{+}23, GLD21, HCL^{+}23, HKCK24, JTC^{+}24, Kud20b, KSR21, LCS^{+}22b, LZ22e, LJWD21, LZ22f, LZ22g, MWDD24, RC22, SR23, TZ23b, VN20, WCZY20, WC24, WSL^{+}21, WS21, YR20a, YLK^{+}21, ZL20a, ZL21a, ZL22b]. pest [TCGTT21]. Petersen [GLW20, LLZ22b, SS24b, Xu24a]. Petrov [BS22a, HLZ20]. Petryshyn [DDE21]. Petviashvili [GLZB20]. PFEM [SSF24]. PGSS [WX21b]. Pharmacokinetic [LY21b, LY21c]. Phase [BJ21, CLZY21, CZ22c, DG21, HXF23, JNY^{+}21, JT24, JM23b, KR20, LZY20, LTC^{+}23, MMDD21, MBLL24, NS22, NKO^{+}23, OBP21, RMSP22, SSJ22, SGI20, TYCK23, TC20a, WHJK23, XCZ22, XAZ^{+}21, YWXL21, ZSS24, Zhu24, vdMvRS21]. phase-field [CLZY21, DG21, HXF23, LZY20, LTC^{+}23, NKO^{+}23, OBP21, WHJK23]. Phase-specific [BJ21]. phase-type [XCZ22]. phenomena [Fab20, SSCG20, TZW22]. phenomenon [PS21a, WHD^{+}23, ZLS23]. phenotypic [CC24, LFA22]. Physical [HD20, SLS24, DXZ^{+}23, FT24, GHP24, HLLS22, LWM^{+}21, LLZB24, LLLS20, LJWD21, OPL22, SY21b, SWS^{+}22b, SCHS22, TAU^{+}22, WLL^{+}24, XRZH22, YWC24, YCZ^{+}23, Yu21, ZGZ^{+}23, ZWZX21]. physics [CQ24, DBVCP23, NYHC24]. physics-informed [CQ24, NYHC24]. phytoplankton [BE22, ZST22, ZLN20]. phytoplankton-zooplankton [ZLN20]. PI [HDP^{+}21, KJL20]. PI-type [KJL20]. PI/PID [HDP^{+}21]. Pick [DMRR22]. PID [FLH^{+}21, HDP^{+}21]. PIDE [GM21a]. Piecewise [KSS23, RFFB23, AHB20, ADLU23, ASS^{+}23b, BRFLGH21, CFP22, HKKL22, HKU21, LLWY23, LLWW22, MFS22, SO21a, SGL21, WLY21, Yun21, ZJC21]. piecewise-affine [LLWW22, MFS22]. piecewise-continuous [HKKL22]. piezoelectric [ASS^{+}23a, JGXL23, KET22, LM22, MZX21, MKKG22, PS21a]. piezoelectric-orthotropic [PS21a]. piezoelectric-type [LM22]. Pinning [FZS21, LK21b, TKY23, WYLL22, WDTF22, WZPL24]. PINNs [GWH24]. Pipe [KSH22, BS22a, BPRR22]. pipeline [ABNS24a, ABNS24b, ABNS24c]. pipelined [HS23b]. pipes [MYC^{+}23]. Pitaevskii [CS24a]. pitchfork [MYW21]. placement [TZ22]. Placing [LG23]. plan [PZZF24]. Planar [WWY24, AHL21, CLCD23, CFR^{+}20, DHS23, GKK^{+}24, GLMT21, GZZ21, HZ20a, HZ23b, HJL^{+}23, JMY22, KFR20, KPS22, LLL23b, LC23, MD22, PWL23, PLH22, QLL20, QLLY21, SLL^{+}22, WZ20a, WHF20, WHKT23, WWWK24, YCW23, ZMD^{+}21, ZCW24b]. Planck [AF24, LM23, UGG^{+}20]. Plane [CGTU20, KWH^{+}23, AS20, BG20a, CSW20, CDD23, DGV22, DHZ23, KMZ22, MM22b, ME21a, QHC21, SSCG20]. plankton [SCD23]. planning [AACGTB23, AACGTB24, DYW^{+}20, dFdOFdSBC20]. planted [ACMT23]. plaque [NE21]. plasma [MN23, SRN^{+}23, VBMB21]. plasticity [LFA22]. plate [AS20, CH24a, KRS20, LC21, WPE20, WSZ22, Yu23]. platelets [NTLNX24]. platelets-reinforced [NTLNX24]. plates [DD20a, Eft21, SS22c, SLHX21]. platoon [WWW^{+}21, ZZX^{+}22]. platypus [GNZ20]. Play [CPMS22]. player [GLLZ24, QSW23, QCW24, Wet21b, Xia21]. Plug [CPMS22]. Plug-and-Play [CPMS22]. plume [TM22a]. plus [LZ22c]. ply [SS22c]. PM [WWZ20b]. PMC [LZ20c]. PMSMs [TlS^{+}22]. POD [FKS23, SV23b]. POD-based [SV23b]. Poidge [WYZ24]. Poidge-convexity [WYZ24]. Poincaré [CLZ24]. point [AS21c, AHL21, BF22, DDE21, ESA20, EBJ20, HBCA21, HSM21, HWXC20, KY22a, KKPG20, LD20b, MMM24, NL20, OH24, RS22a, RY24, SWS21, SCC24, SB24, TCZ22, THBPMC22, WC21, WM24b, WX21b, YR20a, ZS21a, ZZW21b]. point-primitive [TCZ22]. points [BOT24, RBPN20, WYW^{+}20, ZDKC20]. Pointwise [ZSFS22]. Poiseuille [DG20a]. Poisson [CD20b, JZP^{+}20, LHW^{+}23, LLZ22c, MB24, RM23a, RX20, SMW^{+}23, SZP21, STH24, UZ20, VBMB21, YZC^{+}22, YWT20, dSLV20]. Poisson-driven [SZP21]. Pol [MYW21, AGGPO23]. polar [ACS24, DFX24, FX20, WGM^{+}21]. polarising [RMO22]. polarization [ZLU21]. polarized [MS21b]. pole [TZ22]. policies [CHP23, CLD^{+}24]. Policy [AHM23, AM24, BSCL20, CWL^{+}21, JJL23, LX23a, SZXL24, WFKL24, ZRSK21]. political [LKK^{+}23]. pollutants [CSC20]. pollution [HTZ^{+}23, XZH^{+}20]. Polygonal [TM20, LZ24a, Que22, SV21b]. polygons [ZS21a]. polyhedral [DK22c, GM24]. polyhexes [ZZ22b]. polymer [LZY20, LTC^{+}23, SBB^{+}22]. polymers [VBM20]. Polynomial [ST22, ZEA22, BCHRS23, CM22, CGLR23, DP21, FKVZ22, FRR^{+}22, HA20, HZT^{+}21, Jor24, KG22, LX20b, LSCS21, LY23b, MPM23, MKK23, OT20, RFFB23, WJ20, YSL24, YL20a, ZGG^{+}22, ZZ23]. polynomial-based [YSL24]. polynomial-fuzzy-model-based [YL20a]. Polynomially [JHSL21, RDL23]. polynomials [BSG23, CdFP24, DESM^{+}20, DWW24, GHND22, GARD24, JY20, LU24, MJZ21, NNN24, ÖÇÖ23, PP23, RGAM20, UHZ^{+}20, WYG23, YY22]. polyomino [SZZ24]. polypeptide [CDD23]. polytopic [GT22b, LXZ23]. pool [ZRMD21]. popularity [LHQ^{+}23]. Population [LMT24, RSZA23, AALMLM22, Alf21, dJCGVM24, CD23, GLGLH24, GWZ^{+}21, HW24, LZ20a, LD20d, LW22f, MGFF21, NZLW22, PSE24, Shi21, SJZA22, TTC^{+}23, WJL22, WOB21, ZCH20]. populations [BE22, CWY^{+}24a, DNR21, HGD^{+}21, LCS22a, NT23, QCW24, Sek23, SHT^{+}21, WDZG22, Wan24a, YM22b]. poro [BFQ20]. poro-thermoviscoelastic [BFQ20]. poroelastic [SC21]. porous [Azi21, BH22, CKV21, GE21, Has20, HT23, JT24, KRS20, KY21, LZLM20, LTL^{+}20, LWX24, LZY^{+}23b, Ma23, MRR20, MMK^{+}23, MKKG22, MMDD21, MJK22, MS22c, OPP22, OBP21, PJMW22, Pap21, RM20, SN21, SKS21, SGI20, TM22b, YFHC21, ZYY20, ZyZW21]. porous-fluid [LZLM20, LTL^{+}20]. porous-medium [HT23]. port [LJY22]. port-Hamiltonian [LJY22]. portfolio [BLNT22, Chi22, CDM21, KMS^{+}20, KMS^{+}23, MBL20, SG24, XYY23, ZY24b, ZBFW22]. portfolios [AO21b, LOG20, MBL20]. portrait [MBLL24]. posed [CPLH20]. posedness [BPRR22, Ma23, SBT22]. posets [Gas20]. position [KR21, TX21, WWT24, ZDD^{+}22]. positional [CYG20]. positioning [CWY^{+}24a, LXWH22, ZWWH21]. Positive [AGGPO23, MK21b, Que22, CLX22, DCR^{+}20, Gas20, JSM20, KZC23, LW22a, LX20a, LT22, LLWK23, ON21, RH20, SD22a, Sug21, ST22, WQZ^{+}20, WLZF21, WX21a, XZeF20, XZZ22b, YWX20b, ZS21b, Zha23b, ZLJ^{+}21]. positive-valued [ON21]. Positivity [LX20a, INOP20, YTC23b, Yan24]. positivity-preserving [YTC23b]. possessed [Zha23b]. possession [JSCB20]. post [CWL^{+}21, JP20, KWM20a, KWM20b, TCPF24]. post-death [KWM20a, KWM20b]. post-indemnity [CWL^{+}21]. post-processing [TCPF24]. posterior [WCY^{+}24]. posteriori [SY22b, TKF23]. potential [ABTR21, BG20b, GAA^{+}24, HLLW21, Hon24, JLFL20, LLW24, RHN^{+}24, SS23b, XFW^{+}21, YWT20, ZY23a, ZXF20]. potential-based [XFW^{+}21]. potential-Ramsey [ZY23a]. potentially [Kab24]. potentials [ARST21b, ZJJ24]. Power [NYHC24, AA21b, BF22, CMMN20, CZ21b, GSL^{+}22, HOM22, JT20, KMS^{+}23, LWM^{+}21, LHZ^{+}23, LZX^{+}22b, LSC^{+}23, LLQW24, Lyu23, PKG^{+}20, QZZ^{+}23, SSCW22, SWS^{+}22b, WCM^{+}21, WYZX22, WP21b, XYZZ22]. Power-enhanced [NYHC24]. power-GMRES [SSCW22]. power-law [CMMN20]. powerful [DK22a]. powers [HHR21, HCS23, LLG22, LY22b, Miy23, SLZ^{+}20b, WY22, YGHW21]. PPH [JOR^{+}21a]. PPIFE [YCWH23]. PQWs [BG20a]. Practical [CFWL24, DDC22, LT22, QYLW24, ZS21b, ZXZC22, CWZ20, HTF20, KP24]. Prandtl [DO21]. pre [PS21a, YZDC22, ZHJ^{+}21]. pre-assigned [YZDC22]. pre-existing [ZHJ^{+}21]. pre-stressed [PS21a]. preassigned [SYL21]. preassigned-time [CWWZ24]. precaution [CWL^{+}21]. precipitation [OBP21]. Precise [JP22, CZL20, ZDZ24, ZZL^{+}24]. precision [LBX^{+}23, Lóc22, YDXH22, ZMD^{+}21]. precision-adjustable [LBX^{+}23]. preclusion [CLL23]. Preconditioned [CHZ24, HCY22, BSCV24, HWXC20, LLF^{+}23, LGQ20]. preconditioner [HWXC20, LW22g, SWS21, WM24b, WX21b, YAX^{+}21]. Preconditioners [FLZZ24, LMX20]. Preconditioning [LWZ20]. predation [Zha23b]. Predator [KB21, AKR21, AIAFvH21, BLO22, CGB20, CWC20, DWL^{+}24, LHW21, LCML20, Lu22, PGKM24, QLL24, RK22, SZ24a, TTMU20, ZYZ22b]. Predator-prey [KB21, BLO22, DWL^{+}24]. predators [KB21]. predefined-time [AAJ23, ZC23a]. predict [CM20b, YLH22]. predicting [DZL^{+}24]. Prediction [RSXH20, UH24, WWZ20b, ADS23, LDZ21, NGGD23, ZSP24, dAGSFM22]. predictions [NGC21]. predictive [CCL^{+}24, DG23, HL24a]. Preface [BKS21, DBVCP23, Yua20]. Preference [GL23a, BY23b, HRZ^{+}23, QZZ^{+}23, RSR21, TTC^{+}23, YY21c]. preferences [ZY24b]. Preferential [ZH23a, SL20d]. prejudice [TNT23]. premise [LR23]. premium [OLW23, STH24]. premium-based [STH24]. premiums [PV24]. preprojective [Bia21]. Prescribed [CWZL23, GSS24, HQS^{+}24, LXWH22, LFM^{+}22, XXJtC24, YSL21, ZWZ^{+}21, BDPV22, BDH^{+}20, CFR^{+}20, DS24a, FKVZ21, GXZ^{+}24, GL24, HZCY23, INOP20, KPS22, LL23d, QMS22, SYTC24, WZXH22, YLL22, YGHW21, YZ24, ZWW23, ZQ23]. prescribed-performance-based [LL23d].

Prescribed-time [CWZL23, GSS24, ZWZ^{+}21, DS24a, GXZ^{+}24, YGHW21]. presence [BBAS21, CZZP24, FSLZ23, HTF20, ME21b, WCP20, WWW^{+}21]. present [TZ23a]. preservation [FCH21, LSWL20]. preserved [BD20b]. preserves [BBH22]. Preserving [BBH22, REC^{+}23, Alm23, AA21b, ADS21, CHK21, DX24, DBVCP23, GÖRT23, JSCB20, KSW24, LW22c, MV21b, MV23, OKT24, PDK23, PGF20, PW22a, UKA23, WWZ20a, WSY21, WHWW21, XCSW21, YTC23a, YTC23b, Yi23]. pressure [ATDJ22, BFRS24, CFR22, DGP^{+}23, EBGN23, SWS20, Wan23a, WPE20, YOF21]. pressure-correction [SWS20, Wan23a]. pressure-driven [ATDJ22]. pressure-loaded [WPE20]. pressureless [GG22]. prevailing [YTX^{+}24]. prevalence [DKK20]. prevent [KIJK20]. prevention [NGC21]. preview [SRZ22]. Prey [RK22, AKR21, AIAFvH21, BLO22, CGB20, CWC20, DWL^{+}24, KB21, KMZ22, LHW21, LCML20, Lu22, PGKM24, QLL24, SZ24a, TTMU20, ZYZ22b]. prey-one [PGKM24]. prey-parasite [CGB20]. Prey-predator [RK22, SZ24a]. preys [PGKM24]. Price [LWZ22a, YY20b, BBAS21, SNR^{+}22]. priced [SXS22]. Pricing [HANP23, LOS21, WZY21, Wan24b, AT21a, BSdCCGCV21, DP21, GM21a, GW22, HOM22, HL22, JP22, KK23c, KNN20, LH20a, LSCS21, LBJ22]. primacy [WLY20]. Primal [AP22, LAS23]. primal-dual [LAS23]. Prime [LYJM21, LZC22a, SS24e]. Prime-valent [LYJM21]. primes [LY22b]. primitive [LZC22a, TCZ22, WP20]. principal [GLPS22, LWL24, XYlW^{+}24, YZJ^{+}20]. principle [DX24, KLS20, VMR23, WSY21]. principles [DLMP23]. prion [LZ22a]. prior [DHM^{+}20, WHZ^{+}20, WCY^{+}24, XCGN24]. priori [BFQ20]. priorities [CJCL21, LZL20]. Prioritized [LYX23]. prism [LXLP20]. prismatic [ECS21].

Prisoner [SK24b, BY23a, CWY^{+}21, DZ21, Elg20, Elg22, FLHH23, GPH21b, HZX21, HXZ24, HZJL24, LDZ20, LJN^{+}22, LSHH22, LWZW20, LJR20, LL23e, LZD23, LHH24, MRW21, MT21, QCW24, SHT^{+}21, SCT^{+}20, Tak21, Ued23, UU21, WWY^{+}20, WLZW21, YZPL24, YSW^{+}21, ZHQ^{+}24, ZWJ^{+}20, FFL^{+}23b, GLWW22, MHF23, WDW^{+}22, YSYS24, YZZ^{+}22, YYW^{+}23]. Privacy [JSCB20, LRCTR22, MJM20]. Privacy-preserving [JSCB20]. private [Ask22]. Probabilistic [LLWY22, DLL^{+}23, JWY20, LNS^{+}21, QCCW23, TKY23, WLFX22, XBC22a, YSLK23, ZLLL21]. probabilities [CX22, DLS^{+}23, FLR23, HLKC22, WC23, Zha24b, ZM21c, ZLP21]. Probability [VMR23, ABTR21, DL20c, GLY24, Guo22, LWJ21, NK20, NK21b, QLZL22, RS23, SJZA22, ZYySX23]. problem [AP22, ACC^{+}21, AFGG21, AA20, AAA20, AR23, AJIR21, BA20, BHGPC21, Bon22, BM21, BFLR23, BMS21, But21, CM21a, CLX22, CLH24a, CY24, CDL^{+}20, DTK24, DD21b, DS23, DMP21, FQ22, FKK24, FWBY23, GYC24, GK21, Han20, HH20, HBCA21, HT23, HZXZ20, HLZ22, HWW20, KMS^{+}20, KMS^{+}23, KCT^{+}21, KR20, Kwo22a, Kwo22b, LSC^{+}21, LSV21, LaGBT22, LH22, LFA22, LLLL23, LL20a, LWL^{+}20, LLYQ20, LSL^{+}21, LWL24, LLS21b, LZ22g, Lyu23, MGM21, MZ23a, MP20, MBK22, MM20a, NS22, NE21, NWM23, NTNT24, NH21, OMS22, Pap21, SGI20, SW20, SYW21, SKWW21, SY22b, SB24, TBRB24, TTL20, VMZDV20, WHH^{+}20, WJK20, WS21, WGJK21, XZH^{+}20, XJ20, XAZ^{+}21, YSL24, YHB^{+}20, YLH20, YW21, ZM21b, ZL21a, ZL23a, ZL23b, ZSZL20, ZeF21, dFdOFdSBC20, KZ22]. Problems [GLX22, YFW21, AK21, AA24, AACGTB23, AACGTB24, ACGG20, AP20, BFGGM24, BBHM21, BFQ20, BK21, CFF20, CM20a, CRT22, CD20b, CWZ20, CWY^{+}20, CNYC23, CSS^{+}24, CMMN20, CPLH20, Das24, DGP^{+}23, DHR22, DSY22, DTL22, DK22c, EV21, EM21, EBJ20, EFF23, FKS23, FLZZ24, Far22, FZ20a, FHM22, GTSV23, GLT^{+}21, GLWX24, GBF21, Gu20, GWH22, GJWV24, GZLZ24, HHGZ23, HMS20, HL24b, HLLH23, HWXC20, HGZ24, HLLY20, IZ20, JX24, Kar22, KDB^{+}23, KPP22, KSR21, Kum24, Lee23, LW20a, LL23a, LLLC23, LLWY22, LR20, LZRF21, LLCS23, LtTBPA20, LS22a, LZ22f, MV21a, MM22b, MBMA20, dDMMTH22, Mit24, MD24, NYHC24, PWWZ23, QS20, QHC21, RS22a, RNVC21, RMSP22, RPB23, RG22, ST24, SS24c, SR23, SWS21, SV23b, SLTL20, SWX24b, TC22, THW^{+}21, TLPT21, Voe23, VN20, Vyn23, WW20, WFHG20, WM24b, WX21b, XZZ22a, XZ21, XCGN24]. problems [wXhGZ20, wX20b, XLL21a, XLZL23, YR20a, YGM^{+}22, YCEW21, Yan23, YH24a, YAX^{+}21, YA21a, YCWH23, YLC20b, YL23b, Zen21, ZL20c, ZL20b, ZP20, ZC21a, Zha21b, ZL22b, ZZLV23, ZZVL23, ZZL20b, ZZW21b, ZV20, ZV21, ZZW20]. procedure [Eft21, HMEDVR24, LLQG21, MM22a, MM24, Sha21, ZR21]. procedures [CD20a, MMG^{+}23]. Process [AMOS23, AAS23, AS23b, AB20b, BJK^{+}22, Cha23, CL23, DD21b, FZB24, FR23, GK20, GN23, GTK^{+}22, GWZ^{+}20, HVYMS21, JSA21, KRF20, LSC^{+}21, LYN^{+}21, LWZ22c, MS22c, QYD22, SX23b, SMR22, STH24, WHW^{+}21, WL21b, WYLZ24, XYZY20, ZY21a, ZSJ23a]. processes [DP23, DN21, GN20b, GN22, GGBEZRI23, KHNT22, KL21, LZ20a, Nak23, PRP24, SRZK22, ZCLW22, ZG24, dlCJ20]. processing [ARÁSY21, CSV20, GGBEZRI23, HH20, Ily23, MFK22, TCPF24, ZIML20]. Product [YY20b, CHH22, DS24b, HWZZ23, HH22, Hua24, KÖ21, KP23, LLYQ20, LNZ21, MGD21, NSD^{+}23, VZ23, WSC^{+}24, ZT24]. product-type [NSD^{+}23]. production [AVJ24, CH20b, CSZ24, HH20, WGJK21]. products [CSK23, KT20, KT21b, KRT21, TX21]. profile [LC21]. profiles [AO21a, LSZ^{+}23, RFFB23]. program [CCC20b]. Programmable [LL21]. programming [KM22, KRT21, MZLS22]. progressive [ARÁSY21, WL24b]. project [LWZ22a]. Projected [LH20b, ZL22c, LHC21]. Projection [MMS23, Agg22, CCCdLMR23, DL21b, DDCM23, HP23, KS21, LF21, RD22, SWS20, SWW22, TZLL24, WNZ21, ZF21, ZF23a, ZRL^{+}24, ZF23b]. projection-based [ZF21]. projective [BD20a, MAAEF21, PN24, ZSC^{+}21]. prolate [AAL^{+}21]. promote [CWY^{+}24b, CCC20b, FLHH23, HJJL24, ZH23a]. promotes [BY23a, CZW23, DZW21, DZ21, GPH21b, LJN^{+}22, MZ24, MH24, QZWY20, QDW22, SHT^{+}21, SCT^{+}20, TGL^{+}24, WWY^{+}20, WLZW21, WDW^{+}22, XL24b, YZPL24, YY24, ZPZH23]. Promoting [DD21a, LZD23, WCW21, ZRZL24, ZPZH23]. promotion [HWZZ23]. proof [CIHP^{+}21, GS22, ZZ23]. Propagation [TZL24, AZD^{+}24, BS20, BL22, BP23, CLL22, CK20, CWL^{+}21, CLD^{+}24, FM23, GZ20a, GF23, KZWS22, KIJK20, KK24a, LXL22, LLLC23, LQS^{+}23, LXH24, PS22, SSC22, Sha21, SFC23, WW22a, YXZ^{+}21, ZLZ20]. propelled [KXW24, SX23a]. Proper [FSS21, KP22, LW22e, MMOS24, NMS23, GLMT21, MZ23b, QZ21, BFGGM24]. Proper-walk [LW22e, FSS21, QZ21]. Properties [HXL^{+}20, PT21, SKJM21, ABD22, AM20a, AYY24, BM22, BBM21, CWZ21, De 22, FM23, GWSC23, GPHHC23, KRNG24, KP23, LePP24, LWL^{+}20, MS22b, ÖÇÖ23, RY24, SS22a, Sha22, YTGC20, You20]. property [DLL^{+}23, LL23c, XFK22, YXFL23, YWR^{+}24]. Proportional [VSM^{+}21, XKW^{+}23, DXZ^{+}23, LHZ^{+}23, MC22, SXC21, TWL20, WCGW20, WWWZ21, WLPZ22]. proportional-integral [DXZ^{+}23, TWL20]. proportional-integral-based [LHZ^{+}23]. prosocial [QZZ^{+}23]. prospective [WHD^{+}23]. prosperity [CWY^{+}24a]. protected [WQT^{+}24]. protection [CGW^{+}20, MJM20, RK22]. protein [UH24]. proteins [LZ22a, UH24]. Protocol [CHW^{+}24, FLH^{+}21, HLKC22, HCL^{+}23, HKCK24, LDCZ24, LCC^{+}22, LX24a, WYLQ24, ZDC22, ZKLL23, ZD24]. Protocol-based [CHW^{+}24]. protocols [GHS^{+}21, LWD21, LCZW23, YLC20a, Yi23]. provable [JSCB20]. provincial [YWJ23]. provisions [KAT22]. Proximal [LAS23, PX20, DHZ^{+}20]. Prune [LSC^{+}21]. PSE [ZBA^{+}23a, ZBA^{+}23b]. Pseudo [dDMMTH22, UZ20, AS23b, DYT21, GGW21b, HWC20, JZ20, LLWY23, MZL24, RD24, WL23]. pseudo-determinant [AS23b]. Pseudo-feasible [dDMMTH22]. pseudo-inverse [MZL24]. pseudo-spectral [DYT21, JZ20, RD24, WL23]. pseudo-symplectic [GGW21b]. pseudo-sympletic [GGW21b]. pseudo-time [LLWY23]. Pseudo-transient [UZ20]. pseudomatchings [DTDFT22]. pseudorandom [ZH22b]. pseudospectral [CCdlH21, DEHFF24, RPB23, WLPZ22]. PSO [CH20b, JZZ^{+}22]. PSS [HWXC20, SWS21]. psychological [LSHH22]. pth [ZZS^{+}20]. public [BBEM24, CWY^{+}24b, CCC20b, FH24, HWY^{+}22, HJJL24, HRZ^{+}23, JTW^{+}23, LW20c, LS22b, MQW23, PZZF24, QZWY20, QYL^{+}20, QDW22, QCCW23, QZC^{+}24, SYZ^{+}23, SLK^{+}23, WPH20, WCW21, WCXS22, WCY^{+}22, WS23a, WYCL24, Xia21, XL24b, YZY22a, YXZ^{+}21, ZAD21, ZB23, ZRZL24, ZRMD21, ZHZH24]. pull [XBLW23]. pull-based [XBLW23]. pullback [LWX24]. pulse [LX23a, ZZZ^{+}23]. punching [DD20a]. punishing [NCY^{+}20, ZPZH23]. punishment [GDS24, LCS22a, LC22b, LS22b, NCY^{+}20, QZZ^{+}23, QCCW23, SWZZ23, SCT^{+}20, SLK^{+}23, ZAD21]. pure [EW21, JY21, SLK^{+}23, YWY^{+}22, YP22]. pure-feedback [JY21, YWY^{+}22, YP22]. purely [SW24]. Pursuing [RYG21]. pursuit [XBC22a]. puts [LYZ22b]. PVM [PGPCK21]. Pythagorean [BCHRS23, FKVZ21, FKVZ22, SS23a]. Pythagorean-hodograph [FKVZ21].

Q [CHJW23]. Q1 [YHW^{+}23]. QoS [HDP^{+}21]. quad [DYW^{+}20, Han20]. quad-curl [Han20]. quad-rotor [DYW^{+}20]. quadratic [GKK^{+}24, HYT21, HWWW23, LS20, LPP21a, LPP21b, LR21, LD20c, Qi22, SRK23, SJ23b, TLY24, TITN20, TTL20, VZ23, WWY21, WHL21, WJK20, WPJ22, YLZL21, ZY20, ZN20, ZS20a, ZS22d]. Quadratized [BCP23]. Quadrature [DO21, KX23, Zho20, BFS20, KWZX23, LX23b, ML20, MYC^{+}23, MN22, OPS20, QZ22, ZBBK23]. quadrilateral [GLC23]. quadrilaterals [ZWC23]. quadrotor [XWWL23]. quadrotors [SYL21, WGZZ22, YSL21]. quadrupolar [LSZ^{+}23]. quads [ZBBK23]. Qualitative [SWW21a, WJL22, FM23]. Quality [YY20b, LW22f, QDL^{+}21]. quantification [FdLPC^{+}22, KG21]. Quantitative [GNSZ22, CTBB24, SK24b]. quantization [CGG23, CJ22, GZWX23, KGCC20, LCC^{+}22, LXZ23, MPX^{+}20, MSZ22, MTRA22, SC20b, WS23b, XLLH22, YYG20, ZY23b, ZXZ20]. quantization-based [CGG23]. quantizations [LWLA21]. Quantized [YP21, ZPY22, CSS^{+}22, CXY^{+}23, GSS24, LSWL20, LX24a, SWSW21, SCG^{+}20, WCZ20, WSSX22, ZWHY20, ZTW22]. Quantum [Elg20, Elg22, Son24, TZL24, AKPK21, CS24b, FZ20a, LMZ20, SV23a, Wan23b, ZK22]. quarantine [JM23a, KAT22, ZHJ^{+}21]. quartic [BD21, LMS24]. quartically [SUN24]. Quasi [LCXM21, WS21, ZLZ^{+}23a, Ach21, ARST21a, ADLU23, BEIR23, CYHR22, CSV20, DLM20, GCGM21, HCJC21, KKS22b, KS21, LS20, LZL^{+}21, MPM23, QQ23, SWS^{+}22a, Sun22, TCZ^{+}20, WGSS22, XLG^{+}21, YF21b, ZM20]. quasi- [GCGM21, TCZ^{+}20]. quasi-bipartite [SWS^{+}22a]. quasi-compact [WGSS22]. quasi-interpolanting [ADLU23]. quasi-interpolants [BEIR23]. quasi-interpolating [ARST21a]. quasi-interpolation [CSV20, LS20, Sun22]. quasi-linear [QQ23]. quasi-Newton [MPM23]. quasi-Newtonian [ZM20]. quasi-one-dimensional [XLG^{+}21]. quasi-projection [KS21]. quasi-semiregular [YF21b]. quasi-solutions [KKS22b]. Quasi-stability [LCXM21]. Quasi-synchronization [ZLZ^{+}23a, CYHR22, LCXM21]. Quasi-uniform [WS21]. quasicrystals [QHC21]. Quasilinearization [SLTL20, YI22]. quasiperiodic [Zhu24]. Quaternion [Lia21, CWB22, CXYZ23, HD22a, LZH^{+}20, LHZ^{+}21a, LCXM21, LCY^{+}21, MF23, SZCC23, SWX^{+}21, SWX22, URLC21, WHM^{+}21a, WHM^{+}21b, WWWH21, XFK22, YLCL23, ZSC^{+}21, ZJWV24, ZLZ^{+}23a]. quaternion-valued [CWB22, LZH^{+}20, LHZ^{+}21a, LCXM21, LCY^{+}21, SZCC23, SWX^{+}21, SWX22, URLC21, WWWH21, YLCL23, ZLZ^{+}23a]. quaternionic [MP23, PP21]. quaternions [ÖC24]. QUAV [YCM^{+}22]. QUAVs [CXY^{+}23]. quenching [ZLS23]. Queue [GL23b, CA20, Chy22d, SZXL24, ZRSK21]. queueing [AVJ24, BSCL20, MB24, dOSR21]. queues [SCC24]. queuing [KXW24]. quiescent [SWCT21]. quorum [EE22].

R [BTL^{+}22, MDJW20]. rabies [SJZA22]. race [CFF20]. Rachford [BZ21]. radial [CNYC23, Kar22, KY21, NA21, Oru21, Oru22, YLC20b, Zha22b, ZF23b, ZHZZ24]. radiation [FWM^{+}22, KHA22, LW20b, WXLY23, YAX^{+}21]. radiative [Ach21, Azi21, CSG^{+}22, CSG^{+}23, SN21]. radii [BOT24, TGL^{+}24]. radiotherapy [AACGTB23, AACGTB24, dFdOFdSBC20]. radius [CLZ23, CWY^{+}24b, DM21, DM24, DW20, GZ20c, LG24, LZ20b, LLCL23, QLM21, YLY24]. radon [GZRM21, BMMV23, MS21b, RJM23]. Rainbow [GLW20, JWWL20, JMY22, JJC24, QLL20, QLLY21, CMT20, DDS24, MZ22]. Rakotch [RY24]. Ramsey [CF23a, HLLH23, JWMW23, LLW20, LLZ22b, PLH22, WLL21, XLL21a, ZWC23, ZY23a]. Ramsey-type [CF23a]. Randić [BNR20, GGS21, MMMBRS20]. Random [DDH22, NKA20, BS23, BCVV22, CWZZ22, CLM20, CRTV22, DDLW21, DLMP23, FLL^{+}22, FZW23, HGKD22, HLKC22, HLLY20, JW23, Jor21b, Jor23, Jor24, KG22, LX22, LWX24, LXW^{+}20, LTL^{+}21, LZD23, PPS20, PMG20, SLPH22, TWZ^{+}20, WL24a, WWT24, WSL^{+}21, XZ23c, YCL^{+}21, ZY20, ZYLH21, ZYCY22, Zha24b, ZS20b]. Randomized [WL24b, Che21, CH20a, Du24, ZL23b]. randomly [CZ20, KMLW22, KGCC20, KKPS23, LY22d, WB23]. randomness [ACVHC22, LOVHC23]. Range [AR20, ALM20, GLPS22, HZX21, JP22, NGGD23, Zha24b, ZLZ24b]. ranges [WX22, YP22]. rank [BPT23, BBH22, DHM^{+}20, DGMP22, FMN21, LFLH23, LHZ^{+}21b, MY22, PLH^{+}24, SMM^{+}24, SS21c, TZ22, YZJ^{+}20]. rank-1 [MY22]. rank-deficient [SS21c]. rank-one [BPT23]. ranking [TVY21]. ranks [Hu20]. Rannacher [DK22b]. Raphson [AHEGM23, CB21]. rapid [CCH24, DK22a]. rare [YXX^{+}22]. Rashba [SDD22]. ratcheting [JP20]. Rate [SS23c, FLC^{+}22, GZH22, GZCS23, GM21b, HDP^{+}21, HZW21, JDLZ21, LOS21, LW24, MHAT21, RHNT21, SRZK22, SCC24, Wan24b, WY24, XZeF20, XTNF22, ZZCL20]. Rates [GL23b, QY22, EBJ20, JWK22, KH23, LLZ20b, LL20b, Ma21a, NK21a, OH24, QYG^{+}20, RS24, SK24a, SH22, SKSC21, TW21a, WP21b, XXW^{+}20]. rating [WYW^{+}20]. ratio [AIAFvH21, JNY^{+}21, RWGS22, XW22]. ratio-dependent [AIAFvH21, XW22]. Rational [DDNS20, QDW22, BE20, Gas23, LL20b, Liu24, PSSU21, SS23a, SS24a, SLZ^{+}20b, WLZW21, WWZW23, XLD23, YGHW21, Yun21, ZZ23]. rationality [PLW^{+}20, ZWXQ20]. Raviart [WS21]. ray [AHS20, DVS21, HLLW21]. Rayleigh [ASUH21, DMP21, JNY^{+}21, KZS20, SSC22, SNZC20]. Rayleigh-type [SSC22]. RBF [Oru21, Oru22, CD20a, CD20b, Mar24, SV21b]. RBULT [WM24b]. reach [LLWY22]. reach-Avoid [LLWY22]. reachability [TZ22]. Reachable [CH24b, CGZ^{+}22, LCS^{+}22b, WXW^{+}20, ZL22e, FZLF23, FFL23a, FZLJ24, JPZ^{+}21, LLWY23, ZFJ^{+}20]. reaching [XFY20]. Reaction [CGB20, HHY20, IMOR24, Alf24, BBH22, BFLR23, CZLW24, CYHR22, DL21b, DTL22, EHA20, EFF23, GK21, HWD20, HLZ20, HXZ21, HB24, JHO^{+}21, KZWS22, KDB^{+}23, KSR21, LLZ22a, LF21, LWL23b, LR20, LK21b, LGS22, MZ23a, RC22, SMTT23, SWW21b, SSW21, SHYS23, SKH20, SKSC21, TY21, TZW22, VMZDV20, WW20, WJL22, YMRJ22, YGZ^{+}22, ZL20a, ZH20a, ZJC21]. Reaction-Diffusion [HHY20, CGB20, IMOR24, BFLR23, EHA20, GK21, HWD20, HLZ20, HXZ21, HB24, JHO^{+}21, KZWS22, KSR21, LWL23b, RC22, SMTT23, SHYS23, SKH20, SKSC21, VMZDV20, ZL20a]. reactions [DG20a]. Reactive [MHF23, DG20a, Has20]. Reactor [EE22, Zar21]. reactors [XYZY20]. Real [HR20b, RLK^{+}20, LZ22d, Miy23]. Real-time [RLK^{+}20]. rearrangements [BMMV23]. receivers [AB20a, THBPMC22]. reciprocal [FJ24]. reciprocally [TWR20, WHPQ23]. reciprocals [CHY22b, ZQZ21]. reciprocity [LCJ^{+}22, MT21, SWP^{+}22, SYZ^{+}23]. recirculation [LJ23]. recognition [YY21b]. recommender [LZY^{+}23a]. reconstitution [DXZ^{+}23]. Reconstructed [LNC24]. reconstructibility [LLLL23]. reconstructing [XBC22b]. reconstruction [CLCY23, CTBB24, DEHFF24, DL20a, JLOR22, LL23a, MT23, NTT21, RJM23, SWS^{+}22b, WWZ^{+}23, WNZ21, WHJ^{+}24, ZSYX23, ZZW^{+}21c]. reconstructions [PD23]. records [BAL^{+}24, MH24]. recovered [WQZD23]. Recovering [IMS23, AzR22]. Recovery [ZBA^{+}23b, CLCY23, CLC20, FMN21, LZ24d, PX20, TCPF24, WCY20a, WCY20b, XYlW^{+}24, ZH23b, ZGCZ22, ZBA^{+}23a]. RECs [BSdCCGCV21]. rectangle [LSL^{+}20]. rectangular [CH24c, Eft21, KKB23, ME21b, WCY20a, WCY20b, ZZ20a]. recurrence [BAL^{+}24, FdLPC^{+}22]. Recurrent [WYD20, RY24, SCG^{+}20, WCGW20]. Recursive [LWLA21, LWL^{+}23c, TZ22, LX20b, SS21c, ZHG22, ZLC^{+}24b]. Redefining [AKP23]. redistribution [Xia21]. reduce [FB20, HTZ^{+}23]. Reduced [DG23, KMUY21, KKPS23, MP23, WL24a, GJWV24, HREB21, Ily23, ISM^{+}23, KYU22, KNW21, LZHZ24, RD22, RM20, SV23b, TKY23, WWWZ21, WM23, Zar21, ZC23a]. Reduced-order [KKPS23, LZHZ24, TKY23, WM23]. Reduced-quaternionic [MP23]. Reducing [JGC^{+}21, RSZA23, ZCPZ20]. reduction [AZD^{+}24, CM20a, CLL24, CDL^{+}20, DS21d, LJY22, LLWY23, LHZ21d, RF21, UKA23, WZXY20, WJ20, XBC22b, YSL21, Zhu24]. redundant [CHW^{+}24, LL23b]. reference [LCLS20, MCD21]. Refinability [MPRS24]. refined [YWG24]. refinement [CD20b, LTB23]. Refinements [CX22, dOS20]. Reflected [RT20a]. reflection [PS21a, Sha21, SSCG20]. reflectionless [BBM21]. refraction [SSCG20]. refuge [LHW21, LCML20]. regarding [YJL20]. regime [LZ24c, LB20, PRP24, RCCL21, WZY21, ZCJL21, ZZ24, ZC23b]. regime-switching [LB20, WZY21]. regimes [AS22]. Region [ZDZ24, FRW^{+}24, KMZ22, LePP24, LQSZ21, YTC23a, CG20b]. Regional [HKKL22, HXL^{+}20, JM23a, NPE22]. regional-scale [NPE22]. regions [AGGPO23, SV21b, UH24]. regions/proteins [UH24]. regression [AS23b, Che21, CKSP20, KMM^{+}24, LHS21, QYD22, SN21]. regret [NH21]. regular [AG20, BS22a, DFP^{+}24, HZF21, LYhH^{+}24, LHL24, Prz24a, SNMM21, TWB23, ZZ22b, ZDMW23]. regular/irregular [BS22a]. regularity [LX22, LFLP21, ZM21b, ZFS22]. regularization [AYY24, CPLH20, LSZ^{+}23, LWZ20, LMX22, MGJ21, Mit24, OH24, PX20, SR24, SMW^{+}23, SGPZ22, UZ20, WWZ^{+}23, WNZ21, WHJ^{+}24, XCGN24, ZH23b, ZY21b]. Regularized [Du24, CKSP20, GZG^{+}21, KKS22b, qLgZ21, LHS21, MG21, RO21, WDG21, WZZ23, YW21]. regularized-long-wave [qLgZ21]. regularizer [MY22]. regularly [CKW22, JS22]. regulation [JWL^{+}24, LJZL22, LGC^{+}24a, MWZX21, SM21c, SM23, XL24b, ZTLL24, ZDZS21]. regulatory [KR22]. reinfection [GOO20]. reinforce [ZH22b]. reinforced [NTLNX24]. Reinforcement [AHM23, GLL^{+}22, HXWC24, SJ23b, CLH24a, DWS24, GS24, JLZ^{+}22, LJL21, LD23a, LYX23, LTW24, LZZL21, MPC^{+}24, PHS^{+}20, TKY23, Ued23, UU21, WGW^{+}21, XTS^{+}22, ZSZL20]. Reinsurance [GLX22, GLWX24]. rejection [MGM21, SSKS21, XWWL23, YSW^{+}22]. rejections [SKLS22]. relapse [GOO20]. Related [ASS24, DMRR22, FR24, GG22, Hon24, HLM23, HSY24, Nak23, Sha22, THBPMC22, YLH20, ZGWZ22]. relation [AFGG21, OKT24, ZyZW21]. Relations [DESM^{+}20, GAMP20, VPV^{+}21]. Relationship [LCZ^{+}20, ST24, CH22a, LQS^{+}23]. relationships [GP22, LZY^{+}23a, WZLH22]. Relative [VJA^{+}22, MH22, ZY24b]. Relativistic [IHF13, Woo20, CCSPMV23, GST22, KY24, ZLU21]. relaxation [DGV22, He23, LYZ22a, PHL23, YWL20]. Relaxed [LPK22, XSP22, DDPS22, TLDL21, TKF23, WHL21, WCZ22a, ZGLZ22]. relaxes [MT21]. release [HR23, QLKW20]. relevant [LQLF22]. Reliabilities [WMF23]. Reliability [Kun24, LHW^{+}23, LWZ22c, QH21, WMT22, XZ23c, ZR21, CL20, LZX22a, XFY20]. Reliable [LXZ23, LLWW22, MSZ22, PW22b, YGZ^{+}22, ASKA20, DZZ^{+}20, LDX20]. Remark [Cas23, FLS20]. Remarks [BF20, PMTT24, WW22b, DFX24]. Remodelling [WL21b]. removal [BGH24, DHM^{+}20]. Remove [SGPZ22]. Renewable [BSdCCGCV21]. renewal [Cha23, FLHH23, RCCL21]. Renormalization [Miz21]. Rényi [MMMBRS20]. Repeated [QCW24, CZ20, KTZ24, MQW23, PZ20, TG20, Ued23, UU21, WTCC20]. replay [LYX23, LLZB24]. replenishment [CDL^{+}20]. Replicator [NZ20, Wet21b, Wet23, YM22b]. repositioning [KKR21]. Representation [CFA^{+}23, EW21, ACH^{+}22, JT20, WF23b, ZZS^{+}20]. Representations [MS21d, SĆL^{+}21, Kut20, MSMK21].

Reproducing [AS21b, GW21a, AD21a, JY20, LU24, XTNF22]. reproduction [GLGLH24]. reproductive [NS20]. repulsion [TCZ^{+}20]. Reputation [HJJL24, QCCW23, QZC^{+}24, BYCX24, BY23a, BY23b, FHFS24, GWZ^{+}21, HZJL24, JLWX21, PWH20, QZWY20, WCW21, ZWJ^{+}20]. Reputation-based [QCCW23, BY23b, FHFS24]. Reputation-dependent [QZC^{+}24]. reservoir [RMSP22]. residual [GLKrH20, NYHC24, TC20b, ZL21b]. residues [TITN20]. resilience [LZZ22]. Resilient [CC22, LNS^{+}21, LZS^{+}21, LCL22a, ZW22a, FT24, HKL22a, HQK21, LCLS20, LWD21, LSC^{+}23, MZY^{+}22, WSDW24, WYS23, XD23a, XWZC21, ZDGL21, ZKLL23, ZZX^{+}22]. Resistance [ALJ24, CY21b, LXLP20, BJ21, Liu22b, SSP24, SPX20, SPX24, YWFL23]. resistant [CTZ21, FLH^{+}21, IG21, LCZW23]. resistive [CCSPMV23, WGMJ23]. resolution [FZ23, HL23b, OKN^{+}22, RSR21]. resolve [CTWG24]. resolvent [GA21, JS20]. resolves [DGL^{+}23, HWY^{+}22, HC24]. Resolving [KWHS23, RT22]. Resonance [BTP20, DG22b, BF22, BSM23, CH24a, SB24, SZZ24, UUS21, XGRM20, YTLV22]. resonator [AB20a]. resource [BY23b, CGW^{+}20, GCLHT23, HRZ^{+}23, JLL^{+}20, MMR23, MMG^{+}23, SL20d, SWX24a, ZYPL20, ZYZ22a]. resource-epidemic [CGW^{+}20]. resource-limited [GCLHT23]. resources [JM23a, WS22, WDW^{+}22, ZLW^{+}21b]. Respect [BBW21, FF21a, LLZ23, LLSY23, LY24, NT24, RRS21, SS21b, SY23]. respiratory [HTZ^{+}23]. Response [CA20, JZZ^{+}22, AEE21, Alf23, AIAFvH21, BKR22, CFWZ20, EHA20, GAR22, GN21, Har24, LWL21, NNKH21, SWCT21, YLLL20, ZY21a, ZYZ22b]. responses [AS21a, AKR21, EA21]. restarted [AACGTB23, AACGTB24]. restarting [BSCL20]. resting [SSC22]. restoration [LZ22c, LHZ^{+}21b, WZZ23, dSLV20]. Restoring [LCS23]. Restrained [SSA^{+}23]. restricted [Elg20, Elg22, GZRM21, LM21, MTRA22, TJIM20]. restriction [YG24, YCW23]. result [DZZ^{+}20, LLY^{+}21, LW22b, LLY22a, TW21b]. Results [AD21b, KSS23, AC20, BBHM21, BPR21, BLS21, CLZ^{+}22, CLC24, CF23a, FSS20, FFL23a, FSLC22, GHND22, GN20a, GNZ20, GNA22a, GNA22b, HWLL21, HCL^{+}23, ID21, IÖÜP23, KRE20, LSPK20, LY21a, LZZ24b, MZ22, MGD21, WNP20, WLG^{+}21, WZL22, XWMC20, ZFJ^{+}20, ZC23a]. retaliation [QSW23]. retinal [CCG^{+}22]. retirement [JP20, JP23, LWZ22d]. Retraction [CSG^{+}23, NHK^{+}23]. retransmission [WMWS21]. retrial [LHW^{+}23]. retrieving [ALM22, DK22b]. returns [CKW22, PMG20]. revengers [ZZPF23]. revenue [SZXL24]. revenue-cost [SZXL24]. reversal [GL23a]. reverse [CIHP^{+}21, LCWZ24]. reverting [ZY21a]. review [HR20a, SS23d]. Revisited [TWB23, JSCB20, MS21a, YZ23]. revolution [OMS22]. reward [SLK^{+}23]. Rewarding [CHP23]. Rewards [CCC20b, LGD^{+}21]. reweighted [PX20, WCP20, dSLV20]. Reynolds [CU23, ZF23a]. Rezaee [Kop22]. Rezaifar [Kop22]. Rhoades [GN20a]. Rica [LJHL21]. Riccati [BBH21, LK21a, WSZ20]. Riemann [BS20, CLX22, CSZ24, EBGN23, ESA20, Gu20, HZXZ20, HLZ22, JM23b, LC23, MGK23, PGPCK21, QLG^{+}22, SLTL20, ZL23a]. Riemannian [dSAOG21, SSI23, ZN20]. Riesz [QL21, XZZ20, YLS23, ZGZ^{+}24, ZZX20b]. right [NHBJ20]. right-skewed [NHBJ20]. rigid [KZS20, RB21, SSC22, SNZC20, ZTW22]. rigid-flexible [ZTW22]. rigid/free [KZS20]. Rigorous [CD22, PR20]. ring [BTP20, SLJ23]. rings [ZYW21b]. Riordan [SRK^{+}20]. rioting [YR20b]. rise [WDZG22, XLL^{+}21b]. Risk [SG24, ZCLW22, ARMM23, ASUH21, CKW22, DD24b, HSLS20, HGKD22, HWZZ23, JS22, JGC^{+}21, LOG20, LCH20, LWZ22d, OLW23, PRP24, STH24, TZ23a, Xia22, XZ21, ZLW23, ZZ24, ZY24b]. Risk-sensitive [SG24]. risks [CWL^{+}21, CLD^{+}24, Guo22]. RK [ADS21]. RKHS [YL23a]. RL [WS20]. RL-based [WS20]. RLC} [ZC23b, EVB^{+}21]. RLW [WP21b]. RMIL [You20]. rna} [KCJ22]. road [CCH24]. robin [CRT22, GK21, HKCK24, ZKLL23, HKL22b, KSR21]. robin-type [GK21]. robot [SSW22]. robotic [ZTW22]. robotics [LPS21]. robots [CWZL23, CHHC24, HCBY21, MC22, YP21, ZDD^{+}22]. Robust [ARISV20, BTL^{+}22, CZWS20, CLSN24, CZ24b, DZZ23, FJLG22, FMN21, GT22b, EVB^{+}21, HSZ20, HD20, HYT21, JPZ^{+}21, LLLL23, LJWD21, LJW^{+}22, LR23, MZW^{+}21, MFA21, MZSW21, SNRA20, SDW^{+}22, VN20, WCP20, WQT^{+}24, WX21a, ZZ20a, ZDC22, ZFS22, ZH23b, ZLZ24a, ZGCZ22, ZL21c, BDH^{+}23, BJK^{+}22, CHHC24, CZWX21, DLL^{+}23, GK21, HTF20, JYMP^{+}20, JWC20, KRRS22, LHZ^{+}21a, LLY22a, LCJ^{+}22, LZ23, LCL22b, LZ24d, LQNA23, MCT^{+}22, MGFF21, PP21, PKJB22, PGZ^{+}24, SCLD21, SYL21, SYTC24, TCA23, TEGP24, VSM23, WXP21, WLFX22, XZZ22b, XYlW^{+}24, YZJ^{+}20, ZIML20, ZeFWD23, ZZS^{+}22, ZTW22]. Robustness [DXS^{+}20, LSYX21]. rock [DS21c, KT21a, KLH24, SK24b, YXXZ23]. rock-paper-scissor [SK24b]. rock-paper-scissors [KLH24]. ROCOF [DP23]. rod [AR21a]. rod-beam [AR21a]. Roe [EBGN23, PGPCK21]. Roe-type [EBGN23]. Roesser [FRW^{+}24, MTRA22, WLZF21]. rogue [LM20b, ZJG21, ZLA^{+}21]. role [KT21a]. roles [WMWS21]. rolling [DWS24, OMS22, SKWW21, WXWS20]. ROM [LW22g]. ROM-accelerated [LW22g]. Roman [ACS20, AÁC^{+}21, ACSVT22, HCSJ24, IZDR24, SSA^{+}23, YS20]. room [GW20b]. root [BDTP23, DD21b, LWX^{+}22]. root-indices [BDTP23]. Rooted [KZ22]. Rosenau [WP21b]. Rosenbrock [ZF24]. Rossby [WZY20]. Rotary [SS24e]. rotating [DS22b, KZS20, KYU22, PS21a, SCMK23]. Rotation [LHZ21d, ÖC24]. rotational [SWS20]. rotor [DYW^{+}20]. rotors [JA23, UMT22]. rough [AARS23, RSA24]. roughness [KSH22]. round [CM20b, HKCK24, ZKLL23]. round-robin [ZKLL23]. round-robin-like [HKCK24]. route [ZSZL20]. routing [XZH^{+}20]. row [WCZ22a]. RRKPM [LWQW22]. RSS [Zen21]. RSS-like [Zen21]. rubber [NWM23]. Rubenchik [Oru22, Oru21]. Rucklidge [VHB23]. Ruin [Guo22, CX22, TZ23a]. rule [BBEM24, GE20, Kur21, LX23b, ML20, SR24, WCY^{+}22, XYCG21, YZ23, ZL22d]. rules [CKSP20, HZJL24, JL22, LXF20, YWJ^{+}20b, YSW^{+}21, Zho20]. Rumor [LQS^{+}23, TZL24, HCM21, HPC24, KZWS22, ZMJ^{+}23, ZLZ20]. Runge [BMJ20, CCSPMV23, DS21a, RAB24, ZZX20b].

S [CSS^{+}22, GLWL22, LPK22, LR24, LKX20, LR23, NK22, NH22, SJ23a, ZJZD20, ZWZS22]. S-Box [LKX20]. saddle [HWXC20, SWS21, WLY21, WM24b, WX21b]. Safe [CKSP20, GS24, MXF^{+}24]. Safety [QL20, SSW22, WJL^{+}24, JLL^{+}20, ZSYX23]. Safety-critical [WJL^{+}24]. Safety-triggered [SSW22]. Saint [MP20]. Saint-Germain [MP20]. SAIQR [FW24]. sales [MKZ22]. salt [SGPZ22]. same [KB21, ZYC^{+}24]. sample [CSS^{+}22]. sample-data [CSS^{+}22]. Sampled [GY22a, LXW^{+}20, WHCW20, WCX^{+}22, XWDK23, ZDGL21, DZZ^{+}20, GLWL22, HKKL22, HB20, JDLZ21, JHSL21, KKH23, LSPK20, LHFW21, LLT^{+}20, NK22, PHL23, SWX^{+}21, VSM22, WGL^{+}21, XWMC20, XXW^{+}20, ZZH^{+}20, ZLZ^{+}21]. Sampled-Data [XWDK23, GY22a, LXW^{+}20, WHCW20, WCX^{+}22, ZDGL21, DZZ^{+}20, HKKL22, HB20, JDLZ21, JHSL21, KKH23, LSPK20, LHFW21, LLT^{+}20, NK22, PHL23, SWX^{+}21, VSM22, WGL^{+}21, XWMC20, XXW^{+}20, ZZH^{+}20, ZLZ^{+}21]. sampled-data-based [GLWL22]. samples [LFLL21]. Sampling [CCG^{+}22, BSG23, CCZ24, CLL24, CSV20, HL23a, KL22, LSC^{+}21, LNS^{+}21, WCG^{+}20, dRST23]. sandpile [AV21a, AV21b]. sandwich [KET22]. Sarmanov [STH24]. SARS [EA21]. SARS-CoV-2 [EA21]. SARS-CoV-2/cancer [EA21]. satellite [HKU21, XCLW21]. satisfying [DX24]. saturated [AS23a, CTZ21, CWB22, CHW^{+}24, GE21, Has20, HLKC22, LHC^{+}22, MJK22, MHAT21, PJMW22, SK24a, SC21, TM22b, ZXW20]. saturating [RC21]. saturation [ASKA20, CMT20, CL24b, DYX24, DL20c, HLL^{+}21, HYT21, HRA^{+}20, MTRA22, MMHD20, RHNT21, SSKS21, SPZS20, SZ21, SZ22, WXP21, WSS22, WYZX22, WM24a, WZ24, WX21a, XZQ^{+}22, XW23, Yan24, ZYSW22, ZPDL20, ZWYL22]. saturations [WB23]. Saunders [NHBJ20]. SAV [LW22d, ZSS24]. Sawada [ZLA^{+}21]. Sawada-like [ZLA^{+}21]. scaffolds [RC21]. Scalable [MMG^{+}23, GdFPF24, WBY^{+}21, YLPC22]. scalar [CZ20, HLZ21, KK23a, Wan23a]. Scale [YXYZ23, CWY^{+}20, CY21a, CLLA23, EE22, GW21b, HXWC24, JCH^{+}20, KBL^{+}24, LTW22, LLG22, ML22, NPE22, OBP21, SL20a, SNS^{+}20, YLL22, YGHW21, YC21, YTX^{+}24, ZLW23]. scale-free [YTX^{+}24]. scaled [NL20]. scales [AB21b, GE20, MGJ21, SC20a, TZ21]. Scaling [CDBR^{+}22, BGB24, DS21d, RSZA23]. scattered [AS21c, DDNS20]. Scattering [WS24, DG22b, MK21a, MM20a]. scenario [MBLL24, ZBFW22, ZZX^{+}22]. scenarios [PKG^{+}20]. Schaefer [CL23]. Scharfetter [Que22]. Schatten [YHD24]. scheduled [TWB23]. scheduling [AAA20, CYG20, DTP21, HH20, MGM21, QLKW20, SKWW21]. Scheme [DG21, ABPP23, AH21, ADS21, BR20, BBD^{+}22, BCS24, BSCL20, BRMVCD21, CGG23, CT24, CD20b, CZL20, CLZY21, CdR22a, CHHC24, CH24c, CKV21, CCSPMV23, CXGL22, DYT21, DX24, DWS24, DL20a, DK22c, EBGN23, FHM22, FZ23, GLL23, GdFPF24, GMN22, HCTS23, HWCD22, Hu21, IZ20, JSCB20, JOR^{+}21a, KM20, KK24b, LSV21, LPK22, LC22a, LZ23, LHYL24, LC23, LL23d, LU24, LTB23, MCT^{+}22, MWDD24, MSN21, MV21b, MV23, MT23, OPP22, OBP21, OKT24, PGF20, PHS^{+}20, PWZ22, PW22a, QQ23, QL21, RKJ20, RVV24, RO21, RJM23, SKLS22, SNMM21, SMH^{+}20, Shi21, SZ24a, SG21, Sun22, TY22, THW^{+}21, TBRB24, TC20a, TO20, WDG21, WSY21, WLPZ22, WJDL24, WS21, XYM22, YR20a, YTC23a, YTC23b, YLK^{+}21, YLG21, YJ20, ZHL24, Zha22b, ZSFS22, ZHQ^{+}23, ZL23a, ZGZ^{+}23, ZSZL20]. schemes [ABD23, Alm23, AR24, BDH^{+}23, BT22, CCCdLMR23, CG23, CPPL23, ECS21, EFJO20, GÖRT23, GB20, GG22, GCGM21, GI23, DDCM23, He23, HJY24, JY20, JHO^{+}21, KPG22, LZSC22, LZ24c, LtTBPA20, LMÖ23, MTRA22, MDT22, PD23, QS20, REC^{+}23, RST21, SR24, TCGR24, TFL22, Vis23, VN20, WGSS22, WL23, Wan23a, YZLS20, YF21a, YWXL21, Zha21a, ZSS24, ZYX21, ZC22]. Scholes [SO21b, SO21c, SUN24]. Schrödinger [AEL21, BF22, BLS21, CG20a, HLLW21, JZ20, Kud20b, KL24, LM20b, LWNZ22, LZYL21, QWD22, Sun22, YWZT20, YLG21, YTZ20, YWT20, ZJG21, ZSL20, ZLJ20]. Schultz [ZMS23b]. science [SG20a]. scissor [SK24b]. scissors [DS21c, KT21a, KLH24]. Scomber [JC22b]. scoring [WCW21]. Scott [SMTT23]. screening [CKSP20]. Screw [YA21b]. SDDEs [BMY24]. SDEs [FLR23, KMM^{+}24, LCL22c, PS21b, YC21, dRST23]. SDP [LYZ22a]. sea [SLCS20, WWT24, JKJC21]. SEA-based [WWT24]. sea-level [SLCS20]. Search [KMS^{+}20, HR20a, LePP24, LFLH23, WWW24a, You20, ZP20]. searchers [Wet21a]. searching [ALM20, AR20]. seasonal [PSE24, TGDPS22]. seasonalities [TGDPS22]. seasonally [PSE24]. Second [CLCY23, CUA^{+}22, DMD21, LQNA23, MS20, OPP22, AAS23, AMPS20, ARST21a, AFA24, ADB20, BCVV22, BMS21, CKN21, CCPT20, CGJ21, CY21a, DDC22, EHVMM24, FRRS24, GZ20b, HBCA21, HGZ24, JD20, JWL^{+}24, JT24, LSL^{+}20, LJL21, LHYL24, MWDD24, NS22, NI21, PH20, QS20, RS22a, RG22, SK20, TZ24, TCA23, WHH^{+}20, Wan22a, WHY20, WYG23, YSHV^{+}23, YJ20, YCWH23, YWXL21, YZ23, ZL20c, ZL20d, ZZL^{+}24]. second-grade [LSL^{+}20]. Second-order [DMD21, LQNA23, MS20, AFA24, BCVV22, BMS21, CGJ21, CY21a, DDC22, GZ20b, HBCA21, HGZ24, JD20, JWL^{+}24, JT24, LJL21, LHYL24, MWDD24, NI21, QS20, SK20, TZ24, WHH^{+}20, YSHV^{+}23, YCWH23, YZ23, ZL20c, ZL20d, ZZL^{+}24]. second-ordered [TCA23]. secondary [LZ24d, Pop20]. section [DC21, SBB^{+}22]. Secure [DTWJ21, CMYF21, DS24a, DXZ^{+}23, HXWC24, LLLS20, WCM^{+}21, XZZ22b, YXZ^{+}24, ZCS^{+}22, ZY23b]. securities [Ued20]. Security [ZGZ^{+}23, CWL^{+}21, CLD^{+}24, GSL^{+}22, LZ24d, SWS^{+}22b, WLJ^{+}21]. sediment [GDFNNR22]. seed [NPE22]. seeking [CLSN24]. Segel [ALL23]. segment [IMIH22, LLSY23]. Segmentation [PGL^{+}23, HKU21, UBSU21, WSG^{+}21]. Segur [DYA21]. SEI [GM21b, HZJ^{+}21]. Seiches [MRR20]. Seidel [XH23, DS21b, XH22b]. SEIR [CRTVG^{+}23, XYR^{+}23]. SEIRS [SJZA22]. SEIRV [MCSD21]. Selection [DJW20, AKY21, BY23b, BLNT22, BHL23, CCC23, CNYC23, CTWG24, LLY^{+}23, SWP^{+}22, SL20d, WCY^{+}22, WS23a, XYY23, ZCH20, ZZ22a, ZST22, ZH23a, ZBFW22]. selective [HKU21]. Self [FZ20a, ACMT23, BKU20, BSG23, CLCY23, CGW^{+}20, CZWX21, Har22, KAT22, KXW24, LXF20, NKO^{+}23, OPP22, SX23a, WMCZ21]. self-adaptive [CLCY23]. Self-concordance [FZ20a]. self-consistent [BKU20]. self-corrected [NKO^{+}23]. self-initiated [WMCZ21]. self-inversive [BSG23]. self-learning [CZWX21]. self-optimization [LXF20]. self-propelled [KXW24, SX23a]. self-protection [CGW^{+}20]. self-quarantine [KAT22]. self-similar [OPP22]. self-supervised [ACMT23]. self-support [Har22]. selfadjoint [MM20a]. seller [MZW^{+}21]. selling [DD21a]. Semi [DHZ^{+}20, MV21a, MWS23, ASKA20, BT22, BRMVCD21, CZ22c, CCH24, CL24b, CJ21, DdlCS24, DP23, FG24, GZQ^{+}20, GDFNNR22, GY22b, GLMT21, GXZZ20, HHGZ23, HKL22a, HL24a, HH22, IR22, KKPS23, KKS^{+}22a, KTEM22, LLZ^{+}20a, LFLL21, LLY22a, LM24, LR20, LNC22, LTWY24, LZ24d, Liu24, MGM21, MRR20, MBMA20, NK20, NK21b, PHL23, PKJB22, PBKS23, SSKS21, SD22a, SZSZ22, SSW21, SDW^{+}22, WQZ^{+}20, WGL^{+}21, WPZY23, WM23, WHCW20, XCW^{+}22, XYKL20, XZQ^{+}22, XLL^{+}21b, YWFL23, YCWH23, ZYSW22, ZWZS22, ZLU21]. semi- [GLMT21]. Semi-analytical [MV21a, CCH24, LR20]. semi-annular [KTEM22]. semi-arid [XLL^{+}21b]. Semi-blind [DHZ^{+}20]. semi-closed [MRR20]. semi-definite [SD22a]. semi-dihedral [LTWY24, YWFL23]. semi-discretization [IR22]. semi-implicit [BT22, BRMVCD21, CZ22c, FG24, GDFNNR22, PBKS23]. semi-infinite [CJ21, Liu24, MBMA20]. semi-inner-product [HH22]. semi-Lagrangian [PKJB22, PBKS23]. semi-linear [DdlCS24, HHGZ23, SZSZ22, YCWH23]. Semi-Markov [MWS23, CL24b, DP23, GZQ^{+}20, GXZZ20, KKS^{+}22a, LZ24d, PHL23, SSKS21, SSW21, WQZ^{+}20, WPZY23, WM23, WHCW20, XCW^{+}22, ZWZS22]. semi-Markovian [ASKA20, HKL22a, HL24a, KKPS23, LM24, NK20, NK21b, SDW^{+}22, WGL^{+}21, XZQ^{+}22, ZYSW22]. semi-online [MGM21]. semi-relativistic [ZLU21]. semi-smooth [XYKL20]. semi-supervised [LLZ^{+}20a, LFLL21]. semi-synchronous [LLY22a]. semi-tamed [LNC22]. semi-torus [GY22b]. semiarid [TLZ20]. semilinear [AMF20, CY21a, RC22, XZZ22a]. semiregular [YF21b]. sense [YMW24]. sensing [EE22, LPL^{+}22, MJ23]. sensitive [CJCL21, CTZ21, SG24]. Sensitivity [AT21b, AHEGM23]. Sensor [LWL^{+}23c, BMMV23, CHW^{+}24, CL24b, DL20c, KBL^{+}24, LWD21, LLLS20, MSZ22, NGS22, WFYL21, WX21a, WB23, XH22a, XLZ23, ZCZW23, ZXW20]. sentiment [YXZ^{+}21]. separable [HVYMS21, HMS20, LT22, YLS23]. separate [AAA20]. separation [CFR^{+}20, HZ20a, LHZ^{+}21a, SV21a, WZZ24]. September [Ano20k, Ano20v, Ano21l, Ano21x, Ano22l, Ano22x, Ano23l, Ano23x, Ano24l, Ano24x]. SEQIHR [ZHJ^{+}21]. sequence [DL21a, LLSY23, YSLK23]. sequences [AE21, LY20a, RGAM20, SRK^{+}20]. Sequential [ZR21, GWH24, MZW^{+}21]. serendipity [WDC^{+}23]. Series [SS23c, BF22, BCP23, CSW20, CCF24, Ema21, GE20, JT20, KGPS22, PMP21, QDL^{+}21, SV21a, WZY21, XCZ22, Zha24a]. server [MB24, SZXL24]. service [GY22a, GJS^{+}23, LWD21, LZWC22, LJWD21, LWY^{+}24, MB24, SCC24, XXJtC24, YCZ^{+}23, ZBL^{+}23]. Set [YSLK23, AHL21, CH24b, CGZ^{+}22, CLLA23, CLH24a, CSC20, CFPM23, Fen22, FZLF23, FFL23a, FG24, JPZ^{+}21, LLL21, LL23b, LCS^{+}22b, LRS20, MBL20, RZZ21, VJKZ21, WXW^{+}20, WLFX22, Wan23b, WZPL24, WLZD24, ZeFWW20, ZFJ^{+}20, ZeF21, ZL22e]. set-based [CFPM23]. set-membership [LL23b]. sets [FZLJ24, HLP24, LWX24, LX24b, MZW^{+}21, MW20, PMAK22, TT24, TLD22, WLWL22, XT23]. setting [ASV22, BSCV24, LCD^{+}20, vdZGO21]. settling [GY22b, YLL22]. setup [AAA20, GLWW22]. several [CGJ21, KZZ20, KRE20, WLG^{+}21, Wu20a, XL24a, Zha23b]. Severity [BEe21]. sexual [LJHL21]. shadowing [ZZX20a]. shallow [ABPP23, Bab22, CCCdLMR23, CG23, EBGN23, DDCM23, KYU22, KPG22, LLQG21, LC23, MMK^{+}23, TCGR24, WP21b, WSZ22, YTC23b]. shallow-water [TCGR24]. Shannon [HOM22]. Shape [HBWQ20, AF24, CNYC23, CLM20, ECS21, Kar22, LZDX22, LLY^{+}23, MKK23, YY21b, MPRS24]. Shape-adjustable [HBWQ20]. shape-morphing [AF24]. shaped [FN22, GY22b]. shapes [QZZ^{+}23, SH21]. Shapley [IÖÜP24]. share [Wet21a]. sharing [LXF20]. Sharma [Alj21]. Sharp [HT20, CGJ21, JM23b, SMW^{+}23, YJL20, ZWW21b]. sharp-interface [JM23b]. shear [GAR22, KSPC22]. shear-thinning [GAR22, KSPC22]. shearlet [WNZ21]. sheath [SBT22]. sheet [TMA^{+}22]. Sheffer [SRK^{+}20]. shell [Hoj20, VBM20, WSZ22]. shift [BFLR23, LMX20, MQ22, ZH22b]. shift-splitting [LMX20]. Shifted [CGLR23, Han20, RS21, SSCW22, HD22a, UHZ^{+}20, YLL20]. Shifting [WWT24, WX22]. Shilnikov [BRFLGH21]. ships [ZWWH21]. Shishkin [LZ22g, MZ23a, ZL22b]. Shiu [PRP24]. Shock [CK21, BKMO22, WZZ23]. shock-diffusion [WZZ23]. shocks [LMZ20, LHW^{+}23]. Shooting [AT21b, ACGG20]. shop [HH20]. Short [NGC21, CJC20, De 22, FLS20, HZ23b, LWL23b, LOS21, WGLL20, YWR^{+}24]. Short-term [NGC21, CJC20, YWR^{+}24]. short-time [De 22]. Shortest [DTDFT22]. shortfall [OG20]. shot [JTW^{+}23]. shows [KB21]. Shuangbing [RO21]. shuffle [QH21]. shuffle-cubes [QH21]. SI [MK21b]. sic [WMQ21]. side [KB21, YCEW21]. sided [HLZ20, HRA^{+}20, MT24, RFFZ21, SNRA20, YLCZ20, ZLS23]. sides [HPL21]. Sign [MSMK21, PD23, MGS20]. Signal [BY21, ARÁSY21, CJ22, CLCY23, GGBEZRI23, PX20, SWSW21, SRZ22, XZZC22, YSL21]. signals [DLMP23, MF23, NCY^{+}20, WYLW20, ZDX21]. signature [BMMV23]. Signed [IZDR24, RS22b, BBT21, CZLH20, CY24, DP21, GL22, GWW23, HT24, KCJ22, LSWS21, SY22a, Spi22, Sta20, TLCM24, WLZ21, WSC^{+}24, XLZ^{+}24a, YLHZ22, YH24b, YML21]. Significance [AKH^{+}22, NNKH21]. Signless [ZCT22, BOT24, LLZ23]. Signorini [DSY22]. silence [ZLZ20]. silicate [SPX20]. similar [GGL^{+}23, OPP22]. Similarities [LtTBPA20]. similarity [CDBR^{+}22, KK22, TPSP24, YL20b]. SIMO [KLNN23]. Simple [KRRS22, ABD23, BSCV24, Eft21, HC21, HL22, NS21, PGPCK21]. simple-loop [BSCV24]. simplex [LX20b, LMS24, VSM23]. simplicial [DDGN23a, DDGN23b, LXP^{+}22, NLP^{+}22, SG20a, Sha22, YFM^{+}24, ZLP^{+}22]. simplified [NKO^{+}23, SWS21, TWZ^{+}20]. Simulated [LCC^{+}24, UMT22]. simulating [FTML22, PBKS23]. Simulation [AS22, BSRV23, MMDD21, MJW21, XYCG21, XZ23b, Ach21, AKH^{+}22, BCLNV20, CFR22, CK20, CZL20, CSC20, DD20a, GdFPF24, GP22, HSM21, IMIH22, LYKK21, LZY20, LR20, MM22a, MM24, Oru21, Oru22, QPHC21, SNS^{+}20, SBB^{+}22, SLS24, THW^{+}21, TAU^{+}22, VGR^{+}21, VBM20, YWL20, ZHYD20, ZH22a, ZLS23, ZLU21, dlCJ20]. simulations [DdAS22, DC21, LTC^{+}23, LG22, Lu22, NKA20, SMTT23, SRN^{+}23, VBMB21, ZRL^{+}24]. Simulink [FGMS21]. Simulink-based [FGMS21]. Simultaneous [GXZ^{+}24, XbYW20, ZGLZ22, XL24a]. simultaneously [TTL20]. Sina [YXZ^{+}21]. Sina-microblog [YXZ^{+}21]. Sinc [QXG21, YWZ23]. Sinc-collocation [QXG21, YWZ23]. sine [Alm23, CWZ20, YL22]. sine-Gordon [Alm23, YL22]. Single [SD22a, GY22b, JH21, MB24, MJ23, QLKW20, QXSW22, SRKA23, WHZ^{+}20, ZZY20]. single-handed [ZZY20]. single-server [MB24]. single-species [JH21]. Singular [AEL21, BF23, GLX22, ARST21b, AJIR21, AS23b, ADB20, BPT23, BR20, BM20, BGB24, CWH24, CLZ^{+}22, DYT21, FZLF23, FZLJ24, GGW21b, GP23, GLWX24, Iza21, JZP^{+}20, JW21, KLNN23, LLZ20b, LX20a, LBX^{+}23, LR24, LM24, LSM23, LZW^{+}20b, LZZ^{+}23, LX23b, LXZ23, LZZ24b, LYN22, MZSW21, NBTP23, RFFZ21, RSA24, SSKS21, SS24c, SW20, SM21c, TMW20, TWR20, TW21a, TWB23, WXP21, WZC^{+}22, WSDW24, XZ23a, XWZC21, XZQX20, YSL24, YM23a, YTZ22, ZFJ^{+}20, ZJZD20, ZCZ22, ZJWV24, ZM21c, ZZW21b, ZWZS22]. Singular/Continuous [GLX22, GLWX24]. singularities [JL22, KWZX23, KX23, LQLF22, RMCG20]. singularity [AFGG21, WLW20]. singularly [ASR22, DLS^{+}23, GLD21, HCL^{+}23, HKCK24, KSR21, LCS^{+}22b, LZ22e, LZ22f, LZ22g, MWDD24, RC22, SR23, SK20, TZ23b, VN20, WCZY20, WSL^{+}21, WS21, YR20a, YLK^{+}21, ZL20a, ZL21a, ZL22b]. sink [Ach21, AKH^{+}22, TMA^{+}22]. sinks [WLY21]. SIQR [ZLD22]. SIQS [XW22]. SIR [AFG22a, AFG22b, LX23a, ZMJ^{+}23]. SIRS [RP20, ZP21]. SIS [EB22, LP22, RT20b, WW22a, XWLL20]. sitters [MT21]. situ [DTK24]. situation [LCS24]. six [DDNS20, PM20, XLYA21]. six-neuron [XLYA21]. six-tuples [DDNS20]. Sixth [WGSS22]. Sixth-order [WGSS22]. size [BAH^{+}21, JGC^{+}21, RWGS22, SM24]. sizes [GK20]. sizing [LLYQ20]. sketch [ZL21b]. Skew [HANP23, Sin20a]. skew-symmetric [Sin20a]. skewed [NHBJ20]. skipping [ZRSK21]. slab [SKS21]. slant [DD20b]. slave [HB20]. Slice [JL24a, JLOR22]. Sliding [KKS^{+}22a, RMCG20, SRZ22, YCL^{+}21, CZ24b, DMD21, DWS24, GZX20, GHP24, GW23b, GX23, HZL21, HCX^{+}24, IFHA21, LLZ20b, LHW21, LLY^{+}21, LR24, LSM23, LTL^{+}21, LNZ^{+}22, LXK23, LR23, MS20, SJ23a, SWSW21, SCG^{+}20, TMW20, WCG^{+}20, WFYL21, WCW22, WM23, XRZH22, YLK^{+}21, ZZQ^{+}20, ZJZD20, ZCZ22, ZQP^{+}22, ZWY^{+}24, ZJW^{+}22, ZC21b, ZT22b]. sliding-mode [WCG^{+}20]. slip [ATDJ22, Azi21, BH22, DM23, KA22, KA23, ZH22a]. slipping [OMS22]. slope [MTRA22]. sloshing [LZY^{+}23b]. slow [Zhu24]. slung [LCL22b, RLY20]. slung-load [LCL22b, RLY20]. Smagorinsky [HJN^{+}24, ZS22c]. Small [BBK^{+}24, CLZ24, GV23, GZZ21, JW21, QLL20, TX21, YW22a, ZJJ24]. small-gain [JW21]. smaller [KMT^{+}21, WMF23]. smallest [KFR20, WC24]. smart [GMH^{+}20, LCS24, ZIML20, MPRS24]. SMC [GZQ^{+}20, YSW^{+}22]. SMC-based-equivalent-input-disturbance [YSW^{+}22]. smooth [ADLU23, BK21, GKK^{+}24, KZWS22, LDCZ24, RC22, RKRRL21, ŚW21, WW20, XYM22, XYKL20, ZDC22]. smoothed [HKU21, LGK^{+}22, ZWJ^{+}24, KPPK23]. Smoothing [GAA^{+}24, CCC20a, JLOR22, NGGD23]. smoothness [RKJ20]. snap [Bab22]. snap-through [Bab22]. snark [LHZ21d]. snarks [SS24b]. snowdrift [Wet21b, ZH23a]. Sobolev [AOSO20, NA21, PBM20, UBSU21, ZM21b, ZLW21a]. SoC [LZX^{+}22b]. SoC-based [LZX^{+}22b]. Social [LLG^{+}24, PWH20, QYL^{+}20, TZL24, AT24, BSCL20, DZL^{+}24, DGL^{+}23, Fan21, FQHS24, GPH21a, GH21, HH24, HWY^{+}22, HC24, HLD^{+}23, HCH24, HCY21, JGC^{+}21, KSZ23, KK23b, KLP22, LZY^{+}23a, LCS23, LCS24, LZZ21, LJP21, LCJ^{+}22, LCC^{+}24, LZZ^{+}20, LF24, LFY^{+}24, MMHX20, QDW22, QZC^{+}24, SWP^{+}22, SM22, TNT23, WP21a, WBY^{+}21, WS22, YLW20, YY24, ZLF22, ZYZ22a, ZPZH23]. soft [XYCG21]. Software [ZR21, FGMS21]. soil [XYCG21]. soilborne [TCGTT21]. sojourn [LTW22, NK20, NK21b]. sojourn-time [NK20, NK21b]. solar [THBPMC22]. solid [ABD23, ALJ24]. solidification [ST24, XAZ^{+}21]. solids [KPPK23, KK24a]. solitary [GA20, Kud20b]. soliton [Alj21]. Solitons [LZYL21, TCZ^{+}20, ZGHM24]. soluble [SS23b]. Solution [AHEGM23, BB20, LLX21, AAS23, ASAK21, AFA24, ACH^{+}22, BSdCCGCV21, BBH21, BG20a, BS22c, BFS20, BSS21, CGSRG20, CCdS^{+}23, CSZ24, DHR22, DMSS20, DHS23, DD21c, EV21, EM21, ED20, FGMS21, FRRS24, FY23, FJJJ21, FCX21, GLT^{+}21, GYC24, HA20, HZJ^{+}21, Har24, IÖÜP24, JLL^{+}20, JSZ20, Jor21c, KWM20a, KWM20b, Kar22, Kod21, Kud20a, LK21a, LZZ^{+}20, LLZ22c, MUOF22, MSN21, MSB22, MD24, NP20, Pop20, QXG21, QXX21, RS21, RNVC21, SO21a, SM21b, SL20a, SK20, SV23a, SS23b, SV21a, TO20, THBPMC22, UUS21, WF23b, WH21, WHY22, Xy20, XHWM20, YM21, YLS23, ZLLY22, ZyZW21]. Solutions [HD22b, INOP20, JMM20, WYD20, WZGL21, WL21a, Alj21, AB21a, BKU20, BDPV22, BR20, BRB24, BFRS24, CPS24, CD22, EW21, EHVMM24, GA20, GB24, GBF21, GLZB20, GH20b, HW24, HSM21, HAFM21, JT20, Kan20, KKS22b, KY24, KNEM23, Kud20b, KL24, KK22, Kum24, Kwo22a, LLT22, LC21, LL22a, LaGBT22, LPS21, LM20b, LR21, LLZW22, LZ20a, LNO23, qLgZ21, LC23, LGS22, LRS20, LMÖ23, MV21a, MKP20, dDMMTH22, MSE24, Ndo23, OPP22, OiEF22, ÖY21, PS21b, RMO22, RRS21, SS23b, SKJM21, SPM22, Sug21, SFC23, SX23b, SB24, TC22, TXZ24, WW20, WJC^{+}20, WMQ21, Wen20, XLJL22, XYM22, wXhGZ20, wX20b, XLZL23, YSHV^{+}23, YR20b, YWT20, YLZL21, YK22, ZLA^{+}21, ZFS22, ZJC21, Zha23b, ZXB20, ACS24]. Solvability [ESA20, MNZ20, YCEW21, BM20, DDE21, DDR21, KY22a]. solve [AS21b, CLH24a, GLWW22, GWH24, HM20b, JGC^{+}21, WLPZ22, YI22, YY22, ZHL24, ZW21c]. solved [BBD^{+}22, UGG^{+}20]. solver [BS20, BDH^{+}23, BM21, CK20, CY21a, EBGN23, LLF^{+}23, LM23, dSLP20, MCT^{+}22, MGK23, PSXX21, QL21, SS21c, SFC23, XFW^{+}21]. solvers [JM23b, LL23a, PGPCK21, QLG^{+}22]. solvers-based [LL23a]. Solving [AF21, BCVV22, CM22, HVYMS21, QYD22, RS22a, WGJK21, AD21a, AM20b, AF24, BD20b, CD20b, Cav22, CU23, Cha22, CWY^{+}20, CdR22a, CCZ24, CHZ24, CDL^{+}20, DS21b, ER22, EBJ20, GW21a, GC20, GM21a, HMS20, HC21, HCY22, HGZ24, HS23b, JL24a, KGPS22, KKLJ20, LGQ20, LW20a, LLY^{+}23, LBX^{+}23, LW22g, LS22a, LM20d, MPM23, MZL24, Mit24, MN23, NA21, NSD^{+}23, PH20, PKJB22, PW22a, QNX^{+}23, QSL21, RPB23, SSCW22, SG21, SLTL20, SWX24b, SL24, TBRB24, TLDL21, TWWL24, UHZ^{+}20, WZCF21, WPE20, WSZ20, XCGN24, YWZT20, YH24a, YAX^{+}21, YLC20b, ZL21b, ZJJB22, ZZLV23, ZL23b, ZLW23, ZO21, ZJX21, ZJX22a, ZJX22b]. Sombor [CZ24a, CGR21, CRS21, DG22a, Das24, LWZ22b, RRS23, RRS24, Sha22, SD22b, ZYLH21]. Some [Cha23, DH24a, DFX24, FM23, GA21, GLZB20, HWLL21, KRT21, LLL^{+}23a, LY21a, MZ22, SY22a, TLDL21, WNP20, XCZ22, ABTR21, AHS20, AC22, CM21a, CSV20, DDR21, GN22, GB24, GLMT21, Hon24, KP23, KT20, LJAM24, LLP23, LL23c, Ma24, Nak23, NX23, NMS23, ÖÇÖ23, SS24b, TNL23, WL21a, XTNF22, YSHV^{+}23]. Songling [GT22a]. sort [GLW22, Guo23, WMT22]. SOSM [LD21b, SMH^{+}20]. sound [LFLL21]. sounds [LLZ^{+}20a]. Source [DJW20, IMS23, Ach21, AzR22, AKH^{+}22, AAHC21, BKU20, DH23, DH24b, HSM21, KKPG20, Kwo22a, Kwo22b, RC22, SSJ22, TMA^{+}22, WWQ22, ZYY20, Zha21b, ZyZW21]. source/sink [Ach21, AKH^{+}22, TMA^{+}22]. South [JC22b]. SP [LOVHC23]. Space [JZZ^{+}22, LPS21, AzR22, BFS20, BFLR23, BZS21, CPS24, CS24a, CTBB24, DL21b, Elg20, Elg22, FTML22, GYM21, HLM23, JHO^{+}21, KY22b, KKS22b, LZRF21, LLS21b, MNZ20, PS22, QL21, SBT22, SFL^{+}22, WF23b, Wen20, WZN22, Wu20b, XZZ20, XJY24, YWZT20, YLS23, YWZ23, YH24a, YF21a, ZJJ24, ZZX20b, ZLW21a, ZZX23, ZLS23]. Space-filling [LPS21]. space-fractional [CS24a, JHO^{+}21, MNZ20, QL21, ZLS23]. space-time [BFS20, CTBB24, DL21b, HLM23, LZRF21, LLS21b, Wen20, YWZT20, YWZ23, ZLW21a, ZZX23]. spacecraft [JZS^{+}24]. Spaces [BBW21, BBH21, CLUR22, DO21, FGM24, GGW21b, GN20a, Gu20, HH22, HGZ24, Ma21b, OK21, QQLZ22, XBC22b, YHD24]. span [AALMLM22, GZRM21]. Spanning [CCHZ23, DGO22, LZ24a, LXLP20, LY23a, LL20a, MY24, QMS22, SM20, SM21a, SMQ24, YLPC22, YLM24]. spark [TLPY20]. Sparse [Grz22, JWY20, KFM21b, ACH^{+}22, CLCY23, DK22a, FMN21, KFM21a, LLLS20, LWL24, ML20, MY22, PLH^{+}24, PX20, SL20a, TLPY20, WHZ^{+}20, WHHD21, XYY23, XBC22a, XYKL20, ZZVL23, ZLC24a, ZGCZ22, ZZB24]. Sparsity [WNZ21]. Spatial [ANG^{+}24, FKVZ21, GSW^{+}20, WX22, BCHRS23, BY23a, BE22, CWY^{+}21, CWY^{+}24b, DZW21, DNR21, FH24, FMHDB22, FLHH23, HZX21, HXZ24, HZJL24, HDQP23, JLWX21, KRNG24, LCS23, LDZ20, LJN^{+}22, LLS^{+}22, LWZW20, LJR20, LWX^{+}22, LZD23, LS22b, MT21, PSXX21, QZWY20, QYL^{+}20, QDW22, QCCW23, QZC^{+}24, RFFB23, SS23a, SLCS20, SCT^{+}20, SWX24a, TLZ20, WP21a, WS23a, WYCL24, YZY22a, YZPL24, YWJ^{+}20b, ZWS^{+}22, ZH23a, ZHQ^{+}24, ZRZL24, ZRMD21, ZHZH24]. spatially [Alf21, GC24, LZ22c, LCML20, STA24]. spatially-varying [GC24]. Spatio [CD23, WZZ23, AMM21, WS22]. Spatio-temporal [CD23, WZZ23, AMM21, WS22]. Spatiotemporal [CXLM21, CWC20, LLZ22a, YMRJ22, ZST22, CGB20, MMR23, SWW21b]. SPDEs [HLM23]. Special [Yua20, BKS21, CPLH20, DBVCP23, KWZX23, ÖÇÖ23, TLPT21, YL24a]. Species [SBV23, Alf24, BCLNV20, BBH22, DWL^{+}24, DG20a, DE21, GWZ^{+}21, JH21, LW21b, MMR23, MS21a, SP23, WBV24]. specific [BJ21]. specified [HDP^{+}21, JY23]. Spectra [Sta20, BEe21, SZZ22, WL22, YWX20b]. Spectral [Ach21, AMZ21, IR22, JT20, KLNN23, LW22a, SRN^{+}23, Tre23, WWK24, YTGC20, AFL24, AHB20, BA20, BOT24, BMS21, But21, CN23, CLZ23, CJ21, DM21, DM24, DYT21, DW20, EM21, EBJ20, Gas20, GZ20c, HLZ20, HHR21, JZ20, KTEM22, LMZ20, LG24, LZ20b, LWL^{+}20, LLCL23, LYN22, MSN21, QLM21, RD23, RGS23, RD24, SM21b, SJ20, STH24, TZJ20, WL23, XZZ20, XHL20, YWZT20, YWZ23, YH24a, YTZ22, YLY24, ZZX20b, ZLJ20]. spectrum [FKK24, PW22b, RS22b, SW24, ZMNR^{+}22, ZZ22b]. Speed [MD23, BCHRS23, KRRS22, MP22, YHL^{+}20, ZZ22a]. speeds [AKMR21]. SPH [AFG21]. sphere [DM23, NKA20, Zho20]. spherical [AS22, GAMP20, GZRM21, VZ23, Zho20]. spheroidal [GAMP20]. spheroids [AAL^{+}21]. spherulite [LTC^{+}23]. spider [KMZ22, SZ21, SZ22]. spiders [TT20, TT23]. spike [NKO^{+}23]. spillover [Gan23]. spin [Ahm21]. spinal [SBT22]. spine [KY21]. spined [YXFL23]. spinning [Ach21]. spiral [HVZZ23, RHR^{+}21, RHN^{+}24, WYLW20]. spline [AM20b, BEIR23, DK22b, FGM24, GKK^{+}24, GK21, KG22, KP24, LS20, LR20, LMS24, MMM24, Rom21, RG22, TO20, WL24b]. splines [ADLU23, CCC20a, CCC23, DKLV23, LMS24, NGGD23]. Split [CDM21, LCL22c, AACGTB23, AACGTB24, CdR22a, ÖC24, RM23a, YS23]. Split-step [LCL22c, CdR22a]. Splitting [EV21, RT20a, ZL23b, AA24, BZ21, CS24a, CG21, FLZZ24, GB20, JL24a, LMX20, LH20a, LW23a, LLFZ20, LWW23, LG22, MF21, RD24, ZV20]. Splitting-based [ZL23b]. splitting-local [CG21]. splittings [SNMM21]. spore [EKKE20]. spraying [Ach21]. spread [GZ20c, HHW23, HCY21, KLP22, LDZT24, LJHL21, PPS23, ZP21]. spreading [CS20, HSLS20, HLD^{+}23, HCM21, LXP^{+}22, LQS^{+}23, LCC^{+}24, NZLW22, PPZW20, PLQ^{+}22, SH22, TAU^{+}22, WJG^{+}20, YZT^{+}21, ZMJ^{+}23, ZLW^{+}21b]. sprouts [AYKS22]. square [BCVV22, CZ22b, DD21b, HLM^{+}21, Lee20, MZX21, OT20, Pap21, RAB24, WHKT23, WZW^{+}21, YMW24]. square-root [DD21b]. square/fictitious [ZH22a]. squared [Kwo22a]. Squares [YCEW21, BD20a, CDH23, DK22c, HM20b, MSN21, RW21, SS21c, VZ23, ZL23b, REC^{+}23]. SRLW [HWCD22]. Stability [AKDR23, Alf21, Alf24, BRB24, CZ21a, CZ22a, FLZ23, FLWL21, GZ20a, GJ21, GWSC23, HZ23a, HQ22, HSY24, JH21, JLJ23, JLD^{+}24, KR22, KK23a, KMD21, LGL20, LZ24e, PGYW23, QYG^{+}20, QLL24, RT20b, SGI20, SGL21, SKSC21, TWR20, WHC20a, WQZ^{+}20, WHL21, WHPQ23, WHY22, XWLL20, XYR^{+}23, YDSW22, ZF24, ZZL^{+}20a, ZQ24, ZLJ^{+}21, ZZS^{+}22, ANS^{+}20, ASS^{+}23a, AMF20, BMY24, BFRS24, But21, CH24a, CHG23, CZ22c, CXYZ23, CLC24, CZXyS24, CHK21, CF23b, CF23c, DZDL22, DE21, DS22b, DCR^{+}20, DRS20b, DHS23, DDH22, DWZ20, DSS22, DL20b, EM20, ED20, FZW23, FZ20b, FF21b, GGY22a, GLY24, GLWL22, GB21, GM21b, GYC24, GW24, EVB^{+}21, HHZZ20, HZJL24, HY24, HH22, ID21, JWK22, JQJY22, JOR^{+}21a, JW21, KZC23, KK20, KLH24, KLPL20, LSPK20, LPP21a, LPP21b, LKP22, LPKC22, LX20a, LHW21, LKSS21, LCXM21, LXL22, LWL^{+}23a, LZZG23, LZZ24a, LT22, LD20c, qLgZ21]. stability [LFLP21, LQSZ21, LNC22, LZZ24b, Ma21a, Mef21, MG21, MRVC20, PP21, PJ21, PKL23, RAB24, RMO22, RSVL21, RM23b, RBPN20, SS24d, SKS21, Shi21, SWX22, SS22c, SKL23, STA24, ST22, SKH20, TlS^{+}22, TTMU20, TM22b, URLC21, VSM23, WWWH21, WWWZ21, WLFX22, WZPL24, XZeF20, XHL23, XYLF21, XZZC22, YZC^{+}22, YWS24, YMW24, YC21, YL20c, YY21c, ZPSP22, ZZH^{+}20, ZZW21a, ZLF22, ZYC^{+}24, ZH20a, ZLX20, ZZS^{+}20, ZCPZ20, ZS21b, ZYZ22c, Zha22a, ZSFS22, ZXZC22, ZLZ24a, ZDZ24, ZLC20, ZYySX23, Zha23b, ZDX21, ZL21c, dOS20]. stability-based [ZLF22]. stability/stabilization [DZDL22, LZZG23]. stabilizability [YSLK23]. Stabilization [FLC^{+}22, FZ22, GGL^{+}23, JSM20, MS22a, PLRD21, WZZ24, YD22b, ZJZD20, AR21b, ASKA20, ASS^{+}23b, BS22b, CLZ^{+}22, CWB22, CLLA23, CFWL24, CWWZ24, DZDL22, DZZ23, DL21b, DDC22, HYÇK22, HWD20, JDLZ21, JDZH22, JHSL21, LPK22, LLZ21, LF21, LLG22, LZC22b, LJ23, LZZG23, LWL23b, LZHZ20, LM20c, LHFW21, LYXZ21, LD21b, LZZ24b, MD22, MZSW21, RH20, SWX^{+}21, SLZ^{+}20b, TDZ^{+}24, WQZ^{+}20, WLFX22, WX21a, WHL22, XWMC20, XSP22, YZC^{+}22, Yan24, YGHW21, YWG24, YYWL23, YZ23, ZeFWW20, ZZS^{+}20, ZZ20a, ZXC20a, ZF23a, ZQC24, ZL21c, ZeF21]. stabilized [Agg22, CU23, CZ22c, HL24b, LW22d, LHFH22, SWX24b, TLJO20, ZF21, ZS20a, ZS22c, ZS22b]. Stabilizing [YWJ20a]. stable [AR24, CFP22, CLZY21, CHY22a, CKL22, DG21, FG24, JSM20, LHYL24, Luo20, MDSS20, PD23, PNOdV20, RVV24, SZ24a, SK20, WXHZ23, YSYS24, YWXL21, ZSS24, ZGZ^{+}24]. stacked [DWT^{+}21]. Stackelberg [WY23, ZS22d]. stage [GZLZ24, JNY^{+}21, LLYQ20, QLL24, QDL^{+}21, SNMM21, WSG^{+}21, XYKL20]. stage-structured [QLL24]. stages [HH20, RT20b]. staggered [BRMVCD21, DL20a]. Star [HLM^{+}21, CLL23, DYZC20, LXLP20, LY21a, WZL22, YML21]. starlike [AYS20, LG24]. stars [LLW20, XLL24, ZWC23]. Start [GAR22]. Start-up [GAR22]. State [LL23b, MTRA22, MXY^{+}20, NTT21, SWS^{+}22b, ZSM20, ZZCL20, CL22, CWZL23, CGZ^{+}22, CHW^{+}24, CSZ24, DRS20b, DZZ^{+}20, DZX23, FZLF23, FZW23, FSLZ23, GY22a, GSWL23, GLY24, GXZ^{+}24, GM21b, GW21b, GW24, HLLW21, HRLH21, HSZ20, HL23a, HZX22, JNY^{+}21, JQJY22, JW21, LLZ20c, LLL21, LLZ21, LLY^{+}21, LZ22b, LHC^{+}22, LL22b, LM24, LR20, LWG20, LZLZ21, LWL21, LWH22, LFM^{+}22, LZZ^{+}23, MXF^{+}24, MZYW23, NGS22, OPL22, PWZ22, PX23, QLZL22, RD22, Son24, SSW21, SM21c, SLPH22, TWL20, TQZ^{+}24, TW21b, VHB23, WSS22, WC22a, WF23a, WB23, XHWM20, YM23a, YZPL24, YWG24, YWT20, Yoo21, YZ23, ZeFWW20, ZCT22, ZM21c, ZDX21, ZC21c, ZTW22, ZLLX22]. state-dependent [DRS20b, GW24, LZ22b, PWZ22, YZ23, ZC21c]. state-feedback [LZZ^{+}23, OPL22]. state-flipped [DZX23]. state-function [GSWL23]. state-saturated [CHW^{+}24, LHC^{+}22]. States [CFA^{+}23, BF22, Fan21, HGJ21, KPG22, LMT24, LWL^{+}23a, MK21b, SS24d, TWZ^{+}20, WXW^{+}21, ZZX^{+}22, dAGSFM22, LLY^{+}22b]. Static [JGXL23, ZCS^{+}22, DZZ^{+}20, GS21, HRA^{+}20, LBJ22, TWL20, TW21b, ZXC20a]. Stationary [GLL23, HZJ^{+}21, LJH^{+}20, LLFZ20, LWJ21, SK24a, SJZA22, CPPL23, GGBEZRI23, JY20, JH21, RP20, SBB^{+}22, SWW22, ZHJ^{+}21]. Statistical [RS24, ZJC21, ACVHC22, BCVV22, GGBEZRI23, LOVHC23, MJW21]. status [LZ21]. Steady [KPG22, GCGM21, LR20, LWL21, LHFH22, MUOF22, MK21b, SS24d, SX23b, SWX24b, YOF21]. steady-state [LR20, LWL21]. steel [LLYQ20]. steep [YWT20]. Steepened [SSJ22]. Stefan [KR20, NS22, Vyn23, XAZ^{+}21]. Stefan-type [Vyn23]. Steiner [LLSY23, YL20b]. Steklov [MM20a]. stencil [MT23]. stencils [Qi22]. stenoses [PM20, RB21]. stenosis [PPS24, RRR19, RRR20]. step [ADS23, BMJ20, CdR22a, DS21a, KSW24, LCL22c, MMS21, RM23a, SD22a, SK20, TLJO20, TWWL24, WHHD21, Zha22a, ZZVL23, ZS22c, ZS22b, ZDKC20]. stepped [SLHX21]. stepping [GdFPF24, HXF23]. stepsize [LZL^{+}21]. Stepwise [DS20, AYS22, BAH^{+}21]. Stieltjes [ESA20, SWF20]. stiff [CY21a, DHR22, FY23, YMRJ22, ZF24]. stiffness [TM20]. stimuli [RHR^{+}21]. stirred [XYZY20]. stirring [TMA^{+}22]. STL [TGDPS22]. Stochastic [CFF20, GLP20, KKPS21, KDB^{+}23, LXK23, MXY^{+}20, SWX^{+}21, SWX22, TMW20, TLY24, AAJ23, ANS^{+}20, AM20b, AS21d, AB20b, BLS21, BSCL20, CZWS20, CSS^{+}22, CZ21a, CZ22a, CZ22b, CFWZ20, CZS21, CYL^{+}22, CdR22a, CLLA23, CZ24b, CGMFSHR21, CDL^{+}20, DS21a, DD24a, DDLW21, DZDL22, DP21, DRS20b, DN21, DZZ^{+}20, DS21b, ER22, ES24, FLZ23, FSS20, FLWL21, FLC^{+}22, FZ20b, FLH^{+}21, FLLW23, GLGLH24, GNSZ22, GLL23, GXZZ20, GZLY24, EVB^{+}21, HWD20, HZJ^{+}21, HRLH21, HX22, HQ22, HLW^{+}20, HOM22, HWWW23, IÖÜP24, JW23, JH21, JDZH22, JWK22, JZP^{+}20, JQJY22, KY22a, KK23c, LP22, LH20b, LHC21, LNS^{+}21, LLG22, LZZG23, LSCS21, LD20d, LB20, LJH^{+}20, LLD21, LWLA21, LLX21, LWH22, LNC22, Liu22b, LY23c, LH24b, LWJ21, Lu22, LZ24e, MKP20, MMG^{+}23, MSN21, MN22, NGS22, PLRD21, PAY22, PX23, QZH21, QQWJ22, RS23, RAB24, RP20, RHNT21, RM23a, RM23b, RS21, RNVC21]. stochastic [RF21, RD22, RW21, RBPN20, RSA24, SO21b, SO21c, SK24a, SZG21, Shi21, SJZA22, SLZ^{+}20b, SZP21, SWSW21, SGL21, SZZL22, SZSZ22, SWZS20, SDW^{+}22, SSW22, SJ23b, TLHH21, TlS^{+}22, Ten22, TZD22, WHC20a, WZXY20, WZZX21, WWY21, WC23, WY23, WL24a, WLZY21, WH21, WHD^{+}23, WHY22, WYLZ24, XGRM20, XZS20, XKW^{+}23, YZC^{+}22, YLK^{+}21, YLL22, YXMW22, YL22, YMW24, YWS^{+}20, Yi23, YDSW22, YZY22b, YM22b, ZDGL21, ZZX20a, ZY20, ZZS^{+}20, ZZL^{+}20a, ZYW21a, ZCJL21, ZY21a, ZQP^{+}22, ZYZ22b, ZLD22, ZXZ24, ZDZ24, ZYX21, ZCLZ22, ZYySX23, ZS22d, ZZS^{+}22, ZWHY20, ZHJ^{+}21, ZC21b, Zhu24, ZPW22]. stochastic-local [KK23c]. stochastically [GHS^{+}21, KEiS23]. Stochasticity [TAU^{+}22, CLC20]. Stokes [Azi21, BFRS24, BDRM23, CHY22a, DG21, FKS23, HYÇK22, He23, HKL22b, JLJ23, LLS21a, LHFH22, MCT^{+}22, MVO21, RM20, SLZ20a, SY22b, SWX24b, XHL20, YGM^{+}22, YL23b, ZM20, ZRL^{+}24, ZZX21, ZS22b]. Stokes/Darcy [XHL20]. stop [WHD^{+}23]. stop-and-go [WHD^{+}23]. stopping [GYM21]. storage [BO21, JSCB20, QZH21]. Storey [GLP22]. Storing [ALM22]. Störmer [CY21a, Zha20]. strain [ASS^{+}23a, BP23, JA23, KET22, Liu22b, ME21a, NTLNX24, SHWW21, TNTL21, XZSS21]. strains [CTZ21]. strand [ZWZ^{+}22]. Strang [CS24a]. Strategic [BSCL20, GYC^{+}22, JTW^{+}23, YY20b, Elg20, Elg22, INT20, LLW24]. strategies [CD20b, DDGN23a, FLHH23, GZ20a, ISM^{+}23, JM23a, KIJK20, LLY^{+}22b, MZW^{+}21, MCSD21, NGC21, RYG21, Sek23, SK24b, TG20, Ued23, WCS22, XWLL20, ZYPL20, ZHJ^{+}21]. strategists [GWZ^{+}21]. strategy [BY23a, CSS^{+}22, CWZ20, CH24c, DKK20, DZ21, GPH21a, GLWW22, GZH22, GH21, HH24, HRLH21, HWZZ23, IFHA21, INT20, JC22b, JKJC21, LCS23, LLY^{+}23, LLZB24, LSM23, LCH20, LJR20, LNZ^{+}22, MH24, SX23a, SWS^{+}22b, WWY^{+}20, WCZ20, WWW^{+}21, WLZW21, WDTF22, WWZW23, WYS23, Xy20, YXAY21, YP21, YP22, YSW^{+}21, YZY22b, ZLF22, ZCH20, ZHQ^{+}24, ZWJ^{+}20]. strategy-updating [LJR20]. stream [JRP22, MB24]. streaming [OKN^{+}22]. streamline [LZ22e]. StreaMRAK [OKN^{+}22]. streams [NHBJ20]. street [BS23, IMOR24]. strength [XPP^{+}20, ZWHY20, ZBA^{+}23a, ZBA^{+}23b]. stress [DM23, Has20, MCB^{+}21, MS22c, PJMW22, TCPF24, WPE20]. stressed [PS21a, SSC22]. stresses [TC20b]. stretch [LL20a, QZ21]. stretched [CSG^{+}22, CSG^{+}23, SKK^{+}21]. Strict [WSXL22, SYTC24, WZXH22, Yi23]. strict-feedback [Yi23]. strictly [JOR^{+}21a]. String [CF23b, CF23c]. strings [KH23]. strip [DWS24]. strips [CH24a]. striving [DW22]. Strong [BMY24, DD24a, GZZ21, JL20, NMS23, RM23b, ZYX21, DLQY20, GYC24, GLW22, Guo23, HCJC21, LXLP20, LL21, LLLL23, LZZ22, MZ23b, You20, YK22, Zha23a, ZT24]. strongly [ARST21b, HZF21, Hu20, LX23b, LHL24]. Structural [CPPL23, GNZ20, LQSZ21, PJ21, SS22a, GMT20, HXL^{+}20, LCW^{+}23, LWL21, MCB^{+}21, QZ22, SHWW21, WDYS21, ZGG^{+}22]. structurally [GGL^{+}23]. Structure [BZ23a, KSW24, LM20a, LZRG21, AV21a, AV21b, BS22c, BK21, CHH22, Chy22d, Chy22c, DGP^{+}23, DBVCP23, Fan21, GN20a, GLW22, HZZ^{+}21, LY23a, MKKG22, Miy23, PS21a, PW22a, SSC22, SSCG20, SC20b, SZXL24, STH24, WCZ20, WMT22, WHWW21, Xu23, YHSZ20, ZW21b]. structure-preserving [DBVCP23, PW22a, WHWW21]. structured [AALMLM22, FLZZ24, FCX21, Gan23, GGY23, JC22b, JKJC21, LCS22a, LFA22, dSLP20, QLL24, QCW24, SRK23, Shi21, Wan24a]. structures [AK21, AR22, BCLNV20, CK21, Eva20, JGXL23, LZY^{+}23b, LZ24e, MZX21, MZX22, MK21a, NKO^{+}23, Pop20, QL20, SHWW21, SZCP21, XKW^{+}23, ZZY^{+}21]. stubborn [KSZ23]. student [ACMT23]. studies [MMMBRS20]. Study [GK20, GY22b, AMM21, BS22a, BS22c, BM24, BE22, DJWX21, DZZ^{+}21, FB20, FMHDB22, GLP22, HHR21, KCJ22, KZ22, KOP21, LD20c, LJHL21, MVCK24, NGC21, NHK^{+}22, NHK^{+}23, PM20, RB21, SCD23, Sha21, SKL23, SSCG20, VS24, XYCG21, ZZX20a, ZLW21a]. studying [JG21]. Sturm [BBHM21, Bon22, HBCA21, MBMA20, ZL20b]. sub [AD21a, NP20, QXX21]. sub-diffusion [AD21a, NP20]. sub-matrix [QXX21]. subarachnoid [SBT22]. subcell [CH24c]. subconnected [QLZL20]. subdiffusion [WC21]. subdivision [CLUR22, GI23, JY20, JOR^{+}21a, LU24, Rom21, Vis23]. subdivisions [HLSL22]. subgradient [ZL22c]. subgraph [FMT^{+}21]. Subgraphs [HM20a, CLW21, CLWZ21, Prz24a, Prz24b, TNL23, WW23]. subject [AS20, AAA20, Azi21, CJ22, DG20a, DMSS20, DWZZ22, FJLG22, HYT21, HD22b, IFHA21, KGCC20, LZWC22, LTL^{+}22, LNC24, LLZB24, LZZL21, LXW^{+}20, LD21b, LYL22, LFM^{+}22, MV21b, MV23, QZH21, SAB23, SMH^{+}20, WCM^{+}21, WXP21, WPJ22, WYS23, XWL^{+}22, YGHW21, Yi23, ZDGL21, ZL20d, ZSP24, ZLP21]. Sublinear [Che21]. submerged [MMK^{+}23]. submodular [SYW21]. submodularity [CM21b]. submultigraphs [GW23a]. subsequence [FQ22]. subsidy [CLD^{+}24, HH23a, JJL23]. subsonic [LZ24c]. subspace [CG20b, GGW21b, LWL24]. subspaces [GGW21b]. substitution [LZRF21, ZRL^{+}24]. substructure [GLW22, LZRG21]. subsurface [AMBG^{+}20]. subsystems [FLZ23, FZ22, GLY24, JSM20, ZC21c]. subtree [LYhH^{+}24]. subtrees [SLL^{+}22]. Successful [CWY^{+}24a, SC20c]. Successive [CWY^{+}20, YWR^{+}24, AM20b, PWH20, WLPZ22, ZV21]. sufferers [ZZPF23]. Sufficient [DML21, SM20, GGBEZRI23, KEiS23, LL21, Wu20a, ZCW24b]. Sugeno [CFWL24, DMRR22, DL20c, MZSW21, NT24, Yan24]. suggestion [GYC^{+}22]. Sum [HCS23, CLL21, Che23, CHY22b, DDPS22, DDH^{+}24, FQ22, IKÖ24, MZLS22, NTNT24, SJ23b, TLY24, WFKL24, XTS^{+}22, ZCJL21, ZQZ21]. sum-connectivity [Che23]. summation [HSY24, PKL23, PN24]. sums [DF20, LQW20, LZ24b]. sun [RGS23]. Super [SYX22, WLZW21, XJ20, CFP22, CG20a, CWZL23, GNA22a, GNA22b, HCS20, Kun24, WWZW23, ZDC22, ZY21b]. Super-convergence [SYX22]. super-convergent [CG20a]. super-linear [GNA22a, GNA22b]. Super-rational [WLZW21, WWZW23]. super-supraconvergent [CFP22]. super-twisting [CWZL23, ZDC22]. Supercloseness [LZ22e, MZ23a, LZ22g]. superconductors [WS21]. Superconvergence [CN23, SL20b, WHY20, LLS21a, PZS24, SZ23, SZ24a, WS21]. Superconvergent [ARST21a, KK24b, HL24b, RG22, SW20]. supercritical [Zha24b]. Supereulerian [DLM20, DML21, LZ24b]. superiorization [AACGTB23, AACGTB24]. Superlinear [NL20, LHC21]. superposition [WPE20]. superpositioned [SS24b]. supervised [ACMT23, LLZ^{+}20a, LFLL21, SH21]. supervision [HJJL24, MZ24, SWZZ23, ZBL^{+}23]. supplementary [YXAY21]. supply [MKZ22, RST21, She20, YLX^{+}21]. support [Har22, LLZ^{+}20a, LZ22d, Zha24a]. supported [Bab22, BSS21, LWZ22c, ACBLC24]. suppress [SWZZ23]. Suppressing [LCS24]. suppression [Ahm21, FAEZ24, PPS23, Wan21a]. Supra [YTGC20]. Supra-Laplacian [YTGC20]. supraconvergent [CFP22]. sure [BMY24, CFWZ20, GGY22a, RM23b]. Surface [PS22, AKH^{+}22, BM22, BP23, CSG^{+}22, CSG^{+}23, DSZ^{+}23, DD20b, HRLH21, HL23a, HZX22, KET22, KA22, KA23, KKPG20, LD21a, LTL^{+}22, LFM^{+}22, LWY^{+}24, MM22a, MM24, MK21a, NTLNX24, OMS22, SIK22, WL24b, XYZY20, YFHC21, YXXZ23, ZQ23, ZXF20]. surfaces [FKVZ22, HBWQ20, HXF23, KSPC22, MDJW20, RKRRL21, Tre23, YA21b, YL24a]. surfactant [SS23b]. surplus [OLW23]. surplus-dependent [OLW23]. surrogate [Mar24]. surrounding [IMIH22]. survived [YTX^{+}24]. Susceptible [ASAZ23, WQZD23]. susceptible-infected-recovered [WQZD23]. Susceptible-Infected-Susceptible [ASAZ23]. susceptibles [LX23a]. susceptibles-triggered [LX23a]. suspension [ZYySX23]. suspensions [FF21b]. Sustainability [BBAS21, FFL^{+}23b]. sustainable [CHP23, She20, TCGTT21]. sustained [ZB23]. sustains [LWZW20]. SVEIS [WLG^{+}21]. SVIR [GLGLH24]. SVM [SL24]. swarm [BGRM21, LLT22, LS22b, TGL^{+}24, WYCL24]. swarmalator [ANG^{+}24]. sways [Kur21]. sweeping [FZB24]. Swift [LHYL24, WYD20]. swimming [SKK^{+}21]. swing [RLY20]. switch [TO20]. switchable [FN22]. switched [CZK^{+}22, CLWX23, CWWZ24, CWJF21, CZGX22, DZDL22, FLZ23, FJFS21, FZLF23, FZLJ24, FZ22, GZX20, GLY24, GS21, GB21, GLD21, GZWX23, HZW21, HLLS22, HMZZ20, JY21, JPZ^{+}21, JSM20, KZC23, LX20a, LLZ21, LLY^{+}21, LJZL22, LTW22, LZ22b, LWL22, LGC^{+}24a, LT22, LZHZ24, LZHZ20, LZW^{+}20b, LNZ^{+}22, LWH22, LCL22a, LH24b, LZ24d, LLWW22, LZ24e, MFS22, MZFX20, PWZ22, QYG^{+}20, QZH21, QQY^{+}23, SZ24b, SLZ^{+}20b, SWZS20, SZWL21, SM21c, SM23, WXLZ21, WXP21, WCW22, WM23, WLHW23, WM24a, WQT^{+}24, WX21a, XZZW21, XZZC22, YGZ^{+}22, YCL^{+}21, YM23a, YMW24, YGHW21, YWJ20a, YL20c, YD22b, YZY22b, YZ23, ZeFWW20, ZSM20, ZS21b, ZGLZ22, ZXZ20, ZDX21, ZC21c, ZCZW23, ZL22e, ZJLT23]. switches [ZH20a]. switching [CZ20, DL20c, FJLG22, GHP24, HWD20, HZW21, HLW^{+}20, JP23, JWC20, JZP^{+}20, JQJY22, LZ22b, LZWC22, LL22b, LZCL23, LB20, LLD21, LSL^{+}21, LCL22a, MZY^{+}22, NSSA20, PRP24, PWZ22, SLZ^{+}20b, ST22, WCZ20, WQZ^{+}20, WLZF21, WZY21, WJL^{+}24, WSL^{+}21, WCZ^{+}22b, XCW^{+}22, XWWC21, XZZC22, XZQ23, YM23a, YGHW21, YZ23, ZCJL21, ZDZS21, ZQP^{+}22, ZYXL23, ZZ24, ZYX21, ZGLZ22, ZLJ^{+}21, ZHJ^{+}21, ZDX21, ZC21c, ZLP21]. switching-like [GHP24]. sword [Zhu23]. Sylvester [CHZ24, FCX21, HM20b, HCY22, LM20d, NSD^{+}23, XLJL22, YM21, ZW21c]. Sylvester-conjugate [YM21]. Symbolic [KFM21b, KFM21a, SĆL^{+}21]. Symmetric [UU21, YXFL23, YF21b, AACI21, BFGGM24, CET22, CD20a, DS22c, HLX21, KP22, LR21, LL21, LYJM21, LZC22a, qLgZ21, PM21, SD22a, Sin20a, SS24e, TCZ22, Ued23, WDYS21, Zha23a, AK21]. symmetric-conjugate [CET22]. Symmetric-Toeplitz [AK21]. symmetries [AEL21, HSM21, IHF13, Woo20]. Symmetry [NT23, Tak21, CTWG24, DYA21, SS23b, ZL22d]. sympathy [QSW23]. Symplectic [QHC21, QQ23, BIJ22, DD24a, GGW21b, GYM21, JT24, UKA23, WLPZ22, Zha22b]. sympletic [GGW21b]. synaptic [WHW^{+}21]. Synchronization [GWY20, HHY20, RJHC24, SCG^{+}20, SHYS23, WZLC21, WGMJ23, XWM23, YLLL20, AAJ23, Ahm21, BZZ21, CFWZ20, CYHR22, CL24a, CLH24b, CLLA22, DTWJ21, DZZ^{+}21, DL21c, DZX23, FJFS21, FZS21, GBH^{+}21, GXZZ20, GL22, GX23, HKKL22, HLL^{+}21, HP23, HWL24, HB20, HB24, HXWC24, JYMP^{+}20, JWC20, LZH^{+}20, LGL20, LHZ^{+}21a, LCXM21, LCY^{+}21, LM20c, LXW^{+}20, LK21b, LZH21, LHH24, MAAEF21, NPJ^{+}20, PNJ^{+}21, PSE24, PN24, RDKR22, SX23a, SZCC23, SWS^{+}22a, SY21a, TlS^{+}22, TWB22, VHB23, WCG^{+}20, WGL^{+}21, WA22, WZLH22, WLHW23, WB20, WHCW20, WCZ^{+}22b, WCX^{+}22, XPP^{+}20, XYH^{+}24, XLZ^{+}24a, YLX^{+}21, YLHZ22, YWR^{+}24, YZAM20, YZ21, ZSC^{+}21, ZLK^{+}22, ZCS^{+}22, ZC23a, ZLZ^{+}23a, ZHMZ21, ZW22b, ZWHY20, ZWYL22, ZB22]. synchronous [LLY22a, XXJtC24, ZZLV23]. Synergistic [LXF20]. synthesis [CLL24, FFL23a, LNS^{+}21, NK21a, XSP22, ZLZ24b]. synthesizing [QLC^{+}24]. System [MQ22, WSXL22, XZ23b, YXYZ23, ASR22, Ahm21, AR21a, AFGG20, AKR21, AVJ24, AMF20, ADS21, BPG24, BDPV22, BSCL20, CCCdLMR23, CJCL21, CGB20, CKN21, CN23, CXLM21, CM20b, CLCY23, CZL20, CWC20, CLZ24, CZW24, CLL24, DP24, DJWX21, DYA21, DL20c, ESA20, FZLJ24, FJJJ21, GT22a, GWSC23, GDM^{+}20, GL24, GLLZ24, HTF20, HLW^{+}21, HQK21, HKL22b, HLZ22, HDQP23, Hua24, HSL^{+}24, JC22a, JT20, JSM20, KMLW22, KB21, KIB20, KMD21, KK24b, LePP24, LaGBT22, LXCZ21, LR21, LYN^{+}21, LX22, LJ23, LZ24c, LL20b, LCML20, LZL^{+}21, LCL22b, LWX^{+}22, LSC^{+}23, LL23d, LG22, LW21b, MSZ22, MZY^{+}22, MB24, PGKM24, Pap21, PPS23, RC22, RLY20, RBPN20, SCD23, SD22a, SMTT23, SYTC24, ySyS22, SB24, TZW22, TTCY23, TWB23, THBPMC22, TCZ^{+}20, TXZ24, VHB23, WYW^{+}20, WBV22, WWW^{+}22, WZC^{+}22, WLD^{+}22, WM23, WL23, WWW24a, XCLW21, XJY24, YR20b]. system [YTZ20, YWT20, YWJ20a, ZL22a, ZZX24, ZCS21, ZH20a, ZJZD20, ZWWH21, ZJJB22, ZZ22a, ZXZC22, ZLLY22, ZrLkL^{+}23, ZXZ24, ZWY^{+}24, ZCW^{+}24a, ZYySX23, ZGZ^{+}23, ZWFW22, ZZX^{+}22, ZT22b, ZZW^{+}21c, KY24]. system-level [CLL24]. systematic [Guo22, HR20a, SNS^{+}20]. Systems [AKDR23, GLP20, HD20, KFM21b, KSS23, LWL^{+}23c, MWS23, MXY^{+}20, SBV23, WSXL22, XWDK23, YFW21, ZCL^{+}23, AS23a, Alf24, AFGG20, AMZ21, AB21a, AR24, ASKA20, ASS^{+}23b, AS21d, BRFLGH21, BRTR21, BG20b, BRB24, BTP20, CZLH20, CZWS20, CDZW21, CSS^{+}22, CZ22a, CZLW24, CFP22, CL22, CMYF21, CZ22b, CJ22, CWH24, CHZ^{+}20, CP20, CY21a, CNZ^{+}21, CZS21, CYL^{+}22, CLZ^{+}22, CLC^{+}22, CZK^{+}22, CC22, CGZ^{+}22, CLWX23, CCL^{+}24, CTF24, CHHC24, CSS^{+}24, CHW^{+}24, CQ24, CLSN24, CZ24b, CSL24, CZ20, CL24b, CWWZ24, CZXyS24, CHST20, CDR22b, CZWX21, CWJF21, CXGL22, CZGX22, DD24a, DZL^{+}22, DZDL22, DZZ23, DCR^{+}20, DRS20b, DHS23, DXZ^{+}23, DLS^{+}23, DYX24, DDC22, DSS22, DL20b, DW22, DWZZ22, Du24, EW21, EBJ20, FJLG22, FLZ23, FT24, FLZZ24, FRW^{+}24, FLWL21, FLC^{+}22, FZLF23, FFL23a, FZW23, FSLC22, FZ22, FLLW23, FR24, FRR^{+}22, GZX20, GZQ^{+}20, GNSZ22, GSWL23, GZCS23]. systems [GWZZ23, GZL^{+}23, GLY24, GSS24, GLWL22, GC24, GI23, GS21, GB21, GS24, GBCP21, GLTY22, GW21b, GSRL20, GZ20b, GSL^{+}22, GHP24, GMN22, GLD21, GHS^{+}21, GZWX23, GW23b, GJS^{+}23, GW24, GZLY24, GH20b, GT22b, HC20, EVB^{+}21, HREB21, HWD20, HZW21, HQS^{+}24, HKL22a, HL24a, HRLH21, HW24, HT20, HLLS22, HLL22, HLKC22, HL23a, HCL^{+}23, HKCK24, HZX22, HLW^{+}20, HB20, HYT21, HH22, HJW^{+}23, HSY24, HMZZ20, HAFM21, HRA^{+}20, HS23b, INOP20, JY21, JLP23, JDLZ21, JDZH22, JCCC20, JWK22, JZP^{+}20, JQJY22, JW21, JPZ^{+}21, JMC23, JWL^{+}24, JTC^{+}24, JJSY22, JHSL21, KRNG24, KGCC20, KZC23, KMUY21, KKPS21, KKPS23, KKS^{+}22a, KBL^{+}24, KNN20, KFM21a, KWW22, KSW24, Kum23, KKH23, KLPL20, KJL20, LSPK20, LKP22, LPK22, LZY^{+}23a, LPS21, LLZ20b, LCLS20, LMX20, LDX20, LSWL20, LX20a, LWM^{+}21, LWZ21, LLZ21, LLY^{+}21, LJL21, LNS^{+}21, LJZL22, LTW22, LZ22b, LGX22, LHC^{+}22, LLG22]. systems [LZC22b, LZWC22, LHX^{+}22, LJY22, LLY22a, LWL22, LDXZ23, LHZ^{+}23, LWL^{+}23a, LZZG23, LWL23b, LCZW23, LD23a, LHW^{+}23, LL23b, LR24, LM24, LDCZ24, LZZ24a, LNC24, LLZB24, LGC^{+}24a, LZZL21, LT22, LSM23, LLWY22, LZHZ24, LD20c, LZHZ20, LM20c, LZW^{+}20b, LLLS20, LLD21, LSWS21, LTL^{+}21, LYXZ21, LZLZ21, LWL21, LWLA21, LNZ^{+}22, LWH22, LYL22, LFM^{+}22, LW22g, LCL22a, LCS^{+}22b, LZZ^{+}23, LLWK23, LXZ23, LXK23, LH24b, LLQW24, LZZ24b, LLWW22, LJWD21, LJW^{+}22, LR23, LX24a, LZ24e, LY20b, LLT^{+}20, LY22d, MPX^{+}20, MWZX21, MJZ21, ML22, MTFD24, MXF^{+}24, MTRA22, MS20, MR22, MMHD20, MFS22, MD22, MZFX20, MPC^{+}24, MZLS22, MC20, MYY21, Mon20, MZSW21, MZYW23, NK20, NK21b, NK21a, NTT21, NK22, NSSA20, NGS22, OPL22, PHL23, PKL23, PCN^{+}24, PHS^{+}20, PLRD21, PWZ22, PX23, PGYW23, PGZ^{+}24, PKG^{+}20, PGPCK21, QYG^{+}20, QZH21, QQWJ22, QQY^{+}23, QYLW24, QXSW22, QQ23, RRS24, RF21, RD22, RX20, RFFZ21, RJHC24]. systems [SCLD21, SAB23, SS20, SNS^{+}20, SKLS22, SNRA20, SJ23a, SZG21, SMH^{+}20, SGX21, SX24, SZ24b, SWF20, STA24, SPZS20, SC20b, SLZ^{+}20b, SC20a, SZP21, SY21b, SWSW21, SZZL22, SZSZ22, SKJM21, SPM22, SWS^{+}22b, SWZS20, SZWL21, SM21c, SCHS22, SLPH22, ST22, SDW^{+}22, SRZ22, SM23, SKH20, TCGR24, TLHH21, TlS^{+}22, TDZ^{+}24, TMW20, TCA23, TWR20, TW21a, TZ21, TZD22, TTMU20, TTL20, TPTA24, VMZDV20, VSM23, VSM^{+}21, VSM22, WHH^{+}20, WS20, WLQX20, WCZY20, WHLL20, WCZ20, WXW^{+}20, WQZ^{+}20, WJ20, WXLZ21, WXW^{+}21, WFYL21, WCM^{+}21, WZZX21, WXP21, WLY21, WSS22, WA22, WZLH22, WCW22, WDTF22, WXF22, WYZX22, WPZY23, WC23, WZW23, WS23b, WY23, WHPQ23, WLL^{+}24, WJDL24, WYLQ24, WM24a, WQT^{+}24, WJL^{+}24, WZZ24, WFKL24, WHWW21, WLZY21, WPX^{+}20, WCZ^{+}21, WSL^{+}21, WZXH22, WHL22, WSSX22, WWW24b, XD21, XWMC20, XXL^{+}21, XH22a, XLZ23, XZeF20, XCW^{+}22, XZZ22b, XSP22, XTS^{+}22, XWZC21]. systems [XHL23, XWWC21, XXW^{+}20, XZQX20, XZZW21, XLLH22, XWL^{+}22, XZZC22, XZQ^{+}22, XWL23, XZQ23, XZ23c, XLZ^{+}24b, XRZH22, YMRJ22, YYG20, YZC^{+}22, YWY^{+}22, YLC20a, YLCZ20, YY20a, YLK^{+}21, YCL^{+}21, YLL22, YXMW22, YM23a, YWS24, YWC24, Yan24, YGHW21, YW22b, YDXH22, YXW23, YWG24, YL20a, YAX^{+}21, YCZ^{+}23, YXZ^{+}24, Yoo21, YP22, YYWL23, YL20c, YM20, YZ21, YZDC22, YD22b, YMYR21, YZY22b, YZ23, YZ24, ZPSP22, ZMNR^{+}22, ZZH^{+}20, ZZW21a, Zen21, ZDGL21, ZWH23, ZTLL24, ZFJ^{+}20, ZLX20, ZL20d, ZZS^{+}20, ZCPZ20, ZZ20a, ZZL^{+}20a, ZSM20, ZZCW21, ZDZS21, ZL21b, ZZLL21, ZLS21, ZS21b, ZWZ^{+}21, ZCZ22, ZYSW22, ZHG22, ZQP^{+}22, ZW22a, ZDC22, ZD23, ZWW23, ZSP24, ZLZ24b, ZLZ24a, ZDZ24, ZO21, ZPDL20, ZLC20, ZM21c, ZLZ^{+}21, ZJW^{+}22, ZCLZ22, ZPY22, ZGLZ22, ZYySX23, ZY23b, ZXZ20, ZLJ^{+}21, ZZS^{+}22, ZZ20b, ZDX21, ZC21b, ZC21c, ZTW22, ZCZW23, ZQC24, ZWZX21, ZL21c, ZZW^{+}21c, ZCJ22, ZWY^{+}22, ZWZS22, ZL22e]. systems [ZJLT23, Zhu24, ZD24, ZPW22, ZXW20, dOS20]. Szász [AM20a]. Szeged [BDTP23, CH22a, DESTP22, GV24, YJL20]. Szeged-like [DESTP22].

T [Hua24, CSS^{+}22, GLWL22, LPK22, LR24, LR23, NK22, NH22, SJ23a, SWCT21, YI22, ZJZD20, ZWZS22]. T-cells [YI22]. T-product [Hua24]. T-S [CSS^{+}22, LR24, NH22, ZJZD20]. tacit [SM22]. tackle [NLP^{+}22]. Tackling [LWL^{+}23c]. tag [JLWX21]. tag-mediated [JLWX21]. tail [JS22]. tailored [WC21]. tails [CX22]. Takagi [CFWL24, DL20c, MZSW21, Yan24]. Takenaka [SKR22]. tamed [LNC22, LW24]. tangency [MBL20]. tank [XYZY20]. tanks [LZY^{+}23b]. tapered [PPS24, RRR19, RRR20]. tardiness [AAA20]. tardy [DTP21]. Target [XZ23b, AFG22a, AFG22b, HKU21, SRZ22, ZYCY22]. targets [LLWY22]. task [LLCS23]. tasks [CYG20]. Tasso [Alj21]. taxation [HH23a]. taxis [FBT21, IMOR24]. Taylor [BCP23, GE20, JNY^{+}21]. TBiCOR [HCY22]. TCORS [HCY22]. TCP [LCD^{+}20]. TCP/AWM [LCD^{+}20]. TD [LQLF22, LWH22]. TD-BEM [LQLF22]. teacher [ACMT23]. teacher-student [ACMT23]. technique [CM22, CCE^{+}22, CSC20, CTBB24, DYW^{+}20, FAEZ24, IS21, JS20, KWW22, Lee23, LQNA23, NA21, PHL23, RDL23, RG22, SLTL20, VPK20, WXLZ21, YSL24, YI22, ZYC^{+}24, ZZL^{+}20a, ZDKC20]. Techniques [GLX22, BM24, GLWX24, KKLJ20, LLLC23, MPRS24, WCX^{+}22, dAGSFM22]. technological [ZM21a]. technologies [MMG^{+}23]. technology [WWW^{+}22]. telegraph [DMSS20]. telegrapher [CZR21]. tellurite [AHS20]. Temperature [FN22, IMS23, ZLN20]. Temperature-dependent [FN22]. Tempered [DEHFF24, SLZ20c]. Temporal [TZ24, AMM21, CD23, LZZ21, LLLC23, WS22, WZZ23, YFM^{+}24]. temporary [WLG^{+}21]. temptation [LCS24]. tension [MM22a, MM24]. Tensor [DS24b, KSH22, NSD^{+}23, XYlW^{+}24, dSAOG21, BHL23, CHZ24, GZG^{+}21, HLX21, HM20b, HCY22, Hua24, IHF13, JL24a, LYZ22a, LHZ^{+}21b, LM20d, PLH^{+}24, QNX^{+}23, VZ23, WPE20, Woo20, XLJL22, YZJ^{+}20, ZW21c, ZH23b, Zha23a]. tensors [DK22a, DLQY20, Hu20, LL21, LM22, MY22]. Teodorescu [MP23]. Term [IMS23, BZS21, CJC20, DMD21, DFG22, DGO22, FJFS21, GNA22a, GNA22b, GX23, HSM21, LW23a, MT23, NGC21, QXG21, RC22, SSJ22, SZ24b, WW20, WWW24a, YWR^{+}24, ZZX24, ZJZD20]. terminal [ARMM23, LLY^{+}21]. terms [ANS^{+}20, AYY24, FZS21, LMX22, LK21b, Mit24, SMH^{+}20, TCGR24, Zha21b]. ternary [CLZY21]. tessellations [ZS21a]. Test [dDMMTH22, KÖ21, MBL20, ZS21a]. Testing [EFJO20, LWZ22c]. tests [ACVHC22, LOVHC23]. tetrahedra [TPSP24]. tetravalent [CZ21b]. text [DH24b]. textile [DD20a]. texture [HMR23, UBSU21, WZN22]. TGpV [WNZ21]. TGpV-shearlet [WNZ21]. th [AA21b]. their [AKP23, BM24, CP20, CWZ21, CK21, DVS21, GARD24, HCS20, KX23, LaGBT22, ÖÇÖ23, SHW21, SS23d, XLZL23, YWX20b]. theorem [DDE21, JW21, JZ22, KY22a, KÖY24, RY24, TlS^{+}22, TZ21, WJC^{+}20, ZJC21]. theorems [HLX21, SC20a, YY21a]. theoretic [AVJ24, JLL^{+}20, LLW24, SXS22]. Theoretical [GLX22, JRP22, MVCK24, NHK^{+}22, NHK^{+}23, RJM23, XLG^{+}21, Azi21, GMT20, LLCS23, ON21]. Theory [BCVV22, DJW20, KIB20, AR22, BSG23, BSCV24, BP23, CSW20, DdAS22, FC21, FRR^{+}22, Gan23, GXZZ20, GGBEZRI23, HGD^{+}21, JS20, KSPC22, KET22, LD21a, PRP24, SC21, TNTL21, TTL20, VMCG20, WJG^{+}20, WBY^{+}21, XJY24, ZTLW24]. theory-based [GXZZ20]. therapeutic [BJ21]. therapy [ALJ24, NNKH21, VMCG20]. thermal [AEE21, AR21a, ASS^{+}23a, BS22c, FWM^{+}22, FN22, GWH22, KYU22, KR20, KY21, Ma21a, PJMW22, PKG^{+}20, ST24, WWQ22]. Thermo [SS22c, WPJ22]. Thermo-mechanical [SS22c, WPJ22]. thermodynamic [GPRV21, LD20c]. thermodynamically [ABD23]. thermoelastic [Sha21, WPJ22]. thermoelectroelastic [KK24a]. thermohaline [TWZ^{+}20]. thermophysical [AR23]. thermosensitive [HGJ21, XGRM20]. Thermostated [ASAZ23]. thermoviscoelastic [BFQ20]. theta [RM23a]. thick [DG22b, Eft21]. thickness [WXWS20, Yu23]. thin [CH24a, DGV22, Eft21, LC21, RT22, SS23b]. thin-walled [RT22]. THINC [ZHQ^{+}23]. thinning [GAR22, KSPC22]. Third [GB20, SGJ^{+}21, CM22, HCTS23, HJY24, NHK^{+}22, NHK^{+}23, PMTT24, RS22a, SWZZ23, SXC21, SYX22, XLS21, YY24]. third-grade [NHK^{+}22, NHK^{+}23]. third-kind [CM22, SXC21, SYX22]. Third-order [GB20, HCTS23, HJY24, PMTT24]. third-party [YY24]. Thomas [SS24c]. Thomassen [LW22b]. thorough [MUOF22]. Thrackles [Xu21]. Three [ASS^{+}23a, BLO22, GZG^{+}21, RT20a, SS24a, TMA^{+}22, VBM20, AFGG20, BMJ20, CRS21, DWL^{+}24, DGP^{+}23, DJWX21, GPHC24, HPT22, HLZ22, HWL24, KY22b, KFR20, Kwo22b, LMT24, LCS23, LQNA23, SM21b, SSCG20, SB24, TYCK23, TBRB24, TT24, WLY21, Wri24, XbYW20, XLZL23, ZLD22, ZL23c, ZS22b, ZZY24, ZWJ^{+}24]. three-component [DGP^{+}23]. three-coupled [HLZ22]. Three-dimensional [ASS^{+}23a, BLO22, GZG^{+}21, TMA^{+}22, AFGG20, KY22b, LQNA23, SSCG20, TBRB24, WLY21, ZZY24, ZWJ^{+}24]. three-disk [DJWX21]. three-parameter [Wri24]. three-phase [TYCK23]. three-point [SB24]. three-species [DWL^{+}24]. three-step [ZS22b]. three-strategy [LCS23]. three-time [ZLD22]. Threshold [ID21, DH24b, LX23a, MZ24, SZXL24, WCY^{+}24, Xia21, XH22b, XH23, ZHG22]. threshold-based [DH24b, MZ24]. thresholding [WSG^{+}21]. thresholds [KK23b, PPZW20, WQZD23, YZT^{+}21]. throttle [RCG21]. throughflow [TM22b]. Thunsdorff [DMRR22]. Tight [TT20, TZLL24, WHJ^{+}24, ZHG22]. Tikhonov [CPLH20, MGJ21, Mit24, ZY21b]. Time [AKDR23, CCdS^{+}23, CNZ^{+}21, GW23b, JSZ20, JPZ^{+}21, KMS^{+}20, KMS^{+}23, LTL^{+}22, LWL^{+}23c, LWZ22d, MS21a, MZLS22, NI21, QZ22, WJ20, WGLS20, XLZ^{+}24b, YY20a, ZMS23a, AAJ23, AzR22, Alf24, AH21, AC20, AB21b, ASKA20, ASS^{+}23b, BZZ21, BFS20, BFLR23, BFRS24, BZS21, dJCGVM24, CZLH20, CDZW21, Cav22, CL22, CHG23, CWZL23, CWH24, CHZ^{+}20, Che21, CSW20, CYG20, CP20, CYL^{+}22, CLZ^{+}22, CWB22, CZK^{+}22, CL24a, CCL^{+}24, CTF24, CSS^{+}24, CFWL24, CZ24b, CKW22, CL24b, CWWZ24, CZXyS24, CA20, CC24, CCF24, CTBB24, CDL^{+}20, CWJF21, CLLA22, CZGX22, CXY^{+}23, De 22, DJWX21, DZDL22, DK22b, DD21b, DR22, DN21, DMD21, DS24a, DWZ20, DZZ^{+}21, DL21b, DLS^{+}23, DYX24, DDC22, DYW^{+}20, DL20b, DL21c, DL20c, Ema21, FJLG22, FLZ23, FJFS21, FJV21, FLL^{+}22, FZLF23, FFL23a, FZLJ24, FB20, FSLC22, FZ22, GZQ^{+}20, GYC^{+}22, GNSZ22, GWZZ23, GSS24, GdFPF24, GC24]. time [GXZ^{+}24, GE20, GS21, GN20b, GN22, GN23, Gök22, GB21, GLT^{+}21, GW21b, GSRL20, GZ20b, GWY20, GHP24, GLC23, GLD21, GHS^{+}21, Guo22, GX23, GW24, GZLY24, HWD20, HZW21, HHGZ23, HQS^{+}24, Har24, HL24a, HP23, HSM21, HKL22b, HLP24, HHW23, HZ23a, HHZZ20, HLL22, HLM23, HKCK24, HWL24, HWW20, HYT21, HH22, HXF23, HXWC24, HSY24, HMZZ20, JKJC21, JWC20, JDZH22, JCSH24, JWL^{+}24, JSA21, JSM20, JL24b, KMM^{+}24, Kan20, KGCC20, Kar22, KKB23, KPP22, KNN20, KSW24, KLPL20, LSV21, LPP21a, LPP21b, LKP22, LSWL20, LX20a, LCD^{+}20, LWZ21, LLZ21, LLY^{+}21, LHZ^{+}21a, LCXM21, LTW22, LGX22, LHC^{+}22, LLG22, LZWC22, LJY22, LL22b, LM23, LCZW23, LZZ24a, LNC24, LSM23, LLWY22, LLWY23, LZHZ24, LZHZ20, LZRF21, LZS^{+}21, LM20c, LXW^{+}20, LWG20, LLD21, LYXZ21, LZLZ21, LD21b, LNZ^{+}22, LW22g, LZZ^{+}23, LLWK23, LL23d, LXZ23, LZH21, LZZ24b, LLS21b, LLWW22, LJW^{+}22, LX24a, LWY^{+}24, MSZ22]. time [MXF^{+}24, MCT^{+}22, MTRA22, MR22, MHAT21, MMHD20, MFS22, MD22, MDT22, MB24, MP23, MZSW21, Nak23, NP20, DDM24, NK20, NK21b, NH22, OPP22, OPL22, OLW23, PPS20, PKL23, PCN^{+}24, PGZ^{+}24, PN24, QNX^{+}23, QDL^{+}21, QCARL20, QYLW24, RSVL21, RM23a, RVV24, RH20, RG20, RLK^{+}20, SM21b, SAB23, SZCC23, SYL21, SZG21, SMH^{+}20, SUN24, SZ24a, SPZS20, SLZ^{+}20b, SC20a, SZZL22, SKJM21, SMM^{+}24, SLZ20c, SWZS20, SCG^{+}20, SZWL21, SCHS22, SFL^{+}22, SLPH22, SDW^{+}22, SRZ22, SFC23, TlS^{+}22, TYCK23, TDZ^{+}24, TZ24, TMW20, TA21, TZ23a, TZ21, TW21b, TPTA24, TXZ24, VJA^{+}22, VPK20, VHB23, WW20, WS20, WXW^{+}20, WNP20, WZZX21, WGL^{+}21, WWW^{+}21, WA22, Wan22b, WLD^{+}22, WXF22, WM23, WC23, WZW23, WLHW23, WHPQ23, WSDW24, WM24a, WWW24a, WZZ24, WFKL24, WHWW21, WZLC21, WWT24, Wen20, WB20, WZ20b, WZXH22, WHL22, WCX^{+}22, XD21, XXL^{+}21, XH22a]. time [XBLW23, XbYW20, XLJL22, XZeF20, XZ21, XYM22, XZZ22b, XSP22, XCZ22, XTT22, XTS^{+}22, XHL23, XYH^{+}24, XXJtC24, XZQX20, XCSW21, XYLF21, XLLH22, XWL^{+}22, XZZC22, XZQ^{+}22, XWL23, XZQ23, XLZ^{+}24a, YGM^{+}22, YMRJ22, YZC^{+}22, YWY^{+}22, YWZT20, YLCZ20, YCL^{+}21, YLL22, YCM^{+}22, YLHZ22, YWFL23, YWZ23, YWS24, YH24a, YMW24, YGHW21, YW22b, YYF24, YL20a, YXZ^{+}24, YF21a, YYWL23, YL20c, YTLV22, YZDC22, YMYR21, YZ24, YL23b, YL24b, ZPSP22, ZRSK21, ZZH^{+}20, ZZW21a, Zen21, ZL22a, ZT22a, ZWH23, ZTLL24, ZYC^{+}24, ZZQ^{+}20, ZCPZ20, ZSM20, ZLLL21, ZLS21, ZS21b, ZWZ^{+}21, Zha21d, ZYZ22c, Zha22c, ZS22a, ZLK^{+}22, ZSFS22, ZDC22, ZFS22, ZLD22, ZWW23, ZF23a, ZC23a, Zha24a, ZJJ24, ZLZ24a, ZXF20, ZLW21a, ZJW^{+}22, ZGLZ22, ZYySX23, ZZX23, ZLZ^{+}23a, ZLJ20, ZZS^{+}22, ZZ20b, ZWHY20, ZDX21, ZC21b, ZTW22, ZCZW23, ZWZX21, ZL22e, ZJLT23, ZLLX22, ZLU21, dOS20]. time-boundary [QYLW24]. time-delay [FZLJ24, FSLC22, FZ22, GS21, HYT21, LGX22, LXW^{+}20, LLD21, LLWK23, NH22, TMW20, WJ20, WM23, XD21, XZeF20, ZZH^{+}20]. time-delayed [GZLY24, LCZW23, PCN^{+}24]. time-delays [MZSW21, ZZS^{+}22, ZLLX22]. time-dependent [AzR22, BFLR23, Cav22, CKW22, KMM^{+}24, LM23, MP23, SUN24, TYCK23, WNP20, XZZC22, YL23b]. time-domain [ZL22a]. time-efficient [Kar22]. Time-explicit [NI21]. Time-fractional [WGLS20, AH21, GC24, HHGZ23, HSM21, JCSH24, OPP22, RG20, SZ24a, TZ24, VPK20, WW20, WZ20b, XbYW20, XYM22, YWZT20, ZSFS22, ZFS22, ZF23a, ZLU21]. time-harmonic [GLC23, XTT22]. Time-inconsistent [LWZ22d]. time-inhomogeneous [DN21, GN22, GN23]. time-integral [FJV21]. time-integration [MCT^{+}22]. Time-invariant [AKDR23]. time-marching [DK22b, YGM^{+}22]. time-periodic [Zen21]. time-scale [YGHW21]. time-scales [TZ21]. time-space [BZS21, SFL^{+}22, YH24a]. time-stepping [HXF23]. Time-Varying [LWL^{+}23c, KMS^{+}20, KMS^{+}23, LTL^{+}22, XLZ^{+}24b, YY20a, LLWY22, CL22, CHG23, CWZL23, CP20, CYL^{+}22, CLZ^{+}22, CCL^{+}24, CZ24b, CWWZ24, CZGX22, DZDL22, DMD21, FLZ23, FZLF23, FFL23a, GNSZ22, GN20b, GZ20b, GX23, HL24a, HHW23, HWL24, JDZH22, KLPL20, LPP21a, LPP21b, LKP22, LX20a, LHC^{+}22, LLG22, LZWC22, LSM23, LZS^{+}21, LZZ^{+}23, LJW^{+}22, MXF^{+}24, MR22, PKL23, PGZ^{+}24, QNX^{+}23, RSVL21, SAB23, SPZS20, SMM^{+}24, SWZS20, SCG^{+}20, TW21b, TXZ24, WGL^{+}21, WC23, WHPQ23, WZLC21, WWT24, WHL22, WCX^{+}22, XXL^{+}21, XLJL22, XWL^{+}22, XLZ^{+}24a, YLCZ20, YWS24, YMW24, YZ24, YL24b, ZRSK21, ZZW21a, ZWH23, ZYC^{+}24, ZCPZ20, ZSM20, ZYZ22c, ZS22a, ZZ20b, ZL22e, dOS20]. time-varying-delay [XYLF21]. times [AAA20, BY23a, CLC20, QLKW20]. Timoshenko [AEE21, BP23, LD21a, WXHZ23]. TiO [AGC20]. tissue [CS21b, IMIH22]. Tobit [LWL22]. Toda [BFMGA20]. Toeplitz [AK21, AB20a, BSCV24, BBM21, FJJJ21, HV24, MH20, ZJJB22]. token [ZZLK23]. tolerability [Xu24a]. Tolerance [LaGBT22, Kur21, LM20a, YZM24, ZW21b, ZLC^{+}24b]. Tolerance-localized [LaGBT22]. tolerant [AKP23, CL22, GHP24, HZL21, HJW^{+}23, KMLW22, LZDX22, LXWH22, LD23a, LZHZ24, LLQW24, LJW^{+}22, MWZX21, MZFX20, NGS22, PMAK22, SCLD21, SWZS20, WXWS20, WXLZ21, WFYL21, WLL^{+}24, WJDL24, XH22a, XD23a, XCLW21, YWY^{+}22, YLK^{+}21, ZTW22, ZWZX21, ZWY^{+}22]. Too [FR24]. tool [ZMNR^{+}22]. topographic [WZY20]. topography [ABPP23, KPG22]. Topological [BBW21, MCB^{+}21, YLH22, BCR22, CMR20, GG24, LYB^{+}21, Ma21b, SG20a, SHW21]. topologies [FJLG22, HLW^{+}20, LCC^{+}24, SRKA23, WCZ^{+}22b, XD23a, ZDZS21].

Topology [Wan21a, ZZW^{+}21c, CZLH20, CFPM23, DL20c, LNC24, LSL^{+}21, QYLW24, WZLH22, XPP^{+}20, XWWC21, YWXL21, ZWFW22, ZLP21]. Topos [TCG20]. TOPSIS [FZG23]. toroidal [SYL23]. Torrey [BZS21]. torsion [Fre22]. Torsional [SC21, JA23]. torus [GY22b]. Total [ASV22, SO21b, SO21c, AAA20, AG20, BHL23, BY24, CYG20, GZG^{+}21, LWZ20, LZ22c, LW23b, LMZ^{+}21, MZ22, SGPZ22, WHHD21, XCGN24, XYlW^{+}24, YZW20, Zha21d, ZHW22, ZHFL23b]. Touchard [FR23]. Toughness [HDZ22]. tourism [CHP23]. tournaments [LLZC21, YLM24]. toxic [ZST22]. toxic-phytoplankton [ZST22]. toxin [ZST22]. toxins [SCD23]. Traceability [HAK20]. trachomatis [AMA20]. tracing [CH22b, DKKJ21, HHW23]. Tracking [GLP20, LWG20, MXY^{+}20, CZZP24, CZWX21, CZGX22, GZL^{+}23, GSS24, GHS^{+}21, HQS^{+}24, HKL22a, HL24a, HLLS22, JY21, JTC^{+}24, JJSY22, KKPS21, KBL^{+}24, KY22b, LXWH22, LDXZ23, LYX23, LTW24, LL23d, LH24b, LJWD21, LWY^{+}24, MPX^{+}20, MXF^{+}24, MC22, MN22, MJW21, MJ23, PX23, PGZ^{+}24, SX24, SSW22, VSM^{+}21, WLQX20, WSS22, WLD^{+}22, WXF22, WJDL24, WM24a, WPX^{+}20, XD21, XD23a, XWWC21, XXJtC24, XZZW21, XWL23, XWWL23, YY20a, YLL22, YP21, YP22, YZ24, ZWH23, ZLLL21, ZYSW22, ZYCY22, ZW22a, ZD23, ZDZ24, ZPY22, ZY23b, ZWJ^{+}24, ZCJ22, ZJLT23, ZDD^{+}22]. Trade [HH20, KAT22, VSM23, YXX^{+}22]. Trade-off [HH20, KAT22]. trade-offs [VSM23]. trading [HYS24]. traffic [BTL^{+}22, CCH24, CF23b, CF23c, JLL^{+}20, QLG^{+}22, WHD^{+}23, ZL23a]. trafficking [CTBB24]. Trail [LLZC21, YLM24]. Trail-connected [LLZC21]. train [YZJ^{+}20]. Training [VJKZ21, GWH24, JW23]. trajectories [CCZ24, HVZZ23, MJ23]. Trajectory [CZZP24, CTBB24, DYW^{+}20, LXWH22, MC22, QXSW22, XWWL23]. transaction [KMS^{+}20, LYZ22b]. Transcendental [SWFJ20]. transfer [AR23, AGC20, Azi21, DWT^{+}21, FB20, GBF21, HMR23, LLZ20c, LJZL22, LZ22b, LCXJ22, RB21, ST24, Son24, THW^{+}21, Voe23, WWW^{+}22, WF23a, YWL20, ZCT22]. transform [BSST20, BMMV23, BCVV22, CCC20a, DLL^{+}23, De 22, GZRM21, Gu20, KGPS22, Lia21, MSDR21, MS21b, MT24, RFFB23, RJM23, WKM24, Zha21c]. transformation [CZL20, Far22, HSY24, LM20b, LLQG21, LWL^{+}20, MT24, QCARL20, QXG21, ySyS22, YGHW21, YP22, ZJG21]. transformations [RGAM20]. transformed [HHGZ23, PHL23]. transforms [DVS21, MF23, WGM^{+}21, XGF20, ZK22]. Transient [ABNS24a, ABNS24b, ABNS24c, Jan22, CLL22, DGP^{+}23, LLS21a, LLLC23, UZ20, YFHC21, ZC23b]. transition [ÅCL24, HWY^{+}22, HC24, JWK22, JM23b, LLZ20b, LZYL21, NK20, NK21a, QYG^{+}20, SGI20, SKSC21, TW21a, WC23, XXW^{+}20, ZRSK21, ZM21c, ZLP21]. Transitions [PJMW22, TWZ^{+}20]. transitive [CZ21b, DLM20, HCJC21, HLF22, LYJM21, PY21, TYL^{+}20, TCZ22]. Transmission [SRK^{+}21, XZSS21, AYS20, AYS22, BD21, CJ22, CS20, CZR21, DSAY24, DS20, GGW21a, GY22a, HKL22b, ID21, IG21, JM23a, KWM20a, KWM20b, KIJK20, LP22, MHAT21, PS21a, WMWS21, WOB21, WYLZ24, YHB^{+}20, ZYZ22a, ZL22c, ZLC24a]. transmission-blocking [WOB21]. transmission-irregular [DSAY24]. Transport [DG20a, RWGS22, AMBG^{+}20, BJK^{+}22, BGRM21, CKV21, DdAS22, DNR21, FTML22, GDFNNR22, KCT^{+}21, YLL20, MNSR22]. transportation [WYW^{+}20, WGJK21]. transposition [CLWZ21]. transversal [MBK22]. trapezoidal [GE20, JL22]. trapping [SLS24]. Traveling [KL24, MD23, TC22, ZXB20, YR20b, ZZ22a]. travelling [RMO22]. traveltimes [QPHC21]. traverse [JWK22]. treatment [AMA20, AACGTB23, AACGTB24, HCTS23, LQLF22, LZ22a, RVV24, YI22, ZZL22]. Tree [KZ22, LG23, AA20, Gan21, LG24, LL20a, NPE22, SLL^{+}22, WW23, YZW20, ZBFW22]. tree-like [YZW20]. trees [AYS20, BOT24, BNR20, BCR22, CLL21, CWZ21, CDH23, CCHZ23, CLB21, CMR20, CRS21, DG22a, DL21a, DD21c, DGO22, DS22c, DS24c, FLL^{+}22, GGS21, GG24, GWW23, JL21, JGW21, JWMW23, LLZ23, LZ24a, LXLP20, LWZ22b, LY23a, LLSY23, LLCL23, LCWZ24, Lyu23, MY24, MD24, NTNT24, SHW21, SD22b, XY22, YW20, YLPC22, YH24b, ZHW22]. Trefftz [XFW^{+}21]. trends [SLCS20]. Tri [QLKW20, CLZY21]. tri-block [CLZY21]. Tri-criteria [QLKW20]. trial [BBH21]. triangle [CSK23, TLD22, YY21b]. triangle-free [CSK23, TLD22]. triangles [JWWL20]. triangular [BDPV22, CS24c, IKÖ24, LGQ20, LZ22g, SSP24, WYZ24]. triangulated [GLMT21, RiTZ24]. triangulations [LMS24]. Trichiurus [JKJC21]. tridiagonal [BF20, BF23, Cas23, CdFP24, MH20]. Triggered [CNZ^{+}21, CLWX23, MWS23, QQY^{+}23, ZCL^{+}23, ZLLX22, AS23a, BZZ21, CZLH20, CDZW21, CZ22b, CC22, CTF24, CZWX21, CXGL22, DSZ^{+}23, DWZZ22, FSS20, FLLW23, GC24, GSL^{+}22, GZLY24, HRLH21, HLL22, HL23a, HZX22, HJW^{+}23, HSL^{+}24, HMZZ20, JY21, JLP23, JJSY22, KMLW22, KY22b, KJL20, LDX20, LWD21, LZ22b, LGX22, LHC^{+}22, LZC22b, LZWC22, LLY22a, LX23a, LHZ^{+}23, LWL^{+}23a, LCZW23, LR24, LM24, LGC^{+}24a, LSM23, LZW^{+}20b, LTL^{+}21, LYXZ21, LNZ^{+}22, LYL22, LLQW24, LJW^{+}22, LR23, LX24a, LY20b, MYW21, MZY^{+}22, PLRD21, RJHC24, SKLS22, SYL21, SWS^{+}22a, SWX22, SZSZ22, SYTC24, SLPH22, SSW22, TLHH21, TWB22, WFYL21, WLZF21, WWW^{+}21, WCW22, WYZX22, WPZY23, WYLQ24, WSDW24, WPX^{+}20, WCZ^{+}22b, WWW24b, XXL^{+}21, XBLW23, XCW^{+}22, XZY^{+}23, XZQX20, XWWL23, XLZ^{+}24b, XRZH22, YYG20, YWY^{+}22, YXMW22, YCM^{+}22, YWWY23, YM23a, Yan24, YL20a, YCZ^{+}23, Yoo21, YP22, YZ21, YZY22b, YZ24, ZeFWW20, ZZCW21, ZDZS21, ZZLL21, ZYSW22, ZYXL23, ZCW^{+}24a, ZLZ^{+}21, ZCLZ22]. triggered [ZPLL20, ZCZW23, ZCJ22, ZWY^{+}22, ZB22]. triggering [DS24a, LSC^{+}23, LY20b, MTFD24, QZH21, ZXW20]. Trigonometric [BCHRS23, AHB20, BE20, Luo20, Sun22]. trimatrix [Sek23]. trinion [WGM^{+}21]. triopoly [PLW^{+}20]. tripartite [JWMW23, SLL24]. Triple [AÁC^{+}21, CIHP^{+}21, ZCPZ20]. triple-integral [ZCPZ20]. triply [NTLNX24]. trips [CGMFSHR21]. true [Fan21, ZFS22]. truly [YLC20b]. Truncated [HL24a, BA20, KKB23, MY22, MT23, SRK^{+}20, ZM21b]. truncation [Ily23, JP22, LLZ22c, SRZK22]. Trust [CG20b, GLLZ24, Wan24a]. trustworthiness [ZRZL24]. Truth [TCG20]. try [HCL^{+}23, LX24a]. try-once-discard [HCL^{+}23, LX24a]. tryptophan [AMA20]. TSP [SKWW21]. tSSNALM [XYKL20]. tube [Hoj20]. tuberculosis [DKK20, LLY^{+}22b, Zha22c]. tubes [LLWY23]. Tucker [PLH^{+}24]. Tukey [CS21a]. tumbling [GK20]. tumor [ALL23, LXCZ21, MFA21, VBP22]. tumor-immune [LXCZ21]. tumors [ALJ24]. tuning [HCX^{+}24]. tunnel [YXZ^{+}24]. tuples [DDNS20]. Turán [CY24, LLY20, LY21a, LY22a, ZY24a]. turbofan [GZL^{+}23, LCLS20, LCL22a]. Turbulence [Fab20]. Turbulent [LTL^{+}20, CQ24, Miz21]. Turing [CWC20, LGS22, MUMS22, SMTT23, ZS20b]. Turing-Hopf [CWC20]. turning [YR20a]. Tutte [MJZ21]. TV [dSLV20]. TV-based [dSLV20]. TVCS [MZL24]. TVD [MDT22]. TVD-MOOD [MDT22]. twice [LZC22a]. Twin [PMAK22, YR20a]. twisted [CLL23, THZ23]. twisting [CWZL23, ZDC22]. Two [Che22, DS21a, DHS23, GLW24, HQZC24, HJY24, KFÖ23, LG23, MMK^{+}23, NXL21, WHC20b, WSG^{+}21, YZLS20, ZT22a, Zha21d, Zha21a, ZGHM24, ZGZ^{+}24, ZZW20, ASR22, ABNS24a, ABNS24b, ABNS24c, Alf24, AM20b, AKR21, BCLNV20, BS22c, BZS21, BMS21, CM20b, CZ22c, CJ21, CSCZ24, CRTV22, DP24, DE21, DMSS20, ELOY22, EBJ20, FLS20, GWZ^{+}21, GK20, GBF21, GKK^{+}24, DDCM23, GZLZ24, HT24, HLZ20, HZZ^{+}21, HXWC24, JNY^{+}21, JLFL20, JL22, JCSH24, JT20, JT22, JT24, KWH^{+}23, LJAM24, LMX20, LMX22, LZCL23, LZ24c, LXH24, LHZ21c, Liu22b, LWZ22c, LtTBPA20, LZ22f, LZ22g, MS21a, MMDD21, OBP21, PGKM24, PH20, PSXX21, PZS24, PM20, QHC21, QLM21, QWD22, QL21, RS22a, RMSP22, SM21b, ST24, SPX20, SPX24, SS22b, SSJ22, SNMM21, SL20c, SL20b, SW20, SK20, Sta20, Sun22, SLL^{+}22, SFL^{+}22, SJ23b, TNT23, Tak21, TLC22, TZ24, TC20a, TWWL24]. two [TZ23b, Ued23, UZ20, VBMB21, WZ20a, WFHG20, WLY21, WHHD21, WL23, WHKT23, WMF23, WZW24, WZW^{+}21, WSZ22, WPJ22, XYZY20, XYKL20, XAZ^{+}21, YLG21, YCEW21, YHW^{+}23, YXXZ23, YL24a, ZMS23a, Zen21, ZJG21, ZYPL20, ZZY^{+}21, ZL21a, ZL23a, ZZVL23, ZHQ^{+}24, ZJJ24, ZRL^{+}24, ZZX20b, ZZW21b, ZS20a, ZS22c, Zho20, ZC21c, ZCZW23, ZLS23]. two-by-two [LMX20, Zen21]. two-channel [ZCZW23]. two-component [LMX22, UZ20]. two-derivative [ELOY22, ZMS23a]. Two-dimensional [ZGHM24, BCLNV20, BZS21, CJ21, DMSS20, GBF21, JCSH24, PH20, PSXX21, QHC21, QWD22, QL21, Sun22, SFL^{+}22, WFHG20, WSZ22, YLG21, YHW^{+}23, YXXZ23, ZJJ24, ZRL^{+}24, ZZX20b]. Two-disjoint-cycle-cover [HQZC24, NXL21, WHC20b]. Two-facility-location [GLW24]. Two-grid [ZT22a, JLFL20, PZS24, SL20c, SW20, TLC22, TZ24]. two-group [ZYPL20]. two-layer [DDCM23, LXH24, Tak21, WZW24, WZW^{+}21, ZZY^{+}21]. two-layered [PM20]. two-layers [BS22c]. two-level [ZS20a]. two-mode [ZJG21]. two-parameter [ASR22, LZ22f, LZ22g, TZ23b]. Two-parameters [DHS23]. two-person [SJ23b]. two-phase [JNY^{+}21, JT24, MMDD21, RMSP22, SSJ22, TC20a, XAZ^{+}21]. two-point [EBJ20, RS22a]. two-prey [AKR21]. two-round [CM20b]. two-scale [OBP21]. two-side [YCEW21]. two-sided [HLZ20, ZLS23]. two-species [Alf24]. two-sphere [Zho20]. Two-stage [WSG^{+}21, GZLZ24, SNMM21, XYKL20]. Two-step [DS21a, SK20, TWWL24, WHHD21, ZZVL23, ZS22c]. two-strain [Liu22b]. two-strategy [ZHQ^{+}24]. two-time-scale [HXWC24]. two-tree [SLL^{+}22]. two-variable [HZZ^{+}21]. two-way [CRTV22]. type [AM20a, AH21, BG20b, BM24, CRT22, CU23, CLZY21, CF23a, DMRR22, DO21, DYT21, DDH22, DML21, EBGN23, FKK24, FCeFZ21, GGW21b, GG22, GA22, GP22, GK21, GH20b, HV24, HLX21, HWLL21, HGKD22, HCL^{+}23, HLX20, HH22, Hua24, IHF13, Iza21, JD20, KX23, KÖ21, KKS22b, KJL20, LHW21, LWCD24, LLD21, LM22, LZ22e, LSVA20, MKKG22, MN23, NSD^{+}23, OPS20, ÖY21, PBM20, RKJ20, RGS23, SSC22, SSI23, SKLS22, SNMM21, SRZ22, TBRB24, TC20a, URLC21, Vyn23, WHLL20, WGL^{+}21, WW22b, Woo20, XCZ22, YHD24, YY22, ZL20a, ZL21a, ZL22d, ZL23a, ZS22b, dSTV21, AKR21, BSM23]. type-2 [HGKD22, HCL^{+}23, SKLS22, SRZ22]. Type-I [BSM23]. Type-II [BSM23]. Types [HD20, AS20, Bia21, GLW24, INT20, LWZ22c, MYW21, PM20, QLM21, SPX24, TNT23]. typhoid [IG21]. typical [RMCG20].

UAV [WC22b, XDS^{+}20]. UCM [Eva20]. Uganda [NS20]. Uhlenbeck [AB20b, WYLZ24, ZY21a, ZEA22, dlCJ20]. ultimately [MZLS22]. ultimatum [DZW21]. unadjusted [GAA^{+}24]. Unbalanced [TLCM24, BBT21]. unbalancedness [MS21c]. unbounded [GW21b, NNN24]. Uncertain [CL23, KWHS23, LY21c, SHWW21, TY21, XWDK23, YXYZ23, YHSZ20, ASKA20, CZWS20, CZS21, CHHC24, CSS^{+}24, CZ24b, CZWX21, DZZ23, DCR^{+}20, DL20c, FLLW23, GJ21, GGY22a, GGY22b, GSWL23, GSRL20, GSL^{+}22, HLL22, HLKC22, HWW20, HSL^{+}24, JY21, JWC20, JWK22, JSZ20, JPZ^{+}21, KY22b, LLZ20b, LDX20, LHZ^{+}21a, LHX^{+}22, LD23a, LL23b, LR24, LSCS21, LTL^{+}21, Liu21, LY21b, LY22c, LCL22a, LXZ23, MXF^{+}24, MZSW21, NK20, NK21b, OPL22, PP21, PGZ^{+}24, SNRA20, She20, SMH^{+}20, SX24, SLPH22, SKSC21, TDZ^{+}24, TY22, TPTA24, WNP20, WLZF21, WGL^{+}21, WZGL21, WJL^{+}24, WPX^{+}20, WSL^{+}21, WX21a, WZXH22, XD21, XD23a, XZZ22b, XDS^{+}20, XZZW21, XZ23c, XLZ^{+}24b, YLCZ20, Yoo21, YP21, YYWL23, YZDC22, ZS22a, ZHG22, ZD23, ZWW23, ZSP24, ZY23b, ZLZ^{+}23a, ZXZ20, ZLJ^{+}21, ZPLL20, ZZW^{+}21c]. uncertainties [FJLG22, HTF20, HSZ20, Jor21b, LWZ22a, LZZL21, LWL21, QL20, RFFZ21, SXS22, ZWZX21, ZL21c]. Uncertainty [AR22, DLMP23, HKL22a, HP23, IFHA21, KMLW22, KG21, LePP24, MZW^{+}21, MF23, XZQ^{+}22, ZGG^{+}22]. Unconditional [CZ22c, LLS21a, LW22d, LXL22, PZS24, SZ23, SZ24a, LHYL24, Mef21, WS21]. Unconditionally [FG24, CLZY21, YWXL21, ZSS24, ZGZ^{+}24]. unconstrained [ZZW20]. under-actuated [LNZ^{+}22]. underactuated [KY22b]. Understanding [WYLZ24]. underwater [KY22b, LYX23]. undetected [MGFF21]. Undirected [HS23a, ZTYZ23]. unequal [HT24, ZZY^{+}21]. Uneven [WDW^{+}22]. Unfairness [CZW23]. unicyclic [BBT21, CZ24a, Das24, FLL^{+}20, LYhH^{+}24, LD23b, LCWZ24, WYG23, YW20, YD22a, ZCM^{+}24]. unidirectional [LZ20c, LGD^{+}21]. Unified [KPPK23, CS24b, DLL^{+}23, GZLY24, KY21, KFÖ23, LD21b, SMW^{+}23, THW^{+}21]. Uniform [AMF20, But21, CKW22, JZZ^{+}22, OT20, SKH20, ZL22b, AS20, DWW24, GI23, KSR21, LWCD24, LU24, TMA^{+}22, WX20a, WLZD24, WS21, ZLL20, ZJJ24, ZZY24]. Uniformly [SW20, WXHZ23, MZLS22, Mit24, RC22, SR23]. unilateral [GWH22, XJY24]. unimodality [ZZ23]. union [HZ20a, JJC24]. unique [MY24]. Uniquely [BDK23, GPW24, KFR20]. Uniqueness [Kwo22b, CPS24, GYC24, GZRM21, MSE24, PS21b, WJC^{+}20, YSHV^{+}23]. unit [SLJ23, TZLL24]. United [LLY^{+}22b]. uniting [GWSC23]. Unity [CDP23, CD20b, CDLR24, NA21]. Unity-based [CDP23]. univariate [FGM24, XLD23]. universal [AR22, GKT^{+}23, JG21, LBJ22, RPB23, YP22]. Unknown [MXY^{+}20, CL22, CLWX23, CZZP24, GXZ^{+}24, GS24, GSRL20, HLW^{+}21, HLL22, JWK22, JWL^{+}24, JJSY22, LWZ21, LJL21, LZZL21, LH24b, MPX^{+}20, MWZX21, MZFX20, MPC^{+}24, PHS^{+}20, QL20, ST24, WLQX20, Wan22b, WS23b, WPX^{+}20, WWW24b, Yoo21, YZDC22, YZ24, ZWH23, ZTLL24, ZW22a, ZSP24, ZPY22, ZY23b, ZTW22]. unknowns [WSY21]. unlabeled [LFLL21]. unmanned [HZL21, LTL^{+}22, LCL22b, XWWC21, XXJtC24, XWWL23]. unmeasurable [LR23]. unmeasured [CLH24b, WXW^{+}21, Yi23]. unmixing [WHHD21]. unmodeled [WXW^{+}21, XLZ23, XWL23]. Unscented [LWL22]. unstable [FLZ23, FZ22, GLY24, GT22b, YZ23, ZS21b, ZC21c]. unsteady [BS22a, BPRR22, LW22g, RB21, SWS20, SSF24, YOF21, ZF21]. unstructured [BBD^{+}22, BRMVCD21, BDRM23, TKF23]. Unsupervised [JC22a]. unsymmetric [CD20a]. until [TZ23a]. unusual [HD22b]. Unveiling [WZW24]. update [Che21, ySyS22]. updating [LJR20, SRK23, WLZW21, WWZW23, YWJ^{+}20b, YSW^{+}21, YSL21, ZLF22]. uplink [XSP22]. Upper [FCX21, SRZK22, Xu21, CHY22b, DFX24]. ups [LLZ23]. uptake [LNO23]. urban [BS23, QDL^{+}21]. Urysohn [BM24, NRP22]. USA [HXL^{+}20]. usability [MSMK21]. usage [CK20]. use [BAL^{+}24, ES24, JTW^{+}23, ZMS23b]. Using [AT21b, LJL21, MD23, dSTV21, AMA20, AMM21, AMBG^{+}20, AR22, AVJ24, AS21c, ADS23, ACGG20, AHEGM23, AR20, BS22a, BTL^{+}22, BSRV23, BEe21, BJK^{+}22, BLNT22, BGRM21, BO21, BMY24, BS23, CDBR^{+}22, CLL22, CZMX21, CCL^{+}24, CH24c, CLZ24, CUA^{+}22, CGLR23, DESM^{+}20, DYA21, DHM^{+}20, EA24, EFJO20, FGMS21, FLR23, GP23, GTWL21, GK20, GY22b, GS22, GK21, HYÇK22, HDP^{+}21, HKKL22, HR23, HCBY21, JC22a, JNY^{+}21, JMM20, JS20, JY21, JKJC21, KM22, KTEM22, KG21, KJL20, LFLL21, LJAM24, LPP21a, LPP21b, LYKK21, LZY20, LD20a, LCZ^{+}20, LWZ22a, LLP23, LFLH23, LZ20a, LWQW22, MUOF22, MQ22, MC20, MSN21, MSB22, MVO21, MMS21, MV22, MZYW23, NGGD23, PJMW22, QCARL20, QYD22, RS21, RY24, RFFB23, RLK^{+}20, SM21b, SRK23, SWX^{+}21, SS22c, SRGS20, TKY23, TWB23, TJIM20, TKF23, UH24]. using [UUS21, VBM20, WYLW20, WLQX20, WWZ20b, WLD^{+}22, WLPZ22, WNZ21, WCX^{+}22, YJ20, YSW^{+}22, Yoo21, YLC20b, YSL21, ZIML20, ZCPZ20, ZZL^{+}20a, ZSM20, ZXF20, vdMvRS21]. USJSs [LZS^{+}21]. utility [Bri20, LYZ22b, ZEA22]. Utilization [YOF21, ZZCL20]. Uzawa [BM21, SLZ20a].

vacation [AVJ24]. vaccinated [RHNT21]. vaccinating [ZLZ^{+}23b]. Vaccination [WJG^{+}20, FW24, Kab24, LDZT24, MCSD21, WZA^{+}24, WYLZ24, ZYPL20]. vaccination-resource [ZYPL20]. vaccine [GLGLH24, WZA^{+}24]. valency [LZC22a, YF21b]. valent [LYJM21]. validation [Mar24, MGK23, PGR23]. Valuation [BO21, vdZGO21, ZZ24]. value [ASUH21, AJIR21, ACGG20, ASV22, BPT23, BBHM21, CM20a, CRT22, CM22, CSS^{+}24, EM21, EBJ20, Far22, GTSV23, HBCA21, IZ20, JX24, JCH^{+}20, Kar22, RS22a, RPB23, RG22, SO21b, SO21c, SS24c, SJ20, SB24, TZ23a, ZJWV24, ZJX21, ZZW21b, ZJX22a]. value-at-risk [ASUH21]. value-based [CSS^{+}24]. Valued [HHY20, AAJ23, ADV23, CWB22, GJS^{+}23, GX23, HZ23a, HHZZ20, HSZ20, HB20, LZH^{+}20, LHZ^{+}21a, LCXM21, LCY^{+}21, LZH21, ON21, RSVL21, RJHC24, SZCC23, SWX^{+}21, SWX22, URLC21, UBSU21, WWWH21, Wan24b, WZLC21, XYH^{+}24, YYF24, YLCL23, ZLZ^{+}23a, ZQ24]. Values [TCG20, CMR20, DL21a, DQC^{+}21, LBX^{+}23, ZYLH21]. vanishing [BPG24, Xy20, YTZ22]. VaR [BLNT22]. Variable [RYG21, SLS24, YLHZ22, ABPP23, AHEGM23, BPR21, BD20b, CH22a, CSL24, CLM20, DC21, DS22a, DKLV23, FRS22, GZX20, GDB21, HZZ^{+}21, Kan20, Kar22, KR20, PSXX21, PWZ22, QNX^{+}23, RRS21, RY24, TYCK23, TZ24, TM22b, Wan23a, WGLL20, YCEW21, Yu23, ZT22a, ZL22d, ZFS22, ZZLV23, ZGZ^{+}24, ZLW21a, ZF23b, ZC22]. variable-Coefficient [YCEW21, RRS21]. Variable-order [SLS24, PWZ22, WGLL20, ZL22d, ZFS22]. variable-section [DC21]. Variable-time [YLHZ22]. variables [CH24b, LR23, Pop20, PMG20, SV21a, VHB23]. variables-augmented-based [CH24b]. variance [Bri20, Cha23, CDM21]. variances [ZYLH21]. Variant [YY20b, HWXC20]. variates [OG20]. variation [BHL23, GZG^{+}21, LWZ20, LZ22c, LMZ^{+}21, RWGS22, SGPZ22, WHHD21, XCGN24, XYlW^{+}24]. Variational [KWW22, BBHM21, Fen22, GJJC^{+}23, Lee23, LMX22, MNZ20, UUS21, WHWW21, ZS20a, ZS22b]. variational-hemivariational [MNZ20]. Variations [CWZ21, BBB24]. various [AS20, Che22, GLL^{+}22, LCC^{+}24, LZY^{+}23b, TAU^{+}22]. Varying [LLWY22, LWL^{+}23c, VSM23, CL22, CHG23, CWZL23, CP20, CYL^{+}22, CLZ^{+}22, CCL^{+}24, CZ24b, CKW22, CWWZ24, CLL24, CZGX22, DZDL22, DMD21, FLZ23, FZLF23, FFL23a, Fre22, GNSZ22, GC24, GN20b, GZ20b, GX23, EVB^{+}21, HL24a, HHW23, HWL24, JS22, JDZH22, KMS^{+}20, KMS^{+}23, KLPL20, LPP21a, LPP21b, LKP22, LX20a, LHC^{+}22, LLG22, LZWC22, LTL^{+}22, LSM23, LZS^{+}21, LZZ^{+}23, LJW^{+}22, MXF^{+}24, MR22, PKL23, PCN^{+}24, PGZ^{+}24, PSE24, QNX^{+}23, RSVL21, SAB23, SPZS20, SMM^{+}24, SV23b, SWZS20, SCG^{+}20, TA21, TW21b, TXZ24, WGL^{+}21, WHL21, WC23, WHPQ23, WWW24a, WZLC21, WWT24, WHL22, WCX^{+}22, XXL^{+}21, XLJL22, XYLF21, XWL^{+}22, XLZ^{+}24a, XLZ^{+}24b, YLCZ20, YY20a, YWS24, YMW24, YZ24, YL24b, ZRSK21, ZZW21a, ZWH23, ZYC^{+}24, ZCPZ20, ZSM20, ZLS21, ZYZ22c, ZS22a, ZZ20b, ZL22e, dOS20]. vasculature [CS21b]. Vasyunin [DF20]. VDB [LYB^{+}21]. vector [AR24, ALM20, CZ21a, FJ24, GB20, GLP22, GGBEZRI23, LLZ^{+}20a, RHNT21, UBSU21, WPE20, YTZ20, ZGHM24, ZT22b]. vector-borne [RHNT21]. vector-kinetic [AR24]. vector-valued [UBSU21]. vectors [WWZ20a]. vegetation [CWF^{+}21, LLS^{+}22, LWX^{+}22, TLZ20, XLL^{+}21b, ZWS^{+}22]. vehicle [LYX23, WWW^{+}21, XWWL23, ZSZL20, ZZX^{+}22]. vehicles [DSZ^{+}23, HZL21, JJL23, KY22b, LLQW24, LWY^{+}24, XXJtC24]. VEIQS [GZ20a]. velocities [ZL23a]. velocity [DGP^{+}23, HCBY21, LXWH22, LJ23, SSW22, TM22a, ZL20d]. verification [LZ22d]. Verified [SBB^{+}22, Miy23]. Verlet [Zha20]. Vernotte [vdMvRS21]. version [BS20, CG20b]. versions [CP20]. versus [CM22, WL21b]. Vertex [BCR22, LZ24b, WJY21, ASS24, CRDD21, CY21b, CMR20, DESM^{+}20, GG24, MSS24, MMOS24, NMS23, NXL21, SZZ24, WHF20, WSY21, WF23b]. vertex-cover [WF23b]. Vertex-degree-based [BCR22, CMR20, GG24]. vertex-weighted [CRDD21, CY21b]. vertical [IMIH22, MJK22, MMS21, SKS21, ZZLV23]. vertices [CRS21, DD21c, DHZ23, DS22c, GKT^{+}23, GPHC24, HZCY23, HK21, JY23, PMAK22, QMS22, SM24]. Very [ECS21, FY23, ST24, SLCS20]. vessel [IMIH22]. vessels [LXWH22, LTL^{+}22]. via [ABD22, AOSO20, AM20a, AKH^{+}22, ASS^{+}23a, AAHC21, BZZ21, BBHM21, BA20, BG20a, BCVV22, BBM21, CZ21a, CHG23, CH24b, CM22, CZW24, CLH24b, CXGL22, DS24a, DNR21, FKS23, FAEZ24, Fen22, FdLPC^{+}22, GZ20a, GGY22b, GWZZ23, GARD24, GN20a, GHS^{+}21, HWD20, HZL21, HSM21, HVYMS21, HKU21, HL23a, HJW^{+}23, Hua24, HSL^{+}24, HB24, HMR23, JW23, JDLZ21, JGC^{+}21, JW21, JT24, JZ22, KMS^{+}20, KMS^{+}23, Kum24, LSPK20, LDZ21, LPK22, LLZ21, LZDX22, LJY22, LWL^{+}23a, LWL23b, LTC^{+}23, LCLH24, LM24, LGC^{+}24a, LZHZ24, LCY^{+}21, LWZW20, LWG20, LHFW21, LTL^{+}21, LYXZ21, LNZ^{+}22, LM22, LCL22b, LXK23, LMÖ23, MTFD24, MY22, MPC^{+}24, MG21, PHS^{+}20, PN24, RDL23, SWS^{+}22a, SRK^{+}20, SKR22, SSR23, SCG^{+}20, SM23, SHYS23, SL24, TNT23, TMW20, TWR20, TZLL24, UHZ^{+}20, VMCG20, WHZ^{+}20, WCG^{+}20, WJ20, WHL21, WHHD21, WZC^{+}22, WM23, WLHW23, WHPQ23, WGJK21, WGMJ23]. via [XBLW23, XLZ23, XLJL22, XWM23, XLZ^{+}24b, YSLK23, YGHW21, YYWL23, YZY22b, YZ23, YY22, ZZS^{+}20, ZXC20a, ZJZD20, ZHG22, ZDC22, ZYXL23, ZM21c, ZLZ^{+}21, ZGCZ22, ZeFWD23, ZZS^{+}22, ZWHY20, ZWY^{+}22, ZPW22]. Vibration [AEE21, KET22, Bab22, Eft21, EM20, FB20, FZ22, JA23, JGXL23, KMZ22, KG21, LC22a, LM20c, MYW21, MYC^{+}23, SLHX21, Wan21a]. Vicsek [WDZL24, YYL^{+}23]. video [MJW21]. views [ZBFW22]. Vincular [LY20a, AE21]. violence [FR24]. violent [Cro22]. viral [Alf23, ALJ24]. virotherapy [EHA20]. virtual [BKS21, LF21, XZZ22a, ZF21, ZF23a]. virus [CXLM21, KIJK20, LCC^{+}24, LJHL21, NNKH21, XZSS21]. viscid [Kri22]. viscoelastic [ATDJ22, BP23, DC21, EFJO20, GWH22, KKPG20, LSL^{+}20, LC22a, MMS21, ZZY24]. viscoelasticity [LQNA23, YWZ23]. viscoplastic [GAR22]. viscosity [JNY^{+}21, MUOF22, Mef21, SC21]. viscosity-extended [SC21]. viscous [BT22, Fab20, KA22, KA23, PR20]. visibility [BY24, CSK23, Di 22, HXL^{+}20]. visualization [LPS21, MUMS22]. Vlasov [CdR22a, ZLU21, dRST23]. voice [JYMP^{+}20]. volatility [AARS23, HOM22, KK23c, LWZ22a, LSCS21, SO21b, SO21c, Wan24b]. vole [DNR21]. Volterra [AAS23, AFA24, AHB20, AP21, BPG24, BG20a, BM24, BRB24, BFMGA20, CN23, CM22, CUA^{+}22, DYT21, DDH22, EHVMM24, FMP20, GH20b, HWC20, Mar22, MSN21, MSE24, RS21, RSA24, SXC21, SYX22, TC22, WBV22, WH21, Xy20, XKW^{+}23, YTZ22, ZHL24, ZH20b]. volterral [PH20]. Volume [KH23, AB23, ABNS24a, ABNS24b, ABNS24c, BBD^{+}22, BRMVCD21, BDRM23, CCCdLMR23, CG23, CH24c, LC23, LTB23, dSLP20, LYN22, MMDD21, MVO21, PDK23, PGF20, Que22, TZJ20, TKF23, WSY21, Yan23, ZZW21b, REC^{+}23]. volume/finite [BRMVCD21, BDRM23]. Voluntary [CSS^{+}22, SHT^{+}21, ZLZ^{+}23b]. Voronoi [RLK^{+}20, ZS21a]. VOROPACK [RLK^{+}20]. VOROPACK-D [RLK^{+}20]. Vortices [Fab20]. Vries [AF21, YZLS20]. vs [SS22b]. VTOL [DDGBGS^{+}24]. vulnerable [DTWJ21].

Waals [CSZ24]. wait [AAA20]. walk [CRTV22, FSS21, LW22e, NKA20, QZ21, Zha24b]. walks [BS23, FLL^{+}22]. walled [AAHC21, RT22]. walls [SDD22]. Wang [RO21, CG20b]. Wannier [VGdFCBA20]. warehouse [BDKZ21]. warming [CWF^{+}21]. watching [GDM^{+}20]. water [ABPP23, AGC20, BGRM21, CCCdLMR23, CG23, DYA21, DDCM23, KSH22, KYU22, KPG22, LLQG21, LC23, MMK^{+}23, SGI20, TCGR24, WP21b, YTC23b, ZZCL20, ZLN20]. water/ethylene [AGC20]. Watts [CWZ22]. Wave [BP23, KK24a, MD23, WLY20, ADS21, BS20, BFS20, Boc22, CM20a, CLL22, CK20, CEPZ22, DYA21, GA20, HD22b, Jor21c, KMN22, Kud20b, KL24, LM20b, LW22c, LJ23, LLLC23, LWCD24, qLgZ21, LV20, MMK^{+}23, MK21a, MP23, NP20, OKT24, ÖY21, RMO22, RHR^{+}21, RC21, RO21, SSC22, SSJ22, SW20, SFC23, TZ24, TC22, TCZ^{+}20, WDG21, WS21, XbYW20, YR20b, YLG21, YOF21, Zha21a, ZXB20, ZC22]. wave-dynamics [YOF21]. wavefields [GE21]. waveguide [DG22b]. wavelength [YTZ20]. Wavelet [ZK22, HZZ^{+}21, HOM22, KG22, LK21a, SV21a, UUS21, Yu21, Yu23]. wavelet-based [HOM22]. Wavelets [BFS20, AAS23]. waves [BL22, BBM21, CK21, KKPG20, MKKG22, MM22b, Ndo23, Oru21, Oru22, PS22, RHN^{+}24, Sha21, SSCG20, SC21, TTCY23, WZY20, WP21b, ZJG21, ZLA^{+}21, ZZ22a]. wavy [AAHC21]. wavy-walled [AAHC21]. way [CRTV22]. Weak [PWWZ23, QS20, WGLS20, XW23, ABD22, CWZZ22, DG21, KWZX23, KAS21, Kum24, LGC24b, LMÖ23, LW21b, MS21d, MS22b, MZS24, TZ23b, VBP22, WS23a, XTT22, YLC20b, ZL22b, SNMM21]. weakening [Gök22]. Weakly [DMP21, SKL23, BR20, DS22b, DHS23, LX23b, LYN22, RVV24, RSA24, XZ23a, YTZ22, ZZL20b]. Wealth [WCY^{+}22, ARMM23, WLS22, ZLH21]. Wealth-based [WCY^{+}22]. Weather [AS21c]. web [KMZ22]. WeChat [ZJ23]. Weierstrass [MSDR21]. weight [CZ22b, CHH22, DZW21, LSWS21, NH21, OPS20, WC24, XFY20, YW22a, KZ22]. weight-balanced [CZ22b]. Weighted [HWC20, MSK21, dSAOG21, CRDD21, CY21b, DO21, GNSZ22, HCTS23, HWLL21, HCL^{+}23, HJY24, LJW^{+}21, LY23b, LX24a, LCWZ24, MZX21, MSMK21, MS24, OT20, PK20, PN24, PV24, TFL22, TTC^{+}23, WF23a, XCGN24, Yun21, ZL21b, Zha21d, ZYW21b]. weighting [CH24b]. weights [JWC20, TFL22]. Weissenberg [Eva20]. Weizhong [RO21]. welfare [LW22f]. Well [BPRR22, CG23, CKWY23, CCCdLMR23, ECS21, GdFPF24, GBCP21, DDCM23, JL24b, KPG22, LLQG21, LC23, Ma23, QYLW24, SBT22, SWCT21, TCGR24, WCXS22, YTC23b, YWT20]. Well-balanced [CKWY23, CCCdLMR23, ECS21, GdFPF24, GBCP21, DDCM23, KPG22, LLQG21, LC23, TCGR24]. well-being [WCXS22]. well-known [QYLW24]. well-mixed [JL24b]. Well-posedness [BPRR22, Ma23, SBT22]. Wendroff [HCTS23, LL23a]. WENO [ARÁSY21, CH24c, GB20, Hu21, HC21, LZ23, RKJ20]. wheeled [CWZL23, HCBY21, MC22]. Whether [YY20b]. Which [LZ21, FMT^{+}21]. whispering [AAL^{+}21]. white [RX20]. who [Wet21a]. Whole [PP21, SS20]. whose [BF23, RS22b, SY22a, SZZ24, XLS21]. width [GA20, ZZZ^{+}23]. Wiener [BDH^{+}20, CH22a, GV24, GWW23, LWZ22c, Spi22, WL20, YZC^{+}22, ZCM^{+}24]. Wilf [AE21]. Williamson [SKK^{+}21]. willingness [HXZ24]. wind [EKKE20, LHZ^{+}23]. wind-borne [EKKE20]. windup [HRA^{+}20]. wing [DDGBGS^{+}24]. winning [SP23]. Wirtinger [YWS24]. wise [LZ23]. Withdrawal [ABNS24b]. Withdrawn [ABNS24c]. within [LGD^{+}21, PPS23, SK24b, SLS24, SC21, WHLL20]. without [CM20a, CFR^{+}20, CLZ23, DJWX21, GPHC24, HLSL22, HZ20a, HZ23b, HJL^{+}23, OMS22, SM24, WY22, Xu24b, ZL20d, ZFS22, ZZLV23]. Wolfe [You20]. women [NS20]. Woodbury [AS23b]. work [Zha21d]. working [Ili22]. world [WMS^{+}22]. worm [GZ20a]. wormhole [LXL22]. wrapped [LZC21]. wt [GN20a]. wt-distance [GN20a]. Wuwei [CWF^{+}21].

X [AHS20, HLLW21]. X-ray [AHS20, HLLW21]. Xiaofeng [RO21]. xVA [Grz22].

years [Zou23]. Yepremyan [CWZ22]. yield [LCJ^{+}22]. yielding [SSC22]. Yorke [SJTCdlFL21]. Yuan [CG20b]. Yukawa [Woo20, IHF13]. Yukawa-type [Woo20, IHF13].

Z [Hu21]. Zagreb [CLB21, JL21, MD24, SLH20, WZ20a, YD22a]. Zakharov [Oru22, LZ24c, Oru21, SW22]. Zero [BBW21, TG20, WFKL24, AFGG20, AFGG21, IKÖ24, JSZ20, KLPL20, LSPK20, LOS21, SJ23b, TTCY23, WLB22, XTS^{+}22, ZCJL21, Zha22a]. zerocoupon [JSZ20]. Zerodeterminant [TG20]. Zero-Dimensional [BBW21]. Zero-sum [WFKL24, IKÖ24, SJ23b]. Zeroing [SMM^{+}24, QNX^{+}23, XLJL22]. Zeros [BSG23, DESM^{+}20, LSVA20, TS22]. zeta [TS22]. Zhang [FRR^{+}22, SSI23]. zigzag [WL20]. Zika [LJHL21]. Zone [GLP20, HL23a, JZS^{+}24, LNZ^{+}22, PX23, WM24a, WPX^{+}20, ZPDL20]. zones [WLY21, WWW24b]. zonotopic [WWW24a]. zooplankton [ZST22, ZLN20].

References

Alizadeh:2020:BCI

[AA20] Behrooz Alizadeh and Esmaeil Afrashteh. Budgetconstrained inverse median facility location problem on tree networks. Applied Mathematics and Computation, 375(??):Article 125078, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                                       http://www.sciencedirect.
                                                       com/science/article/pii/
                                                       S0096300320300473.

Abiad:2021:ECC

[AA21a] Aida Abiad and Carlos A. Alfaro. Enumeration of cospectral and coinvariant graphs. Applied Mathematics and Computation, 408 (??):Article 126348, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004379.

Amiri:2021:ATP

[AA21b] Abdolreza Amiri and Ioannis K. Argyros. On the approximation of m th power divided differences preserving the local order of convergence. Applied Mathematics and Computation, 410(??):Article 126415, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100504X.

Ali:2024:NMS

[AA24] Rashid Ali and Ali Akgul.

A new matrix splitting generalized iteration method for linear complementarity problems. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323005477.

Allahverdi:2020:NWF

[AAA20] Ali Allahverdi, Harun Aydilek, and Asiye Aydilek. No-wait flowshop scheduling problem with separate setup times to minimize total tardiness subject to makespan. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306800.

Ahangar:2021:TRD

[AÁC^{+}21] H. Abdollahzadeh Ahangar, M. P. Álvarez, M. Chellali, S. M. Sheikholeslami, and J. C. Valenzuela-Tripodoro. Triple Roman domination in graphs. Applied Mathematics and Computation, 391 (??):Article 125444, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304057.

Aragon-Artacho:2023:SMR

[AACGTB23] Francisco J. Aragón-Artacho, Yair Censor, Aviv Gibali, and David Torregrosa-Belén. The superiorization method with restarted perturbations for split minimization problems with an application to radiotherapy treatment planning. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007007.

See corrigendum [AACGTB24].

Aragon-Artacho:2024:CSM

[AACGTB24] Francisco J. Aragón-Artacho, Yair Censor, Aviv Gibali, and David Torregrosa-Belén. Corrigendum to: “The superiorization method with restarted perturbations for split minimization problems with an application to radiotherapy treatment planning” [Appl. Math. Comput. 440 (2023) 127627]. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002522.

See [AACGTB23].

Alazemi:2021:MPI

[AACI21] Abdullah Alazemi, Milica Andelić, and Dragana Cvetković-Ilić. The Moore–Penrose inverse of symmetric matrices with nontrivial equitable partitions. Applied Mathematics and Computation, 400(??): Article 126036, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000849.

Al-Aradi:2022:EDG

[AACJ^{+}22] Ali Al-Aradi, Adolfo Correia, Gabriel Jardim, Danilo de Freitas Naiff, and Yuri Saporito. Extensions of the deep Galerkin method. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003617.

Azizul:2021:IHS

[AAHC21] Fatin M. Azizul, Ammar I. Alsabery, Ishak Hashim, and Ali J. Chamkha. Impact of heat source on combined convection flow inside wavywalled cavity filled with nanofluids via heatline concept. Applied Mathematics and Computation, 393(??): Article 125754, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307074.

Abdurahman:2023:FPT

[AAJ23]  Abdujelil  Abdurahman,

Mairemunisa Abudusaimaiti, and Haijun Jiang. Fixed/ predefined-time lag synchronization of complex-valued BAM neural networks with stochastic perturbations. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008797.

Amodio:2021:AAA

[AAL^{+}21] Pierluigi Amodio, Anton Arnold, Tatiana Levitina, Giuseppina Settanni, and Ewa B. Weinmüller. On the Abramov approach for the approximation of whispering gallery modes in prolate spheroids. Applied Mathematics and Computation, 409(??):Article 125599, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305543.

Abia:2022:NIA

[AALMLM22] L. M. Abia, O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. Numerical integration of an age-structured population model with infinite life span. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004751.

Alos:2023:CFR

[AARS23] Elisa Alòs, Fabio Antonelli, Alessandro Ramponi, and Sergio Scarlatti. CVA in fractional and rough volatility models. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007834.

Ahmadinia:2023:NSM

[AAS23] M. Ahmadinia, H. Afshariarjmand, and M. Salehi. Numerical solution of multidimensional Itô Volterra integral equations by the second kind Chebyshev wavelets and parallel computing process. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001571.

Alberto:2020:ITM

[AB20a] José Alberto and Jose Brox. Inverses of k-Toeplitz matrices with applications to resonator arrays with multiple receivers. Applied Mathematics and Computation, 377(??):Article 125185, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301545.

Ayoubi:2020:PES

[AB20b] Tawfiqullah Ayoubi and Haibo Bao. Persistence and extinction in stochastic delay logistic equation by incorporating Ornstein– Uhlenbeck process. Applied Mathematics and Computation, 386(??):Article 125465, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304252.

Amster:2021:PPS

[AB21a] Pablo Amster and Melanie Bondorevsky. Persistence and periodic solutions in systems of delay differential equations. Applied Mathematics and Computation, 403(??):Article 126193, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002836.

Anderson:2021:MLT

[AB21b] Douglas R. Anderson and Martin Bohner. A multivalued logarithm on time scales. Applied Mathematics and Computation, 397 (??):Article 125954, May 15, 2021. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000023.

Abgrall:2023:HFE

[AB23] Rémi Abgrall and Wasilij Barsukow. A hybrid finite element-finite volume method for conservation laws. Applied Mathematics and Computation, 447(??): ??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000152.

Abbaszadeh:2022:NIM

[ABD22] Mostafa Abbaszadeh, Mostafa Bayat, and Mehdi Dehghan. Numerical investigation of the magnetic properties and behavior of electrically conducting fluids via the local weak form method. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003678.

Abgrall:2023:SGF

[ABD23] Rémi Abgrall, Saray Busto, and Michael Dumbser. A simple and general framework for the construction of thermodynamically compatible schemes for computational fluid and solid mechanics. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007020.

Ahmadi:2024:TMN

[ABNS24a] Seyed Mohammad Ahmadi, Morteza Behbahani-Nejad, and Younes Shekari. Transient modeling of natural gas in pipeline networks by two non-iterative explicit and implicit finite volume methods. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002546.

See withdrawal notice [ABNS24b].

Ahmadi:2024:WNT

[ABNS24b] Seyed Mohammad Ahmadi, Morteza Behbahani-Nejad, and Younes Shekari. Withdrawal notice to “Transient modeling of natural gas in pipeline networks by two non-iterative explicit and implicit finite volume methods” [applied mathematics and computation, 479 (2024) 128792]. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003941.

See [ABNS24a, ABNS24c].

Ahmadi:2024:WTM

[ABNS24c] Seyed Mohammad Ahmadi, Morteza Behbahani-Nejad, and Younes Shekari. Withdrawn: Transient modeling of natural gas in pipeline networks by two noniterative explicit and implicit finite volume methods. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002194.

See withdrawal notice [ABNS24b].

Abreu:2023:GIL

[ABPP23] Eduardo Abreu, Elena Bachini, John Pérez, and Mario Putti. A geometrically intrinsic Lagrangian– Eulerian scheme for 2D shallow water equations with variable topography and discontinuous data. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200844X.

Acu:2021:IPS

[ABTR21] Ana-Maria Acu, Gülen Bascanbaz-Tunca, and Ioan Rasa. Information potential for some probability density functions. Applied Mathematics and Computation, 389(??):Article 125578, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305348.

Ammari:2020:FRL

[AC20] Kaïs Ammari and Boumediène Chentouf. Further results on the long-time behavior of a 2D overhead crane with a boundary delay: Exponential convergence. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306903.

Aretaki:2022:ESC

[AC22] Aik. Aretaki and Ch. Chorianopoulos. On the envelope of some classes of 3 by 3 matrices. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003459.

Aguilar-Canto:2024:MDC

[ACBLC24] Fernando Javier Aguilar-Canto, Carlos Brito-Loeza, and Hiram Calvo. Model discovery of compartmental models with Graph-Supported Neural Networks. Applied Mathematics and Computation, 464(??): ??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005611.

Alcantara:2021:NPD

[ACC^{+}21] Adriano A. Alcântara, Bruno A. Carmo, Haroldo R. Clark, Ronald R. Guardia, and Mauro A. Rincon. On a nonlinear problem with Dirichlet and acoustic boundary conditions. Applied Mathematics and Computation, 411 (??):Article 126514, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006032.

Argyros:2020:EAN

[ACGG20] I. K. Argyros, J. Ceballos, D. González, and J. M. Gutiérrez. Extending the applicability of Newton’s method for a class of boundary value problems using the shooting method. Applied Mathematics and Computation, 384 (??):Article 125378, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303428.

Acharya:2021:SQL

[Ach21] Nilankush Acharya. Spectral quasi linearization simulation on the radiative nanofluid spraying over a permeable inclined spinning disk considering the existence of heat source/ sink. Applied Mathematics and Computation, 411 (??):Article 126547, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006317.

Ashpazzadeh:2022:HMR

[ACH^{+}22] Elmira Ashpazzadeh, Yu-Ming Chu, Mir Sajjad Hashemi, Mahsa Moharrami, and Mustafa Inc. Hermite multiwavelets representation for the sparse solution of nonlinear Abel’s integral equation. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002454.

Aahag:2024:GLB

[ÅCL24] P. Åhag, R. Czyz, and P. H. Lundow. On a generalised Lambert W branch transition function arising from p, q-binomial coefficients. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005167.

Alemanno:2023:HMP

[ACMT23] Francesco Alemanno, Luca Camanzi, Gianluca Manzan, and Daniele Tantari. Hopfield model with planted patterns: a teacher-student self-supervised learning model. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004228.

Ahangar:2020:OID

[ACS20] H. Abdollahzadeh Ahangar, M. Chellali, and S. M. Sheikholeslami. Outer independent double Roman domination. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306095.

Antunes:2024:ICM

[ACS24] Pedro R. S. Antunes, Hernani Calunga, and Pedro Serranho. Improving the conditioning of the Method of Fundamental Solutions for the Helmholtz equation on domains in polar or elliptic coordinates. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004302.

Ahangar:2022:MDR

[ACSVT22] H. Abdollahzadeh Ahangar, M. Chellali, S. M. Sheikholeslami, and J. C. Valenzuela-Tripodoro. Maximal double Roman domination in graphs. Applied Mathematics and Computation, 414 (??):Article 126662, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007463.

AlmarazLuengo:2022:NAA

[ACVHC22] Elena Almaraz Luengo, Marcos Brian Leiva Cerna, Luis Javier García Villalba, and Julio Hernandez-Castro. A new approach to analyze the independence of statistical tests of randomness. Applied Mathematics and Computation, 426(??):1–16, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002004.

Abbaszadeh:2021:NIR

[AD21a] Mostafa Abbaszadeh and Mehdi Dehghan. Numerical investigation of reproducing kernel particle Galerkin method for solving fractional modified distributed-order anomalous sub-diffusion equation with error estimation. Applied Mathematics and Computation, 392(??): Article 125718, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306718.

Ali:2021:MIR

[AD21b] Akbar Ali and Tomislav Doslić. Mostar index: Results and perspectives. Applied Mathematics and Computation, 404(??): Article 126245, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003350.

Arshad:2020:SOA

[ADB20] Sadia Arshad, Ozlem Defterli, and Dumitru Baleanu. A second order accurate approximation for fractional derivatives with singular and non-singular kernel applied to a HIV model. Applied Mathematics and Computation, 374 (??):Article 125061, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300308.

AlMugahwi:2021:BMM

[ADF^{+}21] Mohammed Al Mugahwi, Omar De La Cruz Cabrera, Caterina Fenu, Lothar Reichel, and Giuseppe Rodriguez. Block matrix models for dynamic networks. Applied Mathematics and Computation, 402 (??):Article 126121, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001697.

Andova:2024:DN

[ADKS24] Vesna Andova, Pavel Dimovski, Martin Knor, and Riste Skrekovski. Diameter of nanotori. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005118.

Arandiga:2023:NIQ

[ADLU23] Francesc Aràndiga, Rosa Donat, and Sergio López-Ureña. Nonlinear improvements of quasi-interpolanting splines to approximate piecewise smooth functions. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001157.

Arun:2021:AAP

[ADS21] K. R. Arun, A. J. Das Gupta, and S. Samantaray. Analysis of an asymptotic preserving low mach number accurate IMEX–RK scheme for the wave equation system. Applied Mathematics and Computation, 411(??):Article 126469, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005580.

Arandiga:2023:PSU

[ADS23] Francesc Aràndiga, Rosa Donat, and Daniela Schenone. A 2D prediction step using multiquadric local interpolation with adaptive parameter estimation for image compression. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003338.

Alpay:2023:NCB

[ADV23] Daniel Alpay, Kamal Diki, and Mihaela Vajiac. A note on the complex and bicomplex valued neural networks. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000334.

Auli:2021:WEB

[AE21] Juan S. Auli and Sergi Elizalde. Wilf equivalences between vincular patterns in inversion sequences. Applied Mathematics and Computation, 388(??):Article 125514, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304720.

Abdelrahman:2021:VRT

[AEE21] Alaa A. Abdelrahman, Ismail Esen, and Mohamed A. Eltaher. Vibration response of Timoshenko perforated microbeams under accelerating load and thermal environment. Applied Mathematics and Computation, 407(??):Article 126307, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003969.

Albares:2021:DNL

[AEL21] P. Albares, P. G. Estévez, and J. D. Lejarreta. Derivative non-linear Schrödinger equation: Singular manifold method and Lie symmetries. Applied Mathematics and Computation, 400(??): Article 126089, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001375.

Abrahamsen:2021:SKV

[AF21] Dylan Abrahamsen and Bengt Fornberg. Solving the Korteweg–de Vries equation with Hermite-based finite differences. Applied Mathematics and Computation, 401(??):Article 126101, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001491.

Anderson:2024:FIS

[AF24] William Anderson and Mohammad Farazmand. Fisher information and shapemorphing modes for solving the Fokker–Planck equation in higher dimensions. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006586.

Amirali:2024:SON

[AFA24] Ilhame Amirali, Burcu Fedakar, and Gabil M. Amiraliyev. On the second-order neutral Volterra integrodifferential equation and its numerical solution. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002339.

Antonelli:2021:CBI

[AFG21] L. Antonelli, E. Francomano, and F. Gregoretti. A CUDA-based implementation of an improved SPH method on GPU. Applied Mathematics and Computation, 409 (??):Article 125482, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304410.

Avram:2022:COC

[AFG22a] Florin Avram, Lorenzo Freddi, and Dan Goreac. Corrigendum to “Optimal control of a SIR epidemic with ICU constraints and target objectives”. Applied Mathematics and Computation, 423(??): ??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000984.

See [AFG22b].

Avram:2022:OCS

[AFG22b] Florin Avram, Lorenzo Freddi, and Dan Goreac. Optimal control of a SIR epidemic with ICU constraints and target objectives. Applied Mathematics and Computation, 418 (??):Article 126816, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008997. See

corrigendum [AFG22a].

Algaba:2020:ONF

[AFGG20] A. Algaba, N. Fuentes, E. Gamero, and C. García. Orbital normal forms for a class of three-dimensional systems with an application to Hopf-zero bifurcation analysis of Fitzhugh– Nagumo system. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308859.

Algaba:2021:IPH

[AFGG21] A. Algaba, N. Fuentes, E. Gamero, and C. García. On the integrability problem for the Hopf-zero singularity and its relation with the inverse Jacobi multiplier. Applied Mathematics and Computation, 405(??): Article 126241, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003313.

Agliari:2024:SAH

[AFL24] Elena Agliari, Alberto Fachechi, and Domenico Luongo. A spectral approach to Hebbian-like neural networks. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001619.

Ashrafi:2020:NNR

[AG20] Ali Reza Ashrafi and Ali Ghalavand. Note on non-regular graphs with minimal total irregularity. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308835.

Al-Ghassani:2020:EMP

[AGA20] Asma S. Al-Ghassani and Ziyad AlSharawi. The effect of maps permutation on the global attractor of a periodic Beverton–Holt model. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308975.

Alshehri:2020:NIH

[AGC20] Fahad Alshehri, Jaydeep Goraniya, and Madeleine L. Combrinck. Numerical investigation of heat transfer enhancement of a water/ ethylene glycol mixture with Al_{2}O_{3}–TiO_{2} nanoparticles. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308288.

Arellano-Garcia:2021:MLA

[AGCGSC21] María Evarista Arellano-García, José Ariel Camacho-Gutiérrez, and Selene Solorza-Calderón. Machine learning approach for higherorder interactions detection to ecological communities management. Applied Mathematics and Computation, 411(??):Article 126499, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005889.

Aggul:2022:GDS

[Agg22] Mustafa Aggul. A grad– div stabilized penalty projection algorithm for fluid– fluid interaction. Applied Mathematics and Computation, 414(??):Article 126670, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007542.

Acosta:2023:PIR

[AGGPO23] A. Acosta, R. Gallo, P. García, and D. Peluffo-Ordóñez. Positive invariant regions for a modified Van Der Pol equation modeling heart action. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008001.

Alikhanov:2021:HOL

[AH21] Anatoly A. Alikhanov and Chengming Huang. A high-order L2 type difference scheme for the timefractional diffusion equation. Applied Mathematics and Computation, 411 (??):Article 126545, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006299.

Amiri:2020:SCM

[AHB20] Sadegh Amiri, Mojtaba Hajipour, and Dumitru Baleanu. A spectral collocation method with piecewise trigonometric basis functions for nonlinear Volterra– Fredholm integral equations. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309075.

Aristizabal:2023:SSA

[AHEGM23] Mauricio Aristizabal, José L.

Hernández-Estrada, Manuel Garcia, and Harry Millwater. Solution and sensitivity analysis of nonlinear equations using a hypercomplexvariable Newton–Raphson method. Applied Mathematics and Computation, 451 (??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001509.

An:2021:MGC

[AHL21] Phan Thanh An, Phong Thi Thu Huyen, and Nguyen Thi Le. A modified Graham’s convex hull algorithm for finding the connected orthogonal convex hull of a finite planar point set. Applied Mathematics and Computation, 397 (??):Article 125889, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308420.

Ahmad:2021:LBD

[Ahm21] Israr Ahmad. A Lyapunovbased direct adaptive controller for the suppression and synchronization of a perturbed nuclear spin generator chaotic system. Applied Mathematics and Computation, 395(??): Article 125858, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308110.

Alyazidi:2023:OAP

[AHM23] Nezar M. Alyazidi, Abdalrahman M. Hassanine, and Magdi S. Mahmoud. An online adaptive policy iteration-based reinforcement learning for a class of a nonlinear 3D overhead Crane. Applied Mathematics and Computation, 447(??): ??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008785.

Al-Hadeethi:2020:XRA

[AHS20] Y. Al-Hadeethi and M. I.

Sayyed. X-ray attenuation features of some tellurite glasses evaluated at medical diagnostic energies. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307040.

Arancibia-Ibarra:2021:BAP

[AIAFvH21] Claudio Arancibia-Ibarra, Pablo Aguirre, José Flores, and Peter van Heijster. Bifurcation analysis of a predator–prey model with predator intraspecific interactions and ratio-dependent functional response. Applied Mathematics and Computation, 402(??):Article 126152, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002009.

Amangaliyeva:2021:BVP

[AJIR21] Meiramkul Amangaliyeva, Muvasharkhan Jenaliyev, Sagyndyk Iskakov, and Murat Ramazanov. On a boundary value problem for the heat equation and a singular integral equation associated with it. Applied Mathematics and Computation, 399(??): Article 126009, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000576.

Ahmad:2021:BEA

[AK21] Sk. Safique Ahmad and Prince Kanhya. Backward error analysis and inverse eigenvalue problems for Hankel and Symmetric-Toeplitz structures. Applied Mathematics and Computation, 406(??):Article 126288, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003775.

Abolpour:2023:ECS

[AKDR23] Roozbeh Abolpour, Alireza Khayatian, Maryam Dehghani, and Alireza Rokhsari. An equivalent condition for stability analysis of LTI systems with bounded timeinvariant delay. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006580.

Ali:2022:HNS

[AKH^{+}22] Bagh Ali, Shahid Ali Khan, Ahmed Kadhim Hussein, Thirupathi Thumma, and Sajjad Hussain. Hybrid nanofluids: Significance of gravity modulation, heat source/sink, and magnetohydrodynamic on dynamics of micropolar fluid over an inclined surface via finite element simulation. Applied Mathematics and Computation, 419(??): Article 126878, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009619.

Achleitner:2021:MSF

[AKMR21] Franz Achleitner, Christian Kuehn, Jens M. Melenk, and Alexander Rieder. Metastable speeds in the fractional Allen–Cahn equation. Applied Mathematics and Computation, 408 (??):Article 126329, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004185.

Arulperumjothi:2023:RFC

[AKP23] M. Arulperumjothi, Sandi Klavzar, and S. Prabhu. Redefining fractal cubic networks and determining their metric dimension and faulttolerant metric dimension. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002060.

Aljinovic:2021:OIQ

[AKPK21] Andrea Aglić Aljinović, Domagoj Kovacević, Mate Puljiz, and Ana Zgaljić Keko. On Ostrowski inequality for quantum calculus. Applied Mathematics and Computation, 410 (??):Article 126454, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005439.

Alsakaji:2021:DDM

[AKR21] Hebatallah J. Alsakaji, Soumen Kundu, and Fathalla A. Rihan. Delay differential model of one-predator twoprey system with Monod– Haldane and Holling Type II functional responses. Applied Mathematics and Computation, 397(??):Article 125919, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320308729.

Ahn:2021:IMC

[AKY21] Jun-Geol Ahn, Jin-Gyun Kim, and Hyun-Ik Yang. Interpolation multipoint constraints with selection criteria of degree of freedoms for flexible multibody dynamics. Applied Mathematics and Computation, 409 (??):Article 126361, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004501.

Alfifi:2021:SHB

[Alf21]  H. Y. Alfifi.  Stability

and Hopf bifurcation analysis for the diffusive delay logistic population model with spatially heterogeneous environment. Applied Mathematics and Computation, 408(??):Article 126362, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004513.

Alfifi:2023:EDD

[Alf23] H. Y. Alfifi. Effects of diffusion and delayed immune response on dynamic behavior in a viral model. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007822.

Alfifi:2024:SAH

[Alf24] H. Y. Alfifi. Stability analysis and Hopf bifurcation for two-species reaction– diffusion–advection competition systems with two time delays. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001565.

Aljoudi:2021:ESF

[Alj21] Shorog Aljoudi. Exact solutions of the fractional Sharma–Tasso–Olver equation and the fractional Bogoyavlenskii’s breaking soliton equations. Applied Mathematics and Computation, 405(??):Article 126237, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003271.

Ambegoda-Liyanage:2024:ROV

[ALJ24] Prathibha Ambegoda-Liyanage and Sophia R. J. Jang.

Resistance in oncolytic viral therapy for solid tumors. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000183.

Agosti:2023:IIC

[ALL23] A. Agosti, A. Giotta Lucifero, and S. Luzzi. An image-informed Cahn– Hilliard Keller–Segel multiphase field model for tumor growth with angiogenesis. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000036.

Arnas:2020:DVA

[ALM20] David Arnas, Carl Leake, and Daniele Mortari. The n-dimensional k-vector and its application to orthogonal range searching. Applied Mathematics and Computation, 372(??):Article 125010, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310021.

Agliari:2022:SLR

[ALM22] Elena Agliari, Francesca Elisa Leonelli, and Chiara Marullo. Storing, learning and retrieving biased patterns. Applied Mathematics and Computation, 415(??):Article 126716, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008006.

Almushaira:2023:EEP

[Alm23] Mustafa Almushaira. Efficient energy-preserving eighth-order compact finite difference schemes for the sine-Gordon equation. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002084.

Ahasan:2020:GSM

[AM20a] Mohd. Ahasan and M. Mursaleen. Generalized Szász– Mirakjan type operators via q-calculus and approximation properties. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309087.

Alipour:2020:IAS

[AM20b] Sahar Alipour and Farshid Mirzaee. An iterative algorithm for solving two dimensional nonlinear stochastic integral equations: a combined successive approximations method with bilinear spline interpolation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309397.

Assouli:2024:DPI

[AM24] Mouhcine Assouli and Badr Missaoui. Deep Policy Iteration for high-dimensional mean field games. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003849.

Akinlotan:2020:OCM

[AMA20] Morenikeji Deborah Akinlotan, Daniel G. Mallet, and Robyn P. Araujo. An optimal control model of the treatment of chronic Chlamydia trachomatis infection using a combination treatment with antibiotic and tryptophan. Applied Mathematics and Computation, 375(??):Article 124899, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319308914.

Al-Mamun:2020:CTF

[AMBG^{+}20] A. Al-Mamun, J. Barber, V. Ginting, F. Pereira, and A. Rahunanthan. Contaminant transport forecasting in the subsurface using a Bayesian framework. Applied Mathematics and Computation, 387(??): Article 124980, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309725.

Araujo:2020:USN

[AMF20] Rawlilson O. Araújo, Sheyla S. Marinho, and Julio S. Prates Filho. Uniform stability of a non-autonomous semilinear Bresse system with memory. Applied Mathematics and Computation, 387(??):Article 124418, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319303765.

Aksentijevic:2021:NAS

[AMM21] Aleksandar Aksentijevic, Anja Mihailović, and Dragutin T. Mihailović. A novel approach to the study of spatio-temporal brain dynamics using change-based complexity. Applied Mathematics and Computation, 410(??):Article 126432, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100521X.

Ascione:2023:FEO

[AMOS23] Giacomo Ascione, Farshid Mehrdoust, Giuseppe Orlando, and Oldouz Samimi. Foreign exchange options on Heston–CIR model under Lévy process framework. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000206.

Ali:2020:MPM

[AMPS20] Alsadig Ali, Het Mankad, Felipe Pereira, and Fabrício S. Sousa. The multiscale perturbation method for second order elliptic equations. Applied Mathematics and Computation, 387(??): Article 125023, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630031931015X.

Amparan:2021:SFC

[AMZ21] A. Amparan, S. Marcaida, and I. Zaballa. Spectral filters connecting high order systems. Applied Mathematics and Computation, 391 (??):Article 125672, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306251.

Ansarinasab:2024:SDS

[ANG^{+}24] Sheida Ansarinasab, Fahimeh Nazarimehr, Farnaz Ghassemi, Dibakar Ghosh, and Sajad Jafari. Spatial dynamics of swarmalators’ movements. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300677X.

Anonymous:2020:Aa

[Ano20a] Anonymous. 1 April 2020. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ac

[Ano20b]  Anonymous.  1 August

2020. Applied Mathematics and Computation, 378(??): ??, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:D

[Ano20c] Anonymous. 1 December 2020. Applied Mathematics and Computation, 386 (??):??, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Fa

[Ano20d] Anonymous. 1 February 2020. Applied Mathematics and Computation, 366 (??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Jd

[Ano20e] Anonymous. 1 July 2020.

Applied Mathematics and Computation, 376(??):??, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Jb

[Ano20f] Anonymous. 1 June 2020.

Applied Mathematics and Computation, 374(??):??, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ma

[Ano20g] Anonymous. 1 March 2020. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Mc

[Ano20h] Anonymous. 1 May 2020. Applied Mathematics and Computation, 372(??):??, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Na

[Ano20i] Anonymous. 1 November 2020. Applied Mathematics and Computation, 384 (??):??, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Oa

[Ano20j] Anonymous. 1 October 2020. Applied Mathematics and Computation, 382 (??):??, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Sa

[Ano20k] Anonymous. 1 September 2020. Applied Mathematics and Computation, 380 (??):??, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ab

[Ano20l] Anonymous. 15 April 2020.

Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ad

[Ano20m] Anonymous. 15 August 2020. Applied Mathematics and Computation, 379 (??):??, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Fb

[Ano20n] Anonymous. 15 February 2020. Applied Mathematics and Computation, 367 (??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ja

[Ano20o] Anonymous. 15 January 2020. Applied Mathematics and Computation, 365 (??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Je

[Ano20p] Anonymous. 15 July 2020. Applied Mathematics and Computation, 377(??):??, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Jc

[Ano20q] Anonymous. 15 June 2020. Applied Mathematics and Computation, 375(??):??, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Mb

[Ano20r]  Anonymous.  15 March

2020. Applied Mathematics and Computation, 369 (??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Md

[Ano20s] Anonymous. 15 May 2020. Applied Mathematics and Computation, 373(??):??, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Nb

[Ano20t] Anonymous. 15 November 2020. Applied Mathematics and Computation, 385 (??):??, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Ob

[Ano20u] Anonymous. 15 October 2020. Applied Mathematics and Computation, 383 (??):??, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:Sb

[Ano20v] Anonymous. 15 September 2020. Applied Mathematics and Computation, 381 (??):??, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2020:EBa

[Ano20w]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307441.

Anonymous:2020:EBb

[Ano20x]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 365 (??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319307866.

Anonymous:2020:EBc

[Ano20y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308021.

Anonymous:2020:EBd

[Ano20z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308422.

Anonymous:2020:EBe

[Ano20-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 368 (??):??, March 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319308768.

Anonymous:2020:EBf

[Ano20-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 369 (??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309294.

Anonymous:2020:EBg

[Ano20-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 370 (??):??, April 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309683.

Anonymous:2020:EBh

[Ano20-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 371 (??):??, April 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310070.

Anonymous:2020:EBi

[Ano20-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 372 (??):Article 125052, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300217.

Anonymous:2020:EBj

[Ano20-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 373 (??):Article 125099, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300680.

Anonymous:2020:EBk

[Ano20-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 374 (??):Article 125126, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300953.

Anonymous:2020:EBl

[Ano20-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 375 (??):Article 125172, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301417.

Anonymous:2020:EBm

[Ano20-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 376 (??):Article 125219, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301880.

Anonymous:2020:EBn

[Ano20-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 377 (??):Article 125244, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302137.

Anonymous:2020:EBo

[Ano20-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 378(??):Article 125296, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302654.

Anonymous:2020:EBp

[Ano20-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 379(??):Article 125320, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302861.

Anonymous:2020:EBq

[Ano20-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 380 (??):Article 125348, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030312X.

Anonymous:2020:EBr

[Ano20-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 381 (??):Article 125367, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303313.

Anonymous:2020:EBs

[Ano20-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 382(??):Article 125393, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303568.

Anonymous:2020:EBt

[Ano20-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 383 (??):Article 125415, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303763.

Anonymous:2020:EBu

[Ano20-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 384 (??):Article 125439, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304008.

Anonymous:2020:EBv

[Ano20-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 385 (??):Article 125568, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305245.

Anonymous:2020:EBw

[Ano20-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 386(??):Article 125582, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305385.

Anonymous:2020:EBx

[Ano20-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 387 (??):Article 125550, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320305063.

Anonymous:2021:Aa

[Ano21a] Anonymous. 1 April 2021. Applied Mathematics and Computation, 394(??):??, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ac

[Ano21b]  Anonymous.  1 August

2021. Applied Mathematics and Computation, 402(??): ??, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Da

[Ano21c] Anonymous. 1 December 2021. Applied Mathematics and Computation, 410 (??):??, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Fa

[Ano21d] Anonymous. 1 February 2021. Applied Mathematics and Computation, 390 (??):??, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ja

[Ano21e] Anonymous. 1 January 2021. Applied Mathematics and Computation, 388 (??):??, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Je

[Ano21f] Anonymous. 1 July 2021.

Applied Mathematics and Computation, 400(??):??, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Jc

[Ano21g] Anonymous. 1 June 2021.

Applied Mathematics and Computation, 398(??):??, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ma

[Ano21h] Anonymous. 1 March 2021.

Applied Mathematics and Computation, 392(??):??, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Mc

[Ano21i] Anonymous. 1 May 2021.

Applied Mathematics and Computation, 396(??):??, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Na

[Ano21j] Anonymous. 1 November 2021. Applied Mathematics and Computation, 408 (??):??, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Oa

[Ano21k] Anonymous. 1 October 2021. Applied Mathematics and Computation, 406 (??):??, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Sa

[Ano21l] Anonymous. 1 September 2021. Applied Mathematics and Computation, 404 (??):??, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ab

[Ano21m] Anonymous. 15 April 2021. Applied Mathematics and Computation, 395(??):??, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ad

[Ano21n] Anonymous. 15 August 2021. Applied Mathematics and Computation, 403 (??):??, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Db

[Ano21o] Anonymous. 15 December 2021. Applied Mathematics and Computation, 411 (??):??, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Fb

[Ano21p] Anonymous. 15 February 2021. Applied Mathematics and Computation, 391 (??):??, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Jb

[Ano21q] Anonymous. 15 January 2021. Applied Mathematics and Computation, 389 (??):??, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Jf

[Ano21r] Anonymous. 15 July 2021.

Applied Mathematics and Computation, 401(??):??, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Jd

[Ano21s] Anonymous. 15 June 2021. Applied Mathematics and Computation, 399(??):??, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Mb

[Ano21t]  Anonymous.  15 March

2021. Applied Mathematics and Computation, 393 (??):??, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Md

[Ano21u] Anonymous. 15 May 2021. Applied Mathematics and Computation, 397(??):??, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Nb

[Ano21v] Anonymous. 15 November 2021. Applied Mathematics and Computation, 409 (??):??, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Ob

[Ano21w] Anonymous. 15 October 2021. Applied Mathematics and Computation, 407 (??):??, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:Sb

[Ano21x] Anonymous. 15 September 2021. Applied Mathematics and Computation, 405 (??):??, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2021:EBa

[Ano21y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 388 (??):Article 125657, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306111.

Anonymous:2021:EBb

[Ano21z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 389 (??):Article 125687, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306408.

Anonymous:2021:EBc

[Ano21-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 390 (??):Article 125738, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306913.

Anonymous:2021:EBd

[Ano21-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 391 (??):Article 125774, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030727X.

Anonymous:2021:EBe

[Ano21-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 392 (??):Article 125802, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307554.

Anonymous:2021:EBf

[Ano21-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 393 (??):Article 125840, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307931.

Anonymous:2021:EBg

[Ano21-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 394 (??):Article 125895, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308481.

Anonymous:2021:EBh

[Ano21-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 395 (??):Article 125946, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308997.

Anonymous:2021:EBi

[Ano21-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 396 (??):Article 125971, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000199.

Anonymous:2021:EBj

[Ano21-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 397 (??):Article 126027, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000758.

Anonymous:2021:EBk

[Ano21-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 398 (??):Article 126067, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001156.

Anonymous:2021:EBl

[Ano21-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 399 (??):Article 126110, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001582.

Anonymous:2021:EBm

[Ano21-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 400 (??):Article 126143, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001910.

Anonymous:2021:EBn

[Ano21-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 401 (??):Article 126160, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002502.

Anonymous:2021:EBo

[Ano21-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 402(??):Article 126212, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003027.

Anonymous:2021:EBp

[Ano21-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 403(??):Article 126277, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003660.

Anonymous:2021:EBq

[Ano21-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 404 (??):Article 126294, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003830.

Anonymous:2021:EBr

[Ano21-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 405 (??):Article 126314, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004033.

Anonymous:2021:EBs

[Ano21-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 406(??):Article 126353, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004422.

Anonymous:2021:EBt

[Ano21-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 407 (??):Article 126394, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004835.

Anonymous:2021:EBu

[Ano21-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 408 (??):Article 126429, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100518X.

Anonymous:2021:EBv

[Ano21-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 409 (??):Article 126508, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100597X.

Anonymous:2021:EBw

[Ano21-47]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 410(??):Article 126555, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006391.

Anonymous:2021:EBx

[Ano21-48]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 411 (??):Article 126629, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100713X.

Anonymous:2022:Aa

[Ano22a] Anonymous. 1 April 2022.

Applied Mathematics and Computation, 418(??):??, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ac

[Ano22b]  Anonymous.  1 August

2022. Applied Mathematics and Computation, 426(??): ??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Da

[Ano22c] Anonymous. 1 December 2022. Applied Mathematics and Computation, 434 (??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Fa

[Ano22d] Anonymous. 1 February 2022. Applied Mathematics and Computation, 414 (??):??, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ja

[Ano22e] Anonymous. 1 January 2022. Applied Mathematics and Computation, 412 (??):??, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Je

[Ano22f] Anonymous. 1 July 2022.

Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Jc

[Ano22g] Anonymous. 1 June 2022.

Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ma

[Ano22h] Anonymous. 1 March 2022.

Applied Mathematics and Computation, 416(??):??, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Mc

[Ano22i] Anonymous. 1 May 2022.

Applied Mathematics and Computation, 420(??):??, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Na

[Ano22j] Anonymous. 1 November 2022. Applied Mathematics and Computation, 432 (??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Oa

[Ano22k] Anonymous. 1 October 2022. Applied Mathematics and Computation, 430 (??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Sa

[Ano22l] Anonymous. 1 September 2022. Applied Mathematics and Computation, 428 (??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ab

[Ano22m] Anonymous. 15 April 2022. Applied Mathematics and Computation, 419(??):??, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ad

[Ano22n] Anonymous. 15 August 2022. Applied Mathematics and Computation, 427 (??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Db

[Ano22o] Anonymous. 15 December 2022. Applied Mathematics and Computation, 435 (??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Fb

[Ano22p] Anonymous. 15 February 2022. Applied Mathematics and Computation, 415 (??):??, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Jb

[Ano22q] Anonymous. 15 January 2022. Applied Mathematics and Computation, 413 (??):??, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Jf

[Ano22r] Anonymous. 15 July 2022.

Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Jd

[Ano22s] Anonymous. 15 June 2022.

Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Mb

[Ano22t]  Anonymous.  15 March

2022. Applied Mathematics and Computation, 417 (??):??, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:M

[Ano22u] Anonymous. 15 May 2022. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Nb

[Ano22v] Anonymous. 15 November 2022. Applied Mathematics and Computation, 433 (??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Ob

[Ano22w] Anonymous. 15 October 2022. Applied Mathematics and Computation, 431 (??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:Sb

[Ano22x] Anonymous. 15 September 2022. Applied Mathematics and Computation, 429 (??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2022:EBa

[Ano22y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 412 (??):Article 126675, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007591.

Anonymous:2022:EBb

[Ano22z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 413 (??):Article 126708, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100792X.

Anonymous:2022:EBc

[Ano22-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 414 (??):Article 126744, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008262.

Anonymous:2022:EBd

[Ano22-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 415 (??):Article 126765, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100847X.

Anonymous:2022:EBe

[Ano22-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 416 (??):Article 126803, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008857.

Anonymous:2022:EBf

[Ano22-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 417 (??):Article 126831, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009140.

Anonymous:2022:EBg

[Ano22-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 418 (??):Article 126890, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009735.

Anonymous:2022:EBh

[Ano22-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 419 (??):Article 126920, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322000066.

Anonymous:2022:EBi

[Ano22-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 420 (??):Article 126980, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322000662.

Anonymous:2022:EBj

[Ano22-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 421 (??):??, May 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032200087X.

Anonymous:2022:EBk

[Ano22-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 422 (??):??, June 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322001084.

Anonymous:2022:EBl

[Ano22-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 423 (??):??, June 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322001515.

Anonymous:2022:EBm

[Ano22-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 424 (??):??, July 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322001886.

Anonymous:2022:EBn

[Ano22-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 425 (??):??, July 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322002387.

Anonymous:2022:EBo

[Ano22-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 426 (??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322002715.

Anonymous:2022:EBp

[Ano22-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 427 (??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032200296X.

Anonymous:2022:EBq

[Ano22-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003381.

Anonymous:2022:EBr

[Ano22-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003563.

Anonymous:2022:EBs

[Ano22-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200409X.

Anonymous:2022:EBt

[Ano22-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004581.

Anonymous:2022:EBu

[Ano22-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200501X.

Anonymous:2022:EBv

[Ano22-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005410.

Anonymous:2022:EBw

[Ano22-47]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200563X.

Anonymous:2022:EBx

[Ano22-48]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006038.

Anonymous:2023:Aa

[Ano23a] Anonymous. 1 April 2023.

Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ac

[Ano23b]  Anonymous.  1 August

2023. Applied Mathematics and Computation, 450(??):

??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Da

[Ano23c] Anonymous. 1 December 2023. Applied Mathematics and Computation, 458 (??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Fa

[Ano23d] Anonymous. 1 February 2023. Applied Mathematics and Computation, 438 (??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ja

[Ano23e] Anonymous. 1 January 2023. Applied Mathematics and Computation, 436 (??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Jf

[Ano23f] Anonymous. 1 July 2023. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Jd

[Ano23g] Anonymous. 1 June 2023.

Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ma

[Ano23h] Anonymous. 1 March 2023.

Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Mc

[Ano23i] Anonymous. 1 May 2023.

Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Na

[Ano23j] Anonymous. 1 November 2023. Applied Mathematics and Computation, 456 (??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Oa

[Ano23k] Anonymous. 1 October 2023. Applied Mathematics and Computation, 454 (??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Sa

[Ano23l] Anonymous. 1 September 2023. Applied Mathematics and Computation, 452 (??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ab

[Ano23m] Anonymous. 15 April 2023. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ad

[Ano23n] Anonymous. 15 August 2023. Applied Mathematics and Computation, 451 (??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Db

[Ano23o] Anonymous. 15 December 2023. Applied Mathematics and Computation, 459 (??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Fb

[Ano23p] Anonymous. 15 February 2023. Applied Mathematics and Computation, 439 (??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Jb

[Ano23q] Anonymous. 15 January 2023. Applied Mathematics and Computation, 437 (??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Jg

[Ano23r] Anonymous. 15 July 2023.

Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Je

[Ano23s] Anonymous. 15 June 2023.

Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Mb

[Ano23t]  Anonymous.  15 March

2023. Applied Mathematics and Computation, 441 (??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Md

[Ano23u] Anonymous. 15 May 2023. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Nb

[Ano23v] Anonymous. 15 November 2023. Applied Mathematics and Computation, 457 (??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Ob

[Ano23w] Anonymous. 15 October 2023. Applied Mathematics and Computation, 455 (??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:Sb

[Ano23x] Anonymous. 15 September 2023. Applied Mathematics and Computation, 453 (??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2023:EBa

[Ano23y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 436 (??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032200635X.

Anonymous:2023:EBb

[Ano23z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 437 (??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032200666X.

Anonymous:2023:EBc

[Ano23-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007214.

Anonymous:2023:EBd

[Ano23-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007573.

Anonymous:2023:EBe

[Ano23-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 440 (??):??, March 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007950.

Anonymous:2023:EBf

[Ano23-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 441 (??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322008177.

Anonymous:2023:EBg

[Ano23-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 442 (??):??, April 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322008657.

Anonymous:2023:EBh

[Ano23-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 443 (??):??, April 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000085.

Anonymous:2023:EBi

[Ano23-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 444 (??):??, May 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000292.

Anonymous:2023:EBj

[Ano23-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 445 (??):??, May 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000656.

Anonymous:2023:EBk

[Ano23-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 446 (??):??, June 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000978.

Anonymous:2023:EBl

[Ano23-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 447 (??):??, June 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001248.

Anonymous:2023:EBm

[Ano23-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 448 (??):??, July 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001443.

Anonymous:2023:EBn

[Ano23-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 449 (??):??, July 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001716.

Anonymous:2023:EBo

[Ano23-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 450 (??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323002308.

Anonymous:2023:EBp

[Ano23-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 451 (??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323002606.

Anonymous:2023:EBq

[Ano23-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300276X.

Anonymous:2023:EBr

[Ano23-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002874.

Anonymous:2023:EBs

[Ano23-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003156.

Anonymous:2023:EBt

[Ano23-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003296.

Anonymous:2023:EBu

[Ano23-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003478.

Anonymous:2023:EBv

[Ano23-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004289.

Anonymous:2023:EBw

[Ano23-47]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004526.

Anonymous:2023:EBx

[Ano23-48]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004769.

Anonymous:2024:Aa

[Ano24a] Anonymous. 1 April 2024.

Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ac

[Ano24b]  Anonymous.  1 August

2024. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Da

[Ano24c] Anonymous. 1 December 2024. Applied Mathematics and Computation, 482 (??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Fa

[Ano24d] Anonymous. 1 February 2024. Applied Mathematics and Computation, 462 (??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ja

[Ano24e] Anonymous. 1 January 2024. Applied Mathematics and Computation, 460 (??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Je

[Ano24f] Anonymous. 1 July 2024. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Jc

[Ano24g] Anonymous. 1 June 2024. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ma

[Ano24h] Anonymous. 1 March 2024.

Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Mc

[Ano24i] Anonymous. 1 May 2024.

Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Na

[Ano24j] Anonymous. 1 November 2024. Applied Mathematics and Computation, 480 (??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Oa

[Ano24k] Anonymous. 1 October 2024. Applied Mathematics and Computation, 478 (??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Sa

[Ano24l] Anonymous. 1 September 2024. Applied Mathematics and Computation, 476 (??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ab

[Ano24m] Anonymous. 15 April 2024. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ad

[Ano24n] Anonymous. 15 August 2024. Applied Mathematics and Computation, 475 (??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Db

[Ano24o] Anonymous. 15 December 2024. Applied Mathematics and Computation, 483 (??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Fb

[Ano24p] Anonymous. 15 February 2024. Applied Mathematics and Computation, 463 (??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Jb

[Ano24q] Anonymous. 15 January 2024. Applied Mathematics and Computation, 461 (??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Jf

[Ano24r] Anonymous. 15 July 2024.

Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Jd

[Ano24s] Anonymous. 15 June 2024.

Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Mb

[Ano24t]  Anonymous.  15 March

2024. Applied Mathematics and Computation, 465 (??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Md

[Ano24u] Anonymous. 15 May 2024.

Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Nb

[Ano24v] Anonymous. 15 November 2024. Applied Mathematics and Computation, 481 (??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Ob

[Ano24w] Anonymous. 15 October 2024. Applied Mathematics and Computation, 479 (??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:Sb

[Ano24x] Anonymous. 15 September 2024. Applied Mathematics and Computation, 477 (??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2024:EBa

[Ano24y]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 460 (??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323004964.

Anonymous:2024:EBb

[Ano24z]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 461 (??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323005234.

Anonymous:2024:EBc

[Ano24-27]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005544.

Anonymous:2024:EBd

[Ano24-28]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005866.

Anonymous:2024:EBe

[Ano24-29]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 464 (??):??, March 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006070.

Anonymous:2024:EBf

[Ano24-30]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 465 (??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006239.

Anonymous:2024:EBg

[Ano24-31]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 466 (??):??, April 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006823.

Anonymous:2024:EBh

[Ano24-32]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 467 (??):??, April 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000122.

Anonymous:2024:EBi

[Ano24-33]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 468 (??):??, May 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000389.

Anonymous:2024:EBj

[Ano24-34]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 469 (??):??, May 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000717.

Anonymous:2024:EBk

[Ano24-35]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 470 (??):??, June 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000948.

Anonymous:2024:EBl

[Ano24-36]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 471 (??):??, June 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324001322.

Anonymous:2024:EBm

[Ano24-37]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 472 (??):??, July 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324001656.

Anonymous:2024:EBn

[Ano24-38]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 473 (??):??, July 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324002054.

Anonymous:2024:EBo

[Ano24-39]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 474 (??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324002388.

Anonymous:2024:EBp

[Ano24-40]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 475 (??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324002698.

Anonymous:2024:EBq

[Ano24-41]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003102.

Anonymous:2024:EBr

[Ano24-42]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003308.

Anonymous:2024:EBs

[Ano24-43]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003552.

Anonymous:2024:EBt

[Ano24-44]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003990.

Anonymous:2024:EBu

[Ano24-45]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004259.

Anonymous:2024:EBv

[Ano24-46]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004594.

Anonymous:2024:EBw

[Ano24-47]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004740.

Anonymous:2024:EBx

[Ano24-48]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004971.

Ali:2020:DSA

[ANS^{+}20] M. Syed Ali, Govindasamy Narayanan, Vineet Shekher, Hamed Alsulami, and Tareq Saeed. Dynamic stability analysis of stochastic fractional-order memristor fuzzy BAM neural networks with delay and leakage terms. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308884.

Andersson:2021:DLA

[AO21a] Kristoffer Andersson and Cornelis W. Oosterlee. A deep learning approach for computations of exposure profiles for highdimensional Bermudan options. Applied Mathematics and Computation, 408 (??):Article 126332, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004215.

Andersson:2021:DLC

[AO21b] Kristoffer Andersson and Cornelis W. Oosterlee. Deep learning for CVA computations of large portfolios of financial derivatives. Applied Mathematics and Computation, 409 (??):Article 126399, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004884.

Abolarinwa:2020:LSM

[AOSO20] A. Abolarinwa, N. K. Oladejo, S. O. Salawu, and C. A. Onate. Logarithmic-Sobolev and multilinear Hölder’s inequalities via heat flow monotonicity formulas. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306320.

Arshad:2020:MME

[AP20] Muhammad Arshad and Eun-Jae Park. Multiscale mortar expanded mixed discretization of nonlinear elliptic problems. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309245.

Appleby:2021:GFP

[AP21] John A. D. Appleby and Denis D. Patterson. Growth and fluctuation in perturbed nonlinear Volterra equations. Applied Mathematics and Computation, 396(??): Article 125938, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308912.

Acharya:2022:PHF

[AP22] Sanjib Kumar Acharya and Kamana Porwal. Primal hybrid finite element method for the linear elasticity problem. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005367.

Arnas:2020:RSM

[AR20] David Arnas and Marcos Rodríguez. Range searching in multidimensional databases using navigation metadata. Applied Mathematics and Computation, 386(??):Article 125510, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304689.

Ahn:2021:RBS

[AR21a] Jeongho Ahn and Zachary Rail. A rod-beam system with dynamic contact and thermal exchange condition. Applied Mathematics and Computation, 388 (??):Article 125542, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304987.

Akbas:2021:MGD

[AR21b] Mine Akbas and Leo G. Rebholz. Modular grad-div stabilization for the incompressible non-isothermal fluid flows. Applied Mathematics and Computation, 393(??):Article 125748, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307013.

Alazwari:2022:UAL

[AR22] Mashhour A. Alazwari and Singiresu S. Rao. Uncertainty analysis of large structures using universal grey number theory. Applied Mathematics and Computation, 416(??):Article 126735, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008171.

Alpar:2023:DTC

[AR23] Sultan Alpar and Bolatbek Rysbaiuly. Determination of thermophysical characteristics in a nonlinear inverse heat transfer problem. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007275.

Anandan:2024:ECS

[AR24] Megala Anandan and S. V. Raghurama Rao. Entropy conserving/stable schemes for a vector-kinetic model of hyperbolic systems. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005799.

Amat:2021:CAW

[ARÁSY21] Sergio Amat, Juan Ruiz-Álvarez, Chi-Wang Shu, and Dionisio F. Yáñez. Cellaverage WENO with progressive order of accuracy close to discontinuities with applications to signal processing. Applied Mathematics and Computation, 403(??):Article 126131, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100179X.

Ahmadian:2020:RNA

[ARISV20] D. Ahmadian, O. Farkhondeh Rouz, K. Ivaz, and A. Safdari-Vaighani. Robust numerical algorithm to the European option with illiquid markets. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930685X.

Afhami:2023:CAO

[ARMM23] Bahareh Afhami, Mohsen Rezapour, Mohsen Madadi, and Vahed Maroufy. A comonotonic approximation to optimal terminal wealth under a multivariate Merton model with correlated jump risk. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008761.

Allouch:2021:SMB

[ARST21a] C. Allouch, S. Remogna, D. Sbibih, and M. Tahrichi. Superconvergent methods based on quasi-interpolating operators for Fredholm integral equations of the second kind. Applied Mathematics and Computation, 404(??): Article 126227, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003179.

Altybay:2021:HES

[ARST21b] Arshyn Altybay, Michael Ruzhansky, Mohammed Elamine Sebih, and Niyaz Tokmagambetov. The heat equation with strongly singular potentials. Applied Mathematics and Computation, 399(??):Article 126006, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000540.

Adhikari:2020:PIA

[AS20] Balakrishna Adhikari and B. N. Singh. Parametric instability analysis of laminated composite plate subject to various types of non-uniform periodic inplane edge load. Applied Mathematics and Computation, 373(??):Article 125026, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319310185.

Ackley:2021:LDN

[AS21a] Matthew Ackley and Peter Stechlinski. Lexicographic derivatives of nonsmooth glucose-insulin kinetics under normal and artificial pancreatic responses. Applied Mathematics and Computation, 395 (??):Article 125876, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308298.

Allahviranloo:2021:RKM

[AS21b] Tofigh Allahviranloo and Hussein Sahihi. Reproducing kernel method to solve fractional delay differential equations. Applied Mathematics and Computation, 400(??):Article 126095, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001430.

Amirfakhrian:2021:WDD

[AS21c] Majid Amirfakhrian and Faramarz Samavati. Weather daily data approximation using point adaptive ellipsoidal neighborhood in scattered data interpolation methods. Applied Mathematics and Computation, 392(??):Article 125717, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306706.

Arthi:2021:CHO

[AS21d] G. Arthi and K. Suganya. Controllability of higher order stochastic fractional control delay systems involving damping behavior. Applied Mathematics and Computation, 410(??):Article 126439, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005282.

Abbas:2022:SDF

[AS22] Suhail Abbas and Abdullah Shah. Simulation of different flow regimes in a narrow-gap spherical Couette flow. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000157.

Adloo:2023:OBA

[AS23a] Hassan Adloo and Mohammad Hossein Shafiei. Observer-based adaptive event-triggered control method for input-saturated systems. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004046.

Ameli:2023:SWP

[AS23b] Siavash Ameli and Shawn C. Shadden. A singular Woodbury and pseudodeterminant matrix identities and application to Gaussian process regression. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002011.

Amin:2021:CAN

[ASAK21] Rohul Amin, Kamal Shah, Muhammad Asif, and Imran Khan. A computational algorithm for the numerical solution of fractional order delay differential equations. Applied Mathematics and Computation, 402 (??):Article 125863, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030816X.

Alrebdi:2023:TSI

[ASAZ23] H. I. Alrebdi, Andre Steklain, Edgard P. M. Amorim, and Euaggelos Zotos. Thermostated susceptible-infected-susceptible epidemic model. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200769X.

Askar:2021:CDC

[Ask21] S. S. Askar. On complex dynamics of Cournot-Bertrand game with asymmetric market information. Applied Mathematics and Computation, 393(??): Article 125823, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307761.

Askar:2022:DCD

[Ask22] S. S. Askar. On the dynamics of Cournot duopoly game with private firms: Investigations and analysis. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004283.

Aravindh:2020:FTR

[ASKA20] D. Aravindh, R. Sakthivel, Fanchao Kong, and S. Marshal Anthoni. Finitetime reliable stabilization of uncertain semi-Markovian jump systems with input saturation. Applied Mathematics and Computation, 384(??):Article 125388, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303519.

Aarthika:2022:CAT

[ASR22] K. Aarthika, V. Shanthi, and Higinio Ramos. A computational approach for a two-parameter singularly perturbed system of partial differential equations with discontinuous coefficients. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004830.

Alshenawy:2023:TDN

[ASS^{+}23a] Reda Alshenawy, Saeid Sahmani, Babak Safaei, Yasser Elmoghazy, Ali Al-Alwan, and Muneerah Al Nuwairan. Threedimensional nonlinear stability analysis of axialthermal-electrical loaded FG piezoelectric microshells via MKM strain gradient formulations. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006968.

Aravinth:2023:IOF

[ASS^{+}23b] N. Aravinth, T. Satheesh, R. Sakthivel, G. Ran, and A. Mohammadzadeh. Input-output finite-time stabilization of periodic piecewise systems with multiple disturbances. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002497.

Ahmadi:2024:LPR

[ASS24] Azam Sadat Ahmadi, Nasrin Soltankhah, and Babak Samadi. Limited packings: Related vertex partitions and duality issues. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000857.

Ahmed:2021:MRB

[ASUH21] Dilan Ahmed, Fazlollah Soleymani, Malik Zaka Ullah, and Hataw Hasan. Managing the risk based on entropic value-at-risk under a normal-Rayleigh distribution. Applied Mathematics and Computation, 402(??):Article 126129, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001776.

Arregui:2022:TVA

[ASV22] Iñigo Arregui, Roberta Simonella, and Carlos Vázquez. Total value adjustment for European options in a multi-currency setting. Applied Mathematics and Computation, 413(??):Article 126647, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007311.

Attipoe:2021:CMF

[AT21a] David Sena Attipoe and Antoine Tambue. Convergence of the mimetic finite difference and fitted mimetic finite difference method for options pricing. Applied Mathematics and Computation, 401 (??):Article 126060, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001089.

Aydogmus:2021:MMS

[AT21b] Ozgur Aydogmus and Ali Hakan Tor. A modified multiple shooting algorithm for parameter estimation in ODEs using adjoint sensitivity analysis. Applied Mathematics and Computation, 390 (??):Article 125644, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305981.

Arefin:2024:CIS

[AT24] Md. Rajib Arefin and Jun Tanimoto. Coupling injunctive social norms with evolutionary games. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300632X.

An:2022:EPD

[ATDJ22] Shujuan An, Kai Tian, Zhaodong Ding, and Yongjun Jian. Electroosmotic and pressure-driven slip flow of fractional viscoelastic fluids in microchannels. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001576.

Alfaro:2021:CSS

[AV21a] Carlos A. Alfaro and Ralihe R. Villagrán. Corrigendum to “The structure of sandpile groups of outerplanar graphs” [Applied Mathematics and Computation 395 (2021) 125861]. Applied Mathematics and Computation, 398(??):Article 126022, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000709. See

[AV21b].

Alfaro:2021:SSG

[AV21b] Carlos A. Alfaro and Ralihe R. Villagrán. The structure of sandpile groups of outerplanar graphs. Applied Mathematics and Computation, 395(??):Article 125861, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308146. See

corrigendum [AV21a].

Ambika:2024:MAP

[AVJ24] K. Ambika, K. V. Vijayashree, and B. Janani. Modelling and analysis of production management system using vacation queueing theoretic approach. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003175.

Arandiga:2022:AIM

[AY22] Francesc Aràndiga and Dionisio F. Yáñez. Adaptive interpolation with maximum order close to discontinuities. Applied Mathematics and Computation, 418 (??):Article 126795, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008778.

Al-Yakoob:2022:HDS

[AYKS22] Salem Al-Yakoob, Ali Kanso, and Dragan Stevanović. On Hosoya’s dormants and sprouts. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003071.

Al-Yakoob:2020:TIS

[AYS20] Salem Al-Yakoob and Dragan Stevanović. On transmission irregular starlike trees. Applied Mathematics and Computation, 380(??): Article 125257, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302265.

Al-Yakoob:2022:STI

[AYS22] Salem Al-Yakoob and Dragan Stevanović. On stepwise transmission irregular graphs. Applied Mathematics and Computation, 413 (??):Article 126607, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006913.

Ariizumi:2024:CPL

[AYY24] Shumpei Ariizumi, Yuya Yamakawa, and Nobuo Yamashita. Convergence properties of Levenberg– Marquardt methods with generalized regularization terms. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005349.

Ai:2024:GAP

[AZD^{+}24] Jiali Ai, Chi Zhai, Hongyu Du, Yi Dang, Jindong Dai, and Wei Sun. Grid anisotropy of propagation fronts in cellular automata and its reduction methods. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004326.

Aziz:2021:GSF

[Azi21] Taha Aziz. Generalized Stokes’ flow and radiative heat transfer model of a non-Newtonian fluid in a Darcy porous medium subject to Navier’s slip conditions on the penetrable porous boundary: Group theoretical and compatibility analysis. Applied Mathematics and Computation, 400(??):Article 126048, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000965.

Aida-zade:2022:RST

[AzR22] Kamil Aida-zade and Anar Rahimov. On recovering space or time-dependent source functions for a parabolic equation with nonlocal conditions. Applied Mathematics and Computation, 419(??):Article 126849, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009322.

Bas:2020:DSP

[BA20] Erdal Bas and Bahar Acay.

The direct spectral problem via local derivative including truncated Mittag-Leffler function. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307799.

Babaei:2022:FVS

[Bab22] Hadi Babaei. Free vibration and snap-through instability of FG–CNTRC shallow arches supported on nonlinear elastic foundation. Applied Mathematics and Computation, 413 (??):Article 126606, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006901.

Buyantogtokh:2021:MSS

[BAH^{+}21]  Lkhagva  Buyantogtokh,

Enkhbayar Azjargal, Batmend Horoldagva, Shiikhar Dorjsembe, and Damchaa Adiyanyam. On the maximum size of stepwise irregular graphs. Applied Mathematics and Computation, 392(??):Article 125683, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306366.

Boaretto:2024:UER

[BAL^{+}24] B. R. R. Boaretto, A. C. Andreani, S. R. Lopes, T. L. Prado, and E. E. N. Macau. The use of entropy of recurrence microstates and artificial intelligence to detect cardiac arrhythmia in ECG records. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002108.

Bazighifan:2020:OCF

[Baz20] Omar Bazighifan. On the oscillation of certain fourthorder differential equations with p-Laplacian like operator. Applied Mathematics and Computation, 386(??):Article 125475, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304343.

Berezansky:2020:SEL

[BB20] Leonid Berezansky and Elena Braverman. Solution estimates for linear differential equations with delay. Applied Mathematics and Computation, 372(??): Article 124962, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309543.

Bairagi:2021:BFM

[BBAS21] Nandadulal Bairagi, Santanu Bhattacharya, Pierre Auger, and Biswajit Sarkar. Bioeconomics fishery model in presence of infection: Sustainability and demandprice perspectives. Applied Mathematics and Computation, 405(??):Article 126225, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003155.

Baudon:2024:IBV

[BBB24] Olivier Baudon, Julien Bensmail, and Morgan Boivin. Interplays between variations of arbitrarily partitionable graphs under minimality constraints. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002236.

Bisi:2022:NMB

[BBD^{+}22] Marzia Bisi, Walter Boscheri, Giacomo Dimarco, Maria Groppi, and Giorgio Martalò. A new mixed Boltzmann–BGK model for mixtures solved with an IMEX finite volume scheme on unstructured meshes. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004908.

Bonomo-Braberman:2020:NHK

[BBDVPV20] Flavia Bonomo-Braberman, Mitre C. Dourado, Mario Valencia-Pabon, and Juan C. Vera. A note on homomorphisms of Kneser hypergraphs. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307568.

Bahbouhi:2024:EIC

[BBEM24] Jalal Eddine Bahbouhi, Saif Islam Bouderba, Abdelali Elkouay, and Najem Moussa. Effects of individual and collective decision rule on cooperation in public goods game.

Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000250.

Behr:2021:GTS

[BBH21] Maximilian Behr, Peter Benner, and Jan Heiland. Galerkin trial spaces and Davison-maki methods for the numerical solution of differential Riccati equations. Applied Mathematics and Computation, 410 (??):Article 126401, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004902.

Banaji:2022:ASC

[BBH22] Murad Banaji, Balázs Boros, and Josef Hofbauer. Adding species to chemical reaction networks: Preserving rank preserves nondegenerate behaviours. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200193X.

Barilla:2021:ERD

[BBHM21] David Barilla, Martin Bohner, Shapour Heidarkhani, and Shahin Moradi. Existence results for dynamic Sturm–Liouville boundary value problems via variational methods. Applied Mathematics and Computation, 409(??):Article 125614, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320305622.

Bednarz:2024:SDL

[BBK^{+}24] Monika Bednarz, Agnieszka Burkot, Jakub Kwaśny, Kamil Pawłowski, and Angelika Ryngier. Small dense on-line arbitrarily partitionable graphs. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000547.

Burrage:2021:RPT

[BBM21] Kevin Burrage, Pamela Burrage, and Shev Mac-Namara. The reflectionless properties of Toeplitz waves and Hankel waves: an analysis via Bessel functions. Applied Mathematics and Computation, 389 (??):Article 125576, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305324.

Belardo:2021:LEU

[BBT21] Francesco Belardo, Maurizio Brunetti, and Vilmar Trevisan. Locating eigenvalues of unbalanced unicyclic signed graphs. Applied Mathematics and Computation, 400(??):Article 126082, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001302.

Berghammer:2021:ACZ

[BBW21] Rudolf Berghammer, Steffen Börm, and Michael Winter. Algorithmic counting of zero-dimensional finite topological spaces with respect to the covering dimension. Applied Mathematics and Computation, 389(??):Article 125523, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304811.

Benitez:2020:CCE

[BCCI20] Julio Benítez, Silvia Carpitella, Antonella Certa, and Joaquín Izquierdo. Constrained consistency enforcement in AHP. Applied Mathematics and Computation, 380(??): Article 125273, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302423.

Bay:2023:HIS

[BCHRS23] Thierry Bay, Isabelle Cattiaux-Huillard, Lucia Romani, and Laura Saini. On G^{1} and G^{2} Hermite interpolation by spatial Algebraic– Trigonometric Pythagorean Hodograph curves with polynomial parametric speed. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004095.

Balsa-Canto:2020:TDM

[BCLNV20] Eva Balsa-Canto, Alejandro López-Núñez, and Carlos Vázquez. A twodimensional multi-species model for different Listeria monocytogenes biofilm structures and its numerical simulation. Applied Mathematics and Computation, 384(??):Article 125383, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303465.

Borri:2023:QTS

[BCP23] Alessandro Borri, Francesco Carravetta, and Pasquale Palumbo. Quadratized Taylor series methods for ODE numerical integration. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300406X.

Bermudo:2022:VDB

[BCR22] Sergio Bermudo, Roberto Cruz, and Juan Rada. Vertex-degree-based topological indices of oriented trees. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004696.

Blanes:2024:GEM

[BCS24] S. Blanes, F. Casas, and L. Shaw. Generalized extrapolation methods based on compositions of a basic 2nd-order scheme. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001358.

Burgos:2022:SRF

[BCVV22] C. Burgos, J.-C. Cortés, L. Villafuerte, and R. J. Villanueva. Solving random fractional second-order linear equations via the mean square Laplace transform: Theory and statistical computing. Applied Mathematics and Computation, 418 (??):Article 126846, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009292.

Belachew:2020:HPN

[BD20a] Melisew Tefera Belachew and Nicoletta Del Buono. Hybrid projective nonnegative matrix factorization based on α-divergence and the alternating least squares algorithm. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308173.

Bora:2020:GPM

[BD20b] Aniruddha Bora and Weizhong Dai. Gradient preserved method for solving heat conduction equation with variable coefficients in double layers. Applied Mathematics and Computation, 386(??):Article 125516, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304744.

Bezhaev:2021:QTI

[BD21] Anatoly Yu. Bezhaev and Andrey A. Dobrynin. On quartic transmission irregular graphs. Applied Mathematics and Computation, 399(??):Article 126049, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000977.

Bessy:2020:MWI

[BDH^{+}20] Stéphane Bessy, François Dross, Katarína Hrináková, Martin Knor, and Riste Skrekovski. Maximal Wiener index for graphs with prescribed number of blocks. Applied Mathematics and Computation, 380 (??):Article 125274, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302435.

Baty:2023:RES

[BDH^{+}23] Hubert Baty, Florence Drui, Philippe Helluy, Emmanuel Franck, Christian Klingenberg, and Lukas Thanhäuser. A robust and efficient solver based on kinetic schemes for magnetohydrodynamics (MHD) equations. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007263.

Bresar:2023:UCG

[BDK23] Bostjan Bresar, Tanja Dravec, and Elzbieta Kleszcz. Uniquely colorable graphs up to automorphisms. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001765.

Balk:2021:ECE

[BDKZ21] Bert M. Balk, M. B. M. (René) De Koster, Christian Kaps, and José L. Zofío. An evaluation of crossefficiency methods: With an application to warehouse performance. Applied Mathematics and Computation, 406(??):Article 126261, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003507.

Balbuena:2020:ICL

[BDMB20] C. Balbuena, C. Dalfó, and B. Martínez-Barona. Identifying codes in line digraphs. Applied Mathematics and Computation, 383 (??):Article 125357, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303210.

Bastinec:2022:DID

[BDPV22] J. Bastinec, J. Diblík, S. Pinelas, and J. Vala. Determining the initial data generating solutions with prescribed behaviour of a triangular system of linear discrete equations. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006196.

Busto:2023:ALE

[BDRM23] S. Busto, M. Dumbser, and L. Río-Martín. An arbitrary-Lagrangian–Eulerian hybrid finite volume/finite element method on moving unstructured meshes for the Navier–Stokes equations. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006130.

Brezovnik:2023:SMR

[BDTP23] Simon Brezovnik, Matthias Dehmer, Niko Tratnik, and Petra Zigert Pletersek. Szeged and Mostar rootindices of graphs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008049.

Berrut:2020:PML

[BE20] Jean-Paul Berrut and Giacomo Elefante. A periodic map for linear barycentric rational trigonometric interpolation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309166.

Bordj:2022:MAI

[BE22] Naziha Bordj and Nadjia El Saadi. Moment approximation of individual-based models. application to the study of the spatial dynamics of phytoplankton populations. Applied Mathematics and Computation, 412(??):Article 126594, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006780.

Berraih:2021:SCA

[BEe21] Sid Ahmed Berraih, Sidi Mohammed El Amine Debbal, and Nour elhouda Baakek yettou. Severity cardiac analysis using the higher-order spectra. Applied Mathematics and Computation, 409 (??):Article 126389, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004781.

Barrera:2023:LDS

[BEIR23] D. Barrera, S. Eddargani, M. J. Ibáñez, and S. Remogna. Low-degree spline quasi-interpolants in the Bernstein basis. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003193.

Bebiano:2020:RAT

[BF20] Natália Bebiano and Susana Furtado. Remarks on anti-tridiagonal matrices. Applied Mathematics and Computation, 373 (??):Article 125008, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310008.

Barrera-Figueroa:2022:PSA

[BF22] Víctor Barrera-Figueroa. A power series analysis of bound and resonance states of onedimensional Schrödinger operators with finite point interactions. Applied Mathematics and Computation, 417(??):Article 126774, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008560.

Bueno:2023:SMW

[BF23] M. I. Bueno and Susana Furtado. Singular matrices whose Moore–Penrose inverse is tridiagonal. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003235.

Bandera:2024:CIP

[BFGGM24] Alejandro Bandera, Soledad Fernández-García, and Macarena Gómez-Mármol. On the computation of intrinsic Proper Generalized Decomposition modes of parametric symmetric elliptic problems on Grassmann manifolds. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000511.

Brdar:2023:BNE

[BFLR23] Mirjana Brdar, Sebastian Franz, Lars Ludwig, and Hans-Görg Roos. A balanced norm error estimation for the time-dependent reaction-diffusion problem with shift in space. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005811.

Branquinho:2020:MTV

[BFMGA20] Amílcar Branquinho, Ana Foulquié-Moreno, and Juan C. García-Ardila. Matrix Toda and Volterra lattices. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307143.

Bazarra:2020:PEA

[BFQ20] N. Bazarra, J. R. Fernández, and R. Quintanilla. An a priori error analysis of porothermoviscoelastic problems. Applied Mathematics and Computation, 379 (??):Article 125268, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030237X.

Breckling:2024:NLT

[BFRS24] Sean Breckling, Joseph Fiordilino, Jorge Reyes, and Sidney Shields. A note on the long-time stability of pressure solutions to the 2D Navier Stokes equations. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400300X.

Bertoluzza:2020:WCQ

[BFS20] S. Bertoluzza, S. Falletta, and L. Scuderi. Wavelets and convolution quadrature for the efficient solution of a 2D space-time BIE for the wave equation. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307180.

Beiglo:2020:NSN

[BG20a] H. Beiglo and M. Gachpazan. Numerical solution of nonlinear mixed Volterra–Fredholm integral equations in complex plane via PQWs. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308203.

Beldzinski:2020:NTE

[BG20b] Michał Bełdziński and Marek Galewski. Nash-type equilibria for systems of nonpotential equations. Applied Mathematics and Computation, 385(??):Article 125456, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320304173.

Boos:2024:LMM

[BGB24] Everton Boos, Douglas S. Gonçalves, and Fermín S. V. Bazán. Levenberg– Marquardt method with singular scaling and applications. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001607.

Boyer:2024:DPR

[BGH24] Geoffrey Boyer, Wayne Goddard, and Michael A. Henning. Domination parameters and the removal of matchings. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002224.

Bachstein:2022:CCG

[BGHX22] Anna Bachstein, Wayne Goddard, Michael A. Henning, and John Xue. Compelling colorings: a generalization of the dominator chromatic number. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002673.

Bohaienko:2021:IFW

[BGRM21] Vsevolod Bohaienko, Anatolij Gladky, Mykhailo Romashchenko, and Tetiana Matiash. Identification of fractional water transport model with ψ-Caputo derivatives using particle swarm optimization algorithm. Applied Mathematics and Computation, 390 (??):Article 125665, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306184.

Bru:2022:CMC

[BGS22] Rafael Bru, Maria T. Gassó, and Máximo Santana. Combined matrices and conditioning. Applied Mathematics and Computation, 412 (??):Article 126549, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006330.

Badday:2022:MIF

[BH22] Alaa Jabbar Badday and Akil J. Harfash. Magnetohydrodynamic instability of fluid flow in a porous channel with slip boundary conditions. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004374.

Bengochea:2021:NFP

[BHGPC21] Abimael Bengochea, Antonio Hernández-Garduño, and Ernesto Pérez-Chavela. New families of periodic orbits in the 4-body problem emanating from a kite configuration. Applied Mathematics and Computation, 398(??):Article 125961, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000096.

Bae:2022:CCB

[BHK^{+}22] Hyeong-Ohk Bae, Seung-Yeal Ha, Myeongju Kang, Hyuncheul Lim, Chanho Min, and Jane Yoo. A constrained consensus based optimization algorithm and its application to finance. Applied Mathematics and Computation, 416(??):Article 126726, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008080.

Boutaayamou:2023:OBO

[BHL23] Idriss Boutaayamou, Aissam Hadri, and Amine Laghrib. An optimal bilevel optimization model for the generalized total variation and anisotropic tensor parameters selection. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005847.

Banerjee:2021:PID

[BHP21] S. Banerjee, Michael A.

Henning, and D. Pradhan. Perfect Italian domination in cographs. Applied Mathematics and Computation, 391(??):Article 125703, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306561.

Bermudo:2020:DHC

[BHR20] Sergio Bermudo, Robinson A. Higuita, and Juan Rada. Domination in hexagonal chains. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308094.

Bialkowski:2021:APD

[Bia21] Jerzy Białkowski. An algorithm for the periodicity of deformed preprojective algebras of Dynkin types E6, E7, and E8. Applied Mathematics and Computation, 409(??):Article 126289, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003787.

Bras:2022:CSG

[BIJ22] Michal Braś, Giuseppe Izzo, and Zdzislaw Jackiewicz. Construction of G — or G(ϵ)-symplectic general linear methods. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002788.

Byun:2021:PSC

[BJ21] Jong Hyuk Byun and Il Hyo Jung. Phase-specific cancerimmune model considering acquired resistance to therapeutic agents. Applied Mathematics and Computation, 391(??):Article 125555, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305117.

Bhatti:2022:LGA

[BJK^{+}22] Muhammad Mubashir Bhatti, Shen Jun, Chaudry Masood Khalique, Anwar Shahid, Liu Fasheng, and Mohamed S. Mohamed. Lie group analysis and robust computational approach to examine mass transport process using Jeffrey fluid model. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000224.

Bot:2021:FBD

[BK21] Radu Ioan Bot and Laura Kanzler. A forwardbackward dynamical approach for nonsmooth problems with block structure coupled by a smooth function. Applied Mathematics and Computation, 394 (??):Article 125822, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030775X.

Benitez:2024:CEO

[BKK24] Julio Benítez, Waldemar W. Koczkodaj, and Adam Kowalczyk. Computationally efficient orthogonalization for pairwise comparisons method. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001231.

Bukovsek:2022:EGS

[BKMO22] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Extreme generators of shock induced copulas. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002880.

Bera:2022:DHI

[BKR22] Sovan Bera, Subhas Khajanchi, and Tapan Kumar Roy. Dynamics of an HTLV-I infection model with delayed CTLs immune response. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002806.

Brugnano:2021:PVS

[BKS21] Luigi Brugnano, Dmitry E.

Kvasov, and Yaroslav D. Sergeyev. Preface to the virtual special issue recent developments in applied mathematics and computation. Applied Mathematics and Computation, 409 (??):Article 126023, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000710.

Babadjanova:2020:PSD

[BKU20] Aygul Babadjanova, Thomas Kriecherbauer, and Gayrat Urazboev. The periodic solutions of the discrete modified KdV equation with a self-consistent source. Applied Mathematics and Computation, 376 (??):Article 125136, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301053.

Belevtsov:2022:FAF

[BL22] Nikita S. Belevtsov and Stanislav Yu. Lukashchuk. A fast algorithm for fractional Helmholtz equation with application to electromagnetic waves propagation. Applied Mathematics and Computation, 416(??): Article 126728, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008109.

Bodnar:2022:BPS

[BLNT22] Taras Bodnar, Mathias Lindholm, Vilhelm Niklasson, and Erik Thorsén. Bayesian portfolio selection using VaR and CVaR. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002041.

Bi:2022:TDP

[BLO22] Zhimin Bi, Shutang Liu, and Miao Ouyang. Threedimensional pattern dynamics of a fractional predatorprey model with crossdiffusion and herd behavior. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000418.

Bedo:2023:GCK

[BLOP23] Marcos Bedo, João V. S.

Leite, Rodolfo A. Oliveira, and Fábio Protti. Geodetic convexity and Kneser graphs. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001339.

Breckner:2021:OCR

[BLS21] Brigitte E. Breckner, Hannelore Lisei, and Gheorghe Ionut Simon. Optimal control results for a class of stochastic Schrödinger equations. Applied Mathematics and Computation, 407(??):Article 126310, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003994.

Belmouloud:2020:SCN

[BM20] Imane Belmouloud and Ameur Memou. On the solvability of a class of nonlinear singular parabolic equation with integral boundary condition. Applied Mathematics and Computation, 373 (??):Article 124999, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309919.

Bouchlaghem:2021:UAM

[BM21] Mohammed Bouchlaghem and El Bekkaye Mermri. A Uzawa algorithm with multigrid solver for a bilateral obstacle problem. Applied Mathematics and Computation, 389 (??):Article 125553, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305099.

Ban:2022:CBA

[BM22] Youxue Ban and Changwen Mi. On the competition between adhesive and surface effects in the nanocontact properties of an exponentially graded coating. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004386.

Bhat:2024:CSD

[BM24] Imtiyaz Ahmad Bhat and Lakshmi Narayan Mishra. A comparative study of discretization techniques for augmented Urysohn type nonlinear functional Volterra integral equations and their convergence analysis. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000274.

Bellini:2021:GLA

[BMDE21] Emanuele Bellini, Nadir Murru, Antonio J. Di Scala, and Michele Elia. Group law on affine conics and applications to cryptography. Applied Mathematics and Computation, 409 (??):Article 125537, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304938.

Bujanda:2020:NFS

[BMJ20] B. Bujanda, M. J. Moreta, and J. C. Jorge. New fractional step Runge–Kutta– Nyström methods up to order three. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307350.

Bruni:2023:RTI

[BMMV23] Vittoria Bruni, Silvia Marconi, Giuseppina Monteverde, and Domenico Vitulano. Radon transform of image monotonic rearrangements as feature for noise sensor signature. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003429.

Buterin:2021:ISP

[BMS21] S. A. Buterin, M. A. Malyugina, and C.-T. Shieh. An inverse spectral problem for second-order functionaldifferential pencils with two delays. Applied Mathematics and Computation, 411 (??):Article 126475, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005646.

Botija-Munoz:2024:ENA

[BMY24] Ulises Botija-Munoz and Chenggui Yuan. Explicit numerical approximations for SDDEs in finite and infinite horizons using the adaptive EM method: Strong convergence and almost sure exponential stability. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400314X.

Bermudo:2020:ETR

[BNR20] Sergio Bermudo, Juan E.

Nápoles, and Juan Rada. Extremal trees for the Randić index with given domination number. Applied Mathematics and Computation, 375(??):Article 125122, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300916.

Boonstra:2021:VES

[BO21] Boris C. Boonstra and Cornelis W. Oosterlee. Valuation of electricity storage contracts using the COS method. Applied Mathematics and Computation, 410(??):Article 126416, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005051.

Bocquet:2022:OWA

[Boc22] S. Bocquet. Ocean wave autocorrelation function. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001989.

Bondarenko:2022:FDA

[Bon22] Natalia P. Bondarenko. Finite-difference approximation of the inverse Sturm–Liouville problem with frozen argument. Applied Mathematics and Computation, 413(??):Article 126653, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007372.

Belardo:2024:LPS

[BOT24] Francesco Belardo, Elismar R. Oliveira, and Vilmar Trevisan. Limit points of (signless) Laplacian spectral radii of linear trees. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002807.

Boyina:2023:WPA

[BP23] Kalyan Boyina and Raghu Piska. Wave propagation analysis in viscoelastic Timoshenko nanobeams under surface and magnetic field effects based on nonlocal strain gradient theory. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006543.

Bagomghaleh:2024:NAF

[BPG24] Shadi Malek Bagomghaleh, Saeed Pishbin, and Gholamhossein Gholami. Numerical and analytical findings on the Volterra integralalgebraic index-1 system with vanishing delays. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006185.

Bonettini:2021:NCR

[BPR21] S. Bonettini, M. Prato, and S. Rebegoldi. New convergence results for the inexact variable metric forward– backward method. Applied Mathematics and Computation, 392(??):Article 125719, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030672X.

Benes:2022:NBC

[BPRR22] Michal Benes, Igor Pazanin, Marko Radulović, and Borja Rukavina. Nonzero boundary condition for the unsteady micropolar pipe flow: Well-posedness and asymptotics. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002582.

Baglama:2023:NRO

[BPT23] James Baglama, Vasilije Perović, and Timothy Toolan. Note on a rank-one modification of the singular value decomposition. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003399.

Behera:2020:OMB

[BR20] S. Behera and S. Saha Ray. An operational matrix based scheme for numerical solutions of nonlinear weakly singular partial integro-differential equations. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307635.

Boykov:2024:SSS

[BRB24] I. V. Boykov, V. A. Roudnev, and A. I. Boykova. Stability of solutions of systems of Volterra integral equations. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002005.

Barajas-Ramirez:2021:GSC

[BRFLGH21] Juan Gonzalo Barajas-Ramírez, Arturo Franco-López, and Hugo G. González-Hernández. Generating Shilnikov chaos in 3D piecewise linear systems. Applied Mathematics and Computation, 395(??):Article 125877, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308304.

Brito:2020:DMB

[Bri20] Irene Brito. A decision model based on expected utility, entropy and variance. Applied Mathematics and Computation, 379 (??):Article 125285, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030254X.

Busto:2021:SIH

[BRMVCD21] S. Busto, L. Río-Martín, M. E. Vázquez-Cendón, and M. Dumbser. A semiimplicit hybrid finite volume/finite element scheme for all Mach number flows on staggered unstructured meshes. Applied Mathematics and Computation, 402(??):Article 126117, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100165X.

Basheer:2021:NAA

[BRTR21] Ambreen Basheer, Muhammad Rehan, Muhammad Tufail, and Muhammad Ahsan Razaq. A novel approach for adaptive H_{∞} leader-following consensus of higher-order locally Lipschitz multi-agent systems. Applied Mathematics and Computation, 395(??): Article 125749, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307025.

Bagherpoorfard:2020:MMV

[BS20] Mina Bagherpoorfard and Ali R. Soheili. Moving mesh version of wave propagation algorithm based on augmented Riemann solver. Applied Mathematics and Computation, 375(??): Article 125087, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300564.

Bahmani:2022:NSU

[BS22a] Erfan Bahmani and Ali Shokri. Numerical study of the unsteady 2D coupled magneto-hydrodynamic equations on regular/irregular pipe using direct meshless local Petrov–Galerkin method. Applied Mathematics and Computation, 417 (??):Article 126769, March 15, 2022. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008511.

Batugedara:2022:NEG

[BS22b] Yasasya Batugedara and Kyle J. Schwiebert. Note on the effect of grad–div stabilization on calculating drag and lift coefficients. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005082.

Bekezhanova:2022:ECT

[BS22c] Victoria B. Bekezhanova and Irina V. Stepanova. Evaporation convection in two-layers binary mixtures: equations, structure of solution, study of gravity and thermal diffusion effects on the motion. Applied Mathematics and Computation, 414 (??):Article 126424, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005130.

Bowater:2023:EAP

[BS23] David Bowater and Emmanuel Stefanakis. Extending the Adapted PageRank Algorithm centrality model for urban street networks using non-local random walks. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000577.

Bu:2020:SBS

[BSCL20] Qihui Bu, Yun Sun, Xudong Chai, and Liwei Liu. Strategic behavior and social optimization in a clearing queueing system with N-policy and stochastic restarting scheme. Applied Mathematics and Computation, 381(??):Article 125309, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302757.

Bogoya:2024:FTE

[BSCV24] M. Bogoya, S. Serra-Capizzano, and P. Vassalos. Fast Toeplitz eigenvalue computations, joining interpolation-extrapolation matrix-less algorithms and simple-loop theory: the preconditioned setting. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006525.

Baamonde-Seoane:2021:NSN

[BSdCCGCV21] María A. Baamonde-Seoane, María del Carmen Calvo-Garrido, Michael Coulon, and Carlos Vázquez. Numerical solution of a nonlinear PDE model for pricing renewable energy certificates (RECs). Applied Mathematics and Computation, 404(??):Article 126199, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321002897.

Bharanedhar:2023:ZSI

[BSG23] S. V. Bharanedhar, A. Antony Selvan, and Riya Ghosh. Zeros of self-inversive polynomials with an application to sampling theory. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200621X.

Baysal:2023:ICR

[BSM23] Veli Baysal, Ramazan Solmaz, and Jun Ma. Investigation of chaotic resonance in Type-I and Type-II Morris–Lecar neurons.

Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001091.

Bartwal:2023:SHC

[BSRV23] Naman Bartwal, Shantanu Shahane, Somnath Roy, and Surya Pratap Vanka. Simulation of heat conduction in complex domains of multimaterial composites using a meshless method. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003776.

Blaszczyk:2021:ESF

[BSS21] Tomasz Blaszczyk, Jaroslaw Siedlecki, and HongGuang Sun. An exact solution of fractional Euler– Bernoulli equation for a beam with fixed-supported and fixed-free ends. Applied Mathematics and Computation, 396(??):Article 125932, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308857.

Beltrametti:2020:MPA

[BSST20] M. C. Beltrametti, J. R. Sendra, J. Sendra, and M. Torrente. Moore– Penrose approach in the Hough transform framework. Applied Mathematics and Computation, 375(??): Article 125083, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300527.

Baksalary:2021:FMP

[BT21] Oskar Maria Baksalary and Götz Trenkler. On formulae for the Moore–Penrose inverse of a columnwise partitioned matrix. Applied Mathematics and Computation, 403(??):Article 125913, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308663.

Boscheri:2022:HOS

[BT22] Walter Boscheri and Maurizio Tavelli. High order semi-implicit schemes for viscous compressible flows in 3D. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005318.

Bakibillah:2022:RET

[BTL^{+}22] A. S. M. Bakibillah, Yong Hwa Tan, Junn Yong Loo, Chee Pin Tan, M. A. S. Kamal, and Ziyuan Pu. Robust estimation of traffic density with missing data using an adaptiver extended Kalman filter. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000017.

Brezovnik:2020:RGC

[BTP20] Simon Brezovnik, Niko Tratnik, and Petra Zigert Pletersek. Resonance graphs of catacondensed even ring systems. Applied Mathematics and Computation, 374(??):Article 125064, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300333.

Buterin:2021:USI

[But21] Sergey Buterin. Uniform stability of the inverse spectral problem for a convolution integro-differential operator. Applied Mathematics and Computation, 390 (??):Article 125592, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305476.

Baysal:2021:CSI

[BY21] Veli Baysal and Ergin Yilmaz. Chaotic signal induced delay decay in Hodgkin–Huxley neuron. Applied Mathematics and Computation, 411 (??):Article 126540, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100624X.

Bi:2023:BRC

[BY23a] Yan Bi and Hui Yang. Based on reputation consistent strategy times promotes cooperation in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008864.

Bin:2023:CER

[BY23b] Liu Bin and Wu Yue. Coevolution of reputationbased preference selection and resource allocation with multigame on interdependent networks. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002977.

Bresar:2024:LTM

[BY24] Bostjan Bresar and Ismael G. Yero. Lower (total) mutual-visibility number in graphs. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005805.

Bai:2024:EGE

[BYCX24] Xi Bai, Ye Ye, Tong Chen, and Nenggang Xie. The evolutionary game of emotions considering the influence of reputation. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001814.

Bian:2021:PDR

[BZ21] Fengmiao Bian and Xiaoqun Zhang. A parameterized Douglas–Rachford splitting algorithm for nonconvex optimization. Applied Mathematics and Computation, 410(??):Article 126425, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005142.

Ba:2023:SCD

[BZ23a] Lina Ba and Heping Zhang. Structure connectivity of data center networks. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001911.

Bi:2023:DIC

[BZ23b] Huimin Bi and Xin Zhang. Defective incidence coloring of graphs. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008499.

Bu:2021:FEF

[BZS21] Weiping Bu, Yanmin Zhao, and Chen Shen. Fast and efficient finite difference/finite element method for the two-dimensional multi-term time-space fractional Bloch–Torrey equation. Applied Mathematics and Computation, 398 (??):Article 125985, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000333.

Bao:2021:FTS

[BZZ21] Yuangui Bao, Yijun Zhang, and Baoyong Zhang. Fixedtime synchronization of coupled memristive neural networks via event-triggered control. Applied Mathematics and Computation, 411 (??):Article 126542, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006263.

Chydzinski:2020:RTQ

[CA20] Andrzej Chydzinski and Blazej Adamczyk. Response time of the queue with the dropping function. Applied Mathematics and Computation, 377 (??):Article 125164, July 15, 2020. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301338.

Castillo:2023:RET

[Cas23] K. Castillo. Remark on the eigenvalues of a tridiagonal matrix in biogeography. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006208.

Cavoretto:2022:ALB

[Cav22] Roberto Cavoretto. Adaptive LOOCV-based kernel methods for solving timedependent BVPs. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003022.

Cayli:2020:CMI

[Çay20] Gül Deniz Çayli. Construction methods for idempotent nullnorms on bounded lattices. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307386.

Casella:2021:CIG

[CB21] Francesco Casella and Bernhard Bachmann. On the choice of initial guesses for the Newton–Raphson algorithm. Applied Mathematics and Computation, 398 (??):Article 125991, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000394.

Chen:2022:RFN

[CC22] Qi-Xin Chen and Xiao-Heng Chang. Resilient filter of nonlinear network systems with dynamic eventtriggered mechanism and hybrid cyber attack. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004933.

Cirne:2024:DTM

[CC24] Diego Cirne and Paulo R. A. Campos. A discretetime model of phenotypic evolution. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002467.

Campagna:2020:CEB

[CCC20a] Rosanna Campagna, Costanza Conti, and Salvatore Cuomo. Computational error bounds for Laplace transform inversion based on smoothing splines. Applied Mathematics and Computation, 383(??):Article 125376, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303404.

Cheng:2020:RBP

[CCC20b] Fei Cheng, Tong Chen, and Qiao Chen. Rewards based on public loyalty program promote cooperation in public goods game. Applied Mathematics and Computation, 378 (??):Article 125180, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301491.

Campagna:2023:LAA

[CCC23] Rosanna Campagna, Costanza Conti, and Salvatore Cuomo. A linear algebra approach to HP-splines frequency parameter selection. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004101.

Caballero-Cardenas:2023:IIE

[CCCdLMR23] C. Caballero-Cárdenas, M. J.

Castro, T. Morales de Luna, and M. L. Muñoz-Ruiz. Implicit and implicit-explicit Lagrange-projection finite volume schemes exactly well-balanced for 1D shallow water system. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008529.

Cayama:2021:PMO

[CCdlH21] Jorge Cayama, Carlota M.

Cuesta, and Francisco de la Hoz. A pseudospectral method for the one-dimensional fractional Laplacian on R. Applied Mathematics and Computation, 389(??):Article 125577, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320305336.

Calabro:2023:TDS

[CCdS^{+}23] Francesco Calabrò, Salvatore Cuomo, Daniela di Serafino, Giuseppe Izzo, and Eleonora Messina. Time discretization in the solution of parabolic PDEs with ANNs. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003995.

Cheung:2022:IOT

[CCE^{+}22] Siu Wun Cheung, Eric Chung, Yalchin Efendiev, Wing Tat Leung, and Sai-Mang Pun. Iterative oversampling technique for constraint energy minimizing generalized multiscale finite element method in the mixed formulation. Applied Mathematics and Computation, 415 (??):Article 126622, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007062.

Coelho:2024:ECT

[CCF24] C. Coelho, M. Fernanda P. Costa, and L. L. Ferrás. Enhancing continuous time series modelling with a latent ODE-LSTM approach. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001991.

Cagini:2022:IRO

[CCG^{+}22] C. Cagini, D. Costarelli, R. Gujar, M. Lupidi, G. A. Lutty, M. Seracini, and G. Vinti. Improvement of retinal OCT angiograms by sampling Kantorovich algorithm in the assessment of retinal and choroidal perfusion. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002302.

Chan:2022:IBI

[CCGV^{+}22] Eunice Y. S. Chan, Robert M.

Corless, Laureano González-Vega, J. Rafael Sendra, and Juana Sendra. Inner Bohemian inverses. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000315.

Chen:2024:RSA

[CCH24] Jie Chen, Jinde Cao, and Maobin Hu. A rapid semi-analytical approach for modeling traffic flow on changing road conditions and its application. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400393X.

Chen:2023:STM

[CCHZ23] Guantao Chen, Yuan Chen, Zhiquan Hu, and Shunzhe Zhang. Spanning trees with at most k leaves in 2connected K_{1,r}-free graphs. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000115.

Chen:2024:LMP

[CCL^{+}24] Zhong Chen, Xiaofang Chen, Jinping Liu, Lihui Cen, and Weihua Gui. Learning model predictive control of nonlinear systems with time-varying parameters using Koopman operator. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000493.

Chandrasekaran:2020:OCA

[CCPT20] E. Chandrasekaran, George E.

Chatzarakis, G. Palani, and E. Thandapani. Oscillation criteria for advanced difference equations of second order. Applied Mathematics and Computation, 372(??): Article 124963, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309555.

Cordero-Carrion:2023:NER

[CCSPMV23] Isabel Cordero-Carrión, Samuel Santos-Pérez, and Clara Martínez-Vidallach. Numerical evolution of the resistive relativistic magnetohydrodynamic equations: a minimally implicit Runge–Kutta scheme. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008426.

Chen:2024:ATS

[CCZ24] Xingyu Chen, Jianhuan Cen, and Qingsong Zou. Adaptive trajectories sampling for solving PDEs with deep learning methods. Applied Mathematics and Computation, 481 (??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324003898.

Cavoretto:2020:APM

[CD20a] Roberto Cavoretto and Alessandra De Rossi. Adaptive procedures for meshfree RBF unsymmetric and symmetric collocation methods. Applied Mathematics and Computation, 382 (??):Article 125354, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303180.

Cavoretto:2020:EIR

[CD20b] Roberto Cavoretto and Alessandra De Rossi. Error indicators and refinement strategies for solving Poisson problems through a RBF partition of unity collocation scheme. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308161.

Church:2022:RCP

[CD22] Kevin E. M. Church and Gabriel William Duchesne. Rigorous continuation of periodic solutions for impulsive delay differential equations. Applied Mathematics and Computation, 415 (??):Article 126733, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008158.

Chapwanya:2023:STD

[CD23] Michael Chapwanya and Phindile Dumani. Spatiotemporal dynamics of microbial population under nutrient-limiting conditions. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003077.

Caraballo:2022:SCM

[CDBR^{+}22] L. E. Caraballo, J. M.

Díaz-Báñez, F. Rodríguez, V. Sánchez-Canales, and I. Ventura. Scaling and compressing melodies using geometric similarity measures. Applied Mathematics and Computation, 426(??):

??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002144.

Cheng:2023:CDP

[CDD23] Xiao-Sheng Cheng, Qingying Deng, and Yuanan Diao. Constructions of DNA and polypeptide cages based on plane graphs and odd crossing π -junctions. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008414.

Conte:2021:MMC

[CDDP21] Dajana Conte, Raffaele D’Ambrosio, Maria Pia D’Arienzo, and Beatrice Paternoster. Multivalue mixed collocation methods. Applied Mathematics and Computation, 409 (??):Article 126346, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004355.

Castillo:2024:CPI

[CdFP24] K. Castillo, C. M. da Fonseca, and J. Petronilho. On Chebyshev polynomials and the inertia of certain tridiagonal matrices. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006665.

Chen:2023:FSL

[CDH23] Yuan Chen, Guowei Dai, and Zhiquan Hu. P_{k}- factors in squares and line graphs of trees. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004137.

Cui:2020:SMI

[CDL^{+}20] Ligang Cui, Jie Deng, Rui Liu, Dongyang Xu, Yajun Zhang, and Maozeng Xu. A stochastic multiitem replenishment and delivery problem with leadtime reduction initiatives and the solving methodologies. Applied Mathematics and Computation, 374 (??):Article 125055, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300242.

Cavoretto:2024:NBC

[CDLR24] Roberto Cavoretto, Alessandra De Rossi, Sandro Lancellotti, and Federico Romaniello. Node-bound communities for partition of unity interpolation on graphs. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006719.

Corsaro:2021:SBI

[CDM21] Stefania Corsaro, Valentina De Simone, and Zelda Marino. Split Bregman iteration for multi-period mean variance portfolio optimization. Applied Mathematics and Computation, 392(??): Article 125715, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306688.

Cavoretto:2023:LPU

[CDP23] R. Cavoretto, A. De Rossi, and E. Perracchione. Learning with partition of unitybased kriging estimators. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001078.

Chen:2022:FSS

[CdR22a] Xingyuan Chen and Gonçalo dos Reis. A flexible split-step scheme for solving McKean-Vlasov stochastic differential equations. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002545.

Cruz:2022:AHS

[CDR22b] Roberto Cruz, Frank Duque, and Juan Rada. Anacondensed hexagonal systems. Applied Mathematics and Computation, 418 (??):Article 126798, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008808.

Coclite:2022:CMC

[CdTPP22] A. Coclite, M. D. de Tullio, G. Pascazio, and T. Politi. Characterization of microcapsules deformation in branching channels. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005197.

Cai:2021:FTL

[CDZW21] Yuliang Cai, Jing Dai, Huaguang Zhang, and Yingchun Wang. Fixed-time leader-following/containment consensus of nonlinear multi-agent systems based on event-triggered mechanism. Applied Mathematics and Computation, 396(??): Article 125881, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308341.

Chung:2022:CMM

[CEPZ22] Eric Chung, Yalchin Efendiev, Sai-Mang Pun, and Zecheng Zhang. Computational multiscale method for parabolic wave approximations in heterogeneous media. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001308.

Casas:2022:HOI

[CET22] F. Casas and A. Escorihuela-Tomàs. High order integrators obtained by linear combinations of symmetricconjugate compositions. Applied Mathematics and Computation, 414(??):Article 126700, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007840.

Chiba:2023:RTR

[CF23a] Shuya Chiba and Michitaka Furuya. Ramseytype results for path covers and path partitions. II. Digraphs. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003740.

Corli:2023:CSS

[CF23b] Andrea Corli and Haitao Fan. Corrigendum to “String stability in traffic flows” Applied Mathematics and Computation 443 (2022) Article Number 127775. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000516.

See [CF23c].

Corli:2023:SST

[CF23c] Andrea Corli and Haitao Fan. String stability in traffic flows. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008438.

See corrigendum [CF23b].

Coltraro:2023:RCS

[CFA^{+}23] Franco Coltraro, Josep Fontana, Jaume Amorós, Maria Alberich-Carramiñana, Júlia Borràs, and Carme Torras. A representation of cloth states based on a derivative of the Gauss linking integral. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300334X.

Cacace:2020:SHD

[CFF20] S. Cacace, R. Ferretti, and A. Festa. Stochastic hybrid differential games and match race problems. Applied Mathematics and Computation, 372(??): Article 124966, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309580.

Chen:2020:DHC

[CFL^{+}20] Ye Chen, Suohai Fan, Hong-Jian Lai, Huimin Song, and Murong Xu. Decomposition and r-hued coloring of K_{4}(7)-minor free graphs. Applied Mathematics and Computation, 384 (??):Article 125206, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301752.

Cheng:2021:EEG

[CFLL21] Tao Cheng, Lihua Feng, Weijun Liu, and Lu Lu. On the extensional eigenvalues of graphs. Applied Mathematics and Computation, 408(??):Article 126365, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004549.

Carvalho:2022:NSP

[CFP22] S. Carvalho, J. A. Ferreira, and G. Pena. Nonlinear systems of parabolic IBVP: a stable supersupraconvergent fully discrete piecewise linear FEM. Applied Mathematics and Computation, 419(??):Article 126857, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009401.

Cortellessa:2023:ELS

[CFPM23] Davide Cortellessa, Nicola Ferro, Simona Perotto, and Stefano Micheletti. Enhancing level set-based topology optimization with anisotropic graded meshes. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000723.

Chen:2020:CSP

[CFR^{+}20] Min Chen, Yingying Fan, André Raspaud, Wai Chee Shiu, and Weifan Wang. Choosability with separation of planar graphs without prescribed cycles. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307489.

Campos:2022:DPI

[CFR22] G. C. M. Campos, J. A. Ferreira, and G. Romanazzi.

Density-pressure IBVP: Numerical analysis, simulation and cell dynamics in a colonic crypt. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001230.

Chen:2024:PSM

[CFWL24] Tianrui Chen, Zezhou Fan, Wenhua Wang, and Wenxue Li. Practical stabilization of multi-links highly nonlinear Takagi–Sugeno fuzzy complex networks with Lévy noise based on aperiodically intermittent discrete-time observation control. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002200.

Chen:2020:ASE

[CFWZ20] Siya Chen, Jianwen Feng, Jingyi Wang, and Yi Zhao. Almost sure exponential synchronization of driveresponse stochastic memristive neural networks. Applied Mathematics and Computation, 383(??):Article 125360, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320303246.

Castillo:2020:CSC

[CG20a] Paul Castillo and Sergio Gómez. Conservative super-convergent and hybrid discontinuous Galerkin methods applied to nonlinear Schrödinger equations. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309427.

Costa:2020:SVW

[CG20b] Carina Moreira Costa and Geovani Nunes Grapiglia. A subspace version of the Wang–Yuan Augmented Lagrangian-Trust Region method for equality constrained optimization. Applied Mathematics and Computation, 387(??):Article 124861, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308537.

Castillo:2021:IDS

[CG21] Paul Castillo and Sergio Gómez. An interpolatory directional splittinglocal discontinuous Galerkin method with application to pattern formation in 2D/ 3D. Applied Mathematics and Computation, 397 (??):Article 125984, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000321.

Carlino:2023:WBF

[CG23] Michele Giuliano Carlino and Elena Gaburro. Well balanced finite volume schemes for shallow water equations on manifolds. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007445.

Chakraborty:2020:RDP

[CGB20] Bhaskar Chakraborty, Santu Ghorai, and Nandadulal Bairagi. Reaction-diffusion predator–prey-parasite system and spatiotemporal complexity. Applied Mathematics and Computation, 386(??):Article 125518, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304768.

Callegaro:2023:FQB

[CGG23] Giorgia Callegaro, Alessandro Gnoatto, and Martino Grasselli. A fully quantization-based scheme for FBSDEs. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007251.

Chatzarakis:2021:SOC

[CGJ21] George E. Chatzarakis, Said R. Grace, and Irena Jadlovská. On the sharp oscillation criteria for halflinear second-order differential equations with several delay arguments. Applied Mathematics and Computation, 397(??):Article 125915, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308687.

Ciallella:2023:SBP

[CGLR23] Mirco Ciallella, Elena Gaburro, Marco Lorini, and Mario Ricchiuto. Shifted boundary polynomial corrections for compressible flows: high order on curved domains using linear meshes. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007664.

Cruz-Garcia:2021:MHB

[CGMFSHR21] S. Cruz-García, F. Martínez-Farías, A. S. Santillán-Hernández, and E. Rangel. Mathematical home burglary model with stochastic long crime trips and patrolling: Applied to Mexico City. Applied Mathematics and Computation, 396(??): Article 125865, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308183.

Cruz:2021:SIC

[CGR21] Roberto Cruz, Ivan Gutman, and Juan Rada. Sombor index of chemical graphs. Applied Mathematics and Computation, 399(??):Article 126018, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000667.

Cabrales:2020:NSA

[CGSRG20] Roberto Carlos Cabrales, Juan Vicente Gutiérrez-Santacreu, and José Rafael Rodríguez-Galván. Numerical solution for an aggregation equation with degenerate diffusion. Applied Mathematics and Computation, 377(??):Article 125145, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301144.

Catana:2020:PAP

[CGTU20] J. C. Catana, A. García, J. Tejel, and J. Urrutia. Plane augmentation of plane graphs to meet parity constraints. Applied Mathematics and Computation, 386(??):Article 125513, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304719.

Chen:2020:EHS

[CGW^{+}20] Xiaolong Chen, Kai Gong, Ruijie Wang, Shimin Cai, and Wei Wang. Effects of heterogeneous selfprotection awareness on resource-epidemic coevolution dynamics. Applied Mathematics and Computation, 385(??):Article 125428, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320303891.

Chen:2022:RSB

[CGZ^{+}22] Ruihan Chen, Moxuan Guo, Song Zhu, Yongqiang Qi, Min Wang, and Junhao Hu. Reachable set bounding for linear systems with mixed delays and state constraints. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001692.

Chen:2020:EER

[CH20a] Jia-Qi Chen and Zheng-Da Huang. On the error estimate of the randomized double block Kaczmarz method. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319308999.

Cheng:2020:AMC

[CH20b] Maolin Cheng and Yun Han.

Application of a modified CES production function model based on improved PSO algorithm. Applied Mathematics and Computation, 387(??):Article 125178, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301478.

Cambie:2022:RBV

[CH22a] Stijn Cambie and John Haslegrave. On the relationship between variable Wiener index and variable Szeged index. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003940.

Chen:2022:MIC

[CH22b] Yahong Chen and He Huang. Modeling the impacts of contact tracing on an epidemic with asymptomatic infection. Applied Mathematics and Computation, 416(??):Article 126754, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008365.

Cao:2024:DMC

[CH24a] Tianxiao Cao and Yuda Hu. Dual-mode coupling resonance and dynamic stability of axially moving ferromagnetic thin plate strips in alternating magnetic field. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005775.

Chang:2024:RSE

[CH24b] Xu-Kang Chang and Yong He. Reachable set estimation of delayed Markovian jump neural networks via variables-augmented-based free-weighting-matrices method. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002984.

Chen:2024:WFV

[CH24c] Li Li Chen and Cong Huang. WENO finite volume scheme using subcell strategy on rectangular meshes. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000791.

Chang:2022:FMC

[Cha22]  Shuenn-Yih Chang.  A

family of matrix coefficient formulas for solving ordinary differential equations. Applied Mathematics and Computation, 418 (??):Article 126811, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008948.

Chadjiconstantinidis:2023:SBR

[Cha23] Stathis Chadjiconstantinidis. Some bounds for the renewal function and the variance of the renewal process. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005719.

Chehreghani:2021:SUT

[Che21] Mostafa Haghir Chehreghani. Sublinear update time randomized algorithms for dynamic graph regression. Applied Mathematics and Computation, 410 (??):Article 126434, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005233.

Cheng:2022:TDC

[Che22]  Dongqin Cheng.  Two

disjoint cycles of various lengths in alternating group graph. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004817.

Chen:2023:GSC

[Che23] Xiaohong Chen. General sum-connectivity index of a graph and its line graph. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008475.

Chang:2023:ESN

[CHG23] Xu-Kang Chang, Yong He, and Zhen-Man Gao. Exponential stability of neural networks with a time-varying delay via a cubic function negativedetermination lemma. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006750.

Chen:2022:WIB

[CHH22] Qipin Chen, Wenrui Hao, and Juncai He. A weight initialization based on the linear product structure for neural networks. Applied Mathematics and Computation, 415 (??):Article 126722, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008067.

Chen:2024:NNC

[CHHC24] Zhengquan Chen, Yandong Hou, Ruirui Huang, and Qianshuai Cheng. Neural network compensatorbased robust iterative learning control scheme for mobile robots nonlinear systems with disturbances and uncertain parameters. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000213.

Chiu:2022:ALP

[Chi22]  Wan-Yi Chiu.  Another

look at portfolio optimization with mental accounts. Applied Mathematics and Computation, 419 (??):Article 126851, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009346.

Chen:2023:IDC

[CHJW23] Zhouyang Chen, Zhi Hu, Wei Jin, and Tingzeng Wu. On the ( Q -)index of 2domination critical graphs. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006312.

Chudzik:2021:CGA

[CHK21] Erik Chudzik, Christiane Helzel, and David Kerkmann. The Cartesian Grid Active Flux Method: Linear stability and bound preserving limiting. Applied Mathematics and Computation, 393(??):Article 125501, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304598.

Chen:2024:NEC

[CHL24] Meirun Chen, Michel Habib, and Cheng-Kuan Lin. A novel edge connectivity based on edge partition for hypercube and folded hypercube. Applied Mathematics and Computation, 470(??):??, June 1, 2024.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324000304.

Chica:2023:RPA

[CHP23] Manuel Chica, Juan M. Hernández, and Matjaz Perc. Rewarding policies in an asymmetric game for sustainable tourism. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003521.

Coll:2020:MPO

[CHST20] C. Coll, A. Herrero, E. Sánchez, and N. Thome. On the minus partial order in control systems. Applied Mathematics and Computation, 386(??):Article 125529, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304859.

Chen:2024:PBF

[CHW^{+}24] Weilu Chen, Jun Hu, Zhihui Wu, Xiaojian Yi, and Hongjian Liu. Protocolbased fault detection for state-saturated systems with sensor nonlinearities and redundant channels. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001905.

Chen:2022:EAD

[CHY22a] Yaoyao Chen, Yunqing Huang, and Nianyu Yi. Error analysis of a decoupled, linear and stable finite element method for Cahn–Hilliard–Navier– Stokes equations. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000145.

Chiu:2022:EUB

[CHY22b] Sunben Chiu, Yuqing He, and Pingzhi Yuan. The exact upper bound for the sum of reciprocals of least common multiples. Applied Mathematics and Computation, 416(??):Article 126756, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008389.

Chydzinski:2022:SDL

[Chy22c] Andrzej Chydzinski. On the structure of data losses induced by an overflowed buffer. Applied Mathematics and Computation, 415 (??):Article 126724, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007876.

Chydzinski:2022:PFS

[Chy22d] Andrzej Chydzinski. Perflow structure of losses in a finite-buffer queue. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002892.

Che:2020:FIO

[CHZ^{+}20] Haochi Che, Jun Huang, Xudong Zhao, Xiang Ma, and Ning Xu. Functional interval observer for discretetime systems with disturbances. Applied Mathematics and Computation, 383 (??):Article 125352, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303167.

Chen:2024:PBB

[CHZ24] Qi-Xing Chen, Guang-Xin Huang, and Ming-Yue Zhang. Preconditioned BiCGSTAB and Bi-CRSTAB methods for solving the Sylvester tensor equation. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006380.

Cvetkovic-Ilic:2021:APM

[CIHP^{+}21] Dragana S. Cvetković-Ilić, Clemens Hofstadler, Jamal Hossein Poor, Jovana Milošević, Clemens G. Raab, and Georg Regensburger. Algebraic proof methods for identities of matrices and operators: Improvements of Hartwig’s triple reverse order law. Applied Mathematics and Computation, 409 (??):Article 126357, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100446X.

Chen:2020:DCA

[CJ20] Xuzhou Chen and Jun Ji.

A divide-and-conquer approach for the computation of the Moore–Penrose inverses. Applied Mathematics and Computation, 379(??):Article 125265, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302344.

Chi:2021:FDL

[CJ21] Xiaoqing Chi and Xiaoyun Jiang. Finite difference Laguerre–Legendre spectral method for the two-dimensional generalized Oldroyd-B fluid on a semiinfinite domain. Applied Mathematics and Computation, 402(??):Article 126138, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001867.

Chang:2022:OBF

[CJ22] Xiao-Heng Chang and Xue Jin. Observer-based fuzzy feedback control for nonlinear systems subject to transmission signal quantization. Applied Mathematics and Computation, 414 (??):Article 126657, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007414.

Caro:2020:PCM

[CJC20] Eduardo Caro, Jesús Juan, and Javier Cara. Periodically correlated models for short-term electricity load forecasting. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306344.

Chai:2021:MPM

[CJCL21] Xudong Chai, Tao Jiang, Baoxian Chang, and Liwei Liu. On a multiple priorities matching system with heterogeneous delay sensitive individuals. Applied Mathematics and Computation, 395(??): Article 125873, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308262.

Chamakuri:2020:CMS

[CK20] Nagaiah Chamakuri and Philipp Kügler. A coupled monodomain solver with optimal memory usage for the simulation of cardiac wave propagation. Applied Mathematics and Computation, 378(??):Article 125212, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301818.

Chugainova:2021:SWI

[CK21] A. P. Chugainova and A. G. Kulikovskii. Shock waves in an incompressible anisotropic elastoplastic medium with hardening and their structures. Applied Mathematics and Computation, 401(??):Article 126077, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001259.

Cho:2022:ASN

[CKL22] Junhyun Cho, Yejin Kim, and Sungchul Lee. An accurate and stable numerical method for option hedge parameters. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003502.

Chakraborty:2021:AMS

[CKN21] Samiran Chakraborty, Kapil Kant, and Gnaneshwar Nelakanti. Approximation methods for system of linear Fredholm integral equations of second kind. Applied Mathematics and Computation, 403 (??):Article 126173, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002630.

Chen:2020:SFS

[CKSP20] Huangyue Chen, Lingchen Kong, Pan Shang, and Shanshan Pan. Safe feature screening rules for the regularized Huber regression. Applied Mathematics and Computation, 386 (??):Article 125500, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304586.

Chernogorova:2021:EFD

[CKV21] Tatiana P. Chernogorova, Miglena N. Koleva, and Lubin G. Vulkov. Exponential finite difference scheme for transport equations with discontinuous coefficients in porous media. Applied Mathematics and Computation, 392(??): Article 125691, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306445.

Cheng:2022:UAE

[CKW22] Ming Cheng, Dimitrios G.

Konstantinides, and Dingcheng Wang. Uniform asymptotic estimates in a timedependent risk model with general investment returns and multivariate regularly varying claims. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005100.

Chertock:2023:WBN

[CKWY23] Alina Chertock, Alexander Kurganov, Tong Wu, and Jun Yan. Well-balanced numerical method for atmospheric flow equations with gravity. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006609.

Chen:2020:PBA

[CL20] Shin-Guang Chen and Yi-Kuei Lin. A permutationand-backtrack approach for reliability evaluation in multistate information networks. Applied Mathematics and Computation, 373 (??):Article 125024, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310161.

Chan:2022:OBF

[CL22] Joseph Chang Lun Chan and Tae H. Lee. Observerbased fault-tolerant control for non-infinitely observable descriptor systems with unknown time-varying state and input delays. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003046.

Chen:2023:UGS

[CL23] Dan Chen and Yang Liu.

Uncertain Gordon–Schaefer model driven by Liu process. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001807.

Chen:2024:CSD

[CL24a] Wei-Wei Chen and Hong-Li Li. Complete synchronization of delayed discrete-time fractional-order competitive neural networks. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003333.

Cheng:2024:AFT

[CL24b] Guifang Cheng and Hao Liu. Asynchronous finitetime extended dissipative filter on linear neutral semi-Markov jump systems under sensor saturation and dual-channel DoS attacks. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001954.

Cheng:2021:MAZ

[CLB21] Kun Cheng, Muhuo Liu, and Francesco Belardo. The minimal augmented Zagreb index of k-apex trees for k ∈ {1, 2, 3}. Applied Mathematics and Computation, 402(??):Article 126139, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001879.

Cuenda:2020:CRT

[CLC20] Sara Cuenda, Marta Llorente, and José A. Capitán. Collapse and recovery times in non-linear harvesting with demographic stochasticity. Applied Mathematics and Computation, 380(??): Article 125236, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302058.

Chen:2022:LFN

[CLC^{+}22] Liping Chen, Xiaomin Li, YangQuan Chen, Ranchao Wu, António M. Lopes, and Suoliang Ge. Leaderfollower non-fragile consensus of delayed fractionalorder nonlinear multi-agent systems. Applied Mathematics and Computation, 414 (??):Article 126688, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007724.

Chen:2024:NRS

[CLC24] Yun Chen, Yaqi Li, and Gang Chen. New results on stability analysis for a class of generalized delayed neural networks. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000018.

Cai:2023:LIC

[CLCD23] Jiansheng Cai, Wenwen Li, Wenjing Cai, and Matthias Dehmer. List injective coloring of planar graphs. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006932.

Che:2023:SOS

[CLCY23] Haitao Che, Kaiping Liu, Haibin Chen, and Hong Yan. Second order selfadaptive dynamical system for sparse signal reconstruction and applications to image recovery. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001881.

Cui:2024:ISP

[CLD^{+}24] Guang-Hai Cui, Jun-Li Li, Kun-Xiang Dong, Xing Jin, Hong-Yong Yang, and Zhen Wang. Influence of subsidy policies against insurances on controlling the propagation of epidemic security risks in networks. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002583.

Chen:2024:LSD

[CLH24a] Mujia Chen, Sihao Liu, and Weihua He. Learn to solve dominating set problem with GNN and reinforcement learning. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001899.

Chu:2024:ESC

[CLH24b] Yao Chu, Xiaodi Li, and Xiuping Han. Exponential synchronization of complex networks with unmeasured coupling delays via impulsive observer and impulsive control. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003047.

Chen:2021:CGT

[CLL21] Xiaodan Chen, Xiuyu Li, and Wenshui Lin. On connected graphs and trees with maximal inverse sum indeg index. Applied Mathematics and Computation, 392(??):Article 125731, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306846.

Chai:2022:ATW

[CLL22] Yingbin Chai, Wei Li, and Zuyuan Liu. Analysis of transient wave propagation dynamics using the enriched finite element method with interpolation cover functions. Applied Mathematics and Computation, 412 (??):Article 126564, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006482.

Chang:2023:IMP

[CLL23] Caibing Chang, Xianfu Li, and Yan Liu. Integer k-matching preclusion of twisted cubes and (n, s)-star graphs. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200710X.

Cheon:2024:SLI

[CLL24] Seunghee Cheon, Soobum Lee, and Jaehun Lee. A system-level interface sampling and reduction method for component mode synthesis with varying parameters. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002431.

Cui:2022:FTS

[CLLA22] Qian Cui, Lulu Li, Jianquan Lu, and Abdulaziz Alofi. Finite-time synchronization of complex dynamical networks under delayed impulsive effects. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003642.

Chen:2023:ASS

[CLLA23] Haodong Chen, Lulu Li, Jianquan Lu, and Sultan M. Alghamdi. Asymptotical set stabilization of large-scale logical networks with stochastic delays and the application in finitefield networks. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002217.

Chiu:2020:VRS

[CLM20] Sung Nok Chiu, Leevan Ling, and Michael McCourt. On variable and random shape Gaussian interpolations. Applied Mathematics and Computation, 377 (??):Article 125159, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301284.

Chen:2024:RDN

[CLSN24] Yong Chen, Jiarui Li, Miaoping Sun, and Fuxi Niu. Robust distributed Nash equilibrium seeking for high-order systems with disturbances and coupling constraints. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002650.

Conti:2022:AOE

[CLUR22] Costanza Conti, Sergio López-Ureña, and Lucia Romani. Annihilation operators for exponential spaces in subdivision. Applied Mathematics and Computation, 418(??):Article 126796, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100878X.

Cao:2021:ECF

[CLW21] Mengyu Cao, Benjian Lv, and Kaishun Wang. On even-cycle-free subgraphs of the doubled Johnson graphs. Applied Mathematics and Computation, 403(??):Article 125763, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320307165.

Chen:2023:ETO

[CLWX23] Xiang Chen, Shi Li, Ronghao Wang, and Zhengrong Xiang. Event-triggered output feedback adaptive control for nonlinear switched interconnected systems with unknown control coefficients. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000231.

Cao:2021:EEC

[CLWZ21] Mengyu Cao, Benjian Lv, Kaishun Wang, and Sanming Zhou. Extremal even-cycle-free subgraphs of the complete transposition graphs. Applied Mathematics and Computation, 405(??):Article 126223, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003131.

Chen:2022:EID

[CLX22] Jing Chen, Erbo Li, and Yushan Xue. Evolution of initial discontinuities in the Riemann problem for the Jaulent–Miodek equation with positive dispersion. Applied Mathematics and Computation, 419(??): Article 126869, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009528.

Chen:2022:FRS

[CLZ^{+}22] Wenbin Chen, Junwei Lu, Guangming Zhuang, Fang Gao, Zhengqiang Zhang, and Shengyuan Xu. Further results on stabilization for neutral singular Markovian jump systems with mixed interval time-varying delays. Applied Mathematics and Computation, 420(??): Article 126884, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100967X.

Chen:2023:SRG

[CLZ23] Ming-Zhu Chen, A-Ming Liu, and Xiao-Dong Zhang. On the A_{α} -spectral radius of graphs without linear forests. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001741.

Cheng:2024:CDA

[CLZ24] Huidong Cheng, Wei Li, and Tonghua Zhang. Complex dynamic analysis of a big fish-small fish system by using the Poincaré map. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004338.

Chen:2021:ELD

[CLZY21] Chuanjun Chen, Xi Li, Jun Zhang, and Xiaofeng Yang. Efficient linear, decoupled, and unconditionally stable scheme for a ternary Cahn–Hilliard type Nakazawa–Ohta phase-field model for tri-block copolymers. Applied Mathematics and Computation, 388 (??):Article 125463, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304239.

Cano:2020:MGM

[CM20a] B. Cano and M. J. Moreta. A modified Gautschi’s method without order reduction when integrating boundary value nonlinear wave problems. Applied Mathematics and Computation, 373 (??):Article 125022, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310148.

Charcon:2020:MAS

[CM20b] D. Y. Charcon and L. H. A. Monteiro. A multi-agent system to predict the outcome of a two-round election. Applied Mathematics and Computation, 386 (??):Article 125481, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304409.

Caldarola:2021:MII

[CM21a] Fabio Caldarola and Mario Maiolo. A mathematical investigation on the invariance problem of some hydraulic indices. Applied Mathematics and Computation, 409(??):Article 125726, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306792.

Correa-Morris:2021:MPS

[CM21b] Jyrko Correa-Morris. The median partition and submodularity. Applied Mathematics and Computation, 410(??):Article 126450, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005397.

Chen:2022:STK

[CM22] Hao Chen and Junjie Ma. Solving the third-kind Volterra integral equation via the boundary value technique: Lagrange polynomial versus fractional interpolation. Applied Mathematics and Computation, 414 (??):Article 126685, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007694.

Cindea:2020:BOC

[CMMN20] Nicolae Cîndea, Andaluzia Matei, Sorin Micu, and Constantin Nita. Boundary optimal control for antiplane contact problems with power-law friction. Applied Mathematics and Computation, 386 (??):Article 125448, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304094.

Cruz:2020:EVV

[CMR20] Roberto Cruz, Juan Monsalve, and Juan Rada. Extremal values of vertexdegree-based topological indices of chemical trees. Applied Mathematics and Computation, 380(??):Article 125281, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302502.

Cao:2020:NRS

[CMT20] Shujuan Cao, Yuede Ma, and Zhenyu Taoqiu. A note on rainbow saturation number of paths. Applied Mathematics and Computation, 378(??):Article 125204, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301739.

Chang:2021:EBS

[CMYF21] Beibei Chang, Xiaowu Mu, Zhe Yang, and Jianyin Fang. Event-based secure consensus of muti-agent systems under asynchronous DoS attacks. Applied Mathematics and Computation, 401(??):Article 126120, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001685.

Chakraborty:2023:SSV

[CN23] Samiran Chakraborty and Gnaneshwar Nelakanti. Superconvergence of system of Volterra integral equations by spectral approximation method. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007330.

Chen:2023:SBR

[CNYC23] Chuin-Shan Chen, Amir Noorizadegan, D. L. Young, and C. S. Chen. On the selection of a better radial basis function and its shape parameter in interpolation problems. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007810.

Chen:2021:MBA

[CNZ^{+}21] Zhongyu Chen, Ben Niu, Xudong Zhao, Liang Zhang, and Ning Xu. Model-based adaptive event-triggered control of nonlinear continuoustime systems. Applied Mathematics and Computation, 408(??):Article 126330, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321004197.

Choi:2023:LED

[COPW23] Jinwon Choi, Suil O., Jooyeon Park, and Zhiwen Wang. Laplacian eigenvalue distribution of a graph with given independence number. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001121.

Chen:2020:NVB

[CP20] Jun Chen and Ju H. Park.

New versions of Bessel-Legendre inequality and their applications to systems with time-varying delay. Applied Mathematics and Computation, 375(??): Article 125060, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300291.

Cui:2020:SMT

[CPLH20] Jingjing Cui, Guohua Peng, Quan Lu, and Zhengge Huang. A special modified Tikhonov regularization matrix for discrete illposed problems. Applied Mathematics and Computation, 377(??):Article 125165, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030134X.

Cascarano:2022:PPG

[CPMS22] Pasquale Cascarano, Elena Loli Piccolomini, Elena Morotti, and Andrea Sebastiani. Plug-and-play gradient-based denoisers applied to CT image enhancement. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000534.

Clain:2023:SSO

[CPPL23] Stéphane Clain, Rui M. S. Pereira, Paulo A. Pereira, and Diogo Lopes. Structural schemes for one dimension stationary equations. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003764.

Caballero:2024:EUS

[CPS24] Josefa Caballero, Lukasz Płociniczak, and Kishin Sadarangani. Existence and uniqueness of solutions in the Lipschitz space of a functional equation and its application to the behavior of the paradise fish. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002595.

Chen:2024:PID

[CQ24] Nan Chen and Di Qi.

A physics-informed datadriven algorithm for ensemble forecast of complex turbulent systems. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006495.

Chang:2021:MVW

[CRDD21] Caibing Chang, Haizhen Ren, Zijian Deng, and Bo Deng. The ρ-moments of vertex-weighted graphs. Applied Mathematics and Computation, 400(??):Article 126070, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001181.

Crokidakis:2022:MIC

[Cro22] Nuno Crokidakis. Modeling the impact of civilian firearm ownership in the evolution of violent crimes. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003307.

Cruz:2021:SIT

[CRS21] Roberto Cruz, Juan Rada, and José M. Sigarreta. Sombor index of trees with at most three branch vertices. Applied Mathematics and Computation, 409 (??):Article 126414, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005038.

Cardoso:2022:ACF

[CRT22] Isolda E. Cardoso, Sabrina D. Roscani, and Domingo A. Tarzia. About the convergence of a family of initial boundary value problems for a fractional diffusion equation of robin type. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004490.

Curado:2022:NCM

[CRTV22] Manuel Curado, Rocio Rodriguez, Leandro Tortosa, and Jose F. Vicent. A new centrality measure in dense networks based on two-way random walk betweenness. Applied Mathematics and Computation, 412(??):Article 126560, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006445.

Castaneda:2023:MSE

[CRTVG^{+}23] Antonio Rafael Selva Castañeda, Erick Eduardo Ramirez-Torres, Luis Eugenio Valdés-García, Hilda María Morandeira-Padrón, Diana Sedal Yanez, Juan I. Montijano, and Luis Enrique Bergues Cabrales. Modified SEIR epidemic model including asymptomatic and hospitalized cases with correct demographic evolution. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002916.

Coll:2020:ESI

[CS20] Carmen Coll and Elena Sánchez. Epidemic spreading by indirect transmission in a compartmental farm. Applied Mathematics and Computation, 386 (??):Article 125473, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030432X.

Cerdeira:2021:CNG

[CS21a] J. Orestes Cerdeira and Pedro C. Silva. A centrality notion for graphs based on Tukey depth. Applied Mathematics and Computation, 409(??):Article 126409, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321004987.

Conte:2021:MMG

[CS21b] Martina Conte and Christina Surulescu. Mathematical modeling of glioma invasion: acid- and vasculature mediated go-or-grow dichotomy and the influence of tissue anisotropy. Applied Mathematics and Computation, 407(??):Article 126305, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003945.

Cai:2024:FSS

[CS24a] Yao-Yuan Cai and Hai-Wei Sun. A fast Strang splitting method with mass conservation for the space-fractional Gross–Pitaevskii equation. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400047X.

Cardoso:2024:HHI

[CS24b] J. L. Cardoso and Enas M.

Shehata. Hermite-Hadamard inequalities for quantum integrals: a unified approach. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005143.

Chen:2024:GDI

[CS24c] Huanyin Chen and Marjan Sheibani. The g-Drazin inverses of anti-triangular block operator matrices. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005374.

Coco:2020:LSM

[CSC20] A. Coco, M. Semplice, and S. Serra Capizzano. A level-set multigrid technique for nonlinear diffusion in the numerical simulation of marble degradation under chemical pollutants. Applied Mathematics and Computation, 386(??):Article 125503, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304616.

Cicone:2024:OTF

[CSCZ24] Antonio Cicone, Stefano Serra-Capizzano, and Haomin Zhou. One or two frequencies? The Iterative Filtering answers. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004915.

Chu:2022:CIC

[CSG^{+}22] Yu-Ming Chu, B. M. Shankaralingappa, B. J. Gireesha, Faris Alzahrani, M. Ijaz Khan, and Sami Ullah Khan. Combined impact of Cattaneo–Christov double diffusion and radiative heat flux on bioconvective flow of Maxwell liquid configured by a stretched nano-material surface. Applied Mathematics and Computation, 419(??): Article 126883, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009668. See

retraction notice [CSG^{+}23].

Chu:2023:RNC

[CSG^{+}23] Yu-Ming Chu, B. M.

Shankaralingappa, B. J. Gireesha, Faris Alzahrani, M. Ijaz Khan, and Sami Ullah Khan. Retraction notice to “Combined impact of Cattaneo-Christov double diffusion and radiative heat flux on bio-convective flow of Maxwell liquid configured by a stretched nanomaterial surface” [Applied Mathematics and Computation 419 (15 April 2022) / 126883]. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001650.

See [CSG^{+}22].

Cicerone:2023:MVC

[CSK23] Serafino Cicerone, Gabriele Di Stefano, and Sandi Klavzar. On the mutual visibility in Cartesian products and triangle-free graphs. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006920.

Chen:2024:VGI

[CSL24] Wu-Hua Chen, Hao Sun, and Xiaomei Lu. A variable gain impulsive observer for perturbed Lipschitz nonlinear systems with delayed discrete measurements. Applied Mathematics and Computation, 473(??): ??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001139.

Cordero:2021:GCA

[CSQT21] Alicia Cordero, Pablo Soto-Quiros, and Juan R. Torregrosa. A general class of arbitrary order iterative methods for computing generalized inverses. Applied Mathematics and Computation, 409(??):Article 126381, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321004707.

Cai:2022:VDS

[CSS^{+}22] Xiao Cai, Kaibo Shi, Kun She, Shouming Zhong, Ohmin Kwon, and Yiqian Tang. Voluntary defense strategy and quantized sample-data control for T-S fuzzy networked control systems with stochastic cyberattacks and its application. Applied Mathematics and Computation, 423(??): ??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000613.

Chen:2024:OVB

[CSS^{+}24] Xin Chen, Yifu Song, Yu Shao, Jian Wang, Liu He, and Yuefen Chen. Optimistic value-based optimal control problems with uncertain discrete-time noncausal systems. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004708.

Costarelli:2020:CBS

[CSV20] Danilo Costarelli, Marco Seracini, and Gianluca Vinti. A comparison between the sampling Kantorovich algorithm for digital image processing with some interpolation and quasi-interpolation methods. Applied Mathematics and Computation, 374 (??):Article 125046, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300151.

Chen:2020:GEP

[CSW20] Shijian Chen, Pengjian Shang, and Yue Wu. Generalized entropy plane based on large deviations theory for financial time series. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307118.

Chhatria:2024:LBR

[CSZ24] Balakrishna Chhatria, T. Raja Sekhar, and Dia Zeidan. Limiting behaviour of the Riemann solution to a macroscopic production model with van der Waals equation of state. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005738.

Chou:2021:NMO

[CT21] Jui-Sheng Chou and Dinh-Nhat Truong. A novel metaheuristic optimizer inspired by behavior of jellyfish in ocean. Applied Mathematics and Computation, 389 (??):Article 125535, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304914.

Carlini:2024:SGL

[CT24] Elisabetta Carlini and Silvia Tozza. A scheme for the game p-Laplacian and its application to image inpainting. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300468X.

Corradi:2024:HOA

[CTBB24] Eloina Corradi, Maurizio Tavelli, Marie-Laure Baudet, and Walter Boscheri. A high order accurate space-time trajectory reconstruction technique for quantitative particle trafficking analysis. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003631.

Chen:2024:MMB

[CTF24] Boxun Chen, Ze Tang, and Jianwen Feng. Matrix measure-based eventtriggered consensus of multiagent systems with hybrid time delays. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005313.

Curado:2021:IMP

[CTV21] Manuel Curado, Leandro Tortosa, and Jose F. Vicent. Identifying mobility patterns by means of centrality algorithms in multiplex networks. Applied Mathematics and Computation, 406(??):Article 126269, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003581.

Chen:2024:DSI

[CTWG24] Wei Chen, Boyu Tao, Sheng Wang, and Lin Geng. Diverse selection intensities resolve the cooperation dilemma induced by breaking the symmetry between interaction and learning. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400420X.

Chen:2021:GDD

[CTZ21] Wei Chen, Zhidong Teng, and Long Zhang. Global dynamics for a drug-sensitive and drug-resistant mixed strains of HIV infection model with saturated incidence and distributed delays. Applied Mathematics and Computation, 406 (??):Article 126284, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003738.

Cengizci:2023:SFF

[CU23] Süleyman Cengizci and Ömür Ugur. A stabilized FEM formulation with discontinuity-capturing for solving burgers’-type equations at high Reynolds numbers. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007731.

Chu:2022:NIV

[CUA^{+}22] Yu-Ming Chu, Saif Ullah, Muzaher Ali, Ghulam Fatima Tuzzahrah, and Taj Munir. Numerical investigation of Volterra integral equations of second kind using optimal homotopy asymptotic method. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003782.

Christara:2022:PPL

[CW22] Christina C. Christara and Ruining Wu. Penalty and penalty-like methods for nonlinear HJB PDEs. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001011.

Chen:2023:LAC

[CW23] Jijuan Chen and Tao Wang. On linear 2arboricity of certain graphs. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007603.

Chen:2022:FTS

[CWB22] Yuan Chen, Jianwei Wu, and Haibo Bao. Finitetime stabilization for delayed quaternion-valued coupled neural networks with saturated impulse. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001679.

Chen:2020:SPI

[CWC20] Mengxin Chen, Ranchao Wu, and Liping Chen. Spatiotemporal patterns induced by Turing and Turing-Hopf bifurcations in a predator–prey system. Applied Mathematics and Computation, 380(??): Article 125300, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302617.

Chen:2021:EGW

[CWF^{+}21] Zheng Chen, Yong-Ping Wu, Guo-Lin Feng, Zhong-Hua Qian, and Gui-Quan Sun. Effects of global warming on pattern dynamics of vegetation: Wuwei in China as a case. Applied Mathematics and Computation, 390(??):Article 125666, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306196.

Chang:2024:NLA

[CWH24] Xiao-Heng Chang, Xiao-Yan Wang, and Li-Wei Hou. New LMI approach to H_{∞} control of discrete-time singular systems. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001759.

Cui:2021:FLF

[CWJF21] Jiahao Cui, Ruihua Wang, Ticao Jiao, and Shumin Fei. A foreseeable Lyapunov function approach for H_{∞} asynchronous filtering of discrete-time switched systems. Applied Mathematics and Computation, 395(??): Article 125882, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308353.

Cui:2021:IPD

[CWL^{+}21] Guang-Hai Cui, Zhen Wang, Jun-Li Li, Xing Jin, and Zhi-Wang Zhang. Influence of precaution and dynamic post-indemnity based insurance policy on controlling the propagation of epidemic security risks in networks. Applied Mathematics and Computation, 392(??): Article 125720, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306731.

Cheng:2024:FPT

[CWWZ24] Tingting Cheng, Leimin Wang, Zhouchao Wei, and Guodong Zhang. Fixed/ preassigned-time stabilization of discontinuous switched systems with time-varying delays. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002315.

Chen:2020:SLN

[CWY^{+}20] Xiao-Ping Chen, Wei Wei, Xi Yang, Hao Liu, and Xiao-Ming Pan. Successive linear Newton interpolation methods for solving the large-scale nonlinear eigenvalue problems. Applied Mathematics and Computation, 387(??):Article 124663, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306551.

Chen:2021:EEE

[CWY^{+}21] Wei Chen, Jianwei Wang, Fengyuan Yu, Jialu He, Wenshu Xu, and Rong Wang. Effects of emotion on the evolution of cooperation in a spatial prisoner’s dilemma game. Applied Mathematics and Computation, 411 (??):Article 126497, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005865.

Chen:2024:SIP

[CWY^{+}24a] Wei Chen, Jianwei Wang, Fengyuan Yu, Jialu He, Wenshu Xu, and Wenhui Dai. Successful initial positioning of non-cooperative individuals in cooperative populations effectively hinders cooperation prosperity. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004903.

Chen:2024:HIR

[CWY^{+}24b] Wei Chen, Jianwei Wang, Fengyuan Yu, Wenshu Xu, and Wenhui Dai. Heterogeneous interaction radius based on emotional dynamics can promote cooperation in spatial public goods games. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001115.

Chen:2020:MSE

[CWZ20] Huiling Chen, Mingjing Wang, and Xuehua Zhao. A multi-strategy enhanced sine cosine algorithm for global optimization and constrained practical engineering problems. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308641.

Chen:2021:VET

[CWZ21] Ya-Hong Chen, Hua Wang, and Xiao-Dong Zhang. Variations of the eccentricity and their properties in trees. Applied Mathematics and Computation, 405(??):

Article 126258, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003477.

Chen:2022:NCB

[CWZ22] Pingge Chen, Biao Wu, and Qianju Zhang. A note on a conjecture of Bene Watts– Norin–Yepremyan for Lagrangian. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322002296.

Chang:2023:PTF

[CWZL23] Shaoping Chang, Yijing Wang, Zhiqiang Zuo, and Xiaoyuan Luo. Prescribedtime formation control for wheeled mobile robots with time-varying supertwisting extended state observer. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003582.

Chai:2022:WGM

[CWZZ22] Shimin Chai, Yu Wang, Wenju Zhao, and Yongkui Zou. A C^{0} weak Galerkin method for linear Cahn– Hilliard–Cook equation with random initial condition. Applied Mathematics and Computation, 414(??):Article 126659, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007438.

Chadjiconstantinidis:2022:RBT

[CX22] Stathis Chadjiconstantinidis and Panos Xenos. Refinements of bounds for tails of compound distributions and ruin probabilities. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000340.

Cui:2022:DET

[CXGL22] Lili Cui, Xiangpeng Xie, Hongyan Guo, and Yanhong Luo. Dynamic eventtriggered distributed guaranteed cost FTC scheme for nonlinear interconnected systems via ADP approach. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001667.

Chang:2021:SDI

[CXLM21] Zhengbo Chang, Xiaoyan Xing, Siyu Liu, and Xinzhu Meng. Spatiotemporal dynamics for an impulsive eco-epidemiological system driven by canine distemper virus. Applied Mathematics and Computation, 402(??):Article 126135, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001831.

Cui:2023:FTD

[CXY^{+}23] Guozeng Cui, Hui Xu, Jinpeng Yu, Jiali Ma, and Ze Li. Fixed-time distributed adaptive attitude control for multiple QUAVs with quantized input. Applied Mathematics and Computation, 449(??): ??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001029.

Chen:2023:DAM

[CXYZ23] Yonghui Chen, Yu Xue, Xiaona Yang, and Xian Zhang. A direct analysis method to Lagrangian global exponential stability for quaternion memristive neural networks with mixed delays. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007056.

Chen:2021:GSC

[CY21a] Hao Chen and Yeru Yang.

Generalized Störmer–Cowell methods with efficient iterative solver for large-scale second-order stiff semilinear systems. Applied Mathematics and Computation, 400 (??):Article 126062, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001107.

Chen:2021:RDV

[CY21b] Wuxian Chen and Weigen Yan. Resistance distances in vertex-weighted complete multipartite graphs. Applied Mathematics and Computation, 409(??):Article 126382, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004719.

Chen:2024:TPF

[CY24] Fan Chen and Xiying Yuan.

Turán problem for K_{4}^{−}- free signed graphs. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002753.

Chen:2020:NAS

[CYG20] Rubing Chen, Jinjiang Yuan, and Zhichao Geng. ND-agent scheduling of linear-deteriorating tasks with positional due indices to minimize total completion time and maximum cost. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306897.

Chen:2022:AQS

[CYHR22] Wei Chen, Yongguang Yu, Xudong Hai, and Guojian Ren. Adaptive quasisynchronization control of heterogeneous fractionalorder coupled neural networks with reaction–diffusion. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200220X.

Chen:2022:FDN

[CYL^{+}22] Jiaxi Chen, Zehua Yu, Sanyang Liu, Junmin Li, and Jin Xie. Fully distributed neural control of periodically time-varying parameterized stochastic nonlinear multi-agent systems with hybrid-order dynamics. Applied Mathematics and Computation, 426 (??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322002016.

Chen:2021:ECD

[CYW21] Wei Chen, Zhihu Yang, and Te Wu. Evolution of cooperation driven by collective interdependence on multilayer networks. Applied Mathematics and Computation, 388(??):Article 125532, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304884.

Chen:2023:MGM

[CYZ23] Wang Chen, Xinmin Yang, and Yong Zhao. Memory gradient method for multiobjective optimization. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008591.

Cheng:2020:NLF

[CZ20] Jun Cheng and Yang Zhan. Nonstationary l_{2}–l_{∞} filtering for Markov switching repeated scalar nonlinear systems with randomly occurring nonlinearities. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307064.

Cao:2021:SAN

[CZ21a] Wenping Cao and Quanxin Zhu. Stability analysis of neutral stochastic delay differential equations via the vector Lyapunov function method. Applied Mathematics and Computation, 405(??):Article 126257, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003465.

Cui:2021:CTH

[CZ21b] Li Cui and Jin-Xin Zhou. A classification of tetravalent half-arc-transitive metacirculants of 2-power orders. Applied Mathematics and Computation, 392(??): Article 125755, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307086.

Cao:2022:SSN

[CZ22a] Wenping Cao and Quanxin Zhu. Stability of stochastic nonlinear delay systems with delayed impulses. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000364.

Chang:2022:EBD

[CZ22b] Beibei Chang and Chuanxi Zhu. Edge-based dynamic event-triggered mean square consensus control for stochastic multi-agent systems under weight-balanced digraph. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002843.

Chen:2022:USO

[CZ22c] Chuanjun Chen and Tong Zhang. Unconditional stability and optimal error estimates of first order semi-implicit stabilized finite element method for two phase magnetohydrodynamic diffuse interface model. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003125.

Chen:2024:EUG

[CZ24a] Meng Chen and Yan Zhu. Extremal unicyclic graphs of Sombor index. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300543X.

Chen:2024:RSM

[CZ24b] Yi Chen and Juan Zhou. Robust H_{∞} sliding mode control for delay-dependent uncertain T –S fuzzy descriptor stochastic Markovian jump systems with mode-dependent time-varying delays. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003424.

Cui:2022:NNB

[CZGX22] Di Cui, Wencheng Zou, Jian Guo, and Zhengrong Xiang. Neural network-based adaptive finite-time tracking control of switched nonlinear systems with timevarying delay. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002909.

Chen:2022:OBH

[CZK^{+}22] Zongjie Chen, Yigang Zhang, Qingkai Kong, Ting Fang, and Jing Wang. Observerbased H_{∞} control for persistent dwell-time switched networked nonlinear systems under packet dropout. Applied Mathematics and Computation, 415(??):Article 126679, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007633.

Chen:2020:HOC

[CZL20] Changkai Chen, Xiaohua Zhang, and Zhang Liu. A high-order compact finite difference scheme and precise integration method based on modified Hopf– Cole transformation for numerical simulation of ndimensional Burgers’ system. Applied Mathematics and Computation, 372(??): Article 125009, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031931001X.

Cai:2020:DBF

[CZLH20] Yuliang Cai, Huaguang Zhang, Yang Liu, and Qiang He. Distributed bipartite finite-time event-triggered output consensus for heterogeneous linear multiagent systems under directed signed communication topology. Applied Mathematics and Computation, 378(??):Article 125162, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301314.

Cao:2024:PFR

[CZLW24] Yan Cao, Wei-Jie Zhou, Xiao-Zhen Liu, and Kai-Ning Wu. Passivity of fractional reaction–diffusion systems. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002327.

Chen:2021:FKM

[CZMX21] Li Chen, Shuisheng Zhou, Jiajun Ma, and Mingliang Xu. Fast kernel k-means clustering using incomplete Cholesky factorization. Applied Mathematics and Computation, 402 (??):Article 126037, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000850.

Cveticanin:2021:NLT

[CZR21] Stevan M. Cvetićanin, Dusan Zorica, and Milan R. Rapaić. Non-local telegrapher’s equation as a transmission line model. Applied Mathematics and Computation, 390(??):Article 125602, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305579.

Chen:2021:OCU

[CZS21] Xin Chen, Yuanguo Zhu, and Linxue Sheng. Optimal control for uncertain stochastic dynamic systems with jump and application to an advertising model. Applied Mathematics and Computation, 407 (??):Article 126337, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004264.

Chen:2023:UPE

[CZW23] Wei Chen, Qianlong Zhu, and Te Wu. Unfairness promotes the evolution of cooperation. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200652X.

Cheng:2024:DOC

[CZW24] Ling Cheng, Sirui Zhang, and Yingchun Wang. Distributed optimal capacity allocation of integrated energy system via modified ADMM. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005386.

Cai:2020:RCU

[CZWS20] Xiao Cai, Shouming Zhong, Jun Wang, and Kaibo Shi. Robust H_{∞} control for uncertain delayed T –S fuzzy systems with stochastic packet dropouts. Applied Mathematics and Computation, 385(??):Article 125432, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303933.

Cui:2021:ETD

[CZWX21] Lili Cui, Yong Zhang, Xiaowei Wang, and Xiangpeng Xie. Event-triggered distributed self-learning robust tracking control for uncertain nonlinear interconnected systems. Applied Mathematics and Computation, 395(??):Article 125871, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308249.

Cheng:2024:NFT

[CZXyS24] Mengqing Cheng, Junsheng Zhao, Xiangpeng Xie, and Zong yao Sun. A novel finite-time stability criteria and controller design for nonlinear impulsive systems. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003370.

Chen:2024:TTC

[CZZP24] Jiqing Chen, Haiyan Zhang, Tongtong Zhu, and Shangtao Pan. Trajectory tracking control of a manipulator based on an immune algorithm-optimized neural network in the presence of unknown backlash-like hysteresis. Applied Mathematics and Computation, 470 (??):??, June 1, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324000249.

deAmezaga:2022:CPB

[dAGSFM22] Claudio Sánchez Pérez de Amézaga, Víctor M. García-Suárez, and Juan L. Fernández-Martínez. Classification and prediction of bulk densities of states and chemical attributes with machine learning techniques. Applied Mathematics and Computation, 412 (??):Article 126587, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006718.

Das:2024:OPS

[Das24] Kinkar Chandra Das. Open problems on Sombor index of unicyclic and bicyclic graphs. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032400119X.

Dong:2023:EGO

[DBS^{+}23] Yanni Dong, Hajo Broersma, Changwu Song, Pengfei Wan, and Shenggui Zhang. The effect of graph operations on the degreebased entropy. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006075.

Dumbser:2023:PSI

[DBVCP23] Michael Dumbser, Saray Busto, M. Elena Vázquez-Cendón, and Ilya Peshkov. Preface for the special issue “Hyperbolic PDE in computational physics: Advanced mathematical models and structure-preserving numerics”. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001637.

Dang:2021:FMN

[DC21] Rongqi Dang and Yiming Chen. Fractional modelling and numerical simulations of variable-section viscoelastic arches. Applied Mathematics and Computation, 409(??):Article 126376, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004653.

Dhakal:2020:CCI

[DCCM20] Sandeep Dhakal, Raymond Chiong, Manuel Chica, and Richard H. Middleton. Climate change induced migration and the evolution of cooperation. Applied Mathematics and Computation, 377(??):Article 125090, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030059X.

Dianavinnarasi:2020:DDS

[DCR^{+}20] J. Dianavinnarasi, Yang Cao, R. Raja, G. Rajchakit, and C. P. Lim. Delaydependent stability criteria of delayed positive systems with uncertain control inputs: Application in mosquito-borne morbidities control. Applied Mathematics and Computation, 382(??):Article 125210, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030179X.

Dimitrienko:2020:MST

[DD20a] Yu. I. Dimitrienko and I. D.

Dimitrienko. Modeling and simulation of textile composite plates under a punching load. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306526.

Duldul:2020:DSH

[DD20b] Bahar Uyar Düldül and Mustafa Düldül. Derivation of slant helices on a given surface through new approaches. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307349.

Deutsch:2021:PAS

[DD21a] Yael Deutsch and Israel David. Promoting assets selling through advertisement channels. Applied Mathematics and Computation, 395(??):Article 125836, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030789X.

DiNardo:2021:CAF

[DD21b] Elvira Di Nardo and Giuseppe D’Onofrio. A cumulant approach for the firstpassage-time problem of the Feller square-root process. Applied Mathematics and Computation, 391 (??):Article 125707, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306603.

Dimitrov:2021:SCA

[DD21c] Darko Dimitrov and Zhibin Du. A solution of the conjecture about big vertices of minimal-ABC trees. Applied Mathematics and Computation, 397(??):Article 125818, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307712.

DAmbrosio:2024:SBE

[DD24a] Raffaele D’Ambrosio and Stefano Di Giovacchino. Strong backward error analysis of symplectic integrators for stochastic Hamiltonian systems. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006574.

DAmico:2024:DBR

[DD24b] Guglielmo D’Amico and Riccardo De Blasis. Dividend based risk measures: a Markov chain approach. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000833.

Dalmolin:2022:ATE

[DdAS22] D. Dalmolin, F. S. de Azevedo, and E. Sauter. Advances in the theory of existence and numerical simulations for the one-dimensional transport equation. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200265X.

Dou:2022:PAF

[DDC22] Wenhui Dou, Shihong Ding, and Xiangyong Chen. Practical adaptive finite-time stabilization for a class of second-order systems. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004143.

Grosso:2023:WBI

[DDCM23] A. Del Grosso, M. Castro Díaz, C. Chalons, and T. Morales de Luna. On well-balanced implicitexplicit Lagrange-projection schemes for two-layer shallow water equations. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007706.

Deep:2021:APF

[DDE21] Amar Deep, Deepmala, and R. Ezzati. Application of Petryshyn’s fixed point theorem to solvability for functional integral equations. Applied Mathematics and Computation, 395(??): Article 125878, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308316.

Duran-Delfin:2024:MPB

[DDGBGS^{+}24] J. E. Durán-Delfín, C. D.

García-Beltrán, M. E. Guerrero-Sánchez, G. Valencia-Palomo, and O. Hernández-González. Modeling and passivity-based control for a convertible fixed-wing VTOL. Applied Mathematics and Computation, 461 (??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323004678.

DellAccio:2023:ESS

[DDGN23a] Francesco Dell’Accio, Filomena Di Tommaso, Allal Guessab, and Federico Nudo. Enrichment strategies for the simplicial linear finite elements. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001923.

DellAccio:2023:GCE

[DDGN23b] Francesco Dell’Accio, Filomena Di Tommaso, Allal Guessab, and Federico Nudo. A general class of enriched methods for the simplicial linear finite elements. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003181.

Diop:2022:RIE

[DDH22] Amadou Diop, Moustapha Dieye, and Bipan Hazarika. Random integrodifferential equations of Volterra type with delay: attractiveness and stability. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003757.

Doan:2024:CFC

[DDH^{+}24] Trung Duy Doan, Thi Thanh Chau Do, Pham Hoang Ha, Ngoc Diep Pham, and Ingo Schiermeyer. The conflict-free connection number and the minimum degree-sum of graphs. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004429.

DAmbrosio:2024:MBI

[DdlCS24] Raffaele D’Ambrosio, Hugo de la Cruz, and Carmela Scalone. A Magnus-based integrator for Brownian parametric semi-linear oscillators. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000821.

Dai:2021:MLE

[DDLW21] Min Dai, Jinqiao Duan, Junjun Liao, and Xiangjun Wang. Maximum likelihood estimation of stochastic differential equations with random effects driven by fractional Brownian motion. Applied Mathematics and Computation, 397 (??):Article 125927, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308808.

Nardo:2024:OGB

[DDM24] Elvira Di Nardo, Giuseppe D’Onofrio, and Tommaso Martini. Orthogonal gammabased expansion for the CIR’s first passage time distribution. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003722.

Mejia-de-Dios:2022:PFS

[dDMMTH22] Jesús-Adolfo Mejía de Dios, Efrén Mezura-Montes, and Porfirio Toledo-Hernández. Pseudo-feasible solutions in evolutionary bilevel optimization: Test problems and performance assessment. Applied Mathematics and Computation, 412 (??):Article 126577, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006615.

DellAccio:2020:RHI

[DDNS20] Francesco Dell’Accio, Filomena Di Tommaso, Otheman Nouisser, and Najoua Siar. Rational Hermite interpolation on six-tuples and scattered data. Applied Mathematics and Computation, 386(??):Article 125452, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304136.

Dailly:2022:NSD

[DDPS22] Antoine Dailly, Éric Duchêne, Aline Parreau, and Elzbieta Sidorowicz. The neighbour sum distinguishing relaxed edge colouring. Applied Mathematics and Computation, 419(??): Article 126864, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009474.

Deep:2021:NMS

[DDR21] Amar Deep, Deepmala, and Mohsen Rabbani. A numerical method for solvability of some non-linear functional integral equations. Applied Mathematics and Computation, 402 (??):Article 125637, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305919.

Doan:2024:GLR

[DDS24] Trung Duy Doan, Thi Thanh Chau Do, and Ingo Schiermeyer. Graphs with large (1,2)-rainbow connection numbers. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003473.

DeMartino:2022:CST

[De 22] Antonino De Martino. On the Clifford shorttime Fourier transform and its properties. Applied Mathematics and Computation, 418(??):Article 126812, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100895X.

Deressa:2021:PBS

[DE21] Chernet Tuge Deressa and Dinka Tilahun Etefa. Parameter based stability analysis of generalized mathematical model with delay of competition between two species. Applied Mathematics and Computation, 394(??):Article 125791, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030744X.

Dahy:2024:TFJ

[DEHFF24] Sayed A. Dahy, H. M. El-Hawary, Alaa Fahim, and Amal A. Farhat. Tempered fractional Jacobi–Müntz basis for image reconstruction application and highorder pseudospectral tempered fractional differential matrices. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004156.

Dehmer:2020:RBZ

[DESM^{+}20] Matthias Dehmer, Frank Emmert-Streib, Abbe Mowshowitz, Aleksandar Ilić, Zengqiang Chen, Guihai Yu, Lihua Feng, Modjtaba Ghorbani, Kurt Varmuza, and Jin Tao. Relations and bounds for the zeros of graph polynomials using vertex orbits. Applied Mathematics and Computation, 380(??): Article 125239, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302083.

Dehmer:2022:SLE

[DESTP22] Matthias Dehmer, Frank Emmert-Streib, Niko Tratnik, and Petra Zigert Pletersek. Szeged-like entropies of graphs. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200399X.

Buono:2023:BLA

[DESZ23] Nicoletta Del Buono, Flavia Esposito, Laura Selicato, and Rafał Zdunek. Bilevel algorithm for optimizing hyperparameters in penalized nonnegative matrix factorization. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003533.

Draziotis:2020:AFV

[DF20] Konstantinos A. Draziotis and George Fikioris. Asymptotic formulas for Vasyunin cotangent sums. Applied Mathematics and Computation, 383(??):Article 125380, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303441.

deFreitas:2020:OMA

[dFdOFdSBC20] Juliana Campos de Freitas, Helenice de Oliveira Florentino, Antone dos Santos Benedito, and Daniela Renata Cantane. Optimization model applied to radiotherapy planning problem with dose intensity and beam choice. Applied Mathematics and Computation, 387(??):Article 124786, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307787.

Dmytryshyn:2022:DFP

[DFG22] Andrii Dmytryshyn, Massimiliano Fasi, and Mårten Gulliksson. The dynamical functional particle method for multi-term linear matrix equations. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200532X.

Dettlaff:2024:NAB

[DFP^{+}24] Magda Dettlaff, Hanna Furmańczyk, Iztok Peterin, Adriana Roux, and Radosław Ziemann. A new approach to b-coloring of regular graphs. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003758.

Du:2024:SRN

[DFX24] Dingyi Du, Chunhong Fu, and Qingxiang Xu. Some remarks on the norm upper bounds associated with the generalized polar decompositions of matrices. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006392.

Debnath:2020:TRS

[DG20a] Sudip Debnath and Koeli Ghoshal. Transport of reactive species in oscillatory Couette–Poiseuille flows subject to homogeneous and heterogeneous reactions. Applied Mathematics and Computation, 385 (??):Article 125387, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303507.

Dvorak:2020:NCA

[DG20b] Tomás Dvorák and Mei-Mei Gu. Neighbor connectivity of k-ary n-cubes. Applied Mathematics and Computation, 379(??):Article 125237, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030206X.

Dehghan:2021:NAF

[DG21] Mehdi Dehghan and Zeinab Gharibi. Numerical analysis of fully discrete energy stable weak Galerkin finite element scheme for a coupled Cahn–Hilliard– Navier–Stokes phase-field model. Applied Mathematics and Computation, 410(??):Article 126487, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005762.

Das:2022:SIT

[DG22a] Kinkar Chandra Das and Ivan Gutman. On Sombor index of trees. Applied Mathematics and Computation, 412(??):Article 126575, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321006597.

Delitsyn:2022:RSW

[DG22b] Andrey Delitsyn and Denis S. Grebenkov. Resonance scattering in a waveguide with identical thick perforated barriers. Applied Mathematics and Computation, 412 (??):Article 126592, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006767.

Dietze:2023:ROM

[DG23] Saskia Dietze and Martin A. Grepl. Reduced order model predictive control for parametrized parabolic partial differential equations. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002138.

Du:2023:ADE

[DGL^{+}23] Chunpeng Du, Keyu Guo, Yikang Lu, Haoyu Jin, and Lei Shi. Aspiration driven exit-option resolves social dilemmas in the network. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006907.

Dratman:2022:RDM

[DGMP22] Ezequiel Dratman, Luciano N. Grippo, Verónica Moyano, and Adrián Pastine. On the rank of the distance matrix of graphs. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004684.

Dong:2022:ENS

[DGO22] Fengming Dong, Jun Ge, and Zhangdong Ouyang. Express the number of spanning trees in term of degrees. Applied Mathematics and Computation, 415 (??):Article 126697, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007815.

Daude:2023:HCO

[DGP^{+}23] F. Daude, P. Galon, S. Potapov, A. Beccantini, and G. Mianné. A hyperbolic conservative one-velocity onepressure barotropic threecomponent model for fasttransient fluid-structure interaction problems. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000887.

Dhaouadi:2022:HRM

[DGV22] Firas Dhaouadi, Sergey Gavrilyuk, and Jean-Paul Vila. Hyperbolic relaxation models for thin films down an inclined plane. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004520.

Du:2023:ADF

[DH23] Zhongjie Du and Chuanjiang He. Anisotropic diffusion with fuzzy-based source for binarization of degraded document images. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007524.

Du:2024:SEC

[DH24a] Zhibin Du and Yinhao Huang. Some extensions of Collatz (periodic) conjecture. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002145.

Du:2024:NDE

[DH24b] Zhongjie Du and Chuanjiang He. Nonlinear diffusion equation with a dynamic threshold-based source for text binarization. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004144.

Ding:2020:CNR

[DHM^{+}20] Meng Ding, Ting-Zhu Huang, Tian-Hui Ma, Xi-Le Zhao, and Jing-Hua Yang. Cauchy noise removal using groupbased low-rank prior. Applied Mathematics and Computation, 372(??):Article 124971, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300319309634.

Diaz:2023:AOD

[DHM23] Jorge P. Díaz, Sigrid B. Heineken, and Patricia M. Morillas. Approximate oblique dual frames. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001844.

Deeb:2022:PBP

[DHR22] Ahmad Deeb, Aziz Hamdouni, and Dina Razafindralandy. Performance of Borel–Padé–Laplace integrator for the solution of stiff and non-stiff problems. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002028.

Diblik:2023:TPF

[DHS23] Josef Diblík, Hana Halfarová, and Jan Safarík. Two-parameters formulas for general solution to planar weakly delayed linear discrete systems with multiple delays, equivalent nondelayed systems, and conditional stability. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004393.

Dou:2020:SBI

[DHZ^{+}20] Hong-Xia Dou, Ting-Zhu Huang, Xi-Le Zhao, Jie Huang, and Jun Liu. Semiblind image deblurring by a proximal alternating minimization method with convergence guarantees. Applied Mathematics and Computation, 377 (??):Article 125168, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301375.

Ding:2023:MPG

[DHZ23] Zongpeng Ding, Yuanqiu Huang, and Licheng Zhang. Maximal 1plane graphs with dominating vertices. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003752.

DiStefano:2022:MVG

[Di 22] Gabriele Di Stefano. Mutual visibility in graphs. Applied Mathematics and Computation, 419(??):Article 126850, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009334.

Cabral-Garcia:2024:CMF

[dJCGVM24] Gabriela de Jesús Cabral-García and José Villa-Morales. Conditional moments of the first-passage time of a crowed population. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000419.

Deng:2020:ISS

[DJW20] Xinyang Deng, Wen Jiang, and Zhen Wang. An information source selection model based on evolutionary game theory. Applied Mathematics and Computation, 385(??):Article 125362, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030326X.

Deng:2021:CSD

[DJWX21] Shuning Deng, Jinchen Ji, Guilin Wen, and Huidong Xu. A comparative study of the dynamics of a three-disk dynamo system with and without time delay. Applied Mathematics and Computation, 399(??): Article 126016, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000643.

Dehdezi:2022:RPI

[DK22a] Eisa Khosravi Dehdezi and Saeed Karimi. A rapid and powerful iterative method for computing inverses of sparse tensors with applications. Applied Mathematics and Computation, 415 (??):Article 126720, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008043.

Deswal:2022:RTM

[DK22b] Komal Deswal and Devendra Kumar. Rannacher time-marching with orthogonal spline collocation method for retrieving the discontinuous behavior of hedging parameters. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002430.

Dong:2022:LSB

[DK22c] Cheng Dong and Tong Kang. A least squares based diamond scheme for 3D heterogeneous and anisotropic diffusion problems on polyhedral meshes. Applied Mathematics and Computation, 418(??):Article 126847, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009309.

Das:2020:IMA

[DKK20] Dhiraj Kumar Das, Subhas Khajanchi, and T. K. Kar. The impact of the media awareness and optimal strategy on the prevalence of tuberculosis. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307246.

Das:2021:ECT

[DKKJ21] Dhiraj Kumar Das, Anupam Khatua, T. K. Kar, and Soovoojeet Jana. The effectiveness of contact tracing in mitigating COVID-19 outbreak: a model-based analysis in the context of India. Applied Mathematics and Computation, 404(??): Article 126207, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002976.

Dimitrova:2023:GDV

[DKLV23] Dimitrina S. Dimitrova, Vladimir K. Kaishev, Andrea Lattuada, and Richard J. Verrall. Geometrically designed variable knot splines in generalized (non- )linear models. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005677.

Dong:2020:ENS

[DL20a] Jian Dong and Ding Fang Li. An effect non-staggered central scheme based on new hydrostatic reconstruction. Applied Mathematics and Computation, 372(??): Article 124992, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309841.

Du:2020:FTS

[DL20b] Feifei Du and Jun-Guo Lu. Finite-time stability of neutral fractional order time delay systems with Lipschitz nonlinearities. Applied Mathematics and Computation, 375(??): Article 125079, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300485.

Duan:2020:FTD

[DL20c] Ruirui Duan and Junmin Li. Finite-time distributed H_{∞} filtering for Takagi–Sugeno fuzzy system with uncertain probability sensor saturation under switching network topology: Non-PDC approach. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309531.

Deng:2021:EVM

[DL21a] Kecai Deng and Shuchao Li. On the extremal values for the Mostar index of trees with given degree sequence. Applied Mathematics and Computation, 390 (??):Article 125598, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305531.

Dong:2021:STF

[DL21b] Ziming Dong and Hong Li. A space-time finite element method based on local projection stabilization in space and discontinuous Galerkin method in time for convection-diffusion-reaction equations. Applied Mathematics and Computation, 397(??):Article 125937, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308900.

Du:2021:NCF

[DL21c] Feifei Du and Jun-Guo Lu.

New criterion for finitetime synchronization of fractional order memristorbased neural networks with time delay. Applied Mathematics and Computation, 389(??):Article 125616, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305701.

delaCruz:2020:EPS

[dlCJ20] H. de la Cruz and J. C.

Jimenez. Exact pathwise simulation of multidimensional Ornstein–Uhlenbeck processes. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930726X.

Dai:2023:UTM

[DLL^{+}23] Shaoyu Dai, Bowen Li, Jianquan Lu, Jie Zhong, and Yang Liu. A unified transform method for general robust property of probabilistic Boolean control networks. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003065.

Dong:2020:SPQ

[DLM20] Changchang Dong, Juan Liu, and Jixiang Meng. Supereulerian 3-path-quasi-transitive digraphs. Applied Mathematics and Computation, 372(??):Article 124964, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309567.

Dang:2023:UPR

[DLMP23] Pei Dang, Chuang Li, Weixiong Mai, and Wenliang Pan. Uncertainty principles for random signals. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000024.

Ding:2020:ETS

[DLQY20] Weiyang Ding, Jinjie Liu, Liqun Qi, and Hong Yan. Elasticity M-tensors and the strong ellipticity condition. Applied Mathematics and Computation, 373 (??):Article 124982, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309749.

Dong:2023:FTA

[DLS^{+}23] Rumin Dong, Ziwei Li, Hao Shen, Jing Wang, and Lei Su. Finite-time asynchronous H_{∞} control for Markov jump singularly perturbed systems with partially known probabilities. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003624.

Das:2021:SRB

[DM21] Joyentanuj Das and Sumit Mohanty. On the spectral radius of bi-block graphs with given independence number α. Applied Mathematics and Computation, 402(??):Article 125912, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320308651.

Devi:2023:CSF

[DM23] P. Naga Lakshmi Devi and Phani Kumar Meduri. Couple stress fluid past a contaminated fluid sphere with slip condition. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000140.

Das:2024:MSR

[DM24] Joyentanuj Das and Sumit Mohanty. Maximization of the spectral radius of block graphs with a given dissociation number. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005933.

Ding:2021:SOS

[DMD21] Chen Ding, Li Ma, and Shihong Ding. Second-order sliding mode controller design with mismatched term and time-varying output constraint. Applied Mathematics and Computation, 407(??):Article 126331, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004203.

Dong:2021:SOT

[DML21] Changchang Dong, Jixiang Meng, and Juan Liu. Sufficient Ore type condition for a digraph to be supereulerian. Applied Mathematics and Computation, 410(??):Article 126470, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005592.

Dudarev:2020:EMI

[DMNV20] Vladimir V. Dudarev, Roman M. Mnukhin, Rostislav D. Nedin, and Alexander O. Vatulyan. Effect of material inhomogeneity on characteristics of a functionally graded hollow cylinder. Applied Mathematics and Computation, 382 (??):Article 125333, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030299X.

Dostalik:2021:WNA

[DMP21] Mark Dostalík, Ctirad Matyska, and Vít Prusa. Weakly nonlinear analysis of Rayleigh–Bénard convection problem in extended Boussinesq approximation. Applied Mathematics and Computation, 408 (??):Article 126374, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100463X.

DeMarchi:2021:MAF

[DMPP21] S. De Marchi, F. Marchetti, E. Perracchione, and D. Poggiali. Multivariate approximation at fake nodes. Applied Mathematics and Computation, 391(??):Article 125628, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305828.

Daraby:2022:RTT

[DMRR22] B. Daraby, R. Mesiar, F. Rostampour, and A. Rahimi. Related Thunsdorff type and Frank–Pick type inequalities for Sugeno integral. Applied Mathematics and Computation, 414 (??):Article 126683, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007670.

Devi:2020:LOA

[DMSS20] Vinita Devi, Rahul Kumar Maurya, Somveer Singh, and Vineet Kumar Singh. Lagrange’s operational approach for the approximate solution of two-dimensional hyperbolic telegraph equation subject to Dirichlet boundary conditions. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930709X.

Ding:2021:MCA

[DN21] Kailin Ding and Ning Ning. Markov chain approximation and measure change for timeinhomogeneous stochastic processes. Applied Mathematics and Computation, 392(??):Article 125732, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306858.

Donadello:2021:MMV

[DNR21] Carlotta Donadello, Thi Nhu Thao Nguyen, and Ulrich Razafison. On the mathematical modeling of vole populations spatial dynamics via transport equations on a graph. Applied Mathematics and Computation, 396(??):Article 125885, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308389.

DeBonis:2021:QMI

[DO21] Maria Carmela De Bonis and Donatella Occorsio. Quadrature methods for integro-differential equations of Prandtl’s type in weighted spaces of continuous functions. Applied Mathematics and Computation, 393(??):Article 125721, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320306743.

Dolgopolik:2021:ADM

[Dol21] Maksim V. Dolgopolik.

The alternating direction method of multipliers for finding the distance between ellipsoids. Applied Mathematics and Computation, 409(??):Article 126387, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004768.

deOliveira:2020:FRS

[dOS20] Fúlvia S. S. de Oliveira and Fernando O. Souza. Further refinements in stability conditions for timevarying delay systems. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308586.

Souza:2021:FQM

[dOSR21] Matheus de Oliveira Souza and Pablo M. Rodriguez. On a fractional queueing model with catastrophes. Applied Mathematics and Computation, 410 (??):Article 126468, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005579.

Dias:2021:OPP

[DP21] Fabio S. Dias and Gareth W. Peters. Option pricing with polynomial chaos expansion stochastic bridge interpolators and signed path dependence. Applied Mathematics and Computation, 411(??):Article 126484, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005737.

DAmico:2023:RHO

[DP23] Guglielmo D’Amico and Filippo Petroni. ROCOF of higher order for semi-Markov processes. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007871.

Dakic:2024:GNN

[DP24] Jelena Dakić and Marko D. Petković. Gradient neural network model for the system of two linear matrix equations and applications. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003916.

Dong:2021:MVD

[DQC^{+}21] Yanni Dong, Shengning Qiao, Bing Chen, Pengfei Wan, and Shenggui Zhang. Maximum values of degreebased entropies of bipartite graphs. Applied Mathematics and Computation, 401 (??):Article 126094, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001429.

Diegel:2022:CDA

[DR22] Amanda E. Diegel and Leo G. Rebholz. Continuous data assimilation and long-time accuracy in a C^{0} interior penalty method for the Cahn–Hilliard equation. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200128X.

Das:2020:MGA

[DRS20a] Kinkar Chandra Das, José M.

Rodríguez, and José M.

Sigarreta. On the maximal general ABC index of graphs with given maximum degree. Applied Mathematics and Computation, 386(??):Article 125531, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304872.

Diblik:2020:LSS

[DRS20b] J. Diblík, A. Rodkina, and Z. Smarda. On local stability of stochastic delay nonlinear discrete systems with state-dependent noise. Applied Mathematics and Computation, 374 (??):Article 125019, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310112.

dosReis:2023:ISM

[dRST23] Gonçalo dos Reis, Greig Smith, and Peter Tankov. Importance sampling for McKean–Vlasov SDEs. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323002473.

Dobrynin:2020:STI

[DS20] Andrey A. Dobrynin and Reza Sharafdini. Stepwise transmission irregular graphs. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309415.

DAmbrosio:2021:TSR

[DS21a] Raffaele D’Ambrosio and Carmela Scalone. Twostep Runge–Kutta methods for stochastic differential equations. Applied Mathematics and Computation, 403(??):Article 125930, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320308833.

Du:2021:DSB

[DS21b] Kui Du and Xiao-Hui Sun. A doubly stochastic block Gauss–Seidel algorithm for solving linear equations. Applied Mathematics and Computation, 408(??):Article 126373, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321004628.

Duteil:2021:KAC

[DS21c] Nastassia Pouradier Duteil and Francesco Salvarani. Kinetic approach to the collective dynamics of the rock–paper–scissors binary game. Applied Mathematics and Computation, 388 (??):Article 125496, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304549.

Dzemyda:2021:GMS

[DS21d] Gintautas Dzemyda and Martynas Sabaliauskas. Geometric multidimensional scaling: a new approach for data dimensionality reduction. Applied Mathematics and Computation, 409 (??):Article 125561, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305178.

Deka:2022:CHO

[DS22a] Dharmaraj Deka and Shuvam Sen. Compact higher order discretization of 3D generalized convection diffusion equation with variable coefficients in nonuniform grids. Applied Mathematics and Computation, 413(??):Article 126652, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007360.

Dhiman:2022:LWN

[DS22b] Joginder Singh Dhiman and Sumixal Sood. Linear and weakly non-linear stability analysis of oscillatory convection in rotating ferrofluid layer. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003137.

Du:2022:ESD

[DS22c] Jianwei Du and Xiaoling Sun. Extremal symmetric division deg index of molecular trees and molecular graphs with fixed number of pendant vertices. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005124.

Dimitrov:2023:III

[DS23] Darko Dimitrov and Dragan Stevanović. On the σ_{t}-irregularity and the inverse irregularity problem. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007779.

Ding:2024:NPT

[DS24a] Lusong Ding and Weiwei Sun. Neuroadaptive prescribed-time secure control for nonlinear interconnected NCSs via multiple triggering against DoS attacks. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000341.

Dolgov:2024:TPA

[DS24b] Sergey Dolgov and Dmitry Savostyanov. Tensor product approach to modelling epidemics on networks. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004599.

Du:2024:BID

[DS24c] Jianwei Du and Xiaoling Sun. On bond incident degree index of chemical trees with a fixed order and a fixed number of leaves. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005593.

Alves:2021:RWC

[dSAOG21] Charlan Dellon da Silva Alves, Paulo Roberto Oliveira, and Ronaldo Malheiros Gregório. L^{α} Riemannian weighted centers of mass applied to compose an image filter to diffusion tensor imaging. Applied Mathematics and Computation, 390(??):Article 125603, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305580.

Damnjanovic:2024:TIG

[DSAY24] Ivan Damnjanović, Dragan Stevanović, and Salem Al-Yakoob. On transmissionirregular graphs and long pendent paths. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003795.

Lourenco:2020:OSF

[dSLP20] Marcos Antonio de Souza Lourenço and Elie Luis Martínez Padilla. An octree structured finite volume based solver. Applied Mathematics and Computation, 365 (??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319307131.

diSerafino:2020:AII

[dSLV20] Daniela di Serafino, Germana Landi, and Marco Viola. ACQUIRE: an inexact iteratively reweighted norm approach for TV-based Poisson image restoration. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306708.

Dragicevic:2022:IOC

[DSS22] Davor Dragicević, Adina Luminita Sasu, and Bogdan Sasu. Input-output criteria for stability and expansiveness of dynamical systems. Applied Mathematics and Computation, 414 (??):Article 126574, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006585.

diSerafino:2021:UGD

[dSTV21] Daniela di Serafino, Gerardo Toraldo, and Marco Viola. Using gradient directions to get global convergence of Newton-type methods. Applied Mathematics and Computation, 409 (??):Article 125612, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305671.

Ding:2022:EFG

[DSY22] Rui Ding, Quan Shen, and Yuan Yao. The elementfree Galerkin method for the dynamic Signorini contact problems with friction in elastic materials. Applied Mathematics and Computation, 415 (??):Article 126696, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007803.

Dong:2023:NFB

[DSZ^{+}23] Sheng Dong, Zhipeng Shen, Lu Zhou, Haomiao Yu, and Guibing Zhu. Nonlinear feedback-based eventtriggered output-feedback control for marine surface vehicles under deferred output constraints. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002539.

Doslic:2022:SPP

[DTDFT22] Tomislav Doslić, Meysam Taheri-Dehkordi, and Gholam Hossein Fath-Tabar. Shortest perfect pseudomatchings in fullerene graphs. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001126.

Deswal:2024:HOC

[DTK24] Komal Deswal, Komal Taneja, and Devendra Kumar. Higher-order compact finite difference method for a free boundary problem of ductal carcinoma in situ. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006707.

Dong:2022:CAC

[DTL22] W. B. Dong, H. S. Tang, and Y. J. Liu. Convergence analysis on computation of coupled advection-diffusion-reaction problems. Applied Mathematics and Computation, 420(??):Article 126876, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009590.

DellaCroce:2021:PMS

[DTP21] Federico Della Croce, Vincent T’kindt, and Olivier Ploton. Parallel machine scheduling with minimum number of tardy jobs: Approximation and exponential algorithms. Applied Mathematics and Computation, 397(??):Article 125888, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308419.

Ding:2021:SSC

[DTWJ21] Dong Ding, Ze Tang, Yan Wang, and Zhicheng Ji. Secure synchronization of complex networks under deception attacks against vulnerable nodes. Applied Mathematics and Computation, 399(??):Article 126017, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000655.

Du:2024:RRI

[Du24]  Kui Du.  Regularized

randomized iterative algorithms for factorized linear systems. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006379.

DAmbra:2023:EBA

[DVC23] Pasqua D’Ambra, Panayot S. Vassilevski, and Luisa Cutillo. Extending bootstrap AMG for clustering of attributed graphs. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000735.

Derevtsov:2021:GAR

[DVS21] Evgeny Yu. Derevtsov, Yuriy S. Volkov, and Thomas Schuster. Generalized attenuated ray transforms and their integral angular moments. Applied Mathematics and Computation, 409(??):Article 125494, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304525.

Duan:2020:SRC

[DW20] Cunxiang Duan and Ligong Wang. The α-spectral radius of f-connected general hypergraphs. Applied Mathematics and Computation, 382(??):Article 125336, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303027.

Du:2022:EDC

[DW22] Jinming Du and Ziren Wu.

Evolutionary dynamics of cooperation in dynamic networked systems with active striving mechanism. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003691.

Dai:2024:EIC

[DWL^{+}24] Hui Dai, Xiaoyue Wang, Yikang Lu, Yunxiang Hou, and Lei Shi. The effect of intraspecific cooperation in a three-species cyclic predator-prey model. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000468.

Ding:2024:RLI

[DWS24] Hongfei Ding, Yudong Wang, and Hao Shen. A reinforcement learning integral sliding mode control scheme against lumped disturbances in hot strip rolling. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005763.

Deng:2021:DTL

[DWT^{+}21] Ziwei Deng, Zhuoyue Wang, Zhaohui Tang, Keke Huang, and Hongqiu Zhu. A deep transfer learning method based on stacked autoencoder for cross-domain fault diagnosis. Applied Mathematics and Computation, 408(??):Article 126318, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004070.

Duan:2024:CPU

[DWW24] Cunxiang Duan, Ligong Wang, and Yulong Wei. The characteristic polynomials of uniform hypercycles with length four. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002820.

Dong:2020:NMB

[DWZ20] Zeyu Dong, Xin Wang, and Xian Zhang. A nonsingular M-matrix-based global exponential stability analysis of higherorder delayed discretetime Cohen–Grossberg neural networks. Applied Mathematics and Computation, 385(??):Article 125401, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303635.

Du:2022:ETB

[DWZZ22] Yingxue Du, Yijing Wang, Zhiqiang Zuo, and Wentao Zhang. Event-triggered bipartite consensus for multi-agent systems subject to multiplicative and additive noises. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003095.

Deng:2024:ENN

[DX24] Dingwen Deng and Xiaohong Xiong. Explicit, non-negativity-preserving and maximum-principle-satisfying finite difference scheme for the nonlinear Fisher’s equation. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006367.

Deng:2020:RMA

[DXS^{+}20] ZhengHong Deng, Jiwei Xu, Qun Song, Bin Hu, Tao Wu, and Panfei Huang. Robustness of multi-agent formation based on natural connectivity. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306289.

Dong:2023:API

[DXZ^{+}23] Lewei Dong, Huiling Xu, Liming Zhang, Zhengcai Li, and Yuqing Chen. Adjustable proportionalintegral multivariable observerbased FDI attack dynamic reconstitution and secure control for cyberphysical systems. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200830X.

Devi:2021:OSI

[DYA21] Munesh Devi, Shalini Yadav, and Rajan Arora. Optimal system, invariance analysis of fourth-order nonlinear Ablowitz–Kaup–Newell-Segur water wave dynamical equation using lie symmetry approach. Applied Mathematics and Computation, 404(??):Article 126230, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003209.

Deng:2021:HAP

[DYT21] Guoting Deng, Yin Yang, and Emran Tohidi. High accurate pseudo-spectral Galerkin scheme for pantograph type Volterra integrodifferential equations with singular kernels. Applied Mathematics and Computation, 396(??):Article 125866, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308195.

Du:2020:ATP

[DYW^{+}20] Haibo Du, Bo Yu, Jiajia Wei, Jun Zhang, Di Wu, and Weiqing Tao. Attitude trajectory planning and attitude control for quad-rotor aircraft based on finite-time control technique. Applied Mathematics and Computation, 386 (??):Article 125493, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304513.

Dong:2024:EBA

[DYX24] Bolei Dong, Liping Yan, and Yuanqing Xia. Event-based appointed-time cooperative formation for linear multiagent systems with actuator saturation. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003023.

Deng:2020:NCS

[DYZC20] Xingchao Deng, Qingye Yao, Yanbing Zhang, and Xudong Cui. A note on a conjecture of star chromatic index for outerplanar graphs. Applied Mathematics and Computation, 384 (??):Article 125353, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303179.

Deng:2021:MBP

[DZ21] Yunsheng Deng and Jihui Zhang. Memory-based prisoner’s dilemma game with history optimal strategy learning promotes cooperation on interdependent networks. Applied Mathematics and Computation, 390 (??):Article 125675, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306287.

Deng:2022:ISS

[DZDL22] Yalin Deng, Huasheng Zhang, Yuzhen Dai, and Yuanen Li. Interval stability/stabilization for linear stochastic switched systems with time-varying delay. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002752.

Dai:2022:FDF

[DZL^{+}22] Shifang Dai, Lijuan Zha, Jinliang Liu, Xiangpeng Xie, and Engang Tian. Fault detection filter design for networked systems with cyber attacks. Applied Mathematics and Computation, 412(??):Article 126593, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006779.

Ding:2024:MSF

[DZL^{+}24] Ning Ding, Yu Zhu, Xinyan Liu, Dapeng Dong, and Yang Wang. A modified social force model for crowd evacuation considering collision predicting behaviors. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006173.

Deng:2021:CSU

[DZW21] Lili Deng, Xingxing Zhang, and Cheng Wang. Coevolution of spatial ultimatum game and link weight promotes fairness. Applied Mathematics and Computation, 392(??):Article 125698, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306512.

Du:2023:SFA

[DZX23] Leihao Du, Zhipeng Zhang, and Chengyi Xia. A stateflipped approach to complete synchronization of Boolean networks. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008566.

Dong:2020:NRR

[DZZ^{+}20] Shiyu Dong, Hong Zhu, Shouming Zhong, Kaibo Shi, Jun Cheng, and Wei Kang. New result on reliable H∞ performance state estimation for memory static neural networks with stochastic sampled-data communication. Applied Mathematics and Computation, 364(??):

??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306113.

Dong:2021:NSF

[DZZ^{+}21] Shiyu Dong, Hong Zhu, Shouming Zhong, Kaibo Shi, and Yajuan Liu. New study on fixed-time synchronization control of delayed inertial memristive neural networks. Applied Mathematics and Computation, 399(??):Article 126035, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000837.

Di:2023:RSD

[DZZ23] Ying Di, Jin-Xi Zhang, and Xuefeng Zhang. Robust stabilization of descriptor fractional-order interval systems with uncertain derivative matrices. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300245X.

Elaiw:2021:GDS

[EA21] A. M. Elaiw and A. D. Al Agha. Global dynamics of SARS-CoV-2/cancer model with immune responses. Applied Mathematics and Computation, 408 (??):Article 126364, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004537.

Eidous:2024:KME

[EA24] Omar M. Eidous and Enas A. Ananbeh. Kernel method for estimating overlapping coefficient using numerical integration methods. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005088.

Estrada:2022:NSM

[EB22] Ernesto Estrada and Paolo Bartesaghi. From networked SIS model to the Gompertz function. Applied Mathematics and Computation, 419(??):Article 126882, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009656.

Echeverribar:2023:ERT

[EBGN23] I. Echeverribar, P. Brufau, and P. García-Navarro. Extension of a Roe-type Riemann solver scheme to model non-hydrostatic pressure shallow flows. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007147.

Erfani:2020:EEG

[EBJ20] S. Erfani, E. Babolian, and S. Javadi. Error estimates of generalized spectral iterative methods with accurate convergence rates for solving systems of fractional twopoint boundary value problems. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306307.

Escalante:2021:VHO

[ECS21] C. Escalante, M. J. Castro, and M. Semplice. Very high order wellbalanced schemes for nonprismatic one-dimensional channels with arbitrary shape. Applied Mathematics and Computation, 398 (??):Article 125993, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000412.

Erman:2020:CSA

[ED20] Sertaç Erman and Ali Demir. On the construction and stability analysis of the solution of linear fractional differential equation. Applied Mathematics and Computation, 386(??):Article 125425, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303866.

Emerenini:2022:RSM

[EE22] Blessing O. Emerenini and Hermann J. Eberl. Reactor scale modeling of quorum sensing induced biofilm dispersal. Applied Mathematics and Computation, 418 (??):Article 126792, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008742.

Etangsale:2023:FHI

[EFF23] Grégory Etangsale, Marwan Fahs, and Vincent Fontaine. Families of hybridized interior penalty discontinuous Galerkin methods for locally degenerate advection– diffusion–reaction problems. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300293X.

Evans:2020:TVN

[EFJO20] J. D. Evans, H. L. França, I. L. Palhares Junior, and C. M. Oishi. Testing viscoelastic numerical schemes using the Oldroyd-B fluid in Newtonian kinematics. Applied Mathematics and Computation, 387(??):Article 125106, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320300758.

Eftekhari:2021:SFE

[Eft21] S. A. Eftekhari. A simple finite element procedure for free vibration of rectangular thin and thick plates. Applied Mathematics and Computation, 401 (??):Article 126104, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001521.

Egidi:2023:FMN

[EGMY23] Nadaniela Egidi, Josephin Giacomini, Pierluigi Maponi, and Michael Youssef. An FFT method for the numerical differentiation. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000255.

Elaiw:2020:GDR

[EHA20] A. M. Elaiw, A. D.

Hobiny, and A. D. Al Agha. Global dynamics of reaction-diffusion oncolytic M1 virotherapy with immune response. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307507.

Ezquerro:2024:EAS

[EHVMM24] J. A. Ezquerro, M. A.

Hernández-Verón, Á. A. Magreñán, and A. Moysi. On the existence and the approximation of solutions of Volterra integral equations of the second kind. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400290X.

ElJarroudi:2020:MMN

[EKKE20] Mustapha El Jarroudi, Hasan Karjoun, Louis Kouadio, and Moussa El Jarroudi. Mathematical modelling of non-local spore dispersion of wind-borne pathogens causing fungal diseases. Applied Mathematics and Computation, 376 (??):Article 125107, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030076X.

Elgazzar:2020:QPD

[Elg20] Ahmed S. Elgazzar. Quantum prisoner’s dilemma in a restricted one-parameter strategic space. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309191.

Elgazzar:2022:CQP

[Elg22] Ahmed S. Elgazzar. Corrigendum to “Quantum prisoner’s dilemma in a restricted one-parameter strategic space” [Applied Mathematics and Computation 370 (2020) 124927]. Applied Mathematics and Computation, 418(??):Article 126810, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008936.

Ehigie:2022:EFT

[ELOY22] Julius O. Ehigie, Vu Thai Luan, Solomon A. Okunuga, and Xiong You. Exponentially fitted two-derivative DIRK methods for oscillatory differential equations. Applied Mathematics and Computation, 418 (??):Article 126770, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008523.

Eltaher:2020:NSV

[EM20] Mohamed A. Eltaher and Nazira Mohamed. Nonlinear stability and vibration of imperfect CNTs by doublet mechanics. Applied Mathematics and Computation, 382(??):Article 125311, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302770.

Egidi:2021:SMS

[EM21] Nadaniela Egidi and Pierluigi Maponi. A spectral method for the solution of boundary value problems. Applied Mathematics and Computation, 409 (??):Article 125812, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320307657.

Emami:2021:DGA

[Ema21] Babak Emami. A differential geometric approach to time series forecasting. Applied Mathematics and Computation, 402 (??):Article 126150, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001983.

Eftekhari:2022:NEO

[ER22] Tahereh Eftekhari and Jalil Rashidinia. A novel and efficient operational matrix for solving nonlinear stochastic differential equations driven by multi-fractional Gaussian noise. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002922.

Escot:2023:ELE

[ES23] Lorenzo Escot and Julio E.

Sandubete. Estimating Lyapunov exponents on a noisy environment by global and local Jacobian indirect algorithms. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005720.

Epebinu:2024:IFI

[ES24] Abayomi Dennis Epebinu and Tomasz Szostok. Inequalities for fractional integral with the use of stochastic orderings. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006501.

El-Sayed:2020:SCS

[ESA20] A. M. A. El-Sayed and Reda Gamal Ahmed. Solvability of a coupled system of functional integrodifferential equations with infinite point and Riemann– Stieltjes integral conditions. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309105.

Estrada:2021:ICN

[Est21] Ernesto Estrada. Informational cost and networks navigability. Applied Mathematics and Computation, 397(??):Article 125914, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308675.

Efendiev:2021:SMS

[EV21] Yalchin Efendiev and Petr N. Vabishchevich. Splitting methods for solution decomposition in nonstationary problems. Applied Mathematics and Computation, 397(??):Article 125785, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307384.

Evans:2020:HWN

[Eva20] J. D. Evans. High Weissenberg number boundary layer structures for UCM fluids. Applied Mathematics and Computation, 387(??):Article 124952, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309440.

Hamdi:2021:RSS

[EVB^{+}21] Issam El Hamdi, Alessandro N. Vargas, Hassane Bouzahir, Ricardo C. L. F. Oliveira, and Leonardo Acho. Robust stability of stochastic systems with varying delays: Application to RLC circuit with intermittent closedloop. Applied Mathematics and Computation, 411 (??):Article 126541, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006251.

Elshenhab:2021:RSL

[EW21] Ahmed M. Elshenhab and Xing Tao Wang. Representation of solutions of linear differential systems with pure delay and multiple delays with linear parts given by non-permutable matrices. Applied Mathematics and Computation, 410(??):Article 126443, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005324.

Fabrizio:2020:TPV

[Fab20] Mauro Fabrizio. Turbulence phenomena for viscous fluids: Vortices and instability. Applied Mathematics and Computation, 376(??): Article 125094, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300631.

Fawzy:2024:MSF

[FAEZ24] Doaa M. Fawzy, Ayman A. Arafa, A. Elsaid, and W. K. Zahra. Mosquito suppression via Filippov incompatible insect technique. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003692.

Fang:2021:ICS

[Fan21] Aili Fang. The influence of communication structure on opinion dynamics in social networks with multiple true states. Applied Mathematics and Computation, 406(??):Article 126262, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003519.

Fareo:2022:NTB

[Far22]  A. G. Fareo.  A note

on the transformation of boundary value problems to initial value problems: the iterative transformation method. Applied Mathematics and Computation, 415 (??):Article 126692, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007761.

Feyzollahzadeh:2020:MTM

[FB20] Mahdi Feyzollahzadeh and Mahdi Bamdad. A modified transfer matrix method to reduce the calculation time: a case study on beam vibration. Applied Mathematics and Computation, 378(??):Article 125238, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302071.

Frischmuth:2021:MIH

[FBT21] Kurt Frischmuth, Alexander V. Budyansky, and Vyacheslav G. Tsybulin. Modeling of invasion on a heterogeneous habitat: taxis and multistability. Applied Mathematics and Computation, 410(??):Article 126456, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005452.

Fiaschi:2021:NAG

[FC21] Lorenzo Fiaschi and Marco Cococcioni. Non-Archimedean game theory: a numerical approach. Applied Mathematics and Computation, 409(??):Article 125356, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303209.

Fu:2021:MEF

[FCeFZ21] Shihua Fu, Daizhan Cheng, Jun e Feng, and Jianli Zhao. Matrix expression of finite Boolean-type algebras. Applied Mathematics and Computation, 395(??): Article 125880, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030833X.

Frasca-Caccia:2021:NPM

[FCH21] Gianluca Frasca-Caccia and Peter E. Hydon. Numerical preservation of multiple local conservation laws. Applied Mathematics and Computation, 403 (??):Article 126203, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002939.

Fu:2021:UBL

[FCX21] Chunhong Fu, Jiajia Chen, and Qingxiang Xu. Upper bounds and lower bounds for the Frobenius norm of the solution to certain structured Sylvester equation. Applied Mathematics and Computation, 399(??): Article 126007, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000552.

Froguel:2022:ECR

[FdLPC^{+}22] Lucas Belasque Froguel, Thiago de Lima Prado, Gilberto Corso, Gustavo Zampier dos Santos Lima, and Sergio Roberto Lopes. Efficient computation of recurrence quantification analysis via microstates. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002491.

Feng:2022:EDI

[Fen22] Shu Feng. Effective document image binarization via a convex variational level set model. Applied Mathematics and Computation, 419(??):Article 126861, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009449.

Fahad:2021:OCC

[FF21a] Hafiz Muhammad Fahad and Arran Fernandez. Operational calculus for Caputo fractional calculus with respect to functions and the associated fractional differential equations. Applied Mathematics and Computation, 409 (??):Article 126400, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004896.

Fusi:2021:FSS

[FF21b] Lorenzo Fusi and Angiolo Farina. Flow stability of suspensions. Applied Mathematics and Computation, 410(??):Article 126466, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005555.

Feng:2023:IRR

[FFL23a] Bo Feng, Zhiguang Feng, and Peng Li. Improved results on reachable set synthesis of Markovian jump systems with timevarying delays: General asynchronous control approaches. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000358.

Feng:2023:NIS

[FFL^{+}23b] Tian-Jiao Feng, Song-Jia Fan, Cong Li, Yi Tao, and Xiu-Deng Zheng. Noiseinduced sustainability of cooperation in Prisoner’s dilemma game. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006762.

Frolkovic:2024:USH

[FG24] Peter Frolkovic and Nikola Gajdosová. Unconditionally stable higher order semiimplicit level set method for advection equations. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300629X.

Figueroa:2024:EAL

[FGM24] Silvano Figueroa, Eduardo M. Garau, and Pedro Morin. Error analysis for local coarsening in univariate spline spaces. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000882.

Falcone:2021:SBS

[FGMS21] Alberto Falcone, Alfredo Garro, Marat S. Mukhametzhanov, and Yaroslav D. Sergeyev. A Simulinkbased software solution using the Infinity Computer methodology for higher order differentiation. Applied Mathematics and Computation, 409 (??):Article 125606, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305610.

Flores:2024:ECS

[FH24] Lucas S. Flores and The Anh Han. Evolution of commitment in the spatial public goods game through institutional incentives. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001188.

Feng:2024:EGR

[FHFS24] Kehuan Feng, Songlin Han, Minyu Feng, and Attila Szolnoki. An evolutionary game with reputation-based imitation-mutation dynamics. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000900.

Feng:2022:HOC

[FHM22] Qiwei Feng, Bin Han, and Peter Minev. A high order compact finite difference scheme for elliptic interface problems with discontinuous and high-contrast coefficients. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003885.

Furtado:2024:EAP

[FJ24] Susana Furtado and Charles R. Johnson. Efficiency analysis for the Perron vector of a reciprocal matrix. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003746.

Fang:2021:PBS

[FJFS21] Tian Fang, Shiyu Jiao, Dongmei Fu, and Lei Su. Passivity-based synchronization for Markov switched neural networks with time delays and the inertial term. Applied Mathematics and Computation, 394(??):Article 125786, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307396.

Fu:2021:FAF

[FJJJ21] Yaru Fu, Xiaoyu Jiang, Zhaolin Jiang, and Seongtae Jhang. Fast algorithms for finding the solution of CUPL–Toeplitz linear system from Markov chain. Applied Mathematics and Computation, 396(??): Article 125859, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308122.

Fan:2022:RFT

[FJLG22] Yanyan Fan, Zhenlin Jin, Xiaoyuan Luo, and Baosu Guo. Robust finite-time consensus control for Euler– Lagrange multi-agent systems subject to switching topologies and uncertainties. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004416.

Feng:2021:GBM

[FJV21] Runhuan Feng, Pingping Jiang, and Hans Volkmer. Geometric Brownian motion with affine drift and its time-integral. Applied Mathematics and Computation, 395(??):Article 125874, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320308274.

Faught:2024:NGT

[FKK24] J. Nolan Faught, Mark Kempton, and Adam Knudson. A Nordhaus–Gaddum type problem for the normalized Laplacian spectrum and graph Cheeger constant. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003813.

Falcone:2023:AOC

[FKS23] Maurizio Falcone, Gerhard Kirsten, and Luca Saluzzi. Approximation of optimal control problems for the Navier–Stokes equation via multilinear HJB-POD. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007901.

Farouki:2021:SCC

[FKVZ21] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Zagar. Spatial C^{2} closed loops of prescribed arc length defined by Pythagorean-hodograph curves. Applied Mathematics and Computation, 391 (??):Article 125653, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030607X.

Farouki:2022:CPM

[FKVZ22] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Zagar. On the construction of polynomial minimal surfaces with Pythagorean normals. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005136.

Feng:2022:SGD

[FLC^{+}22] Lichao Feng, Qiumei Liu, Jinde Cao, Chunyan Zhang, and Fawaz Alsaadi. Stabilization in general decay rate of discrete feedback control for non-autonomous Markov jump stochastic systems. Applied Mathematics and Computation, 417(??): Article 126771, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008535.

Fu:2021:ORO

[FLH^{+}21] Haijing Fu, Jiahui Li, Fei Han, Nan Hou, and Hongli Dong. Outlierresistant observer-based H_{∞} PID control under stochastic communication protocol. Applied Mathematics and Computation, 411(??):Article 126535, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006184.

Fu:2023:EBR

[FLHH23] Xiao Fu, Xuesong Liu, Yuhan Hu, and Shaoxu Huang. Emotion-based renewal strategies to promote cooperation in spatial prisoner’s dilemma games. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002990.

Feng:2020:MHI

[FLL^{+}20] Lihua Feng, Ziyuan Li, Weijun Liu, Lu Lu, and Dragan Stevanović. Minimal Harary index of unicyclic graphs with diameter at most

4. Applied Mathematics and Computation, 381(??): Article 125315, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302812.

Feng:2022:ATR

[FLL^{+}22] Lihua Feng, Weijun Liu, Lu Lu, Wei Wang, and Guihai Yu. The access time of random walks on trees with given partition. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002478.

Fu:2023:DET

[FLLW23] Yingying Fu, Jing Li, Xiaobo Li, and Shuiyan Wu. Dynamic eventtriggered adaptive control for uncertain stochastic nonlinear systems. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008682.

Flandoli:2023:NCP

[FLR23] Franco Flandoli, Dejun Luo, and Cristiano Ricci. Numerical computation of probabilities for nonlinear SDEs in high dimension using Kolmogorov equation. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200594X.

Feng:2020:SRG

[FLS20] Lihua Feng, Lu Lu, and Dragan Stevanović. A short remark on graphs with two main eigenvalues. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308501.

Feng:2021:SAN

[FLWL21] Lichao Feng, Lei Liu, Zhihui Wu, and Qiumei Liu. Stability analysis for nonlinear Markov jump neutral stochastic functional differential systems. Applied Mathematics and Computation, 394(??):Article 125782, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307359.

Fan:2023:SAD

[FLZ23] Lina Fan, Yuan Lv, and Quanxin Zhu. Stability analysis of discrete-time switched stochastic nonautonomous systems with external inputs and timevarying delays under partially unstable subsystems. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008542.

Fan:2024:PBM

[FLZZ24] Hongtao Fan, Yajing Li, Hongbing Zhang, and Xinyun Zhu. Preconditioners based on matrix splitting for the structured systems from elliptic PDE-constrained optimization problems. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005106.

Farago:2023:SQP

[FM23] I. Faragó and R. Mosleh.

Some qualitative properties of the discrete models for malaria propagation. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007019.

Fotsa-Mbogne:2022:NON

[FMHDB22] David Jaures Fotsa-Mbogne, Duplex Elvis Houpa-Danga, and David Bekolle. Nonlinear observer of the nonspatial dynamics of anthracnose: Case study of coffee berry disease. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004507.

Fornasier:2021:RRL

[FMN21] Massimo Fornasier, Johannes Maly, and Valeriya Naumova. Robust recovery of lowrank matrices with nonorthogonal sparse decomposition from incomplete measurements. Applied Mathematics and Computation, 392(??):Article 125702, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030655X.

Feher:2020:AOL

[FMP20] Áron Fehér, Lőrinc Márton, and Mihály Pituk. Asymptotically ordinary linear Volterra difference equations with infinite delay. Applied Mathematics and Computation, 386 (??):Article 125499, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304574.

Fabrici:2021:NHG

[FMT^{+}21] Igor Fabrici, Tomás Madaras, Mária Timková, Nico Van Cleemput, and Carol T. Zamfirescu. Non-hamiltonian graphs in which every edge-contracted subgraph is Hamiltonian. Applied Mathematics and Computation, 392(??):Article 125714, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306676.

Feng:2022:TDS

[FN22] Huolei Feng and Yushan Ni. Temperature-dependent switchable thermal bifunctions in different diamondshaped devices. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000923.

Fang:2022:EPS

[FQ22] Jin-Hui Fang and Guoyou Qian. An extremal problem in subsequence sum. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000832.

Fu:2024:ASF

[FQHS24] Libi Fu, Huigui Qin, Yangjian He, and Yongqian Shi. Application of the social force modelling method to evacuation dynamics involving pedestrians with disabilities. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300323004666.

Freud:2023:BTC

[FR23] Thomas Freud and Pablo M. Rodriguez. The Bell– Touchard counting process. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008098.

Fu:2024:TLT

[FR24] Feng Fu and Daniel N.

Rockmore. Too little, too late — a dynamical systems model for gun-related violence and intervention. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006641.

Frego:2022:CFP

[Fre22]  Marco Frego.  Closed

form parametrisation of 3D clothoids by arclength with both linear varying curvature and torsion. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321009905.

Furtula:2022:GZZ

[FRR^{+}22] Boris Furtula, Slavko Radenković, Izudin Redzepović, Niko Tratnik, and Petra Zigert Pletersek. The generalized Zhang–Zhang polynomial of benzenoid systems — theory and applications. Applied Mathematics and Computation, 418 (??):Article 126822, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100905X.

Fermo:2024:ANI

[FRRS24] Luisa Fermo, Lothar Reichel, Giuseppe Rodriguez, and Miodrag M. Spalević. Averaged Nyström interpolants for the solution of Fredholm integral equations of the second kind. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006513.

Fernandez:2022:LFD

[FRS22] Arran Fernandez, Joel E. Restrepo, and Durvudkhan Suragan. On linear fractional differential equations with variable coefficients. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004441.

Fang:2024:FRA

[FRW^{+}24] Jiankang Fang, Chengcheng Ren, Hai Wang, Vladimir Stojanovic, and Shuping He. Finite-region asynchronous H_{∞} filtering for 2-D Markov jump systems in Roesser model. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000456.

Fu:2022:NRA

[FSLC22] Xiuwen Fu, Zhaoliang Sheng, Chong Lin, and Bing Chen. New results on admissibility and dissipativity analysis of descriptor time-delay systems. Applied Mathematics and Computation, 419(??): Article 126860, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009437.

Fu:2023:DSE

[FSLZ23] Xiaoyu Fu, Xinmin Song, Xiyu Liu, and Min Zhang. Distributed state estimation with state equality constraints in the presence of packet dropping. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002126.

Feng:2020:FRE

[FSS20] Zongying Feng, Hanyong Shao, and Lin Shao. Further results on eventtriggered H_{∞} networked control for neural networks with stochastic cyberattacks. Applied Mathematics and Computation, 386(??):Article 125431, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303921.

Fiedorowicz:2021:PCP

[FSS21] Anna Fiedorowicz, Elzbieta Sidorowicz, and Éric Sopena. Proper connection and proper-walk connection of digraphs. Applied Mathematics and Computation, 410(??):Article 126253, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100343X.

Fan:2024:FAR

[FT24] Xianrui Fan and Shaocheng Tong. Fuzzy adaptive resilient decentralized control of nonlinear interconnected cyber-physical systems under false data injection attacks. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000031.

Feng:2022:ISD

[FTML22] Libo Feng, Ian Turner, Timothy Moroney, and Fawang Liu. An investigation of space distributedorder models for simulating anomalous transport in a binary medium. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004970.

Fu:2024:DBN

[FW24] Xinjie Fu and JinRong Wang. Dynamic behaviors and non-instantaneous impulsive vaccination of an SAIQR model on complex networks. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323005945.

Feng:2023:HMD

[FWBY23] Jinhua Feng, Shixi Wang, Hai Bi, and Yidu Yang. An hp-mixed discontinuous Galerkin method for the biharmonic eigenvalue problem. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001388.

Farooq:2022:CNT

[FWM^{+}22] Umar Farooq, Hassan Waqas, Taseer Muhammad, Muhammad Imran, and Ali Saleh Alshomrani. Computation of nonlinear thermal radiation in magnetized nanofluid flow with entropy generation. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321009838.

Fu:2020:OPB

[FX20] Chunhong Fu and Qingxiang Xu. An optimal perturbation bound for the partial isometry associated to the generalized polar decomposition. Applied Mathematics and Computation, 372(??): Article 124987, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309798.

Fortin:2023:VHO

[FY23] A. Fortin and D. Yakoubi. A very high order discontinuous Galerkin method for the numerical solution of stiff DDEs. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008359.

Faybusovich:2020:SCM

[FZ20a] Leonid Faybusovich and Cunlu Zhou. Self-concordance and matrix monotonicity with applications to quantum entanglement problems. Applied Mathematics and Computation, 375(??): Article 125071, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300400.

Fu:2020:ESN

[FZ20b] Xiaozheng Fu and Quanxin Zhu. Exponential stability of neutral stochastic delay differential equation with delay-dependent impulses. Applied Mathematics and Computation, 377 (??):Article 125146, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320301156.

Fu:2022:SST

[FZ22] Teng Fu and Yusheng Zhou. Stabilization of switched time-delay systems with only unstable subsystems: a new approach based on a vibration model of 1.5 degrees of freedom. Applied Mathematics and Computation, 415 (??):Article 126740, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321008225.

Frolkovic:2023:HRC

[FZ23] Peter Frolkovic and Michal Zeravý. High resolution compact implicit numerical scheme for conservation laws. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007883.

Faiz:2024:NFM

[FZB24] Zakaria Faiz, Shengda Zeng, and Hicham Benaissa. A novel fractional Moreau’s sweeping process with applications. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003783.

Fan:2023:GCA

[FZG23] Ruilin Fan, Hanlu Zhang, and Yi Gao. The global cooperation in asteroid mining based on AHP, entropy and TOPSIS. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006099.

Feng:2023:ERS

[FZLF23] Zhiguang Feng, Xinyue Zhang, James Lam, and Chenchen Fan. Estimation of reachable set for switched singular systems with time-varying delay and state jump. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003016.

Feng:2024:NMR

[FZLJ24] Zhiguang Feng, Xinyue Zhang, Jason J. R. Liu, and Zhengyi Jiang. A new method of reachable sets estimation for the nonlinear switched singular system with impulsive performance and time-delay. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003904.

Fu:2021:ESM

[FZS21] Qianhua Fu, Shouming Zhong, and Kaibo Shi. Exponential synchronization of memristive neural networks with inertial and nonlinear coupling terms: Pinning impulsive control approaches. Applied Mathematics and Computation, 402(??):Article 126169, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002599.

Feng:2023:NSS

[FZW23] Likang Feng, Weihai Zhang, and Zhaojing Wu. Noise-to-state stability of random impulsive delay systems with multiple random impulses. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005914.

GaziKarakoc:2020:ACA

[GA20] Seydi Battal GaziKarakoc and Khalid K. Ali. Analytical and computational approaches on solitary wave solutions of the generalized equal width equation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309257.

Gok:2021:SBR

[GA21] Gülistan Kaya Gök and Ali Reza Ashrafi. Some bounds for the resolvent energy. Applied Mathematics and Computation, 397 (??):Article 125958, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000060.

Gokcer:2022:AKT

[GA22] Türkan Yeliz Gökçer and Ismail Aslan. Approximation by Kantorovichtype max-min operators and its applications. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000972.

Ghaderi:2024:SUL

[GAA^{+}24] Susan Ghaderi, Masoud Ahookhosh, Adam Arany, Alexander Skupin, Panagiotis Patrinos, and Yves Moreau. Smoothing unadjusted Langevin algorithms for nonsmooth composite potential functions. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005465.

Garcia-Ancona:2020:RAS

[GAMP20] R. García-Ancona, J. Morais, and R. Michael Porter. Relations among spheroidal and spherical harmonics. Applied Mathematics and Computation, 384 (??):Article 125147, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301168.

Ganie:2021:DLE

[Gan21] Hilal A. Ganie. On the distance Laplacian energy ordering of a tree. Applied Mathematics and Computation, 394(??):Article 125762, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307153.

Ganai:2023:PTS

[Gan23] Tinku Ganai. Perturbation theory of structured matrix pencils with no spillover. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003867.

Garimella:2022:SSF

[GAR22] Sai Manikiran Garimella, Mohan Anand, and Kumbakonam R. Rajagopal. Start-up shear flow of a shear-thinning fluid that approximates the response of viscoplastic fluids. Applied Mathematics and Computation, 412(??):Article 126571, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100655X.

Ghorbani:2024:FTC

[GARD24] Modjtaba Ghorbani, Razie Alidehi-Ravandi, and Matthias Dehmer. Fullerenes via their counting polynomials. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006008.

Gasiorek:2020:ACS

[Gas20] Marcin Gasiorek. On algorithmic Coxeter spectral analysis of positive posets. Applied Mathematics and Computation, 386(??):Article 125507, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304653.

Gasiorek:2023:CRM

[Gas23] Marcin Gasiorek. Congruence of rational matrices defined by an integer matrix. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007111.

Gande:2020:TOW

[GB20] Naga Raju Gande and Ashlesha A. Bhise. Thirdorder WENO schemes with kinetic flux vector splitting. Applied Mathematics and Computation, 378 (??):Article 125203, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301727.

Goksu:2021:FTS

[GB21] Gökhan Göksu and Ulviye Baser. Finite-time stability for switched linear systems by Jordan decomposition. Applied Mathematics and Computation, 395(??): Article 125853, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308067.

Golmankhaneh:2024:ESS

[GB24] Alireza Khalili Golmankhaneh and Donatella Bongiorno. Exact solutions of some fractal differential equations. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400105X.

Gomez-Bueno:2021:HOW

[GBCP21] Irene Gómez-Bueno, Manuel J.

Castro, and Carlos Parés. High-order well-balanced methods for systems of balance laws: a controlbased approach. Applied Mathematics and Computation, 394(??):Article 125820, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307736.

Granados:2021:EMM

[GBF21] J. M. Granados, C. A. Bustamante, and W. F. Florez. Extending meshless method of approximate particular solutions (MAPS) to two-dimensional convection heat transfer problems. Applied Mathematics and Computation, 390 (??):Article 125484, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304434.

Ghorban:2021:LES

[GBH^{+}21] Samira Hossein Ghorban, Fatemeh Baharifard, Bardyaa Hesaam, Mina Zarei, and Hamid Sarbazi-Azad. Linearization error in synchronization of Kuramoto oscillators. Applied Mathematics and Computation, 411(??):Article 126464, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005531.

Ghanbari:2020:CMS

[GC20] Behzad Ghanbari and Changbum Chun. A constructive method for solving the equation Xa = b in R^{n}: a generalization of division in R^{n}. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306654.

Ge:2024:ETC

[GC24] Fudong Ge and YangQuan Chen. Event-triggered control for boundary controlled time-fractional diffusion systems with spatiallyvarying coefficients. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002911.

Gascon:2021:NSQ

[GCGM21] Ll. Gascón, J. M. Corberán, and J. A. García-Manrique. Numerical schemes for quasi-1D steady nozzle flows. Applied Mathematics and Computation, 400(??): Article 126072, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100120X.

Gomez-Corral:2023:MEM

[GCLHT23] A. Gómez-Corral, M. J. Lopez-Herrero, and D. Taipe. A Markovian epidemic model in a resource-limited environment. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004216.

Garres-Diaz:2021:FED

[GDB21] José Garres-Díaz and Luca Bonaventura. Flexible and efficient discretizations of multilayer models with variable density. Applied Mathematics and Computation, 402(??):Article 126097, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001454.

Garres-Diaz:2022:SIA

[GDFNNR22] J. Garres-Díaz, E. D. Fernández-Nieto, and G. Narbona-Reina. A semi-implicit approach for sediment transport models with gravitational effects. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000248.

Gatti:2024:SWB

[GdFPF24] Federico Gatti, Carlo de Falco, Simona Perotto, and Luca Formaggia. A scalable well-balanced numerical scheme for the simulation of fast landslides with efficient time stepping. Applied Mathematics and Computation, 468(??): ??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300694X.

Ghorbani:2020:WSG

[GDM^{+}20] Modjtaba Ghorbani, Matthias Dehmer, Hamidreza Maimani, Sheyda Maddah, Maryam Roozbayani, and Frank Emmert-Streib. The watching system as a generalization of identifying code. Applied Mathematics and Computation, 380(??): Article 125302, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302460.

Guo:2024:ECI

[GDS24] Tian Guo, Chunpeng Du, and Lei Shi. Evolution of cooperation on interdependent networks: the impact of asymmetric punishment. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005192.

Georgiev:2020:TSM

[GE20] Svetlin G. Georgiev and Inci M. Erhan. The Taylor series method and trapezoidal rule on time scales. Applied Mathematics and Computation, 378 (??):Article 125200, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301697.

Galina:2021:DIA

[GE21] Reshetova Galina and Romenski Evgeniy. Diffuse interface approach to modeling wavefields in a saturated porous medium. Applied Mathematics and Computation, 398(??):Article 125978, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000266.

Grim:2023:CCB

[GF23] Anna Grim and Pedro Felzenszwalb. Convex combination belief propagation. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006464.

Garg:2022:GTS

[GG22] Naveen Kumar Garg and G. D. Veerappa Gowda. Godunov-type schemes for the pressureless gas dynamics and related models. Applied Mathematics and Computation, 418 (??):Article 126790, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008729.

Gao:2024:ETE

[GG24] Wei Gao and Yubin Gao.

The extremal trees for exponential vertex-degree-based topological indices. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001061.

Gutierrez-Gutierrez:2023:NSC

[GGBEZRI23] Jesús Gutiérrez-Gutiérrez, Íñigo Barasoain-Echepare, Marta Zárraga-Rodríguez, and Xabier Insausti. Necessary and sufficient conditions for AR vector processes to be stationary: Applications in information theory and in statistical signal processing. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200892X.

Gao:2023:SSB

[GGL^{+}23] Zilin Gao, Changyuan Guo, Yongfu Li, Lizhi Liu, and Weimin Luo. Stabilization of a structurally balanced complex network with similar nodes of different dimensions. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004071.

Gao:2021:MRE

[GGS21] Yubin Gao, Wei Gao, and Yanling Shao. The minimal Randić energy of trees with given diameter. Applied Mathematics and Computation, 411 (??):Article 126489, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005786.

Ganegoda:2021:ADM

[GGW21a] Naleen Ganegoda, Thomas Götz, and Karunia Putra Wijaya. An agedependent model for dengue transmission: Analysis and comparison to field data. Applied Mathematics and Computation, 388 (??):Article 125538, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030494X.

Gao:2021:BSC

[GGW21b] You Gao, Jinru Gao, and Gang Wang. Bounds on subspace codes based on subspaces of type (s, 0, 0, 0) in pseudo-sympletic spaces and singular pseudo-symplectic spaces. Applied Mathematics and Computation, 407 (??):Article 126322, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004112.

Gao:2022:ASSb

[GGY22a] Yin Gao, Jinwu Gao, and Xiangfeng Yang. The almost sure stability for uncertain delay differential equations based on normal Lipschitz conditions. Applied Mathematics and Computation, 420(??): Article 126903, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009863.

Gao:2022:PEU

[GGY22b] Yin Gao, Jinwu Gao, and Xiangfeng Yang. Parameter estimation in uncertain delay differential equations via the method of moments. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200385X.

Guo:2023:MCN

[GGY23] Pei-Chang Guo, Shi-Chen Gao, and Yong-Qing Yang. Monotone convergence of Newton-like iteration for a structured nonlinear eigenproblem. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006063.

Goddard:2020:IDC

[GH20a] Wayne Goddard and Michael A.

Henning. Independent domination, colorings and the fractional idomatic number of a graph. Applied Mathematics and Computation, 382(??):Article 125340, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303064.

Gyori:2020:EEL

[GH20b] István Győri and László Horváth. Explicit estimates and limit formulae for the solutions of linear delay functional differential systems with nonnegative Volterra type operators. Applied Mathematics and Computation, 385 (??):Article 125451, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304124.

Guo:2021:NEG

[GH21] Peilian Guo and Changda Han. Nash equilibrium and group strategy consensus of networked evolutionary game with coupled social groups. Applied Mathematics and Computation, 409 (??):Article 126380, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004690.

Ghorbani:2022:NRP

[GHND22] Modjtaba Ghorbani, Mardjan Hakimi-Nezhaad, and Matthias Dehmer. Novel results on partial Hosoya polynomials: an application in chemistry. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004532.

Guan:2024:FTS

[GHP24] Xinyu Guan, Yanyan Hu, and Kaixiang Peng. Finitetime switching-like sliding mode fault-tolerant control for discrete-time cyberphysical systems under DoS attacks and intermittent faults. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400002X.

Guo:2021:FTC

[GHS^{+}21] Wanli Guo, Wennuo He, Lili Shi, Wen Sun, and Xiaoqing Lu. Fixed-time consensus tracking for nonlinear stochastically disturbed multi-agent systems via discontinuous protocols. Applied Mathematics and Computation, 400(??): Article 126046, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000941.

Gerot:2023:BNU

[GI23] Cédric Gérot and Ioannis Ivrissimtzis. Bivariate non-uniform subdivision schemes based on Lsystems. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003259.

Gao:2021:SMU

[GJ21] Yin Gao and Lifen Jia. Stability in mean for uncertain delay differential equations based on new Lipschitz conditions. Applied Mathematics and Computation, 399(??):Article 126050, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000989.

Gimeno:2023:NIH

[GJJC^{+}23] Joan Gimeno, Àngel Jorba, Marc Jorba-Cuscó, Narcís Miguel, and Maorong Zou. Numerical integration of high-order variational equations of ODEs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008116.

Guo:2023:IFS

[GJS^{+}23] Jin Guo, Ruizhe Jia, Ruinan Su, Yanlong Zhao, and Yong Song. Identification of FIR systems with binary-valued observations against denial-of-service attacks. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001583.

Guo:2024:AAE

[GJWV24] Zhenwei Guo, Tongsong Jiang, Gang Wang, and V. I. Vasil’ev. Algebraic algorithms for eigen-problems of a reduced biquaternion matrix and applications. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005271.

Ghosh:2020:SDK

[GK20] Sudeshna Ghosh and Manish Kumar. Study of drafting, kissing and tumbling process of two particles with different sizes and densities using immersed boundary method in a confined medium. Applied Mathematics and Computation, 386(??):Article 125411, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303672.

Gupta:2021:RSD

[GK21] Aastha Gupta and Aditya Kaushik. A robust spline difference method for robintype reaction-diffusion problem using grid equidistribution. Applied Mathematics and Computation, 390 (??):Article 125597, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030552X.

Groselj:2024:SIS

[GKK^{+}24] Jan Groselj, Mario Kapl, Marjeta Knez, Thomas Takacs, and Vito Vitrih. C^{1}- smooth isogeometric spline functions of general degree over planar mixed meshes: the case of two quadratic mesh elements. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004472.

Ghalavand:2023:LEC

[GKT^{+}23] Ali Ghalavand, Sandi Klavzar, Mostafa Tavakoli, Mardjan Hakimi-Nezhaad, and Freydoon Rahbarnia. Leap eccentric connectivity index in graphs with universal vertices. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005938.

Guo:2022:BLF

[GL22] Ying Guo and Yuze Li. Bipartite leader-following synchronization of fractionalorder delayed multilayer signed networks by adaptive and impulsive controllers. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003174.

Gao:2023:PRE

[GL23a] Shiping Gao and Nan Li.

Preference reversal and the evolution of cooperation. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006415.

Guo:2023:ECR

[GL23b] Chunyang Guo and Yuanyuan Liu. Explicit convergence rates for the M/G/1 queue under perturbation. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003727.

Guo:2024:EBA

[GL24] Ming-Juan Guo and Yuan-Xin Li. Estimator-based adaptive prescribed performance cooperative bipartite containment control of nonlinear multiagent system against DoS attacks. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000572.

Gui:2023:EQO

[GLC23] Qiang Gui, Wei Li, and Yingbin Chai. The enriched quadrilateral overlapping finite elements for time-harmonic acoustics. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300187X.

Guo:2021:EDB

[GLD21] Yaxiao Guo, Junmin Li, and Ruirui Duan. Extended dissipativity-based control for persistent dwelltime switched singularly perturbed systems and its application to electronic circuits. Applied Mathematics and Computation, 402 (??):Article 126114, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001624.

Gamboa:2024:EPB

[GLGLH24] M. Gamboa, M. López-García, and M. J. Lopez-Herrero. On the exact and population bi-dimensional reproduction numbers in a stochastic SVIR model with imperfect vaccine. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006951.

Go:2020:FCB

[GLKrH20] Myeong-Seok Go, Jae Hyuk Lim, Jin-Gyun Kim, and Ki ryoung Hwang. A family of Craig–Bampton methods considering residual mode compensation. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308148.

Geng:2022:RLE

[GLL^{+}22] Yini Geng, Yifan Liu, Yikang Lu, Chen Shen, and Lei Shi. Reinforcement learning explains various conditional cooperation. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002569.

Gao:2023:SDM

[GLL23] Shuaibin Gao, Xiaotong Li, and Zhuoqi Liu. Stationary distribution of the Milstein scheme for stochastic differential delay equations with first-order convergence. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003934.

Guo:2024:ETP

[GLLZ24] Ruqiang Guo, Linjie Liu, Yuyuan Liu, and Liang Zhang. Evolution of trust in the N-player trust game with the margin system. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001218.

Gu:2021:NSP

[GLMT21] Ruijuan Gu, Hui Lei, Yulai Ma, and Zhenyu Taoqiu. Note on (semi-)proper orientation of some triangulated planar graphs. Applied Mathematics and Computation, 392(??):Article 125723, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306767.

Guo:2020:OBA

[GLP20] Xiyue Guo, Hongjing Liang, and Yingnan Pan. Observerbased adaptive fuzzy tracking control for stochastic nonlinear multi-agent systems with dead-zone input. Applied Mathematics and Computation, 379 (??):Article 125269, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302381.

Goncalves:2022:SLS

[GLP22] M. L. N. Gonçalves, F. S.

Lima, and L. F. Prudente. A study of Liu–Storey conjugate gradient methods for vector optimization. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001837.

Gu:2022:MHL

[GLPS22] Ran Gu, Hui Lei, Yuejian Peng, and Yongtang Shi. Maximisers of the hypergraph Lagrangian outside the principal range. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004921.

Gong:2021:NSF

[GLT^{+}21] Zhaohua Gong, Chongyang Liu, Kok Lay Teo, Song Wang, and Yonghong Wu. Numerical solution of free final time fractional optimal control problems. Applied Mathematics and Computation, 405(??):Article 126270, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003593.

Gong:2022:OCN

[GLTY22] Zhaohua Gong, Chongyang Liu, Kok Lay Teo, and Xiaopeng Yi. Optimal control of nonlinear fractional systems with multiple pantograph-delays. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001783.

Gao:2020:RDN

[GLW20] Zhipeng Gao, Hui Lei, and Kui Wang. Rainbow domination numbers of generalized Petersen graphs. Applied Mathematics and Computation, 382(??):Article 125341, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303076.

Guo:2022:SSC

[GLW22] Jia Guo, Mei Lu, and Xin Wang. The (strong) structure connectivity and (strong) substructure connectivity of the (n, k) - bubble-sort network. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200162X.

Gai:2024:TFL

[GLW24] Ling Gai, Mengpei Liang, and Chenhao Wang. Twofacility-location games with mixed types of agents. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006483.

Ge:2022:ISC

[GLWL22] Chao Ge, Zengshuai Liu, Lei Wang, and Yajuan Liu. Improved stability criteria of T–S fuzzy systems with sampled-data-based dissipative control. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001333.

Gao:2022:ASSa

[GLWW22] Lei Gao, Yaotang Li, Zhen Wang, and Rui-Wu Wang. Asymmetric strategy setup solve the Prisoner’s Dilemma of the evolution of mutualism. Applied Mathematics and Computation, 412(??):Article 126590, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006743.

Goreac:2024:LTD

[GLWX24] Dan Goreac, Juan Li, Pangbo Wang, and Boxiang Xu. Linearisation techniques and the dual algorithm for a class of mixed singular/continuous control problems in reinsurance. Part II: Numerical aspects. Applied Mathematics and Computation, 473(??): ??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001279.

GOREAC:2022:LTD

[GLX22] Dan Goreac, Juan Li, and Boxiang Xu. Linearisation techniques and the dual algorithm for a class of mixed singular/continuous control problems in reinsurance. Part I: Theoretical aspects. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003952.

Gao:2024:ISS

[GLY24] Lijun Gao, Zihan Liu, and Suo Yang. Input-to-state stability in probability of constrained impulsive switched delayed systems with unstable subsystems. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002819.

Ghorbani:2021:EEE

[GLZA21] Modjtaba Ghorbani, Xueliang Li, Samaneh Zangi, and Najaf Amraei. On the eigenvalue and energy of extended adjacency matrix. Applied Mathematics and Computation, 397 (??):Article 125939, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308924.

Guan:2020:SLS

[GLZB20] Xue Guan, Wenjun Liu, Qin Zhou, and Anjan Biswas. Some lump solutions for a generalized (3 + 1)-dimensional Kadomtsev– Petviashvili equation. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307490.

Ghosh:2021:HEP

[GM21a] Abhijit Ghosh and Chittaranjan Mishra. Highly efficient parallel algorithms for solving the Bates PIDE for pricing options on a GPU. Applied Mathematics and Computation, 409 (??):Article 126411, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005002.

Gomez:2021:GSA

[GM21b] Miller Cerón Gómez and Eduardo Ibarguen Mondragon. Global stability analysis for a SEI model with nonlinear incidence rate and asymptomatic infectious state. Applied Mathematics and Computation, 402(??):Article 126130, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001788.

Gaspoz:2024:INP

[GM24] Sébastien Gaspoz and Riccardo W. Maffucci. Independence numbers of polyhedral graphs. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005180.

Ge:2020:BSL

[GMH^{+}20] Quanbo Ge, Jinyan Ma, Hongli He, Hong Li, and Guoqiang Zhang. A basic smart linear Kalman filter with online performance evaluation based on observable degree. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319305958.

Guedes:2022:ECD

[GMN22] Priscila F. S. Guedes, Eduardo M. A. M. Mendes, and Erivelton Nepomuceno. Effective computational discretization scheme for nonlinear dynamical systems. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002818.

Giacopelli:2020:GTD

[GMT20] G. Giacopelli, M. Migliore, and D. Tegolo. Graphtheoretical derivation of brain structural connectivity. Applied Mathematics and Computation, 377(??): Article 125150, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301193.

Ghosh:2020:ELR

[GN20a] S. K. Ghosh and C. Nahak. An extension of Lakzian–Rhoades results in the structure of ordered b-metric spaces via wtdistance with an application. Applied Mathematics and Computation, 378 (??):Article 125197, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301661.

Giorno:2020:CBD

[GN20b] Virginia Giorno and Amelia G.

Nobile. On a class of birth-death processes with time-varying intensity functions. Applied Mathematics and Computation, 379 (??):Article 125255, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302241.

Geist:2021:PNC

[GN21] Barbara Katharina Geist and Helmut Neufeld. Particlenumber conservation, causality and linear response in biomedical imaging. Applied Mathematics and Computation, 409(??):Article 126418, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005075.

Giorno:2022:SIE

[GN22] Virginia Giorno and Amelia G. Nobile. On some integral equations for the evaluation of first-passage-time densities of time-inhomogeneous birth-death processes. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000790.

Giorno:2023:TID

[GN23] Virginia Giorno and Amelia G. Nobile. On a timeinhomogeneous diffusion process with discontinuous drift. Applied Mathematics and Computation, 451 (??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001819.

Grace:2022:CNO

[GNA22a] Said R. Grace, Shekhar Singh Negi, and Syed Abbas. Corrigendum to “New oscillatory results for nonlinear delay dynamic equations with super-linear neutral term”: Applied Mathematics and Computation 412 (2022) 126576 Page 1–

12. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001436.

See [GNA22b].

Grace:2022:NOR

[GNA22b] Said R. Grace, Shekhar Singh Negi, and Syed Abbas. New oscillatory results for nonlinear delay dynamic equations with super-linear neutral term. Applied Mathematics and Computation, 412(??):Article 126576, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006603.

See corrigendum [GNA22a].

Gao:2022:QAI

[GNSZ22] Ming Gao, Yichun Niu, Li Sheng, and Donghua Zhou. Quantitative analysis of incipient fault detectability for time-varying stochastic systems based on weighted moving average approach. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200546X.

Goedgebeur:2020:SCR

[GNZ20] Jan Goedgebeur, Addie Neyt, and Carol T. Zamfirescu. Structural and computational results on platypus graphs. Applied Mathematics and Computation, 386(??):Article 125491, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304495.

Gokce:2022:DIB

[Gök22] Aytül Gökçe. A dynamic interplay between Allee effect and time delay in a mathematical model with weakening memory. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003800.

Ghosh:2020:MAR

[GOO20] M. Ghosh, S. Olaniyi, and O. S. Obabiyi. Mathematical analysis of reinfection and relapse in malaria dynamics. Applied Mathematics and Computation, 373 (??):Article 125044, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300138.

Gaburro:2023:HOE

[GÖRT23] Elena Gaburro, Philipp Öffner, Mario Ricchiuto, and Davide Torlo. High order entropy preserving ADER-DG schemes. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007160.

Guaca:2022:MNS

[GP22] Denis Cajas Guaca and Elaine Cristina Catapani Poletti. Modeling and numerical simulation of dissolved oxygen and biochemical oxygen demand concentrations with Holling type III kinetic relationships. Applied Mathematics and Computation, 415 (??):Article 126690, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007748.

Gazonas:2023:NIC

[GP23] George A. Gazonas and Brian M. Powers. A numerical investigation of crack behavior near a fixed boundary using singular integral equation and finite element methods. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004356.

Gao:2021:EDC

[GPH21a] Liyan Gao, Qiuhui Pan, and Mingfeng He. Effects of defensive cooperation strategy on the evolution of cooperation in social dilemma. Applied Mathematics and Computation, 399(??):Article 126047, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000953.

Gao:2021:EBD

[GPH21b] Liyan Gao, Qiuhui Pan, and Mingfeng He. Environmentalbased defensive promotes cooperation in the prisoner’s dilemma game. Applied Mathematics and Computation, 401(??):Article 126074, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001223.

Guzman-Pro:2024:OFP

[GPHC24] Santiago Guzmán-Pro and César Hernández-Cruz. Orientations without forbidden patterns on three vertices. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003734.

Guzman-Pro:2023:DHP

[GPHHC23] Santiago Guzmán-Pro, Pavol Hell, and César Hernández-Cruz. Describing hereditary properties by forbidden circular orderings. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006294.

Ghosh:2021:NTA

[GPRV21] P. Ghosh, J. Pateras, V. Rangachari, and A. Vaidya.

A network thermodynamic analysis of amyloid aggregation along competing pathways. Applied Mathematics and Computation, 393(??): Article 125778, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307311.

Grzelec:2024:NUE

[GPW24] Igor Grzelec, Monika Pilśniak, and Mariusz Woźniak. A note on uniquely embeddable 2-factors. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006744.

Guo:2021:POB

[GR21] Cui Guo and Hong Seo Ryoo. On Pareto-optimal Boolean logical patterns for numerical data. Applied Mathematics and Computation, 403(??):Article 126153, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321002010.

Ganie:2023:NDL

[GRD23] Hilal A. Ganie, Bilal Ahmad Rather, and Kinkar Chandra Das. On the normalized distance Laplacian eigenvalues of graphs. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006889.

Gouvea:2022:INI

[GRMQ22] Alessandra M. M. M.

Gouvêa, Nicolás Rubido, Elbert E. N. Macau, and Marcos G. Quiles. Importance of numerical implementation and clustering analysis in force-directed algorithms for accurate community detection. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003848.

Grzelak:2022:SGM

[Grz22] Lech A. Grzelak. Sparse grid method for highly efficient computation of exposures for xVA. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005203.

Gholami:2021:FTS

[GS21] Hadi Gholami and Mohammad Hossein Shafiei. Finitetime H_{∞} static and dynamic output feedback control for a class of switched nonlinear time-delay systems. Applied Mathematics and Computation, 389 (??):Article 125557, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305130.

Glendinning:2022:CBC

[GS22] P. A. Glendinning and D. J. W. Simpson. Chaos in the border-collision normal form: a computerassisted proof using induced maps and invariant expanding cones. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004313.

Golmisheh:2024:HOF

[GS24] Fatemeh Mahdavi Golmisheh and Saeed Shamaghdari. Heterogeneous optimal formation control of nonlinear multi-agent systems with unknown dynamics by safe reinforcement learning. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300471X.

Gu:2022:ETS

[GSL^{+}22] Yang Gu, Zhenxing Sun, Li-Wei Li, Ju H. Park, and Mouquan Shen. Eventtriggered security adaptive control of uncertain multiarea power systems with cyber attacks. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004180.

Gu:2020:FTC

[GSRL20] Yang Gu, Mouquan Shen, Yuesheng Ren, and Hongxia Liu. H_{∞} finite-time control of unknown uncertain systems with actuator failure. Applied Mathematics and Computation, 383 (??):Article 125375, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303398.

Gao:2024:PTA

[GSS24] Yu Gao, Wei Sun, and Shun-Feng Su. Prescribedtime adaptive output feedback tracking control for nonlinear systems with quantized input. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005398.

Gil:2022:CAE

[GST22] A. Gil, J. Segura, and N. M. Temme. Complete asymptotic expansions for the relativistic Fermi–Dirac integral. Applied Mathematics and Computation, 412 (??):Article 126618, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007025.

Guo:2020:SDE

[GSW^{+}20] Zun-Guang Guo, Gui-Quan Sun, Zhen Wang, Zhen Jin, Li Li, and Can Li. Spatial dynamics of an epidemic model with nonlocal infection. Applied Mathematics and Computation, 377 (??):Article 125158, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301272.

Gao:2023:ACU

[GSWL23] Lihong Gao, Zhibao Song, Zhen Wang, and Ping Li. Adaptive control for uncertain nonlinear systems with state-function constraints and input delay. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003697.

Galias:2022:SSE

[GT22a] Zbigniew Galias and Warwick Tucker. The Songling system has exactly four limit cycles. Applied Mathematics and Computation, 415(??):Article 126691, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100775X.

Gyurkovics:2022:REP

[GT22b] Éva Gyurkovics and Tibor Takács. Robust energyto-peak filter design for a class of unstable polytopic systems with a macroeconomic application. Applied Mathematics and Computation, 420(??):Article 126729, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008110.

Gomez:2022:LME

[GTK^{+}22] Ledys Llasmin Salazar Gomez, Soledad Torres, Jozef Kiseľák, Felix Fuders, Naoyuki Ishimura, Yasukazu Yoshizawa, and Milan Stehlík. Long memory estimation in a non-Gaussian bivariate process. Applied Mathematics and Computation, 420(??): Article 126871, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009541.

Golmankhaneh:2023:IVP

[GTSV23] Alireza Khalili Golmankhaneh, Inés Tejado, Hamdullah Sevli, and Juan E. Nápoles Valdés. On initial value problems of fractal delay equations. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323001492.

Ge:2021:INO

[GTWL21] Quanbo Ge, Shuaishuai Tang, Mengmeng Wang, and Zhenyu Lu. Improved nonlinear observable degree analysis using data fusion. Applied Mathematics and Computation, 392(??): Article 125613, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305683.

Gu:2020:CHT

[Gu20]  Longfei Gu.  The Clif-

ford Hilbert transform and Riemann–Hilbert problems associated to the Dirac operator in Lebesgue spaces. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306782.

Guo:2022:RPC

[Guo22] Fenglong Guo. Ruin probability of a continuous-time model with dependence between insurance and financial risks caused by systematic factors. Applied Mathematics and Computation, 413(??):Article 126634, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007189.

Guo:2023:SMC

[Guo23]  Jia Guo.  On strong

Menger connectivity of (n, k)-bubble-sort networks. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004010.

Ghorbani:2023:GSD

[GV23] Modjtaba Ghorbani and Zahra Vaziri. Graphs with small distance-based complexities. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003570.

Ghorbani:2024:SWC

[GV24] Modjtaba Ghorbani and Zahra Vaziri. On the Szeged and Wiener complexities in graphs. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000043.

Gao:2020:ECI

[GW20a] Yunshu Gao and Xue Wang.

The edge condition for independent cycles with chords in bipartite graphs. Applied Mathematics and Computation, 377(??):Article 125163, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301326.

Guan:2020:CEP

[GW20b] Junbiao Guan and Kaihua Wang. Cooperative evolution in pedestrian room evacuation considering different individual behaviors. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308574.

Geng:2021:RKF

[GW21a] F. Z. Geng and X. Y.

Wu. Reproducing kernel function-based Filon and Levin methods for solving highly oscillatory integral. Applied Mathematics and Computation, 397 (??):Article 125980, May 15, 2021. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100028X.

Grow:2021:BSS

[GW21b] David Grow and Nick Wintz. Bilinear state systems on an unbounded time scale. Applied Mathematics and Computation, 397 (??):Article 125917, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308705.

Glau:2022:DPP

[GW22] Kathrin Glau and Linus Wunderlich. The deep parametric PDE method and applications to option pricing. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004295.

Grzelec:2023:DML

[GW23a] Igor Grzelec and Mariusz Woźniak. On decomposing multigraphs into locally irregular submultigraphs. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002187.

Guo:2023:DSM

[GW23b] Xinchen Guo and Guoliang Wei. Distributed sliding mode consensus control for multiple discretetime Euler–Lagrange systems. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000474.

Guo:2024:FTS

[GW24] Mengyu Guo and Peiguang Wang. Finite-time stability of non-instantaneous impulsive systems with double state-dependent delays. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002856.

Guo:2022:DVU

[GWH22] Furi Guo, JinRong Wang, and Jiangfeng Han. Dynamic viscoelastic unilateral constrained contact problems with thermal effects. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001205.

Guo:2024:AEB

[GWH24] Jia Guo, Haifeng Wang, and Chenping Hou. An adaptive energy-based sequential method for training PINNs to solve gradient flow equations. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003515.

Gao:2023:SAH

[GWSC23] Lijun Gao, Zhenyue Wang, Tao Sun, and Zhengbao Cao. Stability analysis for hybrid deterministic system under delay-dependent impulses uniting properties of edges. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008724.

Guo:2023:DCM

[GWW23] Songlin Guo, Wei Wang, and Chuanming Wang. Disproof of a conjecture on the minimum Wiener index of signed trees. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006518.

Gu:2020:SFO

[GWY20] Yajuan Gu, Hu Wang, and Yongguang Yu. Synchronization for fractionalorder discrete-time neural networks with time delays. Applied Mathematics and Computation, 372(??): Article 124995, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309877.

Gu:2020:EGD

[GWZ^{+}20] Cuiling Gu, Xianjia Wang, Jinhua Zhao, Rui Ding, and Qilong He. Evolutionary game dynamics of Moran process with fuzzy payoffs and its application. Applied Mathematics and Computation, 378 (??):Article 125227, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030196X.

Gao:2021:CDB

[GWZ^{+}21] Hongyu Gao, Juan Wang, Fan Zhang, Xiaopeng Li, and Chengyi Xia. Cooperation dynamics based on reputation in the mixed population with two species of strategists. Applied Mathematics and Computation, 410(??):Article 126433, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005221.

Gao:2023:FFT

[GWZZ23] Shuo Gao, Guoguang Wen, Xiaoqin Zhai, and Peng Zheng. Finite-/fixed-time bipartite consensus for first-order multi-agent systems via impulsive control. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008086.

Guo:2023:OBS

[GX23] Runan Guo and Shengyuan Xu. Observer-based sliding mode synchronization control of complex-valued neural networks with inertial term and mixed time-varying delays. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008293.

Geng:2024:SSF

[GXZ^{+}24] Jiahui Geng, Ning Xu, Xudong Zhao, Ze Tang, and Jiancheng Zhang. Simultaneous state and fault estimation: a prescribedtime unknown input observer approach. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003837.

Guo:2020:GTB

[GXZZ20] Beibei Guo, Yu Xiao, Chiping Zhang, and Yong Zhao. Graph theory-based adaptive intermittent synchronization for stochastic delayed complex networks with semi-Markov jump. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307313.

Gao:2022:SDD

[GY22a] Rui Gao and Guang-Hong Yang. Sampled-data distributed state estimation with multiple transmission channels under denial-of-service attacks. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003034.

Ghosh:2022:SGS

[GY22b] Sudeshna Ghosh and Pooja Yadav. Study of gravitational settling of single semi-torus shaped particle using immersed boundary method. Applied Mathematics and Computation, 413(??):Article 126643, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100727X.

Gao:2022:EDC

[GYC^{+}22] Fengqiang Gao, Yuyue Yan, Zhihao Chen, Linxiao Zheng, and Huan Ren. Effect of density control in partially observable asymmetric-exit evacuation under guidance: Strategic suggestion under time delay. Applied Mathematics and Computation, 418 (??):Article 126838, April 1, 2022. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009218.

Gong:2024:AHA

[GYC24] Yujie Gong, Guangwei Yuan, and Xia Cui. Analysis on a high accuracy fully implicit solution for strong nonlinear diffusion problem — convergence, stability, and uniqueness. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006689.

Gao:2021:SDL

[GYM21] You Gao, Yun-Fei Yao, and He Ma. On the stopping distance of LDPC codes based on symplectic space over finite fields. Applied Mathematics and Computation, 391(??):Article 125625, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305798.

Gao:2020:SAC

[GZ20a] Qingwu Gao and Jun Zhuang. Stability analysis and control strategies for worm attack in mobile networks via a VEIQS propagation model. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319305764.

Gu:2020:PCS

[GZ20b] Da-Ke Gu and Da-Wei Zhang. Parametric control to second-order linear time-varying systems based on dynamic compensator and multi-objective optimization. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306733.

Guo:2020:ADS

[GZ20c] Haiyan Guo and Bo Zhou. On adjacency-distance spectral radius and spread of graphs. Applied Mathematics and Computation, 369 (??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319308112.

Gao:2023:EAA

[GZCS23] Shanshan Gao, Shenggui Zhang, Xinzhuang Chen, and Xiaodi Song. Effects of adding arcs on the consensus convergence rate of leader–follower multiagent systems. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002552.

Guo:2021:TDF

[GZG^{+}21] Lin Guo, Xi-Le Zhao, Xian-Ming Gu, Yong-Liang Zhao, Yu-Bang Zheng, and Ting-Zhu Huang. Threedimensional fractional total variation regularized tensor optimized model for image deblurring. Applied Mathematics and Computation, 404(??):Article 126224, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003143.

Gao:2022:ECR

[GZH22] Mengyao Gao, Xuelin Zhang, and Guodong Han. The exponential convergence rate of Kaczmarz’s algorithm and an acceleration strategy for ART. Applied Mathematics and Computation, 420(??):Article 126885, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009681.

Gao:2023:OTA

[GZL^{+}23] Yuxuan Gao, Ying Zhao, Yu Liu, Zhuo-Rui Pan, and Rui Wang. Output tracking anti-disturbance control for turbofan systems. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003089.

Guo:2024:OMB

[GZLY24] Xiyue Guo, Huaguang Zhang, Xin Liu, and Xiaohui Yue. Output-mask-based adaptive NN control for stochastic timedelayed multi-agent systems with a unified eventtriggered approach. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001978.

Guo:2024:TSM

[GZLZ24] Wenxiu Guo, Hua Zheng, Xiaoping Lu, and Yongxiong Zhang. On the twostage multisplitting iteration methods for linear complementarity problems. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002133.

Gao:2020:SSM

[GZQ^{+}20] Meng Gao, Lihua Zhang, Wenhai Qi, Jinde Cao, Jun Cheng, Yonggui Kao, Yunliang Wei, and Xiaoyu Yan. SMC for semi-Markov jump T –S fuzzy systems with time delay. Applied Mathematics and Computation, 374(??):Article 125001, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309932.

Gouia-Zarrad:2021:NIU

[GZRM21] Rim Gouia-Zarrad, Souvik Roy, and Sunghwan Moon. Numerical inversion and uniqueness of a spherical radon transform restricted with a fixed angular span. Applied Mathematics and Computation, 408 (??):Article 126338, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004276.

Guo:2023:DCH

[GZWX23] Shiyu Guo, Xudong Zhao, Huanqing Wang, and Ning Xu. Distributed consensus of heterogeneous switched nonlinear multiagent systems with input quantization and DoS attacks. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002965.

Gao:2020:DOI

[GZX20] Rui Gao, Ding Zhai, and Xiangpeng Xie. On the design of output information-based sliding mode controllers for switched descriptor systems: Linear sliding variable approach. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306721.

Guo:2021:SEC

[GZZ21] Yirong Guo, Xia Zhang, and Xinmiao Zhang. Strong edge-colorings of planar graphs with small girth. Applied Mathematics and Computation, 394(??):Article 125796, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307499.

Hajishafieiha:2020:NCP

[HA20] J. Hajishafieiha and S. Abbasbandy. A new class of polynomial functions for approximate solution of generalized Benjamin–Bona– Mahony–Burgers (gBBMB) equations. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930757X.

Huseynov:2021:EAS

[HAFM21] Ismail T. Huseynov, Arzu Ahmadova, Arran Fernandez, and Nazim I. Mahmudov. Explicit analytical solutions of incommensurate fractional differential equation systems. Applied Mathematics and Computation, 390 (??):Article 125590, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305452.

Henning:2020:TCD

[HAK20] Michael A. Henning, Nawarat Ananchuen, and Pawaton Kaemawichanurat. Traceability of connected domination critical graphs. Applied Mathematics and Computation, 386(??):Article 125455, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320304161.

Han:2020:SII

[Han20] Jiayu Han. Shifted inverse iteration based multigrid methods for the quad-curl eigenvalue problem. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307623.

Hussain:2023:PAO

[HANP23] Sultan Hussain, Hifsa Arif, Muhammad Noorullah, and Athanasios A. Pantelous. Pricing American options under Azzalini Ito–McKean skew Brownian motions. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002096.

Hartung:2022:NAD

[Har22] Ferenc Hartung. On numerical approximation of a delay differential equation with impulsive self-support condition. Applied Mathematics and Computation, 418 (??):Article 126818, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009012.

Harrison:2024:NSF

[Har24] Peter G. Harrison. On the numerical solution of functional equations with application to response time distributions. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001097.

Hassan:2020:EGA

[Has20] Anthony R. Hassan. The entropy generation analysis of a reactive hydromagnetic couple stress fluid flow through a saturated porous channel. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308355.

Huang:2020:MSS

[HB20] Yao Huang and Haibo Bao. Master-slave synchronization of complex-valued delayed chaotic Lur’e systems with sampled-data control. Applied Mathematics and Computation, 379(??):Article 125261, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302307.

Huang:2024:OSR

[HB24] Zhuoyuan Huang and Haibo Bao. Output synchronization of reaction-diffusion neural networks with multiple output couplings via generalized intermittent control. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002832.

Heidarkhani:2021:CPA

[HBCA21] Shapour Heidarkhani, Martin Bohner, Giuseppe Caristi, and Farahnaz Ayazi. A critical point approach for a second-order dynamic Sturm–Liouville boundary value problem with p-Laplacian. Applied Mathematics and Computation, 409(??):Article 125521, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304793.

Hu:2020:SAG

[HBWQ20] Gang Hu, Cuicui Bo, Guo Wei, and Xinqiang Qin. Shape-adjustable generalized Bézier surfaces: Construction and it is geometric continuity conditions. Applied Mathematics and Computation, 378 (??):Article 125215, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301843.

Hajiketabi:2020:NIB

[HC20] M. Hajiketabi and F. Casas.

Numerical integrators based on the Magnus expansion for nonlinear dynamical systems. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308367.

Huang:2021:SWA

[HC21] Cong Huang and Li Li Chen. A simple WENO– AO method for solving hyperbolic conservation laws. Applied Mathematics and Computation, 395(??):Article 125856, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308092.

He:2024:PBG

[HC24] Jialu He and Lei Cui. The persistence-based game transition resolves the social dilemma. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002534.

Hou:2021:DFC

[HCBY21] Rui Hou, Lizhi Cui, Xuhui Bu, and Junqi Yang. Distributed formation control for multiple non-holonomic wheeled mobile robots with velocity constraint by using improved data-driven iterative learning. Applied Mathematics and Computation, 395(??):Article 125829, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307827.

Hu:2024:ODS

[HCH24] Haibo Hu, Wenhao Chen, and Yixuan Hu. Opinion dynamics in social networks under the influence of mass media. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004375.

Hernandez-Cruz:2021:PSQ

[HCJC21] César Hernández-Cruz and Manuel Alejandro Juárez-Camacho. Pancyclicity in strong k-quasi-transitive digraphs of large diameter. Applied Mathematics and Computation, 408 (??):Article 126319, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004082.

Hu:2023:NRC

[HCL^{+}23] Yue Hu, Chenxiao Cai, SeungHoon Lee, YongGwon Lee, and Oh-Min Kwon.

New results on H_{∞} control for interval type-2 fuzzy singularly perturbed systems with fading channel: the weighted tryonce-discard protocol case. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300108X.

Huang:2021:MCD

[HCM21] He Huang, Yahong Chen, and Yefeng Ma. Modeling the competitive diffusions of rumor and knowledge and the impacts on epidemic spreading. Applied Mathematics and Computation, 388(??):Article 125536, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304926.

He:2020:NSI

[HCS20] Baiying He, Liangyun Chen, and Bing Sun. New super integrable hierarchies associated with osp(2|2) and spo(2|2) and their applications. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308598.

Hu:2023:SPL

[HCS23] Mingying Hu, Haiyan Chen, and Wenwen Sun. Sum of powers of the Laplacian eigenvalues and the Kirchhoff index of a graph. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000528.

Hao:2024:RDG

[HCSJ24] Guoliang Hao, Xiaodan Chen, Seyed Mahmoud Sheikholeslami, and Haining Jiang. Roman [1, 2]- domination of graphs. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006896.

Hao:2023:TOW

[HCTS23] Tianchu Hao, Yaming Chen, Lingyan Tang, and Songhe Song. A thirdorder weighted nonlinear scheme for hyperbolic conservation laws with inverse Lax–Wendroff boundary treatment. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007652.

Hu:2024:APT

[HCX^{+}24] Yinlong Hu, Xuhao Cai, Jin Xu, Shang Shi, Yonghui Sun, and Haiping Du. Analytical parameter tuning for a class of extended disturbance observers and sliding mode control. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006963.

Huang:2021:ISD

[HCY21] He Huang, Yahong Chen, and Zhijun Yan. Impacts of social distancing on the spread of infectious diseases with asymptomatic infection: a mathematical model. Applied Mathematics and Computation, 398 (??):Article 125983, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100031X.

Huang:2022:PTT

[HCY22] Guang-Xin Huang, Qi-Xing Chen, and Feng Yin. Preconditioned TBi-COR and TCORS algorithms for solving the Sylvester tensor equation. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000704.

Huang:2020:RDC

[HD20] Xin Huang and Jiuxiang Dong. A robust dynamic compensation approach for cyber-physical systems against multiple types of actuator attacks. Applied Mathematics and Computation, 380(??): Article 125284, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302538.

Hosny:2022:NQD

[HD22a] Khalid M. Hosny and Mohamed M. Darwish. Novel quaternion discrete shifted Gegenbauer moments of fractional-orders for color image analysis. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000121.

Hussein:2022:SNH

[HD22b] S. O. Hussein and Taysir E. Dyhoum. Solutions for nonhomogeneous wave equations subject to unusual and Neumann boundary conditions. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003599.

Han:2023:AFG

[HD23] Xiaohui Han and Jianping Dong. Applications of fractional gradient descent method with adaptive momentum in BP neural networks. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001133.

Halder:2021:QAJ

[HDP^{+}21] Kaushik Halder, Saptarshi Das, Deepak Kumar Panda, Sourav Das, and Amitava Gupta. QoS aware joint observer and networked PI/ PID controller design using LMIs under specified rate of packet dropouts. Applied Mathematics and Computation, 401(??):Article 126125, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001739.

Huang:2023:FBM

[HDQP23] Wenting Huang, Xiaofang Duan, Lijuan Qin, and Junpyo Park. Fitnessbased mobility enhances the maintenance of biodiversity in the spatial system of cyclic competition. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003041.

Huang:2022:TNL

[HDZ22] Xueyi Huang, Kinkar Chandra Das, and Shunlai Zhu. Toughness and normalized Laplacian eigenvalues of graphs. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200159X.

He:2023:NMB

[He23] Yunhui He. Novel massbased multigrid relaxation schemes for the Stokes equations. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007354.

Hong:2021:APD

[HGD^{+}21] Lijun Hong, Yini Geng, Chunpeng Du, Chen Shen, and Lei Shi. Average payoffdriven or imitation? A new evidence from evolutionary game theory in finite populations. Applied Mathematics and Computation, 394 (??):Article 125784, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307372.

Hussain:2021:CST

[HGJ21] Iqtadar Hussain, Dibakar Ghosh, and Sajad Jafari. Chimera states in a thermosensitive FitzHugh– Nagumo neuronal network. Applied Mathematics and Computation, 410 (??):Article 126461, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005506.

Hosseinpour:2022:HHO

[HGKD22] Mahsa Hosseinpour, Sehraneh Ghaemi, Sohrab Khanmohammadi, and Sabalan Daneshvar. A hybrid high-order type-2 FCM improved random forest classification method for breast cancer risk assessment. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001242.

Huang:2024:NSO

[HGZ24] Qin Huang, Rongfang Gong, and Ye Zhang. A new second-order dynamical method for solving linear inverse problems in Hilbert spaces. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001140.

Hasani:2020:BOF

[HH20] Ali Hasani and Seyed Mohammad Hassan Hosseini. A bi-objective flexible flow shop scheduling problem with machine-dependent processing stages: Tradeoff between production costs and energy consumption. Applied Mathematics and Computation, 386 (??):Article 125533, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304896.

Huang:2022:BTI

[HH22] Yi-Bo Huang and Yong He. Bessel-type inequality in semi-inner-product spaces and its application to stability analysis of discretetime systems with distributed delays. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002211.

Han:2023:ILS

[HH23a] Dun Han and Youxin He. The impact of labor subsidy, taxation and corruption on individual behavior. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004162.

Holub:2023:PCD

[HH23b] Premysl Holub and Jakub Hofman. On S-packing colourings of distance graphs D(1, t) and D(1, 2, t). Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000243.

Hao:2024:IDC

[HH24] Weijuan Hao and Yuhan Hu. The implications of deep cooperation strategy for the evolution of cooperation in social dilemmas. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400050X.

Han:2023:LTF

[HHGZ23] Yuxin Han, Xin Huang, Wei Gu, and Bolong Zheng. Linearized transformed L1 finite element methods for semi-linear time-fractional parabolic problems. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004113.

Hartman:2021:CSD

[HHR21] David Hartman, Milan Hladík, and David Ríha. Computing the spectral decomposition of interval matrices and a study on interval matrix powers. Applied Mathematics and Computation, 403 (??):Article 126174, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002642.

Hong:2023:IDC

[HHW23] Xiao Hong, Yuexing Han, and Bing Wang. Impacts of detection and contact tracing on the epidemic spread in time-varying networks. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006749.

Huang:2020:PSC

[HHY20] Yanli Huang, Jie Hou, and Erfu Yang. Passivity and synchronization of coupled reaction-diffusion complexvalued memristive neural networks. Applied Mathematics and Computation, 379(??):Article 125271, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030240X.

Hu:2020:FTS

[HHZZ20] Taotao Hu, Zheng He, Xiaojun Zhang, and Shouming Zhong. Finite-time stability for fractional-order complex-valued neural networks with time delay. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307076.

Hu:2024:RIP

[HJJL24] Qi Hu, Tao Jin, Yulian Jiang, and Xingwen Liu. Reputation incentives with public supervision promote cooperation in evolutionary games. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006148.

Hou:2023:LPG

[HJL^{+}23] Jianfeng Hou, Yindong Jin, Heng Li, Lianying Miao, and Qian Zhao. On L(p, q) -labelling of planar graphs without cycles of length four. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300053X.

Huang:2024:GSM

[HJN^{+}24] Shen C. Huang, Adam Johnson, Monika Neda, Jorge Reyes, and Hossein Tehrani. A generalization of the Smagorinsky model. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000171.

Huang:2023:AET

[HJW^{+}23] Jingyang Huang, Wei Jia, Tongxin Wan, Shuyi Xiao, Liqi Wang, and Jiuxiang Dong. Adaptive eventtriggered fault-tolerant consensus of linear heterogeneous multiagent systems via hierarchical approach. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000784.

Huang:2024:TCT

[HJY24] Xiaoqian Huang, Yanqun Jiang, and Huanhuan Yang. Two classes of third-order weighted compact nonlinear schemes for Hamilton–Jacobi equations. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000262.

Henning:2021:CDC

[HK21] Michael A. Henning and Pawaton Kaemawichanurat. Connected domination critical graphs with a block having maximum number of cut vertices. Applied Mathematics and Computation, 406(??):Article 126248, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003386.

Hu:2024:OFC

[HKCK24] Yue Hu, Oh-Min Kwon, Chenxiao Cai, and Yeong-Jae Kim. Output feedback H_{∞} control for discrete time singularly perturbed systems with Markov lossy network: the round-robin-like protocol case. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005076.

Han:2022:RSD

[HKKL22] S. Y. Han, S. K. Kommuri, O. M. Kwon, and S. M. Lee. Regional sampled-data synchronization of chaotic neural networks using piecewisecontinuous delay dependent Lyapunov functional. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000807.

Harshavarthini:2022:UDE

[HKL22a] S. Harshavarthini, O. M. Kwon, and S. M. Lee. Uncertainty and disturbance estimator-based resilient tracking control design for fuzzy semi-Markovian jump systems. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002077.

Hoang:2022:NTD

[HKL22b] Thi-Thao-Phuong Hoang, Hemanta Kunwar, and Hyesuk Lee. Nonconforming time discretization based on Robin transmission conditions for the Stokes–Darcy system. Applied Mathematics and Computation, 413 (??):Article 126602, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100686X.

Hnatushenko:2021:SIS

[HKU21] Volodymyr V. Hnatushenko, Peter I. Kogut, and Mykola V. Uvarov. On satellite image segmentation via piecewise constant approximation of selective smoothed target mapping. Applied Mathematics and Computation, 389(??):Article 125615, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305695.

Huang:2022:SEN

[HL22] Min Huang and Guo Luo.

A simple and efficient numerical method for pricing discretely monitored early-exercise options. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000716.

Hu:2023:AFD

[HL23a] Yifan Hu and Wenhui Liu. Adaptive fuzzy dynamic surface control for nonstrict-feedback nonlinear state constrained systems with input dead-zone via event-triggered sampling. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001546.

Huo:2023:GBH

[HL23b] Zhixin Huo and Jiequan Li. A GRP-based high resolution ghost fluid method for compressible multimedium fluid flows I: Onedimensional case. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200580X.

Harshavarthini:2024:TPT

[HL24a] S. Harshavarthini and S. M.

Lee. Truncated predictive tracking control design for semi-Markovian jump systems with time-varying input delays. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001589.

Hou:2024:MSS

[HL24b] Huanyang Hou and Xiaolin Li. A meshless superconvergent stabilized collocation method for linear and nonlinear elliptic problems with accuracy analysis. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002625.

Hou:2023:IDS

[HLD^{+}23] Yunxiang Hou, Yikang Lu, Yuting Dong, Libin Jin, and Lei Shi. Impact of different social attitudes on epidemic spreading in activity-driven networks. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300019X.

Huang:2022:FTD

[HLF22] Zheng Huang, Weijun Liu, and Lihua Feng. sflag-transitive designs for s ≥ 4. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003162.

Hu:2022:ODF

[HLKC22] Jun Hu, Jiaxing Li, Yonggui Kao, and Dongyan Chen. Optimal distributed filtering for nonlinear saturated systems with random access protocol and missing measurements: the uncertain probabilities case. Applied Mathematics and Computation, 418(??):Article 126844, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009279.

He:2021:LSN

[HLL^{+}21] Zhilong He, Chuandong Li, Yi Li, Zhengran Cao, and Xiaoyu Zhang. Local synchronization of nonlinear dynamical networks with hybrid impulsive saturation control inputs. Applied Mathematics and Computation, 410(??):Article 126452, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005415.

Hu:2022:AFT

[HLL22] Yifan Hu, Wenhui Liu, and Guobao Liu. Adaptive finite-time event-triggered control for uncertain nonlinearly parameterized systems with unknown control direction and actuator failures. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005252.

Hu:2023:ARP

[HLLH23] Wenjie Hu, Yibo Li, Huiqing Liu, and Xiaolan Hu. Anti-Ramsey problems in the Mycielskian of a cycle. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300323004368.

Hou:2022:ATC

[HLLS22] Linlin Hou, Yao Li, Wende Luo, and Haibin Sun. Adaptive tracking control of switched cyberphysical systems with cyberattacks. Applied Mathematics and Computation, 415(??):Article 126721, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321008055.

Han:2021:GSX

[HLLW21] Huan Han, Fanghui Li, Tingjian Luo, and Zhengping Wang. Ground state for the X-ray free electron laser Schrödinger equation with harmonic potential. Applied Mathematics and Computation, 401 (??):Article 126113, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001612.

Hwang:2020:DDA

[HLLY20] Yoongu Hwang, Jangwoon Lee, Jeehyun Lee, and Myoungho Yoon. A domain decomposition algorithm for optimal control problems governed by elliptic PDEs with random inputs. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306666.

Holub:2021:SEC

[HLM^{+}21] Premysl Holub, Borut Luzar, Erika Mihaliková, Martina Mockovciaková, and Roman Soták. Star edge-coloring of square grids. Applied Mathematics and Computation, 392(??): Article 125741, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306949.

Hu:2023:BGF

[HLM23] Yaozhong Hu, Juan Li, and Chao Mi. BSDEs generated by fractional space-time noise and related SPDEs. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001480.

Hoang:2024:OHC

[HLP24] Nam-Dũng Hoang, Nguyen Kieu Linh, and Hoang Xuan Phu. Octagonal and hexadecagonal cut algorithms for finding the convex hull of finite sets with linear time complexity. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003928.

Hernandez-Lorenzana:2022:KCD

[HLSL22] Felipe Hernández-Lorenzana and Rocío Sánchez-López. H-kernels in H-colored digraphs without (ξ1, ξ, ξ2)- H-subdivisions of C⃗_{3}. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001734.

Huang:2020:LCS

[HLW^{+}20] Xin Huang, Yamin Liu, Yang Wang, Jianping Zhou, Muyun Fang, and Zhen Wang. L_{2}–L_{∞} consensus of stochastic delayed multi-agent systems with ADT switching interaction topologies. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307921.

Han:2021:ADD

[HLW^{+}21] Jian Han, Xiuhua Liu, Xinjiang Wei, Huifeng Zhang, and Xin Hu. Adjustable dimension descriptor observer based fault estimation of nonlinear system with unknown input. Applied Mathematics and Computation, 396(??): Article 125899, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308523.

Huang:2020:NGT

[HLX20] Xing Huang, Huiqiu Lin, and Jie Xue. The Nordhaus– Gaddum type inequalities of A_{α}-matrix. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307088.

He:2021:NTI

[HLX21] Jun He, Yanmin Liu, and Guangjun Xu. New Stype inclusion theorems for the M-eigenvalues of a 4th-order partially symmetric tensor with applications. Applied Mathematics and Computation, 398 (??):Article 125992, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000400.

Hao:2020:EES

[HLZ20] Zhaopeng Hao, Guang Lin, and Zhongqiang Zhang. Error estimates of a spectral Petrov–Galerkin method for two-sided fractional reaction-diffusion equations. Applied Mathematics and Computation, 374(??):Article 125045, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030014X.

Hu:2021:FSM

[HLZ21] Zhi Hu, Dongdai Lin, and Chang-An Zhao. Fast scalar multiplication of degenerate divisors for hyperelliptic curve cryptosystems. Applied Mathematics and Computation, 404(??): Article 126239, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003295.

Hu:2022:RHP

[HLZ22] Beibei Hu, Ji Lin, and Ling Zhang. On the Riemann– Hilbert problem for the integrable three-coupled Hirota system with a 4 × 4 matrix Lax pair. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002764.

Han:2020:ESC

[HM20a] Miaomiao Han and Zhengke Miao. Eulerian subgraphs and S-connectivity of graphs. Applied Mathematics and Computation, 382(??):Article 125323, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302897.

Huang:2020:GLS

[HM20b] Baohua Huang and Changfeng Ma. Global least squares methods based on tensor form to solve a class of generalized Sylvester tensor equations. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308847.

Hernandez-Maldonado:2024:FPC

[HMEDVR24] Victor Miguel Hernández-Maldonado, Arturo Erdely, Martín Díaz-Viera, and Leonardo Rios. Fast procedure to compute empirical and Bernstein copulas. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002881.

Huska:2023:GTT

[HMR23] Martin Huska, Serena Morigi, and Giuseppe Recupero. Geometric texture transfer via local geometric descriptors. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300200X.

Hoto:2020:PAS

[HMS20] R. S. V. Hoto, L. C. Matioli, and P. S. M. Santos. A penalty algorithm for solving convex separable knapsack problems. Applied Mathematics and Computation, 387(??):Article 124855, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308471.

Huo:2020:ETA

[HMZZ20] Xin Huo, Li Ma, Xudong Zhao, and Guangdeng Zong. Event-triggered adaptive fuzzy output feedback control of MIMO switched nonlinear systems with average dwell time. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306575.

Hojjat:2020:NCS

[Hoj20] Mohammad Hojjat. Nanofluids as coolant in a shell and tube heat exchanger: ANN modeling and multiobjective optimization. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307027.

Huang:2022:HES

[HOM22] Chun-Sung Huang, John G.

O’Hara, and Sure Mataramvura. Highly efficient Shannon wavelet-based pricing of power options under the double exponential jump framework with stochastic jump intensity and volatility. Applied Mathematics and Computation, 414 (??):Article 126669, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007530.

Honarbakhsh:2024:CFE

[Hon24] B. Honarbakhsh. Closedform expressions for some of the integrals related to the method of Kobayashi potential. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004314.

Hu:2020:LDE

[HP20] Sinan Hu and Yuejian Peng. Lagrangian densities of enlargements of matchings in hypergraphs. Applied Mathematics and Computation, 374(??):Article 125068, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300370.

He:2023:FMP

[HP23] Jin-Man He and Li-Jun Pei. Function matrix projection synchronization for the multi-time delayed fractional order memristorbased neural networks with parameter uncertainty. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002795.

Huang:2024:CDK

[HPC24] He Huang, Jialin Pan, and Yahong Chen. The competitive diffusion of knowledge and rumor in a multiplex network: a mathematical model. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001917.

Herrera:2021:IGD

[HPL21] Luis H. Herrera and Pablo Pérez-Lantero. On the intersection graph of the disks with diameters the sides of a convex ngon. Applied Mathematics and Computation, 411 (??):Article 126472, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005610.

Henning:2022:BPD

[HPT22] Michael A. Henning, Monika Pilśniak, and Elzbieta Tumidajewicz. Bounds on the paired domination number of graphs with minimum degree at least three. Applied Mathematics and Computation, 417(??):Article 126782, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100864X.

Hu:2022:SAG

[HQ22] Limi Hu and Xiaoling Qiu. Stability analysis of game models with fixed and stochastic delays. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005471.

He:2021:HBA

[HQK21] Hangfeng He, Wenhai Qi, and Yonggui Kao. HMMbased adaptive attackresilient control for Markov jump system and application to an aircraft model. Applied Mathematics and Computation, 392 (??):Article 125668, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306214.

Han:2024:PTB

[HQS^{+}24] Chenglin Han, Kaiyu Qin, Mengji Shi, Yan Liu, Weihao Li, and Boxian Lin. Prescribed time bipartite output consensus tracking for heterogeneous multi-agent systems with external disturbances. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000225.

Hao:2024:TDC

[HQZC24] Rong-Xia Hao, Xiao-Wen Qin, Hui Zhang, and Jou-Ming Chang. Two-disjoint-cycle-cover pancyclicity of data center networks. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001887.

Hassan:2020:OFR

[HR20a] Bryar A. Hassan and Tarik A. Rashid. Operational framework for recent advances in backtracking search optimisation algorithm: a systematic review and performance evaluation. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309117.

Hauenstein:2020:RMA

[HR20b] Jonathan D. Hauenstein and Margaret H. Regan. Real monodromy action. Applied Mathematics and Computation, 373 (??):Article 124983, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309750.

Hiremath:2023:AMA

[HR23] Varun Hiremath and Jose E. Roman. Acoustic modal analysis with heat release fluctuations using nonlinear eigensolvers. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004186.

Hussain:2020:NAS

[HRA^{+}20] Muntazir Hussain, Muhammad Rehan, Shakeel Ahmed, Tanveer Abbas, and Muhammad Tufail. A novel approach for static antiwindup compensation of one-sided Lipschitz systems under input saturation. Applied Mathematics and Computation, 380(??): Article 125229, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301983.

Hammadi:2021:DDR

[HREB21] Youssef Hammadi, David Ryckelynck, and Amin El-Bakkali. Data-driven reduced bond graph for nonlinear multiphysics dynamic systems. Applied Mathematics and Computation, 409 (??):Article 126359, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004483.

He:2021:ADS

[HRLH21] Miao He, Taotao Rong, Junmin Li, and Chao He. Adaptive dynamic surface full state constraints control for stochastic Markov jump systems based on eventtriggered strategy. Applied Mathematics and Computation, 392(??):Article 125563, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305191.

Huang:2023:ECP

[HRZ^{+}23] Yongchao Huang, Tianyu Ren, Junjun Zheng, Wenyi Liu, and Mengshu Zhang. Evolution of cooperation in public goods games with dynamic resource allocation: a fairness preference perspective. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000139.

Heng:2023:ACH

[HS23a] Pei Heng and Yi Sun. Algorithms for convex hull finding in undirected graphical models. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323000218.

Huynh:2023:CHP

[HS23b] Viet Q. H. Huynh and Hiroshi Suito. Communicationhiding pipelined BiCGSafe methods for solving large linear systems. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000371.

Huang:2024:GCE

[HSL^{+}24] Tao Huang, Yiyu Shao, Liwei Li, Yajuan Liu, and Mouquan Shen. Guaranteed cost event-triggered H_{∞} control of uncertain linear system via output disturbance observer. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001243.

Han:2020:HIR

[HSLS20] Dun Han, Qi Shao, Dandan Li, and Mei Sun. How the individuals’ risk aversion affect the epidemic spreading. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308860.

Hejazi:2021:ANL

[HSM21] S. Reza Hejazi, Elaheh Saberi, and Fatemeh Mohammadizadeh. Anisotropic non-linear time-fractional diffusion equation with a source term: Classification via Lie point symmetries, analytic solutions and numerical simulation. Applied Mathematics and Computation, 391 (??):Article 125652, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306068.

Huang:2024:SAD

[HSY24] Yi-Bo Huang, Zhihuan Song, and Wei Yu. Stability analysis of discretetime systems with arbitrary delay kernels based on kernel-related summation inequality and model transformation. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002121.

Hu:2020:RSE

[HSZ20] Binxin Hu, Qiankun Song, and Zhenjiang Zhao. Robust state estimation for fractional-order complexvalued delayed neural networks with interval parameter uncertainties: LMI approach. Applied Mathematics and Computation, 373 (??):Article 125033, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300023.

Herrero:2020:SPO

[HT20] A. Herrero and N. Thome.

Sharp partial order and linear autonomous systems. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307283.

Hidalgo:2023:NAP

[HT23] Arturo Hidalgo and Lourdes Tello. A numerical approach to a 2D porous-medium mathematical model: Application to an atherosclerosis problem. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007433.

Haemers:2024:SGT

[HT24] Willem H. Haemers and Hatice Topcu. The signed graphs with two eigenvalues unequal to ±1. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005179.

Haghighi:2020:PAN

[HTF20] Payam Haghighi, Babak Tavassoli, and Alireza Farhadi. A practical approach to networked control design for robust H_{∞} performance in the presence of uncertainties in both communication and system. Applied Mathematics and Computation, 381(??): Article 125308, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302745.

He:2023:MOC

[HTZ^{+}23] Sha He, Sanyi Tang, Qimin Zhang, Libin Rong, and Robert A. Cheke. Modelling optimal control of air pollution to reduce respiratory diseases. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003922.

Hu:2020:BSO

[Hu20]  Shenglong Hu.  Bounds

on strongly orthogonal ranks of tensors. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307155.

Hu:2021:NWZ

[Hu21] Fuxing Hu. A note on WENO–Z scheme. Applied Mathematics and Computation, 396(??):Article 125886, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308390.

Huang:2024:CGT

[Hua24] Baohua Huang. Conjugate gradient-type method for the tensor linear system via the t-product and its application in the calculation of Moore–Penrose inverse. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000997.

Hariprasad:2024:CDM

[HV24] M. Hariprasad and Murugesan Venkatapathi. Circulant decomposition of a matrix and the eigenvalues of Toeplitz type matrices. Applied Mathematics and Computation, 468(??): ??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006422.

Hernandez-Veron:2021:SNI

[HVYMS21] M. A. Hernández-Verón, Sonia Yadav, Eulalia Martínez, and Sukhjit Singh. Solving nonlinear integral equations with non-separable kernel via a high-order iterative process. Applied Mathematics and Computation, 409 (??):Article 126385, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004744.

Huzak:2023:FAD

[HVZZ23] Renato Huzak, Domagoj Vlah, Darko Zubrinić, and Vesna Zupanović. Fractal analysis of degenerate spiral trajectories of a class of ordinary differential equations. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006439.

Huang:2020:CN

[HW20] Yuanqiu Huang and Yuxi Wang. The crossing number of K_{5,n+1}\e. Applied Mathematics and Computation, 376(??):Article 125075, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300448.

He:2024:APS

[HW24] Hua He and Wendi Wang.

Asymptotically periodic solutions of fractional order systems with applications to population models. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002297.

He:2020:WPA

[HWC20] Bing He, Qi-Ru Wang, and Jun-Fei Cao. Weighted S^{p}- pseudo S-asymptotic periodicity and applications to Volterra integral equations. Applied Mathematics and Computation, 380(??): Article 125275, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302447.

Han:2023:ICH

[HWC23] Zhimin Han, Yi Wang, and Jinde Cao. Impact of contact heterogeneity on initial growth behavior of an epidemic: Complex networkbased approach. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300190X.

He:2022:NAH

[HWCD22] Yuyu He, Xiaofeng Wang, Hong Cheng, and Yaqing Deng. Numerical analysis of a high-order accurate compact finite difference scheme for the SRLW equation. Applied Mathematics and Computation, 418 (??):Article 126837, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321009206.

Han:2020:FTS

[HWD20] Xin-Xin Han, Kai-Ning Wu, and Xiaohua Ding. Finite-time stabilization for stochastic reaction-diffusion systems with Markovian switching via boundary control. Applied Mathematics and Computation, 385 (??):Article 125422, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303830.

Hua:2024:NMG

[HWL24] Wentao Hua, Yantao Wang, and Chunyan Liu. New method for global exponential synchronization of multi-link memristive neural networks with three kinds of time-varying delays. Applied Mathematics and Computation, 471(??): ??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000651.

He:2021:SNG

[HWLL21] Changxiang He, Wenyan Wang, Yuying Li, and Lele Liu. Some Nordhaus– Gaddum type results of A_{α}-eigenvalues of weighted graphs. Applied Mathematics and Computation, 393 (??):Article 125761, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307141.

Huang:2020:KOI

[HWW20] Zongyuan Huang, Haiyang Wang, and Zhen Wu. A kind of optimal investment problem under inflation and uncertain time horizon. Applied Mathematics and Computation, 375(??): Article 125084, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300539.

Huang:2023:LQM

[HWWW23] Pengyan Huang, Guangchen Wang, Wencan Wang, and Yu Wang. A linearquadratic mean-field game of backward stochastic differential equation with partial information and common noise. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000681.

Huang:2020:EPV

[HWXC20] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. An efficient preconditioned variant of the PSS preconditioner for generalized saddle point problems. Applied Mathematics and Computation, 376(??): Article 125110, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300795.

He:2022:PTM

[HWY^{+}22] Jialu He, Jianwei Wang, Fengyuan Yu, Wei Chen, and Wenshu Xu. The persistence and transition of multiple public goods games resolves the social dilemma. Applied Mathematics and Computation, 418(??):Article 126858, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009413.

Hou:2023:IRA

[HWZZ23] Pengwen Hou, Jun Wang, Qian Zhang, and Shuhua Zhang. Implications of risk aversion behavior on the green product promotion strategy under manufacturer encroachment.

Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000802.

He:2022:BAS

[HX22] Danhua He and Liguang Xu. Boundedness analysis of stochastic delay differential equations with Lévy noise. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321009851.

Huang:2023:ATS

[HXF23] Shijie Huang, Xufeng Xiao, and Xinlong Feng. An adaptive time-stepping method for the phase-field molecular beam epitaxial growth model on evolving surfaces. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006956.

Hu:2020:PSA

[HXL^{+}20] Jun Hu, Chengyi Xia, Huijia Li, Peican Zhu, and Wenjun Xiong. Properties and structural analyses of USA’s regional electricity market: a visibility graph network approach. Applied Mathematics and Computation, 385 (??):Article 125434, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303957.

Huang:2024:RLB

[HXWC24] He Huang, Jiawei Xu, Jing Wang, and Xiangyong Chen. Reinforcement learning-based secure synchronization for twotime-scale complex dynamical networks with malicious attacks. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003011.

Hou:2021:BCF

[HXZ21] Mimi Hou, Xuan-Xuan Xi, and Xian-Feng Zhou. Boundary control of a fractional reaction-diffusion equation coupled with fractional ordinary differential equations with delay. Applied Mathematics and Computation, 406 (??):Article 126260, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003490.

Han:2024:MBA

[HXZ24] Xu Han, Haoxiang Xia, and Xiaowei Zhao. Memorybased adaptive interaction willingness enhances cooperation in spatial prisoner’s dilemma. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002558.

Hu:2024:EGS

[HY24] Zuopeng Hu and Yanlong Yang. Existence and generic stability of open-loop Nash equilibria for noncooperative fuzzy differential games. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005404.

Hacat:2022:EEO

[HYÇK22] Gülnur Haçat, Fikriye Yilmaz, Aytekin Çibik, and Songül Kaya. Error estimates for the optimal control of Navier–Stokes equations using curvature based stabilization. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003149.

He:2024:IMP

[HYS24] Fuli He, Ali Yarahmadi, and Fazlollah Soleymani. Investigation of multivariate pairs trading under copula approach with mixture distribution. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001073.

Huang:2021:RCI

[HYT21] Jun Huang, Lin Yang, and Hieu Trinh. Robust control for incremental quadratic constrained nonlinear time-delay systems subject to actuator saturation. Applied Mathematics and Computation, 405(??): Article 126271, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100360X.

Hou:2020:CUS

[HZ20a] Jianfeng Hou and Hongguo Zhu. Choosability with union separation of planar graphs without cycles of length 4. Applied Mathematics and Computation, 386(??):Article 125477, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304367.

Hu:2020:BBB

[HZ20b] Zhiquan Hu and Yan Zou. Bipancyclicism and bipanconnectivity in bipartite graphs. Applied Mathematics and Computation, 377 (??):Article 125149, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301181.

Hou:2023:SHB

[HZ23a] Hu-Shuang Hou and Hua Zhang. Stability and Hopf bifurcation of fractional complex-valued BAM neural networks with multiple time delays. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001558.

Hou:2023:CPG

[HZ23b] Jianfeng Hou and Hongguo Zhu. On (2, r) -choosability of planar graphs without short cycles. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008980.

He:2023:DPC

[HZCY23] Yujia He, Shurong Zhang, Lin Chen, and Weihua Yang. The disjoint path cover in the data center network HSDC with prescribed vertices in each path. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004319.

He:2021:DSR

[HZF21] Yiqin He, Bicheng Zhang, and Rongquan Feng. Directed strongly regular Cayley graphs on dihedral groups. Applied Mathematics and Computation, 391 (??):Article 125651, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306056.

Han:2021:SSE

[HZJ^{+}21] Bingtao Han, Baoquan Zhou, Daqing Jiang, Tasawar Hayat, and Ahmed Alsaedi. Stationary solution, extinction and density function for a highdimensional stochastic SEI epidemic model with general distributed delay. Applied Mathematics and Computation, 405(??):Article 126236, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100326X.

Hu:2024:ACS

[HZJL24] Qi Hu, Mengyu Zhou, Yulian Jiang, and Xingwen Liu. Analysis of cooperative stability for reputation evaluation rules in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004387.

Hao:2021:FTC

[HZL21] Li-Ying Hao, Yu-Qing Zhang, and Hui Li. Faulttolerant control via integral sliding mode output feedback for unmanned marine vehicles. Applied Mathematics and Computation, 401 (??):Article 126078, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001260.

Hu:2021:ECN

[HZT^{+}21] Wenyu Hu, Zichang Zhang, Zhihui Tu, Jinhong Huang, and Duanmei Zhou. Efficient computation of the nearest polynomial by linearized alternating direction method. Applied Mathematics and Computation, 395(??):Article 125860, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308134.

Han:2021:FTR

[HZW21] Yunrui Han, Ying Zhao, and Peng Wang. Finitetime rate anti-bump switching control for switched systems. Applied Mathematics and Computation, 401 (??):Article 126086, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S009630032100134X.

Han:2021:ECT

[HZX21] Xu Han, Xiaowei Zhao, and Haoxiang Xia. Evolution of cooperation through aspiration-based adjustment of interaction range in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 393(??): Article 125746, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306998.

Hua:2022:ETA

[HZX22] Yu Hua, Tianping Zhang, and Xiaonan Xia. Eventtriggered adaptive neural command-filter-based dynamic surface control for state constrained nonlinear systems. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005148.

Huang:2024:ENE

[HZXT24] Jiaquan Huang, Yuying Zhu, Chengyi Xia, and Jun Tanimoto. The effect of nonlinear environmental feedback on the outcomes of evolutionary dynamics. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400451X.

Hu:2020:RHP

[HZXZ20] Beibei Hu, Ling Zhang, Tiecheng Xia, and Ning Zhang. On the Riemann– Hilbert problem of the Kundu equation. Applied Mathematics and Computation, 381(??):Article 125262, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302319.

He:2021:WIB

[HZZ^{+}21] Yanfei He, Xingwu Zhang, Tao Zhang, Chenxi Wang, Jia Geng, and Xuefeng Chen. A wavelet immersed boundary method for two-variable coupled fluid-structure interactions. Applied Mathematics and Computation, 405(??):Article 126243, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003337.

Ibrahim:2021:TSR

[ID21] Mahmoud A. Ibrahim and Attila Dénes. Threshold and stability results in a periodic model for malaria transmission with partial immunity in humans. Applied Mathematics and Computation, 392(??): Article 125711, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306640.

Ijaz:2021:AIS

[IFHA21] Salman Ijaz, Chen Fuyang, Mirza Tariq Hamayun, and Haris Anwaar. Adaptive integral-sliding-mode control strategy for maneuvering control of F16 aircraft subject to aerodynamic uncertainty. Applied Mathematics and Computation, 402(??):Article 126053, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001016.

Irena:2021:MDA

[IG21] Tsegaye Kebede Irena and Sunita Gakkhar. Modelling the dynamics of antimicrobial-resistant typhoid infection with environmental transmission. Applied Mathematics and Computation, 401(??):Article 126081, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001296.

Ikhdair:2013:RSY

[IHF13] Sameer M. Ikhdair, Majid Hamzavi, and Babatunde J. Falaye. Relativistic symmetries in Yukawa-type interactions with Coulomblike tensor. Applied Mathematics and Computation, 225(??):775–786, December 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300313010849.

See comments [Woo20].

Izgi:2024:MNM

[IKÖ24] Burhaneddin Izgi, Hale Gonce Kocken, and Murat Özkaya. Matrix norm methods for zero-sum fuzzy matrix games with payoffs of triangular fuzzy numbers. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003357.

Ilic:2022:NAW

[Ili22] Dragana S. Cvetković Ilić. Note on the assumptions in working with generalized inverses. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004337.

Ilyin:2023:LBM

[Ily23] Oleg Ilyin. Lattice Boltzmann model for diffusion equation with reduced truncation errors: Applications to Gaussian filtering and image processing. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002928.

Ismaeel:2022:NSN

[IMIH22] A. M. Ismaeel, M. A. Mansour, F. S. Ibrahim, and F. M. Hady. Numerical simulation for nanofluid extravasation from a vertical segment of a cylindrical vessel into the surrounding tissue at the microscale. Applied Mathematics and Computation, 417 (??):Article 126758, March 15, 2022. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008407.

Inferrera:2024:RDM

[IMOR24] G. Inferrera, C. F. Munafò, F. Oliveri, and P. Rogolino. Reaction-diffusion models of crimo-taxis in a street. Applied Mathematics and Computation, 467(??): ??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006732.

Ilyas:2023:RST

[IMS23] Asim Ilyas, Salman A. Malik, and Summaya Saif. Recovering source term and temperature distribution for nonlocal heat equation. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200683X.

Ilea:2020:SPI

[INOP20] Veronica Ilea, Adela Novac, Diana Otrocol, and Radu Precup. Solutions with a prescribed interval of positivity for differential systems with nonlocal conditions. Applied Mathematics and Computation, 375(??): Article 125092, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300618.

Iwamura:2020:EDS

[INT20] Yoshiro Iwamura, Keisuke Nagashima, and Jun Tanimoto. Evolutionary dynamics of a 3-strategy game: Cooperator, defector and costly cooperative loner strategic types. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308811.

Izgi:2023:EMN

[IÖÜP23] Burhaneddin Izgi, Murat Özkaya, Nazim Kemal Üre, and Matjaz Perc. Extended matrix norm method: Applications to bimatrix games and convergence results. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006270.

Izgi:2024:MNB

[IÖÜP24] Burhaneddin Izgi, Murat Özkaya, Nazim Kemal Üre, and Matjaz Perc. Matrix norm based hybrid Shapley and iterative methods for the solution of stochastic matrix games. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001103.

Innerberger:2023:MOO

[IP23] Michael Innerberger and Dirk Praetorius. MooAFEM: an object oriented Matlab code for higher-order adaptive FEM for (nonlinear) elliptic PDEs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007998.

Izadi:2022:SSD

[IR22] Mohammad Izadi and Pradip Roul. Spectral semidiscretization algorithm for a class of nonlinear parabolic PDEs with applications. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003009.

Izadi:2021:EAT

[IS21] Mohammad Izadi and H. M. Srivastava. An efficient approximation technique applied to a non-linear Lane– Emden pantograph delay differential model. Applied Mathematics and Computation, 401(??):Article 126123, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001715.

Ivagnes:2023:HDD

[ISM^{+}23] Anna Ivagnes, Giovanni Stabile, Andrea Mola, Traian Iliescu, and Gianluigi Rozza. Hybrid datadriven closure strategies for reduced order modeling. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000899.

Iqbal:2020:ENM

[IZ20] Sehar Iqbal and Paul Andries Zegeling. An efficient nonlinear multigrid scheme for 2D boundary value problems. Applied Mathematics and Computation, 372(??): Article 124898, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319308902.

Izadi:2021:DFE

[Iza21] Mohammad Izadi. A discontinuous finite element approximation to singular Lane–Emden type equations. Applied Mathematics and Computation, 401 (??):Article 126115, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001636.

Iurlano:2024:SDR

[IZDR24] Enrico Iurlano, Tatjana Zec, Marko Djukanovic, and Günther R. Raidl. Signed double Roman domination on cubic graphs. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000845.

Jahangiri:2023:SGB

[JA23] M. Jahangiri and M. Asghari. The strain gradientbased torsional vibration analysis of micro-rotors with nonlinear flexuraltorsional coupling. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006154.

Janani:2022:TAD

[Jan22] B. Janani. Transient analysis of differentiated breakdown model. Applied Mathematics and Computation, 417(??):Article 126779, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008614.

Jacobs:2022:UDI

[JC22a] B. A. Jacobs and T. Celik. Unsupervised document image binarization using a system of nonlinear partial differential equations. Applied Mathematics and Computation, 418 (??):Article 126806, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008882.

Jang:2022:OHS

[JC22b] Geunsoo Jang and Giphil Cho. Optimal harvest strategy based on a discrete age-structured model with monthly fishing effort for chub mackerel, Scomber japonicus, in South Korea. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200145X.

Jiang:2020:HBC

[JCCC20] Xiaowei Jiang, Xiangyong Chen, Ming Chi, and Jie Chen. On Hopf bifurcation and control for a delay systems. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308987.

Jiang:2020:MSO

[JCH^{+}20] Wei Jiang, Zhong Chen, Ning Hu, Haiyang Song, and Zhaohong Yang. Multi-scale orthogonal basis method for nonlinear fractional equations with fractional integral boundary value conditions. Applied Mathematics and Computation, 378(??):Article 125151, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030120X.

Jiang:2024:ELA

[JCSH24] Yubing Jiang, Hu Chen, Tao Sun, and Chaobao Huang. Efficient L1-ADI finite difference method for the twodimensional nonlinear timefractional diffusion equation. Applied Mathematics and Computation, 471(??): ??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400081X.

Jadlovska:2020:KTO

[JD20] Irena Jadlovská and Jozef Dzurina. Kneser-type oscillation criteria for secondorder half-linear delay differential equations. Applied Mathematics and Computation, 380(??):Article 125289, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302587.

Jia:2021:GSD

[JDLZ21] Jinping Jia, Hao Dai, Li Li, and Fandi Zhang. Global sampled-data stabilization for a class of nonlinear systems with arbitrarily long input delays via a multi-rate control algorithm. Applied Mathematics and Computation, 392(??): Article 125696, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306494.

Jia:2022:GSL

[JDZH22] Jinping Jia, Hao Dai, Fandi Zhang, and Jianwen Huang. Global stabilization of low-order stochastic nonlinear systems with multiple time-varying delays by a continuous feedback control. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003083.

Joe:2021:BNM

[JG21] Kirbi Joe and Maryam Gooyabadi. A Bayesian nonparametric mixture model for studying universal patterns in color naming. Applied Mathematics and Computation, 395(??):Article 125868, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320308213.

Jiang:2021:RBE

[JGC^{+}21] Luo-Luo Jiang, Jian Gao, Zhi Chen, Wen-Jing Li, and Jürgen Kurths. Reducing the bystander effect via decreasing group size to solve the collective-risk social dilemma. Applied Mathematics and Computation, 410(??):Article 126445, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005348.

Jin:2021:MCT

[JGW21] Zemin Jin, Xueyao Gui, and Kaijun Wang. Multicolorful connectivity of trees. Applied Mathematics and Computation, 402 (??):Article 126147, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001958.

Jiang:2023:SFV

[JGXL23] Wei-Wu Jiang, Xiao-Wei Gao, Bing-Bing Xu, and Jun Lv. Static and forced vibration analysis of layered piezoelectric functionally graded structures based on element differential method. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006221.

Ji:2021:SES

[JH21] Weiming Ji and Guixin Hu. Stability and explicit stationary density of a stochastic single-species model. Applied Mathematics and Computation, 390 (??):Article 125593, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305488.

Jian:2021:FNS

[JHO^{+}21] Huan-Yan Jian, Ting-Zhu Huang, Alexander Ostermann, Xian-Ming Gu, and Yong-Liang Zhao. Fast numerical schemes for nonlinear space-fractional multidelay reaction-diffusion equations by implicit integration factor methods. Applied Mathematics and Computation, 408(??):Article 126360, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004495.

Jung:2021:PPD

[JHSL21] H. I. Jung, S. Y. Han, Satnesh Singh, and S. M. Lee. Polynomially parameter dependent exponential stabilization of sampled-data LPV systems. Applied Mathematics and Computation, 411 (??):Article 126473, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005622.

Jin:2024:RDU

[JJC24] Zemin Jin, Qing Jie, and Zhenxin Cao. Rainbow disjoint union of P_{4} and a matching in complete graphs. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001516.

Jin:2023:EGA

[JJL23] Tao Jin, Yulian Jiang, and Xingwen Liu. Evolutionary game analysis of the impact of dynamic dual credit policy on new energy vehicles after subsidy cancellation. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007457.

Ju:2022:AOF

[JJSY22] Xinxu Ju, Xianglei Jia, Xiaocheng Shi, and Shan’en Yu. Adaptive output feedback event-triggered tracking control for nonlinear systems with unknown control coefficient. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200443X.

Jeong:2021:OHS

[JKJC21] Yong Dam Jeong, Sangil Kim, Il Hyo Jung, and Giphil Cho. Optimal harvesting strategy for hairtail, Trichiurus Lepturus, in Korea Sea using discrete-time age-structured model. Applied Mathematics and Computation, 392 (??):Article 125743, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306962.

Ji:2020:SCF

[JL20] Meng Ji and Xueliang Li.

Strong conflict-free connection of graphs. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306319.

Jiang:2021:MAZ

[JL21] Yisheng Jiang and Mei Lu. Maximal augmented Zagreb index of trees with given diameter. Applied Mathematics and Computation, 395(??):Article 125855, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308080.

Jiang:2022:AHO

[JL22] Senbao Jiang and Xiaofan Li. Arbitrarily highorder trapezoidal rules for functions with fractional singularities in two dimensions. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003101.

Jiang:2024:STS

[JL24a] Zhuling Jiang and Jicheng Li. Slice tensor splitting method for solving tensor equation. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005362.

Juang:2024:EMW

[JL24b] Jonq Juang and Yu-Hao Liang. Epidemic models in well-mixed multiplex networks with distributed time delay. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001541.

Ji:2024:SAB

[JLD^{+}24] Hankang Ji, Yuanyuan Li, Xueying Ding, Sultan M. Alghamdi, and Jianquan Lu. Stability analysis of Boolean networks: an eigenvalue approach. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005301.

Jia:2020:ETG

[JLFL20] Hongen Jia, Yang Li, Guorui Feng, and Kaitai Li. An efficient twogrid method for the Cahn– Hilliard equation with the concentration-dependent mobility and the logarithmic Flory–Huggins bulk potential. Applied Mathematics and Computation, 387(??): Article 124548, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319305314.

Ji:2023:SIN

[JLJ23] Ruihong Ji, Wen Luo, and Liya Jiang. Stability of the 3D incompressible Navier– Stokes equations with fractional horizontal dissipation. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001030.

Jie:2020:GTR

[JLL^{+}20] Yingmo Jie, Charles Zhechao Liu, Mingchu Li, Kim-Kwang Raymond Choo, Ling Chen, and Cheng Guo. Game theoretic resource allocation model for designing effective traffic safety solution against drunk driving. Applied Mathematics and Computation, 376(??): Article 125142, July 1, 2020.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301119.

Jo:2022:IRS

[JLOR22] Gwanghyun Jo, Young Ju Lee, and Iván Ojeda-Ruiz. 2D and 3D image reconstruction from slice data based on a constrained bilateral smoothing and dynamic mode decomposition. Applied Mathematics and Computation, 420(??): Article 126877, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009607.

Jeong:2023:ALK

[JLP23] Juyoung Jeong, Yongdo Lim, and Arumugam Parivallal. An asymmetric Lyapunov-Krasovskii functional approach for event-triggered consensus of multi-agent systems with deception attacks. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006579.

Jian:2021:IRA

[JLWX21] Qing Jian, Xiaopeng Li, Juan Wang, and Chengyi Xia. Impact of reputation assortment on tagmediated altruistic behaviors in the spatial lattice. Applied Mathematics and Computation, 396(??): Article 125928, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030881X.

Jia:2022:ENA

[JLZ^{+}22] Danyang Jia, Tong Li, Yang Zhao, Xiaoqin Zhang, and Zhen Wang. Empty nodes affect conditional cooperation under reinforcement learning. Applied Mathematics and Computation, 413(??):Article 126658, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007426.

Jiang:2023:DAP

[JM23a] Jiehui Jiang and Jie Ma. Dynamic analysis of pandemic cross-regional transmission considering quarantine strategies in the context of limited medical resources. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001273.

Jons:2023:RSP

[JM23b] Steven Jöns and Claus-Dieter Munz. Riemann solvers for phase transition in a compressible sharpinterface method. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200697X.

Jin:2023:IMF

[JMC23] Xiao-Zheng Jin, Yong-Sheng Ma, and Wei-Wei Che. An improved model-free adaptive control for nonlinear systems: an LMI approach. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000796.

Jamilla:2020:SND

[JMM20] Cristeta Jamilla, Renier Mendoza, and István Mező. Solutions of neutral delay differential equations using a generalized Lambert W function. Applied Mathematics and Computation, 382(??):Article 125334, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303003.

Jin:2022:RME

[JMY22] Zemin Jin, Huawei Ma, and Rui Yu. Rainbow matchings in an edge-colored planar bipartite graph. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004301.

Jalaali:2021:IEV

[JNY^{+}21] Bahrul Jalaali, Muhammad Ridlo Erdata Nasution, Kumara Ari Yuana, Deendarlianto, and Okto Dinaryanto. Investigating the effects of viscosity and density ratio on the numerical analysis of Rayleigh– Taylor instability in twophase flow using lattice Boltzmann method: From early stage to equilibrium state. Applied Mathematics and Computation, 411 (??):Article 126490, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005798.

Jimenez:2021:ISB

[JOR^{+}21a] I. Jiménez, P. Ortiz, J. Ruiz, J. C. Trillo, and D. F. Yañez. Improving the stability bound for the PPH nonlinear subdivision scheme for data coming from strictly convex functions. Applied Mathematics and Computation, 399(??): Article 126042, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000904.

Jornet:2021:BHB

[Jor21b] Marc Jornet. Beyond the hypothesis of boundedness for the random coefficient of the Legendre differential equation with uncertainties. Applied Mathematics and Computation, 391 (??):Article 125638, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305920.

Jornet:2021:ESM

[Jor21c] Marc Jornet. Exact solution to a multidimensional wave equation with delay. Applied Mathematics and Computation, 409 (??):Article 126421, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005105.

Jornet:2023:RFB

[Jor23]  Marc Jornet.  On the

random fractional Bateman equations. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003661.

Jornet:2024:GPC

[Jor24] Marc Jornet. Generalized polynomial chaos expansions for the random fractional Bateman equations. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003345.

Jeon:2020:DAA

[JP20] Junkee Jeon and Kyunghyun Park. Dynamic asset allocation with consumption ratcheting post retirement. Applied Mathematics and Computation, 385 (??):Article 125418, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303799.

Junike:2022:POP

[JP22] Gero Junike and Konstantin Pankrashkin. Precise option pricing by the COS method-how to choose the truncation range. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000212.

Jeon:2023:OJS

[JP23] Junkee Jeon and Kyunghyun Park. Optimal job switching and retirement decision. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008451.

Jain:2023:HML

[JPY23] Riya Jain, Amiya K. Pani, and Sangita Yadav. HDG method for linear parabolic integro-differential equations. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300156X.

Jin:2021:RFT

[JPZ^{+}21] Shuowei Jin, Yongheng Pang, Xiaoming Zhou, Aiyun Yan, Wei Wang, and Wenbo Hu. Robust finitetime control and reachable set estimation for uncertain switched neutral systems with time delays and input constraints. Applied Mathematics and Computation, 407(??):Article 126321, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004100.

Ji:2023:IOH

[JPZH23] Jiezhou Ji, Qiuhui Pan, Wenqiang Zhu, and Mingfeng He. The influence of own historical information and environmental historical information on the evolution of cooperation. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000711.

Jiao:2022:NIS

[JQJY22] Ticao Jiao, Xiaomei Qi, Jishun Jiang, and Mingzheng Yu. Noise-input-to-state stability analysis of switching stochastic nonlinear systems with mode-dependent multiple impulses. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004842.

Jafarimoghaddam:2022:TBD

[JRP22] Amin Jafarimoghaddam, Alin Rosca, and Ioan Pop. Theoretical breakthrough in the dynamics of a jet in a freestream flow around a corner. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001217.

Jeet:2020:ACN

[JS20] Kamal Jeet and N. Sukavanam. Approximate controllability of nonlocal and impulsive neutral integrodifferential equations using the resolvent operator theory and an approximating technique. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306824.

Jaune:2022:ARD

[JS22] Egle Jaune and Jonas Siaulys. Asymptotic risk decomposition for regularly varying distributions with tail dependence. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002247.

Joby:2021:FTC

[JSA21] Maya Joby, Srimanta Santra, and S. Marshal Anthoni. Finite-time contractive boundedness of extracorporeal blood circulation process. Applied Mathematics and Computation, 388(??):Article 125527, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320304707.

Ji:2020:PPC

[JSCB20] Yanyan Ji, Bilin Shao, Jinyong Chang, and Genqing Bian. Privacy-preserving certificateless provable data possession scheme for big data storage on cloud, revisited. Applied Mathematics and Computation, 386(??):Article 125478, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304379.

Ju:2020:SSP

[JSM20] Yanhao Ju, Yuangong Sun, and Fanwei Meng. Stabilization of switched positive system with impulse and marginally stable subsystems: a mode-dependent dwell time method. Applied Mathematics and Computation, 383(??):Article 125377, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303416.

Jin:2020:TIA

[JSZ20] Ting Jin, Yun Sun, and Yuanguo Zhu. Time integral about solution of an uncertain fractional order differential equation and application to zero-coupon bond model. Applied Mathematics and Computation, 372 (??):Article 124991, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630031930983X.

Sun:2020:NAC

[jSZyW^{+}20] Yu jie Sun, Hao Zhang, Xing yuan Wang, Xiao qing Wang, and Peng fei Yan. 2D non-adjacent coupled map lattice with q and its applications in image encryption. Applied Mathematics and Computation, 373 (??):Article 125039, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300084.

Jimenez:2020:SPP

[JT20] Nelson Gutiérrez Jiménez and Sergii M. Torba. Spectral parameter power series representation for solutions of linear system of two first order differential equations. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309038.

Jin:2022:DTC

[JT22] Wei Jin and Li Tan. On diameter two Cayley graphs. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005112.

Jones:2024:DEB

[JT24] Giselle Sosa Jones and Catalin Trenchea. Discrete energy balance equation via a symplectic second-order method for two-phase flow in porous media. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003709.

Jin:2024:NEB

[JTC^{+}24] Xiaozheng Jin, Xingcheng Tong, Jing Chi, Xiaoming Wu, and Hai Wang. Nonlinear estimator-based funnel tracking control for a class of perturbed Euler–Lagrange systems. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000663.

Jin:2023:SUP

[JTW^{+}23] Xing Jin, Yuchen Tao, Jingrui Wang, Chao Wang, Yongheng Wang, Zhouyang Zhang, and Zhen Wang.

Strategic use of payoff information in k-hop evolutionary best-shot networked public goods game. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300440X.

Jin:2021:ISS

[JW21] Zhenghong Jin and Zhanxiu Wang. Input-to-state stability of the nonlinear singular systems via small-gain theorem. Applied Mathematics and Computation, 402(??):Article 126171, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002617.

Jentzen:2023:OEA

[JW23] Arnulf Jentzen and Timo Welti. Overall error analysis for the training of deep neural networks via stochastic gradient descent with random initialisation. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000760.

Jia:2020:NFR

[JWC20] You Jia, Huaiqin Wu, and Jinde Cao. Non-fragile robust finite-time synchronization for fractional-order discontinuous complex networks with multi-weights and uncertain couplings under asynchronous switching. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930921X.

Jiang:2022:TAA

[JWK22] Baoping Jiang, Zhengtian Wu, and Hamid Reza Karimi. A traverse algorithm approach to stochastic stability analysis of Markovian jump systems with unknown and uncertain transition rates. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000546.

Jin:2024:FTC

[JWL^{+}24] Zengke Jin, Chaoli Wang, Dong Liang, Zongyu Zuo, and Zhenying Liang. Fixedtime cooperative output regulation for second-order nonlinear multiagent systems with an unknown exosystem. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002303.

Jin:2023:CAR

[JWMW23] Zemin Jin, Yite Wang, Huawei Ma, and Huaping Wang. Computing the anti-Ramsey number for trees in complete tripartite graph. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300320X.

Jin:2020:RTE

[JWWL20] Zemin Jin, Fang Wang, Huaping Wang, and Bihong Lv. Rainbow triangles in edge-colored Kneser graphs. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307167.

Jung:2020:SPM

[JWY20] Yoon Mo Jung, Joyce Jiyoung Whang, and Sangwoon Yun. Sparse probabilistic K-means. Applied Mathematics and Computation, 382(??):Article 125328, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302940.

Ji:2024:DEH

[JX24] Antong Ji and Meizhen Xu. Dependence of eigenvalues for higher odd-order boundary value problems. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006562.

Jeong:2020:NFN

[JY20] Byeongseon Jeong and Jungho Yoon. A new family of non-stationary Hermite subdivision schemes reproducing exponential polynomials. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307556.

Jeong:2021:AET

[JY21] Dong Min Jeong and Sung Jin Yoo. Adaptive event-triggered tracking using nonlinear disturbance observer of arbitrarily switched uncertain nonlinear systems in purefeedback form. Applied Mathematics and Computation, 407(??):Article 126335, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321004240.

Jiang:2023:DCC

[JY23] Suyun Jiang and Jin Yan. Disjoint cycles covering specified vertices in bipartite graphs with partial degrees. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006993.

Jahanshahi:2020:NFO

[JYMP^{+}20] Hadi Jahanshahi, Amin Yousefpour, Jesus M. Munoz-Pacheco, Sezgin Kacar, Viet-Thanh Pham, and Fawaz E. Alsaadi. A new fractional-order hyperchaotic memristor oscillator: Dynamic analysis, robust adaptive synchronization, and its application to voice encryption. Applied Mathematics and Computation, 383 (??):Article 125310, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302769.

Ji:2020:DFD

[JZ20] Bingquan Ji and Luming Zhang. A dissipative finite difference Fourier pseudospectral method for the Klein–Gordon–Schrödinger equations with damping mechanism. Applied Mathematics and Computation, 376(??):Article 125148, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030117X.

Ju:2022:FDL

[JZ22] Chunming Ju and Binlin Zhang. On fractional discrete p-Laplacian equations via Clark’s theorem. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005173.

Jiao:2020:AAS

[JZP^{+}20] Ticao Jiao, Guangdeng Zong, Guochen Pang, Housheng Zhang, and Jishun Jiang. Admissibility analysis of stochastic singular systems with Poisson switching. Applied Mathematics and Computation, 386 (??):Article 125508, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304665.

Javaid:2024:OBA

[JZS^{+}24] Umair Javaid, Ziyang Zhen, Sami Shahid, Dauda Sh Ibrahim, and Salman Ijaz. Observer-based attitude control of spacecraft under actuator dead zone and misalignment faults. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005751.

Ji:2022:UIR

[JZZ^{+}22] Kaipeng Ji, Peng Zhao, Xiaowei Zhou, Yuhong Chen, Zhengyang Dong, Jianguo Zheng, Jianzhong Fu, and Huamin Zhou. Uniform initialization in response space for PSO and its applications. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004258.

Khan:2022:OFS

[KA22] M. Ijaz Khan and Faris Alzahrani. Optimized framework for slip flow of viscous fluid towards a curved surface with viscous dissipation and Joule heating features. Applied Mathematics and Computation, 417(??):Article 126777, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008596. See

corrigendum [KA23].

Khan:2023:COF

[KA23] M. Ijaz Khan and Faris Alzahrani. Corrigendum to Optimized framework for slip flow of viscous fluid towards a curved surface with viscous dissipation and Joule heating features [Applied Mathematics and Computation, 417 (2022) 126777]. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000590.

See [KA22].

Kabir:2024:NVG

[Kab24] K. M. Ariful Kabir. Newborn vaccination game model with potentially periodic behavior. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002273.

Kang:2020:EBS

[Kan20]  Jum-Ran Kang.  Exis-

tence and blow-up of solutions for von karman equations with time delay and variable exponents. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309099.

Karageorghis:2022:TEV

[Kar22] Andreas Karageorghis. A time-efficient variable shape parameter Kansa-radial basis function method for the solution of nonlinear boundary value problems. Applied Mathematics and Computation, 413 (??):Article 126613, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006974.

Khaled-Abad:2021:AWG

[KAS21] Leila Jafarian Khaled-Abad and Rezvan Salehi. Application of weak Galerkin finite element method for nonlinear chemotaxis and haptotaxis models. Applied Mathematics and Computation, 409(??):Article 126436, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005257.

Khan:2022:ITB

[KAT22] Md. Mamun-Ur-Rashid Khan, Md. Rajib Arefin, and Jun Tanimoto. Investigating the trade-off between selfquarantine and forced quarantine provisions to control an epidemic: an evolutionary approach. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004398.

Kaushik:2021:PPS

[KB21] Rajat Kaushik and Sandip Banerjee. Predator-prey system: Prey’s counterattack on juvenile predators shows opposite side of the same ecological coin. Applied Mathematics and Computation, 388 (??):Article 125530, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304860.

Khebbache:2024:CAE

[KBL^{+}24] Hicham Khebbache, Abderrahim Benmicia, Salim Labiod, Naamane Bounar, and Abdesselem Boulkroune. Composite adaptive exponential tracking control for large-scale nonlinear systems with sensor faults. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002157.

Kang:2022:SPS

[KCJ22] Yingli Kang, Xiaoyue Chen, and Ligang Jin. A study on parity signed graphs: the rna number. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003964.

Kovtanyuk:2021:IPE

[KCT^{+}21] Andrey Kovtanyuk, Alexander Chebotarev, Varvara Turova, Irina Sidorenko, and Renée Lampe. An inverse problem for equations of cerebral oxygen transport. Applied Mathematics and Computation, 402(??):Article 126154, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002022.

Keanini:2023:SEG

[KDB^{+}23] Russell G. Keanini, Jerry Dahlberg, Philip Brown, Mehdi Morovati, Hamidreza Moradi, Donald Jacobs, and Peter T. Tkacik. Stochastic estimation of Green’s functions with application to diffusion and advection-diffusion-reaction problems. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003557.

Kiouach:2023:IES

[KEiS23] Driss Kiouach, Salim El Azami El-idrissi, and Yassine Sabbar. An improvement of the extinction sufficient conditions for a higherorder stochastically disturbed AIDS/HIV model. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000462.

Khabaz:2022:VDA

[KET22] Mohamad Khaje Khabaz, S. Ali Eftekhari, and Davood Toghraie. Vibration and dynamic analysis of a cantilever sandwich microbeam integrated with piezoelectric layers based on strain gradient theory and surface effects. Applied Mathematics and Computation, 419(??): Article 126867, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009504.

Kaemawichanurat:2023:PDI

[KF23] Pawaton Kaemawichanurat and Odile Favaron. Partial domination and irredundance numbers in graphs. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003223.

Kofman:2021:CSS

[KFM21a] Ernesto Kofman, Joaquín Fernández, and Denise Marzorati. Compact sparse symbolic Jacobian computation in large systems of ODEs. Applied Mathematics and Computation, 403(??):Article 126181, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100271X.

See corrigendum [KFM21b].

Kofman:2021:CCS

[KFM21b] Ernesto Kofman, Joaquín Fernández, and Denise Marzorati. Corrigendum to “Compact Sparse Symbolic Jacobian Computation in Large Systems of ODEs” [Applied Mathematics and Computation 403 (2021) 126181]. Applied Mathematics and Computation, 406(??):Article 126290, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/S0096300321003799.

See [KFM21a].

Kurt:2023:TUF

[KFÖ23] Cemaliye Kürt, Arran Fernandez, and Mehmet Ali Özarslan. Two unified families of bivariate Mittag-Leffler functions. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008530.

Klocker:2020:LBS

[KFR20] Benedikt Klocker, Herbert Fleischner, and Günther R. Raidl. A lower bound for the smallest uniquely Hamiltonian planar graph with minimum degree three. Applied Mathematics and Computation, 380(??):Article 125233, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302022.

Krishnan:2021:MFA

[KG21] K. V. Vishal Krishnan and Ranjan Ganguli. Multifidelity analysis and uncertainty quantification of beam vibration using cokriging interpolation method. Applied Mathematics and Computation, 398(??):Article 125987, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000357.

Kaur:2022:AWO

[KG22] Navjot Kaur and Kavita Goyal. An adaptive wavelet optimized finite difference B-spline polynomial chaos method for random partial differential equations. Applied Mathematics and Computation, 415 (??):Article 126738, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008201.

Kang:2020:FTC

[KGCC20] Wei Kang, Qingfei Gao, Menglong Cao, and Jun Cheng. Finite-time control for Markovian jump systems subject to randomly occurring quantization. Applied Mathematics and Computation, 385 (??):Article 125402, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303647.

Kerr:2022:NMB

[KGPS22] Gilbert Kerr, Gilberto González-Parra, and Michele Sherman. A new method based on the Laplace transform and Fourier series for solving linear neutral delay differential equations. Applied Mathematics and Computation, 420(??): Article 126914, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009978.

Koyano:2023:VFG

[KH23] Hitoshi Koyano and Morihiro Hayashida. Volume formula and growth rates of the balls of strings under the edit distances. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003715.

Khan:2022:IAN

[KHA22] Sohail A. Khan, T. Hayat, and A. Alsaedi. Irreversibility analysis for nanofluid (NiZnFe_{2}O_{4}–C_{8}H_{18} and MnZnFe_{2}O_{4}–C_{8}H_{18}) flow with radiation effect. Applied Mathematics and Computation, 419(??):Article 126879, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009620.

Kohatsu-Higa:2022:DEJ

[KHNT22] Arturo Kohatsu-Higa, Eulalia Nualart, and Ngoc Khue Tran. Density estimates for jump diffusion processes. Applied Mathematics and Computation, 420(??): Article 126814, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008973.

Korpinar:2020:TAS

[KIB20] Zeliha Korpinar, Mustafa Inc, and Mustafa Bayram. Theory and application for the system of fractional Burger equations with Mittag-Leffler kernel. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307738.

Kim:2020:MAE

[KIJK20] Kwang Su Kim, Malik Muhammad Ibrahim, Il Hyo Jung, and Sangil Kim. Mathematical analysis of the effectiveness of control strategies to prevent the autorun virus transmission propagation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309476.

Kwon:2020:PTE

[KJL20] W. Kwon, Yongsik Jin, and S. M. Lee. PI-type event-triggered H_{∞} filter for networked T –S fuzzy systems using affine matched membership function approach. Applied Mathematics and Computation, 385 (??):Article 125420, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303817.

Kim:2020:CSB

[KK20] Philsu Kim and Dojin Kim. Convergence and stability of a BSLM for advection-diffusion models with Dirichlet boundary conditions. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307362.

Kumar:2022:DBS

[KK22] Raj Kumar and Avneesh Kumar. Dynamical behavior of similarity solutions of CKOEs with conservation law. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000625.

Kaslik:2023:SBS

[KK23a] Eva Kaslik and Emanuel-Attila Kokovics. Stability and bifurcations in scalar differential equations with a general distributed delay. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002692.

Keng:2023:CSN

[KK23b] Ying Ying Keng and Kiam Heong Kwa. Contagion in social networks: On contagion thresholds. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002904.

Kim:2023:SLV

[KK23c] Hyun-Gyoon Kim and Jeong-Hoon Kim. A stochastic-local volatility model with Le′vy jumps for pricing derivatives. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002035.

Kumar:2024:WPE

[KK24a] Amit Kumar and Santosh Kapuria. Wave packet enriched finite element for accurate modelling of thermoelectroelastic wave propagation in functionally graded solids. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002571.

Kumar:2024:SSS

[KK24b] Rakesh Kumar and B. V. Rathish Kumar. Superconvergent scheme for a system of Green nonlinear Fredholm integral equations. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003412.

Keshavarz-Kohjerdi:2023:FHC

[KKB23] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. Finding Hamiltonian cycles of truncated rectangular grid graphs in linear time. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005872.

Kwak:2023:GFC

[KKH23] Dohyeok Kwak, Jung Hoon Kim, and Tomomichi Hagiwara. Generalized framework for computing the L_{∞}- induced norm of sampleddata systems. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005926.

Kansal:2022:EMI

[KKK22] Munish Kansal, Sanjeev Kumar, and Manpreet Kaur. An efficient matrix iteration family for finding the generalized outer inverse. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003666.

Kim:2020:NEA

[KKLJ20] Hyunju Kim, Keon Ho Kim, Seyeon Lee, and Bongsoo Jang. New explicit and accelerated techniques for solving fractional order differential equations. Applied Mathematics and Computation, 379(??):Article 125228, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301971.

Kumhar:2020:GFS

[KKPG20] Raju Kumhar, Santimoy Kundu, Deepak Kr. Pandit, and Shishir Gupta. Green’s function and surface waves in a viscoelastic orthotropic FGM enforced by an impulsive point source. Applied Mathematics and Computation, 382(??):Article 125325, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302915.

Kaviarasan:2021:SFE

[KKPS21] Boomipalagan Kaviarasan, Oh-Min Kwon, Myeong Jin Park, and Rathinasamy Sakthivel. Stochastic faulty estimator-based non-fragile tracking controller for multiagent systems with communication delay. Applied Mathematics and Computation, 392(??):Article 125704, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306573.

Kaviarasan:2023:ROF

[KKPS23] Boomipalagan Kaviarasan, Oh-Min Kwon, Myeong Jin Park, and Rathinasamy Sakthivel. Reduced-order filtering for semi-Markovian jump systems against randomly occurring false data injection attacks. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000012.

Keng:2021:CAD

[KKR21] Ying Ying Keng, Kiam Heong Kwa, and Kurunathan Ratnavelu. Centrality analysis in a drug network and its application to drug repositioning. Applied Mathematics and Computation, 395(??): Article 125870, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308237.

Kavikumar:2022:SMC

[KKS^{+}22a] R. Kavikumar, O. M. Kwon, R. Sakthivel, S. H. Lee, S. G. Choi, and S. Priyanka. Sliding mode control for IT2 fuzzy semi-Markov systems with faults and disturbances. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200114X.

Kokurin:2022:IRG

[KKS22b] M. M. Kokurin, M. Yu. Kokurin, and A. V. Semenova. Iteratively regularized Gauss–Newton type methods for approximating quasi-solutions of irregular nonlinear operator equations in Hilbert space with an application to COVID-19 epidemic dynamics. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003861.

Kolomoitsev:2021:ALC

[KL21] Yurii Kolomoitsev and Tetiana Lomako. Asymptotics of the Lebesgue constants for bivariate approximation processes. Applied Mathematics and Computation, 403(??):Article 126192, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002824.

Kang:2022:MSM

[KL22] Sangwoo Kang and Mikyoung Lim. Monostatic sampling methods in limitedaperture configuration. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002260.

Kudryashov:2024:TWS

[KL24] Nikolay A. Kudryashov and Sofia F. Lavrova. Traveling wave solutions of the derivative nonlinear Schrödinger hierarchy. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002637.

Kubyana:2024:ARP

[KLH24] Mmatlou S. Kubyana, Pietro Landi, and Cang Hui. Adaptive rock-paper-scissors game enhances eco-evolutionary performance at cost of dynamic stability. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000079.

Kumar:2023:SAS

[KLNN23] V. B. Kiran Kumar, A. Lexy, M. N. N. Namboodiri, and A. Noufal. Spectral analysis of singular matrices in SIMO channel. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008281.

Kordonis:2022:NSD

[KLP22] Ioannis Kordonis, Athanasios-Rafail Lagos, and George P. Papavassilopoulos. Nash social distancing games with equity constraints: How inequality aversion affects the spread of epidemics. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005276.

Kwon:2020:AZE

[KLPL20] O. M. Kwon, S. H. Lee, M. J. Park, and S. M. Lee. Augmented zero equality approach to stability for linear systems with timevarying delay. Applied Mathematics and Computation, 381(??):Article 125329, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302952.

Kim:2020:EFH

[KLS20] Jinkyu Kim, Hyeonseok Lee, and Jinwon Shin. Extended framework of Hamilton’s principle applied to Duffing oscillation. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307544.

Kaur:2020:HAC

[KM20] Deepti Kaur and R. K. Mohanty. Highly accurate compact difference scheme for fourth order parabolic equation with Dirichlet and Neumann boundary conditions: Application to good Boussinesq equation. Applied Mathematics and Computation, 378 (??):Article 125202, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301715.

Kang:2022:FDI

[KM22] Haobo Kang and Hongjun Ma. Fault detection and isolation of actuator failures in jet engines using adaptive dynamic programming. Applied Mathematics and Computation, 414 (??):Article 126664, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007487.

Kumar:2021:SCA

[KMD21] Vipin Kumar, Muslim Malik, and Amar Debbouche. Stability and controllability analysis of fractional damped differential system with non-instantaneous impulses. Applied Mathematics and Computation, 391 (??):Article 125633, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305877.

Kaneba:2022:ETC

[KMLW22] Christopher Munyiwa Kaneba, Xiaowu Mu, Xin Li, and Xihui Wu. Event triggered control for fault tolerant control system with actuator failure and randomly occurring parameter uncertainty. Applied Mathematics and Computation, 415 (??):Article 126714, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007980.

Kaluza:2024:DLB

[KMM^{+}24] Andrzej Kałuza, Paweł M.

Morkisz, Bartłomiej Mulewicz, Paweł Przybyłowicz, and Martyna Wiacek. Deep learning-based estimation of time-dependent parameters in Markov models with application to nonlinear regression and SDEs. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003679.

Karakazian:2022:FEM

[KMN22] Hagop Karakazian, Sophie Moufawad, and Nabil Nassif. A finite-element model for the Hasegawa–Mima wave equation. Applied Mathematics and Computation, 412(??):Article 126550, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321006342.

Katsikis:2020:TVM

[KMS^{+}20] Vasilios N. Katsikis, Spyridon D. Mourtas, Predrag S. Stanimirović, Shuai Li, and Xinwei Cao. Time-varying minimum-cost portfolio insurance under transaction costs problem via Beetle Antennae Search Algorithm (BAS). Applied Mathematics and Computation, 385 (??):Article 125453, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304148.

Katsikis:2023:TVM

[KMS^{+}23] Vasilios N. Katsikis, Spyridon D. Mourtas, Predrag S. Stanimirović, Shuai Li, and Xinwei Cao. Timevarying minimum-cost portfolio insurance problem via an adaptive fuzzypower LVI-PDNN. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007688.

Knor:2021:GEM

[KMT^{+}21] Martin Knor, Snjezana Majstorović, Aoden Teo Masa Toshi, Riste Skrekovski, and Ismael G. Yero. Graphs with the edge metric dimension smaller than the metric dimension. Applied Mathematics and Computation, 401(??):Article 126076, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001247.

Karasozen:2021:ROM

[KMUY21] Bülent Karasözen, Gülden Mülayim, Murat Uzunca, and Süleyman Yildiz. Reduced order modelling of nonlinear cross-diffusion systems. Applied Mathematics and Computation, 401 (??):Article 126058, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321001065.

Kawano:2022:DPI

[KMZ22] Alexandre Kawano, Antonino Morassi, and Ramón Zaera. Determination of the prey impact region in a spider orb-web from in-plane vibration. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000339.

Kudenatti:2023:OIH

[KNEM23] Ramesh B. Kudenatti and Noor-E-Misbah. The onset of instability in a hydromagnetic channel flow of Casson fluid: the accurate solutions. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005495.

Kirkby:2020:GCT

[KNN20] J. Lars Kirkby, Dang H. Nguyen, and Duy Nguyen. A general continuous time Markov chain approximation for multi-asset option pricing with systems of correlated diffusions. Applied Mathematics and Computation, 386 (??):Article 125472, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304318.

Kozera:2021:EPF

[KNW21] Ryszard Kozera, Lyle Noakes, and Magdalena Wilkołazka. Exponential parameterization to fit reduced data. Applied Mathematics and Computation, 391(??):Article 125645, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305993.

Karpuz:2021:IPT

[KÖ21] Basak Karpuz and Büsra Özsavas. An improved product type oscillation test for partial difference equations. Applied Mathematics and Computation, 391 (??):Article 125629, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030583X.

Kodnyanko:2021:IBP

[Kod21] Vladimir Kodnyanko. Improved bracketing parabolic method for numerical solution of nonlinear equations. Applied Mathematics and Computation, 400(??): Article 125995, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000436.

Kozitsky:2021:JCC

[KOP21] Yuri Kozitsky, Igor Omelyan, and Krzysztof Pilorz. Jumps and coalescence in the continuum: a numerical study. Applied Mathematics and Computation, 390 (??):Article 125610, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305658.

Koprowski:2022:NPN

[Kop22] Przemysław Koprowski. A note on the paper “A new approach for finding the determinant of matrices” by o. rezaifar and h. rezaee. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003198.

Kilic:2024:IOT

[KÖY24] Nurten Kiliç, Özkan Öcalan, and Mustafa Kemal Yildiz. An improved oscillation theorem for nonlinear delay differential equations. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003138.

Kalinowski:2022:PDA

[KP22] Rafał Kalinowski and Monika Pilśniak. Proper distinguishing arc-colourings of symmetric digraphs. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200025X.

Kelenc:2023:SMP

[KP23] Aleksander Kelenc and Iztok Peterin. On some metric properties of direct-co-direct product. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003211.

Kim:2024:PBS

[KP24] Minho Kim and Jörg Peters. A practical box spline compendium. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005453.

Koellermeier:2022:SSW

[KPG22] Julian Koellermeier and Ernesto Pimentel-García. Steady states and wellbalanced schemes for shallow water moment equations with topography. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002417.

Khebchareon:2022:GMF

[KPP22]  Morrakot  Khebchareon,

Ambit Kumar Pany, and Amiya K. Pani. An H^{1}- Galerkin mixed finite element method for identification of time dependent parameters in parabolic problems. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200131X.

Kincl:2023:UDF

[KPPK23] Ondrej Kincl, Ilya Peshkov, Michal Pavelka, and Václav Klika. Unified description of fluids and solids in Smoothed Particle Hydrodynamics. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006531.

Knez:2022:CPH

[KPS22] Marjeta Knez, Francesca Pelosi, and Maria Lucia Sampoli. Construction of G^{2} planar Hermite interpolants with prescribed arc lengths. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200176X.

Kumar:2020:SPM

[KR20] Abhishek Kumar and Rajeev. A Stefan problem with moving phase change material, variable thermal conductivity and periodic boundary condition. Applied Mathematics and Computation, 386 (??):Article 125490, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030446X.

Klavzar:2021:GPN

[KR21] Sandi Klavzar and Gregor Rus. The general position number of integer lattices. Applied Mathematics and Computation, 390 (??):Article 125664, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306172.

Kaslik:2022:SBF

[KR22] Eva Kaslik and Ileana Rodica Radulescu. Stability and bifurcations in fractionalorder gene regulatory networks. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000029.

Klavzar:2020:LBS

[KRE20] Sandi Klavzar, Indra Rajasingh, and D. Ahima Emilet. A lower bound and several exact results on the d-lucky number. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307520.

Kheybari:2020:ANP

[KRF20] Siamak Kheybari, Fariba Mahdi Rezaie, and Hadis Farazmand. Analytic network process: an overview of applications. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307726.

Krivovichev:2022:CIV

[Kri22] Gerasim V. Krivovichev.

Comparison of inviscid and viscid one-dimensional models of blood flow in arteries. Applied Mathematics and Computation, 418 (??):Article 126856, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009395.

Kaczorowski:2024:EHS

[KRNG24] Jerzy Kaczorowski, Waldemar Ratajczak, Peter Nijkamp, and Krzysztof Górnisiewicz.

Economic hierarchical spatial systems — new properties of Löschian numbers. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004885.

Kolluru:2022:SRC

[KRRS22] Ramesh Kolluru, N. V. Raghavendra, S. V. Raghurama Rao, and G. N. Sekhar. Simple and robust contact-discontinuity capturing central algorithms for high speed compressible flows. Applied Mathematics and Computation, 414 (??):Article 126660, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100744X.

Khademi:2020:CMC

[KRS20] Ramin Khademi, Abolhassan Razminia, and Vladimir I. Shiryaev. Conjugatemixed convection of nanofluid flow over an inclined flat plate in porous media. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307532.

Klavzar:2021:SBP

[KRT21] Sandi Klavzar, Freydoon Rahbarnia, and Mostafa Tavakoli. Some binary products and integer linear programming for k-metric dimension of graphs. Applied Mathematics and Computation, 409(??):Article 126420, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005099.

Kolomoitsev:2021:AMQ

[KS21] Yurii Kolomoitsev and Maria Skopina. Approximation by multivariate quasiprojection operators and Fourier multipliers. Applied Mathematics and Computation, 400(??):Article 125955, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000035.

Kaltenbacher:2022:PRI

[KSH22] Stefan Kaltenbacher, Martin Steinberger, and Martin Horn. Pipe roughness identification of water distribution networks: a tensor method. Applied Mathematics and Computation, 413(??):Article 126601, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006858.

Kar:2022:ATF

[KSPC22] Uttam Kumar Kar, Sayantan Sengupta, Shantanu Pramanik, and Soumik Chakraborty. An analytical theory for the forced convection condensation of shear-thinning fluids onto isothermal horizontal surfaces. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321009929.

Kumar:2021:PUA

[KSR21] Sunil Kumar, Sumit, and Higinio Ramos. Parameteruniform approximation on equidistributed meshes for singularly perturbed parabolic reaction-diffusion problems with Robin boundary conditions. Applied Mathematics and Computation, 392(??): Article 125677, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306305.

Kumar:2023:CRC

[KSS23] Vipin Kumar, Gani Stamov, and Ivanka Stamova. Controllability results for a class of piecewise nonlinear impulsive fractional dynamic systems. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006981.

Kong:2024:SPA

[KSW24] Xinlei Kong, Yinjie Song, and Huibin Wu. Structure preserving algorithms with adaptive time step for Birkhoffian systems. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003825.

Kareeva:2023:ISN

[KSZ23] Yulia Kareeva, Artem Sedakov, and Mengke Zhen. Influence in social networks with stubborn agents: From competition to bargaining. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200858X.

Klavzar:2020:LMD

[KT20] Sandi Klavzar and Mostafa Tavakoli. Local metric dimension of graphs: Generalized hierarchical products and some applications. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930668X.

Kabir:2021:RPN

[KT21a] K. M. Ariful Kabir and Jun Tanimoto. The role of pairwise nonlinear evolutionary dynamics in the rock– paper–scissors game with noise. Applied Mathematics and Computation, 394 (??):Article 125767, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307207.

Klavzar:2021:DDC

[KT21b] Sandi Klavzar and Mostafa Tavakoli. Dominated and dominator colorings over (edge) corona and hierarchical products. Applied Mathematics and Computation, 390(??):Article 125647, February 1, 2021.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306019.

Khodakaram-Tafti:2022:DCB

[KTEM22] Amin Khodakaram-Tafti, Homayoun Emdad, and Mojtaba Mahzoon. Dynamical and chaotic behaviors of natural convection flow in semi-annular cylindrical domains using energy-conserving low-order spectral models. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004891.

Kang:2024:CCA

[KTZ24] Kai Kang, Jinyan Tian, and Boyu Zhang. Cooperation and control in asymmetric repeated games. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000614.

Kudryashov:2020:FIG

[Kud20a] Nikolay A. Kudryashov.

First integrals and general solution of the complex Ginzburg–Landau equation. Applied Mathematics and Computation, 386 (??):Article 125407, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303696.

Kudryashov:2020:HDS

[Kud20b] Nikolay A. Kudryashov. Highly dispersive solitary wave solutions of perturbed nonlinear Schrödinger equations. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309646.

Kumar:2023:ACN

[Kum23]  Surendra Kumar.  On

approximate controllability of non-autonomous measure driven systems with non-instantaneous impulse. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007639.

Kumar:2024:NSB

[Kum24]  Raman Kumar.  Nu-

merical solutions for biharmonic interface problems via weak Galerkin finite element methods. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006653.

Kung:2024:REB

[Kun24] Tzu-Liang Kung. Reliability evaluation for bijectionconnected networks based on the super P_{k}-connectivity. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005660.

Kurokawa:2021:DRS

[Kur21] Shun Kurokawa. Disbandment rule sways the evolution of tolerance. Applied Mathematics and Computation, 392(??):Article 125678, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306317.

Kutsenko:2020:MRM

[Kut20] Anton A. Kutsenko. Matrix representations of multidimensional integral and ergodic operators. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308100.

Kong:2023:PGD

[KWH^{+}23] Jiangxu Kong, Yiqiao Wang, Jiacheng Hu, Yang Wang, and Weifan Wang. Plane graphs of diameter two are 2-optimal. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007858.

Kuang:2023:IMH

[KWHS23] Junxian Kuang, Mingliang Wang, Jiajun Han, and Yuhong Sheng. Improved Milne–Hamming method for resolving high-order uncertain differential equations. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003685.

Kamara:2020:ASP

[KWM20a] Abdul A. Kamara, Xiangjun Wang, and Lagès Nadège Mouanguissa. Analytical solution for postdeath transmission model of Ebola epidemics. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307684.

See corrigendum [KWM20b].

Kamara:2020:CAS

[KWM20b] Abdul A. Kamara, Xiangjun Wang, and Lagès Nadège Mouanguissa. Corrigendum to “Analytical solution for post-death transmission model of Ebola epidemics” [Applied Mathematics and Computation 367 (2020) 124776]. Applied Mathematics and Computation, 374 (??):Article 125029, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310215. See

[KWM20a].

Kwon:2022:ESS

[Kwo22a] Kiwoon Kwon. Exact solutions for source localization problem with minimal squared distance error. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002612.

Kwon:2022:UNL

[Kwo22b] Kiwoon Kwon. Uniqueness and nonuniqueness for the L^{1} minimization source localization problem with three measurements. Applied Mathematics and Computation, 413 (??):Article 126649, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007335.

Kong:2022:VIF

[KWW22] Xinlei Kong, Zhongxin Wang, and Huibin Wu. Variational integrators for forced Lagrangian systems based on the local path fitting technique. Applied Mathematics and Computation, 416(??):Article 126739, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008213.

Kang:2023:EAQ

[KWZX23] Hongchao Kang, Ruoxia Wang, Meijuan Zhang, and Chunzhi Xiang. Efficient and accurate quadrature methods of Fourier integrals with a special oscillator and weak singularities. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005665.

Kang:2023:QFM

[KX23] Hongchao Kang and Qi Xu.

Quadrature formulae of many highly oscillatory Fourier-type integrals with algebraic or logarithmic singularities and their error analysis. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322008268.

Kong:2024:CQM

[KXW24] Decheng Kong, Kai Xue, and Ping Wang. Collective queuing motion of self-propelled particles with leadership and experience. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002479.

Kundu:2021:AAT

[KY21] Balaram Kundu and Se-Jin Yook. An accurate approach for thermal analysis of porous longitudinal, spine and radial fins with all nonlinearity effects — analytical and unified assessment. Applied Mathematics and Computation, 402(??):Article 126124, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001727.

Kazemi:2022:AFP

[KY22a] M. Kazemi and A. R. Yaghoobnia. Application of fixed point theorem to solvability of functional stochastic integral equations. Applied Mathematics and Computation, 417(??): Article 126759, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008419.

Kim:2022:DET

[KY22b] Jin Hoe Kim and Sung Jin Yoo. Distributed eventtriggered adaptive outputfeedback formation tracking of uncertain underactuated underwater vehicles in three-dimensional space. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001321.

Kong:2024:NPS

[KY24] Ci Kong and Xingchen Yu.

A note on the periodic solutions to a relativistic Micro-Electro-Mechanical System. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006458.

Karasozen:2022:IDD

[KYU22] Bülent Karasözen, Süleyman Yildiz, and Murat Uzunca. Intrusive and data-driven reduced order modelling of the rotating thermal shallow water equation. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000108.

Kerivin:2022:CSM

[KZ22] Hervé Kerivin and Jinhua Zhao. A computational study on the Maximum-Weight Bounded-Degree Rooted Tree Problem. Applied Mathematics and Computation, 413(??):Article 126623, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007074.

Kang:2023:GES

[KZC23] Yu Kang, Niankun Zhang, and Guoyong Chen. Global exponential stability of impulsive switched positive nonlinear systems with mode-dependent impulses. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005896.

Kanchana:2020:KLI

[KZS20] C. Kanchana, Yi Zhao, and P. G. Siddheshwar. Küppers–Lortz instability in rotating Rayleigh– Bénard convection bounded by rigid/free isothermal boundaries. Applied Mathematics and Computation, 385(??):Article 125406, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303684.

Ke:2022:DRD

[KZWS22] Yue Ke, Linhe Zhu, Peng Wu, and Lei Shi. Dynamics of a reaction-diffusion rumor propagation model with non-smooth control. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005525.

Kitaev:2020:DSI

[KZZ20] Sergey Kitaev, Philip B.

Zhang, and Xutong Zhang. Distributions of several infinite families of mesh patterns. Applied Mathematics and Computation, 372(??): Article 124984, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309762.

Leela-apiradee:2022:TLC

[LaGBT22] Worrawate Leela-apiradee, Artur Gorka, Kanokwan Burimas, and Phantipa Thipwiwatpotjana. Tolerancelocalized and control-localized solutions of interval linear equations system and their application to course assignment problem. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000169.

Linh:2022:FEA

[LAH22] Nguyen Kieu Linh, Phan Thanh An, and Tran Van Hoai. A fast and efficient algorithm for determining the connected orthogonal convex hulls. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002570.

Li:2023:PNP

[LAS23] Jingwang Li, Qing An, and Housheng Su. Proximal nested primal-dual gradient algorithms for distributed constraint-coupled composite optimization. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008694.

Liu:2020:OHS

[LB20] Meng Liu and Chuanzhi Bai. Optimal harvesting of a stochastic mutualism model with regimeswitching. Applied Mathematics and Computation, 373(??):Article 125040, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300096.

Lokeshwar:2022:ENN

[LBJ22] Vikranth Lokeshwar, Vikram Bharadwaj, and Shashi Jain. Explainable neural network for pricing and universal static hedging of contingent claims. Applied Mathematics and Computation, 417(??):Article 126775, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008572.

Li:2023:PPA

[LBX^{+}23] Chuanying Li, Roberto Barrio, Xiong Xiao, Peibing Du, Hao Jiang, Zhe Quan, and Kenli Li. PACF: a precision-adjustable computational framework for solving singular values. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006841.

Lee:2021:APS

[LC21] Eun-Ho Lee and Woocheol Choi. Asymptotic profile of solutions to the heat equation on thin plate with boundary heating. Applied Mathematics and Computation, 408 (??):Article 126356, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004458.

Li:2022:NAC

[LC22a] Qing Li and Huanzhen Chen. Numerical analysis for compact difference scheme of fractional viscoelastic beam vibration models. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002259.

Liu:2022:EIA

[LC22b] Linjie Liu and Xiaojie Chen. Effects of interconnections among corruption, institutional punishment, and economic factors on the evolution of cooperation. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001357.

Linh:2023:WBF

[LC23] Nguyen Ba Hoai Linh and Dao Huy Cuong. A well-balanced finite volume scheme based on planar Riemann solutions for 2D shallow water equations with bathymetry. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003363.

Lin:2022:FDF

[LCC^{+}22] An Lin, Jun Cheng, Jinde Cao, Hailing Wang, and Ahmed Alsaedi. Fault detection filtering for MNNs with dynamic quantization and improved protocol. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005343.

Li:2024:SDV

[LCC^{+}24] Kun Li, Zhiyu Chen, Rui Cong, Jianlei Zhang, and Zhenlin Wei. Simulated dynamics of virus spreading on social networks with various topologies. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000523.

Li:2020:TAN

[LCD^{+}20] Zanhua Li, Xiangyong Chen, Shihong Ding, Yang Liu, and Jianlong Qiu. TCP/AWM network congestion algorithm with funnel control and arbitrary setting time. Applied Mathematics and Computation, 385(??):Article 125410, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303726.

Liu:2020:DRM

[LCH20] Zhang Liu, Ping Chen, and Yijun Hu. On the dual risk model with diffusion under a mixed dividend strategy. Applied Mathematics and Computation, 376(??): Article 125115, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300849.

Li:2022:OMN

[LCJ^{+}22] Wen-Jing Li, Zhi Chen, Ke-Zhong Jin, Jun Wang, Lin Yuan, Changgui Gu, Luo-Luo Jiang, and Matjaž Perc. Options for mobility and network reciprocity to jointly yield robust cooperation in social dilemmas. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005306.

Liu:2022:RHC

[LCL22a] Fan Liu, Mou Chen, and Tao Li. Resilient H_{∞} control for uncertain turbofan linear switched systems with hybrid switching mechanism and disturbance observer. Applied Mathematics and Computation, 413 (??):Article 126597, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006810.

Liu:2022:DOB

[LCL22b] Lijun Liu, Mou Chen, and Tao Li. Disturbance observer-based robust coordination control for unmanned autonomous helicopter slung-load system via coupling analysis method. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002223.

Liu:2022:SSB

[LCL22c] Yufen Liu, Wanrong Cao, and Yuelin Li. Splitstep balanced θ-method for SDEs with non-globally Lipschitz continuous coefficients. Applied Mathematics and Computation, 413 (??):Article 126437, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005269.

Liao:2023:HNK

[LCL23] Jiaqi Liao, Mengyu Cao, and Mei Lu. On the P_{3} -hull numbers of q-Kneser graphs and Grassmann graphs. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006105.

Li:2024:DDP

[LCLH24] Shuai Li, Jinde Cao, Heng Liu, and Chengdai Huang.

Delay-dependent parameters bifurcation in a fractional neural network via geometric methods. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400273X.

Li:2020:ADR

[LCLS20] Yankai Li, Mou Chen, Tao Li, and Peng Shi. Antidisturbance reference mode resilient dynamic output feedback control for turbofan systems. Applied Mathematics and Computation, 378(??):Article 125183, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301521.

Liu:2020:ODP

[LCML20] Guodong Liu, Zhengbo Chang, Xinzhu Meng, and Siyu Liu. Optimality for a diffusive predator–prey system in a spatially heterogeneous environment incorporating a prey refuge. Applied Mathematics and Computation, 384(??):Article 125385, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320303489.

Lee:2022:MPS

[LCS22a] Hsuan-Wei Lee, Colin Cleveland, and Attila Szolnoki. Mercenary punishment in structured populations. Applied Mathematics and Computation, 417(??): Article 126797, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008791.

Liu:2022:RSE

[LCS^{+}22b] Guobao Liu, Xiangyong Chen, Zhongyu Shen, Yajuan Liu, and Xianglei Jia. Reachable set estimation for continuous delayed singularly perturbed systems with bounded disturbances. Applied Mathematics and Computation, 416(??):Article 126751, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100833X.

Lee:2023:RSC

[LCS23] Hsuan-Wei Lee, Colin Cleveland, and Attila Szolnoki. Restoring spatial cooperation with myopic agents in a three-strategy social dilemma. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003958.

Lee:2024:SDI

[LCS24] Hsuan-Wei Lee, Colin Cleveland, and Attila Szolnoki. Suppressing defection by increasing temptation: the impact of smart cooperators on a social dilemma situation. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003254.

Li:2023:IMT

[LCW^{+}23] Qi Li, Le Cheng, Wei Wang, Xianghua Li, Shudong Li, and Peican Zhu. Influence maximization through exploring structural information. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007895.

Lv:2024:WRC

[LCWZ24] Yan Lv, Zhouyang Chen, Tingzeng Wu, and Peng-Li Zhang. On the weighted reverse cover cost of trees and unicyclic graphs with given diameter. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001462.

Liu:2022:FMC

[LCXJ22] Yi Liu, Xiaoqing Chi, Huanying Xu, and Xiaoyun Jiang. Fast method and convergence analysis for the magnetohydrodynamic flow and heat transfer of fractional Maxwell fluid. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003290.

Li:2021:QSQ

[LCXM21] Ruoxia Li, Jinde Cao, Changfeng Xue, and R. Manivannan. Quasi-stability and quasi-synchronization control of quaternion-valued fractional-order discretetime memristive neural networks. Applied Mathematics and Computation, 395(??):

Article 125851, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308043.

Lin:2021:GES

[LCY^{+}21] Dongyuan Lin, Xiaofeng Chen, Guoping Yu, Zhongshan Li, and Yannan Xia. Global exponential synchronization via nonlinear feedback control for delayed inertial memristorbased quaternion-valued neural networks with impulses. Applied Mathematics and Computation, 401 (??):Article 126093, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001417.

Li:2020:RBB

[LCZ^{+}20] Yang Li, Maoyin Chen, Xiaoping Zheng, Zhan Dou, and Yuan Cheng. Relationship between behavior aggressiveness and pedestrian dynamics using behavior-based cellular automata model. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309336.

Li:2023:ORI

[LCZW23] Xin Li, Kaijie Cheng, Liangkuan Zhu, and Guoliang Wei. Outlier-resistant interval observer design for multirate time-delayed systems under the adaptive event-triggered protocols. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008815.

Li:2020:OEI

[LD20a] Mingwu Li and Harry Dankowicz. Optimization with equality and inequality constraints using parameter continuation. Applied Mathematics and Computation, 375(??):Article 125058, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300278.

Li:2020:EAM

[LD20b] Xiaolin Li and Haiyun Dong. Error analysis of the meshless finite point method. Applied Mathematics and Computation, 382(??):Article 125326, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302927.

Lima:2020:MMD

[LD20c] F. N. Lima and J. M.

De Sousa. Mathematical methods of diagonalization of quadratic forms applied to the study of stability of thermodynamic systems. Applied Mathematics and Computation, 378 (??):Article 125176, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301454.

Liu:2020:ASH

[LD20d] Meng Liu and Meiling Deng. Analysis of a stochastic hybrid population model with Allee effect. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319305740.

Lal:2021:DAB

[LD21a] Roshan Lal and Chinika Dangi. Dynamic analysis of bi-directional functionally graded Timoshenko nanobeam on the basis of Eringen’s nonlocal theory incorporating the surface effect. Applied Mathematics and Computation, 395(??): Article 125857, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308109.

Liu:2021:UCA

[LD21b] Lu Liu and Shihong Ding. A unified control approach to finite-time stabilization of SOSM dynamics subject to an output constraint. Applied Mathematics and Computation, 394 (??):Article 125752, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307050.

Li:2023:PCF

[LD23a] Dongdong Li and Jiuxiang Dong. Performanceconstrained fault-tolerant DSC based on reinforcement learning for nonlinear systems with uncertain parameters. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200827X.

Liu:2023:MMI

[LD23b] Guorong Liu and Kecai Deng. The maximum Mostar indices of unicyclic graphs with given diameter. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007081.

Li:2024:DNS

[LDCZ24] Shenshen Li, Haibo Du, Weile Chen, and Wenwu Zhu. Design of nonsmooth consensus protocol for multi-agent systems under DoS attacks. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005295.

Li:2020:ETA

[LDX20] Zeyang Li, Jiuxiang Dong, and Changjiang Xi. Eventtriggered adaptive reliable guaranteed performance control for uncertain nonlinear systems with abrupt non-affine faults. Applied Mathematics and Computation, 380(??): Article 125256, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302253.

Li:2023:BHF

[LDXZ23] Weixun Li, Xiangyang Du, Jingyu Xiao, and Limin Zhang. Bipartite hybrid formation tracking control for heterogeneous multi-agent systems in multi-group cooperativecompetitive networks. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003028.

Li:2020:LIB

[LDZ20] Jiaqi Li, Jianwu Dang, and Jianlei Zhang. Length of information-based bidirectional choice in spatial prisoner’s dilemma. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930829X.

Lee:2021:LPC

[LDZ21] Yan-Li Lee, Qiang Dong, and Tao Zhou. Link prediction via controlling the leading eigenvector. Applied Mathematics and Computation, 411(??):Article 126517, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006068.

Li:2024:IMD

[LDZT24] Ling Li, Gaogao Dong, Huaiping Zhu, and Lixin Tian. Impact of multiple doses of vaccination on epidemiological spread in multiple networks. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000894.

Lee:2020:CCM

[Lee20]  Kiseok Lee.  Cheby-

shev collocation method for the constant mobility Cahn–Hilliard equation in a square domain. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309233.

Lee:2023:VTM

[Lee23] Walter C. Simo Tao Lee.

A variational technique of mollification applied to backward heat conduction problems. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000863.

Lee:2024:SAC

[LePP24] Ho Sub Lee, Chan eun Park, and PooGyeon Park. A search algorithm for the compact uncertainty region of system elements with interval and affine constraint properties. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002844.

Li:2021:LPS

[LF21] Yang Li and Minfu Feng. A local projection stabilization virtual element method for convection-diffusion-reaction equation. Applied Mathematics and Computation, 411(??):Article 126536, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321006202.

Liu:2024:AOE

[LF24] Ying Liu and Aili Fang.

Analysis of opinion evolution based on non-Bayesian social learning. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005684.

Li:2022:LEC

[LFA22] Lin Lin Li, Cláudia Pio Ferreira, and Bedreddine Ainseba. Local exact controllability of an age structured problem modelling phenotypic plasticity in mosquito behaviour. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001114.

Liao:2023:IIU

[LFLH23] Shenghai Liao, Shujun Fu, Yuliang Li, and Hongbin Han. Image inpainting using non-convex low rank decomposition and multidirectional search. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002175.

Lang:2021:AUL

[LFLL21] Rongling Lang, Ya Fan, Guoliang Liu, and Guodong Liu. Analysis of unlabeled lung sound samples using semi-supervised convolutional neural networks. Applied Mathematics and Computation, 411 (??):Article 126511, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006007.

Liu:2021:DAM

[LFLP21] Li Liu, Zhenbin Fan, Gang Li, and Sergey Piskarev. Discrete almost maximal regularity and stability for fractional differential equations in L^{p}([0, 1], Ω). Applied Mathematics and Computation, 389(??):Article 125574, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305300.

Liu:2022:PPD

[LFM^{+}22] Wei Liu, Shiqi Fei, Qian Ma, Huanyu Zhao, and Shengyuan Xu. Prescribed performance dynamic surface control for nonlinear systems subject to partial and full state constraints. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003927.

Lv:2024:DDE

[LFY^{+}24] Xijian Lv, Dongmei Fan, Junxian Yang, Qiang Li, and Li Zhou. Delay differential equation modeling of social contagion with higher-order interactions. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006331.

Lteif:2022:NCH

[LG22] Ralph Lteif and Stéphane Gerbi. A new class of higher-ordered/extended Boussinesq system for efficient numerical simulations by splitting operators. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004477.

Li:2023:PTE

[LG23] Hui Li and Yunshu Gao. Placing two edge-disjoint copies of a tree into a bipartite graph. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300214X.

Li:2024:ESR

[LG24] Xin Li and Ji-Ming Guo. The effect on the spectral radius by attaching a pendant starlike tree. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002790.

Li:2024:ORS

[LGC^{+}24a] Lili Li, Qingjun Guo, Bingqi Chen, Wei Yue, Zhongchang Liu, and Hong Sang. Output regulation for switched systems via dual distributed event-triggered mechanisms under multiple deception attacks. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006343.

Li:2024:WGF

[LGC24b] Wenjuan Li, Fuzheng Gao, and Jintao Cui. A weak Galerkin finite element method for nonlinear convection–diffusion equation. Applied Mathematics and Computation, 461 (??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323004848.

Liu:2021:IUR

[LGD^{+}21] Yifan Liu, Yini Geng, Chunpeng Du, Kaipeng Hu, Chen Shen, Riccardo Pansini, and Lei Shi. The interface of unidirectional rewards: Enhanced cooperation within interdependent networks. Applied Mathematics and Computation, 402(??):Article 126151, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001995.

Liu:2022:CBS

[LGK^{+}22] Mingyang Liu, Guangjun Gao, Boo Cheong Khoo, Zhenhu He, and Chen Jiang.

A cell-based smoothed finite element model for non-Newtonian blood flow. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322005549.

Li:2020:SSC

[LGL20] Wenhua Li, Xingbao Gao, and Ruoxia Li. Stability and synchronization control of inertial neural networks with mixed delays. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307714.

Lang:2020:BTP

[LGQ20] Chao Lang, Rong Gao, and Chujun Qiu. On block triangular preconditioned iteration methods for solving the Helmholtz equation. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306873.

Liu:2022:DPS

[LGS22] Haicheng Liu, Bin Ge, and Jihong Shen. Dynamics of periodic solutions in the reaction–diffusion glycolysis model: Mathematical mechanisms of Turing pattern formation. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003988.

Li:2022:DET

[LGX22] Min Li, Jian Guo, and Zhengrong Xiang. Dynamic event-triggered control design for a class of pnormal nonlinear time-delay systems with actuator failures. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000285.

Li:2020:ISM

[LH20a] Hongshan Li and Zhongyi Huang. An iterative splitting method for pricing European options under the Heston model. Applied Mathematics and Computation, 387(??):Article 125424, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303854.

Li:2020:PEM

[LH20b] Min Li and Chengming Huang. Projected Euler– Maruyama method for stochastic delay differential equations under a global monotonicity condition. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307258.

Li:2022:ANN

[LH22] Yixin Li and Xianliang Hu. Artificial neural network approximations of Cauchy inverse problem for linear PDEs. Applied Mathematics and Computation, 414 (??):Article 126678, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007621.

Liu:2024:BCE

[LH24a] Jinfeng Liu and Fei Huang. Bipartite cacti with extremal matching energy. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400376X.

Liu:2024:ATC

[LH24b] Yanli Liu and Li-Ying Hao.

Adaptive tracking control for constrained nonlinear nonstrict-feedback switched stochastic systems with unknown control directions. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001383.

Li:2021:CPE

[LHC21] Min Li, Chengming Huang, and Ziheng Chen. Compensated projected Euler– Maruyama method for stochastic differential equations with superlinear jumps. Applied Mathematics and Computation, 393(??):Article 125760, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030713X.

Li:2022:DFT

[LHC^{+}22] Jiaxing Li, Jun Hu, Jun Cheng, Yunliang Wei, and Hui Yu. Distributed filtering for time-varying statesaturated systems with packet disorders: an eventtriggered case. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004854.

Lai:2023:HEM

[LHET23] Qiang Lai, Genwen Hu, Uğur Erkan, and Abdurrahim Toktas. Highefficiency medical image encryption method based on 2D logistic–Gaussian hyperchaotic map. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008062.

Lu:2022:SDF

[LHFH22] Xiaoli Lu, Pengzhan Huang, Xinlong Feng, and Yinnian He. A stabilized difference finite element method for the 3D steady Stokes equations. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003447.

Liu:2021:CID

[LHFW21] An Liu, Xia Huang, Yingjie Fan, and Zhen Wang. A control-interval-dependent functional for exponential stabilization of neural networks via intermittent sampled-data control. Applied Mathematics and Computation, 411 (??):Article 126494, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100583X.

Luo:2024:ECG

[LHH24] Yijun Luo, Changwei Huang, and Wenchen Han. The evolution of cooperation and global synchronization in the evolutionary Kuramoto dilemma combined with the prisoner’s dilemma. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400434X.

Liu:2024:MDS

[LHL24] Weijun Liu, Yueli Han, and Lu Lu. On minimal directed strongly regular Cayley graphs over dihedral groups.

Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001176.

Liao:2023:CEU

[LHQ^{+}23] Hui-Min Liao, Qing-Yi Hao, Jia-Li Qian, Chao-Yun Wu, Ning Guo, and Xiang Ling. Cooperative evolution under the joint influence of local popularity and global popularity. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200707X.

Liu:2021:LPR

[LHS21] Lu Liu, Wei Huang, and Li Shen. Learning performance of regularized regression with multiscale kernels based on Markov observations. Applied Mathematics and Computation, 409 (??):Article 126386, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004756.

Li:2021:GASa

[LHW21] Wenxiu Li, Lihong Huang, and Jiafu Wang. Global asymptotical stability and sliding bifurcation analysis of a general Filippovtype predator–prey model with a refuge. Applied Mathematics and Computation, 405(??):Article 126263, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003520.

Li:2023:RAC

[LHW^{+}23] Mingjia Li, Linmin Hu, Shaomin Wu, Bing Zhao, and Yan Wang. Reliability assessment for consecutivek-out-of-n: F retrial systems under Poisson shocks. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000826.

Li:2022:LFB

[LHX^{+}22] Baoxing Li, Tao Han, Bo Xiao, Xi-Sheng Zhan, and Huaicheng Yan. Leaderfollowing bipartite consensus of multiple uncertain Euler–Lagrange systems under deception attacks. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003010.

Li:2024:SOL

[LHYL24] Yaoda Li, Zhibin Han, Yajun Yin, and Wen Li. A second-order linear and unconditional energystable scheme for Swift– Hohenberg equations. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400211X.

Li:2021:NSM

[LHZ^{+}21a] Hong-Li Li, Cheng Hu, Long Zhang, Haijun Jiang, and Jinde Cao. Non-separation method-based robust finitetime synchronization of uncertain fractional-order quaternion-valued neural networks. Applied Mathematics and Computation, 409(??):Article 126377, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004665.

Lin:2021:NNC

[LHZ^{+}21b] Jie Lin, Ting-Zhu Huang, Xi-Le Zhao, Tian-Hui Ma, Tai-Xiang Jiang, and Yu-Bang Zheng. A novel nonconvex low-rank tensor approximation model for hyperspectral image restoration. Applied Mathematics and Computation, 408 (??):Article 126342, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004318.

Liu:2021:NAB

[LHZ21c] Hongyan Liu, Jin Huang, and Wei Zhang. Numerical algorithm based on extended barycentric Lagrange interpolant for two dimensional integrodifferential equations. Applied Mathematics and Computation, 396(??):Article 125931, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308845.

Liu:2021:RSB

[LHZ21d] Siyan Liu, Rong-Xia Hao, and Cun-Quan Zhang. Rotation snark, Berge– Fulkerson conjecture and Catlin’s 4-flow reduction. Applied Mathematics and Computation, 410(??):Article 126441, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321005300.

Li:2023:DME

[LHZ^{+}23] Baihua Li, Shaoyu Hu, Qishui Zhong, Kaibo Shi, and Shouming Zhong. Dynamic memory event-triggered proportional-integral-based H_{∞} load frequency control for multi-area wind power systems. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002394.

Lian:2021:QFF

[Lia21]  Pan Lian.  Quaternion

and fractional Fourier transform in higher dimension. Applied Mathematics and Computation, 389 (??):Article 125585, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305415.

Liu:2021:GME

[Liu21] Z. Liu. Generalized moment estimation for uncertain differential equations. Applied Mathematics and Computation, 392(??): Article 125724, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306779.

Liu:2022:MFI

[Liu22a] Hong-Zhun Liu. A modification to the first integral method and its applications. Applied Mathematics and Computation, 419(??): Article 126855, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009383.

Liu:2022:EPD

[Liu22b] Yue Liu. Extinction, persistence and density function analysis of a stochastic two-strain disease model with drug resistance mutation. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004672.

Liu:2024:ORF

[Liu24] Jianqiang Liu. Orthonormal rational functions on a semiinfinite interval. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003485.

Li:2023:BOF

[LJ23] Ruicheng Li and Feng-Fei Jin. Boundary output feedback stabilization for a cascaded wave PDE–ODE system with velocity recirculation and matched disturbance. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008955.

Laxmi:2024:NCE

[LJAM24] Parik Laxmi, Shilpi Jain, Praveen Agarwal, and Gradimir V. Milovanović. Numerical calculation of the extension of k-beta function and some new extensions by using two parameter k-Mittag-Leffler function. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003187.

Liu:2020:SDS

[LJH^{+}20] Qun Liu, Daqing Jiang, Tasawar Hayat, Ahmed Alsaedi, and Bashir Ahmad. Stationary distribution of a stochastic cholera model between communities linked by migration. Applied Mathematics and Computation, 373 (??):Article 125021, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310136.

Luo:2021:ICP

[LJHL21] Xiao-Feng Luo, Zhen Jin, Daihai He, and Li Li. The impact of contact patterns of sexual networks on Zika virus spread: a case study in Costa Rica. Applied Mathematics and Computation, 393(??):Article 125765, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307189.

Li:2021:OCC

[LJL21] Jun Li, Lianghao Ji, and Huaqing Li. Optimal consensus control for unknown second-order multi-agent systems: Using modelfree reinforcement learning method. Applied Mathematics and Computation, 410(??):Article 126451, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005403.

Li:2022:ABA

[LJN^{+}22] Xiaoyu Li, Danyang Jia, Xiaotong Niu, Chen Liu, Peican Zhu, Dujuan Liu, and Chen Chu. Abilitybased asymmetrical fitness calculation promotes cooperation in spatial prisoner’s dilemma game. Applied Mathematics and Computation, 412(??):Article 126572, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006561.

Li:2021:LMM

[LJP21] Wen-Jing Li, Luo-Luo Jiang, and Matjaz Perc. A limited mobility of minorities facilitates cooperation in social dilemmas. Applied Mathematics and Computation, 391 (??):Article 125705, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306585.

Liu:2020:ESU

[LJR20] Run-Ran Liu, Chun-Xiao Jia, and Zhihai Rong. Effects of strategy-updating cost on evolutionary spatial prisoner’s dilemma game. Applied Mathematics and Computation, 386 (??):Article 125445, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304069.

Li:2021:WNC

[LJW^{+}21] Shudong Li, Laiyuan Jiang, Xiaobo Wu, Weihong Han, Dawei Zhao, and Zhen Wang. A weighted network community detection algorithm based on deep learning. Applied Mathematics and Computation, 401 (??):Article 126012, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000606.

Lu:2022:RAE

[LJW^{+}22] Shao-Yu Lü, Xiao-Zheng Jin, Xiao-Ming Wu, Li-Jian Ding, and Jing Chi. Robust adaptive eventtriggered fault-tolerant control for time-varying systems against perturbations and faulty actuators. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200217X.

Lu:2021:RAE

[LJWD21] Shao-Yu Lü, Xiao-Zheng Jin, Hai Wang, and Chao Deng. Robust adaptive estimation and tracking control for perturbed cyberphysical systems against denial of service. Applied Mathematics and Computation, 404(??):Article 126255, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003453.

Li:2022:MOR

[LJY22] Yanpeng Li, Yaolin Jiang, and Ping Yang. Model order reduction of port-Hamiltonian systems with inhomogeneous initial conditions via approximate finite-time Gramians. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000455.

Li:2022:AOR

[LJZL22] Lili Li, Mingzhe Ju, Ying Zhao, and Tianhe Liu. Almost output regulation for switched linear systems with bumpless transfer control under destabilizing behaviors. Applied Mathematics and Computation, 412 (??):Article 126578, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006627.

Lal:2021:AFB

[LK21a] Shyam Lal and Priya Kumari. Approximation of functions with bounded derivative and solution of Riccati differential equations by Jacobi wavelet operational matrix. Applied Mathematics and Computation, 394(??):Article 125834, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307876.

Liu:2021:AIP

[LK21b] Xiaonan Liu and Yonggui Kao. Aperiodically intermittent pinning outer synchronization control for delayed complex dynamical networks with reaction– diffusion terms. Applied Mathematics and Computation, 410(??):Article 126406, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004951.

Lee:2023:CAD

[LKK^{+}23] Seungpeel Lee, Jina Kim, Dongjae Kim, Ki Joon Kim, and Eunil Park. Computational approaches to developing the implicit media bias dataset: Assessing political orientations of nonpolitical news articles. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003880.

Lee:2022:GMI

[LKP22] Jun Hui Lee, Jung Hoon Kim, and PooGyeon Park. A generalized multipleintegral inequality based on free matrices: Application to stability analysis of time-varying delay systems. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003629.

Li:2021:GASb

[LKSS21] Hui Li, YongGui Kao, Ivanka Stamova, and Chuntao Shao. Global asymptotic stability and Sasymptotic ω-periodicity of impulsive non-autonomous fractional-order neural networks. Applied Mathematics and Computation, 410 (??):Article 126459, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005488.

Liu:2020:CCB

[LKX20] Hongjun Liu, Abdurahman Kadir, and Chengbo Xu. Cryptanalysis and constructing S-Box based on chaotic map and backtracking. Applied Mathematics and Computation, 376(??): Article 125153, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301223.

Lin:2020:OCM

[LL20a] Lan Lin and Yixun Lin. Optimality computation of the minimum stretch spanning tree problem. Applied Mathematics and Computation, 386(??):Article 125502, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304604.

Liu:2020:HBC

[LL20b] Yujuan Liu and Qiong Lu. Hopf bifurcations in 3D competitive system with mixing exponential and rational growth rates. Applied Mathematics and Computation, 378 (??):Article 125209, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301788.

Li:2021:PSC

[LL21] Suhua Li and Yaotang Li. Programmable sufficient conditions for the strong ellipticity of partially symmetric tensors. Applied Mathematics and Computation, 403(??):Article 126134, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100182X.

Lee:2022:NSA

[LL22a] Dongsun Lee and Chaeyoung Lee. Numerical solutions of the Allen– Cahn equation with the p-Laplacian. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005094.

Li:2022:NFA

[LL22b] Qiang Li and Jinling Liang.

Non-fragile asynchronous state estimation for Markovian switching CVNs with partly accessible mode detection: the discrete-time case. Applied Mathematics and Computation, 412 (??):Article 126583, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006676.

Li:2023:LWS

[LL23a] Ang Li and Jiequan Li.

Lax–Wendroff solvers-based Hermite reconstruction for hyperbolic problems. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300084X.

Li:2023:SED

[LL23b] Meiyu Li and Jinling Liang.

State estimation for 2-D uncertain systems with redundant channels and deception attacks: a set-membership method. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003594.

Liao:2023:PSG

[LL23c] Qianfen Liao and Weijun Liu. GCI -property of some groups. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200649X.

Liu:2023:PPB

[LL23d] Hui Liu and Xiaohua Li. A prescribed-performance-based adaptive finite-time tracking control scheme circumventing the dependence on the system initial condition. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000814.

Lu:2023:DEI

[LL23e] Wen Lu and Shu Liang. Direct emotional interaction in prisoner’s dilemma game. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300423X.

Liu:2023:DSR

[LLCL23] Kang Liu, Dan Li, Yuanyuan Chen, and Yuanjing Li. Distance spectral radius of the complements of trees with fixed parameters. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008578.

Liu:2023:GTA

[LLCS23] Chunxia Liu, Kaihong Lu, Xiaojie Chen, and Attila Szolnoki. Game-theoretical approach for task allocation problems with constraints. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004204.

Liu:2021:ABA

[LLD21] Jiamin Liu, Zhao-Yan Li, and Feiqi Deng. Asymptotic behavior analysis of Markovian switching neutral-type stochastic time-delay systems. Applied Mathematics and Computation, 404(??): Article 126205, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002952.

Ladecky:2023:OPF

[LLF^{+}23] Martin Ladecký, Richard J. Leute, Ali Falsafi, Ivana Pultarová, Lars Pastewka, Till Junge, and Jan Zeman. An optimal preconditioned FFT-accelerated finite element solver for homogenization. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000048.

Liu:2020:SSI

[LLFZ20] Zhongyun Liu, Zhen Li, Carla Ferreira, and Yulin Zhang. Stationary splitting iterative methods for the matrix equation AXB = C. Applied Mathematics and Computation, 378 (??):Article 125195, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301648.

Li:2022:DSL

[LLG22] Huijuan Li, Wuquan Li, and Jianzhong Gu. Decentralized stabilization of largescale stochastic nonlinear systems with time-varying powers. Applied Mathematics and Computation, 418 (??):Article 126787, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008699.

Li:2024:SCH

[LLG^{+}24] Jiachen Li, Wenjie Li, Feng Gao, Meng Cai, Zengping Zhang, Xiaoyang Liu, and Wei Wang. Social contagions on higher-order community networks. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002935.

Luo:2021:FCM

[LLH21] Cheng Luo, Bao-Qing Liu, and Hu-Shuang Hou. Fractional chaotic maps with q-deformation. Applied Mathematics and Computation, 393(??):Article 125759, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307128.

Li:2021:CSC

[LLL21] Yalu Li, Haitao Li, and Yuanyuan Li. Constrained set controllability of logical control networks with state constraints and its applications. Applied Mathematics and Computation, 405(??): Article 126259, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003489.

Lang:2023:SMP

[LLL^{+}23a] Rongling Lang, Hui Lei, Siyan Li, Xiaopan Lian, and Susu Wang. Some Maderperfect graph classes. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001376.

Liang:2023:LHC

[LLL23b] Lingmei Liang, Fengxia Liu, and Hong-Jian Lai.

On list r-hued coloring of outer-1-planar graphs. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007299.

Li:2023:NIE

[LLLC23] Yancheng Li, Cong Liu, Wei Li, and Yingbin Chai. Numerical investigation of the element-free Galerkin method (EFGM) with appropriate temporal discretization techniques for transient wave propagation problems. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008232.

Li:2023:RMS

[LLLL23] Xi Li, Yang Liu, Jungang Lou, and Jianquan Lu. Robust minimal strong reconstructibility problem of Boolean control networks. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003788.

Liu:2020:OFS

[LLLS20] Cheng-Qian Liu, Xiao-Jian Li, Yue Long, and Jie Sun. Output feedback secure control for cyber-physical systems against sparse sensor attacks. Applied Mathematics and Computation, 384 (??):Article 125384, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303477.

Li:2023:ADA

[LLP23] Jin Li, Zhilin Li, and Kejia Pan. Accurate derivatives approximations and applications to some elliptic PDEs using HOC methods. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004344.

Li:2021:WBA

[LLQG21] Gang Li, Jiaojiao Li, Shouguo Qian, and Jinmei Gao. A well-balanced ADER discontinuous Galerkin method based on differential transformation procedure for shallow water equations. Applied Mathematics and Computation, 395(??): Article 125848, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308018.

Liu:2024:MBE

[LLQW24] Xinghua Liu, Yuru Liang, Siwei Qiao, and Peng Wang. Memory-based event-triggered fault-tolerant load frequency control of multi-area power systems with electric vehicles. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001085.

Li:2021:UCS

[LLS21a] Zhenzhen Li, Minghao Li, and Dongyang Shi. Unconditional convergence and superconvergence analysis for the transient Stokes equations with damping. Applied Mathematics and Computation, 389(??):Article 125572, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305282.

Lopushansky:2021:NIP

[LLS21b] Andriy Lopushansky, Oleh Lopushansky, and Sergii Sharyn. Nonlinear inverse problem of control diffusivity parameter determination for a space-time fractional diffusion equation. Applied Mathematics and Computation, 390 (??):Article 125589, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305440.

Liang:2022:NIB

[LLS^{+}22] Juan Liang, Chen Liu, Gui-Quan Sun, Li Li, Lai Zhang, Meiting Hou, Hao Wang, and Zhen Wang. Nonlocal interactions between vegetation induce spatial patterning. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001473.

Li:2023:ETG

[LLSY23] Shuchao Li, Xin Liu, Wanting Sun, and Lixia Yan. Extremal trees of a given degree sequence or segment sequence with respect to average Steiner 3-eccentricity. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006300.

Luo:2020:NAG

[LLT^{+}20] Jinnan Luo, Xinzhi Liu, Wenhong Tian, Shouming Zhong, Kaibo Shi, and Jun Cheng. A new approach to generalized dissipativity analysis for fuzzy systems with coupling memory sampled-data control. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307660.

Lachowicz:2022:AKE

[LLT22] Mirosław Lachowicz, Henryk Leszczyński, and Krzysztof A. Topolski. Approximations of kinetic equations of swarm formation: Convergence and exact solutions. Applied Mathematics and Computation, 417(??): Article 126778, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008602.

Li:2020:MRG

[LLW20] Yan Li, Yusheng Li, and Ye Wang. Minimal Ramsey graphs on deleting stars for generalized fans and books. Applied Mathematics and Computation, 372(??): Article 125006, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309981.

Liu:2024:EMP

[LLW24] Aixin Liu, Haitao Li, and Lin Wang. Exploring multipotential games in strategic form: a graph theoretic approach. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001486.

Liu:2023:NFP

[LLWK23] Jason J. R. Liu, James Lam, Xiaomei Wang, and Ka-Wai Kwok. Non-fragile PD control of linear time-delay positive discrete-time systems. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001856.

Lou:2022:ROF

[LLWW22] Fangzheng Lou, Feng Li, Jing Wang, and Yan Wang. Reliable output feedback control for persistent dwelltime switched piecewiseaffine systems against deception attacks. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002053.

Liao:2022:PRA

[LLWY22] Wei Liao, Taotao Liang, Xiaohui Wei, and Qiaozhi Yin. Probabilistic reachavoid problems in nondeterministic systems with timevarying targets and obstacles. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001400.

Liao:2023:DRM

[LLWY23] Wei Liao, Taotao Liang, Chen Wang, and Dayong Yang. A dimensionality reduction method for computing reachable tubes based on piecewise pseudo-time dependent Hamilton–Jacobi equation. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007640.

Liu:2021:SFB

[LLX21] Jingmei Liu, Xiao Liang, and Juanjuan Xu. Solution to the forward and backward stochastic difference equations with asymmetric information and application. Applied Mathematics and Computation, 390 (??):Article 125594, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030549X.

Li:2020:TNS

[LLY20] Jia-Yun Li, Sha-Sha Li, and Jian-Hua Yin. On Turán number for S_{ℓ_{1}} ∪ S_{ℓ_{2}}. Applied Mathematics and Computation, 385 (??):Article 125400, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303623.

Li:2021:NRT

[LLY^{+}21] Jiahao Li, Yu Liu, Jinyong Yu, Yiming Sun, and Mengmeng Liu. A new result of terminal sliding mode finite-time state and fault estimation for a class of descriptor switched systems. Applied Mathematics and Computation, 402 (??):Article 126085, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001338.

Li:2022:NRS

[LLY22a] Jiahao Li, Yu Liu, and Jinyong Yu. A new result on semi-synchronous event-triggered backstepping robust control for a class of non-Lipschitzian networked systems. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001138.

Li:2022:GAT

[LLY^{+}22b] Yong Li, Xianning Liu, Yiyi Yuan, Jiang Li, and Lianwen Wang. Global analysis of tuberculosis dynamical model and optimal control strategies based on case data in the United States.

Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000698.

Li:2023:ASS

[LLY^{+}23] Yang Li, Dejun Liu, Zhexu Yin, Yun Chen, and Jin Meng. Adaptive selection strategy of shape parameters for LRBF for solving partial differential equations. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007172.

Liu:2020:MSM

[LLYQ20] Guoli Liu, Feng Li, Xianyan Yang, and Shuang Qiu. The multi-stage multiproduct batch-sizing problem in the steel industry. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308227.

Lang:2020:GBS

[LLZ^{+}20a] Rongling Lang, Ruibo Lu, Chenqian Zhao, Honglei Qin, and Guodong Liu. Graph-based semisupervised one class support vector machine for detecting abnormal lung sounds. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319304667.

Li:2020:AAD

[LLZ20b] Min Li, Ming Liu, and Yingchun Zhang. Asynchronous adaptive dynamic output feedback sliding mode control for singular Markovian jump systems with actuator faults and uncertain transition rates. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309506.

Li:2020:LST

[LLZ20c] Yipeng Li, Xiaogang Liu, and Shenggui Zhang. Laplacian state transfer in Qgraph. Applied Mathematics and Computation, 384 (??):Article 125370, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303349.

Li:2021:FTS

[LLZ21] Xueling Li, Xiangze Lin, and Yun Zou. Finitetime stabilization of highorder output-constrained switched systems via state feedback. Applied Mathematics and Computation, 403(??):Article 125935, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320308882.

Lei:2022:SFD

[LLZ22a] Yanfang Lei, Junmin Li, and Ailiang Zhao. Spatiotemporal fault detection, estimation and control for nonlinear reaction– diffusion equations. Applied Mathematics and Computation, 418(??):Article 126859, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009425.

Liu:2022:ARN

[LLZ22b] Huiqing Liu, Mei Lu, and Shunzhe Zhang. Anti-Ramsey numbers for cycles in the generalized Petersen graphs. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003514.

Liu:2022:ATA

[LLZ22c] Jinpeng Liu, Yuanyuan Liu, and Yiqiang Q. Zhao. Augmented truncation approximations to the solution of Poisson’s equation for Markov chains. Applied Mathematics and Computation, 414(??):Article 126610, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006949.

LIU:2023:EBB

[LLZ23] Huiqing Liu, Mei Lu, and Jing Zhao. Extremal (balanced) blow-ups of trees with respect to the signless Laplacian index. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002679.

Lv:2024:FCE

[LLZ24] Shaojie Lv, Jiaying Li, and Changheng Zhao. Fixation of cooperation in evolutionary games with environmental feedbacks. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004181.

Li:2024:IKF

[LLZB24] Xin Li, Anzhi Lei, Liangkuan Zhu, and Mingfei Ban. Improving Kalman filter for cyber physical systems subject to replay attacks: an attack-detection-based compensation strategy. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006136.

Liu:2021:TCT

[LLZC21] Juan Liu, Qinghai Liu, Xindong Zhang, and Xing Chen. Trail-connected tournaments. Applied Mathematics and Computation, 389(??):Article 125595, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305506.

Liu:2022:HID

[LLZL22] Juan Liu, Shupeng Li, Xindong Zhang, and Hong-Jian Lai. Hamiltonian index of directed multigraph. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001588.

Liang:2022:ESD

[LLZW22] Bo Liang, Qingchun Li, Jihong Zhang, and Ying Wang. Existence of solutions to a doubly degenerate fourth-order parabolic equation. Applied Mathematics and Computation, 413(??):Article 126650, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007347.

Lei:2020:SFT

[LM20a] Yafei Lei and Jixiang Meng. Structure faulttolerance of arrangement graphs. Applied Mathematics and Computation, 381 (??):Article 125287, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302563.

Li:2020:EGD

[LM20b] Bang-Qing Li and Yu-Lan Ma. Extended generalized Darboux transformation to hybrid rogue wave and breather solutions for a nonlinear Schrödinger equation. Applied Mathematics and Computation, 386 (??):Article 125469, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304288.

Liu:2020:CVB

[LM20c] Xianggang Liu and Li Ma. Chaotic vibration, bifurcation, stabilization and synchronization control for fractional discrete-time systems. Applied Mathematics and Computation, 385 (??):Article 125423, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303842.

Lv:2020:MCA

[LM20d] Changqing Lv and Changfeng Ma. A modified CG algorithm for solving generalized coupled Sylvester tensor equations. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306915.

Liu:2021:REC

[LM21] Xuemei Liu and Jixiang Meng. The krestricted edge-connectivity of the data center network DCell. Applied Mathematics and Computation, 396(??): Article 125941, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308948.

Liu:2022:CEP

[LM22] Xifu Liu and Changxin Mo.

Calculating C-eigenpairs of piezoelectric-type tensors via a Z-eigenpair method. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002089.

Li:2023:ANN

[LM23] Yao Li and Caleb Meredith. Artificial neural network solver for timedependent Fokker–Planck equations. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003545.

Li:2024:DET

[LM24] Xin Li and Xiaowu Mu. Dynamic event-triggered fuzzy control for nonlinear singular semi-Markovian jump systems via state decomposition approach. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006355.

Lukacova-Medvidova:2023:CDG

[LMÖ23] Mária Lukácová-Medvid’ová and Philipp Öffner. Convergence of discontinuous Galerkin schemes for the Euler equations via dissipative weak solutions. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005823.

Liu:2023:CCD

[LMS23] Xuemei Liu, Jixiang Meng, and Eminjan Sabir. Component connectivity of the data center network DCell.

Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008906.

Lyche:2024:LSS

[LMS24] Tom Lyche, Carla Manni, and Hendrik Speleers. A local simplex spline basis for C^{3} quartic splines on arbitrary triangulations. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300499X.

Lachowicz:2024:PET

[LMT24] Mirosław Lachowicz, Milena Matusik, and Krzysztof A. Topolski. Population of entities with three individual states and asymmetric interactions. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005647.

Li:2020:CEP

[LMX20] Chengliang Li, Changfeng Ma, and Xiaofang Xu.

A class of efficient parameterized shift-splitting preconditioners for block two-by-two linear systems. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308070.

Li:2022:FVI

[LMX22] Xiao Li, Xiaoying Meng, and Bo Xiong. A fractional variational image denoising model with two-component regularization terms. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002521.

Lattanzio:2020:NID

[LMZ20] Corrado Lattanzio, Pierangelo Marcati, and Delyan Zhelyazov. Numerical investigations of dispersive shocks and spectral analysis for linearized quantum hydrodynamics. Applied Mathematics and Computation, 385 (??):Article 125450, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304112.

Liu:2021:ENC

[LMZ^{+}21] Jingjing Liu, Ruijie Ma, Xiaoyang Zeng, Wanquan Liu, Mingyu Wang, and Hui Chen. An efficient nonconvex total variation approach for image deblurring and denoising. Applied Mathematics and Computation, 397(??):Article 125977, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000254.

Liu:2022:CSS

[LNC22] Yulong Liu, Yuanling Niu, and Xiujun Cheng. Convergence and stability of the semi-tamed Milstein method for commutative stochastic differential equations with non-globally Lipschitz continuous coefficients. Applied Mathematics and Computation, 414 (??):Article 126680, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007645.

Li:2024:FTC

[LNC24] Hongchao Li, Guowei Niu, and Yining Chen. Fixedtime consensus of leaderfollowing multi-agent systems subject to failed follower: Reconstructed topology approach. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004168.

Lin:2023:ASN

[LNO23] Wenting Lin, Xin Ning, and Zhonghui Ou. Analytical solutions of the nitrogen uptake model with Michaelis– Menten flux. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006440.

Li:2021:RCS

[LNS^{+}21] Xiaoqing Li, Sing Kiong Nguang, Kun She, Jun Cheng, and Shouming Zhong. Resilient controller synthesis for Markovian jump systems with probabilistic faults and gain fluctuations under stochastic sampling operational mechanism. Applied Mathematics and Computation, 392(??): Article 125623, March 1, 2021. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305774.

Liu:2021:APP

[LNZ21] Fengxia Liu, Xirinay Nurmamat, and Panpan Zhang. Arbitrary partitionability of product graphs. Applied Mathematics and Computation, 408(??):Article 126219, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100309X.

Liu:2022:AFT

[LNZ^{+}22] Shanlin Liu, Ben Niu, Guangdeng Zong, Xudong Zhao, and Ning Xu. Adaptive fixed-time hierarchical sliding mode control for switched underactuated systems with deadzone constraints via eventtriggered strategy. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200515X.

Loczi:2022:GHP

[Lóc22] Lajos Lóczi. Guaranteedand high-precision evaluation of the Lambert W function. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004805.

Leitao:2020:MFC

[LOG20] Álvaro Leitao and Luis Ortiz-Gracia. Model-free computation of risk contributions in credit portfolios. Applied Mathematics and Computation, 382 (??):Article 125351, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303155.

Lopez:2021:MMJ

[LOS21] Oscar López, Gerardo Oleaga, and Alejandra Sánchez. Markov-modulated jump-diffusion models for the short rate: Pricing of zero coupon bonds and convexity adjustment. Applied Mathematics and Computation, 395(??):Article 125854, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308079.

Luengo:2023:FAS

[LOVHC23] Elena Almaraz Luengo, Bittor Alaña Olivares, Luis Javier García Villalba, and Julio Hernandez-Castro. Further analysis of the statistical independence of the NIST SP 800-22 randomness tests. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003910.

Lanconelli:2022:NMS

[LP22] Alberto Lanconelli and Berk Tan Perçin. On a new method for the stochastic perturbation of the disease transmission coefficient in SIS models. Applied Mathematics and Computation, 413(??):Article 126600, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006846.

Lee:2022:ROB

[LPK22] Won Il Lee, Bum Yong Park, and Sung Hyun Kim. Relaxed observer-based stabilization and dissipativity conditions of T–S fuzzy systems with nonhomogeneous Markov jumps via non-PDC scheme. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200529X.

Lee:2022:LCS

[LPKC22] S. H. Lee, M. J. Park, O. M. Kwon, and S. G. Choi. Less conservative stability criteria for general neural networks through novel delay-dependent functional. Applied Mathematics and Computation, 420(??): Article 126886, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009693.

Liu:2021:GAE

[LPL21a] Chang Liu, Yingui Pan, and Jianping Li. On the geometric-arithmetic Estrada index of graphs. Applied Mathematics and Computation, 391(??):Article 125700, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306536.

Lu:2021:BGE

[LPL21b] Jian Lu, Xiang-Feng Pan, and Huiqing Liu. Bicyclic graphs with extremal cover cost. Applied Mathematics and Computation, 405(??): Article 126235, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003258.

Liang:2022:FCM

[LPL^{+}22] Junying Liang, Haipeng Peng, Lixiang Li, Fenghua Tong, and Yixian Yang. Flexible construction of measurement matrices in compressed sensing based on extensions of incidence matrices of combinatorial designs. Applied Mathematics and Computation, 420 (??):Article 126901, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100984X.

Lee:2021:EIS

[LPP21a] Tae H. Lee, Myeong Jin Park, and Ju H. Park. Erratum to “An improved stability criterion of neural networks with timevarying delays in the form of quadratic function using novel geometrybased conditions” [Appl. Math. Compt. 404 (2021) 126226]. Applied Mathematics and Computation, 410(??):Article 126390, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004793.

See [LPP21b].

Lee:2021:ISC

[LPP21b] Tae H. Lee, Myeong Jin Park, and Ju H. Park. An improved stability criterion of neural networks with time-varying delays in the form of quadratic function using novel geometrybased conditions. Applied Mathematics and Computation, 404(??):Article 126226, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003167. See

erratum [LPP21a].

Lopez:2024:GLM

[LPP24] José L. López, Pedro J. Pagola, and Pablo Palacios. A generalization of the Laplace’s method for integrals. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400448X.

Lera:2021:SFC

[LPS21] Daniela Lera, Mikhail Posypkin, and Yaroslav D. Sergeyev. Space-filling curves for numerical approximation and visualization of solutions to systems of nonlinear inequalities with applications in robotics. Applied Mathematics and Computation, 390(??):Article 125660, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305762.

Lansky:2023:IOC

[LPS23] Petr Lansky, Federico Polito, and Laura Sacerdote. Input– output consistency in integrate and fire interconnected neurons. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007032.

Lin:2024:FAM

[LQ24] Cuiyun Lin and Tao Qian.

Frequency analysis with multiple kernels and complete dictionary. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300646X.

Li:2021:GBN

[LQL21] Yinkui Li, Xiaoxiao Qin, and Wen Li. The generalized burning number of graphs. Applied Mathematics and Computation, 411 (??):Article 126306, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003957.

Lei:2022:CCR

[LQLF22] Weidong Lei, Xiaofei Qin, Hongjun Li, and Youhua Fan. Causality condition relevant functionsorientated analytical treatment on singularities in 3D TD-BEM. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001977.

Luo:2023:SOA

[LQNA23] Man Luo, Wenlin Qiu, Omid Nikan, and Zakieh Avazzadeh. Secondorder accurate, robust and efficient ADI Galerkin technique for the threedimensional nonlocal heat model arising in viscoelasticity. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007159.

Li:2023:RSD

[LQS^{+}23] Dandan Li, Wenqi Qian, Xiaoxiao Sun, Dun Han, and Mei Sun. Rumor spreading in a dual-relationship network with diverse propagation abilities. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004022.

Liu:2021:SSB

[LQSZ21] Yan Liu, Xulong Qin, Jincheng Shi, and Wenjing Zhi. Structural stability of the Boussinesq fluid interfacing with a Darcy fluid in a bounded region in R^{2}. Applied Mathematics and Computation, 411 (??):Article 126488, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005774.

Li:2020:EPS

[LQW20] Wei Li, Zhongmei Qin, and Yao Wang. Enumeration of permanental sums of lattice graphs. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309063.

Lin:2020:CBS

[LR20] Ji Lin and Sergiy Reutskiy. A cubic B-spline semi-analytical algorithm for simulation of 3D steadystate convection–diffusion– reaction problems. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309361.

Li:2021:ISD

[LR21] Yongjun Li and Valery G. Romanovski. Isochronous solutions of a 3-dim symmetric quadratic system. Applied Mathematics and Computation, 405(??):Article 126250, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003404.

Li:2022:NCG

[LR22] Ting Li and Han Ren.

The number of 4-cycles in a graph. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002855.

Lu:2023:RDE

[LR23] Xuchen Lu and Junchao Ren. Robust (Q, S, R) − αdissipative event-triggered sliding mode control of discrete T–S fuzzy descriptor systems with partial unmeasurable premise variables. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001790.

Li:2024:AET

[LR24] Ze Li and Junchao Ren.

Adaptive event-triggered non-fragile sliding mode control for uncertain T-S fuzzy singular systems with passive constraint. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001012.

Laeuchli:2022:ACM

[LRCTR22] Jesse Laeuchli, Yunior Ramírez-Cruz, and Rolando Trujillo-Rasua. Analysis of centrality measures under differential privacy models. Applied Mathematics and Computation, 412 (??):Article 126546, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006305.

Llibre:2020:DEG

[LRS20] Jaume Llibre, Rafael Ramírez, and Natalia Sadovskaia. Differential equations with a given set of solutions. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306514.

Lamberti:2020:MQS

[LS20] Paola Lamberti and Antonia Saponaro. Multilevel quadratic spline quasiinterpolation. Applied Mathematics and Computation, 373(??):Article 125047, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300163.

Luo:2022:DMI

[LS22a] Xin Luo and Min Sun. Development of modal interval algorithm for solving continuous minimax problems. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200056X.

Lv:2022:PSI

[LS22b] Shaojie Lv and Feifei Song.

Particle swarm intelligence and the evolution of cooperation in the spatial public goods game with punishment. Applied Mathematics and Computation, 412(??):Article 126586, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006706.

Lavor:2021:ISP

[LSC^{+}21] Carlile Lavor, Michael Souza, Luiz M. Carvalho, Douglas S. Gonçalves, and Antonio Mucherino. Improving the sampling process in the interval Branchand-Prune algorithm for the discretizable molecular distance geometry problem. Applied Mathematics and Computation, 389 (??):Article 125586, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305312.

Liu:2023:AMB

[LSC^{+}23] Xingyue Liu, Kaibo Shi, Jun Cheng, Shiping Wen, and Yajuan Liu. Adaptive memory-based eventtriggering resilient LFC for power system under DoS attack. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002102.

Lin:2021:NSP

[LSCS21] Y.-T. Lin, Y.-T. Shih, C.-S. Chien, and Q. Sheng. A note on stochastic polynomial chaos expansions for uncertain volatility and Asian option pricing. Applied Mathematics and Computation, 393(??):Article 125764, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307177.

Li:2022:PDG

[LSHH22] Dandan Li, Xiaoxiao Sun, Youxin He, and Dun Han. On prisoner’s dilemma game with psychological bias and memory learning. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004647.

Li:2020:EDL

[LSL^{+}20] Junfeng Li, Xinhui Si, Botong Li, Limei Cao, and Peipei Zhang. The effects of depletion layer for electroosmotic flow of fractional second-grade viscoelastic fluid in a micro-rectangle channel. Applied Mathematics and Computation, 385 (??):Article 125409, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303714.

Liu:2021:CPF

[LSL^{+}21] Yansheng Liu, Mengjin Song, Haitao Li, Yalu Li, and Wenying Hou. Containment problem of finite-field networks with fixed and switching topology. Applied Mathematics and Computation, 411 (??):Article 126519, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006081.

Liang:2023:ASM

[LSM23] Tiantian Liang, Shengli Shi, and Yuechao Ma. Asynchronous sliding mode control of continuous-time singular Markov jump systems with time-varying delay under event-triggered strategy. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001169.

Lee:2020:IRS

[LSPK20] S. H. Lee, P. Selvaraj, M. J. Park, and O. M. Kwon. Improved results on H_{∞} stability analysis of sampleddata systems via loopedfunctionals and zero equalities. Applied Mathematics and Computation, 373 (??):Article 125003, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309956.

Le:2021:TDS

[LSV21] Van Chien Le, Marián Slodicka, and Karel Van Bockstal. A time discrete scheme for an electromagnetic contact problem with moving conductor. Applied Mathematics and Computation, 404(??): Article 125997, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100045X.

Lobo-Segura:2020:LZC

[LSVA20] J. Lobo-Segura and M. A.

Villalobos-Arias. Location of the zeros of certain parametric families of functions of generalized Fresnel integral type. Applied Mathematics and Computation, 381(??): Article 125253, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302228.

Li:2020:FTC

[LSWL20] Ping Li, Zhibao Song, Zhen Wang, and Wenhui Liu. Fixed-time consensus for disturbed multiple Euler–Lagrange systems with connectivity preservation and quantized input. Applied Mathematics and Computation, 380(??): Article 125303, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302691.

Liu:2021:CMA

[LSWS21] Bo Liu, Housheng Su, Licheng Wu, and Xixi Shen. Controllability for multiagent systems with matrixweight-based signed network. Applied Mathematics and Computation, 411 (??):Article 126520, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006093.

Liu:2021:DRF

[LSYX21] Yuanyuan Liu, Zhongkui Sun, Xiaoli Yang, and Wei Xu. Dynamical robustness and firing modes in multilayer memristive neural networks of nonidentical neurons. Applied Mathematics and Computation, 409 (??):Article 126384, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004732.

Landi:2023:ABR

[LSZ^{+}23] G. Landi, G. V. Spinelli, F. Zama, D. Chillura Martino, P. Conte, P. Lo Meo, and V. Bortolotti. An automatic L_{1} -based regularization method for the analysis of FFC dispersion profiles with quadrupolar peaks. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008773.

Liang:2022:PSS

[LT22] Mengqian Liang and Yazhou Tian. Practical stability of switched homogeneous positive nonlinear systems: Max-separable Lyapunov function method. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002648.

Loubere:2023:MLC

[LTB23] Raphaël Loubère, Rodolphe Turpault, and Alexandre Bourriaud. A MOODlike compact high order finite volume scheme with adaptive mesh refinement. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008608.

Li:2023:MSS

[LTC^{+}23] Weidong Li, How Wei Benjamin Teo, Kaijuan Chen, Jun Zeng, Kun Zhou, and Hejun Du. Mesoscale simulations of spherulite growth during isothermal crystallization of polymer melts via an enhanced 3D phase-field model. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000425.

Li:2020:ICPb

[LTL^{+}20] Zhiyong Li, Tingting Tang, Yu Liu, Elias J. G. Arcondoulis, and Yannian Yang. Implementation of compressible porous-fluid coupling method in an aerodynamics and aeroacoustics code — Part II: Turbulent flow. Applied Mathematics and Computation, 373 (??):Article 124988, May 15, 2020. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309804.

Liu:2021:EDS

[LTL^{+}21] Yu-An Liu, Shengdao Tang, Yufan Liu, Qingkai Kong, and Jing Wang. Extended dissipative sliding mode control for nonlinear networked control systems via event-triggered mechanism with random uncertain measurement. Applied Mathematics and Computation, 396(??):Article 125901, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308547.

Li:2022:TVF

[LTL^{+}22] Lili Li, Yulong Tuo, Tieshan Li, Meijuan Tong, and Shasha Wang. Timevarying formation control of multiple unmanned surface vessels with heterogeneous hydrodynamics subject to actuator attacks. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200073X.

Llorente:2020:SDT

[LtTBPA20] Víctor J. Llorente, Jan H. M. ten Thije Boonkkamp, Antonio Pascau, and Martijn J. H. Anthonissen. Similarities and differences of two exponential schemes for convection-diffusion problems: the FV–CF and ENATE schemes. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306927.

Li:2022:APM

[LTW22] Jiajia Li, Xin Tian, and Guoliang Wei. Asynchronous partially modedependent control for switched larger-scale nonlinear systems with bounded sojourn time. Applied Mathematics and Computation, 418 (??):Article 126809, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008912.

Li:2024:EMO

[LTW24] Jiajia Li, Xin Tian, and Guoliang Wei. Energy management for object tracking under the energyharvesting: Hierarchical reinforcement learning method.

Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003199.

Liu:2024:ADC

[LTWY24] Weijun Liu, Jianxiong Tang, Jiaqiu Wang, and Jing Yang. Algebraic degree of Cayley graphs over dicyclic and semi-dihedral groups. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005581.

Lu:2022:DAN

[Lu22] Chun Lu. Dynamical analysis and numerical simulations on a Crowley–Martin predator–prey model in stochastic environment. Applied Mathematics and Computation, 413(??):Article 126641, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007256.

Lopez-Urena:2024:UNL

[LU24]  Sergio López-Ureña.  A

uniform non-linear subdivision scheme reproducing polynomials at any nonuniform grid. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003503.

Luo:2020:GMT

[Luo20]  Yidong Luo.  Galerkin

method with trigonometric basis on stable numerical differentiation. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930904X.

Lobo:2020:GAO

[LV20] Jervin Zen Lobo and Y. S. Valaulikar. Group analysis of the one dimensional wave equation with delay. Applied Mathematics and Computation, 378 (??):Article 125193, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301624.

Li:2020:NKF

[LW20a] X. Y. Li and B. Y. Wu. A new kernel functions based approach for solving 1-D interface problems. Applied Mathematics and Computation, 380(??): Article 125276, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302459.

Lin:2020:IER

[LW20b] Hairong Lin and Chunhua Wang. Influences of electromagnetic radiation distribution on chaotic dynamics of a neural network. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930832X.

Lv:2020:IHI

[LW20c] Shaojie Lv and Xianjia Wang. The impact of heterogeneous investments on the evolution of cooperation in public goods game with exclusion. Applied Mathematics and Computation, 372(??):Article 124960, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031930952X.

Li:2021:INC

[LW21a] Zhonghua Li and Baoyindureng Wu. The kindependence number of tconnected graphs. Applied Mathematics and Computation, 409(??):Article 126412, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321005014.

Luo:2021:GDB

[LW21b] Demou Luo and Qiru Wang. Global dynamics of a Beddington-DeAngelis amensalism system with weak Allee effect on the first species. Applied Mathematics and Computation, 408 (??):Article 126368, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004574.

Lei:2022:SDG

[LW22a] Xingyu Lei and Jianfeng Wang. Spectral determination of graphs with one positive anti-adjacency eigenvalue. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000819.

Li:2022:ETR

[LW22b] Lingxi Li and Tao Wang. An extension of Thomassen’s result on choosability. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001849.

Li:2022:EPM

[LW22c] Songxin Li and Yongke Wu.

Energy-preserving mixed finite element methods for the elastic wave equation. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000492.

Li:2022:UCA

[LW22d] Yaxiang Li and Jiangxing Wang. Unconditional convergence analysis of stabilized FEM–SAV method for Cahn–Hilliard equation. Applied Mathematics and Computation, 419(??): Article 126880, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321009632.

Li:2022:PWC

[LW22e] Zhenzhen Li and Baoyindureng Wu. Properwalk connection of Hamiltonian digraphs. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002442.

Liu:2022:EQP

[LW22f] Fanglin Liu and Bin Wu. Environmental quality and population welfare in Markovian eco-evolutionary dynamics. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003836.

Liu:2022:RAP

[LW22g] Jun Liu and Zhu Wang. A ROM-accelerated parallelin-time preconditioner for solving all-at-once systems in unsteady convectiondiffusion PDEs. Applied Mathematics and Computation, 416(??):Article 126750, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008328.

Li:2023:LTA

[LW23a] Xicui Li and Bin Wang.

Long term analysis of splitting methods for chargedparticle dynamics. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007500.

Li:2023:NCG

[LW23b] Yinkui Li and Liqun Wei.

Note for the conjecture on the generalized 4connectivity of total graphs of the complete bipartite graph. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003946.

Lord:2024:CET

[LW24] Gabriel Lord and Mengchao Wang. Convergence of a exponential tamed method for a general interest rate model. Applied Mathematics and Computation, 467 (??):??, April 15, 2024. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006720.

Li:2024:IUE

[LWCD24] Jiyong Li, Xianfen Wang, Qianyu Chen, and Shuo Deng. Improved uniform error bounds of a Lawson-type exponential wave integrator method for the Klein– Gordon–Dirac equation. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003382.

Li:2021:DRI

[LWD21] Xin Li, Guoliang Wei, and Derui Ding. Distributed resilient interval estimation for sensor networks under aperiodic denial-of-service attacks and adaptive event-triggered protocols. Applied Mathematics and Computation, 409 (??):Article 126371, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004604.

Liu:2020:TCD

[LWG20] Lizhi Liu, Yinhe Wang, and Zilin Gao. Tracking control for the dynamic links of discretetime complex dynamical network via state observer. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308495.

Liu:2022:ATC

[LWH22] Yanli Liu, Runzhi Wang, and Li-Ying Hao. Adaptive TD control of fullstate-constrained nonlinear stochastic switched systems. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002193.

Lu:2021:SDP

[LWJ21] Minmin Lu, Yan Wang, and Daqing Jiang. Stationary distribution and probability density function analysis of a stochastic HIV model with cell-to-cell infection. Applied Mathematics and Computation, 410(??):Article 126483, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005725.

Liu:2020:ILH

[LWL^{+}20] Ling Liu, Xiao-Yong Wen, Nan Liu, Tao Jiang, and Jin-Yun Yuan. An integrable lattice hierarchy associated with a 4 × 4 matrix spectral problem: N-fold Darboux transformation and dynamical properties. Applied Mathematics and Computation, 387(??):Article 124525, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319305041.

Liu:2021:IBC

[LWL21] Yisi Liu, Xiaojun Wang, and Yunlong Li. An improved Bayesian collocation method for steadystate response analysis of structural dynamic systems with large interval uncertainties. Applied Mathematics and Computation, 411(??):Article 126523, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006123.

Li:2022:UTK

[LWL22] Jiajia Li, Guoliang Wei, and Wangyan Li. Unscented Tobit Kalman filtering for switched nonlinear systems with censored measurement. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004015.

Li:2023:ESN

[LWL^{+}23a] Mingyue Li, Mingzhu Wang, Wenlu Liu, Shuchen Wu, and Xiaodi Li. Exponential stability of nonlinear systems via event-triggered impulsive control based on partial states. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004435.

Li:2023:MLS

[LWL23b] Xing-Yu Li, Kai-Ning Wu, and Xiao-Zhen Liu. Mittag-Leffler stabilization for short memory fractional reactiondiffusion systems via intermittent boundary control. Applied Mathematics and Computation, 449(??): ??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001285.

Liu:2023:DRF

[LWL^{+}23c] Dan Liu, Zidong Wang, Yurong Liu, Changfeng Xue, and Fuad E. Alsaadi. Distributed recursive filtering for time-varying systems with dynamic bias over sensor networks: Tackling packet disorders. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007378.

Liu:2024:EAF

[LWL24] Yong-Jin Liu, Yuqi Wan, and Lanyu Lin. An efficient algorithm for Fantopeconstrained sparse principal subspace estimation problem. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001802.

Liu:2021:RFS

[LWLA21] Dan Liu, Zidong Wang, Yurong Liu, and Fuad E. Alsaadi. Recursive filtering for stochastic parameter systems with measurement quantizations and packet disorders. Applied Mathematics and Computation, 398(??):Article 125960, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000084.

Li:2021:CAE

[LWM^{+}21] Lei Li, Wenting Wang, Qiang Ma, Kunpeng Pan, Xin Liu, Lin Lin, and Jian Li. Cyber attack estimation and detection for cyber-physical power systems. Applied Mathematics and Computation, 400(??): Article 126056, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001041.

Li:2022:FLG

[LWNZ22] Meng Li, Yifan Wei, Binqian Niu, and Yong-Liang Zhao. Fast L2-1_{σ} Galerkin FEMs for generalized nonlinear coupled Schrödinger equations with Caputo derivatives. Applied Mathematics and Computation, 416(??):Article 126734, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100816X.

Liu:2022:EAF

[LWQW22] Zheng Liu, Gaofeng Wei, Shaopeng Qin, and Zhiming Wang. The elastoplastic analysis of functionally graded materials using a meshfree RRKPM. Applied Mathematics and Computation, 413(??):Article 126651, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007359.

Liu:2023:HSF

[LWW23] Weijun Liu, Jiaqiu Wang, and Yongjiang Wu. HSsplitting fields of abelian mixed Cayley graphs. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003119.

Liu:2022:PFS

[LWX^{+}22] Chen Liu, Fang-Guang Wang, Qiang Xue, Li Li, and Zhen Wang. Pattern formation of a spatial vegetation system with root hydrotropism. Applied Mathematics and Computation, 420(??):Article 126913, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009966.

Li:2024:CSP

[LWX24] Yangrong Li, Fengling Wang, and Huan Xia. Continuity-sets of pullback random attractors for discrete porous media equations with colored noise. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006033.

Luo:2024:OBF

[LWY^{+}24] Peng Luo, Defeng Wu, Andre S. Yamashita, Na Feng, and Yang Yang. Observerbased fixed-time dynamic surface tracking control for autonomous surface vehicles under actuator constraints and denialof-service attacks. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005726.

Li:2020:PPC

[LWZ20] Hongyi Li, Chaojie Wang, and Di Zhao. Preconditioning for PDE-constrained optimization with total variation regularization. Applied Mathematics and Computation, 386(??):Article 125470, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030429X.

Li:2021:FTA

[LWZ21] Zheng Li, Fang Wang, and Ruitai Zhu. Finitetime adaptive neural control of nonlinear systems with unknown output hysteresis. Applied Mathematics and Computation, 403(??):Article 126175, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002654.

Li:2022:POI

[LWZ22a] Nan Li, Song Wang, and Kai Zhang. Price options on investment project expansion under commodity price and volatility uncertainties using a novel finite difference method. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000236.

Li:2022:ESI

[LWZ22b] Shuchao Li, Zheng Wang, and Minjie Zhang. On the extremal Sombor index of trees with a given diameter. Applied Mathematics and Computation, 416(??): Article 126731, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008134.

Liu:2022:RET

[LWZ22c] Di Liu, Shaoping Wang, and Chao Zhang. Reliability estimation from two types of accelerated testing data based on an artificial neural network supported Wiener process. Applied Mathematics and Computation, 417(??):Article 126757, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008390.

Luo:2022:TIL

[LWZ22d] Shangzhen Luo, Mingming Wang, and Wei Zhu. Time-inconsistent life-cycle consumption and retirement choice with mortality risk. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004702.

Liu:2020:DSC

[LWZW20] Yan-Ping Liu, Lin Wang, Feng Zhang, and Rui-Wu Wang. Diffusion sustains cooperation via forming diverse spatial patterns in prisoner’s dilemma game. Applied Mathematics and Computation, 375(??): Article 125077, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300461.

Li:2020:PES

[LX20a] Shuo Li and Zhengrong Xiang. Positivity, exponential stability and disturbance attenuation performance for singular switched positive systems with timevarying distributed delays. Applied Mathematics and Computation, 372(??): Article 124981, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300319309737.

Lin:2020:NRF

[LX20b] ShaoZhong Lin and ZhiQiang Xie. A new recursive formula for integration of polynomial over simplex. Applied Mathematics and Computation, 376(??): Article 125140, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301090.

Li:2022:BRA

[LX22] Fuzhi Li and Dongmei Xu. Backward regularity of attractors for lattice FitzHugh–Nagumo system with double random coefficients. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003794.

Li:2023:AHS

[LX23a] Qian Li and Yanni Xiao.

Analysis of a hybrid SIR model combining the fixedmoments pulse interventions with susceptiblestriggered threshold policy. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300323002515.

Liu:2023:EQR

[LX23b] Guidong Liu and Shuhuang Xiang. An efficient quadrature rule for weakly and strongly singular integrals. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300070X.

Lu:2024:EPQ

[LX24a] Ji-Jing Lu and Jun Xiong. Energy-to-peak quantized filtering for T –S fuzzy systems with event-triggered-based weighted try-once-discard protocol: the finite-time case. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004363.

Luo:2024:NCS

[LX24b] Zuwen Luo and Kexiang Xu. The number of connected sets in Apollonian networks.

Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003448.

Li:2021:CDT

[LXCZ21] Jianquan Li, Xin Xie, Yuming Chen, and Dian Zhang. Complex dynamics of a tumor-immune system with antigenicity. Applied Mathematics and Computation, 400(??):Article 126052, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001004.

Li:2020:SES

[LXF20] Cong Li, Hedong Xu, and Suohai Fan. Synergistic effects of self-optimization and imitation rules on the evolution of cooperation in the investor sharing game. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309142.

Li:2021:ECG

[LXF21] Cong Li, Hedong Xu, and Suohai Fan. Evolutionary compromise game on assortative mixing networks. Applied Mathematics and Computation, 390 (??):Article 125681, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306342.

Li:2024:MIE

[LXH24] Dandan Li, Weijie Xie, and Dun Han. Multiinformation and epidemic coupling propagation considering indirect contact on two-layer networks. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001723.

Liu:2023:SBC

[LXK23] Mingyu Liu, Jing Xie, and Yonggui Kao. Stochastic bounded consensus for multi-agent systems with fractional Brownian motions via sliding mode control. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000486.

Li:2022:USO

[LXL22] Xindong Li, Wenwen Xu, and Wei Liu. Unconditional stability and optimal error analysis of mass conservative characteristic mixed FEM for wormhole propagation. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200248X.

Li:2020:RDB

[LXLP20] Zhemin Li, Zheng Xie, Jianping Li, and Yingui Pan. Resistance distance-based graph invariants and spanning trees of graphs derived from the strong prism of a star. Applied Mathematics and Computation, 382(??):Article 125335, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303015.

Li:2022:CSD

[LXP^{+}22] WenYao Li, Xiaoyu Xue, Liming Pan, Tao Lin, and Wei Wang. Competing spreading dynamics in simplicial complex. Applied Mathematics and Computation, 412 (??):Article 126595, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006792.

Liu:2020:SDE

[LXW^{+}20] Yamin Liu, Zuxing Xuan, Zhen Wang, Jianping Zhou, and Yajuan Liu. Sampleddata exponential synchronization of time-delay neural networks subject to random controller gain perturbations. Applied Mathematics and Computation, 385 (??):Article 125429, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303908.

Li:2022:PPT

[LXWH22] Ming-Yang Li, Wen-Bo Xie, Yu-Long Wang, and Xin Hu. Prescribed performance trajectory tracking faulttolerant control for dynamic positioning vessels under velocity constraints. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004222.

Liu:2023:RIC

[LXZ23] Rong-Rong Liu, Jun Xiong, and Zaolin Zhang. Reliable induced L_{∞} control for polytopic uncertain continuous-time singular systems with dynamic quantization. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008335.

Lin:2020:VPI

[LY20a] Zhicong Lin and Sherry H. F. Yan. Vincular patterns in inversion sequences. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306642.

Luo:2020:DET

[LY20b] Shengping Luo and Dan Ye.

Double event-triggered control for linear multi-agent systems with augmented dynamic triggering mechanisms. Applied Mathematics and Computation, 386(??):Article 125522, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030480X.

Li:2021:SNR

[LY21a] Jiayun Li and Jianhua Yin. Some new results on the Turán number of star forests. Applied Mathematics and Computation, 407 (??):Article 126320, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004094.

Liu:2021:PMB

[LY21b] Z. Liu and Y. Yang. Pharmacokinetic model based on multifactor uncertain differential equation. Applied Mathematics and Computation, 392(??):Article 125722, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306755.

Liu:2021:UPM

[LY21c] Z. Liu and Y. Yang. Uncertain pharmacokinetic model based on uncertain differential equation. Applied Mathematics and Computation, 404(??):Article 126118, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001661.

Li:2022:TNK

[LY22a] Guangming Li and Jianhua Yin. On the Turán numbers of kKr in lpartite graphs. Applied Mathematics and Computation, 417(??):Article 126791, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008730.

Liu:2022:DAP

[LY22b] Huafeng Liu and Jun Yue. Diophantine approximation over primes with different powers. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000261.

Liu:2022:MEP

[LY22c] Zhe Liu and Ying Yang.

Moment estimation for parameters in high-order uncertain differential equations. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004738.

Lv:2022:ACK

[LY22d] Yuan-Wei Lv and Guang-Hong Yang. An adaptive cubature Kalman filter for nonlinear systems against randomly occurring injection attacks. Applied Mathematics and Computation, 418(??):Article 126834, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009176.

Li:2023:CST

[LY23a] Danyi Li and Weigen Yan. Counting spanning trees with a Kekulé structure in linear hexagonal chains. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002941.

Liu:2023:CEW

[LY23b] Linghai Liu and Weigen Yan. Combinatorial explanation of the weighted Laplacian characteristic polynomial of a graph and applications. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003569.

Liu:2023:NAS

[LY23c] Xinfei Liu and Xiaoyuan Yang. Numerical approximation of the stochastic equation driven by the fractional noise. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002229.

Li:2024:EGG

[LY24] Fengwei Li and Qingfang Ye. Extremal graphs with given parameters in respect of general ABS index. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004351.

Li:2021:EVT

[LYB^{+}21] Fengwei Li, Qingfang Ye, Hajo Broersma, Ruixuan Ye, and Xiaoyan Zhang. Extremality of VDB topological indices over fbenzenoids with given order. Applied Mathematics and Computation, 393(??): Article 125757, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307104.

Li:2024:AEM

[LYhH^{+}24] Long Li, Yu Yang, Zhi hao Hui, Bang-Bang Jin, Hua Wang, Asfand Fahad, and Heng Zhang. Algorithms for enumerating multiple leaf-distance granular regular α-subtree of unicyclic and edge-disjoint bicyclic graphs. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005039.

Li:2021:PVS

[LYJM21] Jing Jian Li, Jing Yang, Ran Ju, and Hongping Ma. Prime-valent symmetric graphs admitting alternating transitive group. Applied Mathematics and Computation, 407(??):Article 126334, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321004239.

Lee:2021:MSD

[LYKK21] Hyun Geun Lee, Junxiang Yang, Sangkwon Kim, and Junseok Kim. Modeling and simulation of droplet evaporation using a modified Cahn–Hilliard equation. Applied Mathematics and Computation, 390 (??):Article 125591, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305464.

Liu:2022:DET

[LYL22] Mengmeng Liu, Jinyong Yu, and Yu Liu. Dynamic event-triggered asynchronous fault detection for Markov jump systems with partially accessible hidden information and subject to aperiodic DoS attacks. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003915.

Li:2021:MFC

[LYN^{+}21] Xiang Li, Peng Yu, Xiao-Dong Niu, De-Cai Li, and Hiroshi Yamaguchi. A magnetic field coupling lattice Boltzmann model and its application on the merging process of multiple-ferrofluid-droplet system. Applied Mathematics and Computation, 393(??):Article 125769, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307220.

Lu:2022:CAJ

[LYN22] Jiashu Lu, Mengna Yang, and Yufeng Nie. Convergence analysis of Jacobi spectral collocation methods for weakly singular nonlocal diffusion equations with volume constraints. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004192.

Lyu:2023:LAP

[Lyu23] Yijia Lyu. Labeling algorithm for power domination problem of trees. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003326.

Li:2023:PER

[LYX23] Ting Li, Dongsheng Yang, and Xiangpeng Xie. Prioritized experience replay based reinforcement learning for adaptive tracking control of autonomous underwater vehicle. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008025.

Liu:2021:ESC

[LYXZ21] Bin Liu, Meng Yang, Bo Xu, and Guohua Zhang. Exponential stabilization of continuous-time dynamical systems via time and event triggered aperiodic intermittent control. Applied Mathematics and Computation, 398(??):Article 125713, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306664.

Li:2022:SRM

[LYZ22a] Li Li, Xihong Yan, and Xinzhen Zhang. An SDP relaxation method for Perron pairs of a nonnegative tensor. Applied Mathematics and Computation, 423(??): ??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321009498.

Lu:2022:OEA

[LYZ22b] Xiaoping Lu, Dong Yan, and Song-Ping Zhu. Optimal exercise of American puts with transaction costs under utility maximization. Applied Mathematics and Computation, 415 (??):Article 126684, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007682.

Lin:2020:ESP

[LZ20a] Fubiao Lin and Qianhong Zhang. Exact solutions of population balance equations for breakage and growth processes, using group analysis. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307829.

Lin:2020:SRG

[LZ20b] Hongying Lin and Bo Zhou. The α-spectral radius of general hypergraphs. Applied Mathematics and Computation, 386(??):Article 125449, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304100.

Lin:2020:GND

[LZ20c] Shangwei Lin and Wenli Zhang. The 1-good-neighbor diagnosability of unidirectional hypercubes under the PMC model. Applied Mathematics and Computation, 375(??):Article 125091, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300606.

Lin:2021:WNS

[LZ21] Hongying Lin and Bo Zhou.

Which numbers are status differences? Applied Mathematics and Computation, 399(??):Article 126004, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000527.

Li:2022:MMA

[LZ22a] Huixia Li and Hongyong Zhao. Mathematical model of Alzheimer’s disease with prion proteins interactions and treatment. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004519.

Li:2022:BTB

[LZ22b] Jinghan Li and Jun Zhao. Bumpless transfer based event-triggered control for switched linear systems with state-dependent switching. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003708.

Li:2022:SAH

[LZ22c] Rong Li and Bing Zheng. A spatially adaptive hybrid total variation model for image restoration under Gaussian plus impulse noise. Applied Mathematics and Computation, 419(??): Article 126862, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009450.

Li:2022:VMS

[LZ22d] Zhe Li and Chunlei Zhang. The verification of multiplicity support of a defective eigenvalue of a real matrix. Applied Mathematics and Computation, 414 (??):Article 126671, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007554.

Liu:2022:SLS

[LZ22e] Xiaowei Liu and Jin Zhang.

Supercloseness of linear streamline diffusion finite element method on bakhvalov-type mesh for singularly perturbed convectiondiffusion equation in 1D. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003320.

Lv:2022:AFE

[LZ22f] Yanhui Lv and Jin Zhang.

Analysis of finite element method in balanced norms for two-parameter singularly perturbed problems. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003897.

Lv:2022:CSF

[LZ22g] Yanhui Lv and Jin Zhang. Convergence and supercloseness of a finite element method for a two-parameter singularly perturbed problem on Shishkin triangular mesh. Applied Mathematics and Computation, 416(??): Article 126753, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008353.

Li:2023:REC

[LZ23] Ruo Li and Wei Zhong. A robust and efficient component-wise WENO scheme for Euler equations. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006567.

Lai:2024:EST

[LZ24a] Jingchao Lai and Rongkun Zhu. Enumeration of spanning trees containing a perfect matching in linear polygonal chains. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003436.

Li:2024:VDS

[LZ24b] Jiaqi Li and Yi Zhang. Vertex degree sums for supereulerian bipartite digraphs. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400225X.

Li:2024:ATC

[LZ24c] Jiyong Li and Lu Zhao.

Analysis of two conservative fourth-order compact finite difference schemes for the Klein–Gordon– Zakharov system in the subsonic limit regime. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004575.

Liu:2024:ESR

[LZ24d] Yu-An Liu and Li Zhang.

Enhancing secondary recovery in microgrid: Eventdriven robust security control with semi-Markov switched dynamics. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002996.

Lu:2024:SSN

[LZ24e] Boliang Lu and Quanxin Zhu. Stability of switched neutral stochastic functional systems with different structures under high nonlinearity. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002182.

Liu:2021:DNW

[LZC21] Juan Liu, Xindong Zhang, and Xing Chen. The domination number of wrapped butterfly digraphs. Applied Mathematics and Computation, 409(??):Article 126407, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004963.

Li:2022:LPS

[LZC22a] Jing Jian Li, Xiao Qian Zhang, and Jun Yan Chen. On locally primitive symmetric graph of valency twice prime. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000601.

Li:2022:ETC

[LZC22b] Ping Li, Wu Zhao, and Jun Cheng. Event-triggered control for exponential stabilization of impulsive dynamical systems. Applied Mathematics and Computation, 413(??):Article 126608, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006925.

Li:2023:ADAa

[LZCL23] Jiaqi Li, Jianlei Zhang, Zengqiang Chen, and Qun Liu. Aspiration drives adaptive switching between two different payoff matrices. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000504.

Lu:2023:PEA

[LZD23] Shounan Lu, Ge Zhu, and Jianhua Dai. Promoting effect of adaptive interaction based on random neighbors to cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001297.

Li:2022:IEB

[LZDX22] Tao Li, Zhixuan Zhang, Zhuxiang Dai, and Liuyong Xu. Intermediate-estimator-based fault tolerant shape control via PDFs. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005501.

Li:2020:GML

[LZH^{+}20] Hong-Li Li, Long Zhang, Cheng Hu, Haijun Jiang, and Jinde Cao. Global Mittag-Leffler synchronization of fractional-order delayed quaternion-valued neural networks: Direct quaternion approach. Applied Mathematics and Computation, 373(??):Article 125020, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300319310124.

Long:2021:FTS

[LZH21] Changqing Long, Guodong Zhang, and Junhao Hu. Fixed-time synchronization for delayed inertial complex-valued neural networks. Applied Mathematics and Computation, 405(??): Article 126272, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003611.

Lin:2020:FTS

[LZHZ20] Xiangze Lin, Wanli Zhang, Shuaiting Huang, and Enlai Zheng. Finite-time stabilization of input-delay switched systems. Applied Mathematics and Computation, 375(??):Article 125062, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030031X.

Liao:2024:DOF

[LZHZ24] Fang Liao, Yanzheng Zhu, Xiao He, and Donghua Zhou. Dynamic output feedback based fault tolerant control for continuoustime switched affine systems via reduced-order observer. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003369.

Liu:2020:DPP

[LZL20] Fang Liu, Shu-Cai Zou, and Qing Li. Deriving priorities from pairwise comparison matrices with a novel consistency index. Applied Mathematics and Computation, 374 (??):Article 125059, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030028X.

Liu:2021:QFO

[LZL^{+}21] Jianjun Liu, Rui Zhai, Yuhan Liu, Wenliang Li, Bingzhe Wang, and Liyuan Huang. A quasi fractional order gradient descent method with adaptive stepsize and its application in system identification. Applied Mathematics and Computation, 393(??): Article 125797, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307505.

Li:2020:ICPa

[LZLM20] Zhiyong Li, Huaibao Zhang, Yu Liu, and James M. Mc-Donough. Implementation of compressible porous-fluid coupling method in an aerodynamics and aeroacoustics code part I: Laminar flow. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306745.

Liu:2021:ESN

[LZLZ21] Chong Liu, Huaguang Zhang, Yanhong Luo, and Kun Zhang. Echo state network-based online optimal control for discrete-time nonlinear systems. Applied Mathematics and Computation, 409(??):Article 126324, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004136.

Lin:2021:NMS

[LZRF21] Ji Lin, Yuhui Zhang, Sergiy Reutskiy, and Wenjie Feng. A novel meshless spacetime backward substitution method and its application to nonhomogeneous advection-diffusion problems. Applied Mathematics and Computation, 398 (??):Article 125964, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000126.

Li:2021:SSC

[LZRG21] Xiaowang Li, Shuming Zhou, Xiangyu Ren, and Xia Guo. Structure and substructure connectivity of alternating group graphs. Applied Mathematics and Computation, 391 (??):Article 125639, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305932.

Lin:2021:RHD

[LZS^{+}21] Yuqian Lin, Guangming Zhuang, Wei Sun, Junsheng Zhao, and Yuming Chu. Resilient H_{∞} dynamic output feedback controller design for USJSs with timevarying delays. Applied Mathematics and Computation, 395(??):Article 125875, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308286.

Li:2022:HON

[LZSC22] Lili Li, Dan Zhao, Mianfu She, and Xiaoli Chen. On high order numerical schemes for fractional differential equations by block-by-block approach. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001825.

Li:2020:FIN

[LZW^{+}20a] Shudong Li, Dawei Zhao, Xiaobo Wu, Zhihong Tian, Aiping Li, and Zhen Wang. Functional immunization of networks based on message passing. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307209.

Liu:2020:ETC

[LZW^{+}20b] Yanhong Liu, Huimin Zhi, Jumei Wei, Xunlin Zhu, and Quanxin Zhu. Eventtriggered control for linear continuous switched singular systems. Applied Mathematics and Computation, 374(??):Article 125038, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300072.

Li:2022:FDE

[LZWC22] Weihua Li, Huaguang Zhang, Wei Wang, and Zhengbao Cao. Fully distributed event-triggered time-varying formation control of multi-agent systems subject to modeswitching denial-of-service attacks. Applied Mathematics and Computation, 414 (??):Article 126645, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007293.

Li:2022:NRE

[LZX22a] Xiang-Jun Li, Xue-Qian Zeng, and Jun-Ming Xu. Note on reliability evaluation of arrangement graphs. Applied Mathematics and Computation, 418(??):Article 126845, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009280.

Liu:2022:CBE

[LZX^{+}22b] Xinrui Liu, Mingchao Zhang, Xiangpeng Xie, Liang Zhao, and Qiuye Sun. Consensusbased energy management of multi-microgrid: an improved SoC-based power coordinated control method. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001709.

Li:2020:NSP

[LZY20] Qiang Li, Tong Zhang, and Jinyun Yuan. Numerical simulation of polymer crystal growth under flow field using a coupled phase-field and lattice Boltzmann method. Applied Mathematics and Computation, 387(??):Article 124302, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319302309.

Lee:2023:PRS

[LZY^{+}23a] Yan-Li Lee, Tao Zhou, Kexin Yang, Yajun Du, and Liming Pan. Personalized recommender systems based on social relationships and historical behaviors. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006233.

Liu:2023:PAI

[LZY^{+}23b] Jun Liu, Quansheng Zang, Fan Yang, Jing Zhang, and Gao Lin. Performance assessment of internal porous structures on liquid sloshing in various 3D tanks by multi-domain IGABEM. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006944.

Lou:2021:SDI

[LZYL21] Yu Lou, Yi Zhang, Rusuo Ye, and Miao Li. Solitons and dynamics for the integrable nonlocal pair-transition-coupled nonlinear Schrödinger equation. Applied Mathematics and Computation, 409 (??):Article 126417, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005063.

Liu:2020:MBS

[LZZ^{+}20] Chuang Liu, Nan Zhou, Xiu-Xiu Zhan, Gui-Quan Sun, and Zi-Ke Zhang. Markov-based solution for information diffusion on adaptive social networks. Applied Mathematics and Computation, 380(??):Article 125286, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302551.

Li:2021:ESE

[LZZ21] Huichun Li, Xue Zhang, and Chengli Zhao. Explaining social events through community evolution on temporal networks. Applied Mathematics and Computation, 404(??):Article 126148, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100196X.

Liu:2022:GSR

[LZZ22] Huiqing Liu, Ruiting Zhang, and Shunzhe Zhang. On the global strong resilience of fault Hamiltonian graphs. Applied Mathematics and Computation, 418(??):Article 126841, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009243.

Liu:2023:ATV

[LZZ^{+}23] Yiqun Liu, Guangming Zhuang, Junsheng Zhao, Junwei Lu, and Zekun Wang. H^{∞} admissibilization for time-varying delayed nonlinear singular impulsive jump systems based on memory state-feedback control. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000905.

Li:2024:EES

[LZZ24a] Zhao-Yan Li, Qianqian Zhang, and Bin Zhou. On exponential and L_{2}- exponential stability of continuous-time delay-difference systems. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004107.

Long:2024:NRS

[LZZ24b] Shaohua Long, Yu Zhang, and Shouming Zhong. New results on the stability and stabilization for singular neutral systems with time delay. Applied Mathematics and Computation, 473(??):

??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001152.

Li:2023:ISS

[LZZG23] Yuanen Li, Huasheng Zhang, Tingting Zhang, and Han Geng. Interval stability/ stabilization and H_{∞} feedback control for linear impulsive stochastic systems. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006269.

Liang:2021:IRL

[LZZL21] Yuling Liang, Huaguang Zhang, Juan Zhang, and Yanhong Luo. Integral reinforcement learning-based guaranteed cost control for unknown nonlinear systems subject to input constraints and uncertainties. Applied Mathematics and Computation, 408 (??):Article 126336, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004252.

Ma:2021:ASD

[Ma21a] Liangliang Ma. Asymptotic stability and decay rates for the 2D magnetic Bénard fluid equations with mixed partial dissipation, magnetic diffusion and thermal diffusivity. Applied Mathematics and Computation, 399(??): Article 125998, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000461.

Ma:2021:IMC

[Ma21b] Liwen Ma. Important matrix computations in finite topological spaces. Applied Mathematics and Computation, 395(??):Article 125808, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030761X.

Ma:2023:GWP

[Ma23]  Caochuan Ma.  Global

well-posedness and optimal decay estimate for the incompressible porous medium equation near a nontrivial equilibrium. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007482.

Ma:2024:EFS

[Ma24] Jun Ma. Energy function for some maps and nonlinear oscillators. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005489.

Mahmoud:2021:BPS

[MAAEF21] Gamal M. Mahmoud, Tarek Aboelenen, Tarek M. Abed-Elhameed, and Ahmed A. Farghaly. On boundedness and projective synchronization of distributed order neural networks. Applied Mathematics and Computation, 404(??):Article 126198, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002885.

Martire:2022:VIE

[Mar22] Antonio Luciano Martire.

Volterra integral equations: an approach based on Lipschitz-continuity. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322005707.

Marchetti:2024:FSC

[Mar24]  F. Marchetti.  A fast

surrogate cross validation algorithm for meshfree RBF collocation approaches. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004041.

Mirchevski:2024:CFA

[MB24] Stefan Mirchevski and Verica Bakeva. Cost function analysis of a singleserver queueing system with Poisson input stream and erlang- k service time. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002017.

Manuel:2022:GTP

[MBK22] Paul Manuel, Bostjan Bresar, and Sandi Klavzar. The geodesic-transversal problem. Applied Mathematics and Computation, 413 (??):Article 126621, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007050.

Manuel:2023:GPG

[MBK23] Paul Manuel, Bostjan Bresar, and Sandi Klavzar. Geodesic packing in graphs. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323004460.

Muhinyuza:2020:TLT

[MBL20] Stanislas Muhinyuza, Taras Bodnar, and Mathias Lindholm. A test on the location of the tangency portfolio on the set of feasible portfolios. Applied Mathematics and Computation, 386(??):Article 125519, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030477X.

Mendonca:2024:EGD

[MBLL24] J. P. Mendonça, Arthur A.

Brum, M. L. Lyra, and Sérgio A. Lira. Evolutionary game dynamics and the phase portrait diversity in a pandemic scenario. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002212.

Mebirouk:2020:AES

[MBMA20] AbdelMouemin Mebirouk, Sabria Bouheroum-Mentri, and Lidia Aceto. Approximation of eigenvalues of Sturm–Liouville problems defined on a semiinfinite domain. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930815X.

Miranda-Colorado:2020:PIC

[MC20] Roger Miranda-Colorado. Parameter identification of conservative Hamiltonian systems using first integrals. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308525.

Miranda-Colorado:2022:OBP

[MC22] Roger Miranda-Colorado.

Observer-based proportional integral derivative control for trajectory tracking of wheeled mobile robots with kinematic disturbances. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004465.

Mitjana:2021:TGS

[MCB^{+}21] F. Mitjana, S. Cafieri, F. Bugarin, S. Segonds, F. Castanie, and P. Duysinx. Topological gradient in structural optimization under stress and buckling constraints. Applied Mathematics and Computation, 409 (??):Article 126032, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000801.

Musehane:2021:EDC

[MCD21] Ndivhuwo M. Musehane, Madeleine L. Combrinck, and Laurent N. Dala. Eulerian derivation of the conservation equation for energy in a non-inertial frame of reference in arbitrary motion. Applied Mathematics and Computation, 399(??):

Article 126002, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000503.

Ma:2022:OBF

[MCD22] Yong-Sheng Ma, Wei-Wei Che, and Chao Deng. Observer-based fuzzy containment control for nonlinear networked mass under dos attacks. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000273.

Meng:2021:AEV

[MCSD21] Xueyu Meng, Zhiqiang Cai, Shubin Si, and Dongli Duan. Analysis of epidemic vaccination strategies on heterogeneous networks: Based on SEIRV model and evolutionary game. Applied Mathematics and Computation, 403 (??):Article 126172, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002629.

Maia:2022:EER

[MCT^{+}22] A. A. G. Maia, D. F. Cavalca, J. T. Tomita, F. P. Costa, and C. Bringhenti. Evaluation of an effective and robust implicit time-integration numerical scheme for Navier– Stokes equations in a CFD solver for compressible flows. Applied Mathematics and Computation, 413 (??):Article 126612, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006962.

Mei:2022:OFF

[MD22] Keqi Mei and Shihong Ding. Output-feedback finite-time stabilization of a class of constrained planar systems. Applied Mathematics and Computation, 412(??):Article 126573, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006573.

Moreland:2023:ETW

[MD23] Heather L. Moreland and Jack Dockery. Estimation of the traveling wave speed in pancreatic islets using high period periodic orbits. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005562.

Mondal:2024:CSO

[MD24] Sourav Mondal and Kinkar Chandra Das. Complete solution to open problems on exponential augmented Zagreb index of chemical trees. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004442.

Mozaffari:2020:FES

[MDJW20] Mohammad Reza Mozaffari, Fatemeh Dadkhah, Josef Jablonsky, and Peter Fernandes Wanke. Finding efficient surfaces in DEA-R models. Applied Mathematics and Computation, 386(??):Article 125497, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304550.

Maurya:2020:ESL

[MDSS20] Rahul Kumar Maurya, Vinita Devi, Nikhil Srivastava, and Vineet Kumar Singh. An efficient and stable Lagrangian matrix approach to Abel integral and integro-differential equations. Applied Mathematics and Computation, 374 (??):Article 125005, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031930997X.

Michel-Dansac:2022:TMS

[MDT22] Victor Michel-Dansac and Andrea Thomann. TVD-MOOD schemes based on implicit-explicit time integration. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004714.

Masrar:2021:NPS

[ME21a] Yassin Masrar and Mohamed Ettaouil. A nonlinear plane strain finite element analysis for multilayer elastomeric bearings. Applied Mathematics and Computation, 404(??): Article 126218, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003088.

Mehrez:2021:HEE

[ME21b] Zouhaier Mehrez and Afif El Cafsi. Heat exchange enhancement of ferrofluid flow into rectangular channel in the presence of a magnetic field. Applied Mathematics and Computation, 391 (??):Article 125634, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305889.

Meften:2021:CUS

[Mef21] Ghazi Abed Meften. Conditional and unconditional stability for double diffusive convection when the viscosity has a maximum. Applied Mathematics and Computation, 392(??):Article 125694, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306470.

Mandal:2021:OSE

[MF21] Parna Mandal and István Faragó. Operator splitting and error analysis in malaria modeling. Applied Mathematics and Computation, 410(??):Article 126446, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100535X.

Morais:2023:HLC

[MF23] J. Morais and M. Ferreira.

Hyperbolic linear canonical transforms of quaternion signals and uncertainty. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001406.

Moussa:2021:RDA

[MFA21] Kaouther Moussa, Mirko Fiacchini, and Mazen Alamir. Robust domain of attraction estimation for a tumor growth model. Applied Mathematics and Computation, 410(??):Article 126482, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005713.

Marzorati:2022:ECP

[MFK22] Denise Marzorati, Joaquin Fernández, and Ernesto Kofman. Efficient connection processing in equationbased object-oriented models. Applied Mathematics and Computation, 418 (??):Article 126842, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321009255.

Mei:2022:FTF

[MFS22] Zhen Mei, Ting Fang, and Hao Shen. Finite-time l_{2}–l_{∞} filtering for persistent dwelltime switched piecewiseaffine systems against deception attacks. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001722.

Mittal:2021:IRL

[MG21] Gaurav Mittal and Ankik Kumar Giri. Iteratively regularized Landweber iteration method: Convergence analysis via Hölder stability. Applied Mathematics and Computation, 392(??): Article 125744, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306974.

Maddah:2021:NRI

[MGD21] Sheyda Maddah, Modjtaba Ghorbani, and Matthias Dehmer. New results of identifying codes in product graphs. Applied Mathematics and Computation, 410(??):Article 126438, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005270.

Martinez-Guerra:2021:ARE

[MGFF21]  Rafael  Martínez-Guerra

and Juan Pablo Flores-Flores. An algorithm for the robust estimation of the COVID-19 pandemic’s population by considering undetected individuals. Applied Mathematics and Computation, 405(??): Article 126273, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003623.

Mekoth:2021:FTR

[MGJ21] Chitra Mekoth, Santhosh George, and P. Jidesh. Fractional Tikhonov regularization method in Hilbert scales. Applied Mathematics and Computation, 392(??): Article 125701, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306548.

Mckenna:2023:NRS

[MGK23] James Mckenna, Vassilis Glenis, and Chris Kilsby. A new Riemann solver for modelling bridges in flood flows — development and experimental validation. Applied Mathematics and Computation, 447(??): ??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000395.

Ma:2021:SOA

[MGM21] Ran Ma, Sainan Guo, and Cuixia Miao. A semionline algorithm and its competitive analysis for parallel-machine scheduling problem with rejection. Applied Mathematics and Computation, 392(??): Article 125670, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306238.

Ma:2020:CIM

[MGS20] Haifeng Ma, Xiaoshuang Gao, and Predrag S. Stanimirović. Characterizations, iterative method, sign pattern and perturbation analysis for the DMP inverse with its applications. Applied Mathematics and Computation, 378 (??):Article 125196, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030165X.

Marrero:2020:IFE

[MH20] J. Abderramán Marrero and D. A. Aiat Hadj. Improving formulas for the eigenvalues of finite block-Toeplitz tridiagonal matrices. Applied Mathematics and Computation, 382 (??):Article 125324, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302903.

Matejas:2022:HRA

[MH22] Josip Matejas and Vjeran Hari. The high relative accuracy of the HZ method. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004325.

Mo:2024:LHP

[MH24] Fei Mo and Wenchen Han. Long homogeneous payoff records with the latest strategy promotes the cooperation. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002509.

Mehmood:2021:GTD

[MHAT21] Mehwish Mehmood, Muhammad Hamid, Shumaila Ashraf, and Zhenfu Tian. Galerkin time discretization for transmission dynamics of HBV with nonlinear saturated incidence rate. Applied Mathematics and Computation, 410 (??):Article 126481, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005701.

Molnar:2023:RMI

[MHF23] Grant Molnar, Caroline Hammond, and Feng Fu. Reactive means in the iterated Prisoner’s dilemma. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003703.

Mittal:2024:NIF

[Mit24]  Gaurav Mittal.  Non-

stationary iterated frozen Tikhonov regularization with uniformly convex penalty terms for solving inverse problems. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006884.

Miyajima:2023:FVC

[Miy23]  Shinya Miyajima.  Fast

verified computation for real powers of large matrices with Kronecker structure. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002242.

Mizerski:2021:RGA

[Miz21] Krzysztof A. Mizerski.

Renormalization group analysis of the turbulent hydromagnetic dynamo: Effect of anisotropy. Applied Mathematics and Computation, 405(??):Article 126252, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003428.

Muszkieta:2023:CSA

[MJ23] Monika Muszkieta and Joanna Janczura. A compressed sensing approach to interpolation of fractional Brownian trajectories for a single particle tracking experiment. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000693.

Meena:2022:MCF

[MJK22] Om Prakash Meena, Pranitha Janapatla, and Ganesh Kumar K. Mixed convection flow over a vertical cone saturated porous medium with double dispersion effect. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001564.

Mengibaev:2020:IID

[MJM20] Ulugbek Mengibaev, Xiaodan Jia, and Yeqing Ma. The impact of interactive dependence on privacy protection behavior based on evolutionary game. Applied Mathematics and Computation, 379 (??):Article 125231, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302009.

Muszkieta:2021:STF

[MJW21] Monika Muszkieta, Joanna Janczura, and Aleksander Weron. Simulation and tracking of fractional particles motion. From microscopy video to statistical analysis. A Brownian bridge approach. Applied Mathematics and Computation, 396(??):Article 125902, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308559.

Ma:2021:TPF

[MJZ21] Tianlong Ma, Xian’an Jin, and Fuji Zhang. Tutte polynomials of fanlike graphs with applications in benzenoid systems. Applied Mathematics and Computation, 411 (??):Article 126496, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005853.

Manna:2021:DBM

[MK21a] Santanu Manna and Akash Kumar. Dynamic behavior of multi-layer heterogeneous composite magnetoelastic structures for surface wave scattering. Applied Mathematics and Computation, 397(??):Article 125922, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308754.

Mohan:2021:PSS

[MK21b] Nishith Mohan and Nitu Kumari. Positive steady states of a SI epidemic model with cross diffusion. Applied Mathematics and Computation, 410 (??):Article 126423, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005129.

Moon:2023:GLP

[MKK23] Hwan Pyo Moon, Soo Hyun Kim, and Song-Hwa Kwon. Gauss–Legendre polynomial basis for the shape control of polynomial curves. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001649.

Maity:2022:EBM

[MKKG22] Manisha Maity, Santimoy Kundu, Raju Kumhar, and Shishir Gupta. An electromechanical based model for Love-type waves in anisotropic-porous-piezoelectric composite structure with interfacial imperfections. Applied Mathematics and Computation, 418(??):Article 126783, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008651.

Meimaris:2020:CFA

[MKP20] Antonios T. Meimaris, Ioannis A. Kougioumtzoglou, and Athanasios A. Pantelous. Closed-form approximate solutions for a class of coupled nonlinear stochastic differential equations. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306617.

Mu:2022:DCS

[MKZ22] Xiuqing Mu, Kai Kang, and Jing Zhang. Dualchannel supply chain coordination considering credit sales competition. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004945.

Ma:2020:SFJ

[ML20] Junjie Ma and Huilan Liu. A sparse fractional Jacobi– Galerkin–Levin quadrature rule for highly oscillatory integrals. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307672.

Ma:2022:DDF

[ML22] Zhen-Lei Ma and Xiao-Jian Li. Data-driven fault detection for large-scale network systems: a mixed optimization approach. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002181.

Meng:2020:DGM

[MM20a] Jian Meng and Liquan Mei. Discontinuous Galerkin methods of the non-selfadjoint Steklov eigenvalue problem in inverse scattering. Applied Mathematics and Computation, 381(??): Article 125307, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302733.

Moaaz:2020:NOC

[MM20b] Osama Moaaz and Ali Muhib. New oscillation criteria for nonlinear delay differential equations of fourthorder. Applied Mathematics and Computation, 377 (??):Article 125192, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301612.

Mackolil:2022:CSS

[MM22a] Joby Mackolil and B. Mahanthesh. Computational simulation of surface tension and gravitation-induced convective flow of a nanoliquid with cross-diffusion: an optimization procedure. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001928.

See [MM24].

Martins:2022:APW

[MM22b] Nuno F. M. Martins and Pedro Mota. An adapted plane waves method for heat conduction problems. Applied Mathematics and Computation, 415(??):Article 126689, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007736.

Mackolil:2024:CCS

[MM24] Joby Mackolil and B. Mahanthesh. Corrigendum to “Computational simulation of surface tension and gravitation-induced convective flow of a nanoliquid with cross-diffusion: an optimization procedure” [applied mathematics and computation 425 (2022) 127108]. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000195.

See [MM22a].

Mazlumi:2021:STP

[MMDD21] Fatemeh Mazlumi, Mehdi Mosharaf-Dehkordi, and Morteza Dejam. Simulation of two-phase incompressible fluid flow in highly heterogeneous porous media by considering localization assumption in multiscale finite volume method. Applied Mathematics and Computation, 390 (??):Article 125649, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306032.

Merino:2023:SMA

[MMG^{+}23] S. Merino, J. Martínez, F. Guzmán, F. J. Sánchez, R. Guzmán, M. Sidrach de Cardona, and J. D. Lara. Scalable method for administration of resource technologies under stochastic procedures. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200724X.

Mei:2020:FTC

[MMHD20] Keqi Mei, Li Ma, Runxin He, and Shihong Ding. Finite-time controller design of multiple integrator nonlinear systems with input saturation. Applied Mathematics and Computation, 372(??):Article 124986, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309786.

Mao:2020:MMC

[MMHX20] Fubing Mao, Lijia Ma, Qiang He, and Gaoxi Xiao. Match making in complex social networks. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309208.

Magdalena:2023:TLS

[MMK^{+}23] Ikha Magdalena, Indriana Marcela, Nadhira Karima, Gabriel Jonathan, Dhemi Harlan, and M. Bagus Adityawan. Two layer shallow water equations for wave attenuation of a submerged porous breakwater. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002655.

Marsala:2024:SFP

[MMM24] Michelangelo Marsala, Angelos Mantzaflaris, and Bernard Mourrain. G^{1} spline functions for point cloud fitting. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004484.

Martinez-Martinez:2020:CAS

[MMMBRS20] C. T. Martínez-Martínez, J. A. Méndez-Bermúdez, José M. Rodríguez, and José M. Sigarreta. Computational and analytical studies of the Randić index in Erdös–Rényi models. Applied Mathematics and Computation, 377(??): Article 125137, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301065.

Madaras:2024:PIV

[MMOS24] Tomás Madaras, Daniela Matisová, Alfréd Onderko, and Zuzana Sárosiová. Proper interval vertex colorings of graphs. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002741.

Marangi:2023:ORA

[MMR23] Carmela Marangi, Angela Martiradonna, and Stefania Ragni. Optimal resource allocation for spatiotemporal control of invasive species. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006877.

Moosavi:2021:AVN

[MMS21] Rouhollah Moosavi, Reza Moltafet, and Younes Shekari. Analysis of viscoelastic non-Newtonian fluid over a vertical forward-facing step using the Maxwell fractional model. Applied Mathematics and Computation, 401 (??):Article 126119, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001673.

Ma:2023:PBG

[MMS23] Haifeng Ma, Dijana Mosić, and Predrag S. Stanimirović. Perturbation bounds for the group inverse and its oblique projection. Applied Mathematics and Computation, 449(??): ??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001327.

Mehboob:2022:CAA

[MMUS22] Hira Mehboob, Khadija Maqbool, Hameed Ullah, and Abdul Majeed Siddiqui. Computational analysis of an axisymmetric flow of Jeffrey fluid in a permeable micro channel. Applied Mathematics and Computation, 418 (??):Article 126826, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009097.

Moore:2022:ADT

[MN22] Ryleigh A. Moore and Akil Narayan. Adaptive density tracking by quadrature for stochastic differential equations. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003721.

Moufawad:2023:NTM

[MN23] Sophie M. Moufawad and Nabil R. Nassif. Newton type methods for solving a Hasegawa–Mima plasma model. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003107.

Mathiyalagan:2022:OBO

[MNSR22] K. Mathiyalagan, A. Shree Nidhi, H. Su, and T. Renugadevi. Observer and boundary output feedback control for coupled ODEtransport PDE. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001801.

Migorski:2020:SPV

[MNZ20] Stanisław Migórski, Van Thien Nguyen, and Shengda Zeng. Solvability of parabolic variational-hemivariational inequalities involving spacefractional Laplacian. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306605.

Monteiro:2020:MGL

[Mon20] L. H. A. Monteiro. More guns, less crime? a dynamical systems approach. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307969.

Maciejewski:2020:NIG

[MP20] Andrzej J. Maciejewski and Maria Przybylska. Non-integrability of generalised Charlier and Saint-Germain problem. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930712X.

Masroor:2022:ABO

[MP22] Suhaib Masroor and Chen Peng. An agent based observer model of the networked DC drives for speed coordination. Applied Mathematics and Computation, 413(??):Article 126633, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007177.

Morais:2023:RQM

[MP23] J. Morais and R. Michael Porter. Reduced-quaternionic Mathieu functions, timedependent moisil-teodorescu operators, and the imaginarytime wave equation. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006610.

Meng:2024:OBC

[MPC^{+}24] Hao Meng, Denghao Pang, Jinde Cao, Yechen Guo, and Azmat Ullah Khan Niazi. Optimal bipartite consensus control for heterogeneous unknown multiagent systems via reinforcement learning. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002492.

Macias:2023:LCQ

[MPM23] E. M. Macías, R. Pérez, and H. J. Martínez. On the local convergence of a quasi-Newton method for solving matrix polynomial equations. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007469.

Morigi:2024:SSM

[MPRS24] Serena Morigi, Enza Pellegrino, Sara Remogna, and Maria Lucia Sampoli. SMART — Shape Modelling, Approximation and Refinability Towards new computational techniques. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006288.

Mijanovic:2023:NIB

[MPW23] Andjela Mijanović, Bozidar V. Popović, and Viktor Witkovský. A numerical inversion of the bivariate characteristic function. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200875X.

Ma:2020:CFB

[MPX^{+}20] Jiali Ma, Ju H. Park, Shengyuan Xu, Guozeng Cui, and Zhichun Yang. Command-filter-based adaptive tracking control for nonlinear systems with unknown input quantization and mismatching disturbances. Applied Mathematics and Computation, 377 (??):Article 125161, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301302.

Mi:2022:SIH

[MQ22] Wen Mi and Tao Qian. System identification of Hammerstein models by using backward shift algorithm. Applied Mathematics and Computation, 413 (??):Article 126620, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007049.

Ma:2023:ECN

[MQW23] Xiaojian Ma, Ji Quan, and Xianjia Wang. Evolution of cooperation with nonlinear environment feedback in repeated public goods game. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002254.

Mathiyalagan:2022:OBF

[MR22] K. Mathiyalagan and R. Ragul.

Observer-based finite-time dissipativity for parabolic systems with time-varying delays. Applied Mathematics and Computation, 413 (??):Article 126605, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006895.

Magdalena:2020:SHO

[MRR20] I. Magdalena, H. Q. Rif’atin, and D. E. Reeve. Seiches and harbour oscillations in a porous semi-closed basin. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308276.

Montijano:2020:NSE

[MRVC20] J. I. Montijano, L. Rández, M. Van Daele, and M. Calvo. On the numerical stability of the exponentially fitted methods for first order IVPs. Applied Mathematics and Computation, 379(??):Article 125249, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302186.

Mao:2021:ECI

[MRW21] Yajun Mao, Zhihai Rong, and Zhi-Xi Wu. Effect of collective influence on the evolution of cooperation in evolutionary prisoner’s dilemma games. Applied Mathematics and Computation, 392(??):Article 125679, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306329.

Mathiyalagan:2020:SOS

[MS20] Kalidass Mathiyalagan and G. Sangeetha. Second-order sliding mode control for nonlinear fractional-order systems. Applied Mathematics and Computation, 383(??):Article 125264, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302332.

Matsumoto:2021:TDC

[MS21a] Akio Matsumoto and Ferenc Szidarovszky. Time delays and chaos in two competing species revisited. Applied Mathematics and Computation, 395(??): Article 125862, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308158.

Mertens:2021:IHR

[MS21b] Teppo Mertens and Frank Sommen. Inversions for the Hua–Radon and the polarized Hua–Radon transform. Applied Mathematics and Computation, 411 (??):Article 126525, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006147.

Miklavic:2021:DUG

[MS21c] Stefko Miklavic and Primoz Sparl. Distanceunbalancedness of graphs. Applied Mathematics and Computation, 405(??):Article 126233, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003234.

Mosic:2021:RWG

[MS21d] Dijana Mosić and Predrag S. Stanimirović. Representations for the weak group inverse. Applied Mathematics and Computation, 397(??):Article 125957, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000059.

May:2022:DSN

[MS22a] Sandra May and Florian Streitbürger. DoD stabilization for non-linear hyperbolic conservation laws on cut cell meshes in one dimension. Applied Mathematics and Computation, 419(??):Article 126854, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009371.

Mosic:2022:EPW

[MS22b] Dijana Mosić and Predrag S. Stanimirović. Expressions and properties of weak core inverse. Applied Mathematics and Computation, 415(??):Article 126704, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007888.

Mukherjee:2022:MME

[MS22c] S. Mukherjee and G. C. Shit. Mathematical modeling of electrothermal couple stress nanofluid flow and entropy in a porous microchannel under injection process. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001941.

Miao:2024:BMW

[MS24] Suoxia Miao and Housheng Su. Behaviors of matrixweighted networks with antagonistic interactions. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006598.

Moghadam:2022:NSF

[MSB22] Amin Abrishami Moghadam, Ali R. Soheili, and Amir Saboor Bagherzadeh. Numerical solution of fourth-order BVPs by using Lidstonecollocation method. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001412.

Masse:2021:DWT

[MSDR21] A. Massé, F. Sommen, H. De Ridder, and T. Raeymaekers. Discrete Weierstrass transform in discrete Hermitian Clifford analysis. Applied Mathematics and Computation, 391 (??):Article 125641, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305956.

Mostafazadeh:2024:EUB

[MSE24] Mahdi Mostafazadeh, Sedaghat Shahmorad, and Fevzi Erdogan. Existence, uniqueness and blow-up of solutions for generalized autoconvolution Volterra integral equations. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000808.

Mosic:2021:WCO

[MSK21] Dijana Mosić, Predrag S. Stanimirović, and Vasilios N. Katsikis. Weighted composite outer inverses. Applied Mathematics and Computation, 411(??):Article 126493, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005828.

Ma:2021:SPU

[MSMK21] Haifeng Ma, Predrag S. Stanimirović, Dijana Mosić, and Ivan I. Kyrchei. Sign pattern, usability, representations and perturbation for the core-EP and weighted core-EP inverse. Applied Mathematics and Computation, 404(??): Article 126247, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003374.

Mirzaee:2021:ASS

[MSN21] Farshid Mirzaee, Erfan Solhi, and Shiva Naserifar. Approximate solution of stochastic Volterra integrodifferential equations by using moving least squares scheme and spectral collocation method. Applied Mathematics and Computation, 410(??):Article 126447, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005361.

Macekova:2024:IVC

[MSS24] Mária Maceková, Zuzana Sárosiová, and Roman Soták. Interval vertex coloring. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000316.

Ma:2022:REB

[MSZ22] Zheng Ma, Jiasheng Song, and Jianping Zhou. Reliable event-based dissipative filter design for discretetime system with dynamic quantization and sensor fault. Applied Mathematics and Computation, 418 (??):Article 126823, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009061.

Miyaji:2021:EFS

[MT21] Kohei Miyaji and Jun Tanimoto. The existence of fence-sitters relaxes the spatial prisoner’s dilemma and enhances network reciprocity. Applied Mathematics and Computation, 390 (??):Article 125624, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305786.

Montecinos:2023:ERS

[MT23] Gino I. Montecinos and Eleuterio F. Toro. ENO-ET: a reconstruction scheme based on extended ENO stencil and truncated highestorder term. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322008104.

Meszaros:2024:NIT

[MT24] András Mészáros and Miklós Telek. Numerical inverse transformation of double sided Laplace transform with parameter optimization. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001851.

Ma:2024:DFC

[MTFD24] Zhihan Ma, Ze Tang, Jianwen Feng, and Dong Ding. Distributed formation containment control for multiagent systems via dynamic event-triggering communication mechanism. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004193.

Malik:2022:SOF

[MTRA22] Saddam Hussain Malik, Muhammad Tufail, Muhammad Rehan, and Shakeel Ahmed. State and output feedback local control schemes for nonlinear discrete-time 2-D Roesser systems under saturation, quantization and slope restricted input. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000510.

Mondal:2022:ETP

[MUMS22] Arnab Mondal, Ranjit Kumar Upadhyay, Argha Mondal, and Sanjeev Kumar Sharma. Emergence of Turing patterns and dynamic visualization in excitable neuron model. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000960.

Mean:2022:TDO

[MUOF22] Sovanna Mean, Koichi Unami, Hisashi Okamoto, and Masayuki Fujihara. A thorough description of onedimensional steady open channel flows using the notion of viscosity solution. Applied Mathematics and Computation, 415 (??):Article 126730, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008122.

M:2021:SAS

[MV21a] Hariprasad M. and Murugesan Venkatapathi. Semianalytical solutions for eigenvalue problems of chains and periodic graphs. Applied Mathematics and Computation, 411(??):Article 126512, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006019.

Mitchell:2021:APN

[MV21b] S. L. Mitchell and M. Vynnycky. An accuracypreserving numerical scheme for parabolic partial differential equations subject to discontinuities in boundary conditions. Applied Mathematics and Computation, 400(??):Article 125979, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000278. See

corrigendum [MV23].

Moskovka:2022:FME

[MV22] Alexej Moskovka and Jan Valdman. Fast MATLAB evaluation of nonlinear energies using FEM in 2D and 3D: Nodal elements. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001345.

Mitchell:2023:CAP

[MV23] S. L. Mitchell and M. Vynnycky. Corrigendum to ‘An accuracy-preserving numerical scheme for parabolic partial differential equations subject to discontinuities in boundary conditions’ [Applied Mathematics and Computation 400 (2021) 125979]. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0096300322008128.

See [MV21b].

Maslovskaya:2024:TNS

[MVCK24] A. G. Maslovskaya, E. M.

Veselova, A. Yu. Chebotarev, and A. E. Kovtanyuk. Theoretical and numerical study of the Landau–Khalatnikov model describing a formation of 2D domain patterns in ferroelectrics. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006409.

Mohammadi:2021:NLN

[MVO21] Masoud Mohammadi, Shidvash Vakilipour, and Scott Ormiston. Newton linearization of the Navier– Stokes equations for flow computations using a fully coupled finite volume method. Applied Mathematics and Computation, 397(??):Article 125916, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308699.

Michalski:2020:ENI

[MW20] Adrian Michalski and Iwona Włoch. On the existence and the number of independent (1, 2)-dominating sets in the G-join of graphs. Applied Mathematics and Computation, 377 (??):Article 125155, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301247.

Manebo:2024:EGB

[MWDD24] Wubeshet Seyoum Manebo, Mesfin Mekuria Woldaregay, Tekle Gemechu Dinka, and Gemechis File Duressa. An equidistributed gridbased second-order scheme for a singularly perturbed Fredholm integro-differential equation with an interior layer. Applied Mathematics and Computation, 464(??): ??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005672.

Mei:2023:AET

[MWS23] Yu Mei, Guanqi Wang, and Hao Shen. Adaptive eventtriggered L_{2}–L_{∞} control of semi-Markov jump distributed parameter systems. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006592.

Ma:2021:AFT

[MWZX21] Dazhong Ma, Tianbiao Wang, Huaguang Zhang, and Xiangpeng Xie. Adaptive fault-tolerant output regulation of linear systems with unknown dynamics and actuator faults. Applied Mathematics and Computation, 402 (??):Article 126132, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001806.

Ma:2024:HOD

[MXF^{+}24] Haoxiang Ma, Shixun Xiong, Zhumu Fu, Fazhan Tao, and Baofeng Ji. Highorder disturbance observerbased safe tracking control for a class of uncertain MIMO nonlinear systems with time-varying full state constraints. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005994.

Min:2020:ATC

[MXY^{+}20] Huifang Min, Shengyuan Xu, Xin Yu, Shumin Fei, and Guozeng Cui. Adaptive tracking control for stochastic nonlinear systems with full-state constraints and unknown covariance noise. Applied Mathematics and Computation, 385 (??):Article 125397, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303520.

Mao:2022:BSR

[MY22] Xianpeng Mao and Yuning Yang. Best sparse rank-1 approximation to higher-order tensors via a truncated exponential induced regularizer. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005070.

Ma:2024:EST

[MY24] Xiaoxu Ma and Yujun Yang. Enumeration of spanning trees containing perfect matchings in hexagonal chains with a unique kink. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001942.

Ma:2023:AHD

[MYC^{+}23] Yongqi Ma, Yunxiang You, Ke Chen, Lili Hu, and Aichun Feng. Application of harmonic differential quadrature (HDQ) method for vibration analysis of pipes conveying fluid. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006865.

Ma:2021:CMM

[MYW21] Xindong Ma, Yue Yu, and Lifeng Wang. Complex mixed-mode vibration types triggered by the pitchfork bifurcation delay in a driven van der Pol–Duffing oscillator. Applied Mathematics and Computation, 411 (??):Article 126522, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006111.

Mo:2021:CCM

[MYY21] Lipo Mo, Xiaolin Yuan, and Yongguang Yu. Containment control for multi-agent systems with fractional Brownian motion. Applied Mathematics and Computation, 398(??):Article 125814, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307670.

Ma:2022:SRT

[MZ22] Yingbin Ma and Wenhan Zhu. Some results on the 3-total-rainbow index. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002120.

Ma:2023:SBN

[MZ23a] Xiaoqi Ma and Jin Zhang. Supercloseness in a balanced norm of the NIPG method on Shishkin mesh for a reaction diffusion problem. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008967.

Ma:2023:GSP

[MZ23b] Yingbin Ma and Xiaoxue Zhang. Graphs with (strong) proper connection numbers m − 3 and m − 4. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000127.

Ma:2024:ETB

[MZ24] Jinlong Ma and Hongfei Zhao. Evaluation and threshold-based mutual supervision promotes the evolution of cooperation on interdependent networks. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003242.

Meng:2020:AFT

[MZFX20] Xin Meng, Ding Zhai, Zhumu Fu, and Xiangpeng Xie. Adaptive fault tolerant control for a class of switched nonlinear systems with unknown control directions. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309051.

Miao:2024:NFM

[MZL24] Peng Miao, Yuhua Zheng, and Shuai Li. A new FXTZNN model for solving TVCS equation and application to pseudo-inverse of a matrix. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323005787.

Ming:2022:NSD

[MZLS22] Zhongyang Ming, Huaguang Zhang, Yuling Liang, and Hanguang Su. Nonzerosum differential games of continuous-time nonlinear systems with uniformly ultimately ϵ-bounded by adaptive dynamic programming. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003228.

Mosic:2024:EMM

[MZS24] Dijana Mosić, Daochang Zhang, and Predrag S. Stanimirović. An extension of the MPD and MP weak group inverses. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005982.

Mu:2021:RNS

[MZSW21] Yunfei Mu, Huaguang Zhang, Hanguang Su, and Yingchun Wang. Robust normalization and H_{∞} stabilization for uncertain Takagi–Sugeno fuzzy singular systems with timedelays. Applied Mathematics and Computation, 388 (??):Article 125534, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304902.

Ma:2021:ROS

[MZW^{+}21] Gang Ma, Junjun Zheng, Ju Wei, Shilei Wang, and Yefan Han. Robust optimization strategies for seller based on uncertainty sets in context of sequential auction. Applied Mathematics and Computation, 390 (??):Article 125650, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306044.

Ma:2021:MHW

[MZX21] Xiao Ma, Bo Zhou, and Shifeng Xue. A meshless Hermite weighted leastsquare method for piezoelectric structures. Applied Mathematics and Computation, 400(??):Article 126073, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001211.

Ma:2022:HIE

[MZX22] Xiao Ma, Bo Zhou, and Shifeng Xue. A Hermite interpolation elementfree Galerkin method for functionally graded structures. Applied Mathematics and Computation, 419(??): Article 126865, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009486.

Ma:2022:ROD

[MZY^{+}22] Yan Ma, Zhenzhen Zhang, Li Yang, Hao Chen, and Yihao Zhang. A resilient optimized dynamic eventtriggered mechanism on networked control system with switching behavior under mixed attacks. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003745.

Mu:2023:NDA

[MZYW23] Yunfei Mu, Huaguang Zhang, Yuqing Yan, and Yingchun Wang. A novel design approach to state and fault estimation for interconnected systems using distributed observer. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001352.

Nikan:2021:LTB

[NA21] O. Nikan and Z. Avazzadeh. A localisation technique based on radial basis function partition of unity for solving Sobolev equation arising in fluid dynamics. Applied Mathematics and Computation, 401 (??):Article 126063, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001119.

Nakatsu:2023:DFR

[Nak23] Tomonori Nakatsu. On density functions related to discrete time maximum of some one-dimensional diffusion processes. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007408.

Nosrati:2023:EFS

[NBTP23] Komeil Nosrati, Juri Belikov, Aleksei Tepljakov, and Eduard Petlenkov. Extended fractional singular Kalman filter. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001194.

Niu:2020:DPC

[NCY^{+}20] He Niu, Yuyou Chen, Hang Ye, Hong Zhang, Yan Li, and Shu Chen. Distinguishing punishing costly signals from nonpunishing costly signals can facilitate the emergence of altruistic punishment. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309373.

Ndogmo:2023:GCE

[Ndo23] J. C. Ndogmo. Group classification and exact solutions of a class of nonlinear waves. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008372.

Nasresfahani:2021:EAF

[NE21] F. Nasresfahani and M. R.

Eslahchi. Error analysis of finite difference/collocation method for the nonlinear coupled parabolic free boundary problem modeling plaque growth in the artery. Applied Mathematics and Computation, 405(??): Article 126221, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003118.

Nadim:2021:STP

[NGC21] Sk Shahid Nadim, Indrajit Ghosh, and Joydev Chattopadhyay. Short-term predictions and prevention strategies for COVID-19: a model-based study. Applied Mathematics and Computation, 404(??):Article 126251, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003416.

Navarro-Garcia:2023:CSR

[NGGD23] Manuel Navarro-García, Vanesa Guerrero, and María Durban. On constrained smoothing and out-of-range prediction using P -splines: a conic optimization approach. Applied Mathematics and Computation, 441 (??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007470.

Niu:2022:FTS

[NGS22] Yichun Niu, Ming Gao, and Li Sheng. Faulttolerant state estimation for stochastic systems over sensor networks with intermittent sensor faults. Applied Mathematics and Computation, 416(??):Article 126723, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008079.

Nguyen:2021:MRI

[NH21] Kien Trung Nguyen and Nguyen Thanh Hung. The minmax regret inverse maximum weight problem. Applied Mathematics and Computation, 407 (??):Article 126328, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004173.

Ning:2022:HOF

[NH22] Jinghua Ning and Changchun Hua. H_{∞} output feedback control for fractional-order T-S fuzzy model with timedelay. Applied Mathematics and Computation, 416(??): Article 126736, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008183.

Naderi:2020:MRS

[NHBJ20] Mehrdad Naderi, Farzane Hashemi, Andriette Bekker, and Ahad Jamalizadeh. Modeling right-skewed financial data streams: a likelihood inference based on the generalized Birnbaum-Saunders mixture model. Applied Mathematics and Computation, 376(??):Article 125109, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300783.

Nazeer:2022:TSM

[NHK^{+}22] Mubbashar Nazeer, Farooq Hussain, M. Ijaz Khan, Asad ur Rehman, Essam Roshdy El-Zahar, Yu-Ming Chu, and M. Y. Malik. Theoretical study of MHD electro-osmotically flow of third-grade fluid in micro channel. Applied Mathematics and Computation, 420 (??):Article 126868, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009516. See

retraction notice [NHK^{+}23].

Nazeer:2023:RNT

[NHK^{+}23] Mubbashar Nazeer, Farooq Hussain, M. Ijaz Khan, Asad ur Rehman, Essam Roshdy El-Zahar, Yu-Ming Chu, and M. Y. Malik. Retraction notice to “Theoretical study of MHD electro-osmotically flow of third-grade fluid in micro channel” [Applied Mathematics and Computation 420(2022) / 126868]. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001674.

See [NHK^{+}22].

Neoh:2021:TEN

[NI21] S. S. Neoh and F. Ismail. Time-explicit numerical methods for Maxwell’s equation in second-order form. Applied Mathematics and Computation, 392(??): Article 125669, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306226.

Nguyen:2020:OBC

[NK20] Khanh Hieu Nguyen and Sung Hyun Kim. Observerbased control design of semi-Markovian jump systems with uncertain probability intensities and modetransition-dependent sojourntime distribution. Applied Mathematics and Computation, 372(??):Article 124968, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309609.

Nguyen:2021:AHO

[NK21a] Ngoc Hoai An Nguyen and Sung Hyun Kim. Asynchronous H_{∞} observerbased control synthesis of nonhomogeneous Markovian jump systems with generalized incomplete transition rates. Applied Mathematics and Computation, 411(??):Article 126532, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006160.

Nguyen:2021:ADC

[NK21b] Ngoc Hoai An Nguyen and Sung Hyun Kim. Asynchronous dissipative control design for semi-Markovian jump systems with uncertain probability distribution functions of sojourntime. Applied Mathematics and Computation, 397 (??):Article 125921, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308742.

Nguyen:2022:ISD

[NK22] Khanh Hieu Nguyen and Sung Hyun Kim. Improved sampled-data control design of T–S fuzzy systems against mismatched fuzzybasis functions. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002235.

Novikov:2020:RWM

[NKA20] A. Novikov, D. Kuzmin, and O. Ahmadi. Random walk methods for Monte Carlo simulations of Brownian diffusion on a sphere. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306629.

Niu:2023:SPF

[NKO^{+}23] Xiao-Dong Niu, Adnan Khan, Yi Ouyang, Mu-Feng Chen, De-Cai Li, and Hiroshi Yamaguchi. A simplified phase-field lattice Boltzmann method with a self-corrected magnetic field for the evolution of spike structures in ferrofluids. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200577X.

Nataj:2020:SCN

[NL20] Sarah Nataj and S. H.

Lui. Superlinear convergence of nonlinear conjugate gradient method and scaled memoryless BFGS method based on assumptions about the initial point. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308215.

Nie:2022:MAT

[NLP^{+}22] Yanyi Nie, Wenyao Li, Liming Pan, Tao Lin, and Wei Wang. Markovian approach to tackle competing pathogens in simplicial complex. Applied Mathematics and Computation, 417(??): Article 126773, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008559.

Nie:2023:SPV

[NMS23] Kairui Nie, Yingbin Ma, and Elzbieta Sidorowicz. (Strong) proper vertex connection of some digraphs. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004125.

Nguetcho:2021:AED

[NNB21] Aurélien Serge Tchakoutio Nguetcho, Guy Merlin Nkeumaleu, and Jean Marie Bilbault. Anharmonic effects on the dynamic behaviors of Klein Gordon models. Applied Mathematics and Computation, 403(??):Article 126136, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001843.

Njagarah:2021:SAT

[NNKH21] John B. H. Njagarah, Farai Nyabadza, Moatlhodi Kgosimore, and Cang Hui. Significance of antiviral therapy and CTL-mediated immune response in containing hepatitis B and C virus infection. Applied Mathematics and Computation, 397(??):Article 125926, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308791.

Nigam:2024:NLI

[NNN24] Ritu Nigam, Nilofar Nahid, and Gnaneshwar Nelakanti. Non-linear integral equations on unbounded domain with global polynomials. Applied Mathematics and Computation, 471(??): ??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000602.

Nandal:2020:NSN

[NP20] Sarita Nandal and Dwijendra Narain Pandey. Numerical solution of nonlinear fourth order fractional sub-diffusion wave equation with time delay. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308926.

Neupane:2022:CRS

[NPE22] Ram C. Neupane, James A. Powell, and Thomas C. Edwards. Connecting regionalscale tree distribution models with seed dispersal kernels. Applied Mathematics and Computation, 412 (??):Article 126591, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006755.

Nazarimehr:2020:MCS

[NPJ^{+}20] Fahimeh Nazarimehr, Shirin Panahi, Mahdi Jalili, Matjaž Perc, Sajad Jafari, and Brigita Ferčec. Multivariable coupling and synchronization in complex networks. Applied Mathematics and Computation, 372(??): Article 124996, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309889.

Nockowska-Rosiak:2022:MDU

[NRP22] Magdalena Nockowska-Rosiak and Christian Pötzsche. Monotonicity and discretization of Urysohn integral operators. Applied Mathematics and Computation, 414 (??):Article 126686, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007700.

Nannyonga:2020:IHE

[NS20] Betty K. Nannyonga and Martin Singull. Impact of health education on knowledge and behaviours toward obstetric fistula among women of reproductive age in Uganda. Applied Mathematics and Computation, 372(??):Article 124997, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309890.

Nelson:2021:DCS

[NS21] Nicholas J. Nelson and Eric Strauss. Dynamical chaos in a simple model of a knuckleball. Applied Mathematics and Computation, 391 (??):Article 125661, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306147.

Nandi:2022:SOA

[NS22] S. Nandi and Y. V. S. S. Sanyasiraju. A second order accurate fixed-grid method for multi-dimensional Stefan problem with moving phase change materials. Applied Mathematics and Computation, 416(??): Article 126719, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008031.

Niu:2023:TPT

[NSD^{+}23] Jing Niu, Tomohiro Sogabe, Lei Du, Tomoya Kemmochi, and Shao-Liang Zhang. Tensor product-type methods for solving Sylvester tensor equations. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003247.

Nithya:2020:DBN

[NSSA20] V. Nithya, R. Sakthivel, R. Sakthivel, and Faris Alzahrani. Dissipativebased non-fragile filtering for fuzzy networked control systems with switching communication channels. Applied Mathematics and Computation, 373 (??):Article 125011, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310033.

Naz:2023:SMH

[NT23] Rehana Naz and Mariano Torrisi. Symmetry methods for a hyperbolic model for a class of populations. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007123.

Negi:2024:NSE

[NT24] Shekhar Singh Negi and Vicenç Torra. A note on Sugeno exponential function with respect to distortion. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000584.

Nguyen:2024:NSG

[NTLNX24] Nam V. Nguyen, Kim Q. Tran, Jaehong Lee, and H. Nguyen-Xuan. Nonlocal strain gradient-based isogeometric analysis of graphene platelets-reinforced functionally graded triply periodic minimal surface nanoplates. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300323006306.

Nguyen-Thu:2024:MSI

[NTNT24] Huong Nguyen-Thu, Kien Trung Nguyen, and Nguyen Thanh Toan. The max-sum inverse median location problem on trees with budget constraint. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004654.

Nguyen:2021:SDR

[NTT21] Cuong M. Nguyen, Chee Pin Tan, and Hieu Trinh. State and delay reconstruction for nonlinear systems with input delays. Applied Mathematics and Computation, 390 (??):Article 125609, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305646.

Nasser:2021:CCI

[NV21] Mohamed M. S. Nasser and Matti Vuorinen. Computation of conformal invariants. Applied Mathematics and Computation, 389 (??):Article 125617, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305713.

Nepal:2023:AFD

[NWM23] Surendra Nepal, Yosief Wondmagegne, and Adrian Muntean. Analysis of a fully discrete approximation to a moving-boundary problem describing rubber exposed to diffusants. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008013.

Niu:2021:BBH

[NX21] Ruichao Niu and Min Xu.

The bipanconnectivity of bipartite hypercube-like networks. Applied Mathematics and Computation, 389 (??):Article 125564, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305208.

Nie:2023:MMD

[NX23] Kairui Nie and Kexiang Xu. Mixed metric dimension of some graphs. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008050.

Niu:2023:NCC

[NXG23] Minyao Niu, Jinghua Xiao, and You Gao. New constructions of constant dimension codes by improved inserting construction. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000541.

Niu:2021:TDC

[NXL21] Ruichao Niu, Min Xu, and Hong-Jian Lai. Twodisjoint-cycle-cover vertex bipancyclicity of the bipartite generalized hypercube. Applied Mathematics and Computation, 400(??): Article 126090, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001387.

Noorizadegan:2024:PER

[NYHC24] A. Noorizadegan, D. L.

Young, Y. C. Hon, and C. S. Chen. Powerenhanced residual network for function approximation and physics-informed inverse problems. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003710.

Nogales:2020:RBI

[NZ20] Jose M. Sanz Nogales and S. Zazo. Replicator based on imitation for finite and arbitrary networked communities. Applied Mathematics and Computation, 378(??):Article 125166, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301351.

Nie:2022:HCB

[NZLW22] Yanyi Nie, Xiaoni Zhong, Tao Lin, and Wei Wang. Homophily in competing behavior spreading among the heterogeneous population with higher-order interactions. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004544.

Olivares:2021:TSI

[OBP21] Manuela Bastidas Olivares, Carina Bringedal, and Iuliu Sorin Pop. A two-scale iterative scheme for a phasefield model for precipitation and dissolution in porous media. Applied Mathematics and Computation, 396(??): Article 125933, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308869.

Ozdemir:2024:MHS

[ÖC24] Zehra Özdemir and Hazal Ceyhan. Multiplicative hyperbolic split quaternions and generating geometric hyperbolical rotation matrices. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003230.

Ozat:2023:LBA

[ÖÇÖ23] Zeynep Özat, Bayram Çekim, and Mehmet Ali Özarslan. Δ_{ω}-Laguerre based Appell polynomials and their properties associated with some special polynomials. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003053.

Ortiz-Gracia:2020:ESC

[OG20]  Luis Ortiz-Gracia.  Ex-

pected shortfall computation with multiple control variates. Applied Mathematics and Computation, 373 (??):Article 125018, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310100.

Obmann:2024:CRC

[OH24] Daniel Obmann and Markus Haltmeier. Convergence rates for critical point regularization. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300324000778.

Onitsuka:2022:ASC

[OiEF22] Masakazu Onitsuka and Iz iddine El-Fassi. On approximate solutions of a class of Clairaut’s equations. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200279X.

Oyetunbi:2021:ACE

[OK21] Dolapo Muhammed Oyetunbi and Abdul Rahim Khan. Approximating common endpoints of multivalued generalized nonexpansive mappings in hyperbolic spaces. Applied Mathematics and Computation, 392 (??):Article 125699, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306524.

Oslandsbotn:2022:SSM

[OKN^{+}22] Andreas Oslandsbotn, Zeljko Kereta, Valeriya Naumova, Yoav Freund, and Alexander Cloninger. StreaMRAK a streaming multi-resolution adaptive kernel algorithm. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001965.

Ovadia:2024:CDR

[OKT24] Oded Ovadia, Adar Kahana, and Eli Turkel. A convolutional dispersion relation preserving scheme for the acoustic wave equation. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004861.

Osatakul:2023:DTR

[OLW23] Dhiti Osatakul, Shuanming Li, and Xueyuan Wu. Discrete-time risk models with surplus-dependent premium corrections. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005690.

Obradovic:2022:PHH

[OMS22] Aleksandar Obradović, Zoran Mitrović, and Slavisa Salinić. On the problem of a heavy homogeneous ball rolling without slipping over a fixed surface of revolution. Applied Mathematics and Computation, 420(??): Article 126906, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009899.

Okhli:2021:CED

[ON21] Kheirolah Okhli and Mehdi Jabbari Nooghabi. On the contaminated exponential distribution: a theoretical Bayesian approach for modeling positive-valued insurance claim data with outliers. Applied Mathematics and Computation, 392(??): Article 125712, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306652.

Oliveira:2022:SFC

[OPL22] Pedro M. Oliveira, Jonathan M. Palma, and Márcio J. Lacerda. H∈ state-feedback control for discrete-time cyber-physical uncertain systems under DoS attacks. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001758.

Okrasinska-Plociniczak:2022:SOS

[OPP22] Hanna Okrasińska-Płociniczak and Lukasz Płociniczak. Second order scheme for self-similar solutions of a time-fractional porous medium equation on the half-line. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322001199.

Orive:2020:EBG

[OPS20] Ramón Orive, Aleksandar V. Pejcev, and Miodrag M. Spalević. The error bounds of Gauss quadrature formulae for the modified weight functions of Chebyshev type. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307982.

Oruc:2021:RBF

[Oru21] Ömer Oruç. A radial basis function finite difference (RBF–FD) method for numerical simulation of interaction of high and low frequency waves: Zakharov–Rubenchik equations. Applied Mathematics and Computation, 394 (??):Article 125787, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307402. See

corrigendum [Oru22].

Oruc:2022:CRB

[Oru22] Ömer Oruç. Corrigendum to ‘A radial basis function finite difference (RBF–FD) method for numerical simulation of interaction of high and low frequency waves: Zakharov–Rubenchik equations’ [Applied Mathematics and Computation 394 (2021) 125787]. Applied Mathematics and Computation, 418(??):Article 126852, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009358. See

[Oru21].

Occorsio:2020:UWA

[OT20] Donatella Occorsio and Woula Themistoclakis. Uniform weighted approximation on the square by polynomial interpolation at Chebyshev nodes. Applied Mathematics and Computation, 385(??):Article 125457, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030415X.

Ozkan:2021:BTM

[ÖY21] Yesim Saglam Özkan and Emrullah Yasar. Breathertype and multi-wave solutions for (2+1)-dimensional nonlocal Gardner equation. Applied Mathematics and Computation, 390 (??):Article 125663, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306160.

Papuc:2021:DPD

[Pap21] Ioan Papuc. On a Dirichlet problem for the Darcy– Forchheimer–Brinkman system with application to lid-driven porous cavity flow with internal square block. Applied Mathematics and Computation, 402 (??):Article 125906, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320308596.

Peng:2022:SHF

[PAY22] Linyu Peng, Noriyoshi Arai, and Kenji Yasuoka. A stochastic Hamiltonian formulation applied to dissipative particle dynamics. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002107.

Park:2023:SIS

[PBKS23] Sangbeom Park, Soyoon Bak, Philsu Kim, and Yunchang Seol. A semiimplicit semi-Lagrangian method for simulating immersed boundary motion under high inertia and elasticity. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004381.

Pany:2020:FEG

[PBM20] Ambit K. Pany, Saumya Bajpai, and Soumyarani Mishra. Finite element Galerkin method for 2D Sobolev equations with Burgers’ type nonlinearity. Applied Mathematics and Computation, 387(??): Article 125113, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300825.

Peixoto:2024:FEN

[PCN^{+}24] Márcia L. C. Peixoto, Pedro H. S. Coutinho, Anh-Tu Nguyen, Thierry-Marie Guerra, and Reinaldo M. Palhares. Fault estimation for nonlinear parametervarying time-delayed systems. Applied Mathematics and Computation, 465 (??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032300574X.

Pandey:2023:SSA

[PD23] Prashant Kumar Pandey and Ritesh Kumar Dubey. Sign stable arbitrary high order reconstructions for constructing non-oscillatory entropy stable schemes. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002680.

Paul:2023:MPF

[PDK23] Jayanta Paul, Ashok Das, and Jitendra Kumar. Moments preserving finite volume approximations for the non-linear collisional fragmentation model. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005689.

Peng:2020:EPF

[PGF20] Gang Peng, Zhiming Gao, and Xinlong Feng. An extremum-preserving finite volume scheme for convection-diffusion equation on general meshes. Applied Mathematics and Computation, 380(??):Article 125301, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302599.

Pal:2024:HBP

[PGKM24] Debjit Pal, Santu Ghorai, Dipak Kesh, and Debasis Mukherjee. Hopf bifurcation and patterns formation in a diffusive two prey-one predator system with fear in preys and help. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003886.

Pang:2023:ISB

[PGL^{+}23] Zhi-Feng Pang, Zhenyan Guan, Yue Li, Ke Chen, and Hong Ge. Image segmentation based on the hybrid bias field correction. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002199.

Pimentel-Garcia:2021:EIP

[PGPCK21] Ernesto Pimentel-García, Carlos Parés, Manuel J. Castro, and Julian Koellermeier. On the efficient implementation of PVM methods and simple Riemann solvers. Application to the Roe method for large hyperbolic systems. Applied Mathematics and Computation, 388 (??):Article 125544, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305002.

Phan:2023:NVH

[PGR23] Duyen T. M. Phan, Sergey L. Gavrilyuk, and Giovanni Russo. Numerical validation of homogeneous multi-fluid models. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007615.

Peng:2023:SCN

[PGYW23] Yan Peng, Chao Ge, Hongjiu Yang, and Hong Wang. Stability and consistency for networked multiagent systems with doublelayered game under multiple DoS attacks. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003983.

Peng:2024:NNB

[PGZ^{+}24] Runlong Peng, Rongwei Guo, Bin Zheng, Zhonghua Miao, and Jin Zhou. Neural network-based robust consensus tracking for uncertain networked Euler– Lagrange systems with time-varying delays and output constraints. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006914.

Pan:2020:EMS

[PH20] Yubin Pan and Jin Huang.

Extrapolation method for solving two-dimensional volterral integral equations of the second kind. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307763.

Pan:2023:NCR

[PHL23] X. Z. Pan, J. J. Huang, and S. M. Lee. A novel convex relaxation technique on affine transformed sampleddata control issue for fuzzy semi-Markov jump systems. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001959.

Peng:2020:NOB

[PHS^{+}20] Zhinan Peng, Jiangping Hu, Kaibo Shi, Rui Luo, Rui Huang, Bijoy Kumar Ghosh, and Jiuke Huang. A novel optimal bipartite consensus control scheme for unknown multiagent systems via modelfree reinforcement learning. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308136.

Park:2021:SSC

[PJ21] Junpyo Park and Bongsoo Jang. Structural stability of coexistence in evolutionary dynamics of cyclic competition. Applied Mathematics and Computation, 394 (??):Article 125794, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307475.

Pan:2022:TBC

[PJMW22] Zhigang Pan, Lan Jia, Yiqiu Mao, and Quan Wang. Transitions and bifurcations in couple stress fluid saturated porous media using a thermal non-equilibrium model. Applied Mathematics and Computation, 415 (??):Article 126727, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008092.

Pikies:2020:CCC

[PK20] Tytus Pikies and Marek Kubale. Chromatic cost coloring of weighted bipartite graphs. Applied Mathematics and Computation, 375(??):Article 125073, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300424.

Pericaro:2020:ONA

[PKG^{+}20] Gislaine A. Periçaro, Elizabeth W. Karas, Clóvis C. Gonzaga, Débora C. Marcílio, Ana Paula Oening, Luiz Carlos Matioli, Daniel H. M. Detzel, Klaus de Geus, and Marcelo R. Bessa. Optimal non-anticipative scenarios for nonlinear hydro-thermal power systems. Applied Mathematics and Computation, 387(??):Article 124820, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308124.

Park:2022:ERA

[PKJB22] Sangbeom Park, Philsu Kim, Yonghyeon Jeon, and Soyoon Bak. An economical robust algorithm for solving 1D coupled Burgers’ equations in a semi-Lagrangian framework. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002594.

Park:2023:EAB

[PKL23] JunMin Park, Nam Kyu Kwon, and Seok Young Lee. Extended affine Bessel summation inequalities: Applications to stability analysis of linear discretetime systems with timevarying delays. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001947.

Pei:2022:IBA

[PLH22] Yifan Pei, Yongxin Lan, and Hua He. Improved bounds for anti-Ramsey numbers of matchings in outer-planar graphs. Applied Mathematics and Computation, 418 (??):Article 126843, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009267.

Pan:2024:LRS

[PLH^{+}24] Chenjian Pan, Chen Ling, Hongjin He, Liqun Qi, and Yanwei Xu. A low-rank and sparse enhanced Tucker decomposition approach for tensor completion. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300601X.

Peng:2022:CBA

[PLQ^{+}22] Xiao-Long Peng, Chun-Yan Li, Hong Qi, Gui-Quan Sun, Zhen Wang, and Yong-Ping Wu. Competition between awareness and epidemic spreading in homogeneous networks with demography. Applied Mathematics and Computation, 420(??): Article 126875, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009589.

Peng:2021:SSD

[PLRD21] Dongxue Peng, Xiaodi Li, R. Rakkiyappan, and Yanhui Ding. Stabilization of stochastic delayed systems: Event-triggered impulsive control. Applied Mathematics and Computation, 401 (??):Article 126054, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001028.

Peng:2020:DHT

[PLW^{+}20] Yu Peng, Qian Lu, Xue Wu, Yueru Zhao, and Yue Xiao. Dynamics of Hotelling triopoly model with bounded rationality. Applied Mathematics and Computation, 373(??): Article 125027, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310197.

Ponalagusamy:2020:STL

[PM20] R. Ponalagusamy and Ramakrishna Manchi. A study on two-layered (K.L– Newtonian) model of blood flow in an artery with six types of mild stenoses. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307593.

Palacios:2021:KCK

[PM21] José Luis Palacios and Greg Markowsky. Kemeny’s constant and the Kirchhoff index for the cluster of highly symmetric graphs. Applied Mathematics and Computation, 406(??):Article 126283, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003726.

Prabhu:2022:TVF

[PMAK22] S. Prabhu, V. Manimozhi, M. Arulperumjothi, and Sandi Klavzar. Twin vertices in fault-tolerant metric sets and fault-tolerant metric dimension of multistage interconnection networks. Applied Mathematics and Computation, 420(??): Article 126897, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009802.

Popovic:2020:LCG

[PMG20] Bozidar V. Popović, Andjela Mijanović, and Ali I. Genç. On linear combination of generalized logistic random variables with an application to financial returns. Applied Mathematics and Computation, 381(??):

Article 125314, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302800.

Parmar:2021:MPM

[PMP21] Rakesh K. Parmar, Gradimir V. Milovanović, and Tibor K. Pogány. Multi-parameter Mathieu, and alternating Mathieu series. Applied Mathematics and Computation, 400(??):Article 126099, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001478.

Prabaharan:2024:RON

[PMTT24] Natarajan Prabaharan, Mayakrishnan Madhan, Ethiraju Thandapani, and Ercan Tunç. Remarks on the oscillation of nonlinear third-order noncanonical delay differential equations. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004119.

Priyanka:2024:NFP

[PN24] K. Sri Raja Priyanka and G. Nagamani. Non-fragile projective synchronization of delayed discrete-time neural networks via generalized weighted summation inequality. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003461.

Panahi:2021:OSC

[PNJ^{+}21] Shirin Panahi, Fahimeh Nazarimehr, Sajad Jafari, Julien C. Sprott, Matjaž Perc, and Robert Repnik. Optimal synchronization of circulant and noncirculant oscillators. Applied Mathematics and Computation, 394(??):Article 125830, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307839.

Pastor:2020:ESC

[PNOdV20] Alejandro Pastor, Juan Carlos Nuño, José Olarrea, and Javier de Vicente. Enabling stable coexistence by modifying the environment. Applied Mathematics and Computation, 380(??): Article 125235, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302046.

Popova:2020:NPS

[Pop20] Evgenija D. Popova. New parameterized solution with application to bounding secondary variables in FE models of structures. Applied Mathematics and Computation, 378(??):Article 125205, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301740.

Pan:2021:NRS

[PP21] Jie Pan and Zhaoya Pan. Novel robust stability criteria for uncertain parameter quaternionic neural networks with mixed delays: Whole quaternionic method. Applied Mathematics and Computation, 407(??):Article 126326, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100415X.

Pucanovic:2023:CPC

[PP23] Zoran Pucanović and Marko Pesović. Chebyshev polynomials and r-circulant matrices. Applied Mathematics and Computation, 437 (??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322005951.

Pachon:2020:CTL

[PPS20] Angelica Pachon, Federico Polito, and Laura Sacerdote. On the continuoustime limit of the Barabási– Albert random graph. Applied Mathematics and Computation, 378(??):Article 125177, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301466.

Piotrowska:2023:NSP

[PPS23] Monika J. Piotrowska, Aleksandra Puchalska, and Konrad Sakowski. On the network suppression of the pathogen spread within the healthcare system. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003387.

Padma:2024:CMM

[PPS24] R. Padma, R. Ponalagusamy, and R. Tamil Selvi. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Appl. Math. Comput. 362 (2019) 124453]. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400078X.

See [RRR19].

Peng:2020:IHA

[PPZW20] Hao Peng, Wangxin Peng, Dandan Zhao, and Wei Wang. Impact of the heterogeneity of adoption thresholds on behavior spreading in complex networks. Applied Mathematics and Computation, 386 (??):Article 125504, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304628.

Pazanin:2020:EVD

[PR20] Igor Pazanin and Marko Radulović. Effects of the viscous dissipation on the Darcy–Brinkman flow: Rigorous derivation of the higher-order asymptotic model. Applied Mathematics and Computation, 386(??):Article 125479, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304380.

Palmowski:2024:GST

[PRP24] Zbigniew Palmowski, Lewis Ramsden, and Apostolos D. Papaioannou. Gerber–Shiu theory for discrete risk processes in a regime switching environment. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006604.

Przybylo:2024:DRG

[Prz24a] Jakub Przybyło. Decomposability of regular graphs to 4 locally irregular subgraphs. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003771.

Przybylo:2024:NDG

[Prz24b] Jakub Przybyło. A note on decomposing graphs to locally almost irregular subgraphs. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000560.

Pal:2021:ART

[PS21a] Mukesh Kumar Pal and Abhishek Kumar Singh. Analysis of reflection and transmission phenomenon at distinct bonding interfaces in a rotating prestressed functionally graded piezoelectric-orthotropic structure. Applied Mathematics and Computation, 409 (??):Article 126398, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004872.

Przybylowicz:2021:EUA

[PS21b] Paweł Przybyłowicz and Michaela Szölgyenyi. Existence, uniqueness, and approximation of solutions of jump-diffusion SDEs with discontinuous drift. Applied Mathematics and Computation, 403 (??):Article 126191, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002812.

Pradhan:2022:SWP

[PS22] Nirakara Pradhan and Sapan Kumar Samal. Surface waves propagation in a homogeneous liquid layer overlying a monoclinic halfspace. Applied Mathematics and Computation, 414 (??):Article 126655, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007396.

Pribylova:2024:DSS

[PSE24] Lenka Pribylová, Deeptajyoti Sen, and Veronika Eclerová. Destabilization of a seasonal synchronization in a population model with a seasonally varying Allee effect. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005003.

Peiris:2021:GRA

[PSSU21] V. Peiris, N. Sharon, N. Sukhorukova, and J. Ugon. Generalised rational approximation and its application to improve deep learning classifiers. Applied Mathematics and Computation, 389(??):Article 125560, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305166.

Pan:2021:EMS

[PSXX21] Kejia Pan, Hai-Wei Sun, Yuan Xu, and Yufeng Xu. An efficient multigrid solver for two-dimensional spatial fractional diffusion equations with variable coefficients. Applied Mathematics and Computation, 402(??):Article 126091, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001399.

Petrov:2021:PMF

[PT21] Miroslav S. Petrov and Todor D. Todorov. Properties of the multidimensional finite elements. Applied Mathematics and Computation, 391(??):Article 125695, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306482.

Psarrakos:2024:FLM

[PV24] Georgios Psarrakos and Polyxeni Vliora. On the flexibility and loading maximization for weighted premiums. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004053.

Poochinapan:2022:NAS

[PW22a] Kanyuta Poochinapan and Ben Wongsaijai. Numerical analysis for solving Allen– Cahn equation in 1D and 2D based on higher-order compact structure-preserving difference scheme. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004489.

Prins:2022:RCE

[PW22b] Peter J. Prins and Sander Wahls. Reliable computation of the eigenvalues of the discrete KdV spectrum. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004350.

Pan:2020:SDB

[PWH20] Qiuhui Pan, Linpeng Wang, and Mingfeng He. Social dilemma based on reputation and successive behavior. Applied Mathematics and Computation, 384 (??):Article 125358, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303222.

Pan:2023:IPG

[PWL23] Chenran Pan, Weifan Wang, and Runrun Liu. IC-planar graphs are odd-10-colorable. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001893.

Pandiya:2021:NPF

[PWSE21] Ridwan Pandiya, Widodo Widodo, Salmah, and Irwan Endrayanto. Non parameter-filled function for global optimization. Applied Mathematics and Computation, 391(??):Article 125642, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305968.

Peng:2023:WGF

[PWWZ23] Hui Peng, Ruishu Wang, Xiuli Wang, and Yongkui Zou. Weak Galerkin finite element method for linear elasticity interface problems. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006622.

Peng:2022:CDS

[PWZ22] Xiao Peng, Yijing Wang, and Zhiqiang Zuo. Codesign of state-dependent switching law and control scheme for variableorder fractional nonlinear switched systems. Applied Mathematics and Computation, 415(??):Article 126725, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007827.

Peng:2020:PMR

[PX20] Bo Peng and Hong-Kun Xu. Proximal methods for reweighted l_{Q}- regularization of sparse signal recovery. Applied Mathematics and Computation, 386(??):Article 125408, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303702.

Peng:2023:ATC

[PX23] Yanru Peng and Shengyuan Xu. Adaptive tracking control for a class of stochastic nonlinear systems with full-state constraints and dead-zone. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002114.

Pan:2021:FET

[PY21] Jiangmin Pan and Fugang Yin. Finite edge-transitive bi-circulants. Applied Mathematics and Computation, 405(??):Article 126268, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100357X.

Parras:2020:DAO

[PZ20] Juan Parras and Santiago Zazo. A distributed algorithm to obtain repeated games equilibria with discounting. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307775.

Pei:2024:USA

[PZS24] Lifang Pei, Chaofeng Zhang, and Dongyang Shi. Unconditional superconvergence analysis of two-grid nonconforming FEMs for the fourth order nonlinear extend Fisher–Kolmogorov equation. Applied Mathematics and Computation, 471(??): ??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000742.

Pan:2024:DBP

[PZZF24] Na Pan, Ziyan Zeng, Yuji Zhang, and Minyu Feng. Defined benefit pension plan inhibit the emergence of cooperation in the public goods games. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002923.

Qin:2020:AAC

[QCARL20] Bo-Wei Qin, Kwok-Wai Chung, Antonio Algaba, and Alejandro J. Rodríguez-Luis. Analytical approximation of cuspidal loops using a nonlinear time transformation method. Applied Mathematics and Computation, 373(??):Article 125042, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300114.

Quan:2023:RBP

[QCCW23] Ji Quan, Shihui Cui, Wenman Chen, and Xianjia Wang. Reputationbased probabilistic punishment on the evolution of cooperation in the spatial public goods game. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007718.

Quan:2024:RPD

[QCW24] Ji Quan, Xinyue Chen, and Xianjia Wang. Repeated prisoner’s dilemma games in multi-player structured populations with crosstalk. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400122X.

Qiao:2021:CNT

[QDL^{+}21] Honghai Qiao, Zhenghong Deng, Huijia Li, Jun Hu, Qun Song, and Chengyi Xia. Complex networks from time series data allow an efficient historical stage division of urban air quality information. Applied Mathematics and Computation, 410(??):Article 126435, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005245.

Quan:2022:RCB

[QDW22] Ji Quan, Xu Dong, and Xianjia Wang. Rational conformity behavior in social learning promotes cooperation in spatial public goods game. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001813.

Qin:2021:RAB

[QH21] Xiao-Wen Qin and Rong-Xia Hao. Reliability analysis based on the dual-CIST in shuffle-cubes. Applied Mathematics and Computation, 397(??):Article 125900, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308535.

Qiao:2021:SAP

[QHC21] Yanfen Qiao, Guolin Hou, and Alatancang Chen. Symplectic approach for plane elasticity problems of two-dimensional octagonal quasicrystals. Applied Mathematics and Computation, 400(??):Article 126043, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000916.

Qi:2022:MSF

[Qi22] Yingyu Qi. Multi-stencils fast marching method for factored eikonal equations with quadratic anisotropy. Applied Mathematics and Computation, 417(??):Article 126776, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008584.

Qiu:2020:SMA

[QL20] Zhiping Qiu and Dongliang Liu. Safety margin analysis of buckling for structures with unknown but bounded uncertainties. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307519.

Qu:2021:FDS

[QL21] Wei Qu and Zhi Li. Fast direct solver for CN–ADI–FV scheme to two-dimensional Riesz space-fractional diffusion equations. Applied Mathematics and Computation, 401(??):Article 126033, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000813.

Quan:2024:GMA

[QLC^{+}24] Hongzheng Quan, Xiujuan Lu, Chenxiao Cai, Yun Zou, and James Lam. A Gramian matrix approach to synthesizing finite-frequency H_{2} controller. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003850.

Qiao:2022:RSC

[QLG^{+}22] Dianliang Qiao, Zhiyang Lin, Mingmin Guo, Xiaoxia Yang, Xiaoyang Li, Peng Zhang, and Xiaoning Zhang. Riemann solvers of a conserved high-order traffic flow model with discontinuous fluxes. Applied Mathematics and Computation, 413(??):Article 126648, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007323.

Ling:2021:OSD

[qLgZ21] Xing qian Ling and Wei guo Zhang. Orbital stability of dn periodic solutions for the generalized symmetric regularized-long-wave equation. Applied Mathematics and Computation, 405(??): Article 126249, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003398.

Qian:2020:TCS

[QLKW20] Jin Qian, Hexiang Lin, Yufeng Kong, and Yuansong Wang. Tri-criteria single machine scheduling model with release times and learning factor. Applied Mathematics and Computation, 387(??):Article 124543, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319305260.

Qin:2020:RNS

[QLL20] Zhongmei Qin, Hui Lei, and Shasha Li. Rainbow numbers for small graphs in planar graphs. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930880X.

Qi:2024:SBC

[QLL24] Haokun Qi, Bing Liu, and Shi Li. Stability, bifurcation, and chaos of a stage-structured predator– prey model under fearinduced and delay. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002455.

Qin:2021:RNP

[QLLY21] Zhongmei Qin, Shasha Li, Yongxin Lan, and Jun Yue. Rainbow numbers for paths in planar graphs. Applied Mathematics and Computation, 397(??):Article 125918, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308717.

Qin:2021:DSR

[QLM21] Rui Qin, Dan Li, and Jixiang Meng. On the D_{α}-spectral radius of two types of graphs. Applied Mathematics and Computation, 396(??):Article 125898, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308511.

Qin:2020:CSG

[QLZL20] Zongrong Qin, Dingjun Lou, Huiquan Zhu, and Jun Liang. Characterization of k-subconnected graphs. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306125.

Qian:2022:DPD

[QLZL22] Wei Qian, Haibo Liu, Yunji Zhao, and Yalong Li. Delayprobability-dependent state estimation for neural networks with hybrid delays. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001023.

Qiao:2022:ESD

[QMS22] Hongwei Qiao, Jixiang Meng, and Eminjan Sabir. Embedding spanning disjoint cycles in enhanced hypercube networks with prescribed vertices in each cycle. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005550.

Qi:2023:FTZ

[QNX^{+}23] Zhaohui Qi, Yingqiang Ning, Lin Xiao, Jiajie Luo, and Xiaopeng Li. Finitetime zeroing neural networks with novel activation function and variable parameter for solving time-varying Lyapunov tensor equation. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002412.

Qiao:2021:AFD

[QPHC21] Baoping Qiao, Ziqiang Pan, Weichuan Huang, and Chengyin Cao. An adaptive finite-difference method for accurate simulation of firstarrival traveltimes in heterogeneous media. Applied Mathematics and Computation, 394(??):Article 125792, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307451.

Qiu:2023:SCD

[QQ23] Zhiping Qiu and Yu Qiu. Symplectic central difference scheme for quasilinear autonomous Birkhoffian systems. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322007329.

Qu:2022:BKA

[QQLZ22] Wei Qu, Tao Qian, Haichou Li, and Kehe Zhu. Best kernel approximation in Bergman spaces. Applied Mathematics and Computation, 416(??):Article 126749, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008316.

Qi:2022:AIC

[QQWJ22] Qingyuan Qi, Zhenghong Qiu, Xianghua Wang, and Zhijian Ji. Asymmetric information control for stochastic systems with different intermittent observations. Applied Mathematics and Computation, 412 (??):Article 126544, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006287.

Qi:2023:ETI

[QQY^{+}23] Yiwen Qi, Ziyu Qu, Zhaohui Yao, Xiujuan Zhao, and Yiwen Tang. Eventtriggered iterative learning control for asynchronously switched systems. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007305.

Qi:2020:WGM

[QS20] Wenya Qi and Lunji Song. Weak Galerkin method with implicit θ-schemes for second-order parabolic problems. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307234.

Qu:2021:NNM

[QSL21] Haidong Qu, Zihang She, and Xuan Liu. Neural network method for solving fractional diffusion equations. Applied Mathematics and Computation, 391 (??):Article 125635, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305890.

Quan:2023:EFB

[QSW23] Ji Quan, Yuang Shi, and Xianjia Wang. Effect of fairness-based sympathy and retaliation on cooperation in multi-player dilemma games. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001479.

Quenjel:2022:PSG

[Que22] El Houssaine Quenjel. Positive Scharfetter–Gummel finite volume method for convection-diffusion equations on polygonal meshes. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001552.

Qin:2022:LCA

[QWD22] Hongyu Qin, Fengyan Wu, and Deng Ding. A linearized compact ADI numerical method for the two-dimensional nonlinear delayed Schrödinger equation. Applied Mathematics and Computation, 412 (??):Article 126580, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006640.

Qiu:2021:NSF

[QXG21] Wenlin Qiu, Da Xu, and Jing Guo. Numerical solution of the fourth-order partial integro-differential equation with multi-term kernels by the sinc-collocation method based on the double exponential transformation. Applied Mathematics and Computation, 392(??): Article 125693, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306469.

Qiu:2022:IAL

[QXSW22] Xing Qiu, Tao Xu, Babak Soltanalizadeh, and Hulin Wu. Identifiability analysis of linear ordinary differential equation systems with a single trajectory. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003344.

Qu:2021:NMM

[QXX21] Hongli Qu, Dongxiu Xie, and Jie Xu. A numerical method on the mixed solution of matrix equation ∑_{i=1}^{t} A_{i}X_{i}B_{i} = E with sub-matrix constraints and its application. Applied Mathematics and Computation, 411 (??):Article 126460, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100549X.

Qian:2022:RAN

[QY22] Yunyou Qian and Dansheng Yu. Rates of approximation by neural network interpolation operators. Applied Mathematics and Computation, 418 (??):Article 126781, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008638.

Qiu:2022:SFI

[QYD22] Renjun Qiu, Liang Yan, and Xiaojun Duan. Solving Fredholm integral equation of the first kind using Gaussian process regression. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001187.

Qi:2020:SDS

[QYG^{+}20] Wenhai Qi, Xu Yang, Xianwen Gao, Jun Cheng, Yonggui Kao, and Yunliang Wei. Stability for delayed switched systems with Markov jump parameters and generally incomplete transition rates. Applied Mathematics and Computation, 365(??):124718, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307106.

Quan:2020:SED

[QYL^{+}20] Ji Quan, Wenjun Yang, Xia Li, Xianjia Wang, and Jian-Bo Yang. Social exclusion with dynamic cost on the evolution of cooperation in spatial public goods games. Applied Mathematics and Computation, 372 (??):Article 124994, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309865.

Qing:2024:PTB

[QYLW24] Nengneng Qing, Yongqing Yang, Xiaoli Luan, and Haiying Wan. Practical time-boundary consensus for fractional-order multiagent systems under wellknown and estimable topology. Applied Mathematics and Computation, 464(??): ??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005696.

Qin:2021:ESP

[QZ21] Zhongmei Qin and Junxue Zhang. Extremal stretch of proper-walk coloring of graphs. Applied Mathematics and Computation, 405 (??):Article 126240, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003301.

Qin:2022:GTQ

[QZ22] Junning Qin and Hongzhi Zhong. A Galerkin time quadrature element formulation for linear structural dynamics. Applied Mathematics and Computation, 413(??):Article 126609, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006937.

Quan:2024:RDS

[QZC^{+}24] Ji Quan, Xiyue Zhang, Wenman Chen, Caixia Tang, and Xianjia Wang. Reputation-dependent social learning on the evolution of cooperation in spatial public goods games. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002170.

Qi:2021:HFS

[QZH21] Yiwen Qi, Xiujuan Zhao, and Jie Huang. H_{∞} filtering for switched systems subject to stochastic cyber attacks: a double adaptive storage event-triggering communication. Applied Mathematics and Computation, 394 (??):Article 125789, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307426.

Quan:2020:IFB

[QZWY20] Ji Quan, Yawen Zhou, Xianjia Wang, and Jian-Bo Yang. Information fusion based on reputation and payoff promotes cooperation in spatial public goods game. Applied Mathematics and Computation, 368(??): ??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307970.

Qian:2023:CPP

[QZZ^{+}23] Jun Qian, Tongda Zhang, Yingfeng Zhang, Yueting Chai, Xiao Sun, and Zhen Wang. The construction of peer punishment preference: how central power shapes prosocial and antisocial punishment behaviors. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007913.

Rahimi:2024:CMS

[RAB24] Vaz’he Rahimi, Davood Ahmadian, and Luca Vincenzo Ballestra. Construction and mean-square stability analysis of a new family of stochastic Runge–Kutta methods. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000420.

Roy:2021:ASU

[RB21] Ashis Kumar Roy and O. Anwar Bég. Asymptotic study of unsteady mass transfer through a rigid artery with multiple irregular stenoses. Applied Mathematics and Computation, 410(??):Article 126485, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005749.

Rengamannar:2020:ESN

[RBPN20] Kaviya Rengamannar, Ganesh Priya Balakrishnan, Muthukumar Palanisamy, and Michal Niezabitowski. Exponential stability of non-linear stochastic delay differential system with generalized delay-dependent impulsive points. Applied Mathematics and Computation, 382(??):Article 125344, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303040.

Rohan:2021:MWD

[RC21] Eduard Rohan and Robert Cimrman. Modelling wave dispersion in fluid saturating periodic scaffolds. Applied Mathematics and Computation, 410 (??):Article 126256, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003349.

Rao:2022:AAF

[RC22] S. Chandra Sekhara Rao and Abhay Kumar Chaturvedi. Analysis of an almost fourth-order parameteruniformly convergent numerical method for singularly perturbed semilinear reaction-diffusion system with non-smooth source term. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000303.

Ramirez-Cobo:2021:AAL

[RCCL21] Pepa Ramírez-Cobo, Emilio Carrizosa, and Rosa E. Lillo. Analysis of an aggregate loss model in a Markov renewal regime. Applied Mathematics and Computation, 396(??):Article 125869, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308225.

Ren:2021:BAN

[RCG21] Weilin Ren, Rongjun Cheng, and Hongxia Ge. Bifurcation analysis for a novel heterogeneous continuum model considering electronic throttle angle changes with memory. Applied Mathematics and Computation, 401 (??):Article 126079, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001272.

Redmann:2022:FSA

[RD22] Martin Redmann and Igor Pontes Duff. Full state approximation by Galerkin projection reduced order models for stochastic and bilinear systems. Applied Mathematics and Computation, 420(??): Article 126561, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006457.

Rakshith:2023:DLS

[RD23] B. R. Rakshith and Kinkar Chandra Das. On distance Laplacian spectral determination of complete multipartite graphs. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008554.

Raviola:2024:PAS

[RD24] Lisandro A. Raviola and Mariano F. De Leo. Performance of affine-splitting pseudo-spectral methods for fractional complex Ginzburg– Landau equations. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005970.

Ramasamy:2022:EHO

[RDKR22] Mohanasubha Ramasamy, Subhasri Devarajan, Suresh Kumarasamy, and Karthikeyan Rajagopal. Effect of higher-order interactions on synchronization of neuron models with electromagnetic induction. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005215.

Ren:2023:FFF

[RDL23] Yingying Ren, Da-Wei Ding, and Yue Long. Finitefrequency fixed-order dynamic output-feedback control via a homogeneous polynomially parameterdependent technique. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007494.

Ramirez:2023:NMP

[REC^{+}23] Luis Ramírez, Laura Edreira, Iván Couceiro, Pablo Ouro, Xesús Nogueira, and Ignasi Colominas. A new Mean Preserving Moving Least Squares method for Arbitrary Order Finite Volume schemes. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008360.

Redmann:2021:OBM

[RF21] Martin Redmann and Melina A. Freitag. Optimization based model order reduction for stochastic systems. Applied Mathematics and Computation, 398(??):Article 125783, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307360.

Romanengo:2023:PPA

[RFFB23] Chiara Romanengo, Ulderico Fugacci, Bianca Falcidieno, and Silvia Biasotti. Piecewise polynomial approximation of spatial curvilinear profiles using the Hough transform. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300382X.

Ren:2021:AAP

[RFFZ21] Junchao Ren, Lihong Feng, Jun Fu, and Tianyu Zhuang. Admissibility analysis and passive output feedback control for one-sided Lipschitz nonlinear singular Markovian jump systems with uncertainties. Applied Mathematics and Computation, 409(??):Article 126405, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100494X.

Roul:2020:HON

[RG20] Pradip Roul and V. M.

K. Prasad Goura. A high order numerical method and its convergence for timefractional fourth order partial differential equations. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307192.

Roul:2022:SBS

[RG22] Pradip Roul and V. M.

K. Prasad Goura. A superconvergent B-spline technique for second order nonlinear boundary value problems. Applied Mathematics and Computation, 414 (??):Article 126615, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006998.

Rolania:2020:DTS

[RGAM20] D. Barrios Rolanía, J. C. García-Ardila, and D. Manrique. On the Darboux transformations and sequences of p-orthogonal polynomials. Applied Mathematics and Computation, 382(??):Article 125337, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303039.

Rather:2023:DLS

[RGS23] Bilal A. Rather, Hilal A. Ganie, and Yilun Shang. Distance Laplacian spectral ordering of sun type graphs. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000164.

Ren:2020:FTS

[RH20] Chengcheng Ren and Shuping He. Finite-time stabilization for positive Markovian jumping neural networks. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930623X.

Rajagopal:2024:EGG

[RHN^{+}24]  Karthikeyan  Rajagopal,

Dorsa Nezhad Hajian, Hayder Natiq, Yuexi Peng, Fatemeh Parastesh, and Sajad Jafari. Effect of Gaussian gradient in the medium’s action potential morphology on spiral waves. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000626.

Ran:2021:ESP

[RHNT21] Xue Ran, Lin Hu, Lin-Fei Nie, and Zhidong Teng. Effects of stochastic perturbation and vaccinated age on a vector-borne epidemic model with saturation incidence rate. Applied Mathematics and Computation, 394(??):Article 125798, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320307517.

Rajagopal:2021:MIC

[RHR^{+}21] Karthikeyan Rajagopal, Iqtadar Hussain, Zahra Rostami, Chunbiao Li, Viet-Thanh Pham, and Sajad Jafari. Magnetic induction can control the effect of external electrical stimuli on the spiral wave. Applied Mathematics and Computation, 390 (??):Article 125608, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305634.

Renders:2024:DTD

[RiTZ24] Jarne Renders, Shin ichi Tokunaga, and Carol T. Zamfirescu. Domination of triangulated discs and maximal outerplanar graphs. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001206.

Ren:2024:SCV

[RJHC24] Yue Ren, Haijun Jiang, Cheng Hu, and Shanshan Chen. Synchronization of complex-valued multi-layer coupled systems by asynchronous intermittent event-triggered mechanisms. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002510.

Roy:2023:RTG

[RJM23] Souvik Roy, Gihyeon Jeon, and Sunghwan Moon. Radon transform with Gaussian beam: Theoretical and numerical reconstruction scheme. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001935.

Revilla:2022:PPD

[RK22] Tomás A. Revilla and Vlastimil Krivan. Prey-predator dynamics with adaptive protection mutualism. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004428.

Rathan:2020:TSI

[RKJ20] Samala Rathan, Rakesh Kumar, and Ameya D. Jagtap. L ^{1}-type smoothness indicators based WENO scheme for nonlinear degenerate parabolic equations. Applied Mathematics and Computation, 375(??): Article 125112, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300813.

Rejwer-Kosinska:2021:ELF

[RKRRL21] Ewa Rejwer-Kosińska, Liliana Rybarska-Rusinek, and Aleksandr Linkov. On evaluation of local fields by fast multipole method employing smooth equivalent/ check surfaces. Applied Mathematics and Computation, 408(??):??, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004525.

Ryu:2020:VDR

[RLK^{+}20] Joonghyun Ryu, Mokwon Lee, Donguk Kim, Josef Kallrath, Kokichi Sugihara, and Deok-Soo Kim. VOROPACK-D: Realtime disk packing algorithm using Voronoi diagram. Applied Mathematics and Computation, 375(??): Article 125076, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030045X.

Ran:2022:DFQ

[RLX22] Jie Ran, Yu-Qin Li, and Yi-Bin Xiong. On the dynamics of fractional qdeformation chaotic map. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001394.

Ren:2020:MAS

[RLY20] Yong Ren, Kun Li, and Hui Ye. Modeling and anti-swing control for a helicopter slung-load system. Applied Mathematics and Computation, 372(??): Article 124990, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309828.

Rybak:2020:DRS

[RM20] Iryna Rybak and Stefan Metzger. A dimensionally reduced Stokes– Darcy model for fluid flow in fractured porous media. Applied Mathematics and Computation, 384 (??):Article 125260, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302290.

Rathinasamy:2023:BSS

[RM23a] Anandaraman Rathinasamy and Pichamuthu Mayavel. The balanced split step theta approximations of stochastic neutral Hopfield neural networks with time delay and Poisson jumps. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002989.

Rathinasamy:2023:SCA

[RM23b] Anandaraman Rathinasamy and Pichamuthu Mayavel. Strong convergence and almost sure exponential stability of balanced numerical approximations to stochastic delay Hopfield neural networks. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006476.

Rodrigues:2020:SMC

[RMCG20] Diego S. Rodrigues, Paulo F. A. Mancera, Tiago Carvalho, and Luiz Fernando Gonçalves. Sliding mode control in a mathematical model to chemoimmunotherapy: the occurrence of typical singularities. Applied Mathematics and Computation, 387(??): Article 124782, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031930774X.

Rahman:2022:CSA

[RMO22] Nizhum Rahman, Robert Marangell, and Dietmar Oelz. Classification and stability analysis of polarising and depolarising travelling wave solutions for a model of collective cell migration. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000406.

Rocha:2022:MPM

[RMSP22] Franciane F. Rocha, Het Mankad, Fabricio S. Sousa, and Felipe Pereira. The multiscale perturbation method for two-phase reservoir flow problems. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009917.

Reddy:2021:ENS

[RNVC21] G. M. M. Reddy, P. Nanda, M. Vynnycky, and J. A. Cuminato. Efficient numerical solution of boundary identification problems: MFS with adaptive stochastic optimization. Applied Mathematics and Computation, 409 (??):Article 126402, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004914.

Rouatbi:2021:CPC

[RO21] Asma Rouatbi and Khaled Omrani. Comments on the paper “A conservative linear difference scheme for the 2D regularized longwave equation”, by Xiaofeng Wang, Weizhong Dai and Shuangbing Guo [Applied Mathematics and Computation, 342 (2019) 55–70]. Applied Mathematics and Computation, 410(??):Article 126455, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005440.

Romani:2021:CBB

[Rom21]  Lucia Romani.  Creat-

ing a bridge between cardinal B r-spline fundamental functions for interpolation and subdivision. Applied Mathematics and Computation, 401 (??):Article 126071, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001193.

Rajasekar:2020:ESD

[RP20] S. P. Rajasekar and M. Pitchaimani.

Ergodic stationary distribution and extinction of a stochastic SIRS epidemic model with logistic growth and nonlinear incidence. Applied Mathematics and Computation, 377 (??):Article 125143, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301120.

Ross:2023:UBP

[RPB23] I. M. Ross, R. J. Proulx, and C. F. Borges. A universal Birkhoff pseudospectral method for solving boundary value problems. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002709.

Rezaifar:2007:NAF

[RR07] Omid Rezaifar and Hossein Rezaee. A new approach for finding the determinant of matrices. Applied Mathematics and Computation, 188(2):1445–1454, May 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300306014974.

See note [Kop22] invalidating claims of linear convergence.

R:2019:MME

[RRR19] Padma R., Ponalagusamy R., and Tamil Selvi R. Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field. Applied Mathematics and Computation, 362(??):Article 124453, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0096300319304175.

See corrigendum [RRR20, PPS24].

R:2020:CMM

[RRR20] Padma R., Ponalagusamy R., and Tamil Selvi R. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Applied Mathematics and Computation, 362 (2019) 124453]. Applied Mathematics and Computation, 373 (??):Article 125031, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310239. See

[RRR19].

Restrepo:2021:ESL

[RRS21] Joel E. Restrepo, Michael Ruzhansky, and Durvudkhan Suragan. Explicit solutions for linear variablecoefficient fractional differential equations with respect to functions. Applied Mathematics and Computation, 403(??):Article 126177, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321002678.

Rada:2023:ISI

[RRS23] Juan Rada, José M. Rodríguez, and José M. Sigarreta. On integral Sombor indices. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002059.

Rada:2024:SIE

[RRS24] Juan Rada, José M. Rodríguez, and José M. Sigarreta. Sombor index and elliptic Sombor index of benzenoid systems. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002261.

Ray:2021:NSS

[RS21] S. Saha Ray and P. Singh. Numerical solution of stochastic Itô–Volterra integral equation by using shifted Jacobi operational matrix method. Applied Mathematics and Computation, 410(??):Article 126440, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005294.

Ramos:2022:SSO

[RS22a] Higinio Ramos and Gurjinder Singh. Solving second order two-point boundary value problems accurately by a third derivative hybrid block integrator. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000467.

Rowlinson:2022:SGW

[RS22b] Peter Rowlinson and Zoran Stanić. Signed graphs whose spectrum is bounded by −2. Applied Mathematics and Computation, 423(??): ??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000777.

RabieiMotlagh:2023:EMS

[RS23] Omid RabieiMotlagh and Leila Soleimani. Effect of mutations on stochastic dynamics of infectious diseases, a probability approach. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001625.

Rossant:2024:SMH

[RS24] Cyrille Rossant and Leila Schneps. Statistical models of hospital patient fatality rates after accidental falls in children. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001504.

Rouz:2024:DWS

[RSA24] O. Farkhondeh Rouz, S. Shahmorad, and D. Ahmadian. Double weakly singular kernels in stochastic Volterra integral equations with application to the rough Heston model. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001929.

Rego:2021:OCP

[RSR21] Leandro Chaves Rêgo, Hugo Victor Silva, and Carlos Diego Rodrigues. Optimizing the cost of preference manipulation in the graph model for conflict resolution. Applied Mathematics and Computation, 392(??): Article 125729, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306822.

Rarita:2021:NSG

[RST21] Luigi Rarità, Ivanka Stamova, and Stefania Tomasiello. Numerical schemes and genetic algorithms for the optimal control of a continuous model of supply chains. Applied Mathematics and Computation, 388 (??):Article 125464, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304240.

Rajchakit:2021:IEC

[RSVL21] G. Rajchakit, R. Sriraman, P. Vignesh, and C. P. Lim. Impulsive effects on Clifford-valued neural networks with time-varying delays: an asymptotic stability analysis. Applied Mathematics and Computation, 407(??):Article 126309, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003982.

Ren:2020:PCO

[RSXH20] Jing Ren, Hao Sun, Genjiu Xu, and Dongshuang Hou. Prediction on the competitive outcome of an enterprise under the adjustment mechanism. Applied Mathematics and Computation, 372(??):Article 124969, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309610.

Rusconi:2023:RMC

[RSZA23] Simone Rusconi, Christina Schenk, Arghir Zarnescu, and Elena Akhmatskaya. Reducing model complexity by means of the optimal scaling: Population balance model for latex particles morphology formation. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008244.

Rieger:2020:BFR

[RT20a] Janosch Rieger and Matthew K. Tam. Backward–forward– reflected–backward splitting for three operator monotone inclusions. Applied Mathematics and Computation, 381(??):Article 125248, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302174.

Rost:2020:SOM

[RT20b] Gergely Röst and Tamás Tekeli. Stability and oscillations of multistage SIS models depend on the number of stages. Applied Mathematics and Computation, 380(??):Article 125259, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302289.

Rukavishnikov:2022:ARE

[RT22] Viktor A. Rukavishnikov and Oleg P. Tkachenko. Approximate resolving equations of mathematical model of a curved thin-walled cylinder. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322000479.

Rawani:2024:NTB

[RVV24] Mukesh Kumar Rawani, Amit Kumar Verma, and Lajja Verma. Numerical treatment of Burgers’ equation based on weakly L-stable generalized time integration formula with the NSFD scheme. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006549.

Ren:2021:LSE

[RW21] Panpan Ren and Jiang-Lun Wu. Least squares estimation for path-distribution dependent stochastic differential equations. Applied Mathematics and Computation, 410(??):Article 126457, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005464.

Roy:2022:TEM

[RWGS22] Nityananda Roy, Karunia Putra Wijaya, Thomas Götz, and S. Sundar. Transport of ellipsoidal microplastic particles in a 3D liddriven cavity under size and aspect ratio variation. Applied Mathematics and Computation, 413 (??):Article 126646, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100730X.

Ren:2020:IPI

[RX20] Zhicong Ren and Wei Xu.

An improved path integration method for nonlinear systems under Poisson white noise excitation. Applied Mathematics and Computation, 373 (??):Article 125036, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300059.

Ro:2024:APB

[RY24] ChungIl Ro and CholHui Yun. Analytical properties and the box-counting dimension of nonlinear hidden variable recurrent fractal interpolation functions constructed by using Rakotch’s fixed point theorem. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400362X.

Ren:2021:VGS

[RYG21] Huan Ren, Yuyue Yan, and Fengqiang Gao. Variable guiding strategies in multiexits evacuation: Pursuing balanced pedestrian densities. Applied Mathematics and Computation, 397 (??):Article 125965, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000138.

Raimundo:2023:MCC

[RYR^{+}23] Silvia Martorano Raimundo, Hyun Mo Yang, Felipe Alves Rubio, David Greenhalgh, and Eduardo Massad. Modeling criminal careers of different levels of offence. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323002424.

Ran:2021:PAP

[RZZ21] Yingli Ran, Ying Zhang, and Zhao Zhang. Parallel approximation for partial set cover. Applied Mathematics and Computation, 408(??):Article 126358, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004471.

Sanjay:2023:EDF

[SAB23] K. Sanjay, R. Vijay Aravind, and P. Balasubramaniam. Extended dissipative filter design for discrete-time interconnected fuzzy systems with timevarying delays subject to cyber attacks. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002400.

Sun:2024:ESS

[SB24] Rongpu Sun and Zhanbing Bai. Existence of solutions to a system of fractional three-point boundary value problem at resonance. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000481.

Semisalov:2022:VSS

[SBB^{+}22] Boris Semisalov, Vasily Belyaev, Luka Bryndin, Arsenii Gorynin, Alexander Blokhin, Sergey Golushko, and Vasily Shapeev. Verified simulation of the stationary polymer fluid flows in the channel with elliptical cross-section. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200368X.

Scoz:2022:WPM

[SBT22] Alessia Scoz, Laura Bertazzi, and Eleuterio F. Toro. On well-posedness of a mathematical model for cerebrospinal fluid in the optic nerve sheath and the spinal subarachnoid space. Applied Mathematics and Computation, 413 (??):Article 126625, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007098.

Stiadle:2023:CES

[SBV23] Thomas I. Stiadle, Alvin Bayliss, and Vladimir A. Volpert. Cyclic ecological systems with an exceptional species. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008074.

Song:2020:NTN

[SC20a] Chuan-Jing Song and Yao Cheng. Noether’s theorems for nonshifted dynamic systems on time scales. Applied Mathematics and Computation, 374 (??):Article 125086, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300552.

Song:2020:CDN

[SC20b] Jia-Sheng Song and Xiao-Heng Chang. H_{∞} controller design of networked control systems with a new quantization structure. Applied Mathematics and Computation, 376(??): Article 125070, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300394.

Szolnoki:2020:BDI

[SC20c] Attila Szolnoki and Xiaojie Chen. Blocking defector invasion by focusing on the most successful partner. Applied Mathematics and Computation, 385 (??):Article 125430, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030391X.

Solorza-Calderon:2021:TWI

[SC21] Selene Solorza-Calderón. Torsional waves of infinite fully saturated poroelastic cylinders within the framework of Biot viscosityextended theory. Applied Mathematics and Computation, 391(??):Article 125636, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320305907.

Singh:2024:CPE

[SCC24] Saroja Kumar Singh, Gabriel M. B. Cruz, and Frederico R. B. Cruz. Change point estimation of service rate in M/M/1/m queues: a Bayesian approach. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005921.

Sajan:2023:NAA

[SCD23] Sajan, Kapil Kumar Choudhary, and Balram Dubey. A non-autonomous approach to study the impact of environmental toxins on nutrient–plankton system.

Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004058.

Sun:2020:SDT

[SCG^{+}20] Bo Sun, Yuting Cao, Zhenyuan Guo, Zheng Yan, and Shiping Wen. Synchronization of discrete-time recurrent neural networks with time-varying delays via quantized sliding mode control. Applied Mathematics and Computation, 375(??): Article 125093, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030062X.

Schinazi:2022:EPU

[Sch22] Rinaldo B. Schinazi. Evolutionary paths under catastrophes. Applied Mathematics and Computation, 418 (??):Article 126808, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008900.

Sun:2022:AFT

[SCHS22] Haibin Sun, Yahui Cui, Linlin Hou, and Kaibo Shi.

Adaptive finite-time control for cyber-physical systems with injection and deception attacks. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003903.

Stanimirovic:2021:RSC

[SĆL^{+}21] Predrag S. Stanimirović, Miroslav Ćirić, Alberto Lastra, Juan Rafael Sendra, and Juana Sendra. Representations and symbolic computation of generalized inverses over fields. Applied Mathematics and Computation, 406(??):Article 126287, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003763.

Sader:2021:DRF

[SCLD21] Malika Sader, Zengqiang Chen, Zhongxin Liu, and Chao Deng. Distributed robust fault-tolerant consensus control for a class of nonlinear multi-agent systems with intermittent communications. Applied Mathematics and Computation, 403(??):Article 126166, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002563.

Solntsev:2023:CRM

[SCMK23] Igor A. Solntsev, Abhishek Chintagunta, Annabel P. Markesteijn, and Sergey A. Karabasov. CABARET on rotating meshes. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000401.

Song:2020:CNP

[SCT^{+}20] Qun Song, Zhaoheng Cao, Rui Tao, Wei Jiang, Chen Liu, and Jinzhuo Liu. Conditional neutral punishment promotes cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 368(??): ??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307908.

Shirilord:2022:SSI

[SD22a] Akbar Shirilord and Mehdi Dehghan. Single step iterative method for linear system of equations with complex symmetric positive semi-definite coefficient matrices. Applied Mathematics and Computation, 426(??): ??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001953.

Sun:2022:SIT

[SD22b] Xiaoling Sun and Jianwei Du. On Sombor index of trees with fixed domination number. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000327.

Shi:2023:CMP

[SD23] Lingjuan Shi and Kai Deng. Counting the maximal and perfect matchings in benzenoid chains. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000917.

Shahu:2022:CDW

[SDD22] Chiranjeev K. Shahu, Sharad Dwivedi, and Shruti Dubey. Curved domain walls in the ferromagnetic nanostructures with Rashba and nonlinear dissipative effects. Applied Mathematics and Computation, 420(??): Article 126894, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009772.

Sun:2022:RFT

[SDW^{+}22] Shaoxin Sun, Xin Dai, Zhiliang Wang, Yu Zhou, and Xiangpeng Xie. Robust finite-time H_{∞} control for Itô stochastic semi-Markovian jump systems with delays. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002557.

Sekiguchi:2023:AST

[Sek23] Takuya Sekiguchi. Abundance of strategies for trimatrix games in finite populations. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300111X.

Sun:2023:LCS

[SFC23] Linlin Sun, Zhuojia Fu, and Zhikang Chen. A localized collocation solver based on fundamental solutions for 3D time harmonic elastic wave propagation analysis. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006737.

Sun:2022:FAT

[SFL^{+}22] Lu-Yao Sun, Zhi-Wei Fang, Siu-Long Lei, Hai-Wei Sun, and Jia-Li Zhang. A fast algorithm for twodimensional distributedorder time-space fractional diffusion equations. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001795.

Serrano:2020:CMS

[SG20a] Daniel Hernández Serrano and Darío Sánchez Gómez. Centrality measures in simplicial complexes: Applications of topological data analysis to network science. Applied Mathematics and Computation, 382 (??):Article 125331, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302976.

Spiteri:2020:EPN

[SG20b] Raymond J. Spiteri and Wenxian Guo. Efficient partitioned numerical integrators for myocardial cell models. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307301.

Singh:2021:NEF

[SG21] Brajesh Kumar Singh and Mukesh Gupta. A new efficient fourth order collocation scheme for solving Burgers’ equation. Applied Mathematics and Computation, 399(??):Article 126011, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100059X.

Shankar:2024:RSB

[SG24] Ravi Shankar and Mayank Goel. Risk-sensitive benchmarked portfolio optimization under non-linear market dynamics. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003874.

Shargatov:2020:SFP

[SGI20] Vladimir A. Shargatov, Sergey V. Gorkunov, and Andrej T. Il’ichev. Stability of finite perturbations of the phase transition interface for one problem of water evaporation in a porous medium. Applied Mathematics and Computation, 378(??):Article 125208, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301776.

Song:2021:TPI

[SGJ^{+}21] Zhao Song, Hao Guo, Danyang Jia, Matjaž Perc, Xuelong Li, and Zhen Wang. Third party interventions mitigate conflicts on interdependent networks. Applied Mathematics and Computation, 403 (??):Article 126178, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100268X.

Song:2021:SEA

[SGL21] Minghui Song, Yidan Geng, and Mingzhu Liu. Stability equivalence among stochastic differential equations and stochastic differential equations with piecewise continuous arguments and corresponding Euler–Maruyama methods. Applied Mathematics and Computation, 400(??): Article 125813, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307669.

Shi:2022:RSP

[SGPZ22] Baoli Shi, Fang Gu, Zhi-Feng Pang, and Yuhua Zeng. Remove the salt and pepper noise based on the high order total variation and the nuclear norm regularization. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200011X.

Shi:2021:CAF

[SGX21] Lin Shi, Kuixiang Gou, and Dongmei Xie. Convergence analysis of firstorder discrete multi-agent systems with cooperativecompetitive mechanisms. Applied Mathematics and Computation, 410(??):Article 126462, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321005518.

Samir:2021:CDO

[SH21] Chafik Samir and Wen Huang. Coordinate descent optimization for oneto-one correspondence and supervised classification of 3D shapes. Applied Mathematics and Computation, 388(??):Article 125539, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304951.

Shao:2022:ESM

[SH22] Qi Shao and Dun Han. Epidemic spreading in metapopulation networks with heterogeneous mobility rates. Applied Mathematics and Computation, 412(??):Article 126559, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006433.

Sharma:2021:CPS

[Sha21] M. D. Sharma. Correct procedure to study reflection in orthotropic thermoelastic medium: Inhomogeneous propagation of waves. Applied Mathematics and Computation, 391 (??):Article 125692, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306457.

Shang:2022:SID

[Sha22] Yilun Shang. Sombor index and degree-related properties of simplicial networks. Applied Mathematics and Computation, 419(??): Article 126881, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009644.

Shen:2020:USS

[She20] Jiayu Shen. An uncertain sustainable supply chain network. Applied Mathematics and Computation, 378(??):Article 125213, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030182X.

Shi:2021:CSF

[Shi21] Chunmei Shi. The convergence and stability of full discretization scheme for stochastic age-structured population models. Applied Mathematics and Computation, 396(??):Article 125867, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308201.

Shi:2021:IBP

[SHT^{+}21] Juan Shi, Die Hu, Rui Tao, Yunchen Peng, Yong Li, and Jinzhuo Liu. Interaction between populations promotes cooperation in voluntary prisoner’s dilemma. Applied Mathematics and Computation, 392(??):Article 125728, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306810.

Song:2021:EGO

[SHW21] Rui Song, Qiongxiang Huang, and Peng Wang. The extremal graphs of order trees and their topological indices. Applied Mathematics and Computation, 398(??):Article 125988, June 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321000369.

Shi:2021:UIM

[SHWW21] Qinghe Shi, Kejun Hu, Lei Wang, and Xiaojun Wang. Uncertain identification method of structural damage for beam-like structures based on strain modes with noises. Applied Mathematics and Computation, 390(??):Article 125682, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306354.

Sun:2023:SFO

[SHYS23] Yuting Sun, Cheng Hu, Juan Yu, and Tingting Shi. Synchronization of fractional-order reactiondiffusion neural networks via mixed boundary control. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001510.

Salahuddin:2022:AEG

[SIK22] T. Salahuddin, Ikram Imtiaz, and Mair Khan. Analysis of entropy generation in AA7072-methanol and AA7072 + AA7075methanol flow near a parabolic surface. Applied Mathematics and Computation, 413(??):Article 126616, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007001.

Singer:2020:AFS

[Sin20a] Sanja Singer. The antitriangular factorization of skew-symmetric matrices. Applied Mathematics and Computation, 381(??): Article 125263, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302320.

Singh:2020:CIN

[Sin20b] Vipin Kumar Singh. On the convergence of inexact Newton-like methods under mild differentiability conditions. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930863X.

Song:2020:EFO

[SJ20] Zhao Song and Jae-Hun Jung. The exact formula of the optimal penalty parameter value of the spectral penalty method for differential equations. Applied Mathematics and Computation, 381(??):Article 125313, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302794.

Shanmugam:2023:ANN

[SJ23a] Lakshmanan Shanmugam and Young Hoon Joo. Adaptive neural networks-based integral sliding mode control for T–S fuzzy model of delayed nonlinear systems. Applied Mathematics and Computation, 450(??): ??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001522.

Sun:2023:RLE

[SJ23b] Zhongshi Sun and Guangyan Jia. Reinforcement learning for exploratory linearquadratic two-person zerosum stochastic differential games. Applied Mathematics and Computation, 442 (??):??, April 1, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322008311.

Shin:2022:NHD

[SJP22] Dongwook Shin, Youngmok Jeon, and Eun-Jae Park. A novel hybrid difference method for an elliptic equation. Applied Mathematics and Computation, 415 (??):Article 126702, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007864.

Silva-Juarez:2021:OKY

[SJTCdlFL21] Alejandro Silva-Juárez, Esteban Tlelo-Cuautle, Luis Gerardo de la Fraga, and Rui Li. Optimization of the Kaplan–Yorke dimension in fractional-order chaotic oscillators by metaheuristics. Applied Mathematics and Computation, 394 (??):Article 125831, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307840.

Shi:2022:SSR

[SJZA22] Zhenfeng Shi, Daqing Jiang, Xinhong Zhang, and Ahmed Alsaedi. A stochastic SEIRS rabies model with population dispersal: Stationary distribution and probability density function. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002636.

Shokri:2020:NFE

[SK20] Ali Shokri and Mohammad Mehdizadeh Khalsaraei. A new family of explicit linear two-step singularly P -stable Obrechkoff methods for the numerical solution of second-order IVPs. Applied Mathematics and Computation, 376(??): Article 125116, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300850.

Selvan:2024:SDD

[SK24a] T. Tamil Selvan and M. Kumar. Stationary distribution of a double epidemic stochastic model driven by saturated incidence rates. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001693.

Shuvo:2024:ECD

[SK24b] Md. Fahimur Rahman Shuvo and K. M. Ariful Kabir. Exploring cyclic dominance in rock-paper-scissor strategies within the Optional Prisoner’s Dilemma: a quantitative analysis with environmental feedback. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002972.

Suriguga:2020:USD

[SKH20] Ma Suriguga, Yonggui Kao, and Abd-Allah Hyder. Uniform stability of delayed impulsive reaction-diffusion systems. Applied Mathematics and Computation, 372 (??):Article 124954, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309464.

Stanimirovic:2021:PCC

[SKJM21] Predrag S. Stanimirović, Vasilios N. Katsikis, Long Jin, and Dijana Mosić. Properties and computation of continuous-time solutions to linear systems. Applied Mathematics and Computation, 405(??): Article 126242, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003325.

Song:2021:BNL

[SKK^{+}21] Ying-Qing Song, Sami Ullah Khan, M Ijaz Khan, Muhammad Awais, Aamar Abbasi, and Qiu-Hong Shi. Bidirectional non-linear stretched flow of Williamson nanofluid with swimming of motile gyrotactic microorganisms. Applied Mathematics and Computation, 411(??):Article 126502, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005919.

Siddheshwar:2023:WNS

[SKL23] P. G. Siddheshwar, C. Kanchana, and D. Laroze. Weakly nonlinear stability analysis and study of chaotic Darcy–Benard convection of a combusting fluid. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200889X.

Selvaraj:2022:DRI

[SKLS22] P. Selvaraj, O. M. Kwon, S. H. Lee, and R. Sakthivel. Disturbance rejections of interval type-2 fuzzy systems under event-triggered control scheme. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003976.

Stanislawski:2022:IFO

[SKR22] Rafał Stanisławski, Kamil Kozioł, and Marek Rydel. Implementation of fractional-order difference via Takenaka–Malmquist functions. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005264.

Shankar:2021:NIS

[SKS21] B. M. Shankar, Jai Kumar, and I. S. Shivakumara. Numerical investigation of the stability of mixed convection in a differentially heated vertical porous slab. Applied Mathematics and Computation, 389(??):Article 125486, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304458.

Suriguga:2021:SHO

[SKSC21] Suriguga, Yonggui Kao, Chuntao Shao, and Xiangyong Chen. Stability of highorder delayed Markovian jumping reaction-diffusion HNNs with uncertain transition rates. Applied Mathematics and Computation, 389(??):Article 125559, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305154.

Su:2021:MAR

[SKWW21] Fuyong Su, Linglu Kong, Hui Wang, and Zhi Wen. Modeling and application for rolling scheduling problem based on TSP. Applied Mathematics and Computation, 407 (??):Article 126333, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004227.

Saray:2020:SMS

[SL20a] Behzad Nemati Saray and Mehrdad Lakestani. On the sparse multi-scale solution of the delay differential equations by an efficient algorithm. Applied Mathematics and Computation, 381(??):Article 125291, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302605.

Shi:2020:SAT

[SL20b] Dongyang Shi and Qian Liu. Superconvergence analysis of a two grid finite element method for Ginzburg– Landau equation. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306836.

Shi:2020:NTG

[SL20c] Xiangyu Shi and Linzhang Lu. A new two-grid nonconforming mixed finite element method for nonlinear Benjamin–Bona– Mahoney equation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930935X.

Sun:2020:CEI

[SL20d] Chengbin Sun and Chao Luo. Co-evolution of influence-based preferential selection and limited resource with multi-games on interdependent networks. Applied Mathematics and Computation, 374 (??):Article 125063, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300321.

Sun:2024:NAS

[SL24] Hongli Sun and Yanfei Lu. A novel approach for solving linear Fredholm integro-differential equations via LS-SVM algorithm. Applied Mathematics and Computation, 470(??): ??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000298.

Shapoval:2020:IVL

[SLCS20] A. Shapoval, J.-L. Le Mouël, V. Courtillot, and M. Shnirman. Influence of very large spatial heterogeneity on estimates of sea-level trends. Applied Mathematics and Computation, 386(??):Article 125485, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304446.

Song:2020:ZEI

[SLH20] Xiaodi Song, Jianping Li, and Weihua He. On Zagreb eccentricity indices of cacti. Applied Mathematics and Computation, 383 (??):Article 125361, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303258.

Song:2021:CMI

[SLHX21] Zhiwei Song, Wei Li, Xiaoqiao He, and De Xie. Comparisons of matched interface and boundary (MIB) method and its interpolation formulation for free vibration analysis of stepped beams and plates. Applied Mathematics and Computation, 394(??):Article 125817, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307700.

Shen:2023:LEU

[SLJ23] Shouqiang Shen, Weijun Liu, and Wei Jin. Laplacian eigenvalues of the unit graph of the ring Z_{n}. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300437X.

Sun:2023:CEP

[SLK^{+}23] Xingping Sun, Mingyuan Li, Hongwei Kang, Yong Shen, and Qingyi Chen. Combined effect of pure punishment and reward in the public goods game. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300022X.

Sun:2022:ESP

[SLL^{+}22] Daoqiang Sun, Long Li, Kai Liu, Hua Wang, and Yu Yang. Enumeration of subtrees of planar twotree networks. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004787.

Shi:2024:HCC

[SLL24] Yaxin Shi, Fengxia Liu, and HongJian Lai. On rhued coloring of connected P_{5}-free tripartite graphs. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003229.

Sun:2022:MDD

[SLPH22] Shaoxin Sun, Ting Li, Yongheng Pang, and Xingxing Hua. Multiple delaydependent event-triggered finite-time H_{∞} filtering for uncertain networked random systems against state and input constraints. Applied Mathematics and Computation, 415(??):Article 126711, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007955.

Sibatov:2024:VOF

[SLS24] Renat T. Sibatov, Pavel E. L’vov, and HongGuang Sun. Variable-order fractional diffusion: Physical interpretation and simulation within the multiple trapping model. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004211.

Su:2020:QTS

[SLTL20] Guangwang Su, Liang Lu, Bo Tang, and Zhenhai Liu. Quasilinearization technique for solving nonlinear Riemann– Liouville fractional-order problems. Applied Mathematics and Computation, 378(??):Article 125199, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301685.

Shu:2020:LPU

[SLZ20a] Yu Shu, Jian Li, and Chong Zhang. A local and parallel Uzawa finite element method for the generalized Navier– Stokes equations. Applied Mathematics and Computation, 387(??):Article 124671, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306630.

Song:2020:GFT

[SLZ^{+}20b] Zhibao Song, Ping Li, Junyong Zhai, Zhen Wang, and Xia Huang. Global fixed-time stabilization for switched stochastic nonlinear systems under rational switching powers. Applied Mathematics and Computation, 387(??):Article 124856, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308483.

Sun:2020:LDG

[SLZ20c] Xiaorui Sun, Can Li, and Fengqun Zhao. Local discontinuous Galerkin methods for the time tempered fractional diffusion equation. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307179.

Sabir:2020:SCG

[SM20] Eminjan Sabir and Jixiang Meng. Sufficient conditions for graphs to be spanning connected. Applied Mathematics and Computation, 378(??):Article 125198, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301673.

Sabir:2021:GCS

[SM21a] Eminjan Sabir and Jixiang Meng. Generalizations of the classics to spanning connectedness. Applied Mathematics and Computation, 403(??):Article 126167, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002575.

Saffarian:2021:NST

[SM21b] Marziyeh Saffarian and Akbar Mohebbi. Numerical solution of two and three dimensional time fractional damped nonlinear Klein–Gordon equation using ADI spectral element method. Applied Mathematics and Computation, 405(??):Article 126182, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002721.

Sun:2021:ORS

[SM21c] Yuchen Sun and Shuping Ma. Output regulation of switched singular systems based on extended state observer approach. Applied Mathematics and Computation, 399(??):Article 126020, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000680.

Song:2022:ETK

[SM22] Le Song and Yinghong Ma.

Evaluating tacit knowledge diffusion with algebra matrix algorithm based social networks. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002090.

Sun:2023:AOR

[SM23] Yuchen Sun and Shuping Ma. Asynchronous output regulation for a class of linear descriptor switched systems via an error-feedback controller. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001601.

Shi:2024:NMS

[SM24] Chao Shi and Tianlong Ma. A note on maximum size of a graph without isolated vertices under the given matching number. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004642.

Shi:2020:NHS

[SMH^{+}20] Shang Shi, Huifang Min, Yinlong Hu, Yonghui Sun, and Bing Wang. A novel hybrid scheme for fixedtime SOSM control of nonlinear uncertain systems subject to mismatched terms. Applied Mathematics and Computation, 386 (??):Article 125511, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304690.

Stanimirovic:2024:ZNN

[SMM^{+}24] Predrag S. Stanimirović, Spyridon D. Mourtas, Dijana Mosić, Vasilios N. Katsikis, Xinwei Cao, and Shuai Li. Zeroing neural network approaches for computing time-varying minimal rank outer inverse. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005817.

Sabir:2024:SLB

[SMQ24] Eminjan Sabir, Jixiang Meng, and Hongwei Qiao. On spanning laceability of bipartite graphs. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003801.

Sutthimat:2022:CFF

[SMR22] Phiraphat Sutthimat, Khamron Mekchay, and Sanae Rujivan. Closed-form formula for conditional moments of generalized nonlinear drift CEV process. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002879.

Singh:2023:HFS

[SMTT23] Satyvir Singh, R. C. Mittal, Shafeeq Rahman Thottoli, and Mohammad Tamsir. High-fidelity simulations for Turing pattern formation in multidimensional Gray–Scott reaction-diffusion system. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002485.

Stanimirovic:2022:GCI

[SMW22] Predrag S. Stanimirović, Dijana Mosić, and Yimin Wei. Generalizations of composite inverses with certain image and/or kernel. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002338.

Shao:2023:CDI

[SMW^{+}23] Yuanzhen Shao, Mark Mc-Gowan, Siwen Wang, Emil Alexov, and Shan Zhao. Convergence of a diffuse interface Poisson–Boltzmann (PB) model to the sharp interface PB model: a unified regularization formulation. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005756.

Sahoo:2021:EGC

[SN21] A. Sahoo and R. Nandkeolyar. Entropy generation in convective radiative flow of a Casson nanofluid in non-Darcy porous medium with Hall current and activation energy: the multiple regression model. Applied Mathematics and Computation, 402(??):Article 125923, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320308766.

Song:2022:LIC

[SNC22] Shu Jiao Song, Xingtian Ning, and Panpan Cheng. Locating and identifying codes in dihedral graphs. Applied Mathematics and Computation, 416(??):Article 126752, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008341.

Shekhar:2021:CTS

[SNMM21] Vaibhav Shekhar, Snigdhashree Nayak, Nachiketa Mishra, and Debasisha Mishra. Convergence of two-stage iterative scheme for K-weak regular splittings of type II. Applied Mathematics and Computation, 410 (??):Article 126471, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005609.

Shahriari:2022:CPA

[SNR^{+}22] Zahra Shahriari, Fahimeh Nazarimehr, Karthikeyan Rajagopal, Sajad Jafari, Matjaž Perc, and Milan Svetec. Cryptocurrency price analysis with ordinal partition networks. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003113.

Shaheen:2020:RGO

[SNRA20] Bilal Shaheen, Muhammad Shahid Nazir, Muhammad Rehan, and Sohaira Ahmad. Robust generalized observer design for uncertain one-sided Lipschitz systems. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319305806.

Schweiger:2020:MSL

[SNS^{+}20] G. Schweiger, H. Nilsson, J. Schoeggl, W. Birk, and A. Posch. Modeling and simulation of large-scale systems: a systematic comparison of modeling paradigms. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307052.

Siddheshwar:2020:RBC

[SNZC20] P. G. Siddheshwar, Shivakumar B. N., Yi Zhao, and Kanchana C. Rayleigh– Bénard convection in a Newtonian liquid bounded by rigid isothermal boundaries. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309348.

Sabermahani:2021:GLH

[SO21a] Sedigheh Sabermahani and Yadollah Ordokhani. General Lagrange-hybrid functions and numerical solution of differential equations containing piecewise constant delays with bibliometric analysis. Applied Mathematics and Computation, 395(??):Article 125847, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308006.

Salvador:2021:CTV

[SO21b] Beatriz Salvador and Cornelis W. Oosterlee. Corrigendum to “Total value adjustment for a stochastic volatility model. A comparison with the Black–Scholes model”. Applied Mathematics and Computation, 406(??):Article 125999, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000473.

See [SO21c].

Salvador:2021:TVA

[SO21c] Beatriz Salvador and Cornelis W. Oosterlee. Total value adjustment for a stochastic volatility model. A comparison with the Black–Scholes model. Applied Mathematics and Computation, 391(??):Article 125489, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304483. See

corrigendum [SO21b].

Song:2024:QST

[Son24] Xingkun Song. Quantum state transfer on integral oriented circulant graphs. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300560X.

Szolnoki:2023:OSC

[SP23] Attila Szolnoki and Matjaz Perc. Oppressed species can form a winning pair in a multi-species ecosystem. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006427.

Spiro:2022:WIS

[Spi22] Sam Spiro. The Wiener index of signed graphs. Applied Mathematics and Computation, 416(??):Article 126755, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008377.

Stanimirovic:2022:ESC

[SPM22] Predrag S. Stanimirović, Marko D. Petković, and Dijana Mosić. Exact solutions and convergence of gradient based dynamical systems for computing outer inverses. Applied Mathematics and Computation, 412(??):Article 126588, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100672X.

Sardar:2020:CRD

[SPX20] Muhammad Shoaib Sardar, Xiang-Feng Pan, and Si-Ao Xu. Computation of resistance distance and Kirchhoff index of the two classes of silicate networks. Applied Mathematics and Computation, 381(??): Article 125283, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302526.

Sardar:2024:CRD

[SPX24] Muhammad Shoaib Sardar, Xiang-Feng Pan, and Shou-Jun Xu. Computation of the resistance distance and the Kirchhoff index for the two types of claw-free cubic graphs. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001425.

Song:2020:AHF

[SPZS20] Shuai Song, Ju H. Park, Baoyong Zhang, and Xiaona Song. Adaptive hybrid fuzzy output feedback control for fractional-order nonlinear systems with time-varying delays and input saturation. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930654X.

Sharma:2023:AHO

[SR23] Amit Sharma and Pratima Rai. Analysis of a higher order uniformly convergent method for singularly perturbed parabolic delay problems. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000759.

Sayana:2024:OPC

[SR24] K. J. Sayana and G. D. Reddy. Optimal parameter choice rule for filter-based regularization schemes. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004090.

Stanimirovic:2020:CMP

[SRGS20] Predrag S. Stanimirović, Falguni Roy, Dharmendra K. Gupta, and Shwetabh Srivastava. Computing the Moore–Penrose inverse using its error bounds. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930949X.

Srivastava:2020:NAL

[SRK^{+}20] H. M. Srivastava, Mumtaz Riyasat, Subuhi Khan, Serkan Araci, and Mehmet Acikgoz. A new approach to Legendre-truncated-exponential-based Sheffer sequences via Riordan arrays. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306757.

Sharon:2021:THN

[SRK^{+}21] Jane Olive Sharon, T. M.

Rajalaxmi, Sandi Klavzar, R. Sundara Rajan, and Indra Rajasingh. Transmission in H-naphtalenic nanosheet. Applied Mathematics and Computation, 406(??):Article 126285, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100374X.

Saha:2023:LSQ

[SRK23] Tanay Saha, Suman Rakshit, and Swanand R. Khare. Linearly structured quadratic model updating using partial incomplete eigendata. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000607.

Sriram:2023:MFE

[SRKA23] Sridevi Sriram, Karthikeyan Rajagopal, Anitha Karthikeyan, and Akif Akgul. Memristive field effect in a single and multilayer neural network with different connection topologies. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003405.

Smedt:2023:NAB

[SRN^{+}23] Kris Smedt, Daniel Ruprecht, Jitse Niesen, Steven Tobias, and Joonas Nättilä. New applications for the Boris Spectral Deferred Correction algorithm for plasma simulations. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007743.

Sun:2022:SMC

[SRZ22] Qingdong Sun, Junchao Ren, and Feng Zhao. Sliding mode control of discrete-time interval type-2 fuzzy Markov jump systems with the preview target signal. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005537.

Satin:2022:UBR

[SRZK22] Y. A. Satin, R. V. Razumchik, A. I. Zeifman, and I. A. Kovalev. Upper bound on the rate of convergence and truncation bound for non-homogeneous birth and death processes on Z. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000959.

Sasu:2020:ACN

[SS20] Adina Luminita Sasu and Bogdan Sasu. Admissibility criteria for nonuniform dichotomic behavior of nonautonomous systems on the whole line. Applied Mathematics and Computation, 378(??):Article 125167, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301363.

Sedlar:2021:BMD

[SS21a] Jelena Sedlar and Riste Skrekovski. Bounds on metric dimensions of graphs with edge disjoint cycles. Applied Mathematics and Computation, 396(??): Article 125908, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308614.

Sedlar:2021:EMM

[SS21b] Jelena Sedlar and Riste Skrekovski. Extremal mixed metric dimension with respect to the cyclomatic number. Applied Mathematics and Computation, 404(??):Article 126238, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003283.

Staub:2021:ERL

[SS21c] Ruben Staub and Stephan N.

Steinmann. Efficient recursive least squares solver for rank-deficient matrices. Applied Mathematics and Computation, 399(??): Article 125996, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000448.

Savic:2022:SPB

[SS22a] Marko Savić and Milos Stojaković. Structural properties of bichromatic non-crossing matchings. Applied Mathematics and Computation, 415 (??):Article 126695, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007797.

Sedlar:2022:MDV

[SS22b] Jelena Sedlar and Riste Skrekovski. Metric dimensions vs. cyclomatic number of graphs with minimum degree at least two. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002272.

Shukla:2022:TMS

[SS22c] Vivekanand Shukla and Jeeoot Singh. Thermomechanical stability analysis of angle-ply plates using meshless method. Applied Mathematics and Computation, 413(??):Article 126644, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007281.

Schrocker:2023:OIS

[SS23a] Hans-Peter Schröcker and Zbynek Sír. Optimal interpolation with spatial rational Pythagorean hodograph curves. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003831.

Sil:2023:NPS

[SS23b] Subhankar Sil and T. Raja Sekhar. Nonclassical potential symmetry analysis and exact solutions for a thin film model of a perfectly soluble anti-surfactant solution. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007317.

Singh:2023:RCM

[SS23c] Uaday Singh and Birendra Singh. On rate of convergence of matrix means of corrected Fourier series. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003855.

Sooknanan:2023:CTM

[SS23d] Joanna Sooknanan and Terence A. R. Seemungal. Criminals and their models — a review of epidemiological models describing criminal behaviour. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003818.

Schrocker:2024:TPR

[SS24a] Hans-Peter Schröcker and Zbynek Sír. Three paths to rational curves with rational arc length. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003035.

Sedlar:2024:NEC

[SS24b] Jelena Sedlar and Riste Skrekovski. Normal 5-edge-coloring of some snarks superpositioned by the Petersen graph. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006628.

Seiler:2024:NIS

[SS24c] Werner M. Seiler and Matthias Seiß. On the numerical integration of singular initial and boundary value problems for generalised Lane–Emden and Thomas–Fermi equations. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300615X.

Shahab:2024:NNB

[SS24d] Muhammad Luthfi Shahab and Hadi Susanto. Neural networks for bifurcation and linear stability analysis of steady states in partial differential equations. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004466.

Suo:2024:RMS

[SS24e] Hao Hai Suo and Shu Jiao Song. Rotary maps on symmetric groups of prime degree. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000055.

Samadi:2023:RCD

[SSA^{+}23] B. Samadi, N. Soltankhah, H. Abdollahzadeh Ahangar, M. Chellali, D. A. Mojdeh, S. M. Sheikholeslami, and J. C. Valenzuela-Tripodoro. Restrained condition on double Roman dominating functions. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006282.

Saha:2022:RTW

[SSC22] Shalini Saha, Abhishek Kumar Singh, and Amares Chattopadhyay. Rayleightype wave propagation in exponentially graded initially stressed composite structure resting on rigid and yielding foundations. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004957.

Singh:2020:MSR

[SSCG20] Pooja Singh, Abhishek Kumar Singh, Amares Chattopadhyay, and Sayantan Guha. Mathematical study on the reflection and refraction phenomena of threedimensional plane waves in a structure with floating frozen layer. Applied Mathematics and Computation, 386(??):Article 125488, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304471.

Shen:2022:SPG

[SSCW22] Zhao-Li Shen, Meng Su, Bruno Carpentieri, and Chun Wen. Shifted power-GMRES method accelerated by extrapolation for solving PageRank with multiple damping factors. Applied Mathematics and Computation, 420(??): Article 126799, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100881X.

Shi:2024:EAF

[SSF24] Kaiwen Shi, Haiyan Su, and Xinlong Feng. Error analysis of a fully discrete PFEM for the 2D/ 3D unsteady incompressible MHD equations. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003205.

Sakai:2023:GCH

[SSI23] Hiroyuki Sakai, Hiroyuki Sato, and Hideaki Iiduka. Global convergence of Hager– Zhang type Riemannian conjugate gradient method. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007536.

Shah:2022:SWT

[SSJ22] Sarswati Shah, Randheer Singh, and Jasobanta Jena. Steepened wave in twophase Chaplygin flows comprising a source term. Applied Mathematics and Computation, 413(??):Article 126656, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007402.

Sakthivel:2021:DRS

[SSKS21] Ramalingam Sakthivel, Rathinasamy Sakthivel, Oh-Min Kwon, and Palanisamy Selvaraj. Disturbance rejection for singular semi-Markov jump neural networks with input saturation. Applied Mathematics and Computation, 407 (??):Article 126301, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003908.

Sajjad:2024:CRD

[SSP24] Wasim Sajjad, Muhammad Shoaib Sardar, and Xiang-Feng Pan. Computation of resistance distance and Kirchhoff index of chain of triangular bipyramid hexahedron. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004824.

Suarez-Serrato:2023:CPA

[SSR23] Pablo Suárez-Serrato and E. I. Velázquez Richards. Contour parametrization via anisotropic mean curvature flow. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322007676.

Sun:2021:NFD

[SSW21] Lin Sun, Lei Su, and Jing Wang. Non-fragile dissipative state estimation for semi-Markov jump inertial neural networks with reaction–diffusion. Applied Mathematics and Computation, 411(??):Article 126404, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321004938.

Sun:2022:STS

[SSW22] Ping Sun, Rui Shan, and Shuoyu Wang. Safetytriggered stochastic tracking control for a cushion robot by constraining velocity considering the estimated internal disturbance. Applied Mathematics and Computation, 416(??): Article 126761, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008432.

Sun:2022:PSP

[ST22] Yuangong Sun and Yazhou Tian. Polynomial stability of positive switching homogeneous systems with different degrees. Applied Mathematics and Computation, 414(??):Article 126699, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007839.

Salva:2024:RBT

[ST24] N. N. Salva and D. A. Tarzia. Relationship between two solidification problems in order to determine unknown thermal coefficients when the heat transfer coefficient is very large. Applied Mathematics and Computation, 468(??): ??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006938.

Stanic:2020:SSG

[Sta20]  Zoran Stanić.  Spectra

of signed graphs with two eigenvalues. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306198.

Slynko:2024:ESN

[STA24] Vitalii Slynko, Osman Tunç, and Ivan Atamas. Exponential stability in the L^{p}-norm of nonlinear coupled hyperbolic spatially inhomogeneous systems. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001048.

Syuhada:2024:CPL

[STH24] Khreshna Syuhada, Venansius Tjahjono, and Arief Hakim. Compound Poisson–Lindley process with Sarmanov dependence structure and its application for premium-based spectral risk forecasting. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006616.

Sugie:2021:NPP

[Sug21] Jitsuro Sugie. Number of positive periodic solutions for first-order nonlinear difference equations with feedback. Applied Mathematics and Computation, 391 (??):Article 125626, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305804.

Sun:2022:CST

[Sun22]  Zhengjie Sun.  A con-

servative scheme for twodimensional Schrödinger equation based on multiquadric trigonometric quasiinterpolation approach. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000820.

Shi:2024:CQC

[SUN24] Lei Shi, Malik Zaka Ullah, and Hemant Kumar Nashine. On the construction of a quartically convergent method for highdimensional Black–Scholes time-dependent PDE. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005490.

Sokhal:2021:FWS

[SV21a] Simran Sokhal and Sag Ram Verma. A Fourier wavelet series solution of partial differential equation through the separation of variables method. Applied Mathematics and Computation, 388(??):Article 125480, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304392.

Sommariva:2021:RMC

[SV21b] A. Sommariva and M. Vianello.

RBF moment computation and meshless cubature on general polygonal regions. Applied Mathematics and Computation, 409 (??):Article 126375, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004641.

Shukla:2023:HCQ

[SV23a] Alok Shukla and Prakash Vedula. A hybrid classicalquantum algorithm for solution of nonlinear ordinary differential equations. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007767.

Strazzullo:2023:PBR

[SV23b] Maria Strazzullo and Fabio Vicini. POD-based reduced order methods for optimal control problems governed by parametric partial differential equation with varying boundary control. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003600.

Schmitz:2023:NCM

[SVW23] Yannick Schmitz, Duygu Vietz, and Egon Wanke. A note on the complexity of kmetric dimension. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003739.

Shi:2020:USA

[SW20] Dongyang Shi and Yanmi Wu. Uniformly superconvergent analysis of an efficient two-grid method for nonlinear bi-wave singular perturbation problem. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307647.

Sliwiak:2021:DDS

[ŚW21] Adam A. Śliwiak and Qiqi Wang. Differentiating densities on smooth manifolds. Applied Mathematics and Computation, 410 (??):Article 126444, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005336.

Shi:2022:HAA

[SW22] Dongyang Shi and Ran Wang. High accuracy analysis of Galerkin finite element method for Klein– Gordon–Zakharov equations. Applied Mathematics and Computation, 415 (??):Article 126701, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007852.

Singh:2024:NGP

[SW24] Ranveer Singh and Hitesh Wankhede. A note on graphs with purely imaginary per-spectrum. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002248.

Sutimin:2021:HHI

[SWCT21] Sutimin, Karunia Putra Wijaya, Joseph Páez Chávez, and Tianhai Tian. An in-host HIV-1 infection model incorporating quiescent and activated CD4^{+} T cells as well as CTL response. Applied Mathematics and Computation, 409 (??):Article 126410, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004999.

Si:2020:CLN

[SWF20] Yuanchao Si, JinRong Wang, and Michal Feckan. Controllability of linear and nonlinear systems governed by Stieltjes differential equations. Applied Mathematics and Computation, 376(??):Article 125139, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301089.

Song:2020:TBD

[SWFJ20] Fanpeng Song, Jianliang Wu, Suohai Fan, and Fei Jing. Transcendental behavior and disturbance behavior favor human development. Applied Mathematics and Computation, 378(??):Article 125182, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030151X.

Shi:2022:CGS

[SWP^{+}22] Zhenyu Shi, Wei Wei, Matjaz Perc, Baifeng Li, and Zhiming Zheng. Coupling group selection and network reciprocity in social dilemmas through multilayer networks. Applied Mathematics and Computation, 418 (??):Article 126835, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009188.

Shen:2020:RPC

[SWS20] Xiaojuan Shen, Yunxia Wang, and Zhiyong Si. A rotational pressure-correction projection methods for unsteady incompressible magnetohydrodynamics equations. Applied Mathematics and Computation, 387(??): Article 124488, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319304679.

Shen:2021:SPP

[SWS21] Hai-Long Shen, Hong-Yu Wu, and Xin-Hui Shao. A simplified PSS preconditioner for non-Hermitian generalized saddle point problems. Applied Mathematics and Computation, 394(??):Article 125810, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307633.

Shi:2022:LFQ

[SWS^{+}22a] Sangli Shi, Zhengxin Wang, Qiang Song, Min Xiao, and Guo-Ping Jiang. Leaderfollowing quasi-bipartite synchronization of coupled heterogeneous harmonic oscillators via eventtriggered control. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002466.

Su:2022:CAA

[SWS^{+}22b] Qingyu Su, Handong Wang, Chaowei Sun, Bo Li, and Jian Li. Cyber-attacks against cyber-physical power systems security: State estimation, attacks reconstruction and defense strategy. Applied Mathematics and Computation, 413 (??):Article 126639, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007232.

Song:2021:OBS

[SWSW21] Xiaona Song, Mi Wang, Shuai Song, and Zhen Wang. Observer-based sliding mode control for stochastic hyperbolic PDE systems with quantized output signal. Applied Mathematics and Computation, 393(??):Article 125643, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030597X.

Shi:2021:QAB

[SWW21a] Yao Shi, Jianhua Wu, and Yan’e Wang. Qualitative analysis for a biological depletion model. Applied Mathematics and Computation, 401(??):Article 126102, July 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001508.

Song:2021:MBM

[SWW21b] Yongli Song, Shuhao Wu, and Hao Wang. Memorybased movement with spatiotemporal distributed delays in diffusion and reaction. Applied Mathematics and Computation, 404(??): Article 126254, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003441.

Si:2022:PMN

[SWW22] Zhiyong Si, Mingyi Wang, and Yunxia Wang. A projection method for the non-stationary incompressible MHD coupled with the heat equations. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002910.

Shu:2021:SSM

[SWX^{+}21] Jinlong Shu, Baowei Wu, Lianglin Xiong, Tao Wu, and Haiyang Zhang. Stochastic stabilization of Markov jump quaternion-valued neural network using sampleddata control. Applied Mathematics and Computation, 400(??):Article 126041, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000898.

Shu:2022:SSC

[SWX22] Jinlong Shu, Baowei Wu, and Lianglin Xiong. Stochastic stability criteria and event-triggered control of delayed markovian jump quaternion-valued neural networks. Applied Mathematics and Computation, 420(??):Article 126904, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009875.

Sun:2024:CEC

[SWX24a] Chengbin Sun, Chaoqian Wang, and Haoxiang Xia. Co-evolution of cooperation and extortion with resource allocation in spatial multigame. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002443.

Sun:2024:ISE

[SWX24b] Fengxin Sun, Jufeng Wang, and Ying Xu. An improved stabilized elementfree Galerkin method for solving steady Stokes flow problems. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005155.

Sun:2020:MIF

[SWZS20] Shaoxin Sun, Yingchun Wang, Huaguang Zhang, and Jiayue Sun. Multiple intermittent fault estimation and tolerant control for switched T –S fuzzy stochastic systems with multiple time-varying delays. Applied Mathematics and Computation, 377 (??):Article 125114, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300837.

Shi:2023:BSE

[SWZZ23] Zhenyu Shi, Wei Wei, Hongwei Zheng, and Zhiming Zheng. Bidirectional supervision: an effective method to suppress corruption and defection under the third party punishment mechanism of donation games. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300139X.

Shang:2023:ACS

[SX23a] Lihui Shang and Zhiqiang Xu. Adaptive control strategy improves synchronization of self-propelled agents. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002710.

Sun:2023:NSG

[SX23b] Ruixue Sun and Yufeng Xu. Numerical solutions of Gelfand equation in steady combustion process. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007421.

Shi:2024:ATC

[SX24] Tianyu Shi and Zhengrong Xiang. Asymptotic tracking control for a class of uncertain nonlinear systems with output constraint. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003060.

Song:2021:CMT

[SXC21] Huiming Song, Yu Xiao, and Minghao Chen. Collocation methods for thirdkind Volterra integral equations with proportional delays. Applied Mathematics and Computation, 388 (??):Article 125509, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304677.

Sun:2022:GTG

[SXS22] Lu Sun, Qingshan Xu, and Yun Song. Gametheoretic genetic-priced optimization of multiple microgrids under uncertainties. Applied Mathematics and Computation, 426(??): ??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001291.

Skrzypek:2021:DSB

[SY21a] Leslaw Skrzypek and Yuncheng You. Dynamics and synchronization of boundary coupled FitzHugh–Nagumo neural networks. Applied Mathematics and Computation, 388(??):Article 125545, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305014.

Song:2021:OAE

[SY21b] Yunbo Song and Dan Ye. Optimizing DoS attack energy with imperfect acknowledgments and energy harvesting constraints in cyber-physical systems. Applied Mathematics and Computation, 393 (??):Article 125821, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307748.

Shao:2022:SSG

[SY22a] Zhenan Shao and Xiying Yuan. Some signed graphs whose eigenvalues are main. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200100X.

Sun:2022:PEE

[SY22b] Lingling Sun and Yidu Yang. The a posteriori error estimates and adaptive computation of nonconforming mixed finite elements for the Stokes eigenvalue problem. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000376.

Sun:2023:EPC

[SY23] Wensheng Sun and Yujun Yang. Extremal pentagonal chains with respect to the Kirchhoff index. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006087.

Shao:2021:ETR

[SYL21] Xingling Shao, Xiaohui Yue, and Jie Li. Eventtriggered robust control for quadrotors with preassigned time performance constraints. Applied Mathematics and Computation, 392 (??):Article 125667, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306202.

Shao:2023:DOT

[SYL23] Zeling Shao, Xiaoxiang Yu, and Zhiguo Li. On the dispersability of odd toroidal grids. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002564.

Sui:2024:ETR

[SYTC24] Shuai Sui, Yuelei Yu, Shaocheng Tong, and C. L. Philip Chen. Event-triggered robust fuzzy adaptive control for non-strict feedback nonlinear system with prescribed performance. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001735.

Shi:2021:MNS

[SYW21] Majun Shi, Zishen Yang, and Wei Wang. Minimum non-submodular cover problem with applications. Applied Mathematics and Computation, 410(??):Article 126442, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005312.

Song:2022:SCA

[SYX22] Huiming Song, Zhanwen Yang, and Yu Xiao. Superconvergence analysis of collocation methods for linear and nonlinear third-kind Volterra integral equations with non-compact operators. Applied Mathematics and Computation, 412 (??):Article 126562, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006469.

Su:2023:EIN

[SYZ^{+}23] Lichen Su, Zhengzhi Yang, Bowen Zhou, Naitian Zhang, and Yumeng Li. Effects of interdependent network reciprocity on the evolution of public cooperation. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001984.

Song:2021:PSN

[SZ21] Feifei Song and Jianjie Zhou. The partite saturation number of spider. Applied Mathematics and Computation, 394 (??):Article 125793, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307463. See

corrigendum [SZ22].

Song:2022:CPS

[SZ22] Feifei Song and Jianjie Zhou. Corrigendum to “The partite saturation number of spider” [Applied Mathematics and Computation 394 (2021) 125793]. Applied Mathematics and Computation, 418(??):Article 126815, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008985. See

[SZ21].

Shi:2023:USA

[SZ23] Dongyang Shi and Sihui Zhang. Unconditional superconvergence analysis for the nonlinear Biflux diffusion equation. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008396.

Shi:2024:USA

[SZ24a] Dongyang Shi and Sihui Zhang. Unconditional superconvergence analysis of an energy-stable L_{1} scheme for coupled nonlinear time-fractional preypredator equations with nonconforming finite element. Applied Mathematics and Computation, 467(??): ??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300663X.

Shi:2024:FDS

[SZ24b] Jiasheng Shi and Jun Zhao. H_{∞} filter design for switched descriptor systems with different derivative term matrices. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004630.

Shang:2023:NCF

[SZCC23] Weiying Shang, Weiwei Zhang, Dingyuan Chen, and Jinde Cao. New criteria of finite time synchronization of fractional-order quaternion-valued neural networks with time delay. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005586.

Stripinis:2021:MEA

[SZCP21] Linas Stripinis, Julius Zilinskas, Leocadio G. Casado, and Remigijus Paulavicius. On MATLAB experience in accelerating DIRECT– GLce algorithm for constrained global optimization through dynamic data structures and parallelization. Applied Mathematics and Computation, 390 (??):Article 125596, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305518.

Sheng:2021:IFD

[SZG21] Li Sheng, Sen Zhang, and Ming Gao. Intermittent fault detection for linear discrete-time stochastic multi-agent systems. Applied Mathematics and Computation, 410 (??):Article 126480, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005695.

Song:2021:CPD

[SZP21] Bo Song, Ya Zhang, and Ju H. Park. H_{∞} control for Poisson-driven stochastic systems. Applied Mathematics and Computation, 392(??):Article 125716, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030669X.

Song:2022:FAE

[SZSZ22] Xiaona Song, Renzhi Zhang, Shuai Song, and Yijun Zhang. Fuzzy adaptiveevent-triggered control for semi-linear parabolic PDE systems with stochastic actuator failures. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002119.

Sun:2021:HCS

[SZWL21] Jiayue Sun, Huaguang Zhang, Yingchun Wang, and Hongjing Liang. H_{∞} control for switched IT2 fuzzy nonlinear systems with multiple time delays applied in hybrid grid-connected generation. Applied Mathematics and Computation, 395(??):Article 125887, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308407.

Sun:2024:HDI

[SZXL24] Wei Sun, Zhiyuan Zhang, Xumeng Xie, and Shiyong Li. Heterogeneity of delay information and revenuecost structure in a multiserver queue with threshold policy: Customer behaviors and optimization. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005829.

Shi:2022:AFS

[SZZ22] Lingjuan Shi, Heping Zhang, and Lifang Zhao. The antiforcing spectra of (4, 6)- fullerenes. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322001746.

Sun:2024:PGW

[SZZ24] Chunhu Sun, Yaxian Zhang, and Heping Zhang. The polyomino graphs whose resonance graphs have a 1-degree vertex. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001760.

Song:2022:DTC

[SZZL22] Gongfei Song, Zimeng Zhang, Yanan Zhu, and Tao Li. Discrete-time control for highly nonlinear neutral stochastic delay systems. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003873.

Tawfik:2021:GFD

[TA21] Ashraf M. Tawfik and Hamdi M. Abdelhamid. Generalized fractional diffusion equation with arbitrary time varying diffusivity. Applied Mathematics and Computation, 410(??):Article 126449, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005385.

Takesue:2021:SBP

[Tak21] Hirofumi Takesue. Symmetry breaking in the prisoner’s dilemma on twolayer dynamic multiplex networks. Applied Mathematics and Computation, 388(??):Article 125543, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304999.

Tatsukawa:2022:SDS

[TAU^{+}22] Yuichi Tatsukawa, Md. Rajib Arefin, Shinobu Utsumi, Kazuki Kuga, and Jun Tanimoto. Stochasticity of disease spreading derived from the microscopic simulation approach for various physical contact networks. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004027.

Temimi:2024:EAS

[TBRB24] H. Temimi, M. Ben-Romdhane, and M. Baccouch. An efficient accurate scheme for solving the threedimensional Bratu-type problem. Applied Mathematics and Computation, 461 (??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032300485X.

Thanh:2020:BVL

[TC20a] Mai Duc Thanh and Dao Huy Cuong. Building a van Leertype numerical scheme for a model of two-phase flows. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307404.

Tiwari:2020:BFR

[TC20b] Sankalp Tiwari and Anindya Chatterjee. Basis functions for residual stresses. Applied Mathematics and Computation, 386(??):Article 125468, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304276.

Tang:2022:TWS

[TC22] Lu Tang and Shanpeng Chen. Traveling wave solutions for the diffusive Lotka–Volterra equations with boundary problems. Applied Mathematics and Computation, 413 (??):Article 126599, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006834.

Thinh:2023:FFO

[TCA23] Vo Duc Thinh, Thai Doan Chuong, and Nguyen Le Hoang Anh. Formulas of firstordered and second-ordered generalization differentials for convex robust systems with applications. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002837.

Tohme:2020:CTV

[TCG20] Fernando Tohmé, Gianluca Caterina, and Rocco Gangle. Computing truth values in the topos of infinite Peirce’s α-existential graphs. Applied Mathematics and Computation, 385 (??):Article 125343, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302964.

Tabernero:2024:HOW

[TCGR24] Víctor González Tabernero, Manuel J. Castro, and J. A. García-Rodríguez. Highorder well-balanced numerical schemes for onedimensional shallow-water systems with Coriolis terms. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006975.

Tankam-Chedjou:2021:OSM

[TCGTT21]  Israël  Tankam-Chedjou,

Frédéric Grognard, Jean Jules Tewa, and Suzanne Touzeau. Optimal and sustainable management of a soilborne banana pest. Applied Mathematics and Computation, 397(??):Article 125883, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308365.

Taraschi:2024:GDC

[TCPF24] G. Taraschi, M. R. Correa, A. S. Pinto, and C. O. Faria. A global H(div)-conforming finite element post-processing for stress recovery in nearly incompressible elasticity. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000596.

Triki:2020:NMW

[TCZ^{+}20] Houria Triki, Amitava Choudhuri, Qin Zhou, Anjan Biswas, and Ali Saleh Alshomrani. Nonautonomous matter wave bright solitons in a quasi-1D Bose– Einstein condensate system with contact repulsion and dipole–dipole attraction. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309439.

Tang:2022:CFT

[TCZ22] Lang Tang, Jing Chen, and Zhilin Zhang. Classification of flag-transitive pointprimitive non-symmetric 2 − (v, k, λ) designs with v < 100. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003526.

Tan:2024:FTS

[TDZ^{+}24] Cheng Tan, Jianying Di, Zhengqiang Zhang, Ziran Chen, and Wing Shing Wong. Finite-time stabilization of mean-field systems with uncertain parameters, multiple disturbances and delays. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400016X.

Toktas:2024:RBL

[TEGP24] Abdurrahim Toktas, Ugur Erkan, Suo Gao, and Chanil Pak. A robust bit-level image encryption based on Bessel map. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300509X.

Teng:2022:GBB

[Ten22]  Long Teng.  Gradient

boosting-based numerical methods for high-dimensional backward stochastic differential equations. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200203X.

Tang:2022:NWE

[TFL22] Shujiang Tang, Yujie Feng, and Mingjun Li. Novel weighted essentially nonoscillatory schemes with adaptive weights. Applied Mathematics and Computation, 420(??):Article 126893, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009760.

Taha:2020:ZDS

[TG20] Mohammad A. Taha and Ayman Ghoneim. Zerodeterminant strategies in repeated asymmetric games. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308549.

Trull:2022:MSS

[TGDPS22] Oscar Trull, J. Carlos García-Díaz, and A. Peiró-Signes. Multiple seasonal STL decomposition with discrete-interval moving seasonalities. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004726.

Tian:2024:PSI

[TGL^{+}24] Yue Tian, Shun Gao, Haihong Li, Qionglin Dai, and Junzhong Yang. Particle swarm intelligence promotes cooperation by adapting interaction radii in co-evolutionary games. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001498.

Torres-Hernandez:2022:AOC

[THBPMC22] A. Torres-Hernandez, F. Brambila-Paz, and R. Montufar-Chaveznava. Acceleration of the order of convergence of a family of fractional fixedpoint methods and its implementation in the solution of a nonlinear algebraic system related to hybrid solar receivers. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003058.

Tao:2021:DUG

[THW^{+}21] Shi Tao, Qing He, Liang Wang, Baiman Chen, Jiechao Chen, and Yousheng Lin. Discrete unified gas kinetic scheme simulation of conjugate heat transfer problems in complex geometries by a ghost-cell interface method. Applied Mathematics and Computation, 404(??):Article 126228, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003180.

Tao:2023:FIC

[THZ23] Zheng Tao, Zhi Hu, and Zijian Zhou. Faster isogeny computation on twisted Hessian curves. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008918.

Tiplea:2020:DQR

[TITN20] Ferucio Laurentiu Tiplea, Sorin Iftene, George Teseleanu, and Anca-Maria Nica. On the distribution of quadratic residues and non-residues modulo composite integers and applications to cryptography. Applied Mathematics and Computation, 372(??): Article 124993, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309853.

Torres-Jimenez:2021:CAE

[TJAJAG21] Jose Torres-Jimenez, Brenda Acevedo-Juárez, and Himer Avila-George. Covering array EXtender. Applied Mathematics and Computation, 402(??):Article 126122, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001703.

Torres-Jimenez:2020:ICA

[TJIM20] Jose Torres-Jimenez and Idelfonso Izquierdo-Marquez. Improved covering arrays using covering perfect hash families with groups of restricted entries. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308185.

Tsoutsanis:2023:RPM

[TKF23] Panagiotis Tsoutsanis, Machavolu Sai Santosh Pavan Kumar, and Pericles S. Farmakis. A relaxed a posteriori MOOD algorithm for multicomponent compressible flows using highorder finite-volume methods on unstructured meshes. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200618X.

Takizawa:2023:DRO

[TKY23] Michiaki Takizawa, Koichi Kobayashi, and Yuh Yamashita. Design of reducedorder and pinning controllers for probabilistic Boolean networks using reinforcement learning. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300323003806.

Tan:2022:FDT

[TLC22] Zhijun Tan, Kang Li, and Yanping Chen. A fully discrete two-grid finite element method for nonlinear hyperbolic integro-differential equation. Applied Mathematics and Computation, 413(??):Article 126596, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006809.

Teng:2024:USB

[TLCM24] Zhaolin Teng, Dan Li, Yuanyuan Chen, and Jixiang Meng. Unbalanced signed bicyclic graphs minimizing the least eigenvalue. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006471.

Tu:2022:MMD

[TLD22] Jianhua Tu, Yuxin Li, and Junfeng Du. Maximal and maximum dissociation sets in general and triangle-free graphs. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001916.

Tian:2021:SRI

[TLDL21] Zhaolu Tian, Xiaojing Li, Yinghui Dong, and Zhongyun Liu. Some relaxed iteration methods for solving matrix equation AXB = C. Applied Mathematics and Computation, 403(??):Article 126189, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002794.

Tan:2021:OFL

[TLHH21] Lihua Tan, Chuandong Li, Junjian Huang, and Tingwen Huang. Output feedback leader-following consensus for nonlinear stochastic multiagent systems: the event-triggered method. Applied Mathematics and Computation, 395 (??):Article 125879, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308328.

Tang:2020:MSD

[TLJO20] Chunming Tang, Shuai Liu, Jinbao Jian, and Xiaomei Ou. A multi-step doubly stabilized bundle method for nonsmooth convex optimization. Applied Mathematics and Computation, 376(??):Article 125154, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301235.

Thi:2021:NDB

[TLPT21] Hoai An Le Thi, Hoai Minh Le, Duy Nhat Phan, and Bach Tran. Novel DCA based algorithms for a special class of nonconvex problems with application in machine learning. Applied Mathematics and Computation, 409 (??):Article 125904, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320308572.

Tong:2020:EAS

[TLPY20] Fenghua Tong, Lixiang Li, Haipeng Peng, and Yixian Yang. An effective algorithm for the spark of sparse binary measurement matrices. Applied Mathematics and Computation, 371(??):

??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309579.

Tan:2022:CSC

[TlS^{+}22] Zilong Tan, Yang liu, Jiayue Sun, Huaguang Zhang, and Xiangpeng Xie. Chaos synchronization control for stochastic nonlinear systems of interior PMSMs based on fixed-time stability theorem. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001990.

Tian:2024:SAL

[TLY24] Xiu-Qin Tian, Shu-Jun Liu, and Xue Yang. Stochastic adaptive linear quadratic nonzero-sum differential games. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002649.

Tian:2020:DDS

[TLZ20] Canrong Tian, Zhi Ling, and Lai Zhang. Delaydriven spatial patterns in a network-organized semiarid vegetation model. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307702.

Thomes:2020:PFE

[TM20] Renan Lima Thomes and Fernando C. M. Menandro. Polygonal finite element: a comparison of the stiffness matrix integration methods. Applied Mathematics and Computation, 375(??): Article 125089, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300588.

Titarev:2022:ALE

[TM22a] V. A. Titarev and A. A. Morozov. Arbitrary Lagrangian– Eulerian discrete velocity method with application to laser-induced plume expansion. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003150.

Tripathi:2022:NSA

[TM22b] Vinit Kumar Tripathi and Amit Mahajan. Nonlinear stability analysis of double diffusive convection in a fluid saturated porous layer with variable gravity and throughflow. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001461.

Thumma:2022:TDN

[TMA^{+}22] Thirupathi Thumma, S. R.

Mishra, M. Ali Abbas, M. M. Bhatti, and Sara I. Abdelsalam. Three-dimensional nanofluid stirring with nonuniform heat source/sink through an elongated sheet. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000133.

Tao:2020:SAS

[TMW20] Ruifeng Tao, Yuechao Ma, and Chunjiao Wang. Stochastic admissibility of singular Markov jump systems with time-delay via sliding mode approach. Applied Mathematics and Computation, 380(??):Article 125282, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302514.

Tian:2023:DGN

[TNL23] Fangyu Tian, Lin Niu, and Xiangwen Li. Decompositions of graphs of nonnegative characteristic with some forbidden subgraphs. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002953.

Takano:2023:DDP

[TNT23] Masanori Takano, Kenichi Nakazato, and Fumiaki Taka. Dynamics of discrimination and prejudice via two types of social contagion. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000851.

Thang:2021:MAB

[TNTL21] Pham Toan Thang, T. Nguyen-Thoi, and Jaehong Lee. Modeling and analysis of bi-directional functionally graded nanobeams based on nonlocal strain gradient theory. Applied Mathematics and Computation, 407(??):Article 126303, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003921.

Tommasini:2020:FSS

[TO20] Daniele Tommasini and David N. Olivieri. Fast switch and spline scheme for accurate inversion of nonlinear functions: the new first choice solution to Kepler’s equation. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306691.

Trujillo-Pino:2024:FNS

[TPSP24] Agustín Trujillo-Pino, Jose Pablo Suárez, and Miguel A. Padrón. Finite number of similarity classes in longest edge bisection of nearly equilateral tetrahedra. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001036.

Truong:2024:NOB

[TPTA24] Hoai Vu Anh Truong, Van Du Phan, Duc Thien Tran, and Kyoung Kwan Ahn. A novel observerbased neural-network finitetime output control for high-order uncertain nonlinear systems. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001711.

Tang:2024:EBA

[TQZ^{+}24] Tianfeng Tang, Gang Qin, Bin Zhang, Jun Cheng, and Jinde Cao. Eventbased asynchronous state estimation for Markov jump memristive neural networks. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001255.

Treinen:2023:SMC

[Tre23] Ray Treinen. Spectral methods for capillary surfaces described by bounded generating curves. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000553.

Travenec:2022:GCZ

[TS22] Igor Travenec and Ladislav Samaj. Generation of offcritical zeros for hypercubic Epstein zeta functions. Applied Mathematics and Computation, 413 (??):Article 126611, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006950.

Tamang:2023:ADM

[TS23] Nutthavuth Tamang and Yi Sun. Application of the dynamic Monte Carlo method to pedestrian evacuation dynamics. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000450.

Tan:2020:GBT

[TT20] Ta Sheng Tan and Wen Chean Teh. Graph burning: Tight bounds on the burning numbers of path forests and spiders. Applied Mathematics and Computation, 385 (??):Article 125447, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304082.

Tan:2023:BDS

[TT23] Ta Sheng Tan and Wen Chean Teh. Burnability of double spiders and path forests. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006488.

Tian:2024:EMD

[TT24] Yuting Tian and Jianhua Tu. Enumerating maximal dissociation sets in three classes of grid graphs. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001838.

Tian:2023:IBA

[TTC^{+}23] Yang Tian, Hui Tian, Yajuan Cui, Xuzhen Zhu, and Qimei Cui. Influence of behavioral adoption preference based on heterogeneous population on multiple weighted networks. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000498.

Tian:2023:DFN

[TTCY23] He-Yuan Tian, Bo Tian, Su-Su Chen, and Dan-Yu Yang. Degenerate fundamental nonlinear waves on the nonzero-zero background for a coupled Hirota system in a birefringent fiber. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006166.

Truhar:2020:PTH

[TTL20] Ninoslav Truhar, Zoran Tomljanović, and Ren-Cang Li. Perturbation theory for Hermitian quadratic eigenvalue problem — damped and simultaneously diagonalizable systems. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309130.

Tiwari:2020:MFE

[TTMU20] Vandana Tiwari, Jai Prakash Tripathi, Swati Mishra, and Ranjit Kumar Upadhyay. Modeling the fear effect and stability of non-equilibrium patterns in mutually interfering predator–prey systems. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309403.

Tortosa:2021:ARN

[TVY21] Leandro Tortosa, Jose F. Vicent, and Gevorg Yeghikyan. An algorithm for ranking the nodes of multiplex networks with data based on the PageRank concept. Applied Mathematics and Computation, 392(??): Article 125676, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306299.

Tian:2021:DFS

[TW21a] Yufeng Tian and Zhanshan Wang. Dissipative filtering for singular Markovian jump systems with generally hybrid transition rates. Applied Mathematics and Computation, 411 (??):Article 126492, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005816.

Tian:2021:NRP

[TW21b] Yufeng Tian and Zhanshan Wang. A new result on H_{∞} performance state estimation for static neural networks with time-varying delays. Applied Mathematics and Computation, 388 (??):Article 125556, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305129.

Tan:2022:ETI

[TWB22] Hailian Tan, Jianwei Wu, and Haibo Bao. Eventtriggered impulsive synchronization of fractionalorder coupled neural networks. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003186.

Tmar:2023:RCN

[TWB23] Zied Tmar, Taieb Wafi, and Mongi Besbes. Revisited control of non regular singular system by using adaptive scheduled gain — case of methanizer. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004150.

Tan:2020:PSE

[TWL20] Guoqiang Tan, Zhanshan Wang, and Cong Li. H_{∞} performance state estimation of delayed static neural networks based on an improved proportionalintegral estimator. Applied Mathematics and Computation, 370(??):??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309002.

Tian:2020:SAC

[TWR20] Yufeng Tian, Yuzhong Wang, and Junchao Ren. Stability analysis and control design of singular Markovian jump systems via a parameter-dependent reciprocally convex matrix inequality. Applied Mathematics and Computation, 386(??):Article 125471, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304306.

Tian:2024:PTS

[TWWL24] Zhaolu Tian, Yudong Wang, Nian-Ci Wu, and Zhongyun Liu. On the parameterized two-step iteration method for solving the matrix equation AXB = C. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005702.

Tesfay:2020:TBM

[TWZ^{+}20] Daniel Tesfay, Pingyuan Wei, Yayun Zheng, Jinqiao Duan, and Jürgen Kurths. Transitions between metastable states in a simplified model for the thermohaline circulation under random fluctuations. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308604.

Tian:2021:GPN

[TX21] Jing Tian and Kexiang Xu. The general position number of Cartesian products involving a factor with small diameter. Applied Mathematics and Computation, 403(??):Article 126206, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002964.

Tu:2024:NAB

[TXZ24] Kairong Tu, Yu Xue, and Xian Zhang. A new approach based on system solutions for passivity analysis of discrete-time memristor-based neural networks with time-varying delays. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000237.

Tang:2021:UCR

[TY21] Han Tang and Xiangfeng Yang. Uncertain chemical reaction equation. Applied Mathematics and Computation, 411(??):Article 126479, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005683.

Tang:2022:MEU

[TY22] Han Tang and Xiangfeng Yang. Moment estimation in uncertain differential equations based on the Milstein scheme. Applied Mathematics and Computation, 418(??):Article 126825, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009085.

Tan:2023:ETD

[TYCK23] Zhijun Tan, Junxiang Yang, Jianjun Chen, and Junseok Kim. An efficient timedependent auxiliary variable approach for the threephase conservative Allen– Cahn fluids. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006725.

Tang:2020:BTV

[TYL^{+}20] Jianxiong Tang, Chaohui Yang, Shangzhao Li, Shaojun Dai, and Weijun Liu. Block transitive 2 − (v, k, 1) designs and PGL_{2}(q) groups. Applied Mathematics and Computation, 374(??):Article 125034, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300035.

Tian:2021:FTS

[TZ21] Xue Tian and Yi Zhang. Fractional time-scales Noether theorem with Caputo Δ derivatives for Hamiltonian systems. Applied Mathematics and Computation, 393(??):Article 125753, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307062.

Tsatsomeros:2022:RRO

[TZ22] Michael J. Tsatsomeros and Faith Zhang. Recursive rank one perturbations for pole placement and cone reachability. Applied Mathematics and Computation, 416(??): Article 126732, March 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008146.

Teng:2023:FTE

[TZ23a] Ye Teng and Zhimin Zhang. Finite-time expected present value of operating costs until ruin in a Cox risk model with periodic observation. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002436.

Toprakseven:2023:EAW

[TZ23b] Suayip Toprakseven and Peng Zhu. Error analysis of a weak Galerkin finite element method for twoparameter singularly perturbed differential equations in the energy and balanced norms. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007512.

Tan:2024:TSO

[TZ24] Zhijun Tan and Yunhua Zeng. Temporal secondorder fully discrete two-grid methods for nonlinear timefractional variable coefficient diffusion-wave equations. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006264.

Tian:2022:INI

[TZD22] Jiayue Tian, Xueyan Zhao, and Feiqi Deng. Incremental Newton’s iterative algorithm for optimal control of Itô stochastic systems. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000443.

Tian:2020:SCM

[TZJ20] Hao Tian, Jing Zhang, and Lili Ju. A spectral collocation method for nonlocal diffusion equations with volume constrained boundary conditions. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309221.

Tan:2024:OOH

[TZL24] Jipeng Tan, Man Zhang, and Fengming Liu. Onlineoffline higher-order rumor propagation model based on quantum cellular automata considering social adaptation. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004897.

Tong:2024:DIU

[TZLL24] Fenghua Tong, Dawei Zhao, Xin Li, and Lixiang Li. Designing incoherent unit norm tight frames via block coordinate descentbased alternating projection. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001590.

Tian:2022:BPC

[TZW22] Huimin Tian, Lingling Zhang, and Xin Wang. Blow-up phenomena in a class of coupled reaction– diffusion system with nonlocal boundary conditions. Applied Mathematics and Computation, 414(??):Article 126667, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007517.

Ullah:2021:SGS

[UBSU21] Fahim Ullah, Noor Badshah, Hassan Shah, and Asmat Ullah. Sobolev gradients for segmentation of vector-valued texture images. Applied Mathematics and Computation, 402 (??):Article 125962, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000102.

Ueda:2020:CKE

[Ued20] Masahiko Ueda. Common knowledge equilibrium of Boolean securities in distributed information market. Applied Mathematics and Computation, 386 (??):Article 125540, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304963.

Ueda:2023:MTS

[Ued23] Masahiko Ueda. Memorytwo strategies forming symmetric mutual reinforcement learning equilibrium in repeated prisoners’ dilemma game. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008876.

Urena:2020:NLF

[UGG^{+}20] Francisco Ureña, Luis Gavete, Ángel García Gómez, Juan José Benito, and Antonio Manuel Vargas. Non-linear Fokker– Planck equation solved with generalized finite differences in 2D and 3D. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307933.

UlKabir:2024:DPI

[UH24] Md Wasi Ul Kabir and Md Tamjidul Hoque. Dis-Predict3.0: Prediction of intrinsically disordered regions/proteins using protein language model. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001024.

Usman:2020:NOM

[UHZ^{+}20] M. Usman, M. Hamid, T. Zubair, R. U. Haq, W. Wang, and M. B. Liu. Novel operational matricesbased method for solving fractional-order delay differential equations via shifted Gegenbauer polynomials. Applied Mathematics and Computation, 372(??): Article 124985, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309774.

Uzunca:2023:LIM

[UK23] Murat Uzunca and Bülent Karasözen. Linearly implicit methods for Allen– Cahn equation. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001534.

Uzunca:2023:GEP

[UKA23] Murat Uzunca, Bülent Karasözen, and Ayhan Aydin. Global energy preserving model reduction for multi-symplectic PDEs. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005574.

Ugarte:2022:FGE

[UMT22] Juan P. Ugarte, J. A. Tenreiro Machado, and Catalina Tobón. Fractional generalization of entropy improves the characterization of rotors in simulated atrial fibrillation. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001618.

Udhayakumar:2021:MTS

[URLC21] K. Udhayakumar, R. Rakkiyappan, Xiaodi Li, and Jinde Cao. Mutiple ψ-type stability of fractional-order quaternion-valued neural networks. Applied Mathematics and Computation, 401(??):Article 126092, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001405.

Usui:2021:SEM

[UU21] Yuki Usui and Masahiko Ueda. Symmetric equilibrium of multi-agent reinforcement learning in repeated prisoner’s dilemma. Applied Mathematics and Computation, 409(??):Article 126370, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004598.

Upadhyay:2021:MRI

[UUS21] Prateep Upadhyay, S. K. Upadhyay, and K. K. Shukla. Magnetic resonance images denoising using a wavelet solution to Laplace equation associated with a new variational model. Applied Mathematics and Computation, 400 (??):Article 126083, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001314.

Ullah:2020:PTG

[UZ20] Sheik Ahmed Ullah and Shan Zhao. Pseudotransient ghost fluid methods for the Poisson–Boltzmann equation with a twocomponent regularization. Applied Mathematics and Computation, 380(??):Article 125267, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302368.

Villa:2020:TDS

[VBM20] Andrea Villa, Luca Barbieri, and Roberto Malgesini. Three dimensional simulation of the dynamics of electro active polymers using shell elements. Applied Mathematics and Computation, 377 (??):Article 125160, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301296.

Villa:2021:DPE

[VBMB21] Andrea Villa, Luca Barbieri, Roberto Malgesini, and Giacomo Buccella. Discretization of Poisson’s equation in two domains with non algebraic interface conditions for plasma simulations. Applied Mathematics and Computation, 403 (??):Article 126179, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002691.

Vaghi:2022:AAB

[VBP22] Cristina Vaghi, Sebastien Benzekry, and Clair Poignard. Asymptotic analysis of a biphase tumor fluid flow: the weak coupling case. Applied Mathematics and Computation, 413 (??):Article 126635, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007190.

vanderMerwe:2021:CDP

[vdMvRS21] A. J. van der Merwe, N. F. J. van Rensburg, and R. H. Sieberhagen. Comparing the dual phase lag, Cattaneo–Vernotte and Fourier heat conduction models using modal analysis. Applied Mathematics and Computation, 396(??): Article 125934, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308870.

vanderZwaard:2021:CAH

[vdZGO21] Thomas van der Zwaard, Lech A. Grzelak, and Cornelis W. Oosterlee. A computational approach to hedging Credit Valuation Adjustment in a jumpdiffusion setting. Applied Mathematics and Computation, 391(??):Article 125671, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030624X.

Vellasco-Gomes:2020:EBW

[VGdFCBA20] Arianne Vellasco-Gomes, Rubens de Figueiredo Camargo, and Alexys Bruno-Alfonso. Energy bands and Wannier functions of the fractional Kronig-Penney model. Applied Mathematics and Computation, 380(??):Article 125266, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302356.

Vieira:2021:AHD

[VGR^{+}21] Marcelo Bernardes Vieira, Gilson Antonio Giraldi, Allan Carlos Amaral Ribeiro, Marcelo Caniato Renhe, and Claudio Esperança. Anisotropic Helmholtz decomposition for controlled fluid simulation. Applied Mathematics and Computation, 411(??):Article 126501, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005907.

Vignesh:2023:MDT

[VHB23] D. Vignesh, Shaobo He, and Santo Banerjee. Modelling discrete time fractional Rucklidge system with complex state variables and its synchronization. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002801.

Viscardi:2023:ODI

[Vis23]  Alberto Viscardi.  Op-

timized dual interpolating subdivision schemes. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003843.

Vadivoo:2022:RCA

[VJA^{+}22] B. S. Vadivoo, G. Jothilakshmi, Y. Almalki, A. Debbouche, and M. Lavanya. Relative controllability analysis of fractional order differential equations with multiple time delays. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002661.

Volna:2021:TSF

[VJKZ21] Eva Volna, Robert Jarusek, Martin Kotyrba, and Jaroslav Zacek. Training set fuzzification based on histogram to increase the performance of a neural network. Applied Mathematics and Computation, 398 (??):Article 125994, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000424.

Vicentin:2020:MMA

[VMCG20] Daniel Chieregato Vicentin, Paulo F. A. Mancera, Tiago Carvalho, and Luiz Fernando Gonçalves. Mathematical model of an antiretroviral therapy to HIV via Filippov theory. Applied Mathematics and Computation, 387(??):Article 125179, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030148X.

Villa-Morales:2023:PDM

[VMR23] José Villa-Morales and Luis Rincón. Probability distributions and the maximum entropy principle. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008748.

Veiga:2024:IEA

[VMT24] Maria Han Veiga, Lorenzo Micalizzi, and Davide Torlo. On improving the efficiency of ADER methods.

Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005957.

Verdiere:2020:IPC

[VMZDV20] Nathalie Verdière, David Manceau, Shousheng Zhu, and Lilianne Denis-Vidal. Inverse problem for a coupling model of reactiondiffusion and ordinary differential equations systems. Application to an epidemiological model. Applied Mathematics and Computation, 375(??):Article 125067, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300369.

Vulanovic:2020:RHS

[VN20] Relja Vulanović and Thái Anh Nhan. Robust hybrid schemes of higher order for singularly perturbed convection-diffusion problems. Applied Mathematics and Computation, 386 (??):Article 125495, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304537.

Voet:2023:FAF

[Voe23] Yannis Voet. On the fast assemblage of finite element matrices with application to nonlinear heat transfer problems. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005902.

Veeresha:2020:ETN

[VPK20] P. Veeresha, D. G. Prakasha, and Devendra Kumar. An efficient technique for nonlinear time-fractional Klein–Fock–Gordon equation. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306290.

Vujosevic:2021:AGI

[VPV^{+}21] Sasa Vujosević, Goran Popivoda, Zana Kovijanić Vukićević, Boris Furtula, and Riste Skrekovski. Arithmetic–geometric index and its relations with geometric–arithmetic index. Applied Mathematics and Computation, 391 (??):Article 125706, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306597.

Vandecasteele:2024:CES

[VS24] Hannes Vandecasteele and Giovanni Samaey. Computational efficiency study of a micro-macro Markov chain Monte Carlo method for molecular dynamics. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001553.

Vijayakumar:2021:PIO

[VSM^{+}21] M. Vijayakumar, R. Sakthivel, Ardashir Mohammadzadeh, S. A. Karthick, and S. Marshal Anthoni. Proportional integral observer based tracking control design for Markov jump systems. Applied Mathematics and Computation, 410(??):Article 126467, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005567.

Visakamoorthi:2022:HMM

[VSM22] B. Visakamoorthi, K. Subramanian, and P. Muthukumar. Hidden Markov model based non-fragile sampled-data control design for mode-dependent fuzzy systems with actuator faults. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005288.

Vieira:2023:GAR

[VSM23] Igor P. Vieira, Fernando O.

Souza, and Leonardo A. Mozelli. Geometric approach to robust stability analysis of linear parametervarying systems: Computational trade-offs between the exact and the simplex convex hulls. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001868.

Vynnycky:2023:BIO

[Vyn23] M. Vynnycky. On boundary immobilization for onedimensional Stefan-type problems with a moving boundary having initially parabolic-logarithmic behaviour. Applied Mathematics and Computation, 444(??):??, May 1, 2023.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300322008712.

Vavpetic:2023:OAS

[VZ23] Ales Vavpetic and Emil Zagar. Optimal approximation of spherical squares by tensor product quadratic Bézier patches. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300365X.

Wang:2022:AFT

[WA22] Huanqing Wang and Yingdong Ai. Adaptive fixedtime control and synchronization for hyperchaotic Lü systems. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004623.

Wang:2021:TOV

[Wan21a] Jiqiang Wang. Topology optimization for vibration suppression at multiple locations: a non-iterative approach. Applied Mathematics and Computation, 398 (??):Article 125990, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000382.

Wang:2021:ENM

[Wan21b] Shaoxin Wang. An efficient numerical method for condition number constrained covariance matrix approximation. Applied Mathematics and Computation, 397 (??):Article 125925, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030878X.

Wang:2022:MCM

[Wan22a] Min Wang. Multistep collocation method for Fredholm integral equations of the second kind. Applied Mathematics and Computation, 420(??):Article 126870, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100953X.

Wang:2022:CCT

[Wan22b] Zhenhua Wang. Consensus of continuous-time agent dynamics with unknown input and communication delays. Applied Mathematics and Computation, 412 (??):Article 126563, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006470.

Wang:2022:MEG

[Wan22c] Zhiwen Wang. The multiplicities of eigenvalues of a graph. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000431.

Wang:2023:NPC

[Wan23a]  Weilong Wang.  Novel

pressure-correction schemes based on scalar auxiliary variable method for the MHD equations. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006245.

Wang:2023:FQJ

[Wan23b] Yupin Wang. Fractional quantum Julia set. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002461.

Wang:2024:ETS

[Wan24a]  Chaoqian Wang.  Evo-

lution of trust in structured populations. Applied Mathematics and Computation, 471(??):??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000675.

Wang:2024:PEC

[Wan24b] Song Wang. Pricing European call options with interval-valued volatility and interest rate. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400170X.

Wu:2020:FTC

[WB20] Xifen Wu and Haibo Bao.

Finite time complete synchronization for fractionalorder multiplex networks. Applied Mathematics and Computation, 377(??):Article 125188, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301570.

Wu:2023:SEM

[WB23] Xifen Wu and Haibo Bao. H_{∞} state estimation for multiplex networks with randomly occurring sensor saturations. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006129.

Wang:2022:AAB

[WBV22] Zihao Wang, A. Bayliss, and V. A. Volpert. Asymptotic analysis of the bistable Lotka–Volterra competitiondiffusion system. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004453.

Wang:2024:CAF

[WBV24] Z. Wang, A. Bayliss, and V. A. Volpert. Competing alliances in a fourspecies cyclic ecosystem. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005659.

Wang:2021:ESO

[WBY^{+}21] Yuyao Wang, Zhan Bu, Huan Yang, Hui-Jia Li, and Jie Cao. An effective and scalable overlapping community detection approach: Integrating social identity model and game theory. Applied Mathematics and Computation, 390 (??):Article 125601, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305567.

Wang:2021:TFP

[WC21] Yihong Wang and Jianxiong Cao. A tailored finite point method for subdiffusion equation with anisotropic and discontinuous diffusivity. Applied Mathematics and Computation, 401(??): Article 125907, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308602.

Wang:2022:EBJ

[WC22a] Licheng Wang and Hui Chen. Event-based joint state and fault estimation: the envelope-constrained H_{∞} criterion. Applied Mathematics and Computation, 419(??):Article 126873, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009565.

Wei:2022:DDB

[WC22b] Lili Wei and Mou Chen. Distributed DETMs-based internal collision avoidance control for UAV formation with lumped disturbances. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004362.

Wu:2022:IGC

[WC22c] Cang Wu and Jianlong Chen. The {1, 2, 3, 1^{m}}- inverses: a generalization of core inverses for matrices. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002284.

Wang:2023:CSMa

[WC23] Jie Wang and Xin Chen.

H_{∞} consensus for stochastic Markov jump multiagent systems with imperfect time-varying transition probabilities and multiplicative noise. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005781.

Wang:2024:SEH

[WC24] Yuxi Wang and Yang Chen.

The smallest eigenvalue of the Hankel matrices associated with a perturbed Jacobi weight. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000870.

Wang:2020:EBS

[WCG^{+}20] Yuxiao Wang, Yuting Cao, Zhenyuan Guo, Tingwen Huang, and Shiping Wen. Event-based sliding-mode synchronization of delayed memristive neural networks via continuous/periodic sampling algorithm. Applied Mathematics and Computation, 383(??):Article 125379, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032030343X.

Wang:2020:PPM

[WCGW20] Yuxiao Wang, Yuting Cao, Zhenyuan Guo, and Shiping Wen. Passivity and passification of memristive recurrent neural networks with multi-proportional delays and impulse. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308306.

Wang:2021:EBS

[WCM^{+}21] Dongji Wang, Fei Chen, Bo Meng, Xingliu Hu, and Jing Wang. Eventbased secure H_{∞} load frequency control for delayed power systems subject to deception attacks. Applied Mathematics and Computation, 394(??):Article 125788, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307414.

Wang:2020:REL

[WCP20] Kuaini Wang, Jinde Cao, and Huimin Pei. Robust extreme learning machine in the presence of outliers by iterative reweighted algorithm. Applied Mathematics and Computation, 377 (??):Article 125186, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301557.

Wagner:2022:POS

[WCS22] Michael Wagner, Charles J.

Colbourn, and Dimitris E. Simos. In-parameter-order strategies for covering perfect hash families. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000388.

Wang:2021:PCR

[WCW21] Le Wang, Tong Chen, and Zhenghong Wu. Promoting cooperation by reputation scoring mechanism based on historical donations in public goods game. Applied Mathematics and Computation, 390 (??):Article 125605, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305609.

Wang:2022:HSM

[WCW22] Haitao Wang, Xiangyong Chen, and Jing Wang. H_{∞} sliding mode control for PDT-switched nonlinear systems under the dynamic event-triggered mechanism. Applied Mathematics and Computation, 412 (??):Article 126474, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005634.

Wang:2023:LCE

[WCWZ23] Jun Wang, Shimin Cai, Wei Wang, and Tao Zhou. Link cooperation effect of cooperative epidemics on complex networks. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006117.

Wu:2022:SDS

[WCX^{+}22] Tao Wu, Jinde Cao, Lianglin Xiong, Haiyang Zhang, and Jinlong Shu. Sampleddata synchronization criteria for Markovian jumping neural networks with additive time-varying delays using new techniques. Applied Mathematics and Computation, 413(??):Article 126604, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006883.

Wang:2022:DIG

[WCXS22] Shengxian Wang, Xiaojie Chen, Zhilong Xiao, and Attila Szolnoki. Decentralized incentives for general wellbeing in networked public goods game. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003824.

Wang:2020:CCR

[WCY20a] Peizhen Wang, Yanping Chen, and Wei Yang. Corrigendum to “Curl recovery for the lowest order rectangular edge element” [Applied Mathematics and Computation 371 (2020) 124897]. Applied Mathematics and Computation, 372 (??):Article 125030, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319310227. See

[WCY20b].

Wang:2020:CRL

[WCY20b] Peizhen Wang, Yanping Chen, and Wei Yang. Curl recovery for the lowest order rectangular edge element. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308896.

See corrigendum [WCY20a].

Wang:2022:WBR

[WCY^{+}22] Jianwei Wang, Wei Chen, Fengyuan Yu, Jialu He, and Wenshu Xu. Wealthbased rule favors cooperation in costly public goods games when individual selection is inevitable. Applied Mathematics and Computation, 414 (??):Article 126668, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007529.

Wang:2024:ECC

[WCY^{+}24] Jianwei Wang, Wei Chen, Fengyuan Yu, Siyuan Zhou, Jialu He, Wenshu Xu, and Wenhui Dai. The emergence of cooperation in the context of prior agreement with threshold and posterior compensation. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001449.

Wang:2020:MHS

[WCZ20] Bo Wang, Jun Cheng, and Xia Zhou. A multiple hierarchical structure strategy to quantized control of Markovian switching systems. Applied Mathematics and Computation, 373 (??):Article 125037, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300060.

Wu:2021:AFN

[WCZ^{+}21] Yuyan Wu, Jun Cheng, Xia Zhou, Jinde Cao, and Mengzhuo Luo. Asynchronous filtering for nonhomogeneous Markov jumping systems with deception attacks. Applied Mathematics and Computation, 394 (??):Article 125790, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307438.

Wu:2022:RGD

[WCZ22a] Nian-Ci Wu, Ling-Xia Cui, and Qian Zuo. On the relaxed greedy deterministic row and column iterative methods. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004131.

Wu:2022:DET

[WCZ^{+}22b] Zhenyu Wu, Jiawei Chen, Xuexi Zhang, Zehui Xiao, Jie Tao, and Xiaofeng Wang. Dynamic eventtriggered synchronization of complex networks with switching topologies: Asynchronous observer-based case. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004878.

Wang:2020:DEB

[WCZY20] Yanqian Wang, Fu Chen, Guangming Zhuang, and Guang Yang. Dynamic event-based mixed H_{∞} and dissipative asynchronous control for Markov jump singularly perturbed systems. Applied Mathematics and Computation, 386 (??):Article 125443, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304045.

Wan:2023:EEA

[WDC^{+}23] Jianjun Wan, Guozheng Dai, Shaochun Chen, Wen Dong, and Shicang Song. Error estimation of anisotropic interpolation for serendipity elements of arbitrary degree. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004411.

Wang:2021:CCL

[WDG21] Xiaofeng Wang, Weizhong Dai, and Shuangbing Guo. Corrigendum to “A conservative linear difference scheme for the 2D regularized long-wave equation” [Appl. Math. Comput. 342 (2019) 55–70]. Applied Mathematics and Computation, 395(??):Article 125909, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308626.

Wang:2022:LFC

[WDTF22] Kun-Peng Wang, Dong Ding, Ze Tang, and Jianwen Feng. Leader-following consensus of nonlinear multiagent systems with hybrid delays: Distributed impulsive pinning strategy. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001175.

Wang:2022:URN

[WDW^{+}22] Zi-Ren Wang, Zheng-Hong Deng, Huan-Bo Wang, HuXiong Li, and Fei-Wang X. Uneven resources network promotes cooperation in the Prisoner’s Dilemma game. Applied Mathematics and Computation, 413 (??):Article 126619, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007037.

Wu:2021:ISF

[WDYS21] Y. H. Wu, C. Y. Dong, H. S. Yang, and F. L. Sun. Isogeometric symmetric FE– BE coupling method for acoustic-structural interaction. Applied Mathematics and Computation, 393(??): Article 125758, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307116.

Wang:2022:RFC

[WDZG22] Xianjia Wang, Rui Ding, Jinhua Zhao, and Cuiling Gu. The rise and fall of cooperation in populations with multiple groups. Applied Mathematics and Computation, 413 (??):Article 126624, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007086.

Wang:2024:ELE

[WDZL24] Chengjie Wang, Juan Deng, Hui Zhao, and Li Li. Effect of Q-learning on the evolution of cooperation behavior in collective motion: an improved Vicsek model. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400417X.

Wen:2020:GBM

[Wen20] Zhenshu Wen. The generalized bifurcation method for deriving exact solutions of nonlinear spacetime fractional partial differential equations. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307271.

Wettergren:2021:GBM

[Wet21a] Thomas A. Wettergren. Game-based modeling of independent searchers who share a common goal. Applied Mathematics and Computation, 391(??):Article 125621, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305750.

Wettergren:2021:RDP

[Wet21b] Thomas A. Wettergren. Replicator dynamics of an n-player snowdrift game with delayed payoffs. Applied Mathematics and Computation, 404(??):Article 126204, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002940.

Wettergren:2023:RDE

[Wet23] Thomas A. Wettergren.

Replicator dynamics of evolutionary games with different delays on costs and benefits. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003971.

Wang:2023:PST

[WF23a] Shixin Wang and Tao Feng.

Perfect state transfer on weighted bi-Cayley graphs over abelian groups. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001753.

Wei:2023:GRH

[WF23b] Wei Wei and Xiangnan Feng. Graphical representation and hierarchical decomposition mechanism for vertex-cover solution space. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004332.

Wang:2020:LMI

[WFHG20] Fajie Wang, Chia-Ming Fan, Qingsong Hua, and Yan Gu. Localized MFS for the inverse Cauchy problems of two-dimensional Laplace and biharmonic equations. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306502.

Wang:2024:ZSG

[WFKL24] Yun Wang, Tian Fang, Qingkai Kong, and Feng Li. Zero-sum gamebased optimal control for discrete-time Markov jump systems: a parallel offpolicy Q-learning method. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006318.

Wang:2021:DET

[WFYL21] Jiaqi Wang, Fang Fang, Xiaojian Yi, and Yajuan Liu. Dynamic event-triggered fault estimation and sliding mode fault-tolerant control for networked control systems with sensor faults. Applied Mathematics and Computation, 389 (??):Article 125558, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305142.

Wu:2021:SPT

[WGJK21] Zhengtian Wu, Qing Gao, Baoping Jiang, and Hamid Reza Karimi. Solving the production transportation problem via a deterministic annealing neural network method. Applied Mathematics and Computation, 411(??):Article 126518, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100607X.

Wang:2021:FTS

[WGL^{+}21] Yao Wang, Jun Guo, Guobao Liu, Junwei Lu, and Fangyuan Li. Finite-time sampled-data synchronization for uncertain neutraltype semi-Markovian jump neural networks with mixed time-varying delays. Applied Mathematics and Computation, 403(??):Article 126197, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002873.

Wu:2020:NFV

[WGLL20] Fei Wu, Renbo Gao, Jie Liu, and Cunbao Li. New fractional variable-order creep model with short memory. Applied Mathematics and Computation, 380(??): Article 125278, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302472.

Wang:2020:WGF

[WGLS20] Xiuping Wang, Fuzheng Gao, Yang Liu, and Zhengjia Sun. A weak Galerkin finite element method for high dimensional timefractional diffusion equation. Applied Mathematics and Computation, 386 (??):Article 125524, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304823.

Wang:2021:CIA

[WGM^{+}21] Chunpeng Wang, Hongling Gao, Bin Ma, Meihong Yang, Jian Li, Zhiqiu Xia, and Qixian Hao. Color image analysis with novel trinion decimal-order polar harmonic transforms. Applied Mathematics and Computation, 410(??):Article 126453, December 1, 2021.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005427.

Wu:2023:SBM

[WGMJ23] Fuqiang Wu, Yitong Guo, Jun Ma, and Wuyin Jin. Synchronization of bursting memristive Josephson junctions via resistive and magnetic coupling. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003004.

Wang:2022:SOQ

[WGSS22] Zhi Wang, Yongbin Ge, Hai-Wei Sun, and Tao Sun. Sixth-order quasicompact difference schemes for 2D and 3D Helmholtz equations. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004210.

Wang:2021:AFD

[WGW^{+}21] Weiping Wang, Junjiang Guo, Zhen Wang, Hao Wang, Jun Cheng, Chunyang Wang, Manman Yuan, Jürgen Kurths, Xiong Luo, and Yang Gao. Abnormal flow detection in industrial control network based on deep reinforcement learning. Applied Mathematics and Computation, 409 (??):Article 126379, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004689.

Wang:2022:DOB

[WGZZ22] Fang Wang, Yali Gao, Chao Zhou, and Qun Zong. Disturbance observerbased backstepping formation control of multiple quadrotors with asymmetric output error constraints. Applied Mathematics and Computation, 415 (??):Article 126693, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007773.

Wen:2021:CMN

[WH21] Xiaoxia Wen and Jin Huang. A combination method for numerical solution of the nonlinear stochastic Itô–Volterra integral equation. Applied Mathematics and Computation, 407(??):Article 126302, October 15, 2021.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100391X.

Wan:2020:SAN

[WHC20a] Fangzhe Wan, Po Hu, and Huabin Chen. Stability analysis of neutral stochastic differential delay equations driven by Lévy noises. Applied Mathematics and Computation, 375(??): Article 125080, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300497.

Wei:2020:TDC

[WHC20b] Chao Wei, Rong-Xia Hao, and Jou-Ming Chang. Twodisjoint-cycle-cover bipancyclicity of balanced hypercubes. Applied Mathematics and Computation, 381(??): Article 125305, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030271X.

Wu:2020:SDH

[WHCW20] Tianyu Wu, Xia Huang, Xiangyong Chen, and Jing Wang. Sampled-data H_{∞} exponential synchronization for delayed semi-Markov jump CDNs: a looped-functional approach. Applied Mathematics and Computation, 377(??):Article 125156, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301259.

Wen:2023:SMS

[WHD^{+}23] Jianghui Wen, Lijiang Hong, Min Dai, Xinping Xiao, and Chaozhong Wu. A stochastic model for stop-and-go phenomenon in traffic oscillation: On the prospective of macro and micro traffic flow. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007093.

Wang:2020:VPP

[WHF20] Yang Wang, Danjun Huang, and Stephen Finbow. On the vertex partition of planar graphs into forests with bounded degree. Applied Mathematics and Computation, 374(??):Article 125032, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320300011.

Wang:2020:BFP

[WHH^{+}20] Wei Wang, Chi Huang, Chuangxia Huang, Jinde Cao, Jianquan Lu, and Li Wang. Bipartite formation problem of secondorder nonlinear multi-agent systems with hybrid impulses. Applied Mathematics and Computation, 370 (??):??, April 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031930918X.

Wang:2021:TSI

[WHHD21] Jin-Ju Wang, Ting-Zhu Huang, Jie Huang, and Liang-Jian Deng. A twostep iterative algorithm for sparse hyperspectral unmixing via total variation. Applied Mathematics and Computation, 401 (??):Article 126059, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001077.

Wu:2024:MIR

[WHJ^{+}24] Tingting Wu, Chaoyan Huang, Shilong Jia, Wei Li, Raymond Chan, Tieyong Zeng, and S. Kevin Zhou. Medical image reconstruction with multilevel deep learning denoiser and tight frame regularization. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400256X.

Wang:2023:FEE

[WHJK23] Jian Wang, Ziwei Han, Wenjing Jiang, and Junseok Kim. A fast, efficient, and explicit phase-field model for 3D mesh denoising. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004083.

Wang:2023:CSMb

[WHKT23] Yiqiao Wang, Jingjing Huo, Jiangxu Kong, and Qiuyue Tan. Coloring the square of maximal planar graphs with diameter two. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004320.

Wang:2021:SAG

[WHL21] Chen-Rui Wang, Yong He, and Wen-Juan Lin. Stability analysis of generalized neural networks with fast-varying delay via a relaxed negativedetermination quadratic function method. Applied Mathematics and Computation, 391(??):Article 125631, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305853.

Wu:2022:FTS

[WHL22] Jie Wu, Xinyi He, and Xiaodi Li. Finite-time stabilization of time-varying nonlinear systems based on a novel differential inequality approach. Applied Mathematics and Computation, 420(??):Article 126895, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009784.

Wang:2020:FFD

[WHLL20] Jiancheng Wang, Shuping He, Xiaoli Luan, and Fei Liu. Fuzzy fault detection of conic-type nonlinear systems within the finite frequency domain. Applied Mathematics and Computation, 378 (??):Article 125181, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301508.

Wang:2021:CFO

[WHM^{+}21a] Chunpeng Wang, Qixian Hao, Bin Ma, Jian Li, and Hongling Gao. Corrigendum to “Fractional-order quaternion exponential moments for color images” [Applied Mathematics and Computation 400 (2021) 126061]. Applied Mathematics and Computation, 402(??):Article 126196, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002861.

See [WHM^{+}21b].

Wang:2021:FOQ

[WHM^{+}21b] Chunpeng Wang, Qixian Hao, Bin Ma, Jian Li, and Hongling Gao. Fractionalorder quaternion exponential moments for color images. Applied Mathematics and Computation, 400(??): Article 126061, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321001090. See

corrigendum [WHM^{+}21a].

Wu:2022:EEM

[WHM22] Fuqiang Wu, Xikui Hu, and Jun Ma. Estimation of the effect of magnetic field on a memristive neuron. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004404.

Wang:2023:SCT

[WHPQ23] Yibo Wang, Changchun Hua, PooGyeon Park, and Cheng Qian. Stability criteria for time-varying delay systems via an improved reciprocally convex inequality lemma. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000875.

Wang:2021:DAS

[WHW^{+}21] Weiping Wang, Chang He, Zhen Wang, Alexander Hramov, Denggui Fan, Manman Yuan, Xiong Luo, and Jürgen Kurths. Dynamic analysis of synaptic loss and synaptic compensation in the process of associative memory ability decline in Alzheimer’s disease. Applied Mathematics and Computation, 408 (??):Article 126372, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004616.

Wei:2021:CSP

[WHWW21] Chunqiu Wei, Lin He, Huibin Wu, and Hairui Wen. A class of structurepreserving discontinuous Galerkin variational time integrators for Birkhoffian systems. Applied Mathematics and Computation, 393(??):Article 125750, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307037.

Wu:2020:SSO

[WHY20] Chao Wu, Yunqing Huang, and Jinyun Yuan. Superconvergence of the second order cubic edge elements with Maxwell’s equations. Applied Mathematics and Computation, 387(??): Article 124410, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S009630031930356X.

Wu:2022:SNS

[WHY22] Hao Wu, Junhao Hu, and Chenggui Yuan. Stability of numerical solution to pantograph stochastic functional differential equations. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004003.

Wang:2020:CSI

[WHZ^{+}20] Yugang Wang, Ting-Zhu Huang, Xi-Le Zhao, Liang-Jian Deng, and Teng-Yu Ji. A convex single image dehazing model via sparse dark channel prior. Applied Mathematics and Computation, 375(??): Article 125085, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300540.

Wang:2020:TDM

[WJ20] Xiaolong Wang and Yaolin Jiang. Time domain model reduction of timedelay systems via orthogonal polynomial expansions. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308082.

Wang:2020:DFB

[WJC^{+}20] Mei Wang, Baoguo Jia, Churong Chen, Xiaojuan Zhu, and Feifei Du. Discrete fractional Bihari inequality and uniqueness theorem of solutions of nabla fractional difference equations with non-Lipschitz nonlinearities. Applied Mathematics and Computation, 376 (??):Article 125118, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300874.

Wang:2024:CFT

[WJDL24] Xi Wang, Yamei Ju, Derui Ding, and Hongjian Liu. Cooperative fault-tolerant tracking control for multiagent systems: a multiple description encoding scheme. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005064.

Wang:2020:VBC

[WJG^{+}20] Xinyu Wang, Danyang Jia, Shupeng Gao, Chengyi Xia, Xuelong Li, and Zhen Wang. Vaccination behavior by coupling the epidemic spreading with the human decision under the game theory. Applied Mathematics and Computation, 380(??): Article 125232, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302010.

Wu:2020:LDD

[WJK20] Zhengtian Wu, Baoping Jiang, and Hamid Reza Karimi. A logarithmic descent direction algorithm for the quadratic knapsack problem. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930846X.

Wang:2022:QAR

[WJL22] Jingjing Wang, Yunfeng Jia, and Fangfang Li. Qualitative analysis on a reaction– diffusion model arising from population dynamics. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002776.

Wang:2024:SCF

[WJL^{+}24] Anqing Wang, Lei Ju, Lu Liu, Haoliang Wang, Nan Gu, Zhouhua Peng, and Dan Wang. Safetycritical formation control of switching uncertain Euler– Lagrange systems with control barrier functions. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003217.

Wang:2021:VDC

[WJY21] Yun Wang, Suyun Jiang, and Jin Yan. Vertex disjoint copies of K_{1,4} in claw-free graphs. Applied Mathematics and Computation, 393(??):Article 125768, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307219.

Wang:2024:EAN

[WKM24] Hong Wang, Hongchao Kang, and Junjie Ma. An efficient and accurate numerical method for the Bessel transform with an irregular oscillator and its error analysis. Applied Mathematics and Computation, 473(??): ??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001437.

Wang:2020:EWI

[WL20] Guangfu Wang and Yajing Liu. The edge-Wiener index of zigzag nanotubes. Applied Mathematics and Computation, 377 (??):Article 125191, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301600.

Wang:2021:SMI

[WL21a] Hongxing Wang and Xiaoji Liu. Solutions of the matrix inequality AXA ≤ ?A in some partial orders. Applied Mathematics and Computation, 396(??): Article 125940, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308936.

Wang:2021:MDD

[WL21b] Jin-Liang Wang and Hui-Feng Li. Memory-dependent derivative versus fractional derivative (II): Remodelling diffusion process. Applied Mathematics and Computation, 391(??):Article 125627, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320305816.

Wei:2022:ESC

[WL22] Wei Wei and Shuchao Li. On the eccentricity spectra of complete multipartite graphs. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001229.

Wang:2023:CAT

[WL23] Xianfen Wang and Jiyong Li. Convergence analysis of two conservative finite difference Fourier pseudospectral schemes for Klein– Gordon–Dirac system. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007068.

Wang:2024:RBS

[WL24a] Guanjie Wang and Qifeng Liao. Reduced basis stochastic Galerkin methods for partial differential equations with random inputs. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005441.

Wu:2024:RPI

[WL24b] Nian-Ci Wu and Cheng-Zhi Liu. Randomized progressive iterative approximation for B-spline curve and surface fittings. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001413.

Wang:2022:AEF

[WLB22] Zhi-Bo Wang, Da-Yan Liu, and Driss Boutat. Algebraic estimation for fractional integrals of noisy acceleration based on the behaviour of fractional derivatives at zero. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003289.

Wang:2022:FTO

[WLD^{+}22] Di Wang, Can Liu, Dawei Ding, Suixiang Gao, and Ming Chu. Finite-time optimal tracking control using augmented error system method. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000996.

Wang:2022:RSS

[WLFX22] Jianjun Wang, Wen Liu, Shihua Fu, and Jianwei Xia. On robust set stability and set stabilization of probabilistic Boolean control networks. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000789.

Wang:2021:NGD

[WLG^{+}21] Lianwen Wang, Zhijun Liu, Caihong Guo, Yong Li, and Xinan Zhang. New global dynamical results and application of several SVEIS epidemic models with temporary immunity. Applied Mathematics and Computation, 390 (??):Article 125648, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306020.

Wang:2023:MPS

[WLHW23] Xiaomin Wang, Feng Li, Xingliu Hu, and Jing Wang. Mixed H_{∞} /passive synchronization for persistent dwell-time switched neural networks via an activation function dividing method. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200786X.

Wang:2021:EDI

[WLJ^{+}21] Zhen Wang, Chaofan Li, Xing Jin, Hong Ding, Guanghai Cui, and Lanping Yu. Evolutionary dynamics of the interdependent security games on complex network. Applied Mathematics and Computation, 399(??): Article 126051, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000990.

Wang:2022:CGC

[WLJ22] Yongwei Wang, Weijun Liu, and Wei Jin. The commuting graph of the cyclic extension of an abelian group. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002132.

Wang:2021:BRN

[WLL21] Ye Wang, Yusheng Li, and Yan Li. Bipartite Ramsey numbers of K_{t,s} in many colors. Applied Mathematics and Computation, 404(??): Article 126220, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003106.

Wang:2022:GCP

[WLL^{+}22] Wei Wang, Wenyao Li, Tao Lin, Tao Wu, Liming Pan, and Yanbing Liu. Generalized k-core percolation on higher-order dependent networks. Applied Mathematics and Computation, 420(??): Article 126793, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008754.

Wang:2024:AAT

[WLL^{+}24] Lanxin Wang, Yue Long, Tieshan Li, Hanqing Yang, and C. L. Philip Chen. Asynchronous attack tolerant control for Markov jump cyber-physical systems under hybrid cyber-attacks. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000559.

Wang:2022:IFS

[WLPZ22] Xinwei Wang, Jie Liu, Haijun Peng, and Xudong Zhao. An iterative framework to solve nonlinear optimal control with proportional delay using successive convexification and symplectic multi-interval pseudospectral scheme. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005227.

Wang:2020:CTN

[WLQX20] Yingchun Wang, Haifeng Li, Xiaojie Qiu, and Xiangpeng Xie. Consensus tracking for nonlinear multiagent systems with unknown disturbance by using model free adaptive iterative learning control. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306939.

Wang:2022:OCA

[WLS22] Lingling Wang, Shaoyong Lai, and Rongmei Sun. Optimal control about multiagent wealth exchange and decision-making competence. Applied Mathematics and Computation, 417(??): Article 126772, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008547.

Wang:2020:NAC

[WLW20] Fang Wang, Lishan Liu, and Yonghong Wu. A numerical algorithm for a class of fractional BVPs with p-Laplacian operator and singularity — the convergence and dependence analysis. Applied Mathematics and Computation, 382 (??):Article 125339, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303052.

Wang:2022:BNF

[WLWL22] Yupin Wang, Xiaodi Li, Da Wang, and Shutang Liu. A brief note on fractal dynamics of fractional Mandelbrot sets. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004271.

Wang:2020:WLP

[WLY20] Carol J. Wang, Junjie Li, and Pengjun Yin. Wave linked partitions and 312avoiding permutations with primacy being 1. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307659.

Wang:2021:PSP

[WLY21] Lei Wang, Qingdu Li, and Xiao-Song Yang. Periodic sinks and periodic saddle orbits induced by heteroclinic bifurcation in three-dimensional piecewise linear systems with two zones. Applied Mathematics and Computation, 404(??): Article 126200, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002903.

Wang:2021:FNS

[WLZ21] Xiao Wang, You Lu, and Shenggui Zhang. Flow number of signed Halin graphs. Applied Mathematics and Computation, 393 (??):Article 125751, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307049.

Wei:2021:CGD

[WLZD21] Liang Wei, Faxu Li, Haixing Zhao, and Bo Deng. The controllability of graphs with diameter n−2. Applied Mathematics and Computation, 407(??):Article 126327, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321004161.

Wei:2024:MDN

[WLZD24] Liang Wei, Faxu Li, Haixing Zhao, and Bo Deng. On the minimum driver node set of k-uniform linear hypertree networks. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001747.

Wang:2021:ETN

[WLZF21] Jinling Wang, Jinling Liang, Cheng-Tang Zhang, and Dongmei Fan. Eventtriggered non-fragile control for uncertain positive Roesser model with PDT switching mechanism. Applied Mathematics and Computation, 406 (??):Article 126266, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003556.

Wang:2021:SRA

[WLZW21] Si-Yi Wang, Yan-Ping Liu, Feng Zhang, and Rui-Wu Wang. Super-rational aspiration induced strategy updating promotes cooperation in the asymmetric prisoner’s dilemma game. Applied Mathematics and Computation, 403 (??):Article 126180, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002708.

Wei:2021:IIS

[WLZY21] Tengda Wei, Ping Lin, Quanxin Zhu, and Qi Yao. Instability of impulsive stochastic systems with application to image encryption. Applied Mathematics and Computation, 402 (??):Article 126098, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001466.

Wang:2022:EDE

[WM22] Zhiwen Wang and Xianhao Meng. Effects on distance energy of complete bipartite graphs by embedding edges. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003496.

Wang:2023:ANF

[WM23] Xingxing Wang and Yuechao Ma. Adaptive non-fragile sliding mode control for switched semi-Markov jump system with time-delay and attack via reduced-order method. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S009630032200738X.

Wang:2024:FTA

[WM24a] Huanqing Wang and Zhu Meng. Fixed-time adaptive neural tracking control for high-order nonlinear switched systems with input saturation and dead-zone. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003655.

Wei:2024:MRP

[WM24b] Jiangchao Wei and Changfeng Ma. A modified RBULT preconditioner for generalized saddle point problems from the hydrodynamic equations. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400287X.

Wang:2021:EAI

[WMCZ21] Huan Wang, Chuang Ma, Han-Shuang Chen, and Hai-Feng Zhang. Effects of asymptomatic infection and self-initiated awareness on the coupled diseaseawareness dynamics in multiplex networks. Applied Mathematics and Computation, 400(??):Article 126084, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001326.

Wang:2023:RTK

[WMF23] Yihong Wang, Jixiang Meng, and Jianxi Fan. Reliabilities for two kinds of graphs with smaller diameters. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000437.

Wang:2021:ASC

[WMQ21] Pan Wang, Tian-Ping Ma, and Feng-Hua Qi. Analytical solutions for the coupled Hirota equations in the firebringent [sic: birefringent] fiber. Applied Mathematics and Computation, 411(??):Article 126495, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005841.

Wang:2022:EDH

[WMS^{+}22] Haiying Wang, Jack Murdoch Moore, Michael Small, Jun Wang, Huijie Yang, and Changgui Gu. Epidemic dynamics on higherdimensional small world networks. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009942.

Wang:2022:REM

[WMT22] Na Wang, Jixiang Meng, and Yingzhi Tian. Reliability evaluation of modified bubble-sort graph networks based on structure fault pattern. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003319.

Wang:2021:DRI

[WMWS21] Haiying Wang, Jack Murdoch Moore, Jun Wang, and Michael Small. The distinct roles of initial transmission and retransmission in the persistence of knowledge in complex networks. Applied Mathematics and Computation, 392(??): Article 125730, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306834.

Wang:2020:GAO

[WNG20] Gang Wang, Min-Yao Niu, and You Gao. Generalized asymptotically optimal codebooks. Applied Mathematics and Computation, 379(??):Article 125230, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301995.

Wang:2020:SRA

[WNP20] Xiao Wang, Yufu Ning, and Zhen Peng. Some results about uncertain differential equations with timedependent delay. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307398.

Wu:2021:SRU

[WNZ21] Tingting Wu, Michael K. Ng, and Xi-Le Zhao. Sparsity reconstruction using nonconvex TGpV-shearlet regularization and constrained projection. Applied Mathematics and Computation, 410(??):Article 126170, December 1, 2021.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002605.

Woldegerima:2021:MAI

[WOB21] Woldegebriel Assefa Woldegerima, Rachid Ouifki, and Jacek Banasiak. Mathematical analysis of the impact of transmission-blocking drugs on the population dynamics of malaria. Applied Mathematics and Computation, 400(??):Article 126005, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000539.

Woods:2020:CRS

[Woo20] R. C. Woods. Comments on “Relativistic symmetries in Yukawa-type interactions with Coulomb-like tensor”. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306678.

See [IHF13].

Wu:2020:FHE

[WP20] Cixuan Wu and Jiangmin Pan. Finite hexavalent edgeprimitive graphs. Applied Mathematics and Computation, 378(??):Article 125207, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301764.

Wang:2021:ECS

[WP21a] Xiaofeng Wang and Matjaz Perc. Emergence of cooperation in spatial social dilemmas with expulsion. Applied Mathematics and Computation, 402 (??):Article 126149, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001971.

Wongsaijai:2021:ODR

[WP21b] Ben Wongsaijai and Kanyuta Poochinapan. Optimal decay rates of the dissipative shallow water waves modeled by coupling the Rosenau–RLW equation and the Rosenau–Burgers equation with power of nonlinearity. Applied Mathematics and Computation, 405(??):Article 126202, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002927.

Weijermars:2020:LSM

[WPE20] Ruud Weijermars, Tri Pham, and Mahmood Ettehad. Linear superposition method (LSM) for solving stress tensor fields and displacement vector fields: Application to multiple pressure-loaded circular holes in an elastic plate with far-field stress. Applied Mathematics and Computation, 381(??): Article 125234, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302034.

Wang:2020:IBA

[WPH20] Mengyao Wang, Qiuhui Pan, and Mingfeng He. The interplay of behaviors and attitudes in public goods game considering environmental investment. Applied Mathematics and Computation, 382 (??):Article 125250, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302198.

Wu:2022:TAT

[WPJ22] B. Wu, D. Peng, and R. Jones. On thermoelastic analysis of two collinear cracks subject to combined quadratic thermomechanical load. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321009887.

Wu:2020:ETA

[WPX^{+}20] Li-Bing Wu, Ju H. Park, Xiang-Peng Xie, Ya-Juan Liu, and Zhi-Chun Yang. Event-triggered adaptive asymptotic tracking control of uncertain nonlinear systems with unknown deadzone constraints. Applied Mathematics and Computation, 386(??):Article 125528, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304732.

Wang:2023:AET

[WPZY23] Qiyi Wang, Li Peng, Huarong Zhao, and Shenhao Yang. Adaptive event-triggered filtering for semi-Markov jump systems under communication constraints. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004447.

Wang:2024:RCP

[WQT^{+}24] Chen Wang, Yiwen Qi, Yiwen Tang, Xin Li, and Ming Ji. Robust control with protected feedback information for switched systems under injection attacks. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001863.

Wang:2020:SSP

[WQZ^{+}20] Huajian Wang, Wenhai Qi, Lihua Zhang, Jun Cheng, and Yonggui Kao. Stability and stabilization for positive systems with semi-Markov switching. Applied Mathematics and Computation, 379(??):Article 125252, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302216.

Wang:2023:ETI

[WQZD23] Ning-Ning Wang, Shui-Han Qiu, Xiao Wen Zhong, and Zeng-Ru Di. Epidemic thresholds identification of susceptible-infected-recovered model based on the Eigen Microstate. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000930.

Wright:2024:NCM

[Wri24] Stephen E. Wright. A note on computing maximum likelihood estimates for the three-parameter asymmetric Laplace distribution. Applied Mathematics and Computation, 464(??):??, March 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005507.

Wang:2020:CMF

[WS20] Xiaoling Wang and Housheng Su. Completely modelfree RL-based consensus of continuous-time multi-agent systems. Applied Mathematics and Computation, 382(??):Article 125312, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302782.

Wu:2021:QUU

[WS21] Yanmi Wu and Dongyang Shi. Quasi-uniform and unconditional superconvergence analysis of Ciarlet– Raviart scheme for the fourth order singularly perturbed bi-wave problem modeling d-wave superconductors. Applied Mathematics and Computation, 397 (??):Article 125924, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308778.

Wang:2022:IGS

[WS22] Chaoqian Wang and Attila Szolnoki. Involution game with spatiotemporal heterogeneity of social resources. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003812.

Wang:2023:ISP

[WS23a] Chaoqian Wang and Attila Szolnoki. Inertia in spatial public goods games under weak selection. Applied Mathematics and Computation, 449(??): ??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001108.

Wang:2023:MFO

[WS23b] Xianming Wang and Mouquan Shen. Model free optimal control of unknown nonaffine nonlinear systems with input quantization and DoS attack. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000838.

Wang:2024:SND

[WS24] Jun Wang and Yuefang Sun. Scattering number of digraphs. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006446.

Wen:2024:HCD

[WSC^{+}24] Chao Wen, Qiang Sun, Hongyan Cai, Huanhuan Guan, and Chao Zhang. The homomorphism and coloring of the direct product of signed graphs. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002480.

Wang:2024:DET

[WSDW24] Suzhen Wang, Yan Sheng, Chengyu Deng, and Yanqian Wang. Dynamic eventtriggered resilient controller design for discrete-time singular Markov jump CPSs under hybrid cyber-attacks. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006756.

Wu:2021:TSI

[WSG^{+}21] Tingting Wu, Jinbo Shao, Xiaoyu Gu, Michael K. Ng, and Tieyong Zeng. Twostage image segmentation based on nonconvex ℓ_{2}–ℓ_{p} approximation and thresholding. Applied Mathematics and Computation, 403(??):Article 126168, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002587.

Wu:2021:EDF

[WSL^{+}21] Jiacheng Wu, Lei Su, Shaoming Li, Jing Wang, and Xiangyong Chen. Extended dissipative filtering for singularly perturbed systems with random uncertain measurement: a doublelayer switching mechanism. Applied Mathematics and Computation, 410 (??):Article 126465, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005543.

Wang:2022:AAT

[WSS22] Le Wang, Wei Sun, and Shun-Feng Su. Adaptive asymptotic tracking control for nonlinear systems with state constraints and input saturation. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004167.

Wu:2022:NBA

[WSSX22] Jing Wu, Wei Sun, Shun-Feng Su, and Jianwei Xia. Neural-based adaptive control for nonlinear systems with quantized input and the output constraint. Applied Mathematics and Computation, 413 (??):Article 126637, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007219.

Wu:2022:IAF

[WSXL22] Dawei Wu, Yonghui Sun, Rongsheng Xia, and Shumin Lu. Improved adaptive fuzzy control for non-strict feedback nonlinear systems: a dynamic compensation system approach. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005446.

Wang:2021:FVS

[WSY21] Jiangfu Wang, Zhiqiang Sheng, and Guangwei Yuan. A finite volume scheme preserving maximum principle with cell-centered and vertex unknowns for diffusion equations on distorted meshes. Applied Mathematics and Computation, 398 (??):Article 125989, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000370.

Wu:2020:NIA

[WSZ20] Ai-Guo Wu, Hui-Jie Sun, and Ying Zhang. A novel iterative algorithm for solving coupled Riccati equations. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930637X.

Wu:2022:CFE

[WSZ22] Rongfang Wu, Xiaoqin Shen, and Jikun Zhao. Conforming finite element methods for two-dimensional linearly elastic shallow shell and clamped plate models. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003332.

Wu:2020:AKB

[WTCC20] Mu-En Wu, Hui-Huang Tsai, Wei-Ho Chung, and Chien-Ming Chen. Analysis of Kelly betting on finite repeated games. Applied Mathematics and Computation, 373(??):Article 125028, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319310203.

Wu:2020:NSC

[Wu20a] Hongwu Wu. A necessary and sufficient condition for the oscillation of first order linear difference equations with several delay arguments. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308513.

Wu:2020:LBF

[Wu20b]  Junru Wu.  On a lin-

earity between fractal dimension and order of fractional calculus in Hölder space. Applied Mathematics and Computation, 385 (??):Article 125433, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303945.

Wang:2020:CED

[WW20] Yuan-Ming Wang and Xin Wen. A compact exponential difference method for multi-term time-fractional convection-reaction-diffusion problems with non-smooth solutions. Applied Mathematics and Computation, 381(??):Article 125316, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302824.

Wang:2022:BPD

[WW22a] Xinhe Wang and Zhen Wang. Bifurcation and propagation dynamics of a discrete pair SIS epidemic model on networks with correlation coefficient. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322005513.

Willmot:2022:RGI

[WW22b] Gordon E. Willmot and Jae-Kyung Woo. Remarks on a generalized inverse Gaussian type integral with applications. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003769.

Wang:2023:IST

[WW23] Tao Wang and Baoyindureng Wu. Induced subgraphs of a tree with constraint degree. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007287.

Wu:2024:SCM

[WWK24] Jiadong Wu, Jing Wang, and Liying Kang. Spectral conditions for matching extension. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004430.

Wang:2022:NTL

[WWQ22] Zhengdao Wang, Yikun Wei, and Yuehong Qian. A novel thermal lattice Boltzmann model with heat source and its application in incompressible flow. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002429.

Wei:2024:SAT

[WWT24] Yangchun Wei, Haoping Wang, and Yang Tian. Shifting asymmetric timevarying BLF-based modelfree hybrid force/position control for 3-DOF SEAbased manipulator with random initial error. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005325.

Wang:2021:NCB

[WWW^{+}21] Weiping Wang, Chunyang Wang, Zhen Wang, Baijing Han, Chang He, Jun Cheng, Xiong Luo, Manman Yuan, and Jürgen Kurths. Nonlinear consensus-based autonomous vehicle platoon control under eventtriggered strategy in the presence of time delays. Applied Mathematics and Computation, 404(??):Article 126246, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003362.

Wang:2022:ADT

[WWW^{+}22] Weiping Wang, Chunyang Wang, Zhen Wang, Manman Yuan, Xiong Luo, Jürgen Kurths, and Yang Gao. Abnormal detection technology of industrial control system based on transfer learning. Applied Mathematics and Computation, 412(??):Article 126539, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006238.

Wang:2024:OSE

[WWW24a] Ziyun Wang, Xianzhe Wang, and Yan Wang. Orthotopesearch-expansion-based extended zonotopic Kalman filter design for a discretetime linear parametervarying system with a dual-noise term. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001474.

Wu:2024:ETA

[WWW24b] Meng Wu, Li-Bing Wu, and Pu Wang. Eventtriggered adaptive leaderless consensus control for nonlinear multi-agent systems with unknown dead-zones and output constraints. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006926.

Wang:2021:IDS

[WWWH21] Huamin Wang, Guoliang Wei, Shiping Wen, and Tingwen Huang. Impulsive disturbance on stability analysis of delayed quaternion-valued neural networks. Applied Mathematics and Computation, 390(??):Article 125680, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306330.

Wang:2024:PPG

[WWWK24] Yang Wang, Jiali Wang, Weifan Wang, and Jiangxu Kong. Partitioning planar graphs into bounded degree forests. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001772.

Wang:2021:NRO

[WWWZ21] Junlan Wang, Xin Wang, Yantao Wang, and Xian Zhang. Non-reduced order method to global h-stability criteria for proportional delay high-order inertial neural networks. Applied Mathematics and Computation, 407(??):Article 126308, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003970.

Wang:2020:LCC

[WWY^{+}20] Jianwei Wang, Rong Wang, Fengyuan Yu, Ziwei Wang, and Qiaochu Li. Learning continuous and consistent strategy promotes cooperation in prisoner’s dilemma game with mixed strategy. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308793.

Wang:2021:LQC

[WWY21] Guangchen Wang, Wencan Wang, and Zhiguo Yan. Linear quadratic control of backward stochastic differential equation with partial information. Applied Mathematics and Computation, 403(??):Article 126164, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100254X.

Wu:2022:LCM

[WWY22] Jian Wu, Li Wang, and Weihua Yang. Learning to compute the metric dimension of graphs. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004246.

Wang:2024:PGD

[WWY24] Tao Wang, Ya-Nan Wang, and Xiaojing Yang. Planar graphs with distance of 3-cycles at least 2 and no cycles of lengths 5, 6, 7. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004077.

Wang:2020:HGD

[WWZ20a] Chao Wang, Hua Wang, and Yuzhi Zhang. The r-Hamming gap and distancegap-preserving mappings from binary vectors to permutations. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306885.

Wang:2020:PDP

[WWZ20b] Yufang Wang, Haiyan Wang, and Shuhua Zhang. Prediction of daily pm _{2.5} concentration in China using data-driven ordinary differential equations. Applied Mathematics and Computation, 375(??):Article 125088, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300576.

Wang:2023:CRR

[WWZ^{+}23] Chengxiang Wang, Xiaoyan Wang, Kequan Zhao, Min Huang, Xianyun Li, and Wei Yu. A cascading ℓ_{0} regularization reconstruction method in nonsubsampled contourlet domain for limited-angle CT. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001820.

Wang:2023:ECD

[WWZW23] Si-Yi Wang, Qing-Lian Wang, Xiao-Wei Zhang, and Rui-Wu Wang. Evolutionary cooperation dynamics of combining imitation and super-rational aspiration induced strategy updating. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300303X.

Wang:2020:ULH

[WX20a] Wen-Huan Wang and Yi-Sai Xue. On the runiform linear hypertrees with extremal Estrada indices. Applied Mathematics and Computation, 377 (??):Article 125144, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301132.

Xu:2020:NAS

[wX20b] Wei wei Xu. Note on analytic solutions of a class of extended constrained matrix maximization problems. Applied Mathematics and Computation, 376(??): Article 125120, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300898.

Wu:2021:RES

[WX21a] Ailong Wu and Xiangru Xing. Robust exponential stabilization of positive uncertain switched neural networks with actuator saturation and sensor faults. Applied Mathematics and Computation, 411 (??):Article 126548, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006329.

Wu:2021:NCP

[WX21b] Hongyu Wu and Shuhuang Xiang. A new constraint preconditioner based on the PGSS iteration method for non-Hermitian generalized saddle point problems. Applied Mathematics and Computation, 396(??): Article 125864, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308171.

Wu:2022:SDN

[WX22] Chufen Wu and Zhaoquan Xu. Spatial dynamics in a non-local lattice competition model under shifting ranges. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003435.

Wang:2022:MFF

[WXF22] Wei Wang, Xiangpeng Xie, and Changyang Feng.

Model-free finite-horizon optimal tracking control of discrete-time linear systems. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200474X.

Xu:2020:ASC

[wXhGZ20] Wei wei Xu, Zhi hao Ge, and Lei Zhu. Analytic solutions of a class of extended constrained matrix maximization problems. Applied Mathematics and Computation, 372(??): Article 124989, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309816.

Wang:2023:UES

[WXHZ23] Xiaofeng Wang, Wenlong Xue, Yong He, and Fu Zheng. Uniformly exponentially stable approximations for Timoshenko beams. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001972.

Wang:2023:FMM

[WXLY23] Rong Wang, Qiuyan Xu, Zhiyong Liu, and Jiye Yang. Fast meshfree methods for nonlinear radiation diffusion equation. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006452.

Wang:2021:AFT

[WXLZ21] Yuanqing Wang, Ning Xu, Yajuan Liu, and Xudong Zhao. Adaptive faulttolerant control for switched nonlinear systems based on command filter technique. Applied Mathematics and Computation, 392(??): Article 125725, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306780.

Wang:2021:OBR

[WXP21] Zhe Wang, Dingyu Xue, and Feng Pan. Observer-based robust control for singular switched fractional order systems subject to actuator saturation. Applied Mathematics and Computation, 411(??):Article 126538, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006226.

Wang:2020:RSE

[WXW^{+}20] Xuelian Wang, Jianwei Xia, Jing Wang, Zhen Wang, and Jian Wang. Reachable set estimation for Markov jump LPV systems with time delays. Applied Mathematics and Computation, 376(??): Article 125117, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300862.

Wang:2021:ANN

[WXW^{+}21] Sanxia Wang, Jianwei Xia, Xueliang Wang, Wenjing Yang, and Linqi Wang. Adaptive neural networks control for MIMO nonlinear systems with unmeasured states and unmodeled dynamics. Applied Mathematics and Computation, 408 (??):Article 126369, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004586.

Wang:2020:DFT

[WXWS20] Yudong Wang, Jianwei Xia, Zhen Wang, and Hao Shen. Design of a faulttolerant output-feedback controller for thickness control in cold rolling mills. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308331.

Wang:2022:DPG

[WY22] Bing Wang and Jianhua Yin. Degree powers of graphs without Bl, s. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005239.

Wang:2023:PBS

[WY23] Yu Wang and Zhiguo Yan. Pareto-based Stackelberg differential game for stochastic systems with multi-followers. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005860.

Wu:2024:GCM

[WY24] Zongmin Wu and Ran Yang.

The Gauss-cos model for the autocorrelation function of fertility rate. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003680.

Wang:2024:ECE

[WYCL24] Xianjia Wang, Zhipeng Yang, Guici Chen, and Yanli Liu. Enhancing cooperative evolution in spatial public goods game by particle swarm optimization based on exploration and q-learning. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000067.

Wang:2020:RSN

[WYD20] Jintao Wang, Lu Yang, and Jinqiao Duan. Recurrent solutions of a nonautonomous modified Swift–Hohenberg equation. Applied Mathematics and Computation, 379(??):Article 125270, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302393.

Wu:2023:SIP

[WYG23] Tingzeng Wu, Yong Yu, and Xing Gao. The second immanantal polynomials of Laplacian matrices of unicyclic graphs. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002072.

Wang:2022:FDP

[WYLL22] Yu Wang, Yujing Yang, Yang Liu, and Jungang Lou. Fault detection and pinning control of Boolean networks. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200306X.

Wang:2024:DAC

[WYLQ24] Xin Wang, DongSheng Yang, Weihua Li, and Jia Qin. Designing adaptive continuous event-triggered consensus protocol for nonlinear multi-agent systems with a nonautonomous leader. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002960.

Wang:2020:CSD

[WYLW20] Xueli Wang, Guoyong Yuan, Jun Liu, and Guangrui Wang. Control of spiral drift by using feedback signals from a circular measuring domain in oscillatory media. Applied Mathematics and Computation, 368(??): ??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307945.

Wu:2024:UDH

[WYLZ24] Shuying Wu, Sanling Yuan, Guijie Lan, and Tonghua Zhang. Understanding the dynamics of hepatitis B transmission: a stochastic model with vaccination and Ornstein– Uhlenbeck process. Applied Mathematics and Computation, 476(??):??, September 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002340.

Wu:2023:EBD

[WYS23] Shuang Wu, Dan Ye, and Xinfeng Shao. Event-based distributed resilient control strategy for microgrids subject to disturbances and hybrid attacks. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004423.

Wang:2020:IIT

[WYW^{+}20] Yongjie Wang, Zhouzhou Yao, Chao Wang, Jiale Ren, and Qiao Chen. The impact of intelligent transportation points system based on Elo rating on emergence of cooperation at Y intersection. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309154.

Wang:2024:PCT

[WYZ24] Bo Wang, Liping Yuan, and Tudor Zamfirescu. Poidgeconvexity in triangular lattices. Applied Mathematics and Computation, 474(??): ??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001280.

Wang:2022:OBD

[WYZX22] Yingchun Wang, Wei Yan, Huaguang Zhang, and Xiangpeng Xie. Observerbased dynamic event-triggered H_{∞} LFC for power systems under actuator saturation and deception attack. Applied Mathematics and Computation, 420(??): Article 126896, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321009796.

Wang:2020:CDZ

[WZ20a] Yiqiao Wang and Lina Zheng. Computation on the difference of Zagreb indices of maximal planar graphs with diameter two. Applied Mathematics and Computation, 377 (??):Article 125187, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301569.

Wu:2020:NHO

[WZ20b] Longyuan Wu and Shuying Zhai. A new high order ADI numerical difference formula for time-fractional convection-diffusion equation. Applied Mathematics and Computation, 387(??): Article 124564, ???? 2020.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319305478.

Wang:2024:LSN

[WZ24] Changxin Wang and Junxue Zhang. Linear saturation numbers of Berge-C_{3} and berge-C_{4}. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001577.

Wang:2024:EVE

[WZA^{+}24] Jingrui Wang, Huizhen Zhang, Tianbo An, Xing Jin, Chao Wang, Jian Zhao, and Zhen Wang. Effect of vaccine efficacy on vaccination behavior with adaptive perception. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000158.

Wang:2022:AMP

[WZC^{+}22] Xin Wang, Guangming Zhuang, Guoliang Chen, Qian Ma, and Junwei Lu. Asynchronous mixed H_{∞} and passive control for fuzzy singular delayed Markovian jump system via hidden Markovian model mechanism. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003277.

Wang:2021:LCC

[WZCF21] Fajie Wang, Qinghai Zhao, Zengtao Chen, and Chia-Ming Fan. Localized Chebyshev collocation method for solving elliptic partial differential equations in arbitrary 2D domains. Applied Mathematics and Computation, 397(??):Article 125903, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308560.

Wang:2021:SLU

[WZGL21] Jian Wang, Yuanguo Zhu, Yajing Gu, and Ziqiang Lu. Solutions of linear uncertain fractional order neutral differential equations. Applied Mathematics and Computation, 407 (??):Article 126323, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004124.

Wu:2022:FRS

[WZL22] Tingzeng Wu, Tian Zhou, and Huazhong Lü. Further results on the star degree of graphs. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001606.

Wei:2021:SAS

[WZLC21] Xiaofeng Wei, Ziye Zhang, Chong Lin, and Jian Chen. Synchronization and anti-synchronization for complex-valued inertial neural networks with timevarying delays. Applied Mathematics and Computation, 403(??):Article 126194, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321002848.

Wang:2022:CSN

[WZLH22] Xin Wang, Shidong Zhai, Guoqiang Luo, and Tao Huang. Cluster synchronization in a network of nonlinear systems with directed topology and competitive relationships. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000170.

Wen:2022:CTD

[WZN22] You-Wei Wen, Mingchao Zhao, and Michael Ng. Cartoon and texture decomposition for color image in opponent color space. Applied Mathematics and Computation, 414 (??):Article 126654, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007384.

Wang:2024:MPC

[WZPL24] Yong Wang, Jie Zhong, Qinyao Pan, and Ning Li. Minimal pinning control for set stability of Boolean networks. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006021.

Wu:2021:ECM

[WZW^{+}21] Yu’e Wu, Zhipeng Zhang, Xinyu Wang, Ming Yan, Qingfeng Zhang, and Shuhua Zhang. Evolution of cooperation in the multigame on a two-layer square network. Applied Mathematics and Computation, 400(??): Article 126088, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001363.

Wang:2023:FTL

[WZW23] Boyu Wang, Yijun Zhang, and Miao Wei. Fixed-time leader-following consensus of multi-agent systems with intermittent control. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005987.

Wang:2024:TLM

[WZW24] Runzhou Wang, Xinsheng Zhang, and Minghu Wang. A two-layer model with partial mapping: Unveiling the interplay between information dissemination and disease diffusion. Applied Mathematics and Computation, 468(??): ??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006768.

Wu:2022:FTA

[WZXH22] Ziwen Wu, Tianping Zhang, Xiaonan Xia, and Yu Hua. Finite-time adaptive neural command filtered control for non-strict feedback uncertain multi-agent systems including prescribed performance and input nonlinearities. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200039X.

Wang:2020:IRM

[WZXY20] Yingchun Wang, Yu Zheng, Xiangpeng Xie, and Jun Yang. An improved reduction method based networked control against false data injection attacks and stochastic input delay. Applied Mathematics and Computation, 385 (??):Article 125421, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303829.

Wang:2020:GEE

[WZY20] Jie Wang, Ruigang Zhang, and Liangui Yang. A Gardner evolution equation for topographic Rossby waves and its mechanical analysis. Applied Mathematics and Computation, 385 (??):Article 125426, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303878.

Wang:2021:PEL

[WZY21] Yayun Wang, Zhimin Zhang, and Wenguang Yu. Pricing equity-linked death benefits by complex Fourier series expansion in a regimeswitching jump diffusion model. Applied Mathematics and Computation, 399 (??):Article 126031, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000795.

Wang:2023:STR

[WZZ23] Yan Wang, Lingxin Zhou, and Xuyuan Zhang. Spatiotemporal regularized shockdiffusion filtering with local entropy for restoration of degraded document images. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006919.

Wang:2024:SCT

[WZZ24] Guoliang Wang, Zhikang Zhu, and Yande Zhang. Stabilization of continuoustime Markovian jump systems: a mode separation but optimization method. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001164.

Wang:2021:FTA

[WZZX21] Yingchun Wang, Jiaxin Zhang, Huaguang Zhang, and Xiangpeng Xie. Finitetime adaptive neural control for nonstrict-feedback stochastic nonlinear systems with input delay and output constraints. Applied Mathematics and Computation, 393(??):Article 125756, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307098.

Xu:2021:AAT

[XAZ^{+}21] Minghan Xu, Saad Akhtar, Ahmad F. Zueter, Mahmoud A. Alzoubi, Laxmi Sushama, and Agus P. Sasmito. Asymptotic analysis of a two-phase Stefan problem in annulus: Application to outward solidification in phase change materials. Applied Mathematics and Computation, 408 (??):Article 126343, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100432X.

Xiao:2022:MOM

[XBC22a] Guiyun Xiao, Zheng-Jian Bai, and Wai-Ki Ching. A modified orthogonal matching pursuit for construction of sparse probabilistic boolean networks. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001278.

Xu:2022:RAR

[XBC22b] Wei-Ru Xu, Natália Bebiano, and Guo-Liang Chen. A reduction algorithm for reconstructing periodic Jacobi matrices in Minkowski spaces. Applied Mathematics and Computation, 419(??):Article 126853, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S009630032100936X.

Xia:2023:CCF

[XBLW23] Xue Xia, Jing Bai, Xiaohe Li, and Guoguang Wen. Containment control for fractional order MASs with nonlinearity and time delay via pull-based event-triggered mechanism. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002631.

Xian:2020:SIT

[XbYW20] Jun Xian, Xiong bin Yan, and Ting Wei. Simultaneous identification of three parameters in a timefractional diffusion-wave equation by a part of boundary Cauchy data. Applied Mathematics and Computation, 384(??):Article 125382, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320303350.

Xie:2024:DIP

[XCGN24] Yujia Xie, Wengu Chen, Huanmin Ge, and Michael K. Ng. Deep image prior and weighted anisotropicisotropic total variation regularization for solving linear inverse problems. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004132.

Xu:2021:CFT

[XCLW21] Xiaofeng Xu, Mou Chen, Tao Li, and Qingxian Wu. Composite fault tolerant attitude control for flexible satellite system under disturbance and input delay. Applied Mathematics and Computation, 409 (??):Article 126419, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005087.

Xu:2021:EHO

[XCSW21] Zhuangzhi Xu, Wenjun Cai, Yongzhong Song, and Yushun Wang. Explicit high-order energypreserving exponential time differencing method for nonlinear Hamiltonian PDEs. Applied Mathematics and Computation, 404(??):Article 126208, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321002988.

Xie:2022:MBE

[XCW^{+}22] Lifei Xie, Jun Cheng, Hailing Wang, Jiange Wang, Mengjie Hu, and Zhidong Zhou. Memory-based eventtriggered asynchronous control for semi-Markov switching systems. Applied Mathematics and Computation, 415(??):Article 126694, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007785.

Xie:2022:SNI

[XCZ22] Jiayi Xie, Zhenyu Cui, and Zhimin Zhang. Some new infinite series expansions for the first passage time densities in a jump diffusion model with phase-type jumps. Applied Mathematics and Computation, 429(??):??, September 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003253.

Xi:2021:AAT

[XD21] Changjiang Xi and Jiuxiang Dong. Adaptive asymptotic tracking control of uncertain nonlinear timedelay systems depended on delay estimation information. Applied Mathematics and Computation, 391 (??):Article 125662, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306159.

Xiao:2023:DOF

[XD23a] Shuyi Xiao and Jiuxiang Dong. Distributed outputfeedback resilient faulttolerant tracking control of uncertain heterogeneous linear MASs under directed topologies and DoS attacks. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008700.

Xing:2023:EDM

[XD23b] Rundan Xing and Zhibin Du. The exponential distance matrix of block graphs. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032200741X.

Xu:2020:MUC

[XDS^{+}20] Jiwei Xu, Zhenghong Deng, Qun Song, Qian Chi, Tao Wu, Yijie Huang, Dan Liu, and Mingyu Gao. Multi-UAV counter-game model based on uncertain information. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306769.

Xie:2022:NCD

[XFK22] Jin-Hua Xie, Yan-Quan Feng, and Young Soo Kwon. Normal Cayley digraphs of generalized quaternion groups with CI-property. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000522.

Xi:2021:NLC

[XFW^{+}21] Qiang Xi, Zhuojia Fu, Wenjie Wu, Hui Wang, and Yong Wang. A novel localized collocation solver based on Trefftz basis for potentialbased inverse electromyography. Applied Mathematics and Computation, 390 (??):Article 125604, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305592.

Xue:2020:GCR

[XFY20] Min Xue, Chao Fu, and Shan-Lin Yang. Group consensus reaching based on a combination of expert weight and expert reliability. Applied Mathematics and Computation, 369(??): ??, March 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630031930894X.

Xu:2020:ANA

[XGF20] Zhenhua Xu, Hongrui Geng, and Chunhua Fang. Asymptotics and numerical approximation of highly oscillatory Hilbert transforms. Applied Mathematics and Computation, 386 (??):Article 125525, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304835.

Xu:2020:DSR

[XGRM20] Ying Xu, Yeye Guo, Guodong Ren, and Jun Ma. Dynamics and stochastic resonance in a thermosensitive neuron. Applied Mathematics and Computation, 385 (??):Article 125427, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032030388X.

Xia:2022:FTA

[XH22a] ZeLiang Xia and Shuping He. Finite-time asynchronous H_{∞} fault-tolerant control for nonlinear hidden Markov jump systems with actuator and sensor faults. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002867.

Xiong:2022:EFI

[XH22b] Zhuang Xiong and Yaoping Hou. Eigenvalue-free interval for Seidel matrices of threshold graphs. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200251X.

See corrigendum [XH23].

Xiong:2023:CEF

[XH23] Zhuang Xiong and Yaoping Hou. Corrigendum to “Eigenvalue-free interval for Seidel matrices of threshold graphs” [Applied Mathematics and Computation 427 (2022) 127177]. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008888.

See [XH22b].

He:2021:AOM

[xHhPjW^{+}21] Zhen xue He, Yu hua Pan, Ke jian Wang, Li min Xiao, and Xiang Wang. Area optimization for MPRM logic circuits based on improved multiple disturbances fireworks algorithm. Applied Mathematics and Computation, 399(??):Article 126008, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000564.

Xue:2020:APM

[XHL20] Dandan Xue, Yanren Hou, and Wenjia Liu. Analysis of the parareal method with spectral deferred correction method for the Stokes/ Darcy equations. Applied Mathematics and Computation, 387(??):Article 124625, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306174.

Xing:2023:LCF

[XHL23] Ying Xing, Xinyi He, and Xiaodi Li. Lyapunov conditions for finite-time stability of disturbed nonlinear impulsive systems. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007366.

Xu:2020:NAF

[XHWM20] Fei Xu, Qiumei Huang, Min Wang, and Hongkun Ma. A novel adaptive finite element method for the ground state solution of Bose–Einstein condensates. Applied Mathematics and Computation, 385 (??):Article 125404, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303660.

Xia:2021:CAA

[Xia21] Ke Xia. The characteristics of average abundance function of multi-player threshold public goods evolutionary game model under redistribution mechanism. Applied Mathematics and Computation, 392(??): Article 125733, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030686X.

Xiao:2022:CBR

[Xia22]  Lin Xiao.  Compound

binomial risk model in a Markovian environment with capital cost and the calculation algorithm. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000558.

Xie:2020:SCC

[XJ20] Zhifu Xie and William Johnson. Super central configurations in the collinear 5-body problem. Applied Mathematics and Computation, 381(??):Article 125194, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301636.

Xu:2024:IJS

[XJY24] Hao Xu, Hehua Ju, and Meng Yu. An innovative joint-space dynamic theory for mobile multi-axis system with unilateral constraint. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400345X.

Xu:2023:PSG

[XKW^{+}23] Libai Xu, Dehan Kong, Lidan Wang, Hong Gu, Toby Kenney, and Ximing Xu. Proportional stochastic generalized Lotka–Volterra model with an application to learning microbial community structures. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001017.

Xie:2024:NIM

[XL24a] Huiqing Xie and Manhong Lu. A new iterative method for simultaneous computation of several eigenpair derivatives of a large matrix. Applied Mathematics and Computation, 472(??): ??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000985.

Xie:2024:RGE

[XL24b] Kai Xie and Tingjin Liu.

The regulation of good and evil promotes cooperation in public goods game. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003059.

Xia:2023:LMU

[XLD23] Peng Xia, Na Lei, and Tian Dong. On the linearization methods for univariate Birkhoff rational interpolation. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008992.

Xu:2021:TAQ

[XLG^{+}21] Ke Xu, Xiang Li, Zhenxun Gao, Chongwen Jiang, and Chun-Hian Lee. Theoretical analysis of quasione-dimensional compressible magnetohydrodynamic channel flow. Applied Mathematics and Computation, 411(??):Article 126500, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005890.

Xiao:2022:IFT

[XLJL22] Lin Xiao, Xiaopeng Li, Lei Jia, and Sai Liu. Improved finite-time solutions to time-varying Sylvester tensor equation via zeroing neural networks. Applied Mathematics and Computation, 416(??):Article 126760, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008420.

Xu:2021:ARP

[XLL21a] Jiale Xu, Mei Lu, and Ke Liu. Anti-Ramsey problems for cycles. Applied Mathematics and Computation, 408(??):Article 126345, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004343.

Xue:2021:IDN

[XLL^{+}21b] Qiang Xue, Chen Liu, Li Li, Gui-Quan Sun, and Zhen Wang. Interactions of diffusion and nonlocal delay give rise to vegetation patterns in semi-arid environments. Applied Mathematics and Computation, 399(??): Article 126038, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000862.

Xi:2024:PSC

[XLL24] Wenying Xi, Wensong Lin, and Yuquan Lin. Packing 2- and 3-stars into cubic graphs. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004563.

Xu:2022:ACF

[XLLH22] Bo Xu, Yanjun Liang, Yuan-Xin Li, and Zhongsheng Hou. Adaptive command filtered fixed-time control of nonlinear systems with input quantization. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002600.

Xue:2021:GWT

[XLS21] Jie Xue, Ruifang Liu, and Jinlong Shu. On graphs whose third largest distance eigenvalue dose not exceed −1. Applied Mathematics and Computation, 402(??):Article 126137, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001855.

Xu:2021:FEB

[XLYA21] Changjin Xu, Zixin Liu, Lingyun Yao, and Chaouki Aouiti. Further exploration on bifurcation of fractional-order six-neuron bi-directional associative memory neural networks with multi-delays. Applied Mathematics and Computation, 410(??):Article 126458, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005476.

Xia:2023:DAC

[XLZ23] Meizhen Xia, Zhucheng Liu, and Tianping Zhang. Distributed adaptive cooperative control via command filters for multiagent systems including input unmodeled dynamics and sensor faults. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003636.

Xu:2024:FTB

[XLZ^{+}24a] Yao Xu, Xinzhi Liu, Lu Zhang, Wenxue Li, Yongbao Wu, and Yang Liu. Fixedtime bipartite synchronization of nonlinear impulsive time-varying signed networks with delays. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003667.

Xu:2024:TVF

[XLZ^{+}24b] Ziqiang Xu, Yun Li, Xisheng Zhan, Huaicheng Yan, and Yiyan Han. Timevarying formation of uncertain nonlinear multiagent systems via adaptive feedback control approach with event-triggered impulsive estimator. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001796.

Xu:2023:ASH

[XLZL23] Xiangtian Xu, Gaosheng Li, Yuqin Zhao, and Tiejun Liu. Analytical solutions for heat conduction problems with three kinds of periodic boundary conditions and their applications. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008037.

Xi:2020:OTC

[XPP^{+}20] Xiaojian Xi, Shirin Panahi, Viet-Thanh Pham, Zhen Wang, Sajad Jafari, and Iqtadar Hussain. Optimum topology and coupling strength for synchronization. Applied Mathematics and Computation, 379(??):Article 125226, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301958.

Xue:2022:ETA

[XRZH22] Yanmei Xue, Wen Ren, Bo-Chao Zheng, and Jinke Han. Event-triggered adaptive sliding mode control of cyber-physical systems under false data injection attack. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004775.

Xie:2022:RSS

[XSP22] Xiangpeng Xie, Xicheng Shen, and Chen Peng. Relaxed stabilization synthesis of discrete-time nonlinear systems with uplink data loss based on a novel online evaluation mechanism. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000352.

Xiao:2023:MNI

[XT23] Junyi Xiao and Jianhua Tu. Maximizing the number of independent sets in claw-free cubic graphs. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008803.

Xu:2022:NRK

[XTNF22] Minqiang Xu, Emran Tohidi, Jing Niu, and Yuzhi Fang. A new reproducing kernel-based collocation method with optimal convergence rate for some classes of BVPs. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004179.

Xin:2022:ORL

[XTS^{+}22] Xilin Xin, Yidong Tu, Vladimir Stojanovic, Hai Wang, Kaibo Shi, Shuping He, and Tianhong Pan. Online reinforcement learning multiplayer non-zero sum games of continuous-time Markov jump linear systems. Applied Mathematics and Computation, 412 (??):Article 126537, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006214.

Xie:2022:WGF

[XTT22] Yingying Xie, Ming Tang, and Chunming Tang. A weak Galerkin finite element method for indefinite time-harmonic Maxwell equations. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005458.

Xu:2020:PMP

[Xu20] Xuefeng Xu. On the perturbation of the Moore– Penrose inverse of a matrix. Applied Mathematics and Computation, 374 (??):Article 124920, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309129.

Xu:2021:NUB

[Xu21] Yian Xu. A new upper bound for Conway’s thrackles. Applied Mathematics and Computation, 389 (??):Article 125573, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305294.

Xu:2023:SP

[Xu23] Yian Xu. On the structure of pentagraphs. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007044.

Xu:2024:FTE

[Xu24a] Liqiong Xu. Fault tolerability evaluation for the component edge-connectivity of folded Petersen networks. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001450.

Xu:2024:CGI

[Xu24b] Yian Xu. Coloring graphs without induced P_{5} or K_{5} − e. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004836.

Xie:2022:RDI

[XW22] Yingkang Xie and Zhen Wang. A ratio-dependent impulsive control of an SIQS epidemic model with non-linear incidence. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001047.

Xu:2023:WSN

[XW23] Tongtong Xu and Baoyindureng Wu. Weak saturation number of a complete bipartite graph. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002667.

Xu:2023:SDC

[XWDK23] Xiao Xu, Li Wang, Zhenbin Du, and Yonggui Kao. H_{∞} sampled-data control for uncertain fuzzy systems under Markovian jump and FBm. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001832.

Xu:2022:FTA

[XWL^{+}22] Jiahong Xu, Lijie Wang, Yang Liu, Jize Sun, and Yingnan Pan. Finite-time adaptive optimal consensus control for multi-agent systems subject to timevarying output constraints. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002508.

Xu:2023:AFF

[XWL23] Ke Xu, Huanqing Wang, and Peter Xiaoping Liu.

Adaptive fuzzy finite-time tracking control of nonlinear systems with unmodeled dynamics. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001613.

Xie:2020:SAC

[XWLL20] Yingkang Xie, Zhen Wang, Junwei Lu, and Yuxia Li. Stability analysis and control strategies for a new SIS epidemic model in heterogeneous networks. Applied Mathematics and Computation, 383(??):Article 125381, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303453.

Xiao:2023:SAF

[XWM23] Shasha Xiao, Zhanshan Wang, and Lei Ma. Synchronization analysis of fractional order delayed BAM neural networks via multi-delay-boundary inequality. Applied Mathematics and Computation, 451 (??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300323002023.

Xia:2020:FRF

[XWMC20] Yude Xia, Jing Wang, Bo Meng, and Xiangyong Chen. Further results on fuzzy sampled-data stabilization of chaotic nonlinear systems. Applied Mathematics and Computation, 379(??):Article 125225, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301946.

Xiong:2021:LDC

[XWWC21] Shixun Xiong, Qingxian Wu, Yuhui Wang, and Mou Chen. An ℓ_{2}– ℓ_{∞} distributed containment coordination tracking of heterogeneous multiunmanned systems with switching directed topology. Applied Mathematics and Computation, 404(??): Article 126080, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001284.

Xu:2023:ETA

[XWWL23] Lin-Xing Xu, Yu-Long Wang, Fei Wang, and Yue Long. Event-triggered active disturbance rejection trajectory tracking control for a quadrotor unmanned aerial vehicle. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001364.

Xing:2021:EBA

[XWZC21] Mingqi Xing, Yanqian Wang, Guangming Zhuang, and Fu Chen. Eventbased asynchronous and resilient filtering for singular Markov jump LPV systems against deception attacks. Applied Mathematics and Computation, 403 (??):Article 126176, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002666.

Xiong:2024:PTF

[XXJtC24] Shi-Xun Xiong, Xiang-Peng Xie, Guo-Ping Jiang, and Meng ting Chen. Prescribed time formation synchronous tracking of unmanned aerial vehicles with denial of service attacks. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001395.

Xia:2021:ETF

[XXL^{+}21] Weifeng Xia, Shengyuan Xu, Junwei Lu, Yongmin Li, Yuming Chu, and Zhengqiang Zhang. Event-triggered filtering for discrete-time Markovian jump systems with additive time-varying delays. Applied Mathematics and Computation, 391 (??):Article 125630, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305841.

Xu:2020:EDB

[XXW^{+}20] Tianshu Xu, Jianwei Xia, Sanxia Wang, Yuxiao Lian, and Huasheng Zhang. Extended dissipativity-based non-fragile sampled-data control of fuzzy Markovian jump systems with incomplete transition rates. Applied Mathematics and Computation, 380(??):Article 125258, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302277.

Xiao-yong:2020:NSN

[Xy20] Zhang Xiao-yong. A new strategy for the numerical solution of nonlinear Volterra integral equations with vanishing delays. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306009.

Xu:2022:EED

[XY22] Hengmeng Xu and Weigen Yan. On eigenvalues and the energy of dendrimer trees. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001370.

Xiang:2021:CMS

[XYCG21] Guangjian Xiang, Deshun Yin, Chenxi Cao, and Yunfei Gao. Creep modelling of soft soil based on the fractional flow rule: Simulation and parameter study. Applied Mathematics and Computation, 403 (??):Article 126190, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321002800.

Xiong:2024:NAB

[XYH^{+}24] Kailong Xiong, Juan Yu, Cheng Hu, Shiping Wen, and Fanchao Kong. Nonseparation analysis-based finite/fixed-time synchronization of fully complexvalued impulsive dynamical networks. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300323006690.

Xiu:2020:TFT

[XYKL20] Xianchao Xiu, Ying Yang, Lingchen Kong, and Wanquan Liu. tSSNALM: a fast two-stage semi-smooth Newton augmented Lagrangian method for sparse CCA. Applied Mathematics and Computation, 383 (??):Article 125272, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302411.

Xu:2021:GAS

[XYLF21] Yao Xu, Jintong Yu, Wenxue Li, and Jiqiang Feng. Global asymptotic stability of fractional-order competitive neural networks with multiple time-varying-delay links. Applied Mathematics and Computation, 389(??):Article 125498, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304562.

Xu:2024:TRP

[XYlW^{+}24] Zhi Xu, Jing-Hua Yang, Chuan long Wang, Fusheng Wang, and Xi hong Yan. Tensor robust principal component analysis with total generalized variation for high-dimensional data recovery. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004417.

Xie:2022:FDS

[XYM22] Yingying Xie, Daopeng Yin, and Liquan Mei. Finite difference scheme on graded meshes to the timefractional neutron diffusion equation with nonsmooth solutions. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005483.

Xu:2023:SAO

[XYR^{+}23] Conghui Xu, Yongguang Yu, Guojian Ren, Yuqin Sun, and Xinhui Si. Stability analysis and optimal control of a fractionalorder generalized SEIR model for the COVID-19 pandemic. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300379X.

Xia:2023:HDS

[XYY23] Siwei Xia, Yuehan Yang, and Hu Yang. Highdimensional sparse portfolio selection with nonnegative constraint. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008347.

Xin:2020:AFB

[XYZY20] Li-Ping Xin, Bo Yu, Lin Zhao, and Jinpeng Yu. Adaptive fuzzy backstepping control for a two continuous stirred tank reactors process based on dynamic surface control approach. Applied Mathematics and Computation, 377 (??):Article 125138, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320301077.

Xu:2022:MOP

[XYZZ22] Chentao Xu, Zhe You, Liwen Zhang, and Minghong Zhou. The maximum outdegree power of complete k-partite oriented graphs. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200488X.

Xie:2021:FTD

[XZ21] Jiayi Xie and Zhimin Zhang. Finite-time dividend problems in a Lévy risk model under periodic observation. Applied Mathematics and Computation, 398 (??):Article 125981, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000291.

Xiang:2023:AFI

[XZ23a] Shuhuang Xiang and Qingyang Zhang. Asymptotics on the Fredholm integral equation with a highly oscillatory and weakly singular kernel. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002813.

Xiao:2023:DMN

[XZ23b] Ya Xiao and Linhua Zhou.

Dynamic modelling and numerical simulation of formation control for intelligent multi-agent system with target geometric configuration. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008943.

Xu:2023:RAU

[XZ23c] Qinqin Xu and Yuanguo Zhu. Reliability analysis of uncertain random systems based on uncertain differential equation. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001777.

Xu:2024:PIG

[XZ24] Leyou Xu and Bo Zhou.

The path-index of a graph. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004812.

Xie:2020:DDD

[XZeF20] Jiyang Xie, Shuqian Zhu, and Jun e Feng. Delaydependent and decay-rate-dependent conditions for exponential mean stability and non-fragile controller design of positive Markov jump linear systems with time-delay. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308264.

Xiao:2020:CPR

[XZH^{+}20] Yiyong Xiao, Xiaorong Zuo, Jiaoying Huang, Abdullah Konak, and Yuchun Xu. The continuous pollution routing problem. Applied Mathematics and Computation, 387(??):Article 125072, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320300412.

Xu:2022:PAF

[XZQ^{+}22] Tianbo Xu, Chunxia Zhu, Wenhai Qi, Jun Cheng, Kaibo Shi, and Liangliang Sun. Passive analysis and finite-time anti-disturbance control for semi-Markovian jump fuzzy systems with saturation and uncertainty. Applied Mathematics and Computation, 424(??):??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001163.

Xu:2023:FTC

[XZQ23] Qiyi Xu, Ning Zhang, and Wenhai Qi. Finite-time control for discrete-time nonlinear Markov switching LPV systems with DoS attacks. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008517.

Xu:2020:ETM

[XZQX20] Qiyi Xu, Yijun Zhang, Wenhai Qi, and Shunyuan Xiao. Event-triggered mixed H_{∞} and passive filtering for discrete-time networked singular Markovian jump systems. Applied Mathematics and Computation, 368(??): ??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307957.

Xu:2020:SMB

[XZS20] Junyan Xu, Tonghua Zhang, and Keying Song. A stochastic model of bacterial infection associated with neutrophils. Applied Mathematics and Computation, 373(??):Article 125025, May 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319310173.

Xue:2021:TDM

[XZSS21] Ling Xue, Hongyu Zhang, Wei Sun, and Caterina Scoglio. Transmission dynamics of multi-strain dengue virus with crossimmunity. Applied Mathematics and Computation, 392(??):Article 125742, March 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306950.

Xiong:2023:ETD

[XZY^{+}23] Menghui Xiong, Baoyong Zhang, Deming Yuan, Yijun Zhang, and Jun Chen. Event-triggered distributed online convex optimization with delayed bandit feedback. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000346.

Xu:2020:GLS

[XZZ20] Yang Xu, Yanming Zhang, and Jingjun Zhao. General linear and spectral Galerkin methods for the Riesz space fractional diffusion equation. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306563.

Xiao:2022:NVE

[XZZ22a] Liuchao Xiao, Meng Zhou, and Jikun Zhao. The nonconforming virtual element method for semilinear elliptic problems. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004763.

Xie:2022:RDS

[XZZ22b] Jiyang Xie, Shuqian Zhu, and Dawei Zhang. A robust distributed secure interval observation approach for uncertain discrete-time positive systems under deception attacks. Applied Mathematics and Computation, 413(??):Article 126638, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007220.

Xu:2022:GES

[XZZC22] Xiaozeng Xu, Hongbin Zhang, Qunxian Zheng, and Wei Chen. Global exponential stability and H_{∞} control of limit cycle for switched affine systems under time-dependent switching signal. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008894.

Xu:2021:ACD

[XZZW21] Ning Xu, Xudong Zhao, Guangdeng Zong, and Yuanqing Wang. Adaptive control design for uncertain switched nonstrict-feedback nonlinear systems to achieve asymptotic tracking performance. Applied Mathematics and Computation, 408 (??):Article 126344, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004331.

Yetgin:2021:HFO

[YA21a] Zeki Yetgin and Hüseyin Abaci. Honey formation optimization framework for design problems. Applied Mathematics and Computation, 394(??):Article 125815, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307682.

Yildirim:2021:SS

[YA21b] Yasemin Yildirim and Erhan Ata. Screw surfaces. Applied Mathematics and Computation, 410 (??):Article 126476, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005658.

Yang:2023:FVM

[Yan23]  Wenming Yang.  A fi-

nite volume method for ferrohydrodynamic problems coupled with microscopic magnetization dynamics. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200772X.

Yang:2024:PET

[Yan24] Gengjiao Yang. The positivity and event-triggered stabilization of Takagi–Sugeno fuzzy systems with actuator saturation. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001127.

Ye:2021:FCP

[YAX^{+}21] Shuai Ye, Hengbin An, Xiaowen Xu, Xinhai Xu, and Xuejun Yang. A filter in constructing the preconditioner for solving linear equation systems of radiation diffusion problems. Applied Mathematics and Computation, 399(??): Article 126010, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000588.

Yin:2021:SLS

[YC21] Wensheng Yin and Jinde Cao. On stability of largescale G-SDEs: a decomposition approach. Applied Mathematics and Computation, 388(??):Article 125466, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304264.

Yang:2021:SAT

[YCEW21] Suxiang Yang, Huanzhen Chen, Vincent J. Ervin, and Hong Wang. Solvability and approximation of two-side conservative fractional diffusion problems with variablecoefficient based on leastsquares. Applied Mathematics and Computation, 406(??):Article 126229, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003192.

Yang:2021:SMC

[YCL^{+}21] Yi Yang, Fei Chen, Jiahong Lang, Xiangyong Chen, and Jing Wang. Sliding mode control of persistent dwell-time switched systems with random data dropouts. Applied Mathematics and Computation, 400(??):Article 126087, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321001351.

Yang:2024:CDC

[YCL24] Jin-Xuan Yang, Zhong-Pan Cao, and Yikang Lu. Contagion dynamics on a compound model. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004629.

Yang:2022:FTA

[YCM^{+}22] Wei Yang, Guozeng Cui, Qian Ma, Jiali Ma, and Chongben Tao. Finite-time adaptive event-triggered command filtered backstepping control for a QUAV. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321009814.

Yu:2023:DCP

[YCW23] Jiahao Yu, Min Chen, and Weifan Wang. 2distance choosability of planar graphs with a restriction for maximum degree. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001182.

Yi:2023:OCA

[YCWH23] Huaming Yi, Yanping Chen, Yang Wang, and Yunqing Huang. Optimal convergence analysis of a linearized second-order BDF-PPIFE method for semi-linear parabolic interface problems. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006555.

Yu:2020:CAC

[YCZ^{+}20] Liqin Yu, Fuyuan Cao, Xingwang Zhao, Xiaodan Yang, and Jiye Liang. Combining attribute content and label information for categorical data ensemble clustering. Applied Mathematics and Computation, 381(??): Article 125280, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302496.

Ye:2023:ETB

[YCZ^{+}23] Hu Ye, Peng Cheng, Xiang Zhang, Shuping He, and Weidong Zhang. Eventtriggered-based H_{∞} control for Markov jump cyberphysical systems against denial-of-service attacks. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001996.

Yang:2022:MFZ

[YD22a] Jiaxiang Yang and Hanyuan Deng. Maximum first Zagreb index of orientations of unicyclic graphs with given matching number. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002156.

Yu:2022:SAM

[YD22b] Peilin Yu and Feiqi Deng. Stabilization analysis of Markovian asynchronous switched systems with input delay and Lévy noise. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000583.

Yu:2022:SAI

[YDSW22] Peilin Yu, Feiqi Deng, Yuanyuan Sun, and Fangzhe Wan. Stability analysis of impulsive stochastic delayed Cohen–Grossberg neural networks driven by Lévy noise. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005185.

Yao:2022:CCN

[YDXH22] Dajie Yao, Chunxia Dou, Xiangpeng Xie, and Songlin Hu. Containment control of non-affine multi-agent systems based on given precision. Applied Mathematics and Computation, 412 (??):Article 126579, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006639.

Yeganeh:2021:CNO

[YF21a] Solmaz Mousavi Yeganeh and Javad Farzi. A class of non-oscillatory direct-space-time schemes for hyperbolic conservation laws. Applied Mathematics and Computation, 399(??):Article 126013, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000618.

Yin:2021:SGV

[YF21b] Fu-Gang Yin and Yan-Quan Feng. Symmetric graphs of valency 4 having a quasi-semiregular automorphism. Applied Mathematics and Computation, 399(??): Article 126014, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100062X.

Ye:2021:ODL

[YFHC21] Zuyang Ye, Qingli Fan, Shibing Huang, and Aiping Cheng. A one-dimensional line element model for transient free surface flow in porous media. Applied Mathematics and Computation, 392(??):Article 125747, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307001.

You:2024:IMF

[YFM^{+}24] Xuemei You, Xiaonan Fan, Yinghong Ma, Zhiyuan Liu, and Ruifeng Zhang. Impact of message fatigue in information-disease coupled dynamics on temporal simplicial networks. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003400.

Yang:2021:CPL

[YFW21] Jian Yang, Michal Feckan, and JinRong Wang. Consensus problems of linear multi-agent systems involving conformable derivative. Applied Mathematics and Computation, 394 (??):Article 125809, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307621.

Yang:2024:NPR

[YG24] Qian Yang and Yunshu Gao. Near packings with restriction on degrees in graphs. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002893.

Yao:2021:GPT

[YGHW21] Hejun Yao, Fangzheng Gao, Jiacai Huang, and Yuqiang Wu. Global prescribed-time stabilization via time-scale transformation for switched nonlinear systems subject to switching rational powers. Applied Mathematics and Computation, 393(??): Article 125766, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307190.

Yadav:2022:NTM

[YGM^{+}22] Vivek S. Yadav, Naveen Ganta, Bikash Mahato, Manoj K. Rajpoot, and Yogesh G. Bhumkar. New time-marching methods for compressible Navier–Stokes equations with applications to aeroacoustics problems. Applied Mathematics and Computation, 419(??): Article 126863, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009462.

Yan:2022:REH

[YGZ^{+}22] Zhilian Yan, Tong Guo, Anqi Zhao, Qingkai Kong, and Jianping Zhou. Reliable exponential H_{∞} filtering for a class of switched reaction–diffusion neural networks. Applied Mathematics and Computation, 414(??):Article 126661, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007451.

Yang:2024:DFA

[YH24a] Yi Yang and Jin Huang.

Double fast algorithm for solving time-space fractional diffusion problems with spectral fractional Laplacian. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300324001875.

Yu:2024:HST

[YH24b] Qian Yu and Yaoping Hou. Harmonic signed trees. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002169.

Yang:2020:MME

[YHB^{+}20] Yidu Yang, Jiayu Han, Hai Bi, Hao Li, and Yu Zhang. Mixed methods for the elastic transmission eigenvalue problem. Applied Mathematics and Computation, 374 (??):Article 125081, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300503.

Yang:2024:EFE

[YHD24] Wenwan Yang, Xiao-Min Huang, and Feifei Du. Embedding and fractional embedding of Bergman-type spaces in the Schatten class. Applied Mathematics and Computation, 472(??):??, July 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001000.

Yu:2020:MOH

[YHL^{+}20] Jianwu Yu, Kaifeng Huang, Hong Luo, Yao Wu, and Xiaobing Long. Manipulate optimal highorder motion parameters to construct high-speed cam curve with optimized dynamic performance. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309452.

Yu:2020:UCS

[YHSZ20] Xiaohui Yu, Mingke He, Hongxia Sun, and Zhen Zhou. Uncertain coalition structure game with payoff of belief structure. Applied Mathematics and Computation, 372(??):Article 125000, May 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309920.

Yang:2023:FQF

[YHW^{+}23] Yi Yang, Jin Huang, Yifei Wang, Ting Deng, and Hu Li. Fast Q1 finite element for two-dimensional integral fractional Laplacian. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008256.

Yuzbasi:2022:BQT

[YI22] Suayip Yüzbasi and Mohammad Izadi. Besselquasilinearization technique to solve the fractional-order HIV-1 infection of CD4+ Tcells considering the impact of antiviral drug treatment. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003939.

Yi:2023:AFC

[Yi23] Jiale Yi. Adaptive fuzzy connectivity-preserving consensus protocols for stochastic strict-feedback nonlinear MASs subject to unmeasured periodic disturbances. Applied Mathematics and Computation, 444(??):??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008931.

Yao:2020:SON

[YJ20] Changhui Yao and Shanghui Jia. A second order numerical scheme for nonlinear Maxwell’s equations using conforming finite element. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309324.

Yao:2020:SBB

[YJL20] Yan Yao, Shengjin Ji, and Guang Li. On the sharp bounds of bicyclic graphs regarding edge Szeged index. Applied Mathematics and Computation, 377 (??):Article 125135, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301041.

Yuan:2022:GES

[YK22] Baoquan Yuan and Xueli Ke. Global existence of strong solutions to the multi-dimensional inhomogeneous incompressible MHD equations. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002326.

Ye:2020:ETF

[YL20a] Dan Ye and Xiehuan Li.

Event-triggered fault detection for continuoustime networked polynomialfuzzy-model-based systems. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307210.

Yu:2020:CSE

[YL20b] Guihai Yu and Xingfu Li. Connective Steiner 3-eccentricity index and network similarity measure. Applied Mathematics and Computation, 386 (??):Article 125446, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304070.

Yu:2020:NSC

[YL20c] Qiang Yu and Hao Lv. The new stability criteria of discrete-time switched systems with an improved mode dependent average dwell time approach. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307222.

Yang:2022:NAA

[YL22] Shuang Yang and Yangrong Li. Numerical attractors and approximations for stochastic or deterministic sine-Gordon lattice equations. Applied Mathematics and Computation, 413 (??):Article 126640, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007244.

Yao:2023:CEE

[YL23a] Binhong Yao and Peixing Li. Covariance estimation error of incomplete functional data under RKHS framework. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007809.

Yue:2023:ECD

[YL23b] Jing Yue and Jian Li. Efficient coupled deep neural networks for the timedependent coupled Stokes– Darcy problems. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005884.

Yu:2024:TSC

[YL24a] Binyan Yu and Yongshun Liang. On two special classes of fractal surfaces with certain Hausdorff and Box dimensions. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006781.

Yue:2024:DDC

[YL24b] Yuanyuan Yue and Qingshan Liu. Distributed dual consensus algorithm for time-varying optimization with coupled equality constraint. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400184X.

Yang:2020:FDH

[YLC20a] Nana Yang, Junmin Li, and Jiaxi Chen. Fully distributed hybrid adaptive learning consensus protocols for a class of non-linearly parameterized multi-agent systems. Applied Mathematics and Computation, 375 (??):Article 125074, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300436.

You:2020:TMM

[YLC20b] Xiangyu You, Wei Li, and Yingbin Chai. A truly meshfree method for solving acoustic problems using local weak form and radial basis functions. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306861.

Yu:2023:MAQ

[YLCL23] Siyi Yu, Hua Li, Xiaofeng Chen, and Dongyuan Lin. Multistability analysis of quaternion-valued neural networks with cosine activation functions. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000188.

Yang:2020:NDU

[YLCZ20] Yawei Yang, Chong Lin, Bing Chen, and Xin Zhao. H_{∞} observer design for uncertain one-sided Lipschitz nonlinear systems with time-varying delay. Applied Mathematics and Computation, 375(??):Article 125066, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300357.

Yang:2021:LEC

[YLG21] Yuna Yang, Hongwei Li, and Xu Guo. A linearized energy-conservative scheme for two-dimensional nonlinear Schrödinger equation with wave operator. Applied Mathematics and Computation, 404(??): Article 126234, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003246.

Yang:2020:NFM

[YLH20] Qingzhi Yang, Yiyong Li, and Pengfei Huang. A novel formulation of the max-cut problem and related algorithm. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309622.

Yu:2022:TIB

[YLH22] Guihai Yu, Xingfu Li, and Deyan He. Topological indices based on 2- or 3eccentricity to predict anti-HIV activity. Applied Mathematics and Computation, 416(??):Article 126748, March 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008304.

Yang:2022:VTI

[YLHZ22] Shiju Yang, Chuandong Li, Xiping He, and Wanli Zhang. Variable-time impulsive control for bipartite synchronization of coupled complex networks with signed graphs. Applied Mathematics and Computation, 420(??):Article 126899, May 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009826.

Yang:2021:AFT

[YLK^{+}21] Chengyu Yang, Fei Li, Qingkai Kong, Xiangyong Chen, and Jian Wang. Asynchronous fault-tolerant control for stochastic jumping singularly perturbed systems: an H_{∞} sliding mode control scheme. Applied Mathematics and Computation, 389(??):Article 125562, January 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030518X.

Yin:2020:CSH

[YLL20] Baoli Yin, Yang Liu, and Hong Li. A class of shifted high-order numerical methods for the fractional mobile/immobile transport equations. Applied Mathematics and Computation, 368(??):??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930791X.

Yang:2022:DFT

[YLL22] Yi Yang, Xiaohua Li, and Xiaoping Liu. Decentralized finite-time connective tracking control with prescribed settling time for p-normal form stochastic large-scale systems. Applied Mathematics and Computation, 412(??):Article 126581, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006652.

Yang:2020:SDR

[YLLL20] Jiaojiao Yang, Jianquan Lu, Jungang Lou, and Yang Liu. Synchronization of drive-response Boolean control networks with impulsive disturbances. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630031930671X.

Yang:2024:ASC

[YLM24] Hong Yang, Juan Liu, and Jixiang Meng. Antidirected spanning closed trail in tournaments. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005052.

Yang:2022:PCM

[YLPC22] Jinn-Shyong Yang, Xiao-Yan Li, Sheng-Lung Peng, and Jou-Ming Chang. Parallel construction of multiple independent spanning trees on highly scalable datacenter networks. Applied Mathematics and Computation, 413(??):Article 126617, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007013.

Yang:2023:FSM

[YLS23] Hong Yang, Cheng-Xue Lao, and Zi-Hang She. Fast solution methods for Riesz space fractional diffusion equations with nonseparable coefficients. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008979.

Yang:2020:ECC

[YLW20] Zhihu Yang, Zhi Li, and Long Wang. Evolution of cooperation in a conformitydriven evolving dynamic social network. Applied Mathematics and Computation, 379(??):Article 125251, August 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302204.

Yan:2021:CSH

[YLX^{+}21] Lizhao Yan, Jian Liu, Fei Xu, Kok Lay Teo, and Mingyong Lai. Control and synchronization of hyperchaos in digital manufacturing supply chain. Applied Mathematics and Computation, 391 (??):Article 125646, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320306007.

Yuan:2024:SRF

[YLY24] Qixuan Yuan, Ruifang Liu, and Jinjiang Yuan. The spectral radius of H_{2k}-free graphs. Applied Mathematics and Computation, 469(??):??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000328.

Yuan:2021:SQM

[YLZL21] Yongxin Yuan, Lina Liu, Huiting Zhang, and Hao Liu. The solutions to the quadratic matrix equation X^{∗}AX + B^{∗}X + D =

0. Applied Mathematics and Computation, 410 (??):Article 126463, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100552X.

Yu:2020:OBA

[YM20] Peng Yu and Yuechao Ma. Observer-based asynchronous control for Markov jump systems. Applied Mathematics and Computation, 377(??):Article 125184, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301533.

Yan:2021:IAG

[YM21] Tongxin Yan and Changfeng Ma. An iterative algorithm for generalized Hamiltonian solution of a class of generalized coupled Sylvesterconjugate matrix equations. Applied Mathematics and Computation, 411 (??):Article 126491, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005804.

Yang:2022:PBH

[YM22a] Hong Yang and Jixiang Meng. Paths in [h, k]- bipartite hypertournaments. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000911.

Yuan:2022:RDH

[YM22b] Hairui Yuan and Xinzhu Meng. Replicator dynamics of the hawk-dove game with different stochastic noises in infinite populations. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003460.

Yang:2023:ETC

[YM23a] Anqing Yang and Shuping Ma. Event-triggered control for switched singular systems with asynchronous switching and state jumps. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006178.

Yuan:2023:OCA

[YM23b] Lili Yuan and Jixiang Meng.

Ore conditions for antistrong digraphs. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003508.

Yuan:2021:MSG

[YML21] Xiying Yuan, Yanqi Mao, and Lele Liu. Maximal signed graphs with odd signed cycles as star complements. Applied Mathematics and Computation, 408 (??):Article 126367, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004562.

Yadav:2022:SPF

[YMRJ22] Vivek S. Yadav, Vikas Maurya, Manoj K. Rajpoot, and Jyoti Jaglan. Spatiotemporal pattern formations in stiff reaction–diffusion systems by new time marching methods. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003733.

Yang:2024:LSP

[YMW24] Liu Yang, Weijun Ma, and Xin Wang. Lagrange stability and passivity in the mean square sense of discrete-time stochastic Markovian switched neural networks with time-varying mixed delays. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002613.

Yuan:2021:CCF

[YMYR21] Xiaolin Yuan, Lipo Mo, Yongguang Yu, and Guojian Ren. Containment control of fractional discretetime multi-agent systems with nonconvex constraints. Applied Mathematics and Computation, 409(??):Article 126378, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300321004677.

Yeoh:2021:UPW

[YOF21] Chin Vern Yeoh, Ean Hin Ooi, and Ji Jinn Foo. Utilization of pressure wavedynamics in accelerating convergence of the lattice-Boltzmann method for steady and unsteady flows. Applied Mathematics and Computation, 411 (??):Article 126498, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005877.

Yoo:2021:DET

[Yoo21] Sung Jin Yoo. Decentralized event-triggered adaptive control of a class of uncertain interconnected nonlinear systems using local state feedback corrupted by unknown injection data. Applied Mathematics and Computation, 399(??):Article 126044, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000928.

Yousif:2020:CPR

[You20] Osman Omer Osman Yousif.

The convergence properties of RMIL+ conjugate gradient method under the strong Wolfe line search. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307696.

Yoo:2021:QFC

[YP21] Sung Jin Yoo and Bong Seok Park. Quantized feedback control strategy for tracking performance guarantee of nonholonomic mobile robots with uncertain nonlinear dynamics. Applied Mathematics and Computation, 407 (??):Article 126349, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004380.

Yoo:2022:UET

[YP22] Sung Jin Yoo and Bong Seok Park. A universal error transformation strategy for distributed eventtriggered formation tracking of pure-feedback nonlinear multiagent systems with communication and avoidance ranges. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004866.

Yadav:2020:HON

[YR20a] Swati Yadav and Pratima Rai. A higher order numerical scheme for singularly perturbed parabolic turning point problems exhibiting twin boundary layers. Applied Mathematics and Computation, 376(??): Article 125095, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300643.

Yang:2020:NPT

[YR20b] Caroline Yang and Nancy Rodriguez. A numerical perspective on traveling wave solutions in a system for rioting activity. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306381.

Yue:2020:NPR

[YS20] Jun Yue and Jiamei Song.

Note on the perfect Roman domination number of graphs. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306770.

Yee:2023:RAE

[YS23] H. C. Yee and Björn Sjögreen. Recent advancement of entropy split methods for compressible gas dynamics and MHD. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006191.

Yadav:2023:AEU

[YSHV^{+}23] Sonia Yadav, Sukhjit Singh, M. A. Hernández-Verón, Eulalia Martínez, Ajay Kumar, and R. P. Badoni. About the existence and uniqueness of solutions for some second-order nonlinear BVPs. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003879.

Yue:2021:PCR

[YSL21] Xiaohui Yue, Xingling Shao, and Jie Li. Prescribed chattering reduction control for quadrotors using aperiodic signal updating. Applied Mathematics and Computation, 405(??): Article 126264, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003532.

Yadav:2024:GPB

[YSL24] Abhishek Yadav, Amit Setia, and Concetta Laurita. A Gegenbauer polynomialbased numerical technique for a singular integral equation of order four and its application to a crack problem. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003394.

Yang:2023:SSI

[YSLK23] Xinrong Yang, Qilong Sun, Haitao Li, and Xiangshan Kong. Set stabilizability of impulsive probabilistic Boolean networks via impulsive sequence design. Applied Mathematics and Computation, 449(??):??, July 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001145.

You:2021:EAU

[YSW^{+}21] Tao You, Liang Shi, Xiaoyu Wang, Muslimbek Mengibaev, Ying Zhang, and Peng Zhang. The effects of aspiration under multiple strategy updating rules on cooperation in prisoner’s dilemma game. Applied Mathematics and Computation, 394 (??):Article 125770, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307232.

Yin:2022:DRU

[YSW^{+}22] Xiang Yin, Jinhua She, Min Wu, Daiki Sato, and Kouhei Ohnishi. Disturbance rejection using SMC-based-equivalent-input-disturbance approach. Applied Mathematics and Computation, 418(??):Article 126839, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100922X.

Sun:2022:UBT

[ySyS22] Ying ying Sun and Wan yi Sun. An update of a Bäcklund transformation and its applications to the Boussinesq system.

Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000509.

Yang:2024:BNF

[YSYS24] Yimei Yang, Hao Sun, Guangjing Yang, and Yanru Sun. Bilateral negotiation facilitates stable coexistence of cooperation with defection in Prisoner’s Dilemma game. Applied Mathematics and Computation, 471(??): ??, June 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000638.

Yan:2023:EIR

[YTC23a] Ruifang Yan, Wei Tong, and Guoxian Chen. An efficient invariant-region-preserving central scheme for hyperbolic conservation laws. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005744.

Yan:2023:MCW

[YTC23b] Ruifang Yan, Wei Tong, and Guoxian Chen. A mass conservative, well balanced and positivitypreserving central scheme for shallow water equations. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008463.

Yang:2020:SPS

[YTGC20] Yong Yang, Lilan Tu, Tianjiao Guo, and Jiabo Chen. Spectral properties of Supra-Laplacian for partially interdependent networks. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307325.

Yu:2022:ADO

[YTLV22] Qiang Yu, Ian Turner, Fawang Liu, and Viktor Vegh. The application of the distributed-order time fractional Bloch model to magnetic resonance imaging. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002624.

Yuhui:2024:CES

[YTX^{+}24] Qiu Yuhui, Lv Tianyang, Zhang Xizhe, Hu Honghua, and Ma Yuanchi. Cooperation emerged and survived in scale-free networks in co-evolution and betrayerprevailing circumstances. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400153X.

Yin:2020:CBB

[YTZ20] Hui-Min Yin, Bo Tian, and Xin-Chao Zhao. Chaotic breathers and breather fission/fusion for a vector nonlinear Schrödinger equation in a birefringent optical fiber or wavelength division multiplexed system. Applied Mathematics and Computation, 368(??): ??, March 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630031930760X.

Yao:2022:HSM

[YTZ22] Guoqing Yao, DongYa Tao, and Chao Zhang. A hybrid spectral method for the nonlinear Volterra integral equations with weakly singular kernel and vanishing delays. Applied Mathematics and Computation, 417(??):Article 126780, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008626.

Yu:2021:DWA

[Yu21] Qiang Yu. A decoupled wavelet approach for multiple physical flow fields of binary nanofluid in doublediffusive convection. Applied Mathematics and Computation, 404(??):Article 126232, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321003222.

Yu:2023:HBW

[Yu23] Qiang Yu. A homotopybased wavelet method for extreme large bending analysis of heterogeneous anisotropic plate with variable thickness on orthotropic foundation. Applied Mathematics and Computation, 439(??):??, February 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007135.

Yuan:2020:PSI

[Yua20]  Jinyun Yuan.  Preface

of special issue on Computational and Industrial Mathematics. Applied Mathematics and Computation, 387(??):Article 125288, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302575.

Yun:2021:WPR

[Yun21] Beong In Yun. C^{2} weighted piecewise rational interpolation. Applied Mathematics and Computation, 408 (??):Article 126366, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004550.

Yang:2020:ELE

[YW20] Xiuwen Yang and Ligong Wang. Extremal Laplacian energy of directed trees, unicyclic digraphs and bicyclic digraphs. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307295.

Yue:2021:ARB

[YW21] Gaocheng Yue and Jintao Wang. The attractors for the regularized Bénard problem with fractional Laplacian. Applied Mathematics and Computation, 391(??):Article 125640, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320305944.

Yang:2022:PEC

[YW22a] Wei Yang and Baoyindureng Wu. On Spacking edge-colorings of graphs with small edge weight. Applied Mathematics and Computation, 418 (??):Article 126840, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009231.

Yao:2022:BCF

[YW22b] Yu Yao and Li-Bing Wu. Backstepping control for fractional discrete-time systems. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005240.

Yang:2024:IDC

[YWC24] Guoli Yang, Yu’e Wu, and Matteo Cavaliere. Informationdriven cooperation on adaptive cyber-physical systems. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006550.

Yang:2023:HTR

[YWFL23] Jing Yang, Yongjiang Wu, Lihua Feng, and Weijun Liu. Hitting time, resistance distance and Kirchhoff index of Cayley graphs over dicyclic and semi-dihedral groups. Applied Mathematics and Computation, 448(??):??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001170.

Yao:2024:ISS

[YWG24] Meng Yao, Guoliang Wei, and Xinchen Guo. Inputto-state stabilization of networked impulsive systems under communication constraints: a refined Lyapunov functional. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005350.

Yin:2020:SBC

[YWJ20a] Qitian Yin, Mao Wang, and He Jing. Stabilizing backstepping controller design for arbitrarily switched complex nonlinear system. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307817.

You:2020:EHU

[YWJ^{+}20b] Tao You, Peng Wang, Danyang Jia, Fei Yang, Xiaodong Cui, and Chen Liu. The effects of heterogeneity of updating rules on cooperation in spatial network. Applied Mathematics and Computation, 372(??): Article 124959, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309518.

Yu:2023:CNA

[YWJ23] Guihai Yu, Yibo Wan, and Caoqing Jiang. Complex network analysis on provincial innovation development in China. Applied Mathematics and Computation, 455(??):??, October 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002722.

Yao:2020:GIM

[YWL20] Xin Yao, Yihe Wang, and Jianxing Leng. A gradually implicit method with deferred correction and adaptive relaxation factors for simulation to nonlinear heat transfer. Applied Mathematics and Computation, 385 (??):Article 125386, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303490.

Yang:2024:SLS

[YWR^{+}24] Dongsheng Yang, Hu Wang, Guojian Ren, Yongguang Yu, and Weiyi Xu. Successive lag synchronization of distributed-order multilayer neural networks with the short-term memory property. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003497.

Yao:2020:EMM

[YWS^{+}20] Wei Yao, Chunhua Wang, Yichuang Sun, Chao Zhou, and Hairong Lin. Exponential multistability of memristive Cohen–Grossberg neural networks with stochastic parameter perturbations. Applied Mathematics and Computation, 386 (??):Article 125483, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304422.

Yang:2024:CWI

[YWS24] Xujun Yang, Xiang Wu, and Qiankun Song. Caputo– Wirtinger integral inequality and its application to stability analysis of fractional-order systems with mixed time-varying delays. Applied Mathematics and Computation, 460 (??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300323004721.

Yin:2020:ECG

[YWT20] Li-Feng Yin, Xing-Ping Wu, and Chun-Lei Tang. Existence and concentration of ground state solutions for critical Schrödinger–Poisson system with steep potential well. Applied Mathematics and Computation, 374 (??):Article 125035, June 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300047.

Yang:2023:DAC

[YWWY23] Yang Yang, Xue Wang, Yingcheng Wu, and Wenbin Yue. Dynamic average consensus control based on event-triggered cloud access. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002230.

Yin:2020:NEM

[YWX^{+}20a] Qian Yin, Zhishuang Wang, Chengyi Xia, Matthias Dehmer, Frank Emmert-Streib, and Zhen Jin. A novel epidemic model considering demographics and intercity commuting on complex dynamical networks. Applied Mathematics and Computation, 386 (??):Article 125517, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304756.

Yuan:2020:CMG

[YWX20b] Bo-Jun Yuan, Yi Wang, and Jing Xu. Characterizing the mixed graphs with exactly one positive eigenvalue and its application to mixed graphs determined by their H-spectra. Applied Mathematics and Computation, 380(??):Article 125279, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320302484.

Yu:2021:FSO

[YWXL21] Qian Yu, Kunyang Wang, Binhu Xia, and Yibao Li. First and second order unconditionally energy stable schemes for topology optimization based on phase field method. Applied Mathematics and Computation, 405(??):Article 126267, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003568.

Yan:2022:FTE

[YWY^{+}22] Yan Yan, Libing Wu, Weijun Yan, Yuhan Hu, Nannan Zhao, and Ming Chen.

Finite-time event-triggered fault-tolerant control for a family of pure-feedback systems. Applied Mathematics and Computation, 426(??): ??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001710.

Yang:2023:STS

[YWZ23] Xuehua Yang, Lijiao Wu, and Haixiang Zhang. A space-time spectral order sinc-collocation method for the fourth-order nonlocal heat model arising in viscoelasticity. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003612.

Yang:2020:CAS

[YWZT20] Yin Yang, Jindi Wang, Shangyou Zhang, and Emran Tohidi. Convergence analysis of space-time Jacobi spectral collocation method for solving timefractional Schrödinger equations. Applied Mathematics and Computation, 387(??): Article 124489, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319304680.

Ye:2021:SSC

[YXAY21] Shuai Ye, Xinhai Xu, Hengbin An, and Xuejun Yang. A supplementary strategy for coarsening in algebraic multigrid. Applied Mathematics and Computation, 394 (??):Article 125795, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320307487.

Yang:2023:SPE

[YXFL23] Da-Wei Yang, Zihao Xu, Yan-Quan Feng, and Jaeun Lee. Symmetric property and edge-disjoint Hamiltonian cycles of the spined cube. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002448.

Yang:2022:ETO

[YXMW22] Yang Yang, Xiaorui Xi, Songtao Miao, and Jinran Wu. Event-triggered output feedback containment control for a class of stochastic nonlinear multi-agent systems. Applied Mathematics and Computation, 418 (??):Article 126817, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009000.

Yao:2023:CGC

[YXW23] Lingling Yao, Dongmei Xie, and Han Wang. Complex group consensus of multi-agent systems with cooperativecompetitive mechanisms: acyclic partition method. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008384.

Yu:2022:EAG

[YXX^{+}22] Guihai Yu, Chao Xiong, Jianxiong Xiao, Deyan He, and Gang Peng. Evolutionary analysis of the global rare earth trade networks. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200323X.

Ye:2023:ELE

[YXXZ23] Zuyang Ye, Yuxuan Xiong, Feng Xiong, and Qihua Zhang. An equivalent line element approach for free surface flow through two-dimensional rock mass including fracture networks. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002503.

Yang:2023:DAF

[YXYZ23] Wenjing Yang, Jianwei Xia, Miao Yu, and Na Zhang. Decentralized adaptive funnel control of uncertain large-scale interconnected nonlinear system. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007627.

Yin:2021:MDE

[YXZ^{+}21] Fulian Yin, Xinyu Xia, Xiaojian Zhang, Mingjia Zhang, Jiahui Lv, and Jianhong Wu. Modelling the dynamic emotional information propagation and guiding the public sentiment in the Chinese Sina-microblog. Applied Mathematics and Computation, 396(??):Article 125884, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308377.

Ye:2024:DFT

[YXZ^{+}24] Zehua Ye, Xinran Xu, Dan Zhang, Jun Cheng, and Huaicheng Yan. Distributed finite-time secure filtering for T –S fuzzy systems under hybrid cyberattacks: Application to tunnel diode circuits. Applied Mathematics and Computation, 480(??):??, November 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003618.

Yang:2020:TVF

[YY20a] Haijiao Yang and Dan Ye. Time-varying formation tracking control for highorder nonlinear multi-agent systems in fixed-time framework. Applied Mathematics and Computation, 377 (??):Article 125119, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300886.

Ye:2020:PQM

[YY20b] Taofeng Ye and Huiqiang Yang. Price and quality management with strategic consumers: Whether to introduce a high or low product variant. Applied Mathematics and Computation, 386(??):Article 125541, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304975.

Yamanaka:2021:ONT

[YY21a] Yusuke Yamanaka and Naoto Yamaoka. Oscillation and nonoscillation theorems for Meissner’s equation. Applied Mathematics and Computation, 388 (??):Article 125526, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304847.

Yang:2021:IMT

[YY21b] Chengzhuan Yang and Qian Yu. Invariant multiscale triangle feature for shape recognition. Applied Mathematics and Computation, 403(??):Article 126096, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001442.

Yu:2021:SDD

[YY21c] Yu Yu and Weisheng Yu.

The stability and duality of dynamic Cournot and Bertrand duopoly model with comprehensive preference. Applied Mathematics and Computation, 395(??): Article 125852, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308055.

Yuzbasi:2022:CMS

[YY22] Suayip Yüzbasi and Gamze Yildirim. A collocation method to solve the parabolic-type partial integrodifferential equations via Pell–Lucas polynomials. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200042X.

Yang:2024:SMM

[YY24] Qianxi Yang and Yanlong Yang. A social monitoring mechanism for third-party judges promotes cooperation in evolutionary games. Applied Mathematics and Computation, 483(??):??, December 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004521.

Yao:2024:LTD

[YYF24] Zichen Yao, Zhanwen Yang, and Yongqiang Fu. Long time decay analysis of complex-valued fractional abstract evolution equations with delay. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004617.

Yan:2020:DBE

[YYG20] Shen Yan, Fan Yang, and Zhou Gu. Derivative-based event-triggered control for networked systems with quantization. Applied Mathematics and Computation, 383(??):Article 125359, October 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303234.

You:2023:MVM

[YYL^{+}23] Feng You, Han-Xin Yang, Yumeng Li, Wenbo Du, and Gang Wang. A modified Vicsek model based on the evolutionary game. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006397.

Yang:2020:GOP

[YYW^{+}20] Xiaoxia Yang, Xiaoli Yang, Qianling Wang, Yuanlei Kang, and Fuquan Pan. Guide optimization in pedestrian emergency evacuation. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307039.

You:2023:CBU

[YYW^{+}23] Tao You, Haochun Yang, Jian Wang, Peng Zhang, Jinchao Chen, and Ying Zhang. Cooperative behavior under the influence of multiple experienced guiders in Prisoner’s dilemma game. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004034.

You:2023:FTS

[YYWL23] Luyao You, Xueyan Yang, Shuchen Wu, and Xiaodi Li. Finite-time stabilization for uncertain nonlinear systems with impulsive disturbance via aperiodic intermittent control. Applied Mathematics and Computation, 443(??):??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008505.

Yu:2021:ETI

[YZ21] Nanxiang Yu and Wei Zhu. Event-triggered impulsive chaotic synchronization of fractional-order differential systems. Applied Mathematics and Computation, 388(??):Article 125554, January 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320305105.

Yuan:2023:FSS

[YZ23] Xiaolan Yuan and Yusheng Zhou. The fastest stabilization of second-order switched systems with all modes unstable via an optimal state-dependent switching rule revisited. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003648.

Yuan:2024:ETA

[YZ24] Guangxia Yuan and Zhengqiang Zhang. Event-triggered adaptive prescribed performance tracking for nonlinear time-varying systems with unknown control directions. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005283.

Yao:2020:EFG

[YZAM20] Zhao Yao, Ping Zhou, Ahmed Alsaedi, and Jun Ma. Energy flow-guided synchronization between chaotic circuits. Applied Mathematics and Computation, 374(??):Article 124998, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309907.

Yan:2022:FTA

[YZC^{+}22] Zhiguo Yan, Min Zhang, Gaizhen Chang, Hui Lv, and Ju H. Park. Finitetime annular domain stability and stabilization of Itô stochastic systems with Wiener noise and Poisson jumps-differential Gronwall inequality approach. Applied Mathematics and Computation, 412 (??):Article 126589, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006731.

Yu:2022:APA

[YZDC22] Zhefeng Yu, Feng Zhao, Shihong Ding, and Xiangyong Chen. Adaptive preassigned finite-time control of uncertain nonlinear systems with unknown control gains. Applied Mathematics and Computation, 417(??): Article 126784, March 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008663.

Yang:2020:LRT

[YZJ^{+}20] Jing-Hua Yang, Xi-Le Zhao, Teng-Yu Ji, Tian-Hui Ma, and Ting-Zhu Huang. Lowrank tensor train for tensor robust principal component analysis. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307751.

Yan:2020:TNL

[YZLS20] Jingye Yan, Hong Zhang, Ziyuan Liu, and Songhe Song. Two novel linearimplicit momentum-conserving schemes for the fractional Korteweg–de Vries equation. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307374.

Yang:2024:LFT

[YZM24] Yayu Yang, Mingzu Zhang, and Jixiang Meng. Link fault tolerance of BC networks and folded hypercubes on h-extra rcomponent edge-connectivity. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300512X.

Yang:2024:ISL

[YZPL24] Zhengzhi Yang, Lei Zheng, Matjaz Perc, and Yumeng Li. Interaction state Qlearning promotes cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005337.

Yang:2021:IHA

[YZT^{+}21] Qiwen Yang, Xuzhen Zhu, Yang Tian, Guanglu Wang, Yuexia Zhang, and Lei Chen. The influence of heterogeneity of adoption thresholds on limited information spreading. Applied Mathematics and Computation, 411 (??):Article 126448, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005373.

Yue:2020:ANT

[YZW20] Jun Yue, Yiping Zhu, and Meiqin Wei. The annihilation number and the total domination number of a tree-like graph. Applied Mathematics and Computation, 380(??):Article 125240, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320302095.

Yang:2022:AMD

[YZY22a] Luhe Yang, Lianzhong Zhang, and Duoxing Yang. Asymmetric micro-dynamics in spatial anonymous public goods game. Applied Mathematics and Computation, 415(??):Article 126737, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008195.

Yuan:2022:AOF

[YZY22b] Manman Yuan, Junyong Zhai, and Hui Ye. Adaptive output feedback control for a class of switched stochastic nonlinear systems via an event-triggered strategy. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000595.

You:2022:IEG

[YZZ^{+}22] Tao You, Hailun Zhang, Ying Zhang, Qing Li, Peng Zhang, and Mei Yang. The influence of experienced guider on cooperative behavior in the Prisoner’s dilemma game. Applied Mathematics and Computation, 426(??):??, August 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001771.

Zhang:2021:CCE

[ZAD21] Boyu Zhang, Xinmiao An, and Yali Dong. Conditional cooperator enhances institutional punishment in public goods game. Applied Mathematics and Computation, 390 (??):Article 125600, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305555.

Zarei:2021:ROM

[Zar21]  M. Zarei.  On a re-

duced order modeling of the nuclear reactor dynamics. Applied Mathematics and Computation, 393(??): Article 125819, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307724.

Zhu:2022:ETS

[ZB22] Sha Zhu and Haibo Bao. Event-triggered synchronization of coupled memristive neural networks.

Applied Mathematics and Computation, 415(??):Article 126715, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007992.

Zhang:2023:CSU

[ZB23] Wei Zhang and Ulrik Brandes. Is cooperation sustained under increased mixing in evolutionary public goods games on networks? Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006774.

Zwick:2023:CRD

[ZBA^{+}23a] B. F. Zwick, G. C. Bourantas, F. Alkhatib, A. Wittek, and K. Miller. Corrigendum to Recovery by discretization corrected particle strength exchange (DC PSE) operators’: Applied Mathematics and Computation 448 (2023) 127923. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002163.

See [ZBA^{+}23b].

Zwick:2023:RDC

[ZBA^{+}23b] B. F. Zwick, G. C. Bourantas, F. Alkhatib, A. Wittek, and K. Miller. Recovery by discretization corrected particle strength exchange (DC PSE) operators. Applied Mathematics and Computation, 448(??): ??, July 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000929.

See corrigendum [ZBA^{+}23a].

Zhou:2023:NQG

[ZBBK23] Jun Zhou, Pieter J. Barendrecht, Michael Barton, and Jirí Kosinka. Numerical quadrature for Gregory quads. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002205.

Zhao:2022:MPP

[ZBFW22] Daping Zhao, Lin Bai, Yong Fang, and Shouyang Wang. Multi-period portfolio selection with investor views based on scenario tree. Applied Mathematics and Computation, 418 (??):Article 126813, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008961.

Zhou:2023:DCU

[ZBL^{+}23] Xin Zhou, Adam Belloum, Michael H. Lees, Tom van Engers, and Cees de Laat. The dynamics of corruption under an optional external supervision service. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003417.

Zhang:2021:CGM

[ZC21a] Zemian Zhang and Xuesong Chen. A conjugate gradient method for distributed optimal control problems with nonhomogeneous Helmholtz equation. Applied Mathematics and Computation, 402(??):Article 126019, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321000679.

Zhou:2021:NFH

[ZC21b] Yaoyao Zhou and Gang Chen. Non-fragile H_{∞} finite-time sliding mode control for stochastic Markovian jump systems with time delay. Applied Mathematics and Computation, 409(??):Article 126383, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004720.

Zhou:2021:OSD

[ZC21c] Yusheng Zhou and Danhong Chen. Optimized state-dependent switching law design for a class of switched nonlinear systems with two unstable subsystems. Applied Mathematics and Computation, 397 (??):Article 125872, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308250.

Zlotnik:2022:HOC

[ZC22] Alexander Zlotnik and Raimondas Ciegis. On higherorder compact ADI schemes for the variable coefficient wave equation. Applied Mathematics and Computation, 412(??):Article 126565, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321006494.

Zhang:2023:NRF

[ZC23a] Guodong Zhang and Jinde Cao. New results on fixed/predefined-time synchronization of delayed fuzzy inertial discontinuous neural networks: Nonreduced order approach. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007391.

Zorica:2023:DGF

[ZC23b] Dusan Zorica and Stevan M. Cvetićanin. Dissipative and generative fractional RLC circuits in the transient regime. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300396X.

Zhang:2023:NOB

[ZCD23] Lei Zhang, Wei-Wei Che, and Jun-Hang Ding. NNsobserver-based fully distributed consensus control for MASs under deception attacks. Applied Mathematics and Computation, 456(??):??, November 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003090.

Zhang:2020:FDS

[ZCH20] Shuai Zhang, Ruaridh Clark, and Yunke Huang. Frequency-dependent strategy selection in a hunting game with a finite population. Applied Mathematics and Computation, 382(??):Article 125355, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320303192.

Zhu:2022:DET

[ZCJ22] Lin Zhu, Wei-Wei Che, and Xiao-Zheng Jin. Dynamic event-triggered tracking control for model-free networked control systems. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004234.

Zhang:2021:CNZ

[ZCJL21] Jiannan Zhang, Ping Chen, Zhuo Jin, and Shuanming Li. On a class of nonzero-sum stochastic differential dividend games with regime switching. Applied Mathematics and Computation, 397(??):Article 125956, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000047.

Zhang:2023:IDE

[ZCL^{+}23] Qiongwen Zhang, Jun Cheng, Daixi Liao, Jinde Cao, and Fawaz E. Alsaadi. Improved dynamic event-triggered control for nonlinear systems with fading channels. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001595.

Zhang:2022:RML

[ZCLW22] Aili Zhang, Ping Chen, Shuanming Li, and Wenyuan Wang. Risk modelling on liquidations with Lévy processes. Applied Mathematics and Computation, 412 (??):Article 126584, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006688.

Zhao:2022:OBD

[ZCLZ22] Guangtong Zhao, Liang Cao, Xiaomeng Li, and Qi Zhou. Observer-based dynamic event-triggered control for nonstrict-feedback stochastic nonlinear multiagent systems. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003630.

Zhang:2024:MWI

[ZCM^{+}24] Shan Zhang, Xun Chen, Zhen-Wei Ma, Xiao-Dong Zhang, and Ya-Hong Chen. The minimum Wiener index of unicyclic graphs with maximum degree. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000535.

Zhang:2020:RCS

[ZCPZ20] Bao-Lin Zhang, Luhua Cheng, Kejia Pan, and Xian-Ming Zhang. Reducing conservatism of stability criteria for linear systems with time-varying delay using an improved tripleintegral inequality. Applied Mathematics and Computation, 380(??):Article 125254, September 1, 2020.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030223X.

Zhan:2021:LBM

[ZCS21] Chengjie Zhan, Zhenhua Chai, and Baochang Shi. A lattice Boltzmann model for the coupled cross-diffusion-fluid system. Applied Mathematics and Computation, 400(??):Article 126105, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321001533.

Zhang:2022:SOF

[ZCS^{+}22] Ziwei Zhang, Zongjie Chen, Zhang Sheng, Dan Li, and Jing Wang. Static output feedback secure synchronization control for Markov jump neural networks under hybrid cyber-attacks. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003484.

Zhang:2022:SLS

[ZCT22] Xiao-Qin Zhang, Shu-Yu Cui, and Gui-Xian Tian. Signless Laplacian state transfer on Q -graphs. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001369.

Zhang:2024:ETC

[ZCW^{+}24a] Dan Zhang, Xuefeng Chen, Hui Wang, Hangli Ren, and Huiling Xu. Eventtriggered H_{∞} control for networked dynamic system with nonideal interconnection. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005428.

Zhang:2024:SCP

[ZCW24b] Ganchao Zhang, Min Chen, and Weifan Wang. A sufficient condition for planar graphs with girth 5 to be (1, 6)-colorable. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001784.

Zhang:2022:ASM

[ZCZ22] Xuefeng Zhang, Shunan Chen, and Jin-Xi Zhang. Adaptive sliding mode consensus control based on neural network for singular fractional order multiagent systems. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005161.

Zhou:2023:DET

[ZCZW23] Xin Zhou, Guici Chen, Song Zhu, and Shiping Wen. Distributed event-triggered finite-time H_{∞} filtering for switched systems on sensor networks with two-channel network attacks and asynchronous modes. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004009.

Zhang:2023:AAT

[ZD23] Haoyue Zhang and Shihong Ding. Adaptive asymptotic tracking control design for high-order uncertain nonlinear systems. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032300036X.

Zhu:2024:OBC

[ZD24] Fanglai Zhu and Wenqing Du. Observer-based consensus of multi-agent systems under odd distributed impulsive control protocol. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006434.

Zhang:2022:RDT

[ZDC22] Weijian Zhang, Haibo Du, and Zhaobi Chu. Robust discrete-time nonsmooth consensus protocol for multi-agent systems via super-twisting algorithm. Applied Mathematics and Computation, 413(??):Article 126636, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321007207.

Zou:2022:DOF

[ZDD^{+}22] Ying Zou, Chao Deng, Lu Dong, Lei Ding, and Ming Lu. Distributed output feedback consensus tracking control of multiple nonholonomic mobile robots with only position information of leader. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000480.

Zeng:2021:SDR

[ZDGL21] Pengyu Zeng, Feiqi Deng, Xiaobin Gao, and Xiaohua Liu. Sampled-data resilient H_{∞} control for networked stochastic systems subject to multiple attacks. Applied Mathematics and Computation, 405(??): Article 126265, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003544.

Zhu:2020:FPO

[ZDKC20] Xiaomin Zhu, Fangfang Dou, Andreas Karageorghis, and C. S. Chen. A fictitious points one-step MPS–MFS technique. Applied Mathematics and Computation, 382(??):Article 125332, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320302988.

Zhu:2023:COR

[ZDMW23] Yanhong Zhu, Shaofei Du, Xuesong Ma, and Furong Wang. A class of orientably-regular embeddings of graphs of order p^{3}. Applied Mathematics and Computation, 442(??): ??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007755.

Zhou:2021:ISS

[ZDX21] Lei Zhou, Hui Ding, and Xiaoqing Xiao. Input-to-state stability of discretetime switched nonlinear systems with generalized switching signals. Applied Mathematics and Computation, 392(??):Article 125727, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320306809.

Zhang:2024:RSA

[ZDZ24] Huasheng Zhang, Yuzhen Dai, and Chenglong Zhu. Region stability analysis and precise tracking control of linear stochastic systems. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005714.

Zhang:2021:COR

[ZDZS21] Juan Zhang, Jing Dai, Huaguang Zhang, and Shaoxin Sun. Cooperative output regulation of heterogeneous linear multi-agent systems based on the eventtriggered distributed control under switching topologies. Applied Mathematics and Computation, 390 (??):Article 125611, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032030566X.

Zhu:2022:PAA

[ZEA22] Yichen Zhu and Marcos Escobar-Anel. Polynomial affine approach to HARA utility maximization with applications to Ornstein–Uhlenbeck 4/2 models. Applied Mathematics and Computation, 418 (??):Article 126836, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100919X.

Zhu:2021:SSP

[ZeF21] Sanmei Zhu and Jun e Feng. The set stabilization problem for Markovian jump Boolean control networks: an average optimal control approach. Applied Mathematics and Computation, 402(??):Article 126133, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321001818.

Zhao:2023:DDB

[ZeFWD23] Rong Zhao, Jun e Feng, Biao Wang, and Renato De Leone. Disturbance decoupling of Boolean networks via robust indistinguishability method. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003892.

Zhang:2020:ETM

[ZeFWW20] Qiliang Zhang, Jun e Feng, Biao Wang, and Peihe Wang. Event-triggered mechanism of designing set stabilization state feedback controller for switched Boolean networks. Applied Mathematics and Computation, 383(??):Article 125372, October 15, 2020.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303362.

Zeng:2021:RLI

[Zen21] Min-Li Zeng. The RSSlike iteration method for block two-by-two linear systems from time-periodic parabolic optimal control problems. Applied Mathematics and Computation, 410(??):Article 126477, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100566X.

Zhang:2021:PBS

[ZF21] Xi Zhang and Minfu Feng.

A projection-based stabilized virtual element method for the unsteady incompressible Brinkman equations. Applied Mathematics and Computation, 408(??):Article 126325, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004148.

Zhang:2023:LPS

[ZF23a] Yadong Zhang and Minfu Feng. A local projection stabilization virtual element method for the time-fractional Burgers equation with high Reynolds numbers. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005835.

Zheng:2023:VPM

[ZF23b] Sanpeng Zheng and Renzhong Feng. A variable projection method for the general radial basis function neural network. Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001789.

Zhan:2024:SAE

[ZF24] Rui Zhan and Jinwei Fang. Stability analysis of explicit exponential Rosenbrock methods for stiff differential equations with constant delay. Applied Mathematics and Computation, 482(??):??, December 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324004399.

Zhang:2020:IRR

[ZFJ^{+}20] Liang Zhang, Zhiguang Feng, Zhengyi Jiang, Ning Zhao, and Yang Yang. Improved results on reachable set estimation of singular systems. Applied Mathematics and Computation, 385 (??):Article 125419, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303805.

Zhang:2022:RFM

[ZFS22] Jiali Zhang, Zhi-Wei Fang, and Hai-Wei Sun. Robust fast method for variableorder time-fractional diffusion equations without regularity assumptions of the true solutions. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003472.

Zhang:2024:MLC

[ZG24] Wenbo Zhang and Wei Gu. Machine learning for a class of partial differential equations with multidelays based on numerical Gaussian processes. Applied Mathematics and Computation, 467(??):??, April 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006677.

Zhao:2022:RSR

[ZGCZ22] Bin Zhao, Pengbo Geng, Wengu Chen, and Zhu Zeng. Robust sparse recovery via a novel convex model. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000091.

Zheng:2022:NSU

[ZGG^{+}22] Yongfeng Zheng, Yan Gu, Liang Gao, Yanzheng Wang, Jinping Qu, and Chuanzeng Zhang. A new structural uncertainty analysis method based on polynomial expansions. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002065.

Zhang:2024:TDV

[ZGHM24] Huan Zhang, Ming Gong, Jingsong He, and Boris Malomed. Two-dimensional vector solitons in Bose– Einstein-condensate mixtures. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000080.

Zhao:2022:SFD

[ZGLZ22] Xiao-Qi Zhao, Shun Guo, Yue Long, and Guang-Xin Zhong. Simultaneous fault detection and control for discrete-time switched systems under relaxed persistent dwell time switching. Applied Mathematics and Computation, 412 (??):Article 126585, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100669X.

Zhang:2022:CDR

[ZGWZ22] Jie Zhang, Daniel Gray, Hua Wang, and Xiao-Dong Zhang. On the combinatorics of derangements and related permutations. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004155.

Zhao:2023:SCS

[ZGZ^{+}23] Younan Zhao, Peng Gu, Fanglai Zhu, Tianyi Liu, and Runjie Shen. Security control scheme for cyber-physical system with a complex network in physical layer against false data injection attacks. Applied Mathematics and Computation, 447(??):??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000772.

Zhang:2024:TFU

[ZGZ^{+}24] Xue Zhang, Xian-Ming Gu, Yong-Liang Zhao, Hu Li, and Chuan-Yun Gu. Two fast and unconditionally stable finite difference methods for Riesz fractional diffusion equations with variable coefficients. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323005040.

Zhang:2020:MSS

[ZH20a] Cun-Hua Zhang and Ye He. Multiple stability switches and Hopf bifurcations induced by the delay in a Lengyel–Epstein chemical reaction system. Applied Mathematics and Computation, 378(??):Article 125201, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301703.

Zhao:2020:EFC

[ZH20b] Longbin Zhao and Chengming Huang. Exponential fitting collocation methods for a class of Volterra integral equations. Applied Mathematics and Computation, 376(??):Article 125121, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300904.

Zhang:2022:LSF

[ZH22a] Rong Zhang and Qiaolin He. The least-square/ fictitious domain method based on Navier slip boundary condition for simulation of flow-particle interaction. Applied Mathematics and Computation, 415 (??):Article 126687, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321007712.

Zheng:2022:BCS

[ZH22b] Jun Zheng and Hanping Hu. Bit cyclic shift method to reinforce digital chaotic maps and its application in pseudorandom number generator. Applied Mathematics and Computation, 420(??): Article 126788, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321008705.

Zhang:2023:PSP

[ZH23a] Lan Zhang and Changwei Huang. Preferential selection to promote cooperation on degree-degree correlation networks in spatial snowdrift games. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002825.

Zhang:2023:RTR

[ZH23b] Shuang Zhang and Le Han. Robust tensor recovery with nonconvex and nonsmooth regularization. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006403.

Zhang:2020:ISV

[Zha20]  Jingjing Zhang.  An

improved Störmer–Verlet method based on exact discretization for nonlinear oscillators. Applied Mathematics and Computation, 386(??):Article 125476, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304355.

Zhang:2021:TCL

[Zha21a]  Gengen Zhang.  Two

conservative and linearlyimplicit compact difference schemes for the nonlinear fourth-order wave equation. Applied Mathematics and Computation, 401 (??):Article 126055, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100103X.

Zhang:2021:MBM

[Zha21b] Li-Li Zhang. A modulusbased multigrid method for nonlinear complementarity problems with application to free boundary problems with nonlinear source terms. Applied Mathematics and Computation, 399(??):

Article 126015, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000631.

Zhang:2021:PIG

[Zha21c] Shuaiqi Zhang. On pathindependent Girsanov transform. Applied Mathematics and Computation, 395(??): Article 125845, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307980.

Zhang:2021:TCA

[Zha21d] Xingong Zhang. Two competitive agents to minimize the weighted total late work and the total completion time. Applied Mathematics and Computation, 406(??):Article 126286, October 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321003751.

Zhang:2022:CCZ

[Zha22a] Gui-Lai Zhang. Convergence, consistency and zero stability of impulsive onestep numerical methods. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001035.

Zhang:2022:MMS

[Zha22b]  Shengliang Zhang.  A

meshless multi-symplectic local radial basis function collocation scheme for the “good” Boussinesq equation. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200371X.

Zhang:2022:DCT

[Zha22c]  Wenjing Zhang.  Dis-

ease clearance of tuberculosis infection: an in-host continuous-time Markov chain model. Applied Mathematics and Computation, 413(??):Article 126614, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006986.

Zhao:2023:CSE

[Zha23a]  Jianxing Zhao.  Condi-

tions of strong ellipticity and calculations of Meigenvalues for a partially symmetric tensor. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004149.

Zhao:2023:LES

[Zha23b] Kaihong Zhao. Local exponential stability of several almost periodic positive solutions for a classical controlled GA-predation ecosystem possessed distributed delays. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006142.

Zhang:2024:ABB

[Zha24a]  Rui Zhang.  Asymmet-

ric beta-binomial GARCH models for time series with bounded support. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324000286.

Zhang:2024:LDP

[Zha24b] Shuxiong Zhang. Large deviation probabilities for the range of a d-dimensional supercritical branching random walk. Applied Mathematics and Computation, 462(??):??, February 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005131.

Zhu:2023:PCA

[ZHFL23a] Wen-Han Zhu, Rong-Xia Hao, Yan-Quan Feng, and Jaeun Lee. The 3path-connectivity of the k-ary n-cube. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005732.

Zhu:2023:PCL

[ZHFL23b] Wen-Han Zhu, Rong-Xia Hao, Yan-Quan Feng, and Jaeun Lee. Path connectivity of line graphs and total graphs of complete bipartite graphs. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300351X.

Zhang:2022:FDU

[ZHG22] Zhi-Hui Zhang, Li-Ying Hao, and Mingjie Guo. Fault detection for uncertain nonlinear systems via recursive observer and tight threshold. Applied Mathematics and Computation, 414(??):Article 126665, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321007499.

Zhou:2021:ESD

[ZHJ^{+}21] Baoquan Zhou, Bingtao Han, Daqing Jiang, Tasawar Hayat, and Ahmed Alsaedi. Ergodic stationary distribution and extinction of a hybrid stochastic SEQIHR epidemic model with media coverage, quarantine strategies and pre-existing immunity under discrete Markov switching. Applied Mathematics and Computation, 410(??):Article 126388, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032100477X.

Zare:2024:CAI

[ZHL24] Farideh Zare, Mohammad Heydari, and Ghasem Barid Loghmani. Convergence analysis of an iterative scheme to solve a family of functional Volterra integral equations. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002601.

Zhang:2021:MAD

[ZHLZ21] Zhi-Gang Zhang, Xiao Hu, Zhao-Ting Liu, and Lu-Tao Zhao. Multi-attribute decision making: an innovative method based on the dynamic credibility of experts. Applied Mathematics and Computation, 393(??): Article 125816, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307694.

Zheng:2021:DHM

[ZHMZ21] Jun Zheng, Hanping Hu, Hao Ming, and Yanxia Zhang. Design of a hybrid model for construction of digital chaos and local synchronization. Applied Mathematics and Computation, 392(??):Article 125673, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300320306263.

Zhou:2020:QRN

[Zho20] Yang Zhou. Quadrature rules with neighborhood of spherical designs on the two-sphere. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319307611.

Zhang:2023:EFD

[ZHQ^{+}23] Xin Zhang, Lintao Huang, Xueyu Qin, Feng Qu, and Chao Yan. An efficient finite difference IFWENO– THINC hybrid scheme for capturing discontinuities. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000589.

Zhang:2024:CES

[ZHQ^{+}24] Yao Zhang, Qing-Yi Hao, Jia-Li Qian, Chao-Yun Wu, Ning Guo, and Xiang Ling. The cooperative evolution in the spatial prisoner’s dilemma game with the local loyalty of two-strategy. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006537.

Zhu:2023:DEE

[Zhu23] Xiaochen Zhu. The dynamic edge environment under interactive diversity is a double-edged sword. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005793.

Zhu:2024:EPR

[Zhu24]  Jinjie Zhu.  Embed-

ding phase reduction for fast-slow systems with noise-induced stochastic quasiperiodic orbits. Applied Mathematics and Computation, 465(??):??, March 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032300591X.

Zhao:2022:IDT

[ZHW22] Shu-Li Zhao, Rong-Xia Hao, and Chao Wei. Internally disjoint trees in the line graph and total graph of the complete bipartite graph. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000765.

Zhao:2024:CCE

[ZHWS24] Qiwen Zhao, Zhixue He, Xiaoyue Wang, and Lei Shi. Cooperation can emerge in the co-evolution of the local environments. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004065.

Zhang:2020:IDN

[ZHYD20] Ying-Nan Zhang, Hong-Qian He, Guo-Fu Yu, and Yi-Jun Dong. Integrable discretizations and numerical simulation for a modified coupled integrable dispersionless equation. Applied Mathematics and Computation, 364(??): ??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306587.

Zou:2024:SPG

[ZHZH24] Kuan Zou, Wenchen Han, Lan Zhang, and Changwei Huang. The spatial public goods game on hypergraphs with heterogeneous investment. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006197.

Zhuang:2024:CMI

[ZHZZ24] Qiao Zhuang, Alfa Heryudono, Fanhai Zeng, and Zhongqiang Zhang. Collocation methods for integral fractional Laplacian and fractional PDEs based on radial basis functions. Applied Mathematics and Computation, 469(??): ??, May 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324000201.

Zalabarria:2020:ORP

[ZIML20] Unai Zalabarria, Eloy Irigoyen, Raquel Martinez, and Andrew Lowe. Online robust R-peaks detection in noisy electrocardiograms using a novel iterative smart processing algorithm. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308318.

Zhu:2023:DMK

[ZJ23] Hongmiao Zhu and Zhen Jin. A dynamics model of knowledge dissemination in a WeChat group from perspective of duplex networks. Applied Mathematics and Computation, 454(??):??, October 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002527.

Zhao:2021:SSP

[ZJC21] Caidi Zhao, Huite Jiang, and Tomás Caraballo. Statistical solutions and piecewise Liouville theorem for the impulsive reaction– diffusion equations on infinite lattices. Applied Mathematics and Computation, 404(??):Article 126103, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009630032100151X.

Zheng:2020:MED

[ZJCA20] Yanpeng Zheng, Xiaoyu Jiang, Xiaoting Chen, and Fawaz Alsaadi. More extensions of a determinant inequality of hartfiel. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308197.

Zhai:2021:CDN

[ZJG21] Yunyun Zhai, Ting Ji, and Xianguo Geng. Coupled derivative nonlinear Schrödinger III equation: Darboux transformation and higher-order rogue waves in a two-mode nonlinear fiber. Applied Mathematics and Computation, 411(??):Article 126551, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006354.

Zhang:2024:IUE

[ZJJ24] Pingrui Zhang, Xiaoyun Jiang, and Junqing Jia. Improved uniform error estimates for the twodimensional nonlinear space fractional Dirac equation with small potentials over long-time dynamics. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006276.

Zhang:2022:IMS

[ZJJB22] Xing Zhang, Xiaoyu Jiang, Zhaolin Jiang, and Heejung Byun. An improvement of methods for solving the CUPL–Toeplitz linear system. Applied Mathematics and Computation, 421(??): ??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322000182.

Zhu:2023:FTA

[ZJLT23] Hao-Yang Zhu, Xiaoyue Jiang, Yuan-Xin Li, and Shaocheng Tong. Finitetime adaptive fuzzy output tracking of switched nonlinear systems with ISD-ADT. Applied Mathematics and Computation, 443(??): ??, April 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322008736.

Zhao:2022:NNB

[ZJW^{+}22] Zhi-Ye Zhao, Xiao-Zheng Jin, Xiao-Ming Wu, Hai Wang, and Jing Chi. Neural network-based fixedtime sliding mode control for a class of nonlinear Euler–Lagrange systems. Applied Mathematics and Computation, 415 (??):Article 126718, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100802X.

Zhang:2024:SVD

[ZJWV24] Dong Zhang, Tongsong Jiang, Gang Wang, and V. I. Vasil’ev. On singular value decomposition and generalized inverse of a commutative quaternion matrix and applications. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004605.

Zhao:2021:CBB

[ZJX21] Jingjun Zhao, Xingzhou Jiang, and Yang Xu. Convergence of block boundary value methods for solving delay differential algebraic equations with index-1 and index-2. Applied Mathematics and Computation, 399(??):Article 126034, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000825.

Zhao:2022:FDM

[ZJX22a] Jingjun Zhao, Xingzhou Jiang, and Yang Xu. Fully discretized methods based on boundary value methods for solving diffusion equations. Applied Mathematics and Computation, 418 (??):Article 126848, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009310.

Zhao:2022:KGB

[ZJX22b] Jingjun Zhao, Xingzhou Jiang, and Yang Xu. A kind of generalized backward differentiation formulae for solving fractional differential equations. Applied Mathematics and Computation, 419(??):Article 126872, April 15, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009553.

Zhang:2020:SSF

[ZJZD20] Di Zhang, Yuanwei Jing, Qingling Zhang, and Georgi M. Dimirovski. Stabilization of singular T-S fuzzy Markovian jump system with mode-dependent derivativeterm coefficient via sliding mode control. Applied Mathematics and Computation, 364(??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306356.

Zhang:2022:WMO

[ZK22] Zhiguo Zhang and Mark A.

Kon. Wavelet matrix operations and quantum transforms. Applied Mathematics and Computation, 428(??):??, September 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322002533.

Zheng:2023:MDE

[ZKLL23] Xiaoyuan Zheng, Yu Kang, Hongyi Li, and Jitao Li. Multiple description encoding-decoding-based resilient filtering for complex networks under the round-robin protocol. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004198.

Zhang:2020:CFE

[ZL20a] Jin Zhang and Xiaowei Liu. Convergence of a finite element method on a Bakhvalov-type mesh for singularly perturbed reaction-diffusion equation. Applied Mathematics and Computation, 385(??):Article 125403, November 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303659.

Zhang:2020:DES

[ZL20b] Maozhu Zhang and Kun Li. Dependence of eigenvalues of Sturm–Liouville problems with eigenparameter dependent boundary conditions. Applied Mathematics and Computation, 378(??):Article 125214, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301831.

Zhang:2020:AEF

[ZL20c] Tao Zhang and Xiaolin Li. Analysis of the elementfree Galerkin method with penalty for general secondorder elliptic problems. Applied Mathematics and Computation, 380(??):Article 125306, September 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302721.

Zhang:2020:NSO

[ZL20d] Yan Zhang and Yang Liu. Nonlinear second-order multi-agent systems subject to antagonistic interactions without velocity constraints. Applied Mathematics and Computation, 364 (??):??, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306599.

Zhang:2021:HOF

[ZL21a] Jin Zhang and Yanhui Lv.

High-order finite element method on a Bakhvalovtype mesh for a singularly perturbed convectiondiffusion problem with two parameters. Applied Mathematics and Computation, 397(??):Article 125953, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000011.

Zhang:2021:CSM

[ZL21b] Yanjun Zhang and Hanyu Li. A count sketch maximal weighted residual Kaczmarz method for solving highly overdetermined linear systems. Applied Mathematics and Computation, 410(??):Article 126486, December 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005750.

Zhu:2021:RSS

[ZL21c] Zhen Zhu and Jun-Guo Lu. Robust stability and stabilization of hybrid fractionalorder multi-dimensional systems with interval uncertainties: an LMI approach. Applied Mathematics and Computation, 401 (??):Article 126075, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001235.

Zeng:2022:CNM

[ZL22a] Yihui Zeng and Zhendong Luo. The Crank–Nicolson mixed finite element method for the improved system of time-domain Maxwell’s equations. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004969.

Zhang:2022:UCW

[ZL22b] J. Zhang and X. Liu. Uniform convergence of a weak Galerkin finite element method on Shishkin mesh for singularly perturbed convection–diffusion problems in 2D. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004209.

Zhang:2022:PSB

[ZL22c] Li Zhang and Shuai Liu.

Projected subgradient based distributed convex optimization with transmission noises. Applied Mathematics and Computation, 418 (??):Article 126794, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008766.

Zhang:2022:LTR

[ZL22d] Zhi-Yong Zhang and Cheng-Bao Liu. Leibniz-type rule of variable-order fractional derivative and application to build Lie symmetry framework. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003423.

Zhu:2022:RSE

[ZL22e] Xingao Zhu and Shutang Liu. Reachable set estimation for continuoustime impulsive switched nonlinear time-varying systems with delay and disturbance. Applied Mathematics and Computation, 420(??): Article 126910, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009930.

Zhang:2023:RPG

[ZL23a] Qinglong Zhang and Shuzhi Liu. The Riemann problem and a Godunov-type scheme for a traffic flow model on two lanes with two velocities. Applied Mathematics and Computation, 436(??):??, January 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005768.

Zhang:2023:SBR

[ZL23b] Yanjun Zhang and Hanyu Li. Splitting-based randomized iterative methods for solving indefinite least squares problem. Applied Mathematics and Computation, 446(??): ??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000619.

Zhang:2023:GTD

[ZL23c] Yuke Zhang and Huiqiu Lin. Graphs with three distinct distance eigenvalues. Applied Mathematics and Computation, 445(??): ??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000176.

Zhang:2021:GLS

[ZLA^{+}21] Run-Fa Zhang, Ming-Chu Li, Mohammed Albishari, Fu-Chang Zheng, and Zhong-Zhou Lan. Generalized lump solutions, classical lump solutions and rogue waves of the (2 + 1)-dimensional Caudrey– Dodd–Gibbon–Kotera–Sawadalike equation. Applied Mathematics and Computation, 403(??):Article 126201, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002915.

Zhao:2020:GES

[ZLC20] Yongshun Zhao, Xiaodi Li, and Jinde Cao. Global exponential stability for impulsive systems with infinite distributed delay based on flexible impulse frequency. Applied Mathematics and Computation, 386(??):Article 125467, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304227.

Zhang:2024:DOF

[ZLC24a] Man Zhang, Chong Lin, and Bing Chen. Dynamic output feedback control for CPSs under transmission delays and sparse malicious attacks. Applied Mathematics and Computation, 473(??): ??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324001267.

Zhou:2024:FTR

[ZLC^{+}24b] Qianru Zhou, Hai Liu, Baolei Cheng, Yan Wang, Yuejuan Han, and Jianxi Fan. Fault tolerance of recursive match networks based on g-good-neighbor fault pattern. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323004873.

Zhang:2022:SSE

[ZLD22] Ge Zhang, Zhiming Li, and Anwarud Din. A stochastic SIQR epidemic model with Lévy jumps and three-time delays. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004039.

Zhang:2022:APB

[ZLDY22] Liming Zhang, Haihong Li, Qionglin Dai, and Junzhong Yang. Adaptive persistence based on environment comparison enhances cooperation in evolutionary games. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321009954.

Zha:2022:ESB

[ZLF22] Jiajing Zha, Cong Li, and Suohai Fan. The effect of stability-based strategy updating on cooperation in evolutionary social dilemmas. Applied Mathematics and Computation, 413 (??):Article 126603, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006871.

Zhong:2021:IDW

[ZLH21] Yue Zhong, Shaoyong Lai, and Chunhua Hu. Investigations to the dynamics of wealth distribution in a kinetic exchange model. Applied Mathematics and Computation, 404(??):Article 126231, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321003210.

Zheng:2020:GAS

[ZLJ20] Minling Zheng, Fawang Liu, and Zhengmeng Jin. The global analysis on the spectral collocation method for time fractional Schrödinger equation. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319306812.

Zheng:2021:SAL

[ZLJ^{+}21] Fuqi Zheng, Jian Liu, Ticao Jiao, Cunshan Zhang, Jianrong Zhao, Jishun Jiang, and Xiaomei Qi. Stability analysis and L_{1}- gain characterization for uncertain Markovian hybrid switching positive systems. Applied Mathematics and Computation, 411 (??):Article 126503, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005920.

Zhang:2022:GML

[ZLK^{+}22] Xiao-Li Zhang, Hong-Li Li, Yonggui Kao, Long Zhang, and Haijun Jiang. Global Mittag-Leffler synchronization of discretetime fractional-order neural networks with time delays. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200491X.

Zhang:2020:PMH

[ZLL20] Yi Zhang, Mei Lu, and Ke Liu. Perfect matching and Hamilton cycle decomposition of complete balanced (k + 1)- partite k-uniform hypergraphs. Applied Mathematics and Computation, 386(??):Article 125492, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320304501.

Zhang:2021:FTO

[ZLLL21] Anguo Zhang, Lulu Li, Yuanyuan Li, and Jianquan Lu. Finite-time output tracking of probabilistic Boolean control networks. Applied Mathematics and Computation, 411 (??):Article 126413, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321005026.

Zou:2022:ETS

[ZLLX22] Cong Zou, Bing Li, Feiyang Liu, and Bingrui Xu. Event-triggered μ-state estimation for Markovian jumping neural networks with mixed time-delays. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001424.

Zhang:2022:SME

[ZLLY22] Huiting Zhang, Lina Liu, Hao Liu, and Yongxin Yuan. The solution of the matrix equation AXB = D and the system of matrix equations AX = C, XB = D with X^{∗}X = I_{p}. Applied Mathematics and Computation, 418(??):Article 126789, April 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321008717.

Zhao:2020:EWT

[ZLN20] Qiuyue Zhao, Shutang Liu, and Xinglong Niu. Effect of water temperature on the dynamic behavior of phytoplankton-zooplankton model. Applied Mathematics and Computation, 378(??):Article 125211, August 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320301806.

Zhu:2021:ACB

[ZLP21] Fengzeng Zhu, Xu Liu, and Li Peng. Adaptive consensus-based distributed H_{∞} filtering with switching topology subject to partial information on transition probabilities. Applied Mathematics and Computation, 411(??):Article 126534, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006172.

Zhao:2022:PSC

[ZLP^{+}22] Dandan Zhao, Runchao Li, Hao Peng, Ming Zhong, and Wei Wang. Percolation on simplicial complexes. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004040.

Zhang:2021:IOB

[ZLS21] Tu Zhang, Liwei Li, and Mouquan Shen. Interval observer-based finite-time control for linear parametervarying systems. Applied Mathematics and Computation, 411(??):Article 126478, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321005671.

Zhu:2023:SEQ

[ZLS23] Lin Zhu, Nabing Liu, and Qin Sheng. A simulation expressivity of the quenching phenomenon in a twosided space-fractional diffusion equation. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005975.

Zubair:2021:HDS

[ZLU21] Tamour Zubair, Tiao Lu, and Muhammad Usman. Higher dimensional semirelativistic time-fractional Vlasov–Maxwell code for numerical simulation based on linear polarization and 2D Landau damping instability. Applied Mathematics and Computation, 401 (??):Article 126100, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100148X.

Zhao:2021:SCG

[ZLW21a] Zhihui Zhao, Hong Li, and Jing Wang. The study of a continuous Galerkin method for Sobolev equation with space-time variable coefficients. Applied Mathematics and Computation, 401 (??):Article 126021, July 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000692.

Zhu:2021:ARE

[ZLW^{+}21b] Xuzhen Zhu, Yuxin Liu, Shengfeng Wang, Ruijie Wang, Xiaolong Chen, and Wei Wang. Allocating resources for epidemic spreading on metapopulation networks. Applied Mathematics and Computation, 411 (??):Article 126531, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006159.

Zhang:2023:SFB

[ZLW23] Aili Zhang, Shuanming Li, and Wenyuan Wang. A scale function based approach for solving integraldifferential equations in insurance risk models. Applied Mathematics and Computation, 450(??):??, August 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323001340.

Zhang:2020:NBS

[ZLX20] Shuo Zhang, Lu Liu, and Dingyu Xue. Nyquist-based stability analysis of noncommensurate fractionalorder delay systems. Applied Mathematics and Computation, 377(??):Article 125111, July 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320300801.

Zhu:2020:DDE

[ZLZ20] Linhe Zhu, Wenshan Liu, and Zhengdi Zhang. Delay differential equations modeling of rumor propagation in both homogeneous and heterogeneous networks with a forced silence function. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309178.

Zhao:2021:LFC

[ZLZ^{+}21] Can Zhao, Xinzhi Liu, Shouming Zhong, Kaibo Shi, Daixi Liao, and Qishui Zhong. Leader-following consensus of multi-agent systems via novel sampleddata event-triggered control. Applied Mathematics and Computation, 395(??): Article 125850, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308031.

Zhao:2023:QSD

[ZLZ^{+}23a] Mingfang Zhao, Hong-Li Li, Long Zhang, Cheng Hu, and Haijun Jiang. Quasi-synchronization of discrete-time fractionalorder quaternion-valued memristive neural networks with time delays and uncertain parameters. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323002643.

Zuo:2023:EEV

[ZLZ^{+}23b] Chao Zuo, Yuting Ling, Fenping Zhu, Xinyu Ma, and Guochun Xiang. Exploring epidemic voluntary vaccinating behavior based on information-driven decisions and benefit-cost analysis. Applied Mathematics and Computation, 447(??): ??, June 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323000747.

Zhang:2024:RAS

[ZLZ24a] Jia-Rui Zhang, Jun-Guo Lu, and Qing-Hao Zhang. Robust asymptotic stability analysis for fractionalorder systems with commensurate time delays: the 1 < β ≤ 2 case. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002285.

Zhang:2024:OBC

[ZLZ24b] Xuefeng Zhang, Yuanda Lv, and Jin-Xi Zhang. H_{∞} observer-based controller synthesis for fractional order systems over finite frequency range. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001681.

Zhang:2020:HHO

[ZM20] Yongchao Zhang and Liquan Mei. A hybrid high-order method for a class of quasi-Newtonian Stokes equations on general meshes. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307337.

Zara:2021:NIT

[ZM21a] M. C. Zara and L. H. A.

Monteiro. The negative impact of technological advancements on mental health: an epidemiological approach. Applied Mathematics and Computation, 396(??):Article 125905, May 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308584.

Zellinger:2021:THM

[ZM21b] Werner Zellinger and Bernhard A. Moser. On the truncated Hausdorff moment problem under Sobolev regularity conditions. Applied Mathematics and Computation, 400(??):Article 126057, July 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321001053.

Zhao:2021:AHC

[ZM21c] Yinghong Zhao and Yuechao Ma. Asynchronous H_{∞} control for hidden singular Markov jump systems with incomplete transition probabilities via state decomposition approach. Applied Mathematics and Computation, 407 (??):Article 126304, October 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321003933.

Zhang:2021:HPC

[ZMD^{+}21] Peng Zhang, Jianmin Ma, Menglan Duan, Ye Yuan, and Jinjia Wang. A highprecision curvature constrained Bernoulli–Euler planar beam element for geometrically nonlinear analysis. Applied Mathematics and Computation, 397 (??):Article 125986, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000345.

Zhang:2023:DAH

[ZMJ^{+}23] Ziyu Zhang, Xuehui Mei, Haijun Jiang, Xupeng Luo, and Yang Xia. Dynamical analysis of Hyper-SIR rumor spreading model. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000565.

Zandi-Mehran:2022:FBT

[ZMNR^{+}22]  Nazanin  Zandi-Mehran,

Fahimeh Nazarimehr, Karthikeyan Rajagopal, Dibakar Ghosh, Sajad Jafari, and Guanrong Chen. FFT bifurcation: a tool for spectrum analyzing of dynamical systems. Applied Mathematics and Computation, 422(??): ??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000728.

Zeifang:2023:TPN

[ZMS23a] Jonas Zeifang, Arjun Thenery Manikantan, and Jochen Schütz. Time parallelism and Newton-adaptivity of the two-derivative deferred correction discontinuous Galerkin method. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323003673.

Zontini:2023:EBC

[ZMS23b] Diego D. Zontini, Maikon L. Mirkoski, and João A. F. Santos. Error bounds in the computation of outer inverses with generalized Schultz iterative methods and its use in computing of Moore–Penrose inverse. Applied Mathematics and Computation, 440(??): ??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322007342.

Zhang:2020:RCQ

[ZN20] Erchuan Zhang and Lyle Noakes. Riemannian cubics in quadratic matrix Lie groups. Applied Mathematics and Computation, 375(??):Article 125082, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300515.

Zhanlav:2021:HOJ

[ZO21] T. Zhanlav and Kh. Otgondorj. Higher order Jarrattlike iterations for solving systems of nonlinear equations. Applied Mathematics and Computation, 395(??): Article 125849, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030802X.

Zou:2023:GAY

[Zou23] Qinmeng Zou. GMRES algorithms over 35 years. Applied Mathematics and Computation, 445(??):??, May 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000383.

Zhang:2020:MBN

[ZP20] Xu Zhang and Zheng Peng.

A modulus-based nonmonotone line search method for nonlinear complementarity problems. Applied Mathematics and Computation, 387(??):Article 125175, ???? 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320301442.

Zhang:2021:LSM

[ZP21] Yuexia Zhang and Dawei Pan. Layered SIRS model of information spread in complex networks. Applied Mathematics and Computation, 411(??):Article 126524, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321006135.

Zhao:2020:ACN

[ZPDL20] Shiyi Zhao, Yingnan Pan, Peihao Du, and Hongjing Liang. Adaptive control for non-affine nonlinear systems with input saturation and output dead zone. Applied Mathematics and Computation, 386 (??):Article 125506, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300320304641.

Zhou:2020:ETC

[ZPLL20] Yu Zhou, Yingnan Pan, Shubo Li, and Hongjing Liang. Event-triggered cooperative containment control for a class of uncertain non-identical networks. Applied Mathematics and Computation, 375(??): Article 125065, June 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300345.

Zada:2022:FTS

[ZPSP22] Akbar Zada, Bakhtawar Pervaiz, Muthaiah Subramanian, and Ioan-Lucian Popa. Finite time stability for nonsingular impulsive first order delay differential systems. Applied Mathematics and Computation, 421(??):??, May 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000297.

Zhuang:2022:ECS

[ZPW22] Jiawei Zhuang, Shiguo Peng, and Yonghua Wang. Exponential consensus of stochastic discrete multiagent systems under DoS attacks via periodically intermittent control: an impulsive framework. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004635.

Zhao:2022:QMF

[ZPY22] Huarong Zhao, Li Peng, and Hongnian Yu. Quantized model-free adaptive iterative learning bipartite consensus tracking for unknown nonlinear multi-agent systems. Applied Mathematics and Computation, 412 (??):Article 126582, January 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006664.

Zhao:2023:HPE

[ZPZH23] Shanshan Zhao, Qiuhui Pan, Wenqiang Zhu, and Mingfeng He. How “punishing evil and promoting good” promotes cooperation in social dilemma. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006853.

Zhang:2023:FHO

[ZQ23] Weijuan Zhang and Jianguo Qian. Flips on homologous orientations of surface graphs with prescribed forbidden facial cycles. Applied Mathematics and Computation, 444(??): ??, May 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322008487.

Zhao:2024:SAC

[ZQ24] Ningning Zhao and Yuanhua Qiao. Stability analysis of Clifford-valued memristorbased neural networks with impulsive disturbances and its application to image encryption. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001826.

Zhou:2024:DOM

[ZQC24] Hua-Cheng Zhou, Jun-Yan Qian, and Rui-Yang Cai. Disturbance observer and Mittag-Leffler stabilization design for multidimensional fractional distributed parameter systems. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400033X.

Zhang:2022:OBS

[ZQP^{+}22] Ning Zhang, Wenhai Qi, Guocheng Pang, Jun Cheng, and Kaibo Shi. Observerbased sliding mode control for fuzzy stochastic switching systems with deception attacks. Applied Mathematics and Computation, 427(??):??, August 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322002314.

Zhao:2021:SRL

[ZQZ21] Jianrong Zhao, Guoyou Qian, and Wei Zhao. On the sum of reciprocals of least common multiples, II. Applied Mathematics and Computation, 399(??): Article 126003, June 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000515.

Zarzour:2021:SPS

[ZR21] Nasir Zarzour and Kamel Rekab. Sequential procedure for software reliability estimation. Applied Mathematics and Computation, 402(??):Article 126116, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321001648.

Zhang:2024:NST

[ZRL^{+}24] Yuhui Zhang, Timon Rabczuk, Ji Lin, Jun Lu, and C. S. Chen. Numerical simulations of twodimensional incompressible Navier–Stokes equations by the backward substitution projection method. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006410.

Zhang:2023:PEL

[ZrLkL^{+}23] Tao Zhang, Zhong rong Lu, Ji ke Liu, Yan mao Chen, and Guang Liu. Parameter estimation of linear fractional-order system from Laplace domain data. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005963.

Zheng:2021:EIP

[ZRMD21] Junjun Zheng, Tianyu Ren, Gang Ma, and Jinhui Dong. The emergence and implementation of pool exclusion in spatial public goods game with heterogeneous abilityto-pay. Applied Mathematics and Computation, 394 (??):Article 125835, April 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307888.

Zeifman:2021:EBM

[ZRSK21] A. I. Zeifman, R. V. Razumchik, Y. A. Satin, and I. A. Kovalev. Ergodicity bounds for the Markovian queue with time-varying transition intensities, batch arrivals and one queue skipping policy. Applied Mathematics and Computation, 395(??): Article 125846, April 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307992.

Zhang:2024:PCT

[ZRZL24] Mengshu Zhang, Tianyu Ren, Xiao-Jun Zeng, and Jia Li. Promoting cooperation through dynamic trustworthiness in spatial public goods games. Applied Mathematics and Computation, 479(??):??, October 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324003643.

Zheng:2020:TLS

[ZS20a] Bo Zheng and Yueqiang Shang. A two-level stabilized quadratic equal-order finite element variational multiscale method for incompressible flows. Applied Mathematics and Computation, 384(??):Article 125373, November 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320303374.

Zheng:2020:TII

[ZS20b] Qianqian Zheng and Jianwei Shen. Turing instability induced by random network in FitzHugh-Nagumo model. Applied Mathematics and Computation, 381(??):Article 125304, September 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320302708.

Zengin:2021:NPI

[ZS21a] Rahman Salim Zengin and Volkan Sezer. A novel point inclusion test for convex polygons based on Voronoi tessellations. Applied Mathematics and Computation, 399(??):Article 126001, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000497.

Zhang:2021:PES

[ZS21b] Jie Zhang and Yuangong Sun. Practical exponential stability of discretetime switched linear positive systems with impulse and all modes unstable. Applied Mathematics and Computation, 409 (??):Article 126408, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004975.

Zhang:2022:ETV

[ZS22a] Guidong Zhang and Yuhong Sheng. Estimating timevarying parameters in uncertain differential equations. Applied Mathematics and Computation, 425(??): ??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001680.

Zheng:2022:TSSb

[ZS22b] Bo Zheng and Yueqiang Shang. A three-step stabilized algorithm for the Navier–Stokes type variational inequality. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005379.

Zheng:2022:TSSa

[ZS22c] Bo Zheng and Yueqiang Shang. A two-step stabilized finite element algorithm for the Smagorinsky model. Applied Mathematics and Computation, 422(??):??, June 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000571.

Zheng:2022:LQP

[ZS22d] Yueyang Zheng and Jingtao Shi. A linear-quadratic partially observed Stackelberg stochastic differential game with application. Applied Mathematics and Computation, 420(??): Article 126819, May 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009024.

Zhang:2021:AQP

[ZSC^{+}21] Weiwei Zhang, Chunlin Sha, Jinde Cao, Guanglan Wang, and Yuan Wang.

Adaptive quaternion projective synchronization of fractional order delayed neural networks in quaternion field. Applied Mathematics and Computation, 400(??): Article 126045, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032100093X.

Zhang:2022:PEE

[ZSFS22] Qifeng Zhang, Cuicui Sun, Zhi-Wei Fang, and Hai-Wei Sun. Pointwise error estimate and stability analysis of fourth-order compact difference scheme for timefractional Burgers’ equation. Applied Mathematics and Computation, 418 (??):Article 126824, April 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009073.

Zhang:2023:MPE

[ZSJ23a] Qinchunxue Zhang, Lan Shu, and Bichuan Jiang. Moran process in evolutionary game dynamics with interval payoffs and its application. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000449.

Zheng:2023:AGM

[ZSJ23b] Ruiling Zheng, Peifeng Su, and Xian’an Jin. Arithmetic-geometric matrix of graphs and its applications. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300322008323.

Zhang:2020:NFE

[ZSL20] Houchao Zhang, Dongyang Shi, and Qingfu Li. Nonconforming finite element method for a generalized nonlinear Schrödinger equation. Applied Mathematics and Computation, 377 (??):Article 125141, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301107.

Zhang:2020:SBD

[ZSM20] Jie Zhang, Yuangong Sun, and Fanwei Meng. State bounding for discrete-time switched nonlinear timevarying systems using ADT method. Applied Mathematics and Computation, 372 (??):Article 125002, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309944.

Zhang:2024:EPU

[ZSP24] Jin Zhang, Jing Shi, and Chen Peng. An extended prediction for uncertain LTI systems subject to input delays and unknown disturbances. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324003862.

Zhang:2024:EUE

[ZSS24] Fan Zhang, Hai-Wei Sun, and Tao Sun. Efficient and unconditionally energy stable exponential-SAV schemes for the phase field crystal equation. Applied Mathematics and Computation, 470(??):??, June 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032400064X.

Zhang:2022:SPS

[ZST22] Feifan Zhang, Jiamin Sun, and Wang Tian. Spatiotemporal pattern selection in a nontoxic-phytoplankton

— toxic-phytoplankton — zooplankton model with toxin avoidance effects. Applied Mathematics and Computation, 423(??):??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000935.

Zhao:2023:ORC

[ZSYX23] Guijie Zhao, Jiayue Sun, Ying Yan, and Xiangpeng Xie. Optimal reconstruction of constrained janbu method with ADP and non-integral safety factor. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003351.

Zhou:2020:GII

[ZSZ20] Jun Zhou, Jun Shen, and Weinian Zhang. Growth in an impulsive integral inequality. Applied Mathematics and Computation, 377 (??):Article 125152, July 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320301211.

Zhou:2020:RLS

[ZSZL20] Bo Zhou, Qiankun Song, Zhenjiang Zhao, and Tangzhi Liu. A reinforcement learning scheme for the equilibrium of the invehicle route choice problem based on congestion game. Applied Mathematics and Computation, 371(??): ??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308872.

Zeng:2022:TGF

[ZT22a] Yunhua Zeng and Zhijun Tan. Two-grid finite element methods for nonlinear time fractional variable coefficient diffusion equations. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322004829.

Zhou:2022:BDS

[ZT22b] Hao Zhou and Sanyi Tang. Bifurcation dynamics on the sliding vector field of a Filippov ecological system. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001382.

Zhu:2024:ECN

[ZT24] Qinze Zhu and Yingzhi Tian. Extra connectivity of networks modeled by the strong product graphs. Applied Mathematics and Computation, 477(??):??, September 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002868.

Zhai:2024:DDO

[ZTLL24] Ganghui Zhai, Engang Tian, Yuqiang Luo, and Dong Liang. Data-driven optimal output regulation for unknown linear discretetime systems based on parameterization approach. Applied Mathematics and Computation, 461(??):??, January 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004691.

Zhang:2024:IAC

[ZTLW24] Qianwei Zhang, Rui Tang, Yilun Lu, and Xinyu Wang. The impact of anxiety on cooperative behavior: a network evolutionary game theory approach. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001930.

Zhou:2022:NNS

[ZTW22] Xingyu Zhou, Yang Tian, and Haoping Wang. Neural network state observerbased robust adaptive faulttolerant quantized iterative learning control for the rigid-flexible coupled robotic systems with unknown time delays. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003605.

Zhang:2023:PPA

[ZTYZ23] Qi Zhang, Rongxia Tang, Zhengan Yao, and Zan-Bo Zhang. A parallel PageRank algorithm for undirected graph. Applied Mathematics and Computation, 459(??):??, December 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004459.

Zheng:2020:CMB

[ZV20] Hua Zheng and Seakweng Vong. On convergence of the modulusbased matrix splitting iteration method for horizontal linear complementarity problems of H_{+}-matrices. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308823.

Zheng:2021:MBS

[ZV21] Hua Zheng and Seakweng Vong. On the modulusbased successive overrelaxation iteration method for horizontal linear complementarity problems arising from hydrodynamic lubrication. Applied Mathematics and Computation, 402(??):Article 126165, August 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321002551.

Zhang:2021:AED

[ZW21a] Bo Zhang and Baoyindureng Wu. Anti-Eulerian digraphs. Applied Mathematics and Computation, 411(??):Article 126513, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321006020.

Zhang:2021:SFT

[ZW21b] Guozhen Zhang and Dajin Wang. The structure fault tolerance of arrangement graphs. Applied Mathematics and Computation, 400 (??):Article 126039, July 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300321000874.

Zhang:2021:DIA

[ZW21c] Xin-Fang Zhang and Qing-Wen Wang. Developing iterative algorithms to solve Sylvester tensor equations. Applied Mathematics and Computation, 409 (??):Article 126403, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300321004926.

Zhang:2022:RDA

[ZW22a] Zhipeng Zhang and Huimin Wang. Resilient decentralized adaptive tracking control for nonlinear interconnected systems with unknown control directions against DoS attacks. Applied Mathematics and Computation, 415 (??):Article 126717, February 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321008018.

Zheng:2022:MLS

[ZW22b] Bibo Zheng and Zhanshan Wang. Mittag-Leffler synchronization of fractionalorder coupled neural networks with mixed delays. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003770.

Zhang:2023:MBR

[ZWC23] Xuemei Zhang, Chunyan Weng, and Yaojun Chen. Multicolor bipartite Ramsey numbers for quadrilaterals and stars. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006506.

Zheng:2022:ITS

[ZWFW22] Yi Zheng, Xiaoqun Wu, Ziye Fan, and Wei Wang. Identifying topology and system parameters of fractionalorder complex dynamical networks. Applied Mathematics and Computation, 414(??):Article 126666, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007505.

Zhai:2023:FTT

[ZWH23] Junchang Zhai, Huanqing Wang, and Zuowei He. Fixed-time tracking control for high-order nonlinear systems with unknown timevarying input delay. Applied Mathematics and Computation, 452(??):??, September 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323001960.

Zhou:2020:FTS

[ZWHY20] Ya Zhou, Xiaoxiao Wan, Chuangxia Huang, and Xinsong Yang. Finite-time stochastic synchronization of dynamic networks with nonlinear coupling strength via quantized intermittent control. Applied Mathematics and Computation, 376 (??):Article 125157, July 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320301260.

Zhu:2020:ICE

[ZWJ^{+}20] Peican Zhu, Xiaoyu Wang, Danyang Jia, Yangming Guo, Shudong Li, and Chen Chu. Investigating the coevolution of node reputation and edge-strategy in prisoner’s dilemma game. Applied Mathematics and Computation, 386(??):Article 125474, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304331.

Zhou:2024:CSF

[ZWJ^{+}24] Guo Zhou, Tiantian Wang, Chen Jiang, Fangcheng Shi, Lei Zhang, Yu Wang, and Buyao Yang. A coupled smoothed finite element method and Lagrangian particle tracking model for three-dimensional dilute particle-laden flows. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032400198X.

Zhao:2021:EFC

[ZWNG21] Jinhua Zhao, Xianjia Wang, Lei Niu, and Cuiling Gu. Environmental feedback and cooperation in climate change dilemma. Applied Mathematics and Computation, 397(??):Article 125963, May 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000114.

Zhang:2022:BAS

[ZWS^{+}22] Hong-Tao Zhang, Yong-Ping Wu, Gui-Quan Sun, Chen Liu, and Guo-Lin Feng. Bifurcation analysis of a spatial vegetation model. Applied Mathematics and Computation, 434(??):??, December 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005331.

Zhuang:2021:NOC

[ZWW21a] Ziwei Zhuang, Qiru Wang, and Hongwu Wu. A new oscillation criterion for first-order delay differential equations by iteration. Applied Mathematics and Computation, 390 (??):Article 125632, February 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320305865.

Zhuang:2021:PSO

[ZWW21b] Ziwei Zhuang, Qiru Wang, and Hongwu Wu. A partly sharp oscillation criterion for first-order delay differential equations. Applied Mathematics and Computation, 409 (??):Article 126422, November 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321005117.

Zhang:2023:AMP

[ZWW23] Yanqi Zhang, Zhenlei Wang, and Xin Wang. Adaptive modified prescribed performance constraint control for uncertain nonlinear discrete-time systems. Applied Mathematics and Computation, 441(??):??, March 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322007846.

Zhang:2021:ADC

[ZWWH21] Huifeng Zhang, Xinjiang Wei, Yongli Wei, and Xin Hu. Anti-disturbance control for dynamic positioning system of ships with disturbances. Applied Mathematics and Computation, 396 (??):Article 125929, May 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300320308821.

Zhang:2020:DCM

[ZWXQ20] Ming Zhang, Guanghui Wang, Jin Xu, and Cunquan Qu. Dynamic contest model with bounded rationality. Applied Mathematics and Computation, 370(??): ??, April 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309014.

Zhu:2022:DFT

[ZWY^{+}22] Zhibin Zhu, Fuyong Wang, Yanhui Yin, Zhongxin Liu, and Zengqiang Chen. Distributed fault-tolerant containment control for a class of non-linear multi-agent systems via event-triggered mechanism. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003241.

Zhang:2024:DFS

[ZWY^{+}24] Jianyu Zhang, Yingying Wang, Songwei Yang, Jiaojiao Li, and Hao Qu. A design of fuzzy sliding mode control for Markovian jumping system with different input matrices. Applied Mathematics and Computation, 463(??):??, February 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005416.

Zhou:2022:OBS

[ZWYL22] Chao Zhou, Chunhua Wang, Wei Yao, and Hairong Lin. Observer-based synchronization of memristive neural networks under DoS attacks and actuator saturation and its application to image encryption. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001643.

Zhang:2021:PTO

[ZWZ^{+}21] Jiancheng Zhang, Zhenhua Wang, Xudong Zhao, Yan Wang, and Ning Xu. Prescribed-time observers of LPV systems: a linear matrix inequality approach. Applied Mathematics and Computation, 398 (??):Article 125982, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0096300321000308.

Zou:2022:NIE

[ZWZ^{+}22] Chengye Zou, Xingyuan Wang, Changjun Zhou, Shujuan Xu, and Chun Huang. A novel image encryption algorithm based on DNA strand exchange and diffusion. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322003654.

Zhu:2022:DME

[ZWZS22] Laixiang Zhu, Yanqian Wang, Guangming Zhuang, and Gongfei Song. Dynamicmemory event-based asynchronous dissipative filtering for T–S fuzzy singular semi-Markov jump systems against multi-cyber attacks. Applied Mathematics and Computation, 431(??):??, October 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S009630032200426X.

Zhu:2021:DFT

[ZWZX21] Baopeng Zhu, Yingchun Wang, Huaguang Zhang, and Xiangpeng Xie. Distributed finite-time fault estimation and fault-tolerant control for cyber-physical systems with matched uncertainties. Applied Mathematics and Computation, 403(??):Article 126195, August 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100285X.

Zhu:2020:TWS

[ZXB20] Wenjing Zhu, Yonghui Xia, and Yuzhen Bai. Traveling wave solutions of the complex Ginzburg–Landau equation with Kerr law nonlinearity. Applied Mathematics and Computation, 382(??):Article 125342, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303088.

Zhang:2020:SNF

[ZXC20a] Zhipeng Zhang, Chengyi Xia, and Zengqiang Chen. On the stabilization of nondeterministic finite automata via static output feedback. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306794.

Zhou:2020:HDC

[ZXC^{+}20b] Hui Zhou, Liufeng Xu, Yang Cui, Rongquan Feng, and Qi Ding. On Hamilton decompositions of Cayley graphs on dihedral groups. Applied Mathematics and Computation, 372 (??):Article 124967, May 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0096300319309592.

Zhao:2020:ETA

[ZXF20] Shubo Zhao, Xufeng Xiao, and Xinlong Feng. An efficient time adaptivity based on chemical potential for surface Cahn– Hilliard equation using finite element approximation. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319308938.

Zuo:2020:OBD

[ZXW20] Zhiqiang Zuo, Pengfei Xie, and Yijing Wang. Outputbased dynamic event-triggering control for sensor saturated systems with external disturbance. Applied Mathematics and Computation, 374(??):Article 125043, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300126.

Zheng:2020:NOD

[ZXZ20] Qunxian Zheng, Shengyuan Xu, and Zhengqiang Zhang. Nonfragile H_{∞} observer design for uncertain nonlinear switched systems with quantization. Applied Mathematics and Computation, 386(??):Article 125435, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303969.

Zhang:2024:DCF

[ZXZ24] Shen Zhang, Juanjuan Xu, and Huanshui Zhang. Decentralized control of forward and backward stochastic difference system with nested asymmetric information. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324002959.

Zhang:2022:PSN

[ZXZC22] He Zhang, Shengyuan Xu, Zhengqiang Zhang, and Yuming Chu. Practical stability of a nonlinear system with delayed control input. Applied Mathematics and Computation, 423(??): ??, June 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322000947.

Zhang:2020:BSO

[ZY20] Huanjun Zhang and Zhiguo Yan. Backward stochastic optimal control with mixed deterministic controller and random controller and its applications in linear-quadratic control. Applied Mathematics and Computation, 369(??):??, March 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319308343.

Zhang:2021:SCM

[ZY21a] Xiaofeng Zhang and Rong Yuan. A stochastic chemostat model with meanreverting Ornstein–Uhlenbeck process and Monod-Haldane response function. Applied Mathematics and Computation, 394(??):Article 125833, April 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0096300320307864.

Zhao:2021:NDM

[ZY21b] Zhenyu Zhao and Lei You.

A numerical differentiation method based on Legendre expansion with super order Tikhonov regularization. Applied Mathematics and Computation, 393(??): Article 125811, March 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307645.

Zhang:2023:PRN

[ZY23a] Liang Zhang and Jianhua Yin. The potential-Ramsey numbers r_{pot}(C_{n}, K_{t}^{−k}) and r_{pot}(P_{n}, K_{t}^{−k}). Applied Mathematics and Computation, 451(??):??, August 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002047.

Zhao:2023:FAS

[ZY23b] Jipeng Zhao and Guang-Hong Yang. Fuzzy adaptive secure tracking control against unknown false data injection attacks for uncertain nonlinear systems with input quantization. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006257.

Zhang:2024:TN

[ZY24a] Liang Zhang and Jianhua Yin. On the Turán number of K_{λ_{1}} ∪ K_{λ_{2}}. Applied Mathematics and Computation, 468(??):??, May 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300323006902.

Zhang:2024:OPR

[ZY24b] Panpan Zhang and Zhiguo Yan. Optimal portfolio with relative performance and CRRA risk preferences in a partially observable financial market. Applied Mathematics and Computation, 481(??):??, November 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324004089.

Zhai:2024:NSC

[ZYC^{+}24] Zhengliang Zhai, Huaicheng Yan, Shiming Chen, Yufang Chang, and Jing Zhou. Novel stability criteria of generalized neural networks with time-varying delay based on the same augmented LKF and bounding technique. Applied Mathematics and Computation, 460(??):??, January 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004587.

Zhang:2022:NPF

[ZYCY22] Xing Zhang, Zhibin Yan, Yunqi Chen, and Yanhua Yuan. A novel particle filter for extended target tracking with random hypersurface model. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001655.

Zhang:2021:EVV

[ZYLH21] Weilin Zhang, Lihua You, Hechao Liu, and Yufei Huang. The expected values and variances for Sombor indices in a general random chain. Applied Mathematics and Computation, 411(??):Article 126521, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100610X.

Zhang:2020:OCS

[ZYPL20] Heting Zhang, Zhanwen Yang, Kasia A. Pawelek, and Shengqiang Liu. Optimal control strategies for a two-group epidemic model with vaccinationresource constraints. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300319309488.

Zhang:2022:ETA

[ZYSW22] Xiaoli Zhang, Yang Yi, Mouquan Shen, and Qin Wang. Event-triggered antidisturbance tracking control for semi-Markovian jump systems with exogenous disturbances and input saturation. Applied Mathematics and Computation, 424(??): ??, July 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001151.

Zhang:2021:CBU

[ZYW21a] Libin Zhang, Zijun Yao, and Bin Wu. Calculating biodiversity under stochastic evolutionary dynamics. Applied Mathematics and Computation, 411 (??):Article 126543, December 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006275.

Zhao:2021:MPI

[ZYW21b] Ruju Zhao, Hua Yao, and Junchao Wei. Moore– Penrose inverses in rings and weighted partial isometries in C^{∗}-algebras. Applied Mathematics and Computation, 395(??):Article 125832, April 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320307852.

Zhao:2021:SCE

[ZYX21] Jingjun Zhao, Yulian Yi, and Yang Xu. Strong convergence of explicit schemes for highly nonlinear stochastic differential equations with Markovian switching. Applied Mathematics and Computation, 398 (??):Article 125959, June 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000072.

Zhang:2023:CND

[ZYXL23] Meijie Zhang, Xinsong Yang, Zhengrong Xiang, and Xiaoyang Liu. Consensus of nonlinear MAS via double nonidentical modedependent event-triggered switching control. Applied Mathematics and Computation, 453(??):??, September 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323002540.

Zhang:2020:FDP

[ZYY20] Xiao Zhang, Chunxiao Yang, and Jinge Yang. Fast diffusion in a porous building material with a nonlocal source. Applied Mathematics and Computation, 382(??):Article 125327, October 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320302939.

Zhao:2023:ADS

[ZYySX23] Junsheng Zhao, Yixuan Yuan, Zong yao Sun, and Xiangpeng Xie. Applications to the dynamics of the suspension system of fast finite time stability in probability of p-norm stochastic nonlinear systems. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003909.

Zhang:2022:ISR

[ZYZ22a] Ningbo Zhang, Qiwen Yang, and Xuzhen Zhu. The impact of social resource allocation on epidemic transmission in complex networks. Applied Mathematics and Computation, 433(??):??, November 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004799.

Zhang:2022:PPM

[ZYZ22b] Shengqiang Zhang, Sanling Yuan, and Tonghua Zhang. A predator–prey model with different response functions to juvenile and adult prey in deterministic and stochastic environments. Applied Mathematics and Computation, 413(??):Article 126598, January 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321006822.

Zhang:2022:ACG

[ZYZ22c] Zhongjie Zhang, Tingting Yu, and Xian Zhang. Algebra criteria for global exponential stability of multiple time-varying delay Cohen–Grossberg neural networks. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005355.

Zhou:2021:ASO

[ZyZW21] Yang Zhou, Li ying Zhang, and Tao Wang. Analytical solution for onedimensional non-Darcy flow with bilinear relation in porous medium caused by line source. Applied Mathematics and Computation, 392(??):Article 125674, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306275.

Zhang:2020:RSR

[ZZ20a] Xuefeng Zhang and Zeli Zhao. Robust stabilization for rectangular descriptor fractional order interval systems with order 0 < α <

1. Applied Mathematics and Computation, 366(??):??, February 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307581.

Zhou:2020:AOF

[ZZ20b] Hanqi Zhou and Junyong Zhai. Adaptive output feedback control for a class of nonlinear time-varying delay systems. Applied Mathematics and Computation, 365(??):??, January 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319306848.

Zhang:2022:LNM

[ZZ22a] Liang Zhang and Xiao-Qiang Zhao. Linear and nonlinear minimal speed selection of traveling waves for a competitive system with nonlocal dispersal. Applied Mathematics and Computation, 435(??):??, December 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322004349.

Zhang:2022:CFS

[ZZ22b] Yaxian Zhang and Heping Zhang. Continuous forcing spectrum of regular hexagonal polyhexes. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322001448.

Zhou:2023:CPP

[ZZ23] Yue Zhou and Yueming Zhong. A combinatorial proof of the parity unimodality of the (m, n)- rational q-Catalan polynomial for m ≠ 3. Applied Mathematics and Computation, 440(??):??, March 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300322006890.

Zhang:2024:EVG

[ZZ24] Zhimin Zhang and Wei Zhong. Efficient valuation of guaranteed minimum accumulation benefits in regime switching jump diffusion models with lapse risk. Applied Mathematics and Computation, 478(??):??, October 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0096300324002947.

Zhu:2024:IEC

[ZZB24] Junlei Zhu, Hongguo Zhu, and Yuehua Bu. Injective edge chromatic number of sparse graphs. Applied Mathematics and Computation, 473(??):??, July 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001401.

Zhang:2020:SFI

[ZZCL20] Meng Zhang, Yi Zhao, Lansun Chen, and Zeyu Li. State feedback impulsive modeling and dynamic analysis of ecological balance in aquaculture water with nutritional utilization rate. Applied Mathematics and Computation, 373 (??):Article 125007, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309993.

Zhang:2021:CCN

[ZZCW21] Juan Zhang, Huaguang Zhang, Yuliang Cai, and Wei Wang. Consensus control for nonlinear multi-agent systems with event-triggered communications. Applied Mathematics and Computation, 408 (??):Article 126341, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300321004306.

Zhou:2022:PCN

[ZZD22] Yihong Zhou, Xiao Zhang, and Feng Ding. Partiallycoupled nonlinear parameter optimization algorithm for a class of multivariate hybrid models. Applied Mathematics and Computation, 414(??):Article 126663, February 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321007475.

Zeng:2020:NIS

[ZZH^{+}20] Hong-Bing Zeng, Zheng-Liang Zhai, Yong He, Kok-Lay Teo, and Wei Wang. New insights on stability of sampled-data systems with time-delay. Applied Mathematics and Computation, 374(??):Article 125041, June 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300102.

Zhang:2020:SAO

[ZZL^{+}20a] Zhiming Zhang, Wei Zheng, H. K. Lam, Shuhuan Wen, Fuchun Sun, and Ping Xie. Stability analysis and output feedback control for stochastic networked systems with multiple communication delays and nonlinearities using fuzzy control technique. Applied Mathematics and Computation, 386(??):Article 125374, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320303386.

Zhao:2020:NEB

[ZZL20b] Ruijuan Zhao, Bing Zheng, and Maolin Liang. A new error bound for linear complementarity problems with weakly chained diagonally dominant B-matrices. Applied Mathematics and Computation, 367(??):??, February 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319307805.

Zuo:2022:BTL

[ZZL22] Hujian Zuo, Weifeng Zhao, and Ping Lin. Boundary treatment of linear multistep methods for hyperbolic conservation laws. Applied Mathematics and Computation, 425(??):??, July 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322001631.

Zheng:2024:IAF

[ZZL^{+}24] Miaozi Zheng, Peng Zhang, Binbin Li, Menglan Duan, Jianmin Ma, and Jinxin Wang. Improvement on the axial force accuracy of the ANCF Euler–Bernoulli beam element with the second-order approximate function of the centerline and precise constraint equations. Applied Mathematics and Computation, 466(??): ??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323006161.

Zhang:2023:ATG

[ZZLK23] Ju Zhang, Jin-Xin Zhou, Yan-Tao Li, and Young Soo Kwon. The automorphisms of 2-token graphs. Applied Mathematics and Computation, 446(??):??, June 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323000413.

Zhang:2021:DEE

[ZZLL21] Juan Zhang, Huaguang Zhang, Weihua Li, and Keqin Li. Distributed edge-event triggered consensus control for multiagent systems by edge-based asynchronous communications. Applied Mathematics and Computation, 397 (??):Article 125920, May 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320308730.

Zhang:2023:MBS

[ZZLV23] Yongxiong Zhang, Hua Zheng, Xiaoping Lu, and Seakweng Vong. Modulusbased synchronous multisplitting iteration methods without auxiliary variable for solving vertical linear complementarity problems. Applied Mathematics and Computation, 458(??):??, December 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323004174.

Zhang:2023:EGR

[ZZPF23] Yuji Zhang, Ziyan Zeng, Bin Pi, and Minyu Feng. An evolutionary game with revengers and sufferers on complex networks. Applied Mathematics and Computation, 457(??):??, November 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323003375.

Zhang:2020:NFO

[ZZQ^{+}20] Lihua Zhang, Yaoyao Zhou, Wenhai Qi, Jinde Cao, Jun Cheng, Yunliang Wei, Xiaoyu Yan, and Shaowu Li. Non-fragile observerbased H_{∞} finite-time sliding mode control. Applied Mathematics and Computation, 375(??):Article 125069, June 15, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300382.

Zhang:2020:PMA

[ZZS^{+}20] Huasheng Zhang, Guangming Zhuang, Wei Sun, Yongmin Li, and Junwei Lu. pth moment asymptotic interval stability and stabilization of linear stochastic systems via generalized H-representation. Applied Mathematics and Computation, 386(??):Article 125520, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320304781.

Zheng:2022:SAR

[ZZS^{+}22] Wei Zheng, Zhiming Zhang, Fuchun Sun, Hak Keung Lam, and Shuhuan Wen. Stability analysis and robust controller design for systems with mixed timedelays and stochastic nonlinearity via cone complementarity linearization. Applied Mathematics and Computation, 430(??):??, October 1, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322003265.

Zhang:2023:TSP

[ZZVL23] Yongxiong Zhang, Hua Zheng, Seakweng Vong, and Xiaoping Lu. A two-step parallel iteration method for large sparse horizontal linear complementarity problems. Applied Mathematics and Computation, 438(??):??, February 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322006828.

Zhu:2020:TMD

[ZZW20] Zhibin Zhu, Dongdong Zhang, and Shuo Wang. Two modified DY conjugate gradient methods for unconstrained optimization problems. Applied Mathematics and Computation, 373 (??):Article 125004, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300319309968.

Zeng:2021:HSC

[ZZW21a] Hong-Bing Zeng, Zheng-Liang Zhai, and Wei Wang. Hierarchical stability conditions of systems with time-varying delay. Applied Mathematics and Computation, 404(??): Article 126222, September 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032100312X.

Zhao:2021:HAA

[ZZW21b] Tengjin Zhao, Zhiyue Zhang, and Tongke Wang. A hybrid asymptotic and augmented compact finite volume method for nonlinear singular two point boundary value problems. Applied Mathematics and Computation, 392(??):Article 125745, March 1, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320306986.

Zhu:2021:TRB

[ZZW^{+}21c] Jun-Wei Zhu, Qiao-Qian Zhou, Li-Bing Wu, Jian-Ming Xu, and Xin Wang. Topology reconstruction based fault identification for uncertain multi-agent systems with application to multi-axis motion control system. Applied Mathematics and Computation, 399(??):Article 126000, June 15, 2021. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321000485.

Zhan:2020:NSL

[ZZX20a] Qingyi Zhan, Zhifang Zhang, and Xiangdong Xie. Numerical study on (ω, Lδ)- lipschitz shadowing of stochastic differential equations. Applied Mathematics and Computation, 376(??):Article 125108, July 1, 2020. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300320300771.

Zhao:2020:IRK

[ZZX20b] Jingjun Zhao, Yanming Zhang, and Yang Xu. Implicit Runge–Kutta and spectral Galerkin methods for the two-dimensional nonlinear Riesz space fractional diffusion equation. Applied Mathematics and Computation, 386(??):Article 125505, December 1, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009630032030463X.

Zhao:2021:LND

[ZZX21] Jingjun Zhao, Wenjiao Zhao, and Yang Xu. Lagrange nodal discontinuous Galerkin method for fractional Navier–Stokes equations. Applied Mathematics and Computation, 391 (??):Article 125697, February 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300320306500.

Zhou:2022:ARC

[ZZX^{+}22] Zepeng Zhou, Fanglai Zhu, Dezhi Xu, Shenghui Guo, and Younan Zhao. Attack resilient control for vehicle platoon system with full states constraint under actuator faulty scenario. Applied Mathematics and Computation, 419(??): Article 126874, April 15, 2022. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321009577.

Zhao:2023:HDG

[ZZX23] Jingjun Zhao, Wenjiao Zhao, and Yang Xu. Hybridizable discontinuous Galerkin methods for spacetime fractional advectiondispersion equations. Applied Mathematics and Computation, 442(??):??, April 1, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S009630032200813X.

Zeng:2024:NCO

[ZZX24] Caijin Zeng, Zhongcheng Zhou, and Chengkang Xie. Null controllability of an ODE-heat system coupled at boundary and internal term. Applied Mathematics and Computation, 475(??):??, August 15, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001966.

Zhang:2020:DDI

[ZZY20] Guang Zhang, Ruixuan Zhang, and Yubin Yan. The diffusion-driven instability and complexity for a single-handed discrete Fisher equation. Applied Mathematics and Computation, 371(??):??, April 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300319309385.

Zhang:2021:CSN

[ZZY^{+}21] Jiaqi Zhang, Ruigang Zhang, Liangui Yang, Quansheng Liu, and Liguo Chen. Coherent structures of nonlinear barotropic-baroclinic interaction in unequal depth two-layer model. Applied Mathematics and Computation, 408(??):Article 126347, November 1, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300321004367.

Zhou:2024:CAF

[ZZY24] Ziyi Zhou, Haixiang Zhang, and Xuehua Yang. CN ADI fast algorithm on non-uniform meshes for the three-dimensional nonlocal evolution equation with multi-memory kernels in viscoelastic dynamics. Applied Mathematics and Computation, 474(??):??, August 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300324001528.

Zong:2023:DBP

[ZZZ^{+}23] Xiju Zong, Ji-Feng Zhang, Jiangtao Zheng, Zengxin Liu, and Guoyao Ma. Design of boundary pulse width modulation controller for a class of heat equations. Applied Mathematics and Computation, 437(??):??, January 15, 2023. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300322005859.

Zhuang:2024:CFG

[ZZZ24] Jiayuan Zhuang, Yuanpeng Zhu, and Jian Zhong. Curve fitting by GLSPIA. Applied Mathematics and Computation, 466(??):??, April 1, 2024. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0096300323005969.